A Complete Bibliography of Publications in Marine Ecology Progress Series: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

25 November 2024 Version 1.20

Title word cross-reference

−20 [LAP^{+}24]. ^{13} [BCC20, HRH22, MBW^{+}20, ZMD^{+}21]. ^{1}5 [MBW^{+}20, BCC20, FLC^{+}24b, HRH22, LNBdB21, SCW^{+}24, ZMD^{+}21]. ^{34} [SB20b]. _{37}^{k} [RBVA23]. _{2} [BJM^{+}22, EHHF^{+}20, GMB^{+}24, HRC21, JMS^{+}22, KVdM^{+}22, LHW^{+}20, MR20, MGM^{+}24]. _{4} [KVdM^{+}22]. : [NYW^{+}22, SST^{+}20a, VS23]. δ [BCC20, FLC^{+}24b, HRH22, LNBdB21, MBW^{+}20, SB20b, ZMD^{+}21]. p [GMB^{+}24, MGM^{+}24].

-influenced [SHH^{+}22].

110^{o} [LHD20]. 182 [Ano22b]. 19 [LGR^{+}22]. 1960s [THS^{+}24]. 1970s [THS^{+}24].

2000s [THS^{+}24]. 2010s [THS^{+}24]. 227 [Ano22a]. 2D [FTC^{+}20].

30-year [ZPVN23]. 3D [FTC^{+}20].

5 [ZLHH20]. 51 [Ano22c].

627 [Ano21d]. 641 [Ano21c]. 645 [Ano20]. 646 [Ano21a]. 654 [Ano21h]. 656 [Ano21b]. 660 [Ano21e, Ano21i]. 666 [Ano21g]. 678 [Ano21f]. 683 [Ano22d]. 688 [Ano22f]. 693 [Ano22c].

95 [Ano22e].

aalge [SFMK^{+}23]. Abalone [ABB^{+}20, GAFS20, HNKK21, MY21]. abandonment [SWHE20]. Abbott [CBWH24]. abbotti [CBWH24]. ability [DBS^{+}20, Lei20]. abiotic [CDS23, FTA^{+}22, HBR^{+}21, RHM^{+}22, ZZYB21]. above [KPS^{+}22]. Absence [RCB24b, DFCM21]. absolute [RLSO23]. Abundance [OÁBP^{+}21, PS23, SSB^{+}20, SF23, BKR^{+}21, BWRP20, BLE^{+}22, CHD^{+}22, CGH^{+}23, DSHM20, DSK^{+}20, DBK^{+}23, FBC^{+}24a, FBP^{+}22, GB20, GKK^{+}22, GMLC22, HKD^{+}20, JDM^{+}20, JHDE20, LCFJ20, MSR^{+}21b, MMS^{+}22, OLM24, OJ24, RKV^{+}21, RWH^{+}24, RBCAV22, RLSO23, SME^{+}23, WC21, WFS23]. abundances [DHE^{+}23, KBWRV21, dlBSL^{+}22]. abundant [BCC20, CLL^{+}22, MLP^{+}24]. Acanthaster [DDB^{+}24]. acanthias [ASY^{+}24]. Acanthurus [LD24]. Acartia [LSSG24, VKdJ24]. accelerates [SGP24]. acceleration [SSF^{+}22]. access [MKR^{+}20]. Acclimation [IBRL^{+}22, UBO^{+}24, DB21, FMD20, JH20, LLL20, RHM^{+}22]. according [KKMW23]. account [DCR^{+}21]. Accounting [SBFP^{+}24, WYZ^{+}20]. accumulation [WSD^{+}20]. accumulations [dBDL^{+}20]. Acetate [CKM^{+}20]. Acetate-utilizing [CKM^{+}20]. achilles [LD24]. acid [Bai22, DHA^{+}20, EHHF^{+}20, FLC^{+}24b, HIT^{+}22, ICH^{+}24, JJV^{+}21, KGG^{+}21, LAP^{+}24, LNBdB21, MPR^{+}23, PSN^{+}20, SDM^{+}22, SHZ^{+}20, SSG^{+}22, ZZLS24]. acid-base [EHHF^{+}20]. acidification [ASBE22, BK23, BMB24, CG22, CKP20, DHA^{+}20, GON^{+}20, HMB^{+}22, HHBP23, KLN^{+}21, LK21, LSS^{+}22, RKV^{+}21, SBSS21, SOSE22, SJSR24, VHG20, WGK22, YDG21, vLBGGT22]. acids [HBE22, MBW^{+}20, SDW20, XCT^{+}20]. Acoustic [AHCT22a, OLA^{+}24, SMH^{+}20, AUS^{+}23, BBPO21, BBH^{+}20, BMS^{+}21, CGM^{+}24, CKKL21, DPDJ^{+}21, DMM^{+}24a, ESG20, FLBP^{+}21, GMS^{+}21, LMS^{+}20, LCR^{+}22, MJJJ20, OWH23, PBAC23, PBS22, PAAEV^{+}24, RYF^{+}24, RS20b, SHS21, SBFP^{+}24, VH20, WSPS^{+}20, ZSWR^{+}22]. acoustics [CDD^{+}20a]. Acropora [IBRL^{+}22, LvAH^{+}24, STF^{+}24, YRB^{+}20]. acroporid [NSM24]. Across [STO^{+}23, AKW^{+}22, ATR^{+}24, Ano23, AR21b, BTB^{+}24, BMOW21, BBH^{+}20, BMP^{+}22, CP22b, DL21, DMM^{+}24b, FN24, GSL^{+}24a, HHO^{+}23, HAB^{+}22, HB24, IBC^{+}21, LBR^{+}24, MHH^{+}20, MDY^{+}21, MDB^{+}21, OVÁS^{+}21, OLG^{+}23, QRC^{+}24, RHSM21, RBB^{+}24, RS21, RC23, RBCAV22, ROMB22, SFL23, SS24, SSFS20, THL^{+}20, THR^{+}23, VIV^{+}20, WBS24, ZHG^{+}21]. act [CGH^{+}23, DDD^{+}23]. acting [RGK22]. Actinia [CSWM24]. active [RCES22, SF23]. activities [HWT^{+}22]. Activity [CGCC24, ADBF21, BP24, CBWH24, DBHCO^{+}22, EMB^{+}22, FSK^{+}21, GVOCSGG21, LFA^{+}23]. acuminata [LSSG24]. Acute [ASS^{+}21b, AR21a, CKP20, SCP^{+}23]. acutorostrata [BMS^{+}21]. acutus [LCP^{+}20]. adaptation [AVFD^{+}21, BWC^{+}22, NGRH20, YRB^{+}20]. adaptions [LCFJ20, YWW^{+}24]. ADD [BMS^{+}21]. addition [KBWRV21, MGH21]. additive [WYO^{+}20]. addressed [HvLM21]. Adélie [LLJ^{+}20, MAS^{+}20, MKR^{+}20, RBW^{+}20]. adhaerens [WSU21]. adipose [LAP^{+}24]. adjacent [AHMVM23]. Adriatic [CCA^{+}20, HGT^{+}21, ZMC^{+}22, ZFD^{+}24]. adult [ABB^{+}20, CSN^{+}23, FSK^{+}21, HNKK21, RLME^{+}21, RS20b, SRHM21, SBJ^{+}24, TCY^{+}22]. adults [VHG20]. advance [CGM^{+}24]. Advances [Ano23, GAPM24]. Advancing [DTHC24, RHN^{+}24]. advantage [NMB^{+}22]. advection [LNKJ^{+}22]. advice [TOMDC23]. Aegean [GG23, KDK^{+}24]. aerial [AKGA^{+}21, BJH^{+}23, CP22b, SFMK^{+}23]. Aerobic [BMS^{+}24, ASY^{+}24, CHJ21]. aff. [STF^{+}24]. affect [CKT^{+}24, CEAO^{+}22, DA24, GBG^{+}23, LCP^{+}20, MPB^{+}24, RVCT21, RHH^{+}20, RTT^{+}21, SdFBM24, WTK^{+}24, WHB^{+}20]. affected [BS20a, CGCC24, SFOPL21]. affecting [Ano23, RBC^{+}24, SMA^{+}20]. affects [ABD^{+}21, AMD22, BAAR20, CCA^{+}20, FLY^{+}24, JWH^{+}21, JPJH23, LAD^{+}21, OJ24, Tha21, TGB^{+}20]. afflicted [VS23]. Africa [BAS23, CKCS20, DMM^{+}24a, DHF24, LMH^{+}24, MBS^{+}24, MLP^{+}24, PSP^{+}20]. African [SAH^{+}24]. after [BSA^{+}24, EWBH24, MLE^{+}22, NF23, NMB^{+}22, OLPGK22, RVV^{+}24, SLO^{+}22, SNE^{+}24, TSS^{+}23, VGA20]. Aftereffects [VSGD24]. against [EB21]. Age [CSK^{+}23, FLC^{+}24a, KSMP20, SHL^{+}21, ARH^{+}20, DMTP22, KRG^{+}20, KWD21, MBW^{+}20, SHZ^{+}20]. Age-specific [SHL^{+}21]. ageing [CTH^{+}20]. Agelas [AGA^{+}22]. Agent [PLH^{+}21, LBMSS23]. Agent-based [PLH^{+}21, LBMSS23]. agents [BAB^{+}24]. aggregates [LBB^{+}21]. aggregating [DDD^{+}23]. Aggregation [HNKK21, NGPGO^{+}23, BGMRGM^{+}20, BBG^{+}22, FPZJAO^{+}22, GNC^{+}23, HNB^{+}24, LGGR^{+}22, MRB^{+}24, MPB^{+}24, WFS23]. aggregations [AKGA^{+}21, CDD^{+}20a, GBE^{+}22]. aging [CBJ^{+}23]. agrees [BZT^{+}23]. ahead [GRB22]. aid [SWP20]. airguns [VKdJ24]. airspeed [KKMW23]. akashiwo [BAS23]. Akumal [REGP22]. al [Ano20, Ano21a, Ano21b, Ano21d, Ano21c, Ano21f, Ano21e, Ano21g, Ano21h, Ano22a, Ano22b, Ano22c, Ano22d, Ano22f, Ano22e, DFRS^{+}24, KCT24]. Alabama [APH^{+}23]. Alaria [BAHG23, MBB^{+}22]. alarm [SST20b]. alascanus [DMTP22]. Alaska [SAMP24, BBR24, CKS^{+}24, DME^{+}23, FYM^{+}24, LKR^{+}21, LHM^{+}23, MGA^{+}20, MSF24, RSEG21, SRG^{+}23, SOS^{+}21a]. Alaskan [FWWH22, SIBM20]. albacares [AAFC^{+}21, FLC^{+}24a, SKTO23]. albatross [CCP^{+}21, FCV^{+}24, GFH^{+}22, MHAL24, MCJ^{+}20, VLG^{+}21]. albatrosses [FRO^{+}21, PGT^{+}21, SSC^{+}23, SWR23]. albimarginatus [TCM^{+}23]. albirostris [SBS^{+}22]. alcyonaceans [SMA^{+}20]. Aldabra [ACvdC^{+}21]. alert [dMGGG21]. Aleutian [SOS^{+}21a]. alewife [OLA^{+}24]. alexandrini [NBW^{+}23]. alfredi [ACC^{+}20, VvDRM20]. alga [BE21, GS20, PLG^{+}23, SE20]. algae [CCC^{+}23, CAY24, CDS23, FDF20, GPCM22, HM23, KLN^{+}21, KHG^{+}21, LHHH^{+}20, OCFEB21, VSG20, VLPN24]. algal [BFC^{+}20, BMPD20, BAS23, FQVMF23, FLCQ21, KKN^{+}20, KGG^{+}21, LAL21, LWP^{+}24, TCM^{+}20, VK21]. algal- [BFC^{+}20]. algal-dominated [FQVMF23]. algal-gardening [VK21]. Aliger [VSGD24]. Alkenone [RBVA23]. Alleged [ASL^{+}22]. allocation [BQM^{+}22, CBHM21, LBH^{+}21b]. allochthonous [RLP^{+}23]. allowed [Ano23]. allows [BSP^{+}22]. along [AHCT22a, AHCT22b, BFQ^{+}23, BDJ^{+}24, BSA^{+}24, CVP^{+}21, GMRMG22, GLM^{+}24, HMW^{+}21, IAG^{+}23, LBH^{+}21a, LHD20, LSC^{+}21, LBH^{+}21c, LBMF^{+}24, MLP^{+}24, RWSR24, TSW^{+}21, VSG20]. Alopias [KKS^{+}20]. alpha [Vir23]. alpinus [HHM^{+}22, HYG^{+}20]. also [CCC^{+}23]. alter [DMBB22, EHHF^{+}20, FNH^{+}23, KKMW23, SFOPL21]. altering [Ano24o]. alternative [GSL20, HBJ^{+}21]. alterniflora [HPH20, NCZ^{+}21, RS21]. alters [CRG^{+}20, FS22, HMKK^{+}20, LPO^{+}21]. alveolata [VDC^{+}24]. alvinocaridid [MON^{+}24]. Alvinocaris [MON^{+}24]. Amazon [ARN^{+}22]. amberjack [TCY^{+}22]. ambient [MLE^{+}22, YXB22]. ambiguity [AUS^{+}23]. ambush [PBT^{+}21]. America [AGB24, RPD^{+}23]. American [BGL^{+}21, CGW21, CCW21, JBWS24, KCB^{+}24, LCP^{+}20, SJSR24, SNH^{+}20, TWC20, WBGC24]. americanus [CGW21, DFCM21, KCB^{+}24, MLB^{+}21, MPSH22, SJSR24, TFCS22, WBGC24]. amid [TPGG^{+}23]. amidst [AMD22]. Amino [ICH^{+}24, SDM^{+}22, FLC^{+}24b, LNBdB21, PSN^{+}20, SDW20, XCT^{+}20]. Ammodytes [BJM^{+}22, LBW21]. ammonium [JRVE^{+}22, KBWRV21]. among [AGB24, AOO^{+}21, BMKB22, BFQ^{+}23, BBMM20, BLCC^{+}20, DKSS20, DMBB22, Fai21, GBE^{+}22, HHO^{+}23, HSNT24, ILON21, ISMM20, JHDE20, LO20, MAPFH22, MJW24, MRH20, OFSJ20, PRCvdK23, PBRF23, RMD24, ŠGH^{+}21, TPGG^{+}23, WRG^{+}24, YRB^{+}20]. among-colony [BFQ^{+}23]. Amphibalanus [JPJ^{+}22]. Amphipod [dOFBSdJR^{+}22, AC20, LFA^{+}23, PFM^{+}21]. amphipods [KHOK22]. amphitrite [JPJ^{+}22]. amplified [RAHM^{+}21]. amplifies [KLN^{+}21]. ampullatus [DFHT23, WIM23]. anadromous [HYG^{+}20, LCP^{+}23]. anaerobic [PFDF^{+}21]. analis [HNB^{+}24]. analyses [CDP^{+}21, FNO^{+}23, GWDR20, HKY^{+}20, HTGG22, JJV^{+}21, KdlVJ^{+}21, LRC20, OCR^{+}24, PAL^{+}22, PSN^{+}20, QFT^{+}24, VMDA22, WYZ^{+}20, ZMC^{+}22]. analysis [APOGG20, AZ24, BTDR20, BLVL24, BWGH23, CDF^{+}20, CG22, CCL^{+}22, CAPS20, DPDJ^{+}21, DSK^{+}22, DDD^{+}23, EB20, FS21, FWWH22, FFMR24, FNL^{+}23, GFT^{+}21, GSN^{+}20, GB20, HRH22, HSA^{+}21, HBJ^{+}21, JPJ^{+}22, LWP^{+}24, LORB^{+}24, MTSO^{+}24, MHK22a, MHH^{+}22, NSKH23, PBB^{+}20, PLC^{+}20, PWvdM^{+}21, PRC^{+}21, RBF^{+}20, SHZ^{+}20, SD22, TGT^{+}23, TMN^{+}24, UIS^{+}24, WWH^{+}22, XCT^{+}20, ZPVN23, dSLD24, dLCN^{+}22, vTLG^{+}21]. analytical [OH22]. ancestral [MCFB20]. Anchoa [AAHW20]. anchoita [CBKM20]. anchoveta [CMT^{+}21]. Anchovy [FBC^{+}24b, AAHW20, CILGD^{+}22, CBKM20, CKCS20, PRC^{+}21, TMPPG24, vdKMC^{+}24]. ancient [MMA^{+}22]. anemone [BLB23, GRB22]. anemones [CSWM24]. angle [LT21]. Anguilla [FJB^{+}23, HMK^{+}22, SCM21]. anguillid [THK^{+}22]. animal [Ano23, HPFB20, KCLJ20]. animals [BD20, ICH^{+}24]. annandalei [NGRH20]. Annelida [CGLC20, HKD^{+}20, PGC21, YWW^{+}24]. annual [ABD^{+}21, BMPD20, BDS^{+}20, DMS^{+}21, DRP^{+}20, MAS^{+}20, OLA^{+}24, UAL24, dDW^{+}22, dSGC^{+}23]. anomalies [DVMH^{+}24]. Anomalous [TMPPG24]. Anoplopoma [CBHM21]. Antarctic [IAG^{+}23, ANP^{+}24, CSD^{+}24, DFP^{+}21b, FJJ^{+}21, GSL^{+}24a, GDAC23, HOK^{+}22, IMM^{+}20, LKB^{+}21, MSL^{+}24, MSR^{+}21b, SBSS21, SNP20, SBRQ21, VGG^{+}23, WRW20, ZZLS24, dLCN^{+}22, vTLG^{+}21]. Antarctica [ORB^{+}21, RBW^{+}20, VH20]. antarcticus [PBO^{+}21]. Anthropocene [RNQ^{+}20]. Anthropogenic [CBVA^{+}21, LD24, MPSH22, HGT^{+}21, MCH^{+}21]. anti [CGGdD24, LK21, RWSR24]. anti-herbivore [RWSR24]. anti-predation [LK21]. anti-predatory [CGGdD24]. antillarum [LvAH^{+}24]. antimicrobial [AGA^{+}22]. antioxidant [VSG20]. antioxidants [CCBD^{+}21]. Antipatharia [MMA^{+}22]. antipodes [MCFB20, MCC^{+}21]. Antipredatory [AEHD21]. Aotearoa [FNM^{+}21, MON^{+}24]. Aotearoa/New [MON^{+}24]. apex [GDCP20, MDS^{+}23, RHSM21]. Apocyclops [NGRH20]. Apparent [GDAC23]. appearance [MMS^{+}22]. appendicularian [DPCL^{+}23]. appetites [BSRVS22]. Application [SRS^{+}20, ACG22, COT^{+}21, OWH23, PCP^{+}20, TGT^{+}23]. applications [FTA^{+}21, HJG21, OH22]. applied [RSS20]. approach [BBS^{+}21, BQM^{+}22, BSH^{+}21b, CLV^{+}20, CCAdA24, CBB^{+}22b, FPG^{+}21, LBMSS23, ZCBC24]. approaches [BMW24, KSS^{+}20, LDW^{+}21, OH22]. approved [PSP^{+}20]. aquaculture [HGK^{+}24, KCLJ20, MVH^{+}21, MVC^{+}23, MTRM21, RAHM^{+}21]. aquatic [BMO^{+}24, BB24, OH22, OWH23]. Arabian [SMA^{+}20]. arbuscula [VHG20]. Arc [BKR^{+}21]. archipelagoes [YMU^{+}21]. archival [FKTK20, TCM^{+}23]. archived [WWH^{+}22]. arctic [SFO^{+}21, Bai22, BMF24, CIDM23, CLC^{+}23, CDP^{+}21, DFHT23, DMS^{+}21, DHJ^{+}22, DB21, DDG^{+}21, FDE^{+}21, GB20, HHM^{+}22, HYG^{+}20, HEKH24, KHL^{+}22, KGM^{+}24, KGG^{+}21, KHG^{+}21, LMRS^{+}24, LHFD^{+}24, LBTE^{+}21, LHHH^{+}20, MDW^{+}22, MBB^{+}22, NSB^{+}23, PGGE21, RB22, SGB^{+}21, SIBM20, SB20b, UBO^{+}24, URSS24, WLA^{+}21, YGMNA22]. Arctic-breeding [LBTE^{+}21, SGB^{+}21]. Arctocephalus [dLCN^{+}22]. Arctoscopus [KI24]. Ardenna [BLCC^{+}20, FNM^{+}21]. area [AU21, ACC^{+}20, AKGA^{+}21, BWC^{+}22, CDF^{+}20, CCS^{+}22, HYG^{+}20, HNB^{+}24, HCP^{+}24, HBG^{+}22, HMK^{+}22, ISMM20, LJ20, LTM20, MSL^{+}24, MBB^{+}23, NSB^{+}23, PBAC23, PTCD22, RFJC^{+}24, RMR^{+}23, RBF^{+}20, RLSO23, SW23b, SMPLF^{+}22, SPFL20, WLB^{+}20, SVT^{+}24]. area-based [HBG^{+}22]. areas [ACA^{+}24, AAFC^{+}21, BSP^{+}22, BPS^{+}24b, DVP^{+}23, FST22, LBTE^{+}21, MCC^{+}21, NCS^{+}23, OvBS^{+}24, PSP^{+}20, RLME^{+}21, SOS^{+}21a, SD22, VDC^{+}24, Whi20, Whi21]. argenteus [AGIS^{+}22]. Argentina [ASS^{+}21b, BFQ^{+}23, CGM23, MBW^{+}20]. Argentine [ASIA22, CBKM20]. argentinus [ASIA22]. Argopecten [GMB^{+}24]. argus [ADBF21, BBRM22]. Argyrosomus [NHFS22, VAMF22]. Arhynchobatidae [SOS^{+}21b]. Ariake [YOI^{+}23]. arms [BLA^{+}24]. arrival [Ano24o]. arthropods [RLP^{+}23]. articulated [GPCM22]. artifacts [BH22]. artificial [BDS^{+}20, HBS^{+}21, MTSO^{+}24, MPKS24, PBT^{+}21, VPB^{+}20]. artisanal [SMPLF^{+}22]. ascidian [ARdlB^{+}22, BSP^{+}21]. ascidians [dSLD24]. Ascothoracida [KZC23]. asexual [LLL20]. Asia [MSL^{+}21]. Asian [RAA^{+}23]. aspect [GMCE^{+}22]. aspects [ZFD^{+}24]. assemblage [CDC21, FST22, GRQ21, GSPH20, HSB20, HEP^{+}24, KMW^{+}21, LFFW21, MBS^{+}24, NMAE^{+}22, REGP22, SPFL20, SRR^{+}20, VGA20]. Assemblages [OLPGK22, AMUPJR^{+}24, AHMVM23, BICQG23, CCR^{+}20, Cha22, DSHF22, DHF24, FAJAO^{+}21, FSSL^{+}20, GSC^{+}22, GS20, GDG^{+}22, KKN^{+}20, LFK^{+}21, MDPL20, MMB22, OJ24, PCR23, RBM22, SDM^{+}22, SCW^{+}24, SME^{+}23, SRG^{+}23, SKI^{+}20]. assembly [BMK^{+}22, GLM^{+}24]. assess [BTDR20, EAS20, GBA24, LNBdB21, RSS20, SRS^{+}20]. assessed [AKGA^{+}21, BASM^{+}24, HBG^{+}22, VMDA22]. Assessing [APOGG20, CC22, CWC^{+}22, GBF^{+}23, HESF24, HRH22, HHS21, JETGM24, MAK^{+}22, PCLQ^{+}20, SCW^{+}24, TKM^{+}22]. Assessment [GBH^{+}20, MGHH21, BMS^{+}24, CLR^{+}21, CSWM24, EMB^{+}22, ECC^{+}21, FRO^{+}21, FLC^{+}24b, HO21, MSC^{+}22, TWC22, YSD^{+}21]. assessments [SBW^{+}22]. assimilation [DDR^{+}21, SJ22, SOSE22]. associated [AWB23, AAP^{+}23, BAB^{+}24, CTTLCG^{+}24, CCJ^{+}23, CM22, CRGMA^{+}21, CBVA^{+}21, DFP^{+}21a, DL21, HSNT24, HOK^{+}22, IMYW24, KMW^{+}21, KNH24, KI24, LBH^{+}21c, MHK22b, PBS22, PBKG22, SKHC23, SdGWN21, TP20, TTA20, VKdJ24, VAMF22, VBB^{+}21, WC21, WRG^{+}24, ZCB20, ZBB^{+}20, dBDL^{+}20]. Association [HIT^{+}22, VK21]. Associations [DKSS20, DKSS21, CDF^{+}20, CPH20, CP22b, DIC^{+}22, GAF20, HNKK21, LSKS24, MDK^{+}21, MPKS24]. assumption [RBF^{+}20]. assumptions [DDR^{+}21, LBMB^{+}20]. Asterias [SKHC23]. asteroid [Hew21, ZDBS20]. Astrophorina [SMS^{+}24]. Asynchronized [TNS^{+}22]. At-sea [DCR^{+}22, WdGR^{+}20, BKA^{+}20, PGT^{+}21]. Atherina [VMCA^{+}20]. Atlantic [ABN^{+}21, BBBT^{+}20, FSS23, GFT^{+}21, GSL^{+}24b, MPR^{+}23, PS23, PDH^{+}24, RBM^{+}21, SCM^{+}23, SMS^{+}24, ASIA22, AHH^{+}20, AMT^{+}21, AGIS^{+}22, AGB^{+}23, BKR^{+}21, BVKF23, BLW^{+}24, BSF24, BSÁG^{+}22, BLCC^{+}20, BGDJ20, BKH^{+}24, CDD^{+}20a, CDD^{+}20b, CFBPH23, CBB^{+}22b, CBKM20, DIC^{+}22, DSK^{+}20, DGM^{+}23a, DSA^{+}24, EAS20, FTA^{+}21, FOO20, FNL^{+}23, FDE^{+}21, FNS^{+}20, GLB^{+}20, GBG^{+}23, GRQ21, GON^{+}20, GMS^{+}21, GBA^{+}22, GBF^{+}23, HBE22, HBJ^{+}21, HCvdHM23, JWH^{+}21, KCGR20, KMW^{+}21, KRMR24, LOF^{+}24, LÁGLL20, LLF^{+}20, LNKJ^{+}22, LCGS^{+}21, MdOO^{+}20, MVC^{+}23, MAH^{+}20, MHF^{+}21, MMMV22, MFCF21, MDY^{+}21, MHÁC^{+}22, MDQ^{+}21, NCS^{+}23, OvBS^{+}24, OVÁS^{+}21, OSJ^{+}22, OJ24, PSÓ^{+}24, PDM22, RMR^{+}23, RLME^{+}21, RYF^{+}24, RGK22, RRR^{+}23, RCB24b, RMD24, SBS20, SFLQ23, SDE^{+}22, SDF^{+}21, SDE^{+}21, SD22, TQG24, VGA20, VMDA22, VIV^{+}20, WIM23, WF21]. atlanticum [SSB^{+}20]. Atlantidae [AMMADDH20]. Atlanto [FSAG22]. Atlanto-Iberian [FSAG22]. atmospheric [Ano23]. Atoll [ACvdC^{+}21, ECvBL21, PBS22]. atresia [GVCSO^{+}24]. Atrina [GVOCSGG21]. attachment [DGM^{+}23b, RC23, WMT20]. attacks [PBS^{+}24]. attempt [MAS^{+}20]. attendance [ODB^{+}24]. attracted [MBM^{+}20]. attraction [JTBS^{+}22, SBS20]. auk [DMS^{+}21]. auklets [STO^{+}23, SWHE20]. auks [CFSH22]. auratus [CHJ21]. Aurelia [LLL20, RL22]. aurita [LBB^{+}21, LLL20]. austral [BBGM23]. Australasian [CHJ21, RMASA20]. Australia [BMP^{+}22, CLM^{+}21, CMTP22, DKS^{+}24, GSPH20, LBH^{+}21c, LMB^{+}20, MRH20, PSBH21, SB20a, SMB^{+}20]. Australian [BLWJ20, DKAB23, GCDA20, GSC^{+}22, NSPH20, RCC^{+}22, VPH^{+}21]. australis [MBW^{+}20]. Austrovenus [FSF^{+}24]. autochthonous [SWP20, UAL24]. automatic [MDH^{+}20]. autonomous [MCL24]. Autumn [JBDH21, ZZLS24, BSF24, OSJ^{+}22]. autumn-spawned [BSF24]. availability [CCRP20, CGR^{+}20, DNE^{+}24, GJM^{+}20, GBF^{+}23, HOK^{+}22, IMM^{+}20, KSW^{+}21, MSL^{+}21, NT24, RC23, RWSR24, SMA^{+}20, SSBA^{+}20, ZZLS24]. Avicennia [LPD^{+}24]. avoid [BMS^{+}21, HBC^{+}20]. avoidance [IMM^{+}20, JTBS^{+}22, TNMN24]. away [KHOK22, SHM^{+}20b]. Azores [PDH^{+}24].

Babitonga [PBAC23]. Baccalaureus [KZC23]. backed [JTBS^{+}22]. background [BCP^{+}22]. backscatter [BMOW21]. bacteria [LRC20, YAZA21, ZCB20]. bacterial [LDM^{+}20, SWP20]. bacterium [YHW^{+}23]. Baffin [DFHT23, SLB^{+}21]. Bahamas [OVO^{+}21, SC23]. Bahía [FPZJAO^{+}22]. bailout [TP20]. bait [HRH22]. baited [BWGH23, FWGD^{+}23, WYZ^{+}20]. Baja [BGMAM^{+}21, CEAO^{+}22]. Baker [Whi21]. Balaena [PDF^{+}23]. Balaenoptera [BMS^{+}21]. balance [BP24, RSS20, ZCF20]. Baldanzi [Ano21a]. baleen [BKR^{+}21, FNL^{+}23]. Bali [NBW^{+}23]. Baltic [HOP^{+}20, KBWRV21, KBE^{+}22, KSMP20, LCP^{+}23, MGB^{+}22, UAL24, Vir23, ZLL^{+}24]. Banc [EHLM^{+}23]. Banderas [FPZJAO^{+}22]. Bank [MMB22, JMR21, MSL^{+}24, RBF^{+}20]. banks [SPFL20]. barbatus [PCLQ^{+}20]. barcoding [TKM^{+}22, VFG24]. Barents [EGK^{+}21, FSN^{+}20, FLY^{+}24, BAAR20, DKSS20, DKSS21, GBSS24, HLA^{+}21, OSJ^{+}22, SS24]. Barile [Ano22a]. barnacle [AR21b, LBH^{+}21b, NDA^{+}22, RWM^{+}22, SAC24, SGP24]. barnacles [LT21]. barnesi [SAC24]. barren [WK21]. barrens [DKS^{+}24]. Barrier [CSDB22, MHK22b, MRH20, RCES22, SMH^{+}20, SYMR21, AR21b, RS21]. barriers [BH22, BNSH21, LMH^{+}24]. basal [RLP^{+}23]. base [EHHF^{+}20]. based [BBS^{+}21, BMPD20, EVVMQ^{+}23, ESG20, FDL^{+}21, GLG21, GB20, HSU^{+}21, HvLM21, HBG^{+}22, LDW^{+}21, LBMSS23, MCL24, MSC^{+}22, MAK^{+}22, MGA^{+}20, MDG^{+}21, NGPGO^{+}23, ORB^{+}21, OAdJA^{+}20, OCR^{+}24, PLH^{+}21, PS23, PHV20, RHN^{+}24, RRR^{+}23, RTG^{+}20, SBS^{+}24, SLH^{+}22, TD21, UIS^{+}24, VDC^{+}24]. baseline [MSTH22]. baselines [FBC^{+}24a]. Basin [CKM^{+}20, HGT^{+}21, FLC^{+}24b, HEKH24, WWH^{+}22]. basin-scale [FLC^{+}24b]. Basin-wide [HGT^{+}21]. basins [OVÁS^{+}21]. basis [HESF24]. bass [BSRVS22, RWR24, TPGG^{+}23]. bassanus [dDW^{+}22]. bat [CLL^{+}22]. bathyal [LFK^{+}21]. bathymetric [LFK^{+}21]. bathypelagic [MPR^{+}23]. Bathyraja [SOS^{+}21b]. batoid [SVT^{+}24]. Bay [BASM^{+}24, CCW21, DNE^{+}24, DMBB22, HLCH23, HKD^{+}20, OAM^{+}24, PBAC23, PDFH20, RH20, SRHM21, SRG^{+}24, SNH^{+}20, TD21, WLB^{+}20, ZSC^{+}22, AAHW20, GMB^{+}24, STF^{+}24, BBL^{+}21a, CDD^{+}20a, DFHT23, GRTK^{+}21, HDL^{+}21, HCS^{+}24, KTR^{+}20, OLA^{+}24, RGL24, SST^{+}20a, SHG^{+}21, SFL23, SLB^{+}21, TCT^{+}23, vdKMC^{+}24]. Bayesian [BSH^{+}21b, KKS^{+}20, RPB21, WYO^{+}20]. bays [LOF^{+}24, MVH^{+}21]. be [Ano23, DDB^{+}24, WH24]. Beach [SNH^{+}20, BNSH21, BNS^{+}24, CGL^{+}22, GSL^{+}24b, GSBO21, LFA^{+}23, RPD^{+}23]. beach-cast [GSBO21]. beaches [BSA^{+}24, SMB^{+}20, TQG24]. Beagle [DRP^{+}20, RBF^{+}20]. beaked [BSFM20, BBSM^{+}20, FPB^{+}21, SBS^{+}22, SSF^{+}22]. beaks [WWH^{+}22, vTLG^{+}21]. bear [BBL^{+}21a, FWA^{+}23, HDL^{+}21, KTR^{+}20, LAP^{+}24]. bearded [OBD^{+}21]. bearing [BAS21, CGW21]. bears [AAP^{+}23, BAAR20, SLB^{+}21]. beats [Ano23]. Beatty [Ano22b]. Beaufort [CGR^{+}20, OBD^{+}21, SFF^{+}20]. bed [KKN^{+}20]. beds [AMD22, KHOK22, MDH^{+}20, MHK22a, PDM22, QMDGM22, STH21, SGP24]. beetle [MGBCGM^{+}22]. before [EWBH24, MLE^{+}22, VGA20]. Behavior [HDA^{+}22, AKU^{+}22, BK23, BSFM20, BKTN21, CKKL21, dOFBSdJR^{+}22, FNO^{+}23, GMLC22, GBF^{+}23, LGR^{+}22, LLJ^{+}20, LCP^{+}23, MPM21, MLE^{+}22, MJJJ20, MAH^{+}20, MAMF21, MCH^{+}21, RB23, SRHM21, SJSR24, SSB^{+}23, SHL^{+}21, TCP^{+}23, TRK^{+}22, WHB^{+}20, WLC22]. Behavioral [CPH20, LKR^{+}21, AWC21, ASS^{+}21b, BH22, BMS^{+}24, BQC^{+}22, GFH^{+}22, JBG22, MDS^{+}21, SJQ^{+}22, SON^{+}20, WGK22]. behaviors [BSF^{+}20, CZS^{+}22, HDM^{+}24]. Behaviour [NSB^{+}23, Ano23, BLCC^{+}20, BKA^{+}20, BOY^{+}23, DFRS^{+}24, EBL^{+}21, EMDR22, FTA^{+}22, FDS^{+}21, FFT^{+}20, FJB^{+}23, FCV^{+}24, KPW^{+}20, KCB^{+}24, MJ23, MPSH22, PGT^{+}21, PJA^{+}21, PBT^{+}21, RBW^{+}20, SWR23, SD22, TNMN24, VAMF22]. Behavioural [DJR^{+}20, HJG21, BMBC21, BRSD20, CCA^{+}20, DCR^{+}22]. behaviours [CGGdD24]. being [Ano24o]. belauensis [CASF21]. bell [RS20b]. below [KPS^{+}22, SWHE20]. beluga [CGR^{+}20, KTW^{+}22]. benefit [YDG21]. Bengal [SST^{+}20a]. Benguela [CEG^{+}21, DSHF22, DFSH23]. Benthic [BMK^{+}22, BSA^{+}24, HDB21, Mon23, MPKS24, BQM^{+}24, CAEG21, CSH^{+}21b, COWM^{+}22a, DNJCH21, DMBB22, EB20, FSSL^{+}20, FP20, GVR^{+}20, HOP^{+}20, HGT^{+}21, JDDF21, LR22, MCK21, MBT^{+}20, MCF^{+}21, MDG^{+}21, MG22, MBR^{+}21, PAF^{+}22, RF20, SBT24, SOSE22, SLO^{+}22, SE20, SB20b, Vir23, WIY^{+}23, ZBB^{+}23, Zue22, vDTS^{+}22, vdRJL20]. benthic-pelagic [JDDF21, RF20]. benthivorous [RCB^{+}24a]. benthopelagic [PBT^{+}21]. benthos [GBC^{+}20]. Bering [CSK^{+}22, FD21, GGT^{+}20, GBA24, LEG^{+}20, OLM24, OBD^{+}21, SOS^{+}21b, SOS^{+}21a, TEK21]. Bermuda [CVG^{+}23]. Bertalanffy [RPB21]. best [GBW^{+}20]. beta [AS24, AHMVM23, AC20, CATF^{+}21, Vir23]. Better [DPDJ^{+}21, Ano23, SJA^{+}21, SAGG^{+}24, SDE^{+}21, WRB21]. between [AS20, Ano23, ARdlB^{+}22, AAFC^{+}21, BBEF^{+}20, BTMS24, BQM^{+}22, BBSM^{+}20, BDS^{+}20, CAV^{+}20, CLD23, CRG^{+}20, CASF21, CILGD^{+}22, CSN^{+}23, DCK^{+}24, DDR^{+}21, DBP^{+}22, DKSS21, DAT^{+}21, DCL21, EVAB^{+}23, FJJ^{+}21, FSK^{+}21, GDAC23, GMLC22, HMB^{+}22, HIT^{+}22, HSBB^{+}21, HOK^{+}20, HEP^{+}24, JNJ^{+}20, JSK19, KKN^{+}20, LPJ^{+}20, LMRS^{+}24, LMS^{+}20, LCJ22, LAGGM^{+}21, MdOO^{+}20, MHH^{+}20, MAOR^{+}21, MGBCGM^{+}22, MHK22b, MHH^{+}22, NTN^{+}24, OLA^{+}24, OYO^{+}20, PRM^{+}21, PGT^{+}21, RSvL21, RH23, RBVA23, SKTO23, SMA^{+}20, SVS^{+}22, TCY^{+}22, YHW^{+}23, ZZX^{+}20, ZZLS24, ZCF20, vdVWF^{+}24]. Beyond [STY^{+}20, TPZ^{+}21, Ano24o]. bi [EGK^{+}21]. bi-directional [EGK^{+}21]. biased [PFM^{+}21]. biases [OCR^{+}24]. bicyclis [AAI^{+}24]. Bight [FSS23, FST22, GNC^{+}23, LLF^{+}20]. Bijagós [EHLM^{+}23]. billed [DCK^{+}24, FDE^{+}21]. billfishes [BGDJ20]. Bio [MTRM21, CVG^{+}23]. bio-logging [CVG^{+}23]. Bio-physical [MTRM21]. Biochemical [NGRH20, GVOCSGG21]. biodeposits [PRD^{+}20, PBC^{+}22, SHZ^{+}20]. Biodiversity [HSSN^{+}21, CSWM24, ECC^{+}21, HTL^{+}20, TAN^{+}22]. bioeconomic [NSKH23]. Bioenergetic [RL22, AU21, FPG^{+}21, SJ22]. bioenergetics [HTT^{+}21, HCvdHM23, MGM^{+}24, RHN^{+}24]. bioenergetics-based [RHN^{+}24]. biofilm [MMLPP^{+}24]. biofilms [KPS^{+}22]. Biogenic [YAZA21]. Biogeographic [AHMVM23, MYSF23, AR21b, WSB^{+}24]. Biogeographical [JHDE20]. Biogeography [PGC21, AC20, DMM^{+}24b, RGK22, SS24]. Biogeophysical [LFK^{+}21]. biologging [Ano23, TCY^{+}22]. Biological [DFRS^{+}24, FNH^{+}23, LSS^{+}22, MDG^{+}21, WK21, AHMVM23, ALL^{+}23, BMB24, EAS20, KTK21, MCH^{+}21, UPK^{+}20, WAA24]. biology [ZCB20]. biomarkers [RHSM21, SDM^{+}22, SFL23]. Biomass [TKP^{+}20, AZ24, BLB23, BBKW20, BHQ^{+}22, CGH^{+}23, DHA^{+}20, GB20, ILON21, KSS^{+}20, LHD20, LGD23, LEG^{+}20, LHHH^{+}20, RSvL21, RS21, SS24, SHT^{+}22, TABM21, Vir23]. biomechanics [GPCM22]. biomedical [WLC22]. biomimicry [MDS^{+}21]. Biophysical [NDA^{+}22, SBJ^{+}24]. biopsies [LAP^{+}24]. bioregions [TMN^{+}24]. biosynthesis [HMT21]. biotic [BMO^{+}24, MPKS24]. bioturbators [VST24]. bird [CILGD^{+}22, JPL^{+}24, SAF^{+}24]. bird-borne [CILGD^{+}22]. birds [GBF^{+}23, RCB^{+}24a]. Birgus [ACvdC^{+}21]. birostris [FPZJAO^{+}22, HGK^{+}22, LGGR^{+}22]. birth [WTK^{+}24]. Biscay [GRTK^{+}21, vdKMC^{+}24]. bivalve [CG22, CMTB^{+}22, FOM^{+}20, KCLJ20, SW23a, SFOPL21, TR24, WC21, WLB^{+}20]. bivalves [AM20, GON^{+}20, SVS^{+}22]. bivoltine [TBS20]. Black [KKMW23, MMA^{+}22, BSB^{+}21, CCL^{+}22, EGK^{+}21, JFR^{+}21, JTBS^{+}22, LPD^{+}24, MIL^{+}20, RBF^{+}22, TPGG^{+}23, VLG^{+}21]. black-backed [JTBS^{+}22]. black-browed [VLG^{+}21]. black-legged [BSB^{+}21, EGK^{+}21, MIL^{+}20]. Black-tailed [KKMW23]. blacklegged [OBW^{+}20]. blacktip [AKGA^{+}21]. Bleaching [SCK23, BMP^{+}22, DeC24, GED20, JBDH21, MDM^{+}20, SSFD^{+}23]. bleeding [WLC22]. blennoides [ASGI24]. blockage [ADF20]. bloody [EHLM^{+}23]. bloom [DL20, DKSS20, DHA^{+}20, GGRV23, LWP^{+}24, MCL24]. blooms [BAS23, DKSS21, DMBB22, HWT^{+}22, LAL21, OSJ^{+}22, YOI^{+}23]. Blubber [KTW^{+}22]. Blue [AZ24, MLVJ23, PJA^{+}21, TFCS22, WLB23, BBEF^{+}20, BBGM23, DCK^{+}24, EMDR22, FSH^{+}21, FKS^{+}22, HAB^{+}22, HCS^{+}24, KSW23, LSSL21, LSS^{+}22, MAOR^{+}21, RGL24, RGK22, SRK^{+}23, SFL23, ZBB^{+}20]. bluefin [AOO^{+}21, BKH^{+}24, DBHCO^{+}22, FSK^{+}21, HKY^{+}20, HIT^{+}22, MMMV22, UIS^{+}24, VMDA22]. Body [CRB^{+}21, IMM^{+}20, LCFJ20, RCC^{+}22, BAS21, CKT^{+}24, CDD^{+}20b, CRGMA^{+}21, DFP^{+}21b, EBL^{+}21, HHY^{+}23, LCP^{+}20, LKB^{+}21, OFSJ20, RBS^{+}24, SDE^{+}22, SWHE20, ZZLS24]. bomb [FLC^{+}24a]. boobies [CAW^{+}22, CILGD^{+}22]. booby [CBWH24]. boom [HHO^{+}23]. boom-or-bust [HHO^{+}23]. Boosted [CMF21]. boreal [CLC^{+}23]. borealis [BMF24]. Boreogadus [BLW^{+}24, DDG^{+}21, MSC^{+}22]. Boring [CGLC20]. borne [CILGD^{+}22]. Both [KPTT23, Ano24o, MJ23]. bottlenose [BRD20, CCA^{+}20, DFHT23, GG23, TPZ^{+}21, TRK^{+}22, WIM23]. Bottom [KNO^{+}21, DGM^{+}23a, GG23, HGT^{+}21, LHFD^{+}24, MDG^{+}21, Sig23, THD^{+}22, vDTS^{+}22]. bottom-set [GG23]. boundaries [MDQ^{+}21]. boundary [Hew21, MMD20]. bouquet [ZLHH20]. bouts [RBW^{+}20]. bowhead [FFT^{+}20, PDF^{+}23]. box [CCL^{+}22, RSS20]. box-balance [RSS20]. brachyuran [MCH^{+}21]. brackish [NGRH20]. Bransfield [ZZLS24]. Brazil [BQM^{+}24, BZT^{+}23, GFDN^{+}24, LCR^{+}22, LTM20, PBAC23]. Brazilian [ARN^{+}22, KVdM^{+}22, MdMR24, TABM21]. breaching [MAMF21]. break [BMOW21]. breakup [HHM^{+}22]. breakwater [MCK21]. Bredning [MKS^{+}20]. breeders [PCP^{+}20]. Breeding [GLW21, ODB^{+}24, OOW^{+}20, BFQ^{+}23, BLCC^{+}20, BBG^{+}22, BHQ^{+}22, CAW^{+}22, CKT^{+}24, CGS^{+}23, CFSH22, DdGA^{+}21, DCK^{+}24, DWJL^{+}21, EVAB^{+}23, FTA^{+}22, GQBB^{+}24, GJM^{+}20, GBF^{+}23, JW22, LBTE^{+}21, LLJ^{+}20, MTCT^{+}23, NCS^{+}23, OBW^{+}20, PLLT^{+}21, PRM^{+}21, PGT^{+}21, PHE^{+}20, RFJC^{+}24, RWH^{+}24, RLME^{+}21, RMD24, SSC^{+}23, SGB^{+}21, STO^{+}23, SDF^{+}21, SWHE20, WdGR^{+}20, WLA^{+}21, dSGC^{+}23]. brevis [CP22a]. Bridging [SAGG^{+}24]. bring [Ano23]. brings [Ano24o]. Britain [DHE^{+}23]. British [KSS^{+}20, JPJH23, MAK^{+}22, RBC^{+}24]. Broad [GMS^{+}21, JMP^{+}24]. Broad-scale [GMS^{+}21, JMP^{+}24]. broadcast [HHO^{+}23]. Brood [SJSR24, HB24]. Brood-grooming [SJSR24]. brooder [BNSH21]. brooders [HHO^{+}23]. brooding [BSF^{+}20, CGW21, MK20]. broods [BAS21]. browed [VLG^{+}21]. brown [BMPD20, CAW^{+}22, CDS23, DB21, GDAC23, LCP^{+}23]. BRUVS [WYZ^{+}20]. bryozoan [AMD22, PDM22]. bryozoans [PEBG^{+}20]. bubble [ANP^{+}24]. bubble-net [ANP^{+}24]. budgets [CSBM20, CCA^{+}20, LHW^{+}20, SBDM^{+}23]. Buena [THD^{+}22]. buffers [EB21, SHH^{+}22]. building [CTTLCG^{+}24, JH20, PFM^{+}21, SDM^{+}23]. Bulk [PSN^{+}20, LNBdB21]. bull [CMF21, EMB^{+}22, LBMF^{+}24]. Bullia [BNS^{+}24]. Burdwood [MSL^{+}24, RBF^{+}20]. burial [AZ24, XWW^{+}21]. burrow [LRC20]. burrowers [WRB21]. burrowing [DBK^{+}23]. burrows [KVdM^{+}22]. business [MBT^{+}20]. bust [HHO^{+}23]. Bycatch [MPF^{+}22, BLCC^{+}20, CWC^{+}22, FNM^{+}21, FRO^{+}21, FCV^{+}24, GRTK^{+}21, MDW^{+}22, RTG^{+}20, Sig23, WSB^{+}24]. Byssal [RC23]. byssus [CG22].

C [ARB^{+}20, BCC20, HRH22, LAP^{+}24, MBW^{+}20, SST^{+}20a, ZMD^{+}21]. C. [BSH21a, HEKH24, MHN24]. CA [HGK^{+}24]. cable [LRC20, YAZA21]. cage [MVH^{+}21, MVC^{+}23]. cahow [CVG^{+}23]. calanoid [JP20]. Calanus [CLC^{+}23, HSU^{+}21, HEKH24, MHN24, RRR^{+}23, TEK21, VKdJ24, WDP^{+}21, WGK22]. calcification [EP23, GPCM22, VLPN24]. Calcified [HM23]. calcifying [SWP20]. calcium [CRLC21]. Caledonia [APG^{+}22, WdGR^{+}20]. calf [ASS^{+}21b, PBS^{+}24]. calibration [BLVL24]. California [AMMADDH20, BGMAM^{+}21, CEAO^{+}22, FST22, GNC^{+}23, KDB^{+}24, SRHM21, SSB^{+}20, AMUPJR^{+}24, AKGA^{+}21, BWRP20, BLGM^{+}24, CSK^{+}23, CFB21, DFB^{+}24, GBW^{+}20, GWA^{+}23, GTH^{+}20, JW22, LO20, LNM22, MJ21, MPB^{+}24, NT24, PRC^{+}21, RJE^{+}22, SHB^{+}20, SE20, YLH20, Zue22]. Californian [CCJ^{+}23]. californianus [LNM22]. California’ss [BSRVS22]. call [RBB^{+}21]. callichirid [SC23]. Callinectes [HAB^{+}22, LSSL21, SRK^{+}23, TFCS22, ZBB^{+}20]. calling [LGR^{+}22, MPM21]. calls [MHH^{+}22]. caloric [HM23]. calves [MBW^{+}20, SDE^{+}22]. camera [DCR^{+}21, SMK23]. cameras [CILGD^{+}22]. campechanus [BDS^{+}20, DSK^{+}22, ESG20, MPKS24]. Camps [Ano20]. can [Ano24o, BML^{+}22, BH22, CBGB24, DDB^{+}24, LSSL21, NP22, NMB^{+}22, SGW^{+}21, YBSH23]. Canada [KSS^{+}20, MCL24, BMF24, BOY^{+}23, BKH^{+}24, FFT^{+}20, FRO^{+}21, JPJH23, LGD23, MAK^{+}22, MVH^{+}21]. Canadian [DFHT23, HYG^{+}20, KTW^{+}22, LHFD^{+}24, MDW^{+}22]. canaliculus [DBS^{+}20, STH21]. Canary [FBC^{+}24b, VMDA22]. Cancer [HRH22, LO20, RJE^{+}22]. candidate [ZLHH20]. cannonball [FSS23]. canopies [BPSB^{+}22]. canopy [BTB^{+}24, FDF20, KKN^{+}20, VDG24]. canopy- [FDF20]. canopy-forming [BTB^{+}24, VDG24]. Cantabrian [AGIS^{+}22]. canyon [HOK^{+}22]. Cap [SMS^{+}24, SNE^{+}24]. capacity [BKT^{+}22, CPD^{+}20, SME^{+}23, UBO^{+}24, WD22]. Cape [BKA^{+}20, DV20, FPB^{+}21, MSC^{+}20]. capelin [FLY^{+}24, MGA^{+}20]. capensis [vdHMA^{+}24]. capitata [HBR^{+}21, RHM^{+}22]. Capitellidae [HKD^{+}20]. captive [DBP^{+}22]. capture [PLC^{+}20, SVS^{+}22, WFS23]. capture-recapture [WFS23]. captures [OLA^{+}24]. Caranx [DFP^{+}21a, DMM^{+}24a, GDCP20, GFDN^{+}24]. carbohydrates [LDM^{+}20]. Carbon [KBP22, PFDF^{+}21, SBDM^{+}23, ALL^{+}23, AZ24, CEAO^{+}22, FCSG21, HMT21, KGG^{+}21, LOS^{+}22, LH22, LRD^{+}20, LHW^{+}20, OCR^{+}24, RLP^{+}23, SFF^{+}20, SMPLF^{+}22, UAL24, WLB23]. carbonate [CSBM20, CRLC21]. carbonation [CEZH21]. carcasses [FCSG21, SSP^{+}22]. Carcharhinus [AKGA^{+}21, EMB^{+}22, LBMF^{+}24, MSH20, TCM^{+}23]. carcharias [LBH^{+}21c, SMB^{+}20]. Carcharodon [LBH^{+}21c, SMB^{+}20]. Carcinus [CRB^{+}21, CGM23]. Caretta [CTH^{+}20, MAH^{+}20]. Caribbean [NMAE^{+}22, AGA^{+}22, ADBF21, BBRM22, CTTLCG^{+}24, CDF^{+}20, CBJ^{+}23, EP23, FBP^{+}22, GBE^{+}22, KDLL21, LCP^{+}20, LvAH^{+}24, MPA^{+}22, MJJJ20, OMW^{+}20, PAF^{+}22, SMPLF^{+}22, SdGWN21]. carnivorous [CCBT^{+}20]. Carolina [AKW^{+}22, FPB^{+}21, MMS^{+}22, GLP^{+}24]. carrion [BHQ^{+}22, TABM21]. Carryover [MK20]. cascades [CSH^{+}21a, EB20, RS20a]. case [CTH^{+}20, GFDN^{+}24, HRC21, KTK21, LFK^{+}21, LNKJ^{+}22, PMH23, PCLQ^{+}20, RJ21, WHN^{+}22]. casitas [RBM22]. cast [GSBO21]. Castellà [Ano20]. CATAIN [SMK23]. catch [BBKW20, ZCF20]. Catchability [BS20a, SBL^{+}23, TdLHL21]. catervarius [MGA^{+}20]. Catomerus [AR21b]. caught [ABN^{+}21]. Caulerpa [AS20, OVÁS^{+}21]. Causal [RSvL21]. cause [BH22, PRC^{+}21]. cause-and-effect [PRC^{+}21]. caused [BSM23]. causes [LCT^{+}20]. cauta [MHAL24]. Cautious [EPHS23]. cavirostris [FPB^{+}21]. Cayman [WSPS^{+}20]. cell [HvdSR^{+}21, LAGGM^{+}21]. census [VDC^{+}24]. centennial [KBE^{+}22]. central [CLD23, CCJ^{+}23, CCR^{+}20, FD21, SMA^{+}20, YLH20, ZFD^{+}24, Zue22, BWRP20]. centre [SCM21]. Centropomus [AWC21]. Centrostephanus [DKS^{+}24]. century [BD20, FDL^{+}21, JBWS24, WFR^{+}23, vdVTW^{+}22]. Cepphus [JW22, JFR^{+}21]. Cerastoderma [dMGGG21]. cervicornis [IBRL^{+}22, YRB^{+}20]. cetacean [DPDJ^{+}21, DCL21, LTM20]. cetaceans [ISMM20, MTB^{+}23]. CF [HKY^{+}20]. CF-IRMS [HKY^{+}20]. cf. [DDB^{+}24]. CH [KVdM^{+}22]. chain [BBEF^{+}20]. chains [Rob21]. Challenges [OA24, Ano24o, BBBT^{+}20, MAS^{+}20]. challenging [NBW^{+}23]. chambers [CAEG21]. change [ARE^{+}24, Ano24o, AR21b, BKM^{+}21, BSM23, BRU^{+}22, BBBT^{+}20, CSK^{+}22, CDP^{+}21, DHE^{+}23, FNM^{+}21, GBW^{+}20, GSL^{+}24a, GBE^{+}22, HEP^{+}24, JETGM24, KDK^{+}24, LCFJ20, LBMB^{+}20, MBG^{+}23, OBD^{+}21, OA24, RHN^{+}24, SFOPL21, TAN^{+}22, TNS^{+}22, VLG^{+}21, vLSB^{+}21, vLBGGT22]. Changes [BMP^{+}22, GWA^{+}23, ICIS23, OLPGK22, TSS^{+}23, WHB^{+}20, vdVTW^{+}22, ARdlB^{+}22, BRC22, BDJ^{+}24, BPS^{+}24a, BSÁG^{+}22, CAV^{+}20, CMT^{+}21, CGR^{+}20, CHJ21, CMTB^{+}22, DV20, FWWH22, GGT^{+}20, HKK^{+}23, HGH^{+}24, HNKK21, KGM^{+}24, KI24, LÁGLL20, LRM^{+}21, MLB^{+}21, MW20, NSPH20, RVSA^{+}21, RDL^{+}21, SHG^{+}21, SBW^{+}22, SFLQ23, TTM^{+}22, TTM^{+}24, WYO^{+}20, WGK22, ZFD^{+}24]. Changing [SSBA^{+}20, AESS24, Ano23, CLF^{+}20, DB21, FYM^{+}24, MSR^{+}21b, PGGE21, Rus20, RBCwD^{+}20, WYO^{+}22, WMKV21, YSD^{+}21, dDW^{+}22]. channel [NCZ^{+}21, DRP^{+}20, RBF^{+}20, vdKMC^{+}24]. channel-mediated [NCZ^{+}21]. Channelling [RLP^{+}23]. char [HHM^{+}22, HYG^{+}20]. characterisation [RRS^{+}23]. Characteristics [LHM^{+}23, Ano23, KDLL21, KWD21, KHG^{+}21, PGW23]. Characterization [MMB22, PAAEV^{+}24, AHCT22a, MMS^{+}20, SLG^{+}24, VOP21]. Characterizing [GTH^{+}20]. Charles [LGM21]. Charonia [SMH^{+}20]. charr [CDP^{+}21, NSB^{+}23]. Chelonia [SBC^{+}21]. Chemical [VMCA^{+}20, AGA^{+}22, SHZ^{+}20]. chemistry [AAFC^{+}21, DMTP22, JHK^{+}22, KNH24, PGM^{+}23, SRS^{+}20, ZXL^{+}24]. chemosynthetic [MON^{+}24]. Chesapeake [AAHW20, HCS^{+}24, OLA^{+}24, RGL24, SFL23, SRG^{+}24]. chick [BDC^{+}22, LGD23, MHAL24, RBW^{+}20]. chick-rearing [BDC^{+}22, RBW^{+}20]. chicks [EVVMQ^{+}23]. Chile [HSA^{+}21]. Chilean [BSF^{+}20, vLSB^{+}21]. China [HLCH23, NYW^{+}22, YWW^{+}24, WLB^{+}20, ZZX^{+}20, ZSC^{+}22]. Chinook [ATLT20, BAB^{+}24, CSK^{+}23, DCG^{+}20, NMB^{+}22, RS20b, SBJ^{+}24, WDJ20]. Chironex [MJK24]. chitobiase [KSS^{+}20]. chiton [IBC^{+}21, MB20]. chlorophyll [OÁBP^{+}21, RVCT21]. chlorophyte [OVÁS^{+}21]. chloroticus [SST20b]. choice [dOFBSdJR^{+}22]. Chondrus [FLCQ21]. chorusing [VAMF22]. chronic [ASS^{+}21b, BG21, SW23a]. Chronically [COWM^{+}22a]. chronologies [FLC^{+}24a, JBGA23]. chronosequences [AZ24]. Chrysophrys [CHJ21]. chrysostoma [SSC^{+}23]. Chthamalus [SAC24, WRS^{+}23]. Chukchi [OBD^{+}21]. chum [FYM^{+}24, LAD^{+}21]. circatidal [TWC20]. circumpolar [FTC^{+}20, WLA^{+}21]. citizen [FFMR24, NJHN24, NRE21]. cladocerans [HLCH23]. clam [CGL^{+}22, GSL20, LCJ22, LSS^{+}22, MNG^{+}20, RPD^{+}23, SCK23]. clams [RVV^{+}24]. classes [BJH^{+}23, GBSS24, MPB^{+}24]. clavata [SLG^{+}24]. cleaner [GBC^{+}20]. clearance [AM20, DPCL^{+}23]. cliff [IHA^{+}22]. Climate [BKM^{+}21, BBBT^{+}20, LPO^{+}21, PRC^{+}21, PHE^{+}20, Tha21, TTA20, ZBB^{+}20, vLSB^{+}21, ARE^{+}24, Ano24o, AR21b, BPS^{+}21, BAH^{+}22, BRU^{+}22, CLD23, CSK^{+}22, CLF^{+}20, DHE^{+}23, FYM^{+}24, FNM^{+}21, FYK^{+}20, GBW^{+}20, GBE^{+}22, GBA24, JSDH20, JETGM24, KDK^{+}24, LÁGLL20, LCFJ20, MHAL24, MSR^{+}21b, MBG^{+}23, OA24, PGGE21, RLME^{+}21, RHN^{+}24, SBW^{+}22, SS24, SFOPL21, TSW^{+}21, WAA^{+}23, WYO^{+}22, WMKV21, WCRTT20, vLBGGT22]. Climate-associated [TTA20]. Climate-change [BBBT^{+}20]. climate-driven [BAH^{+}22, CLD23, SBW^{+}22]. climate-mediated [Ano24o, WAA^{+}23]. climates [AVFD^{+}21, GJK^{+}24]. climber [THD^{+}22]. closed [HNB^{+}24]. closer [MKR^{+}20]. closure [MLE^{+}22]. Clupea [BSF24, FN24, HGH^{+}24, MGM^{+}24, RBC^{+}24]. cnidarian [JPB21]. CO [GMB^{+}24, MGM^{+}24, LCJ22, BMKB22, CGM^{+}24, JNJ^{+}20, MHN24, MGBCGM^{+}22, RH23, RBCAV22, SCM^{+}23, WCS^{+}21, WD22, WF21, YMU^{+}21, BJM^{+}22, HRC21, JMS^{+}22, KVdM^{+}22, MR20]. co-distributed [YMU^{+}21]. co-existing [JNJ^{+}20]. co-infection [RBCAV22]. Co-infesting [LCJ22]. co-limitations [WCS^{+}21]. co-occurrence [CGM^{+}24]. co-occurring [BMKB22, MHN24, MGBCGM^{+}22, RH23, SCM^{+}23, WD22, WF21]. Coast [BWRP20, AAI^{+}24, BFQ^{+}23, CVP^{+}21, DKAB23, FRO^{+}21, GMRMG22, KNO^{+}21, LBMF^{+}24, MLP^{+}24, OBS^{+}22, OCC23, PRM^{+}21, SBS^{+}24, TTA20, WBG^{+}22, ARN^{+}22, LBH^{+}21a]. Coastal [MLP^{+}24, RAHM^{+}21, ASBA^{+}20, BVD^{+}20, BG20, BLGM^{+}24, BSP^{+}22, CATF^{+}21, CAEG21, CRB^{+}21, CEG^{+}21, CRG^{+}20, CFB21, CHDG^{+}23, CSB24, CSWM24, EB20, FFMR24, Fig21, FWGD^{+}23, GCDA20, GVS23, GBG^{+}23, GCGCAS^{+}20, GDCP20, GON^{+}20, GLM^{+}24, HGR^{+}22, HGK^{+}22, HO23, HMW^{+}21, IAG^{+}23, ISMM20, JPJH23, JBWS24, JHK^{+}22, KCLJ20, KKMW23, LPO^{+}21, LOF^{+}24, LGD23, LTM20, MAK^{+}22, MYSF23, MVH^{+}21, MNG^{+}20, NSPH20, NCZ^{+}21, OLG^{+}23, PWvdV20, PWvdM^{+}21, PDH^{+}24, RDT^{+}23, RBFA20, SDM^{+}22, ŠGH^{+}21, STM^{+}24, SMB^{+}20, SJH^{+}22, TWC22, UBO^{+}24, UAL24, VS23, WXZ^{+}24, WLB^{+}20, XCT^{+}20, YOI^{+}23, ZXL^{+}24, dlBSR^{+}20, vdVTW^{+}22, vdVWF^{+}24]. coastline [TABM21]. coasts [CKCS20]. cockle [FSF^{+}24, dMGGG21]. cockles [EHLM^{+}23, HESF24]. coconut [ACvdC^{+}21]. Cod [DV20, BLW^{+}24, BSM23, CDD^{+}20a, DDG^{+}21, FLY^{+}24, GBG^{+}23, HOP^{+}20, JNJ^{+}20, KSW23, KSMP20, KTK21, LNKJ^{+}22, LCGS^{+}21, MSC^{+}22, MVH^{+}21, MVC^{+}23, MGB^{+}22, RCB24b, SBS20, SJH^{+}22, MSC^{+}20]. Cod/Gulf [DV20]. coexistence [BVKF23]. coexisting [EVVMQ^{+}23]. cohabiting [CMdL^{+}20]. Coho [ATLT20, BAB^{+}24, FSDB20, JNY^{+}20]. Cohort [MSV21]. cold [AMUPJR^{+}24, CKS^{+}24, GSC^{+}22, GED20, LEG^{+}20, NSM24, QHT^{+}21]. cold-water [CKS^{+}24, GED20]. colder [OHBJ20]. colder-water [OHBJ20]. Collapse [TCT^{+}23, LCT^{+}20, MRB^{+}24]. collected [LS21, TBH21]. collection [BML^{+}22, KCGR20]. Cololabis [FNH^{+}23]. Colombia [CRLC21]. Colombian [EMGMEV^{+}21]. colonial [DL21, JPB21]. colonies [CFE^{+}22, MKR^{+}20, WPG^{+}24, WdGR^{+}20, WLA^{+}21, YRB^{+}20]. colonisation [BICQG23]. colonization [FLCQ21, GCGCAS^{+}20, SNE^{+}24, WSU21]. colony [BFQ^{+}23, BBG^{+}22, CAW^{+}22, CKT^{+}24, EGK^{+}21, GJM^{+}20, JPMG23, LMHS21, MAS^{+}20, TWC^{+}23]. colony-induced [MAS^{+}20]. Coloration [DHM21]. colour [CRG^{+}20]. columba [JW22]. Columbia [JPJH23, KSS^{+}20, MAK^{+}22, NMB^{+}22, RBC^{+}24]. column [DNE^{+}24, OJPC23]. combat [SBC^{+}21]. combination [ECC^{+}21, HHBP23]. Combined [BSF^{+}20, BKT^{+}22, KSW^{+}21, MIL^{+}20, RCES22, SBSS21, LOF^{+}24]. Combining [CAP^{+}21, CVG^{+}23, MDW^{+}22, CGS^{+}23, DCR^{+}22, BPS^{+}24b]. commensalism [IBC^{+}21]. Comment [BS21, KCT24]. commercial [CLM^{+}21, DdGA^{+}21, FDS^{+}21, FRO^{+}21, HRH22, JLT^{+}21, TAN^{+}22]. commercially [GTH^{+}20, HMW^{+}21, PBO^{+}21]. Common [FMD20, KKS^{+}20, AWC21, CSB24, FNS^{+}20, GRTK^{+}21, HESF24, MGHH21, MPF^{+}22, PBB^{+}20, RMD24, SFMK^{+}23, TPZ^{+}21, WLA^{+}21]. communities [BQM^{+}24, BPS^{+}24a, BGL^{+}21, BTB^{+}24, BAS23, CC22, CKM^{+}20, CSH^{+}21b, CEAO^{+}22, COWM^{+}22a, ENRC^{+}20, FS22, FWGD^{+}23, GVR^{+}20, GPMH^{+}24, GMCE^{+}22, GSN^{+}20, HGT^{+}21, HO23, HSBB^{+}21, HDE21, KLN^{+}21, LDM^{+}20, LHFD^{+}24, MCK21, MJW24, MBR^{+}21, MLAS22, MW20, NSM24, PAAEV^{+}24, RSEG21, RDT^{+}23, SBDM^{+}23, SNP20, SIBM20, TSS^{+}23, TKP^{+}20, VIV^{+}20, WXZ^{+}24, WIY^{+}23, WYO^{+}22, WCRTT20, WCP^{+}21b, XCT^{+}20, XYSF20, ZHG^{+}21, ZCM^{+}21, vDTS^{+}22]. Community [ACG22, KF20, AHH^{+}20, Ano24o, ASS^{+}23, BWRP20, BTA^{+}20, BWGH23, BDJ^{+}24, BMPD20, CAEG21, CDF^{+}20, CIDM23, CGS^{+}23, CSD^{+}24, DPDJ^{+}21, DHA^{+}20, DMBB22, EWF^{+}21, Fig21, FP20, FD21, GBW^{+}20, GCDA20, GWA^{+}23, HSB20, HSSN^{+}21, HDB21, HNKK21, LS21, LCFJ20, MKS^{+}20, MFC^{+}24, MAOR^{+}21, MTSO^{+}24, OJ24, PWvdM^{+}21, PRD^{+}20, RDL^{+}21, RWH^{+}24, SMK23, SFLQ23, SLO^{+}22, SE20, TGT^{+}23, TWC22, UAL24, VGG^{+}23, VH20, dBDL^{+}20, vLSB^{+}21, vdWDH^{+}21]. Community-level [ACG22]. commuting [KKMW23]. comparable [CLGH24]. Comparative [HKD^{+}20, MSC^{+}22, YMU^{+}21, RBF^{+}20]. compared [BFC^{+}20, OHBJ20]. Comparing [BMW24]. Comparison [BGG^{+}22, KSS^{+}20, LMRS^{+}24, PAF^{+}22, SIBM20, CIDM23, CDD^{+}20b, HO23, ILON21, Lei20, MLP^{+}24, NTN^{+}24]. compensating [SSM20]. Competition [AS20, JSK19, LAD^{+}21, GGCL22, HWA^{+}22, KLFP23, NSMJ20, OCFEB21, TFH^{+}22, WSPS^{+}20]. competitors [OCFEB21]. Complex [MSL^{+}24, BS21, FNS^{+}20, GFH^{+}22, OvBS^{+}24, YAZA21]. complexity [BMO^{+}24, CFHCS^{+}24, CC22, HDB21, HAB^{+}22, HHY^{+}23, MSPS22, PCR23]. component [ZLHH20]. composition [AHH^{+}20, BWRP20, BPS^{+}24a, BCP^{+}22, DHA^{+}20, FBC^{+}24c, FSSL^{+}20, GSL^{+}24b, GMRMG22, GVOCSGG21, HKK^{+}23, HDB21, LAP^{+}24, LHHH^{+}20, MJW24, dACMRZ^{+}20, OJ24, PLLT^{+}21, SHZ^{+}20, SNP20, SSG^{+}22, UAL24, VH20, WXZ^{+}24, ZHG^{+}21, dSGC^{+}23]. compositional [CSH^{+}21b]. compositions [HLCH23, ICH^{+}24]. Compound [KdlVJ^{+}21, DSA^{+}24, FWWH22, HMT21, XCT^{+}20]. Compound-specific [KdlVJ^{+}21, DSA^{+}24, FWWH22, HMT21, XCT^{+}20]. Compounded [BPSB^{+}22]. Comprehensive [ECC^{+}21]. compromises [LAP^{+}24]. concentration [BOY^{+}23]. concentrations [BBR24, LSSG24]. conch [VSGD24]. concrete [HTL^{+}20]. Condition [GBG^{+}23, ASH^{+}20, BKT^{+}22, BBRM22, CRB^{+}21, CKT^{+}24, CDD^{+}20b, CRGMA^{+}21, CKCS20, DKS^{+}24, DDD^{+}23, DFP^{+}21b, FLCQ21, FOM^{+}20, GSK^{+}20, GGT^{+}20, LCP^{+}20, LGD23, MVC^{+}23, RHH^{+}20, RBS^{+}24, RCC^{+}22, SBRQ21, WTK^{+}24, WBG^{+}22]. condition-related [FLCQ21]. conditioned [NP22]. conditions [ABC^{+}24, BRC22, BSF24, CMvKdR23, EHLM^{+}23, FTA^{+}22, GBH^{+}20, GRTK^{+}21, GLW21, GB21, GMB^{+}24, HPRP21, HWT^{+}22, ICIS23, KKMW23, LBH^{+}21c, MK20, NBL^{+}22, OLM24, RHM^{+}22, RJ21, ROMB22, SRG^{+}23, SJSR24, SBJ^{+}24, SMB^{+}20, TYL^{+}24, THR^{+}23, UPK^{+}20, WRW20, WHB^{+}20, ZZYB21]. conduct [KDS^{+}20]. Confirmed [DGM^{+}23b]. confirms [LH22]. conflicts [YSD^{+}21]. congener [OHBJ20]. congeneric [JBG22, MGBCGM^{+}22, RC23]. connected [Ano24o, PLH^{+}21]. connections [vLSB^{+}21]. Connectivity [GIDP^{+}21, SKTO23, AFTW24, AFP^{+}24, ABH^{+}21, BMW24, BNSH21, CAP^{+}21, DTHC24, Det24, EB20, FBM^{+}24, FHAP20, GCDP21, GFDN^{+}24, HESF24, HNN^{+}24, JMP^{+}24, LMH^{+}24, MDPL20, MGHH21, MAPFH22, MDY^{+}21, MON^{+}24, MA20, MLP^{+}24, MDQ^{+}21, NDA^{+}22, OvBS^{+}24, QRC^{+}24, RWR24, SVT^{+}24, STM^{+}24, SF23, SWP^{+}24, XW24, vdVWF^{+}24]. connects [KNH24]. consequence [Ano24o]. Consequences [ASH^{+}20, BLA^{+}24, DWJL^{+}21, JETGM24, MTB^{+}23]. consequent [LRM^{+}21]. Conservation [SHG^{+}21, YSP^{+}22, APG^{+}22, DFP^{+}21a, DVR^{+}23, FBC^{+}24a, HGK^{+}24, JLT^{+}21, MSDY23, NRE21, SON^{+}20, SDE^{+}21, vdHMA^{+}24]. considerations [EPHS23]. consistency [HHO^{+}23, RB22]. Consistent [NCS^{+}23, FWGD^{+}23, KKC^{+}24, OLM24]. Conspecific [LFA^{+}23, FSF^{+}24, SSK^{+}23]. conspecifics [BKTN21, KPTT23, SST20b, TdLHL21, VMCA^{+}20]. constant [FN24]. constitutive [VSG20]. constrain [RBM^{+}21]. constrained [LMHS21]. constrains [JPMG23, MBT^{+}20]. constraints [GSRC20, LS21]. construction [WBGC24]. consumer [CBB^{+}22b]. consumers [FDF20, FJJ^{+}21, SMPLF^{+}22, ZPVN23]. consumption [AU21, DCG^{+}20, MDH^{+}21, Rob21, TGB^{+}20]. consumptive [GLG21]. contacts [KPS^{+}22]. contamination [ABD^{+}21]. contemporary [JSK19, MGB^{+}22]. content [CRB^{+}21, DSK^{+}22, HRH22, PWvdM^{+}23, RRS^{+}23, VPB^{+}20, VMDA22, ZZLS24]. contents [ZMC^{+}22]. Context [HMB^{+}22, MNG^{+}20, RVV^{+}24, GON^{+}20]. Context-dependency [MNG^{+}20]. Context-dependent [HMB^{+}22, RVV^{+}24]. continental [BGG^{+}22, DNJCH21, GWA^{+}23, SHM^{+}20a]. contingent [AWC21]. continue [NHFS22]. continuous [MCvdM^{+}24, SBL^{+}23, SLJ^{+}20, TTA20]. contractile [HDM^{+}24]. Contrasted [BTB^{+}24]. Contrasting [MDS^{+}23, MWCR20, PAT^{+}24, ŠGH^{+}21, SCM^{+}23, CCP^{+}21, FDF20, GSC^{+}22, TPZ^{+}21]. contrasts [BMK^{+}22]. contribute [BHQ^{+}22, MBGM21, PBS^{+}24]. contributes [BVD^{+}20, MMD20]. Contribution [FPG^{+}21, GBSS24, HDB21, HvdSR^{+}21, QMDGM22]. Contributions [Ano23, Ano24o, DDR^{+}21, SFF^{+}20]. control [ABC^{+}24, CAY24, DSB^{+}21, DDB^{+}24, MDS^{+}21, RCB24b]. controlled [PMH23]. conventional [CGS^{+}23]. Cook [SAMP24]. cooperative [DVR^{+}23]. cope [DBS^{+}20]. copepod [Fig21, FCSG21, GM20, JP20, LSSG24, OSBA22, SBDM^{+}23, VKdJ24, VIV^{+}20]. copepod-parasite [OSBA22]. Copepoda [LLF^{+}20]. copepods [HLCH23, ICH^{+}24, NGRH20, SFF^{+}20]. Coping [DB21]. Copper [BH22]. Copyright [Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n]. Coral [CLGH24, CGH^{+}23, DML^{+}21, HBR^{+}21, SSK^{+}23, APG^{+}22, AKU^{+}22, BGL^{+}21, BFC^{+}20, BBD^{+}22, BQC^{+}22, BRU^{+}22, BMP^{+}22, CLD23, CGLC20, CRLC21, CPH20, CCL^{+}22, DeC24, DFB^{+}24, DDR^{+}21, DCR^{+}21, DMBJ21, ER20, EP23, EKRT^{+}20, FBC^{+}24a, GSRC20, GMCE^{+}22, GWC^{+}21, GMS^{+}21, GBC^{+}20, GMLC22, HHO^{+}23, HNN^{+}24, ICIS23, JBDH21, JPMG23, JMS^{+}22, KDLL21, KKH22, KETL22, LvAH^{+}24, LS21, LCR^{+}22, MPA^{+}22, MBS^{+}24, MSPS22, MDM^{+}20, MFC^{+}24, MBR^{+}21, NMAE^{+}22, OCFEB21, PGM^{+}23, PAAEV^{+}24, PCM^{+}20, REGP22, RBB^{+}24, RCES22, RBM22, SSFD^{+}23, SHM^{+}20b, SRG^{+}23, SCP^{+}23, SDM^{+}23, SWP^{+}24, TB21, VHG20, WNH^{+}20, XMY^{+}22, YRB^{+}20, BMP^{+}22]. coral-depleted [EP23]. coral-dominated [BFC^{+}20]. coral-dwelling [CPH20]. coralline [CCC^{+}23, GPCM22, HM23, KLN^{+}21, TCM^{+}20, VLPN24]. corallivory [DML^{+}21]. corals [BVKF23, BBKW20, CTTLCG^{+}24, ER20, EVH^{+}21, GED20, JH20, KP21, KZC23, LMHS21, MMA^{+}22, MHK22b, NSM24, NYW^{+}22, RNQ^{+}20]. cordgrass [SAH^{+}24]. cordiformis [JJV^{+}21]. core [GSC^{+}22]. coriacea [BPS^{+}21]. Cormorants [GVR^{+}20]. correlate [MMS^{+}22]. correlated [RC23]. correlates [HPDG22, WK21]. correlations [LCJ22]. Corridors [LMH^{+}24]. Corrigendum [Ano20, Ano21a, Ano21b, Ano21d, Ano21c, Ano21f, Ano21e, Ano22a, Ano22b, Ano22c, Ano22d]. corroborate [UIS^{+}24]. cortisol [KTW^{+}22]. Coryphopterus [BSH21a]. Cost [CP22b, CBB^{+}22a]. Cost-effective [CP22b]. Costa [BPS^{+}21]. costs [DWJL^{+}21, GFK^{+}23, GCCM20, RL24]. could [TTM^{+}22, TTM^{+}24]. counter [CSDB22]. counterparts [WRB21]. coupling [JDDF21, PBT^{+}21, RF20, Zue22]. courtship [MMS^{+}22]. covariance [BB24, CAEG21]. covariate [EPHS23]. covariates [KCB^{+}24, SW23b]. cover [BLWJ20, DDG^{+}21, GSL^{+}24a, IAG^{+}23, LHFD^{+}24, PAF^{+}22, SdFBM24]. covered [BP24]. Covid [LGR^{+}22]. Covid-19 [LGR^{+}22]. cowrie [VOM^{+}20]. cows [WKAH22]. crab [ACvdC^{+}21, BSF^{+}20, CRB^{+}21, CGCC24, CGM23, DBK^{+}23, GSL20, GCCM20, HRH22, HAB^{+}22, HMB^{+}22, HCS^{+}24, KVdM^{+}22, KLFP23, LO20, LSS^{+}22, RGL24, RJE^{+}22, RAA^{+}23, Rob21, RBFA20, ŠGH^{+}21, SFL23, TFCS22, TWC20, WLC22, ZBB^{+}20, ZBB^{+}23, dlBSR^{+}20]. crab-clam [GSL20]. crab-gastropod [RBFA20]. crabs [BMKB22, BBMM20, CCW21, GTH^{+}20, HWA^{+}22, HW20, LSSL21, MCH^{+}21, RGK22, SRK^{+}23, SAH^{+}24, SNH^{+}20, SF23, WRB21, ZSWR^{+}22]. crash [BWC^{+}22, SSC^{+}23]. crash-landing [SSC^{+}23]. crassispina [BKTN21]. Crassostrea [AR21a, GCDP21, MVSV22, OAM^{+}24, PMH23, PRD^{+}20]. crawler [THD^{+}22]. crawling [MLVJ23]. credible [FWA^{+}23]. creeks [RB20]. Crepidula [CP22a, MK20]. crispus [FLCQ21]. criteria [BMS^{+}24]. Critical [BBSM^{+}20, DSA^{+}24, FYM^{+}24, CDS23, DCG^{+}20, FBC^{+}24a, KKH22, Lei20, SW23a]. Critically [FDS^{+}21, HCP^{+}24, LDW^{+}21, PBAC23, SVT^{+}24, LvAH^{+}24]. crocodile [LCP^{+}20]. Crocodylus [LCP^{+}20]. Croix [AEHD21]. cross [CKS^{+}24, LPB^{+}23, WWH^{+}22]. cross-basin [WWH^{+}22]. cross-generational [LPB^{+}23]. cross-shelf [CKS^{+}24]. crown [CSDB22, DML^{+}21, DDB^{+}24]. crown-of-thorns [CSDB22, DML^{+}21, DDB^{+}24]. Crozet [FPG^{+}21]. crucial [Ano24o]. crustacean [KZC23, dlBSL^{+}22]. crustaceans [AC20, FDL^{+}21, TP20]. crustose [VLPN24]. cryptic [OSBA22]. cryptobenthic [BSH21a]. Cryptochiton [MB20]. ctenophores [SCM^{+}23]. Cuba [NMAE^{+}22]. Cubozoa [MJK24]. cucumber [HP23]. cue [SST20b]. cues [CFSH22, dOFBSdJR^{+}22, MHH^{+}22, SBMW22, TP20, VMCA^{+}20]. culture [NP22]. cunner [TPGG^{+}23]. cup [WMT20]. Current [AMMADDH20, CAAG^{+}22, GSC^{+}22, LO20, LNM22, NSPH20, SBDM^{+}23, Ano23, Ano24o, ENF^{+}22, AMUPJR^{+}24, ASS^{+}23, FBC^{+}24b, KDB^{+}24, PRC^{+}21, RJE^{+}22, SHB^{+}20, SSB^{+}20]. currents [JP20, KPS^{+}22, TMPPG24, WDP^{+}21]. cusped [IMYW24]. Cuvier [BSFM20, FPB^{+}21, SSF^{+}22]. Cyanobacterial [DMBB22]. cycle [CEAO^{+}22, GD23, HTT^{+}21, QHT^{+}21]. cyclicity [SWC^{+}23b]. cycling [DDR^{+}21, RAHM^{+}21, RSL^{+}22, WLB23, ZBB^{+}23]. Cymodocea [AS20]. cyprid [JPJ^{+}22]. Cyprinodon [EHHF^{+}20].

Daily [MSC^{+}20, CWD^{+}22, JNJ^{+}20]. damicornis [LHW^{+}20]. damselfishes [CPH20, VK21]. Danish [SNE^{+}24]. d’Arguin [EHLM^{+}23]. darkening [MBT^{+}20]. data [AFP^{+}24, BTDR20, BML^{+}22, BZT^{+}23, BBKW20, CBB^{+}22a, DCR^{+}21, FFMR24, GSL20, HBG^{+}22, KKS^{+}20, MDW^{+}22, MGB^{+}22, MLP^{+}24, NGPGO^{+}23, NBW^{+}23, PCP^{+}20, PWvdM^{+}21, RSS20, RTG^{+}20, SBL^{+}23, SHT^{+}22, TCY^{+}22, VH20, VDC^{+}24]. data-limited [BTDR20, KKS^{+}20]. Davis [DFHT23]. day [BSFM20]. Daya [HLCH23]. Daylength [MBB^{+}22]. dead [BKTN21, SFF^{+}20]. Dealing [OCR^{+}24]. debate [SBS20]. debris [BGGH^{+}22]. Decadal [EP23, BLWJ20, MCJ^{+}20, STO^{+}23]. Decadal-scale [EP23, MCJ^{+}20]. decade [CDD^{+}20a, OBD^{+}21, RJE^{+}22]. decades [BTB^{+}24, GWA^{+}23, PGGE21, PHE^{+}20, vdHMA^{+}24]. decapod [FDL^{+}21]. decision [dOFBSdJR^{+}22]. decision-making [dOFBSdJR^{+}22]. decisions [RWSR24]. decline [BTB^{+}24, BLWJ20, CLGH24, DHE^{+}23, FBP^{+}22, JMR21, SWHE20]. declines [MCJ^{+}20]. declining [OvBS^{+}24, SD22]. decomposition [HSB20, PFDF^{+}21, WF21]. decouples [WRW20]. decrease [XMY^{+}22]. decreases [RS21]. Decreasing [TP20]. deep [BKM^{+}21, BNH^{+}21, CEAO^{+}22, DCR^{+}21, ENRC^{+}20, FBM^{+}24, GBW^{+}20, HSSN^{+}21, HBG^{+}22, KP21, LPJ^{+}20, LR22, MMB22, MLAS22, PAT^{+}24, RBM^{+}21, SSP^{+}22, SRG^{+}23, SFMK^{+}23, TKM^{+}22, VNM^{+}21, VNG^{+}22, WIY^{+}23, ZFD^{+}24, ZCB20]. deep-diving [SFMK^{+}23]. deep-sea [BNH^{+}21, ENRC^{+}20, FBM^{+}24, GBW^{+}20, HBG^{+}22, KP21, LPJ^{+}20, MMB22, PAT^{+}24, RBM^{+}21, SRG^{+}23, TKM^{+}22, WIY^{+}23, ZCB20]. deep-water [DCR^{+}21, MLAS22, VNM^{+}21, VNG^{+}22, ZFD^{+}24]. deepening [ASGI24]. deepsea [BBKW20]. deepwater [JJV^{+}21, MRH20, MTB^{+}23]. Defaunation [COWM^{+}22b]. defences [DML^{+}21]. defenses [AGA^{+}22, RWSR24, SFOPL21, VSG20]. deficit [dMGGG21]. defined [SRG^{+}23]. Defining [VDC^{+}24]. defoliation [KHOK22]. degradation [Ano24o, BAHG23, GSBO21, HGT^{+}21]. Delaware [SNH^{+}20]. deleterious [CSDB22]. delivery [HOK^{+}22, WSD^{+}20]. Delphinus [PBB^{+}20]. delphis [PBB^{+}20]. Delta [CCBT^{+}20]. Demersal [BBL^{+}21b, AGB24, BWGH23, BNH^{+}21, HOP^{+}20, KNO^{+}21, LFK^{+}21, LSKS24, TAN^{+}22, VOP21]. Demographic [VLG^{+}21, BMO^{+}24, Det24, HESF24, KI24, SHG^{+}21]. Demographics [HGK^{+}22, CASF21, SBFP^{+}24]. Demography [WBS24, ER20, JBWS24, LBH^{+}21a, MJ21, SAMP24, SBW^{+}22]. demonstrates [ACA^{+}24, MDS^{+}21]. demonstration [CGM^{+}24]. Demospongiae [SMS^{+}24]. Dendraster [NHH^{+}21]. denitrification [EKRT^{+}20, MMLPP^{+}24, STH21]. Denmark [PLH^{+}21, SLO^{+}22]. denning [LHM^{+}23]. densities [FSF^{+}24, KRMR24, NP22]. density [EPHS23, FYK^{+}20, GGT^{+}20, HCS^{+}24, KRMR24, KLFP23, NRE21, NTN^{+}24, ORB^{+}21, RS21, RYF^{+}24, SBL^{+}23, SBFP^{+}24, VBB^{+}21, WDJ20]. density-dependent [FYK^{+}20]. dentatus [BSF^{+}20]. denticle [DBP^{+}22]. dentine [ZMD^{+}21]. deoxygenation [COWM^{+}22b]. dependencies [MLB^{+}21]. dependency [FMD20, MNG^{+}20]. dependent [ABC^{+}24, BAH^{+}22, DMM^{+}24b, FYK^{+}20, HMB^{+}22, IMM^{+}20, MPA^{+}22, MJ21, RVV^{+}24, SBMW22, SJ22, SSK^{+}23, TTJS20]. depends [PFM^{+}21, THR^{+}23]. depleted [BAS23, EP23]. depletion [LDM^{+}20, WLB^{+}20]. deployments [AJH22]. deposit [HW20]. deposit-feeding [HW20]. deposition [GSBO21, SRK^{+}23]. deposits [BAHG23]. Depot [Ano21b]. depredated [FPG^{+}21]. depredating [GG23]. depredation [CLM^{+}21]. Depth [HYG^{+}20, LR22, MPA^{+}22, BSFM20, BGDJ20, Fai21, GIDP^{+}21, GMCE^{+}22, GDG^{+}22, MTCT^{+}23, PDF^{+}23, SSF^{+}22, Tha21, WKR21, XWW^{+}21]. Depth-dependent [MPA^{+}22]. derived [ASS^{+}21a, BVD^{+}20, BB24, BSH^{+}21b, CSH^{+}21b, KRG^{+}20, OJPC23, PCP^{+}20, RTG^{+}20, SWP^{+}24]. Deriving [JMP^{+}24]. Dermal [DBP^{+}22]. Dermochelys [BPS^{+}21]. describe [EMDR22]. description [OVO^{+}21]. descriptions [BGDJ20]. design [HCP^{+}24, OWH23, OAdJA^{+}20, WCP^{+}21a]. designing [VNM^{+}21]. Desperate [RBB^{+}21]. despite [CLGH24, HPRP21, TCT^{+}23, dSGC^{+}23]. detect [BMB24]. detected [FWWH22, JS20]. detecting [BGG^{+}22]. detection [WCL20]. detections [FWA^{+}23]. detects [HO21, THK^{+}22]. Determinants [FHAP20, CSD^{+}24, MBR^{+}21]. determination [APH^{+}23]. determine [BNSH21, SLH^{+}22]. determined [DGP22, DMM^{+}24a, SHS21]. determines [DHM^{+}24, RH21]. Determining [BSH21a, BJH^{+}23, SW23b, CSN^{+}23, FBC^{+}24a]. deterrent [BMS^{+}21]. detoxification [dFRvdG^{+}22]. Detritivores [RDT^{+}23]. detritus [BVD^{+}20, HSB20, PFDF^{+}21]. develop [BBS^{+}21]. developing [BNS^{+}24, EHHF^{+}20]. development [CMTP22, CP22a, ENF^{+}22, GSC^{+}22, JFR^{+}21, SBW^{+}22, ZDBS20]. developments [Ano23]. device [BMS^{+}21]. devices [DDD^{+}23]. Dewar [KCT24]. Dewar-Fowler [KCT24]. Diadema [LvAH^{+}24]. diatom [AS24, DL20, LDM^{+}20, LBB^{+}21]. Diatoms [Bai22, RSvV^{+}23, Vir23, WSU21]. dictates [LPD^{+}24]. Dictyota [RTT^{+}21]. dieffenbachii [FJB^{+}23]. Diel [HNB^{+}24, HNN^{+}24, LRM^{+}21, CZS^{+}22, GD23, HBS^{+}21, IMM^{+}20, PJA^{+}21]. Diet [CAW^{+}22, CCBD^{+}21, GSL^{+}24b, GMRMG22, HOP^{+}20, MAS^{+}20, PLC^{+}20, WTK^{+}24, BPS^{+}24b, CVG^{+}23, CKT^{+}24, CFFC^{+}23, CGR^{+}20, CDP^{+}21, DSK^{+}20, FOO20, FSAG22, FBC^{+}24c, HSA^{+}21, HMT21, KTW^{+}22, LNM22, MSC^{+}22, MCF^{+}21, OMW^{+}20, PVL^{+}21, PRCvdK23, PCP^{+}20, PDF^{+}23, RAA^{+}23, RMASA20, SD22, TMN^{+}24, WRW20, WYO^{+}20, dSGC^{+}23, vdRJL20]. Dietary [CEG^{+}21, GGCL22, JJV^{+}21, RHSM21, CSN^{+}23, JBGA23, MPR^{+}23, STO^{+}23, TFCS22, WSB^{+}24, WBG^{+}22, ZFD^{+}24]. dietary-induced [STO^{+}23]. diets [BLW^{+}24, CDC^{+}23, CSN^{+}23, CMF21, DKS^{+}24, GBA24, KGG^{+}21, LKR^{+}21]. differ [CKP20, GLG21, KHG^{+}21, RH23]. Differences [AOO^{+}21, CCP^{+}21, EVAB^{+}23, PDF^{+}23, ZZX^{+}20, ARH^{+}20, BQC^{+}22, CTTLCG^{+}24, CCA^{+}21, DCG^{+}20, GSBO21, HO21, JBGA23, JNJ^{+}20, KKN^{+}20, KSW23, LLJ^{+}20, MJW24, MDG^{+}21, dACMRZ^{+}20, PGW23, RHSM21, SC23, TPZ^{+}21, XMY^{+}22, YRB^{+}20, YGMNA22, dSGC^{+}23]. Different [BGL^{+}21, LSC^{+}21, BBEF^{+}20, BJH^{+}23, BAF20, CCC^{+}23, CFHCS^{+}24, EHLM^{+}23, GBSS24, GMLC22, HWT^{+}22, LNZL21, PMH23, PGT^{+}21, RS20a, SCB^{+}20, XWH^{+}22, YHW^{+}23, ZBB^{+}23, vdWDH^{+}21]. Differential [BS20b, CKS^{+}24, DPCL^{+}23, JRVE^{+}22, LK21, CFSH22, FS21]. differentially [MPB^{+}24]. differentiation [BPA21, HOK^{+}20]. differently [ATLT20, MR20]. Differing [MDK^{+}21]. differs [BLBBL20, MGBCGM^{+}22, RBVA23]. diffusion [LNKJ^{+}22]. diffusive [Hew21]. digestion [HHY^{+}23, TGB^{+}20]. digestion-limited [HHY^{+}23]. digital [SEJ^{+}24]. digitalis [BNS^{+}24]. dimensions [OCFEB21]. dimethylsulfoniopropionate [SNP20]. dimidiatus [GBC^{+}20]. dimorphic [DVMH^{+}24]. dinitrogen [EKRT^{+}20]. dinner [RS20b]. dinoflagellate [CP22a, HWT^{+}22, LSSG24, LNZL21, ZSC^{+}22]. Dinophysis [LSSG24]. Diplodus [PSB^{+}23]. Direct [WHRC22, Ano23, BNS^{+}24, LEG^{+}20, OA24]. direct-developing [BNS^{+}24]. directed [CBVA^{+}21]. directional [EGK^{+}21]. directions [OH22]. disappearing [ER20]. disaster [BSA^{+}24]. discharge [MKS^{+}20]. Disclaimer [Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano24l, Ano24m, Ano24n]. Disclosing [SJA^{+}21]. Discontinuous [BSM23]. Discovery [OSBA22, PEBG^{+}20]. Discriminating [PSÓ^{+}24]. Discrimination [AAFC^{+}21, DSA^{+}24]. discus [HNKK21, MY21]. disease [AGA^{+}22, GEY^{+}21, Hew21, MGM^{+}24, SKHC23]. Disentangling [FYK^{+}20, TAN^{+}22]. disparate [CLGH24]. dispersal [ABH^{+}21, BNSH21, CCL^{+}22, EPB23, FOM^{+}20, KHOK22, MAPFH22, MDS^{+}21, MDQ^{+}21, NDA^{+}22, OAdJA^{+}20, PFM^{+}21, QRC^{+}24, RJE^{+}22, SCM21, TMPPG24, YSP^{+}22, ZXL^{+}24, ZZYB21]. dispersing [BPA21, EPB23]. dispersion [BFQ^{+}23]. Dissolved [GBW^{+}20, MSL^{+}21, RHH^{+}20, DDR^{+}21, HvdSR^{+}21]. Dissostichus [LSR^{+}24]. distance [DWJL^{+}21, EPHS23, EPB23, GMS^{+}21, JMP^{+}24, SLBK^{+}20, SGB^{+}21]. distances [MBB^{+}23]. Distinct [RGK22, UIS^{+}24, YHW^{+}23, DCG^{+}20, MdMR24, SSK^{+}23, VPB^{+}20]. Distinguishing [CKKL21, GPP^{+}24]. distributed [YMU^{+}21]. Distribution [AMMADDH20, BZT^{+}23, CCW21, FNS^{+}20, SdGWN21, ABD^{+}21, AFP^{+}24, Ano24o, BMF24, BVKF23, BBEF^{+}20, BAB^{+}24, BKM^{+}21, BRD20, BRSD20, CCS^{+}22, CBWH24, CSD^{+}24, DdGA^{+}21, DCR^{+}21, DNJCH21, EVH^{+}21, EWBH24, FTA^{+}21, FSS23, Fig21, FDS^{+}21, FNL^{+}23, FNH^{+}23, GKK^{+}22, GBA24, HPDG22, JHDE20, JFB^{+}23, KSW23, LBW21, LÁGLL20, MHH^{+}20, MSR^{+}21b, MFC^{+}24, MFCF21, MTSO^{+}24, NSPH20, OAM^{+}24, ORB^{+}21, OJPC23, PBAC23, PS23, QRC^{+}24, RMR^{+}23, RDL^{+}21, RBM^{+}21, RBC^{+}24, SRHM21, SCM21, SSB^{+}20, SAF^{+}24, SBJ^{+}24, SDF^{+}21, Tha21, VDG24, ZBB^{+}20, dlBSR^{+}20]. distributional [SEJ^{+}24, SCM^{+}23]. Distributions [JLT^{+}21, ADF20, AHCT22a, AHCT22b, ASBA^{+}20, FTC^{+}20, GRB22, HMV^{+}22, LR22, PCLQ^{+}20, PTCD22, SRHM21, TKS^{+}20, YAZA21]. Disturbance [WCRTT20, BBS^{+}21, BGGH^{+}22, BPSB^{+}22, CLD23, CDC21, DFB^{+}24, GWC^{+}21, JETGM24, Mon23, MBR^{+}21, NFDF^{+}20, SCP^{+}23, SGP24]. disturbance-recovery [BGGH^{+}22]. dive [BSFM20, BMS^{+}24, FNO^{+}23, MLE^{+}22, PDF^{+}23, SD22]. diver [SAMdL^{+}20]. diver-operated [SAMdL^{+}20]. Divergent [BOY^{+}23, KKC^{+}24, CLR20, CLGH24, COWM^{+}22b, FWWH22]. Diverse [VGG^{+}23, FSSL^{+}20, GFK^{+}23, RCC^{+}24]. diversify [JDDF21]. diversions [GKK^{+}22]. Diversity [Vir23, AS24, AM20, AHMVM23, AKW^{+}22, AC20, BLB23, BBKW20, CATF^{+}21, Cha22, CIDM23, CNC^{+}20, CFB21, CSH^{+}21b, CSB24, CGH^{+}23, DBK^{+}23, ENRC^{+}20, EPB23, FAJAO^{+}21, GLB^{+}20, GIDP^{+}21, HB24, KP21, LGGR^{+}22, MYSF23, MHK22a, SVT^{+}24, SOS^{+}21a, STY^{+}20, SIBM20, SRR^{+}20, TQG24, TKM^{+}22, TMN^{+}24, WXZ^{+}24]. dives [RBW^{+}20, SSF^{+}22]. Diving [MTCT^{+}23, MCFB20, BDC^{+}22, BKA^{+}20, CWC^{+}22, EMDR22, FFT^{+}20, GJM^{+}20, MJJJ20, RBW^{+}20, SFMK^{+}23, WHB^{+}20]. DMSP [YHW^{+}23]. DNA [CVG^{+}23, COT^{+}21, CGS^{+}23, FOO20, FWA^{+}23, ICIS23, MSC^{+}22, MTSO^{+}24, MHK22a, RRS^{+}23, SJS^{+}23, THK^{+}22, TKM^{+}22, TMN^{+}24]. DNA-based [MSC^{+}22]. Do [PK24, CEAO^{+}22, SFOPL21]. document [FWGD^{+}23]. does [DMS^{+}21, DWJL^{+}21, RHH^{+}20, WRS^{+}23]. dollar [NHH^{+}21]. dollars [SF23, UBO^{+}24]. dolphin [CEG^{+}21, EWBH24, GRTK^{+}21, MCLC^{+}23, PBB^{+}20, SON^{+}20, TRK^{+}22]. dolphins [BRD20, BSÁG^{+}22, CNC^{+}20, CCA^{+}20, GG23, OCC23, PBAC23, SBS^{+}22, TPZ^{+}21]. domain [YAZA21]. dominance [RDL^{+}22]. dominant [HOP^{+}20]. dominate [TFH^{+}22]. dominated [ASBA^{+}20, BFC^{+}20, CWD24, FQVMF23, HPDG22]. domination [HDB21]. Donax [CGL^{+}22, RPD^{+}23]. dormancy [HDM^{+}24]. Doryteuthis [JBGA23]. Doubtful [BRD20]. DOV [SAMdL^{+}20]. down [CAY24, MSF24, SAH^{+}24]. downward [WDP^{+}21]. dredged [SRK^{+}23]. drift [KTR^{+}20, MJ23]. drifting [DDD^{+}23, NGPGO^{+}23, Str23]. drive [AS24, BICQG23, DSHM20, GIDP^{+}21, GDG^{+}22]. driven [BAH^{+}22, CLD23, GSL20, HGT^{+}21, KGM^{+}24, LLJ^{+}20, MMP^{+}22, RWM^{+}22, SWR23, SBW^{+}22, YHW^{+}23]. driver [ASPM21, CH20, OSJ^{+}22, UAL24]. Drivers [SCP^{+}23, SRR^{+}20, VDG24, BNH^{+}21, BRSD20, CCR^{+}20, DSHF22, FBC^{+}24b, FPZJAO^{+}22, FP20, LBH^{+}21a, LSC^{+}21, LD24, LFA^{+}23, MSR^{+}21a, MJJJ20, MYSF23, MSPS22, RFJC^{+}24, RBB^{+}24, RCB^{+}24a, SBRQ21, SPFL20, Vir23, WHRC22]. drives [HAB^{+}22, dlBSL^{+}22]. driving [DHF24]. drone [CBB^{+}22a, MTSO^{+}24, VDC^{+}24]. drone-based [VDC^{+}24]. Drones [CCT^{+}23, BSH^{+}21b]. drop [VKdJ24]. drum [BLBBL20]. Dual [YDG21]. dubius [BJM^{+}22]. dugong [CCT^{+}23, SYMR21]. Dungeness [HRH22, LO20, RJE^{+}22]. duration [EPB23, GBG^{+}23, GLG21, MTB21, ZMD^{+}21]. during [AMMADDH20, AMT^{+}21, AEHD21, BAS23, BMP^{+}22, BBGM23, CAAG^{+}22, CGW21, CCMRII22, CZS^{+}22, CMTB^{+}22, DCG^{+}20, DGP22, DVMH^{+}24, DRP^{+}20, DDB^{+}24, FSK^{+}21, FKTK20, GGRV23, HNKK21, HPRP21, Hew21, HWT^{+}22, KNH24, KKMW23, KI24, LT21, LGR^{+}22, LHM^{+}23, MSR^{+}21a, MY21, MG22, OLM24, OCFEB21, PRM^{+}21, PDF^{+}23, RH21, RBS^{+}24, RMD24, SSFD^{+}23, SST^{+}20a, SBRQ21, TNMN24, TSS^{+}23, WdGR^{+}20, WSU21, ZDBS20]. Dutch [BD20, JKvLP20]. dwelling [BQC^{+}22, CPH20]. dynamic [CAEG21, DPP^{+}21, DFRS^{+}24, FTA^{+}22, SWP20, SWR23]. Dynamics [CSBM20, CBJ^{+}23, APOGG20, ACvdC^{+}21, ARN^{+}22, ASBA^{+}20, BTMS24, BSM23, BAHG23, BGGH^{+}22, CLV^{+}20, CNC^{+}20, CSH^{+}21a, Det24, DSK^{+}22, DKSS20, DKSS21, DDG^{+}21, FSN^{+}20, FYK^{+}20, GRB22, GJM^{+}20, HHM^{+}22, HGK^{+}22, HBS^{+}21, KPTT23, KKN^{+}20, KETL22, LBMF^{+}24, LHW^{+}20, MSTH22, MGA^{+}20, MWCR20, MBR^{+}21, MMP^{+}22, MBB^{+}23, NSM24, NGPGO^{+}23, NSKH23, OBW^{+}20, OSBA22, PHV20, PBC^{+}22, QRC^{+}24, RTT^{+}21, SCB^{+}20, SB20a, SRG^{+}24, SSK^{+}23, SDE^{+}21, TR24, URSS24, UAL24, VGA20, WAA^{+}23, WFS23, WLB^{+}20, ZSC^{+}22, ZPVN23, vdVVC^{+}21]. Dyspanopeus [KLFP23].

eagle [DCN^{+}23]. Early [ABC^{+}24, JS20, LSR^{+}24, AFP^{+}24, AAHW20, BRC22, BSF^{+}20, BAB^{+}24, BBGM23, CGLC20, CRÁ^{+}20, CSK^{+}22, CDS23, DCG^{+}20, DeC24, FSDB20, FKTK20, MLB^{+}21, MG22, MBGM21, OÁBP^{+}21, SSK^{+}23, WMKV21, dMGGG21]. Earth [HPFB20]. Earthquake [GAFS20, EWBH24]. East [CKM^{+}20, DKAB23, ENRC^{+}20, FNS^{+}20, RCC^{+}22, TTA20, AHH^{+}20, BPS^{+}21, GSC^{+}22, MLE^{+}22, NSPH20, RVCT21, RBW^{+}20]. Eastern [BGMAM^{+}21, EMGMEV^{+}21, ABWJ^{+}20, AR21a, BLA^{+}24, CTTLCG^{+}24, DNE^{+}24, DKS^{+}24, DMTP22, GCDP21, GSS22, GGT^{+}20, GBA24, ISMM20, JFB^{+}23, KPK^{+}24, LHD20, LBH^{+}21c, MVSV22, PRD^{+}20, SSB^{+}23, SMB^{+}20, TEK21, VDB^{+}24, APOGG20, CGLC20, CRLC21, FFT^{+}20, LGGR^{+}22, MGB^{+}22, dACMRZ^{+}20, PS23, RMR^{+}23]. eating [WYO^{+}20]. Ebro [CCBT^{+}20]. Echeneis [PBKG22]. Echinarachnius [UBO^{+}24]. echinoderm [CLL^{+}22]. echinoderms [ZDBS20]. Echinoid [NP22, CM22]. echinoids [CLR20]. Echinometra [UPK^{+}20]. echo [NGPGO^{+}23]. echo-sounder [NGPGO^{+}23]. echosounder [MCvdM^{+}24]. echosounding [OJPC23]. eco [AHCT22a, AHCT22b]. eco-engineered [AHCT22a, AHCT22b]. Ecological [BLA^{+}24, DCK^{+}24, FBM^{+}24, GVS23, MCK21, MSR^{+}21a, MA20, RBB^{+}24, SLG^{+}24, SWC^{+}23a, THL^{+}20, WKAH22, AKU^{+}22, Ano24o, BMW24, BMO^{+}24, BGL^{+}21, BNH^{+}21, CCAdA24, CGH^{+}23, DSHM20, DDD^{+}23, FTC^{+}20, GPMH^{+}24, GRB22, LOF^{+}24, MDS^{+}23, PAL^{+}22, RBM22, ŠGH^{+}21, SCB^{+}20, SSM20, WJ21]. ecologically [DKAB23, NP22]. ecologies [MdMR24]. Ecology [vTLG^{+}21, Ano23, AAHW20, BBH^{+}20, BSRVS22, BDS^{+}20, BKH^{+}24, BLBBL20, BPS^{+}24b, CVP^{+}21, CVG^{+}23, CCBT^{+}20, CGS^{+}23, CLL^{+}22, DJR^{+}20, DHM^{+}24, DCR^{+}22, DPL^{+}21, DFSH23, EVVMQ^{+}23, EMGMEV^{+}21, FLBP^{+}21, GLP^{+}24, GFT^{+}21, GAPM24, GWDR20, HRH22, JPJH23, KdlVJ^{+}21, KZC23, KSW23, LKB^{+}21, LWP^{+}24, MJK24, OTD^{+}24, PBB^{+}20, PSN^{+}20, PYB^{+}21, PDH^{+}24, QHT^{+}21, RHSM21, Rus20, RBCwD^{+}20, SHB^{+}20, SBS^{+}22, SBS^{+}24, SSB^{+}20, SSB^{+}23, VNM^{+}21, WNH^{+}20, ZMC^{+}22, ZCB20]. ecophysiological [BQM^{+}22, JRVE^{+}22, KHG^{+}21]. Ecosystem [FDF20, NFDF^{+}20, VLG^{+}21, dLCN^{+}22, ASBA^{+}20, BQM^{+}24, BB24, BS20b, BSM23, BSÁG^{+}22, CEG^{+}21, CC22, CGR^{+}20, CDP^{+}21, COWM^{+}22a, DCN^{+}23, DGM^{+}23a, FS22, GBG^{+}23, HSU^{+}21, HSSN^{+}21, HSNT24, HHS21, HvLM21, JDDF21, KF20, KGM^{+}24, LORB^{+}24, MAK^{+}22, MBT^{+}20, MGA^{+}20, MA20, NSKH23, RCB24b, SBB^{+}24, SBT24, SS24, SAGG^{+}24, SGW^{+}21, SDE^{+}21, TMPPG24, TAN^{+}22, TOMDC23, LO20, LNM22]. ecosystem-based [MGA^{+}20]. Ecosystem-level [NFDF^{+}20]. ecosystems [Bai22, BLGM^{+}24, BCP^{+}22, EVH^{+}21, EB20, HDB21, JMM24, JSK19, NCZ^{+}21, OLG^{+}23, OYO^{+}20, RJ21, RBF^{+}20, RHN^{+}24, RSvV^{+}23, RAF^{+}22, RCC^{+}24, SDW20, SHM^{+}20b, TPZ^{+}21, VNG^{+}22]. ecotrophic [KDK^{+}24]. ecotypes [DFP^{+}21b, JNJ^{+}20]. ectoparasites [HNN^{+}24]. ectoparasitic [MHF^{+}21]. ectosymbiotic [ZCB20]. ectotherms [Tha21]. Ecuador [HGK^{+}22]. eddies [GSC^{+}22]. eddy [BB24, BNR20, CAEG21, DL20, HOK^{+}22]. eddy-stimulated [DL20]. Edge [TR24, CKT^{+}24, DPP^{+}21, IBC^{+}21, MKR^{+}20, OHBJ20, PCR23, VDG24]. eDNA [CSWM24, MJK24]. edule [dMGGG21]. edulis [BVD^{+}20, LK21, SSFS20]. Edward [vTLG^{+}21]. eel [FJB^{+}23, THK^{+}22]. Eelgrass [BB24, BBM^{+}24, CAEG21, CEZH21, CDC21, DNE^{+}24, GEY^{+}21, KBP22, LOS^{+}22, MGHH21, MNG^{+}20, MHK22a, MBB^{+}23, OBS^{+}22, PCR23, ROMB22, SLO^{+}22, XWW^{+}21]. eelgrass-clam [MNG^{+}20]. eels [CDS20, HMK^{+}22, SCM21]. Eeyou [DNE^{+}24]. Effect [AM20, AR21a, CSH^{+}21a, DGM^{+}23a, FBC^{+}24c, KKN^{+}20, LDM^{+}20, LBMSS23, LHW^{+}20, MVC^{+}23, MCH^{+}21, PCP^{+}20, PBC^{+}22, RWH^{+}24, ZHG^{+}21, CTH^{+}20, FBKG24, PRC^{+}21, RS20b, SBT24, dSLD24, dFRvdG^{+}22]. effective [CP22b, DVR^{+}23, PBO^{+}21]. Effectiveness [COT^{+}21, LCR^{+}22, TKM^{+}22]. Effects [ALL^{+}23, BK23, BPS^{+}21, BKTN21, BBL^{+}21a, BP24, BRU^{+}22, CMT^{+}21, CP22a, CWD^{+}22, DNE^{+}24, DKAB23, DDG^{+}21, GRQ21, GMCE^{+}22, GON^{+}20, GBC^{+}20, GPCM22, GMB^{+}24, HSB20, HWA^{+}22, HMK^{+}22, KPS^{+}22, LPB^{+}23, LNZL21, LHHH^{+}20, MPM21, MLE^{+}22, MDPL20, MSF24, MGM^{+}24, NYW^{+}22, NBL^{+}22, PRD^{+}20, RAA^{+}23, SHZ^{+}20, SBMW22, SKI^{+}20, TCP^{+}23, WKR21, WGK22, XYSF20, YLH20, vDTS^{+}22, ABB^{+}20, Ano23, ASS^{+}21b, BSF^{+}20, BVKF23, BWGH23, BAF20, BKT^{+}22, BMB24, BPSB^{+}22, CATF^{+}21, CDC21, COWM^{+}22a, CDS23, Det24, EBL^{+}21, FTA^{+}22, FLCQ21, FCSG21, FYK^{+}20, GM20, GLB^{+}20, GFDPASR20, GPMH^{+}24, GAFS20, GS20, GLG21, HSU^{+}21, HMB^{+}22, HTGG22, HHBP23, HCS^{+}24, IBRL^{+}22, JS20, JCGS22, KLN^{+}21, KNH24, LKB^{+}21, LDW^{+}21, MVSV22, MK20, NFDF^{+}20, OA24, PGM^{+}23, PCR23, PCM^{+}20, RKV^{+}21, RHM^{+}22, REGP22, RSEG21, RLME^{+}21, RH20]. effects [SBSS21, SDW20, SW23a, SRK^{+}23, SWR23, SHM^{+}20b, SAH^{+}24, SBB20, THL^{+}20, TYL^{+}24, TR24, WMKV21, WS24, WBG^{+}22]. efficacy [ASY^{+}24, COWM^{+}22b, HNB^{+}24]. efficiency [AJH22, CS22, DSHM20, HSU^{+}21, MTCT^{+}23, SJ22, SVS^{+}22]. effluent [AWB23]. effluents [GLB^{+}20]. effort [KRMR24, LGD23, MTCT^{+}23, PDFH20, ZZX^{+}20]. efforts [HESF24]. egg [BPS^{+}21, CGW21, FBC^{+}24c, SOS^{+}21a, VFG24]. egg-bearing [CGW21]. Eggs [CLR20, CM22, CMT^{+}21, CFB21, HBE22]. eider [MPF^{+}22]. Eisenia [AAI^{+}24]. either [Ano24o]. Elasmobranch [MBS^{+}24, GNC^{+}23, HPDG22, PBKG22, RHN^{+}20, SW23b]. elasmobranchs [DAB^{+}21, MRH20]. elegans [EVVMQ^{+}23]. elegant [EVVMQ^{+}23]. eleginoides [LSR^{+}24]. element [TTJS20]. elemental [BTMS24, CMdL^{+}20, XW24]. elephant [CGGdD24, CSN^{+}23, LNBdB21, NBL^{+}22]. Elevated [EHHF^{+}20, RTT^{+}21, SKHC23, BS20b, HRC21, IBRL^{+}22, MGM^{+}24, SBSS21]. Elevation [DBK^{+}23, HSBB^{+}21]. elevational [WBS24]. elevations [LBR^{+}24]. eliciting [HJG21]. Elkhorn [HGK^{+}24]. elongatus [LBH^{+}21a]. elucidates [BPS^{+}24b]. Elucidating [BCC20]. embayment [HHM^{+}22, MCLC^{+}23]. embryo [TBH21]. embryonic [BJM^{+}22, TBS20]. embryos [SOS^{+}21b, Str23]. emergence [MSR^{+}21a]. emergent [MDW^{+}22, OAdJA^{+}20]. emerging [PYB^{+}21]. emertoni [CCBT^{+}20]. emission [KVdM^{+}22]. empirical [AFP^{+}24, GSL20]. Empirically [PHV20]. employ [TCM^{+}20]. enables [BWC^{+}22, MDB^{+}21]. encapsulated [Str23]. encounter [ASH^{+}20, MHDJ21]. encounters [FSDB20]. encyst [BMKB22]. end [PRC^{+}21]. end-to-end [PRC^{+}21]. Endangered [CVG^{+}23, FDS^{+}21, HCP^{+}24, LDW^{+}21, BML^{+}22, CBWH24, CSK^{+}23, CPD^{+}20, DCR^{+}22, FCV^{+}24, MJJJ20, MCFB20, PBAC23, PVB^{+}20, SVT^{+}24, XWH^{+}22, LvAH^{+}24]. endemic [VOM^{+}20]. endemism [AHMVM23]. endoparasites [KZC23]. Endozoicomonas [KP21]. enduring [UBO^{+}24]. energetic [AAP^{+}23, GCCM20, MVH^{+}21]. energetics [Ano23, GSP^{+}20, TCP^{+}23]. Energy [BQM^{+}22, RL24, BAAR20, CBHM21, CRB^{+}21, CHJ21, FPG^{+}21, JFR^{+}21, MG22, SBS20, SBW^{+}22, WDJ20]. engage [AMT^{+}21]. engineer [JDDF21]. engineered [AHCT22a, AHCT22b]. engineering [THL^{+}20]. England [HO24, TFCS22, WBGC24]. English [vdKMC^{+}24]. Engraulis [CMT^{+}21, CBKM20]. enhance [BLB23, CMO^{+}20]. enhanced [LHD20, YDG21]. enhancement [MCK21, THL^{+}20]. enhancing [SOSE22]. enriched [BSA^{+}24, SNE^{+}24]. enrichment [CBGB24, HPH20, NYW^{+}22, RHH^{+}20, SDW20]. ensemble [SGW^{+}21]. ensemble-model [SGW^{+}21]. entering [MMT^{+}24]. Entrainment [GSC^{+}22]. entry [NMB^{+}22]. environment [CKS^{+}24, ENF^{+}22, FST22, KLA21, RLSO23, SJQ^{+}22, dSGC^{+}23]. Environmental [AS24, BNH^{+}21, BPA21, DFCM21, DSHF22, Fig21, HPDG22, JCGS22, MJJJ20, MAMF21, MBGM21, RFJC^{+}24, SJS^{+}23, STH21, SMA^{+}20, SMB^{+}20, SBRQ21, THK^{+}22, TTM^{+}22, TTM^{+}24, ABB^{+}20, Ano23, Ano24o, ABC^{+}24, ASGI24, AR21a, BVKF23, BTA^{+}20, BGMRGM^{+}20, BAF20, CATF^{+}21, CCR^{+}20, CSH^{+}21b, CBKM20, CMvKdR23, DSB^{+}21, DMCB22, EHLM^{+}23, FBC^{+}24a, FBC^{+}24b, FAJAO^{+}21, FPZJAO^{+}22, GRQ21, GLW21, GVCSO^{+}24, HPRP21, HMK^{+}22, JPB21, KCB^{+}24, LSC^{+}21, LD24, LKB^{+}21, LDW^{+}21, LÁGLL20, LCGS^{+}21, MYSF23, MIL^{+}20, MHÁC^{+}22, MTSO^{+}24, MHK22a, MHH^{+}22, RJ21, RH23, SW23b, SRG^{+}23, SCM^{+}23, SvBS^{+}20, SPFL20, TAN^{+}22, TCM^{+}23, TNS^{+}22, TRK^{+}22, THR^{+}23, VSG20, VLG^{+}21, Vir23, WHRC22, XWH^{+}22, XW24]. environments [CAEG21, DFRS^{+}24, DGM^{+}23a, OYO^{+}20, OWH23, PTCD22, SSFS20, VNM^{+}21, WGK22]. enzyme [CMT^{+}21]. epi [BMOW21]. epi- [BMOW21]. epibenthic [SIBM20]. epibiont [IBC^{+}21, KMW^{+}21]. epibiosis [CAY24]. epidemics [GEY^{+}21]. epidermis [LNBdB21]. epifaunal [BICQG23, CDC21, FSSL^{+}20, PCR23, STY^{+}20]. epigenetic [TPZ^{+}21]. epinephelid [BBH^{+}20]. Epinephelus [MPKS24]. epiphytic [QMDGM22]. equilibrium [SFO^{+}21]. equina [CSWM24]. equipped [Sig23]. era [SEJ^{+}24]. eradications [RJ21]. eras [SFL23]. Eretmochelys [BASM^{+}24, DGM^{+}23b]. erosion [IHA^{+}22]. Erratum [Ano21g, Ano21h, Ano21i, Ano22f, Ano22e]. error [WCL20, WCP^{+}21a]. escape [BMKB22, MPSH22]. escapes [JSDH20]. esculenta [BAHG23, MBB^{+}22]. essential [HBE22]. established [AZ24]. establishment [GB21, PSBH21, PCM^{+}20, SJH^{+}22, XWW^{+}21]. estimate [ABN^{+}21, MHDJ21]. Estimates [AU21, MDH^{+}21, EPHS23, FSAG22, FBC^{+}24c, FCV^{+}24, KSS^{+}20]. Estimating [BLE^{+}22, CBB^{+}22a, FN24, FTC^{+}20, GGT^{+}20, MJ21, RRR^{+}23, SBL^{+}23, BMW24, CBTL20, PAF^{+}22, WFS23]. estimation [DAB^{+}21, KPK^{+}24, SBFP^{+}24]. estuaries [CWD24, GFDN^{+}24, GTH^{+}20, LOS^{+}22, LBMF^{+}24, MBT^{+}20, Whi20]. Estuarine [BAS23, CMO^{+}20, AESS24, ATR^{+}24, ABC^{+}24, AWC21, Bai22, CCJ^{+}23, CSH^{+}21b, COWM^{+}22b, EMB^{+}22, GCDA20, HTGG22, KDGG24, LDW^{+}21, LFFW21, MDS^{+}23, MSH20, MAMF21, MMS^{+}20, NHFS22, PBAC23, PTCD22, RLSO23, ROMB22, SMPLF^{+}22, SHH^{+}22, SFOPL21, TBS20, Whi21, ZSWR^{+}22]. estuarine-dependent [ABC^{+}24]. Estuarine-terrestrial [CMO^{+}20]. Estuary [BTMS24, BAS23, CCW21, AESS24, BMO^{+}24, BLBBL20, CMF21, COWM^{+}22a, HPDG22, HESF24, LAL21, OJ24, PBRF23, TRK^{+}22, VGA20, CMO^{+}20, XYSF20]. Eubalaena [FNL^{+}23, MBW^{+}20]. Eucalanus [JP20]. eukaryote [CS22]. eukaryotic [RKV^{+}21, XYSF20]. Euphausia [LRM^{+}21, MSR^{+}21b, SBSS21, WRW20, ZZLS24]. euphausiid [LRM^{+}21]. Europe [MGHH21]. European [FBC^{+}24c, GFDPASR20, KGG^{+}21, MDB^{+}21, RWR24, SBL^{+}23, SKHC23, TMPPG24, vdVWF^{+}24]. eutrophic [LAL21]. eutrophication [BG20, EKRT^{+}20, VS23]. evacuation [FBC^{+}24c]. evaluate [HNB^{+}24, JBGA23, LCR^{+}22, MDW^{+}22]. evaluated [MHK22a, RYF^{+}24]. Evaluating [BMB24, RBC^{+}24, LCJ22]. Evaluation [Cha22, DSK^{+}22, RGL24, DSA^{+}24]. Evechinus [SST20b]. even [BFC^{+}20]. event [AMMADDH20, CRGMA^{+}21, CMTB^{+}22, Hew21, KNH24, MSH20, THK^{+}22]. events [AMUPJR^{+}24, AVFD^{+}21, CKS^{+}24, JPL^{+}24, KRMR24, RVCT21, RFJC^{+}24, OLPGK22]. Evidence [BBGM23, CS20, GSS22, KCGR20, MRH20, OVO^{+}21, RS20b, Ano24o, BMKB22, DNJCH21, DDD^{+}23, FBP^{+}22, FSK^{+}21, HTL^{+}20, JJV^{+}21, KGG^{+}21, MKR^{+}20, PWvdM^{+}21, SOS^{+}21b, SJH^{+}22, TCY^{+}22, VDC^{+}24, GSL20]. Eviota [DMBJ21]. evolutionary [BNH^{+}21, CS20]. exacerbated [MMP^{+}22]. exacerbates [BJM^{+}22, WS24]. examined [HNB^{+}24]. Examining [LNKJ^{+}22, RBM22, CSB24]. example [BMB24]. excavated [MPKS24]. excavation [WIY^{+}23]. excentricus [NHH^{+}21]. exceptionally [OLM24]. excess [VS23]. exchange [dlBSL^{+}22]. excluding [FCV^{+}24]. exhibit [AZ24, CMdL^{+}20, COWM^{+}22a, GFH^{+}22, MdMR24, SJQ^{+}22]. existing [JNJ^{+}20]. Exocoetidae [PS23]. exoculata [ZCB20]. exogenous [RDT^{+}23]. exopolymer [WSU21]. expanding [CHD^{+}22, SBB20, WBS24, WRB21]. expands [LBR^{+}24]. expansion [AESS24, CS20, DA24, LPD^{+}24, MBB^{+}23, OLM24, TPGG^{+}23, TTM^{+}22, TTM^{+}24, vdKMC^{+}24]. expansions [ASL^{+}22]. expansive [EMB^{+}22]. expected [ARH^{+}20, DSHM20]. experienced [HKY^{+}20]. experiment [BSF24, dBDL^{+}20]. Experimental [LH22, EB20, HTGG22, KHL^{+}22, MGH21, SSP^{+}22, VGG^{+}23]. experimentally [PCP^{+}20]. Experiments [HGK^{+}24, BH22, DKAB23]. explain [ABH^{+}21, BVKF23, FN24, STY^{+}20]. explaining [CG22]. explains [FWA^{+}23, SSFD^{+}23]. explicit [DIC^{+}22, SLJ^{+}20]. exploitation [SFL23]. exploited [CLR^{+}21, PBO^{+}21, SWJ^{+}21]. exploration [JMS^{+}22]. explore [OH22]. Exploring [CSWM24, LGGR^{+}22, SSFS20, WLB^{+}20, ZCBC24, BSRVS22, TKM^{+}22, vdRJL20]. exponential [RS21]. export [CHDG^{+}23].

Exposing [WMKV21]. Exposure [BMM21, AMD22, CCJ^{+}23, CSK^{+}22, FCV^{+}24, GED20, GMCE^{+}22, GLG21, JBG22, MAK^{+}22, MMP^{+}22, RH21, TBS20, ZHG^{+}21]. exposures [BJH^{+}23]. express [NP22]. expression [CMT^{+}21, TEK21, TWC20]. extant [PEBG^{+}20]. extends [MSC^{+}22]. extent [SRG^{+}24, WCRTT20]. external [HWT^{+}22]. extirpation [RCB24b]. extracellular [LSSG24]. Extreme [CSN^{+}23, KRMR24, AMUPJR^{+}24, AVFD^{+}21, CMTB^{+}22, GLW21, MSH20, OOW^{+}20, RLME^{+}21, SW23a, WRG^{+}24]. Extremely [APG^{+}22, RHN^{+}20]. extrinsic [BAB^{+}24, CMdL^{+}20]. eye [CAPS20, GFDN^{+}24]. eyed [MCFB20, MCC^{+}21, YSP^{+}22]. eyes [FS21]. Ezo [MY21].

face [BAH^{+}22, NSPH20, YSD^{+}21]. facilitate [AVFD^{+}21, BS20b, HPFB20, MSDY23, NCZ^{+}21, TTM^{+}22, TTM^{+}24]. facilitates [HNN^{+}24, LvAH^{+}24]. facilitation [ARdlB^{+}22, RS20a, SSK^{+}23]. facing [SD22, WLB23, WDP^{+}21]. Factors [LBMF^{+}24, Ano23, Ano24o, ASGI24, BAB^{+}24, CGL^{+}22, DMCB22, DSA^{+}24, DFCM21, FBC^{+}24a, FYK^{+}20, GWA^{+}23, GRQ21, HBR^{+}21, HMK^{+}22, MPB^{+}24, MIL^{+}20, MAMF21, RBC^{+}24, WSD^{+}20]. Facultative [GM20, DDR^{+}21]. fail [CBTL20]. failed [PGT^{+}21]. Failure [CMTB^{+}22]. failures [ODB^{+}24, Str23]. fall [AMMADDH20, OYO^{+}20, SVB^{+}24, SSP^{+}22]. fall-specific [SSP^{+}22]. falls [HSSN^{+}21]. Family [ASS^{+}21a, CCAdA24, CPH20, PMH23]. Farallon [NBL^{+}22]. Faria [Ano21e]. farm [WBG^{+}22, WBGC24, vESB^{+}23]. farmed [AM20, BAHG23]. farming [MTRM21]. farms [JTBS^{+}22, TBH21]. fascicularis [WNH^{+}20]. fast [CBHM21]. fasting [AAP^{+}23, OTD^{+}24]. fates [QFT^{+}24]. Fatty [MBW^{+}20, Bai22, DHA^{+}20, HBE22, HIT^{+}22, JJV^{+}21, KGG^{+}21, LAP^{+}24, MPR^{+}23, SHZ^{+}20, SSG^{+}22, ZZLS24]. fauna [BSA^{+}24, SJS^{+}23, SLO^{+}22]. faunal [CP22b, MFC^{+}24, OYO^{+}20]. faunal-habitat [CP22b]. feasibility [DGM^{+}23b]. features [CCRP20, CHDG^{+}23, CBKM20, DPP^{+}21, GLM^{+}24, SCM^{+}23]. Fecal [FWA^{+}23]. fecundity [SDE^{+}22]. feedforward [NSKH23]. Feeding [BSB^{+}21, COGGH20, EVVMQ^{+}23, GGN21, SLB^{+}21, VMDA22, AU21, AKU^{+}22, CVP^{+}21, CCT^{+}23, DJR^{+}20, EMGMEV^{+}21, FLCQ21, GM20, HRH22, HW20, IMYW24, JP20, LMRS^{+}24, LSKS24, LHW^{+}20, OTD^{+}24, PSÓ^{+}24, PDF^{+}23, PJA^{+}21, RFJC^{+}24, RWSR24, SBSS21, SHZ^{+}20, SSB^{+}20, Sig23, SSFS20, SD22, VOM^{+}20, YGMNA22, ZMC^{+}22, vdRJL20]. female [BSF^{+}20, BKA^{+}20, PGW23, SBRQ21]. females [RB23, SDE^{+}22]. feminisation [SBC^{+}21]. Fenton [MY21]. ferruginea [DV20]. ferry [MCL24]. ferry-based [MCL24]. fertilisation [LDPH24]. few [PWvdV20]. fiddler [WRB21]. fidelity [ACA^{+}24, DMM^{+}24a, LBTE^{+}21, RWR24]. field [ASS^{+}21a, CRÁ^{+}20, PCP^{+}20, RSS20]. fields [dDW^{+}22, vdWDH^{+}21]. Fifteen [MW20]. Fiji [PLC^{+}20]. fill [GAPM24]. filtering [BVKF23, BBD^{+}22]. filtration [BKT^{+}22]. fimbria [CBHM21]. fin [GVAZ^{+}22, MDK^{+}21, RBVA23]. Fine [ASBA^{+}20, FCV^{+}24, HBS^{+}21, LO20, MDH^{+}20, MTSO^{+}24, PBS22, SMK23, TKS^{+}20, WDJ20, KCT24, KDB^{+}24, MMS^{+}20, OH22, OWH23, PBAC23, PVB^{+}20, SWR23, SHS21, SSBA^{+}20, VDG24]. Fine-resolution [SMK23]. Fine-scale [ASBA^{+}20, FCV^{+}24, HBS^{+}21, LO20, MDH^{+}20, MTSO^{+}24, PBS22, TKS^{+}20, WDJ20, KCT24, MMS^{+}20, OH22, OWH23, PBAC23, PVB^{+}20, SWR23, SHS21, SSBA^{+}20, VDG24]. finer [SHM^{+}20a]. finer-scale [SHM^{+}20a]. fingerprints [BTMS24]. finmarchicus [CLC^{+}23, HSU^{+}21, MHN24, RRR^{+}23]. finned [BBPO21, BMS^{+}24, SBS^{+}24, SJQ^{+}22]. FIRe [GSN^{+}20]. First [CBHM21, GB20, GVS23, NSB^{+}23, VDC^{+}24, WIY^{+}23]. first-time [NSB^{+}23]. first-year [GVS23]. Firth [BHQ^{+}22]. Fish [FFMR24, OLPGK22, OLG^{+}23, ADF20, AMUPJR^{+}24, ABC^{+}24, AWC21, ASH^{+}20, ASBA^{+}20, BS20a, BCC20, BLVL24, BQM^{+}24, BPS^{+}24a, BCP^{+}22, BBL^{+}21b, CRG^{+}20, CFHCS^{+}24, CRÁ^{+}20, CMdL^{+}20, CLR^{+}21, CFB21, CSK^{+}22, CCL^{+}22, CMvKdR23, DFP^{+}21a, DSHM20, DVP^{+}23, DSHF22, DHF24, DDD^{+}23, DDG^{+}21, ENRC^{+}20, ECC^{+}21, FVC^{+}20, FAJAO^{+}21, FSAG22, FNH^{+}23, FD21, GBW^{+}20, GSC^{+}22, GPMH^{+}24, GS20, GRQ21, GSPH20, GIDP^{+}21, GSS22, GBC^{+}20, GBA24, HOP^{+}20, HPDG22, HNN^{+}24, HO23, HBS^{+}21, HRC21, HO24, HEP^{+}24, JDM^{+}20, KNO^{+}21, KKH22, KWD21, LFK^{+}21, LDW^{+}21, LCR^{+}22, LCFJ20, LNKJ^{+}22, MCvdM^{+}24, MCF^{+}21, MDPL20, MYSF23, MFC^{+}24, MMB22, MTSO^{+}24, MHK22a, dACMRZ^{+}20, MLP^{+}24, OBS^{+}22, OCR^{+}24, OCGAP^{+}22, PTLK21, PCLQ^{+}20, PRCvdK23, PSB^{+}23, PWvdV20, PWvdM^{+}21, PWvdM^{+}23, PAAEV^{+}24, RSvL21, RSS20, RLSO23, RCC^{+}24, SBMW22, SMPLF^{+}22]. fish [SRG^{+}24, SWJ^{+}21, SJ22, SME^{+}23, SFLQ23, SWP^{+}24, SLJ^{+}20, SSBA^{+}20, TTA20, TAN^{+}22, TWC22, VGA20, VDC^{+}24, VFG24, WAA^{+}23, WXZ^{+}24, WYO^{+}20, WMKV21, WFR^{+}23, WSB^{+}24, Whi20, WYZ^{+}20, XCT^{+}20, XW24, ZXL^{+}24, ZCF20, dlBSL^{+}22, vdWDH^{+}21]. fish-derived [SWP^{+}24]. fish-eating [WYO^{+}20]. fish-habitat [WYZ^{+}20]. fisheries [AFTW24, BTDR20, BBBT^{+}20, CVP^{+}21, CSK^{+}23, CMTP22, CILGD^{+}22, CWC^{+}22, CCP^{+}21, DdGA^{+}21, DFHT23, DAT^{+}21, FDS^{+}21, FRO^{+}21, HESF24, JLT^{+}21, JSK19, KDK^{+}24, MdOO^{+}20, MDW^{+}22, MSR^{+}21b, MGA^{+}20, PRM^{+}21, WSB^{+}24, YSD^{+}21, ZCF20]. fishers [SCB^{+}20]. fishery [BBRM22, CLM^{+}21, DCR^{+}21, HRH22, HSNT24, JHK^{+}22, MDW^{+}22, MRB^{+}24, MPF^{+}22, MHDJ21, NDA^{+}22, RTG^{+}20, TCT^{+}23]. fishery-independent [RTG^{+}20]. fishes [AESS24, AMRL^{+}24, AKU^{+}22, ALL^{+}23, AGB24, BGG^{+}22, CATF^{+}21, CFFC^{+}23, CMO^{+}20, CGH^{+}23, DFSH23, EAS20, Fai21, FQVMF23, FLBP^{+}21, GAF20, GAPM24, GSRC20, GGT^{+}20, GBC^{+}20, GMLC22, HMW^{+}21, KCRA22, LPJ^{+}20, Lei20, LSKS24, MPR^{+}23, MHK22b, NMAE^{+}22, PS23, PAT^{+}24, PCM^{+}20, SDW20, SFO^{+}21, THL^{+}20, TKM^{+}22, TNS^{+}22, VOP21, WRG^{+}24, Whi21]. Fishing [CRG^{+}20, ASGI24, AOO^{+}21, CBVA^{+}21, CCT^{+}23, FPG^{+}21, FQVMF23, FD21, HSU^{+}21, HO23, JBGA23, MRB^{+}24, MBM^{+}20, SGW^{+}21, TAN^{+}22, TOMDC23, vLBGGT22]. fishing-induced [MRB^{+}24]. fishing-related [ASGI24]. fissus [WRS^{+}23]. fitness [DA24, SF23]. five [GWA^{+}23, PGM^{+}23, WFR^{+}23]. fixation [EKRT^{+}20]. Fjord [SLO^{+}22, LKB^{+}21, SSFS20, SNE^{+}24]. fjords [SJH^{+}22]. flatfish [GVS23, vdVTW^{+}22]. flats [BD20]. flatworm [CS20]. flaw [HDL^{+}21]. fleckeri [MJK24]. fledging [LPJ^{+}21, YSP^{+}22]. Flemish [SMS^{+}24]. fleshy [QMDGM22]. Flexibility [WRW20]. flexible [MDS^{+}23]. flight [Ano23, DWJL^{+}21, GFK^{+}23, KPTT23, KKMW23, SWR23]. floating [NGPGO^{+}23]. Florida [HKD^{+}20, RH20, RBM22, APH^{+}23, BBRM22, DMBB22, FFMR24, GPP^{+}24, GBA^{+}22, HHO^{+}23, HBJ^{+}21, JPMG23, PDFH20, TD21, VSGD24]. flounder [CWD24, DV20, DFCM21, HOP^{+}20, HRC21, MLB^{+}21, TFCS22, WBG^{+}22]. flounders [HTGG22]. flow [ARB^{+}20, AR21b, TR24]. fluctuating [BMM21]. fluctuation [MBGM21]. fluctuations [SWJ^{+}21, WRG^{+}24]. fluid [SWP20]. flume [CAEG21]. flume-chambers [CAEG21]. fluorescence [GSN^{+}20]. fluxes [BDJ^{+}24, HDE21, SFF^{+}20]. fly [AMT^{+}21]. fly-and-forage [AMT^{+}21]. flying [KPTT23, PS23]. focal [CWD^{+}22, KDS^{+}20]. focus [Bai22, MPF^{+}22]. fold [DWJL^{+}21]. follow [KDS^{+}20]. following [AAI^{+}24, GLW21, GWC^{+}21, JMR21, MSH20]. Food [GSRC20, SSP^{+}22, Bai22, BAS21, BWC^{+}22, CLV^{+}20, CGM23, DBS^{+}20, DSA^{+}24, GM20, GB20, GMB^{+}24, HOP^{+}20, HLCH23, HBE22, HW20, IMM^{+}20, IAG^{+}23, KPK^{+}24, LCFJ20, LRD^{+}20, MMD20, MLAS22, NT24, NHH^{+}21, NP22, PWvdV20, PWvdM^{+}23, RC23, Rob21, RBB^{+}21, RCC^{+}24, SDM^{+}22, SC23, SMS^{+}24, SD22, TGB^{+}20, TFH^{+}22, ZZLS24, vdRJL20]. food-conditioned [NP22]. food-web [PWvdM^{+}23, RCC^{+}24, SMS^{+}24]. forage [AMT^{+}21, GBA24]. Foraging [ADF20, BBR24, BDC^{+}22, DdGA^{+}21, HTGG22, MAH^{+}20, MCC^{+}21, PBB^{+}20, PBT^{+}21, ANP^{+}24, Ano23, BKTN21, BSB^{+}21, BLCC^{+}20, BRSD20, BBGM23, BPS^{+}24b, CVG^{+}23, CAAG^{+}22, CCS^{+}22, CCA^{+}21, COM^{+}24, CBTL20, CBB^{+}22b, CKP20, CCP^{+}21, DPP^{+}21, EVAB^{+}23, FTA^{+}22, FFT^{+}20, FNO^{+}23, GDAC23, GWDR20, GJM^{+}20, HKK^{+}23, HSO^{+}22, HOK^{+}22, JPJH23, JCGS22, KPW^{+}20, LJ20, LGR^{+}22, LLJ^{+}20, LGD23, LPD^{+}24, LWP^{+}24, LNBdB21, MJ23, MAS^{+}20, MTCT^{+}23, MTB21, MKR^{+}20, MDB^{+}21, MdMR24, NCS^{+}23, OOW^{+}20, PRM^{+}21, PDF^{+}23, SHB^{+}20, SJQ^{+}22, SSB^{+}23, SWC^{+}23b, SSF^{+}22, SB20b, TNMN24, TRK^{+}22, TWC^{+}23, WHN^{+}22, dDW^{+}22]. foraminifera [ASL^{+}22]. foray [DFRS^{+}24]. forbesii [CZS^{+}22]. force [DHF24]. forcing [Fig21, JPB21, MSF24, SLH^{+}22]. forecast [SGW^{+}21]. Forecasted [HGH^{+}24]. Forensic [APH^{+}23]. forereefs [CFE^{+}22]. forest [AAI^{+}24, BS20b, BSRVS22, CLL^{+}22, DHM21, ENF^{+}22, GPMH^{+}24, GSK^{+}20, MJW24, SLG^{+}24, SE20, TSS^{+}23, WK21, WCRTT20, WCP^{+}21b]. forest-forming [BS20b]. forests [BICQG23, KVdM^{+}22, LHFD^{+}24, NFDF^{+}20, YLH20, Zue22]. forgetting [DeC24]. formation [BNR20, BRSD20, SWP20]. forming [BTB^{+}24, BS20b, FDF20, VDG24]. fornicata [CP22a, MK20]. Forth [BHQ^{+}22]. fossil [PEBG^{+}20]. fouling [SMK23]. Foundation [FS22, CMTB^{+}22, EB21, ZHG^{+}21]. foundational [vdHMA^{+}24]. four [DPCL^{+}23, GGN21, KHL^{+}22, KTW^{+}22, MCH^{+}21, OFSJ20, PGGE21, WdGR^{+}20]. fourfinger [XWH^{+}22]. Fowler [KCT24]. fractionated [DSB^{+}21]. fractionation [KHL^{+}22]. fractioned [SS24]. fragment [dBDL^{+}20]. fragmentation [BICQG23, BBL^{+}21a, LCT^{+}20]. fragmented [TR24]. framework [OWH23, WFS23]. franciscana [PBAC23]. Francisco [CMO^{+}20]. Fraser [CNC^{+}20, JPJH23]. French [BPSB^{+}22, VLPN24]. frequency [GLG21, RVCT21, SBMW22]. frequency-dependent [SBMW22]. Freshening [CSDB22]. Freshwater [NMB^{+}22, GKK^{+}22]. fringe [HEKH24, KVdM^{+}22]. fringing [TSW^{+}21]. front [BDJ^{+}24, BSB^{+}21, GBSS24]. frontal [COM^{+}24]. fronts [BSB^{+}21]. Fucales [HOK^{+}20]. Fuegian [AU21, AFP^{+}24]. fuel [GEY^{+}21, PWvdV20]. full [HTT^{+}21, OLA^{+}24]. fulmars [DdGA^{+}21, DAT^{+}21]. function [AKU^{+}22, CCRP20, CIDM23, CMO^{+}20, COWM^{+}22a, HSSN^{+}21, JMR21, MBT^{+}20, OBS^{+}22, PLLT^{+}21, RBM22]. Functional [COM^{+}24, DGP22, EB21, GLB^{+}20, AS24, BBEF^{+}20, BBSM^{+}20, DTHC24, FP20, HNN^{+}24, HWA^{+}22, HHY^{+}23, LDM^{+}20, dACMRZ^{+}20, NMAE^{+}22, SSG^{+}22, SIBM20, WXZ^{+}24]. Functionally [MR20]. functioning [FDF20, SBT24, vdVTW^{+}22]. functions [DGM^{+}23a, FS22, REGP22]. Fundulus [RB20]. Fundy [SHG^{+}21]. fungal [MGH21]. further [FQVMF23, GVAZ^{+}22, JJV^{+}21, MTRM21]. fuscus [JTBS^{+}22]. future [AVFD^{+}21, GJK^{+}24, NHFS22, OH22, UPK^{+}20]. futures [HBR^{+}21].

Gadiculus [AGIS^{+}22]. gadid [KHL^{+}22, WBG^{+}22]. Gadidae [AGIS^{+}22]. Gadiformes [ASGI24]. Gadus [BLW^{+}24, BSM23, CDD^{+}20a, GBG^{+}23, JNJ^{+}20, KSMP20, LCGS^{+}21, MVC^{+}23, SBS20, SJH^{+}22]. gag [GPP^{+}24, LBMB^{+}20]. gahi [JBGA23]. gain [ICH^{+}24]. Galapagos [AMRL^{+}24, FAJAO^{+}21, CBGB24, Fig21, SAMdL^{+}20]. Galaxea [WNH^{+}20]. Galeorhinus [DIC^{+}22]. galeus [DIC^{+}22]. Galician [AGIS^{+}22]. gametophyte [LDPH24]. gammarid [KKN^{+}20]. gammaridean [KHOK22]. gammarus [SBL^{+}23]. gannet [CCA^{+}21, KCGR20]. gannets [LPJ^{+}21, RMASA20, dDW^{+}22]. gap [SAGG^{+}24]. gaps [GAPM24, KCLJ20]. gardening [VK21]. gas [ESG20, MAPFH22]. gastric [FBC^{+}24c]. gastropod [CP22a, CH20, MR20, RH21, RBFA20, SMH^{+}20, SKI^{+}20]. Gastropoda [AMMADDH20]. gazella [dLCN^{+}22]. GBR [SYMR21]. gelatinous [BBL^{+}21b]. Gene [AR21b, CMT^{+}21, TEK21]. genera [PGM^{+}23]. generalist [GGCL22, RHSM21, vdRJL20]. generalists [GLM^{+}24]. generalized [WYO^{+}20]. generating [GVS23]. generation [FOO20, YAZA21]. generational [LPB^{+}23]. generations [UPK^{+}20]. generic [BBS^{+}21]. genes [PMH23]. Genetic [HBJ^{+}21, KI24, MTK^{+}21, MON^{+}24, SOS^{+}21b, YRB^{+}20, AKW^{+}22, BNSH21, BNS^{+}24, CNC^{+}20, CDP^{+}21, DFB^{+}24, EPB23, HOK^{+}20, HB24, LGGR^{+}22, MGHH21, PAL^{+}22, SOS^{+}21a, TPZ^{+}21, TKM^{+}22]. genetics [PLH^{+}21, YMU^{+}21]. Geniculate [CCC^{+}23]. genomic [LO20, SLG^{+}24]. genotype [IBRL^{+}22, KDLL21]. Genotypic [AVFD^{+}21]. gentoo [MTCT^{+}23]. genus [DMBJ21]. geochemical [GCDP21, OYO^{+}20]. Geographic [BKA^{+}20, LBH^{+}21a, SSFS20, TPGG^{+}23, ZZX^{+}20]. geographical [Ano24o]. geolocators [DAT^{+}21]. geological [BGL^{+}21]. geomorphic [WJ21]. Georgia [MCL24, IMM^{+}20, JNY^{+}20]. Georgian [BDC^{+}22]. germinans [LPD^{+}24]. germination [CCMRII22, XWW^{+}21]. Giant [BSRVS22, ASPM21, BE21, BLWJ20, Det24, DMM^{+}24a, GDCP20, KLA21, LDPH24, NM22, NBW^{+}23, SLG^{+}24, SCK23, VGG^{+}23]. giants [SB20a]. Gibraltar [MMMV22]. gigas [BSRVS22, GSP^{+}20, LK21, OAM^{+}24, PMH23, SCK23, VSGD24]. gilberti [GLP^{+}24]. gillnet [CWC^{+}22, CCT^{+}23]. gillnets [GG23, Sig23]. Glacial [KI24]. glaciale [BG20, BG21]. glacialis [FNL^{+}23, HEKH24, MHN24, TEK21]. glacier [BSB^{+}21]. glaciers [MKS^{+}20]. gladii [JBGA23]. gladius [APOGG20, HTT^{+}21, SSB^{+}23]. glass [BKM^{+}21, HMK^{+}22]. glauca [EMDR22, FSH^{+}21, FKS^{+}22]. gleaned [RCC^{+}24]. Glimpsing [MCL24]. GLMF1 [PLG^{+}23]. Global [Ano23, AC20, LÁGLL20, RPB21, AZ24, Ano24o, CCAdA24, LBMSS23, MW20, NBDM^{+}24, PSN^{+}20, RB22]. globally [CAW^{+}22, CGS^{+}23]. Globicephala [BBPO21, BMS^{+}24, SBS^{+}24]. glomerata [GSP^{+}20]. gobies [BSH21a, BQC^{+}22, DMBJ21]. Gobiidae [DMBJ21]. Going [FQVMF23]. Gorgona [CRLC21]. governance [NDA^{+}22]. gradient [ATR^{+}24, GSL^{+}24a, GLB^{+}20, GLM^{+}24, IAG^{+}23, LSC^{+}21, LKB^{+}21, RWSR24, TSW^{+}21, VSG20]. gradients [CMO^{+}20, NCZ^{+}21, RC23, SMA^{+}20, SSFS20, SRR^{+}20, WBS24]. gradual [NBDM^{+}24]. graeffei [HP23]. grandis [DFB^{+}24]. graph [OAdJA^{+}20]. grapsoid [SAH^{+}24]. gray [CRGMA^{+}21, DME^{+}23, HPRP21, HSO^{+}22, JFB^{+}23, MSC^{+}20, MHDJ21]. Grazer [IBC^{+}21, ASS^{+}23, BK23, Det24, RS21]. grazer-mediated [Det24]. grazers [MR20, NBDM^{+}24]. Grazing [CS22, BPSB^{+}22, CBJ^{+}23, GBC^{+}20, LSSG24, LHD20, MMD20, NM22, RS21, SKI^{+}20, WLB23]. Great [CCW21, CSDB22, MHK22b, MRH20, RCES22, SMH^{+}20, SYMR21]. Greater [CSK^{+}22, BFC^{+}20, HTL^{+}20, SHT^{+}22, TCY^{+}22, GBE^{+}22]. Green [RH20, SHM^{+}20a, BDB^{+}20, CBJ^{+}23, CGM23, GFT^{+}21, GSL^{+}24b, HKK^{+}23, LWP^{+}24, MJ23, MLE^{+}22, PLG^{+}23, PLC^{+}20, QRC^{+}24, RVCT21, RMR^{+}23, SYMR21, STH21, SBC^{+}21, VPH^{+}21, ZSWR^{+}22]. green-lipped [QRC^{+}24, STH21]. Greenland [BTMS24, ENRC^{+}20, FFT^{+}20, HLA^{+}21, LMRS^{+}24, MHN24, MKS^{+}20, MPF^{+}22, SJS^{+}23]. Grey [BHQ^{+}22, FOO20, SSC^{+}23]. grey-headed [SSC^{+}23]. grisea [BLCC^{+}20, FNM^{+}21]. grooming [SJSR24]. Gross [CRLC21, CS22]. ground [CAW^{+}22, LGR^{+}22, PSÓ^{+}24]. groundfish [BBKW20, EWF^{+}21]. grounds [AOO^{+}21, CBTL20, CCT^{+}23, DMS^{+}21, HIT^{+}22, vdVWF^{+}24]. group [PLLT^{+}21, SSG^{+}22]. grouper [BBH^{+}20, FHAP20, GBE^{+}22, LBMB^{+}20, SRS^{+}20, WSPS^{+}20]. growing [CBHM21]. Growth [BSF24, CHJ21, CKCS20, CMvKdR23, EHLM^{+}23, JWH^{+}21, KETL22, LJ20, ZLL^{+}24, ASBE22, ABC^{+}24, AAHW20, ARH^{+}20, ASH^{+}20, BRC22, BVD^{+}20, BG20, BG21, BNH^{+}21, BKT^{+}22, BNR20, CTH^{+}20, CBB^{+}22a, CS22, DCG^{+}20, DV20, DFCM21, FS21, FMD20, FDE^{+}21, GVS23, GSRC20, GB21, GPCM22, HP23, HBR^{+}21, HTT^{+}21, HAB^{+}22, HIT^{+}22, JPMG23, JNJ^{+}20, JNY^{+}20, KCGR20, KSW23, LDW^{+}21, MBB^{+}22, MJ21, MWCR20, NHFS22, NMB^{+}22, PMH23, RHM^{+}22, RPB21, RHH^{+}20, RTT^{+}21, RL22, RC23, Rob21, RSL^{+}22, SBSS21, SLH^{+}22, SMK23, SdFBM24, SJ22, SAH^{+}24, THS^{+}24, TGB^{+}20, VLPN24, YDG21, dFRvdG^{+}22, dMGGG21]. growth-selective [KCGR20]. grylle [JFR^{+}21]. grypus [BHQ^{+}22, FOO20, HPRP21, MHDJ21]. guarantee [WRS^{+}23]. Guiana [MCLC^{+}23]. guianensis [MCLC^{+}23]. guide [VMCA^{+}20]. guillemots [JW22, JFR^{+}21]. Gulf [DV20, AHMVM23, ASH^{+}20, AKGA^{+}21, BMO^{+}24, BTMS24, BLGM^{+}24, CKS^{+}24, DFB^{+}24, DSK^{+}20, ESG20, FHAP20, FNL^{+}23, GBW^{+}20, GSS22, HAB^{+}22, HPH20, LJ20, MDS^{+}23, MYSF23, MGA^{+}20, MSF24, OLA^{+}24, PTCD22, RSL^{+}22, SRS^{+}20, VOP21, WRG^{+}24]. gull [ASS^{+}21b, EVVMQ^{+}23, MIL^{+}20]. gulls [CVP^{+}21, DVMH^{+}24, JTBS^{+}22, KKMW23, PVL^{+}21]. gummy [PBO^{+}21]. gut [FWA^{+}23, HRH22]. Gymnoscopelus [KRG^{+}20]. Gymnothorax [DHM21]. Gyre [DL20, vLBGGT22].

H46 [YHW^{+}23]. Habitat [ASIA22, AGIS^{+}22, ASGI24, BLCC^{+}20, CHD^{+}22, DMBJ21, FSK^{+}21, LCT^{+}20, MCLC^{+}23, ORB^{+}21, SHS21, TPGG^{+}23, VvDRM20, WIM23, ASY^{+}24, BRC22, BMO^{+}24, BMK^{+}22, BICQG23, BCP^{+}22, BZT^{+}23, BKA^{+}20, BDS^{+}20, BOY^{+}23, BGDJ20, CCRP20, CFHCS^{+}24, CBWH24, CLR^{+}21, CPH20, CP22b, CBB^{+}22b, CMO^{+}20, CCP^{+}21, CLF^{+}20, DHM21, DFHT23, DKAB23, DIC^{+}22, ECvBL21, Fai21, FS21, GBE^{+}22, GDAC23, GBF^{+}23, HDB21, HGT^{+}21, HO21, HBG^{+}22, HHY^{+}23, HCS^{+}24, JBWS24, JFB^{+}23, KLA21, LBH^{+}21c, LFK^{+}21, LHM^{+}23, LCR^{+}22, LFFW21, LTM20, LLF^{+}20, MCvdM^{+}24, MDK^{+}21, MLB^{+}21, NT24, NCS^{+}23, NGRH20, NSB^{+}23, OVO^{+}21, OBS^{+}22, OLG^{+}23, PBAC23, PSN^{+}20, PCR23, PGT^{+}21, PDF^{+}23, PAT^{+}24, PSP^{+}20, QHT^{+}21, RGL24, RBF^{+}22, RLSO23, RTG^{+}20, SRHM21, SRG^{+}24, SHH^{+}22, SNH^{+}20, SHT^{+}22, SDM^{+}23, SHL^{+}21, SdGWN21, TCY^{+}22, THR^{+}23, TWC22, TR24]. habitat [VGLL^{+}21, WAA^{+}23, WYZ^{+}20]. Habitat-based [ORB^{+}21]. habitats [AMRL^{+}24, BS21, BFC^{+}20, BSP^{+}22, CSK^{+}22, DKS^{+}24, DBHCO^{+}22, DCR^{+}21, FWGD^{+}23, GAFS20, GSS22, HNN^{+}24, KCLJ20, LR22, MJ23, MON^{+}24, MPKS24, PLLT^{+}21, RF20, TTJS20, VMCA^{+}20, WK21, Whi20]. habits [GGN21, SLB^{+}21, TFCS22, WBG^{+}22, vdRJL20]. haddock [DSK^{+}20, FLY^{+}24]. haemoglobin [MTCT^{+}23]. hair [LNBdB21, SLB^{+}21]. hake [GFDPASR20, MHH^{+}20]. Half [BD20, FDL^{+}21, vdVTW^{+}22]. half-century [FDL^{+}21]. halibut [BTMS24]. Halichoerus [BHQ^{+}22, FOO20, HPRP21, MHDJ21]. Halimeda [CSBM20]. Haliotis [HNKK21, MY21]. Hamilton [Ano22f]. hammerhead [BLGM^{+}24, EMGMEV^{+}21, GGN21, HCP^{+}24]. hammerheads [GLP^{+}24]. Hampshire [CCW21]. Handling [HHY^{+}23, SFOPL21, SWHE20]. Handling- [HHY^{+}23]. hanleyanus [CGL^{+}22, RPD^{+}23]. hannai [HNKK21, MY21]. haplotypes [SVT^{+}24]. Harbor [MMT^{+}24, CLR20]. harbors [SWP20]. Harbour [ATLT20, HJG21, NRE21, SBS^{+}22, SvBS^{+}20, SWC^{+}23b, TMN^{+}24]. hard [LSS^{+}22, OJ24]. hard-substrate [OJ24]. hare [TBS20]. harengus [BSF24]. Harmful [LAL21, LWP^{+}24, BAS23, LSSG24]. harp [KdlVJ^{+}21, KGG^{+}21]. harpacticoids [ACG22]. harvested [ATR^{+}24, JSK19]. harvesting [CEZH21, EHLM^{+}23, FNM^{+}21]. hatchery [CSK^{+}23]. hatchery-origin [CSK^{+}23]. Hatching [GKK^{+}22, BJM^{+}22, CGW21, CMT^{+}21, WHRC22]. hatching-enzyme [CMT^{+}21]. hatching-gene [CMT^{+}21]. hatchlings [BPS^{+}21, MSR^{+}21a, SBC^{+}21]. Hatteras [FPB^{+}21]. hauling [Sig23]. haulouts [TMN^{+}24]. Hawai‘i [OLPGK22, HDA^{+}22, JBDH21, LD24, SON^{+}20]. Hawaiian [MMB22, LFK^{+}21, MSPS22, VOM^{+}20]. hawksbill [BASM^{+}24, DGM^{+}23b, MJJJ20]. headed [SSC^{+}23]. heads [RBM22]. health [KCLJ20]. hearing [SFMK^{+}23]. heat [BB24, BBM^{+}24, CLD23, DBS^{+}20, ICIS23, MHAL24, ODB^{+}24, OA24]. Heating [SSFD^{+}23]. Heatwave [OLPGK22, CMTB^{+}22, KNH24, MSF24, OOW^{+}20, RCB^{+}24a, SSFD^{+}23, SAMP24, TSS^{+}23, WPG^{+}24]. heatwaves [BBM^{+}24, BICQG23, CKT^{+}24, DGP22, KGM^{+}24, NBDM^{+}24, RBS^{+}24, WS24, WH24]. heavily [MCLC^{+}23]. Heavy [WSU21]. Hector [EWBH24, OCC23]. Heermann [DVMH^{+}24, EVVMQ^{+}23]. heermanni [EVVMQ^{+}23]. height [KKMW23]. Helicolenus [KWD21]. Heliocidaris [BKTN21]. helminth [RBCAV22]. Hemigrapsus [RAA^{+}23]. Hemisphere [CPD^{+}20, ZPVN23, SVB^{+}24]. hemispheric [ZBB^{+}20]. Henry [LGM21]. herbicides [KSW^{+}21]. herbivore [FQVMF23, NSMJ20, RWSR24]. herbivores [CAV^{+}20, MR20]. herbivorous [AKU^{+}22, CFFC^{+}23, FMD20, PCM^{+}20, SKI^{+}20]. Herbivory [BFC^{+}20, RTT^{+}21, RH20, SYMR21, VPH^{+}21, YDG21]. Herdmania [BSP^{+}21]. heritage [SLG^{+}24]. Hermaphroditism [GVCSO^{+}24]. hermit [RBFA20]. heros [HMT21]. herring [BAF20, BSF24, DMCB22, FN24, HGH^{+}24, MAK^{+}22, MBM^{+}20, MGM^{+}24, PSÓ^{+}24, RBC^{+}24, VBB^{+}21]. heteroclitus [RB20]. heterogeneity [BGGH^{+}22, MHK22a]. heterogeneous [CKS^{+}24, TOMDC23]. Heterosigma [BAS23]. heterotrophic [CS22, ZSC^{+}22]. hidden [TKM^{+}22]. Hiding [TNMN24]. hierarchical [RPB21]. hierarchy [XCT^{+}20]. High [BG21, BSP^{+}21, DSHM20, DPL^{+}21, FSF^{+}24, HCP^{+}24, JPMG23, JMS^{+}22, VPH^{+}21, APG^{+}22, ARE^{+}24, ACA^{+}24, ARB^{+}20, BJM^{+}22, BPS^{+}24a, DMS^{+}21, DWJL^{+}21, FCV^{+}24, GDAC23, HHM^{+}22, KRMR24, LKR^{+}21, LBTE^{+}21, MBS^{+}24, MSR^{+}21a, ODB^{+}24, SHH^{+}22, STO^{+}23, SBC^{+}21, ZCBC24]. high-density [KRMR24]. high-latitude [MBS^{+}24, ODB^{+}24]. high-temperate [ARB^{+}20]. high-use [ACA^{+}24]. higher [CCJ^{+}23, DVP^{+}23, PBKG22, RB23]. highlight [COM^{+}24, DVR^{+}23, GRTK^{+}21, LR22, MLB^{+}21]. highlighted [RDL^{+}22]. highlights [NCS^{+}23]. Highly [MDM^{+}20, SAMdL^{+}20, SRR^{+}20, TRK^{+}22, UBO^{+}24]. Hiroshima [OAM^{+}24]. hispida [BOY^{+}23, LHM^{+}23, VGLL^{+}21]. histamine [JPJ^{+}22]. Historic [BNSH21, WRB21]. historic-range [WRB21]. Historical [BNS^{+}24, EWF^{+}21, CNC^{+}20, CMTP22, MGB^{+}22, PAL^{+}22]. histories [CTH^{+}20, CBB^{+}22a, CLGH24, XWH^{+}22]. history [AKW^{+}22, Ano23, AR21a, BQM^{+}22, BSH21a, BAF20, BPS^{+}24a, DVMH^{+}24, FS21, FYM^{+}24, FSDB20, FKS^{+}22, KI24, LBH^{+}21a, LSR^{+}24, LPB^{+}23, MHH^{+}20, ROMB22, SRS^{+}20, SSK^{+}23, TCP^{+}23, WH24]. hogfish [FS21]. hoiho [YSP^{+}22]. Hokkaido [ISMM20]. holobiont [DPL^{+}21]. holobionts [ZBB^{+}23]. Holothuria [MMLPP^{+}24]. holothurians [THD^{+}22]. Homarus [CGW21, KCB^{+}24, SJSR24, SBL^{+}23, WBGC24]. home [BBH^{+}20, DFP^{+}21a, GMS^{+}21]. homing [SHM^{+}20a]. Horizon [MTB^{+}23]. Horizontal [FKTK20, ASPM21, FNO^{+}23, SBJ^{+}24, TKS^{+}20, VDB^{+}24]. horneri [GS20, HOK^{+}20, SE20]. horse [GFDN^{+}24, SHG^{+}21]. horse-eye [GFDN^{+}24]. Horsens [SLO^{+}22]. horseshoe [CCW21, SNH^{+}20, TWC20, WLC22]. hosei [CNC^{+}20]. Host [BBMM20, LLF^{+}20, ATR^{+}24, BMKB22, DDR^{+}21, HvdSR^{+}21, IBC^{+}21, IBRL^{+}22, ICH^{+}24, KHOK22, LCJ22, PBKG22, RBCAV22, SMA^{+}20]. Host-switching [BBMM20]. hosted [WKR21, WNH^{+}20]. hosts [RH23]. hotspot [BRSD20, MW20, TFH^{+}22]. hotspots [HLA^{+}21, JLT^{+}21, RTG^{+}20, SWP^{+}24]. Howell [Ano22c]. HP15 [WSU21]. Hudson [BBL^{+}21a, HDL^{+}21, KTR^{+}20]. Humboldt [ASS^{+}23, CAAG^{+}22, SBDM^{+}23]. Humpback [KDB^{+}24, AU21, ANP^{+}24, BKR^{+}21, BMBC21, BZT^{+}23, BPS^{+}24b, CPD^{+}20, HSA^{+}21, HDA^{+}22, LGR^{+}22, MDK^{+}21, OTD^{+}24, PLLT^{+}21, RFJC^{+}24, RCC^{+}22, SHL^{+}21]. Humpback-krill [KDB^{+}24]. hunger [BKTN21]. Hunting [CGGdD24]. hurricane [Mon23, VGA20]. hurricanes [LCP^{+}20, SBB20, WHB^{+}20, VSGD24]. hurry [KCRA22]. husbandry [HBR^{+}21]. hyalinus [BSH21a]. hydraulic [STM^{+}24]. Hydrobates [BGMAM^{+}21]. hydrodynamic [GKK^{+}22, SLH^{+}22]. hydrodynamics [SRHM21]. hydromedusae [MSTH22]. hydrothermal [CGCC24, FBM^{+}24, ZCB20]. hydrozoans [AHMVM23]. Hypanus [MPSH22]. hyperborea [KMW^{+}21]. hyperboreus [HEKH24]. Hyperoodon [DFHT23, WIM23]. hyposalinity [RVV^{+}24]. hypotheses [MJK24, RPD^{+}23]. hypothesis [CCAdA24, CSDB22, KKH22]. Hypoxia [SRG^{+}24, WC21, BSF^{+}20, DMM^{+}24b, EHHF^{+}20, SW23a, WGK22, vDTS^{+}22].

Ian [VSGD24]. Iberia [GGRV23]. Iberian [FSAG22, GMRMG22, CVP^{+}21, TMPPG24]. Ice [KGG^{+}21, BBL^{+}21a, BOY^{+}23, CSD^{+}24, DKSS20, DDG^{+}21, GSL^{+}24a, GB20, HHM^{+}22, IAG^{+}23, JFB^{+}23, KTR^{+}20, KHG^{+}21, LHM^{+}23, LHHH^{+}20, MKR^{+}20, NF23, OBD^{+}21, ORB^{+}21, SdFBM24, WRW20]. Ice-algal [KGG^{+}21]. ice-rafting [NF23]. Icefjord [SJS^{+}23]. Iceland [SBS^{+}22]. Icelandic [LGR^{+}22, SBS^{+}24]. ICES [TOMDC23]. ichthyoplankton [SCM^{+}23, TTA20]. iconic [DMM^{+}24a]. identification [COT^{+}21, MTK^{+}21]. identifications [TCM^{+}20]. identified [BBSM^{+}20, MJJJ20, SSF^{+}22]. identifies [CCMRII22, CWD24]. identify [FOO20, VNG^{+}22]. Identifying [MSPS22, SON^{+}20, ACA^{+}24, PYB^{+}21]. identity [GLG21]. Idioteuthis [JJV^{+}21]. If [BMKB22]. ignis [LSC^{+}21]. ignobilis [DFP^{+}21a, DMM^{+}24a, GDCP20]. ignored [VST24]. II [BBM^{+}24, KSW^{+}21, OTD^{+}24]. II-inhibiting [KSW^{+}21]. Illex [ASIA22]. illumination [BSFM20]. illustrate [PWvdM^{+}23]. Ilulissat [SJS^{+}23]. imagery [CP22b, LS21, VDC^{+}24]. imaging [OLG^{+}23, SME^{+}23]. imbricata [BASM^{+}24, DGM^{+}23b]. immature [ABN^{+}21, DMTP22, EMB^{+}22, LJ20, SMB^{+}20]. Impact [CLV^{+}20, CEZH21, GCDA20, LGR^{+}22, MHN24, SSC^{+}23, BWC^{+}22, CGLC20, GVCSO^{+}24, GTH^{+}20, KNO^{+}21, KDGG24, SOSE22, WCL20]. Impacts [AWB23, AKU^{+}22, CGM23, FST22, MKS^{+}20, MLAS22, SLBK^{+}20, SE20, WIY^{+}23, WH24, Ano23, BKM^{+}21, BAH^{+}22, BBBT^{+}20, BBRM22, DSHM20, DHA^{+}20, FNM^{+}21, GVR^{+}20, HPH20, JPL^{+}24, KDK^{+}24, KSW^{+}21, LWP^{+}24, RCES22, RWSR24, RB20, RSvV^{+}23, SAMP24, SWHE20, TAN^{+}22, WLC22, vLBGGT22, BBM^{+}24]. impair [MHF^{+}21]. Impaired [BJM^{+}22]. imperial [WHN^{+}22]. implication [FBC^{+}24c]. implications [AGA^{+}22, APG^{+}22, ARE^{+}24, CM22, CSBM20, CSK^{+}23, DFP^{+}21a, DPL^{+}21, GBW^{+}20, GBA^{+}22, HCP^{+}24, HBG^{+}22, JHK^{+}22, KCLJ20, LBH^{+}21a, MJW24, MGA^{+}20, MSF24, MNG^{+}20, MLAS22, NCZ^{+}21, NDA^{+}22, RBS^{+}24, SHG^{+}21, SLH^{+}22, SBS20, SBW^{+}22, SBRQ21, SHL^{+}21, TB21, VNM^{+}21, XWW^{+}21, YSP^{+}22, vESB^{+}23]. implies [HSNT24]. Importance [CGM^{+}24, GFDN^{+}24, BGGH^{+}22, COM^{+}24, CBTL20, DVR^{+}23, HHS21, NSMJ20, OCGAP^{+}22, PVL^{+}21, RCB^{+}24a, WKAH22, ZZYB21, vdWDH^{+}21]. important [CGS^{+}23, CWD24, DKAB23, GTH^{+}20, HMW^{+}21, LCFJ20, MKR^{+}20]. imprints [BNS^{+}24]. improve [AJH22, GJK^{+}24, RHN^{+}24]. Improved [SLO^{+}22, DTHC24, EPHS23]. improvements [BRD20, SGW^{+}21]. improves [SNH^{+}20]. Improving [BTDR20]. in-water [PLC^{+}20]. incidence [BMP^{+}22, IBC^{+}21, JPMG23, LLF^{+}20]. Incident [LMHS21]. incidents [FRO^{+}21]. include [PK24]. including [HDM^{+}24]. inclusion [TOMDC23]. incomplete [CTH^{+}20, PFDF^{+}21]. Incorporating [BGDJ20, FBC^{+}24a, MDPL20]. Increase [Sig23, CBGB24, DPP^{+}21, DWJL^{+}21, HBR^{+}21, KVdM^{+}22, MTCT^{+}23, NHFS22, OOW^{+}20, PDM22, RSL^{+}22, SWHE20, WCS^{+}21, YBSH23]. Increased [KDGG24, LGD23, dFRvdG^{+}22, CCJ^{+}23, CBP22, DGM^{+}23a, LHHH^{+}20, PFDF^{+}21, PSP^{+}20]. increases [ADF20, BMM21, BVD^{+}20, HOK^{+}22, LSSG24, MRB^{+}24, RS21, RCB24b, STO^{+}23]. Increasing [ENRC^{+}20, RS21, AR21b, BP24, CCMRII22, STM^{+}24, YRB^{+}20]. increasingly [Ano24o]. independent [HBG^{+}22, RTG^{+}20]. index [MTB21, RBVA23, STO^{+}23]. Indian [ECvBL21, MFC^{+}24, RBF^{+}22, AAFC^{+}21, CHD^{+}22, FLC^{+}24a, GDCP20, LHD20, PBS22, SAC24]. indicate [GSK^{+}20, LDW^{+}21, SHM^{+}20a, dDW^{+}22]. indicates [DMTP22, LBTE^{+}21, RCB^{+}24a]. Indication [BMBC21]. indicator [BBS^{+}21, JPL^{+}24, TD21]. indicators [DFRS^{+}24, JS20, KF20, LORB^{+}24, SB20b]. indices [AUS^{+}23]. indigenous [RDL^{+}22, VOM^{+}20]. indirect [Ano23, SBB20, WHRC22]. Individual [ACC^{+}20, ABD^{+}21, Ano23, CSN^{+}23, GWDR20, GBF^{+}23, HSU^{+}21, KCB^{+}24, LKR^{+}21, LLJ^{+}20, MSV21, PDF^{+}23, RMASA20]. individual-based [HSU^{+}21]. individual-level [KCB^{+}24]. individuals [ABB^{+}20, MAS^{+}20]. Indo [MDH^{+}21]. Indonesia [NBW^{+}23]. induce [JNJ^{+}20, MY21, MPSH22]. induced [BBS^{+}21, BPA21, CSH^{+}21a, MRB^{+}24, MAS^{+}20, STO^{+}23, SFLQ23]. Inducible [SFOPL21, VSG20]. induction [GSN^{+}20]. industrial [PRM^{+}21, SSB^{+}23]. inefficient [MB20]. infauna [HMV^{+}22, Mon23]. Infaunal [CIDM23, GWA^{+}23]. infaunalisation [HGT^{+}21]. infection [RBCAV22]. infections [BKT^{+}22, SdFBM24]. infectious [BAB^{+}24]. infer [CAP^{+}21, GBA24]. inference [KKS^{+}20, RCB24b]. inferred [BKH^{+}24, GKK^{+}22, JHK^{+}22, PLC^{+}20, dLCN^{+}22]. infesting [LCJ22]. inflatus [ZDBS20]. inflow [DHJ^{+}22, OSJ^{+}22]. Influence [AMUPJR^{+}24, CVP^{+}21, CSH^{+}21b, DMCB22, FOM^{+}20, HMV^{+}22, KBWRV21, SvBS^{+}20, VOM^{+}20, WAA^{+}23, ZCM^{+}21, ACA^{+}24, BOY^{+}23, DBHCO^{+}22, GLB^{+}20, GRTK^{+}21, JPB21, KBP22, KPTT23, KWD21, LSC^{+}21, LFK^{+}21, MCK21, MIL^{+}20, RS20a, RDT^{+}23, RGK22, SBJ^{+}24, SDE^{+}22, TBS20, TWC20, OTD^{+}24]. influenced [BMBC21, CAW^{+}22, Hew21, RJ21, SRHM21, SHH^{+}22]. Influences [KDLL21, KCB^{+}24, DBK^{+}23, FBKG24, IBC^{+}21, KLFP23, LLF^{+}20, LLL20, LCGS^{+}21, MBB^{+}22, MA20, SHB^{+}20, SRG^{+}24, TCM^{+}23]. influencing [LBMF^{+}24]. inform [DFHT23, JLT^{+}21, SON^{+}20, WSB^{+}24]. information [CAP^{+}21, KTK21]. ingested [KCGR20]. ingestion [CS22, GDAC23, MAK^{+}22, RBB^{+}21]. Inglefield [MKS^{+}20]. inhabiting [LCP^{+}20, dSGC^{+}23]. inheritance [PCP^{+}20]. inhibiting [KSW^{+}21]. inhibition [BE21]. inhibits [JMS^{+}22]. initial [MG22]. injections [DDB^{+}24]. Injured [SST20b]. injuries [BMP^{+}22]. injury [GCCM20]. Inlet [KSS^{+}20, SAMP24]. inner [EVAB^{+}23, RDL^{+}22, SHG^{+}21]. inorganic [DDR^{+}21, JRVE^{+}22, LAL21, RHH^{+}20, SST^{+}20a, SCW^{+}24, VS23]. input [IAG^{+}23]. inputs [MPR^{+}23]. insensitive [BSP^{+}22]. Inshore [HO24, FHAP20, HSNT24, HEP^{+}24, MHK22b, MLP^{+}24]. inshore-offshore [FHAP20]. inshore-reefs [MHK22b]. insight [FNO^{+}23, GVAZ^{+}22, MSH20, SLB^{+}21]. Insights [SBS^{+}22, AFTW24, CGR^{+}20, DIC^{+}22, FKS^{+}22, GFT^{+}21, GSL^{+}24b, JPJ^{+}22, LWP^{+}24, NSB^{+}23, OMW^{+}20, PAL^{+}22, QHT^{+}21, QRC^{+}24, RBFA20, SHZ^{+}20, ZMC^{+}22].

Instantaneous [SAMdL^{+}20]. insularis [LÁGLL20]. intake [FPG^{+}21]. Integrated [CTH^{+}20, DPDJ^{+}21, GBH^{+}20]. Integrating [VDB^{+}24, SAGG^{+}24]. Integration [AFP^{+}24, BTA^{+}20, CC22, GED20]. integrative [FLBP^{+}21]. intensity [CGCC24, LT21, MDG^{+}21, WCS^{+}21, WCRTT20]. intensive [WLB^{+}20]. Inter [ABD^{+}21, DRP^{+}20, GJM^{+}20, LLJ^{+}20, RLME^{+}21, RMASA20, GSBO21, GWDR20, HDA^{+}22, MRH20, RMR^{+}23, WPM22, dDW^{+}22, dSGC^{+}23]. Inter- [RMASA20, MRH20]. Inter-annual [ABD^{+}21, DRP^{+}20, dDW^{+}22, dSGC^{+}23]. Inter-colony [GJM^{+}20]. Inter-individual [LLJ^{+}20, GWDR20]. inter-island [HDA^{+}22]. inter-nesting [RMR^{+}23]. Inter-population [RLME^{+}21, WPM22]. inter-specific [GSBO21]. interaction [HMB^{+}22, LMS^{+}20, OSBA22, RBFA20, WSU21]. Interactions [CLD23, CILGD^{+}22, BDS^{+}20, CAV^{+}20, CCP^{+}21, DAT^{+}21, dOFBSdJR^{+}22, FLCQ21, FNH^{+}23, GGN21, JHDE20, LFA^{+}23, MSL^{+}24, MDS^{+}23, MNG^{+}20, SDM^{+}22, SBT24, SSK^{+}23, SMS^{+}24, TPGG^{+}23, WJ21, YHW^{+}23, YLH20, YAZA21, vESB^{+}23]. Interactive [LKB^{+}21, TYL^{+}24]. Interannual [AHH^{+}20, BWRP20, BSB^{+}21, HNKK21, HPRP21, JNY^{+}20, MBB^{+}23, SHH^{+}22, vdVVC^{+}21]. Interdecadal [THS^{+}24]. Interdisciplinary [CCL^{+}22]. interface [GFDN^{+}24, TABM21]. interference [NSMJ20]. Interior [NMB^{+}22]. intermediate [RBCAV22, ZCM^{+}21]. Internal [KCT24, LT21, MHF^{+}21]. internesting [AEHD21]. interpreting [CDS23, GON^{+}20]. interrelated [Ano24o]. Interrelationships [MHK22b]. Interspecies [SC23]. Interspecific [BBG^{+}22, XMY^{+}22, YGMNA22, GAF20, TPGG^{+}23]. Intertidal [JDDF21, BK23, BMM21, BTB^{+}24, BMPD20, BGGH^{+}22, CCMRII22, Cha22, CS20, CGM23, CSWM24, CWD24, DGM^{+}23a, GBH^{+}20, HMV^{+}22, LT21, MG22, RH21, RVV^{+}24, ROMB22, SAC24, SHT^{+}22, SGP24, TKP^{+}20, YMU^{+}21, ZCBC24]. intertidal-dominated [CWD24]. intervals [AEHD21]. Intra [GAF20, WPM22, dDW^{+}22, RMASA20]. Intra- [GAF20, WPM22, dDW^{+}22]. intra-individual [RMASA20]. intraguild [FSAG22]. Intraspecific [NSMJ20, RWSR24, GGCL22, KLFP23, MRH20]. Intrinsic [BAB^{+}24, MIL^{+}20]. introduced [Ano24o, LCJ22, OCR^{+}24, VOM^{+}20]. introduction [BBMM20, HBJ^{+}21, DTHC24]. intrusions [WRG^{+}24]. inundation [DGM^{+}23a]. invader [Ano24o]. invasion [CGM23, NCZ^{+}21]. invasions [GRB22, WAA24]. invasive [Ano24o, AMD22, BVKF23, CAPS20, DSHM20, GRB22, GS20, GMS^{+}21, HWA^{+}22, KDGG24, LK21, NCZ^{+}21, NBDM^{+}24, PDM22, RJ21, RAA^{+}23, TGB^{+}20]. Invertebrate [SRHM21, DHJ^{+}22, DL21, FDF20, FSSL^{+}20, HSB20, MW20, OCR^{+}24, VGG^{+}23, WDJ20, vDTS^{+}22]. invertebrates [BLB23, DMBB22, EB21, MG22]. investigate [RWR24]. investigated [MTSO^{+}24]. Investigating [DKS^{+}24, FBC^{+}24b, HTT^{+}21, JTBS^{+}22]. investment [ROMB22]. involved [MAMF21]. Ireland [BBPO21, DHE^{+}23, RWR24]. Irish [FLCQ21]. Irma [VSGD24]. Irminger [PJA^{+}21]. IRMS [HKY^{+}20]. irradiance [BG21, KLN^{+}21, LHHH^{+}20]. irradians [GMB^{+}24]. irrigation [SBC^{+}21, YBSH23]. irrorata [GBA21]. Island [CRLC21, OCC23, SC23, HDA^{+}22, MA20, PBS22, RS21, ZCM^{+}21, DCR^{+}22, MDK^{+}21, NYW^{+}22, OLPGK22, SSC^{+}23]. island-associated [PBS22]. Islands [HKK^{+}23, TKS^{+}20, GMLC22, LCP^{+}20, MYSF23, dACMRZ^{+}20, RJ21, SHM^{+}20a, WSD^{+}20, AEHD21, BDC^{+}22, EVAB^{+}23, FPG^{+}21, FAJAO^{+}21, Fig21, MMB22, NBL^{+}22, THD^{+}22, VMDA22, WSPS^{+}20, vTLG^{+}21]. isolated [GIDP^{+}21, PVB^{+}20, SW23b, SHM^{+}20a]. Isolation [DFB^{+}24, HHBP23, JMP^{+}24]. isolation-by-distance [JMP^{+}24]. isopod [LCJ22]. Isoprene [HDE21]. isotope [APOGG20, BKH^{+}24, BPS^{+}24b, CKM^{+}20, CAPS20, DSK^{+}22, DFSH23, EVVMQ^{+}23, FS21, FWWH22, FNL^{+}23, GFT^{+}21, GVAZ^{+}22, GB20, GWDR20, HKY^{+}20, HRH22, JBGA23, KHL^{+}22, KdlVJ^{+}21, KGG^{+}21, LH22, LWP^{+}24, LNBdB21, OCC23, OCR^{+}24, OTD^{+}24, PBB^{+}20, PSN^{+}20, PLC^{+}20, PWvdM^{+}21, RDL^{+}21, SBS^{+}24, VMDA22, VNG^{+}22, WYO^{+}20, XCT^{+}20, ZMC^{+}22, dLCN^{+}22, vTLG^{+}21]. isotope-inferred [BKH^{+}24]. isotopes [BCC20, CVG^{+}23, CBTL20, DCR^{+}22, DSA^{+}24, GPP^{+}24, GDAC23, HMT21, HMW^{+}21, LKR^{+}21, LPD^{+}24, MBW^{+}20, PVB^{+}20, PWvdM^{+}23, RVSA^{+}21, SDW20, SDM^{+}22, SLB^{+}21, SB20b, WWH^{+}22, ZMD^{+}21]. Isotopic [DCL21, VOP21, BDC^{+}22, CFFC^{+}23, CC22, CCA^{+}21, DHM^{+}24, HSA^{+}21, ICH^{+}24, MON^{+}24, PCP^{+}20, RBF^{+}20, RMD24, XW24, YGMNA22]. Istchee [DNE^{+}24]. Isthmus [CLR20]. Istiophorus [BGDJ20]. Isurus [VDB^{+}24].

jack [GFDN^{+}24]. James [DNE^{+}24]. Jan [WIM23]. Japan [AAI^{+}24, AOO^{+}21, FSK^{+}21, FKTK20, HKK^{+}23, HOK^{+}20, ISMM20, KNO^{+}21, OAM^{+}24, STF^{+}24, TKS^{+}20, YOI^{+}23]. Japanese [FYK^{+}20, NTN^{+}24, THK^{+}22, YMU^{+}21]. japonica [HMK^{+}22, ILON21, ZZX^{+}20]. japonicus [KI24, NHFS22, THS^{+}24]. Jardines [NMAE^{+}22]. jellyfish [FSS23, LLL20, RL22]. jellyfishes [DPL^{+}21]. Jiaozhou [ZSC^{+}22]. Johnson [Ano21c]. Joint [URSS24]. Joseph [RH20]. jurisdictions [CBVA^{+}21, YSD^{+}21]. Jusufovski [Ano21d]. Juvenile [BASM^{+}24, OBD^{+}21, VGA20, dMGGG21, AHCT22a, AHCT22b, CBHM21, CFHCS^{+}24, CMdL^{+}20, CCW21, COGGH20, CMF21, DML^{+}21, DGM^{+}23b, FS21, FYM^{+}24, FSDB20, GFT^{+}21, GSL^{+}24b, GBG^{+}23, GSS22, HBR^{+}21, HO24, HCS^{+}24, JNJ^{+}20, JNY^{+}20, LPJ^{+}21, LWP^{+}24, LSSL21, MCvdM^{+}24, MJJJ20, MVH^{+}21, MHF^{+}21, MMS^{+}22, NM22, RHM^{+}22, RWR24, SHB^{+}20, SKTO23, TFCS22, TPGG^{+}23, TCT^{+}23, VDC^{+}24, WDJ20, Whi21, ZXL^{+}24]. juveniles [DBS^{+}20, FSF^{+}24, HTGG22, KCGR20, KETL22, PK24, THS^{+}24].

Kaartvedt [DFRS^{+}24]. Kaikōura [EWBH24]. Kara [DSB^{+}21]. Karenia [CP22a]. Kelp [KNH24, AAI^{+}24, AVFD^{+}21, ARdlB^{+}22, AMD22, ASS^{+}21b, BSF^{+}20, BE21, BSRVS22, BAHG23, BLWJ20, CLL^{+}22, DHM21, Det24, FBKG24, GPMH^{+}24, GSBO21, GSK^{+}20, HSSN^{+}21, KMW^{+}21, KDGG24, LHFD^{+}24, LCT^{+}20, LDPH24, LPB^{+}23, MJ23, MBB^{+}22, MJW24, NM22, NFDF^{+}20, PFDF^{+}21, PDM22, SE20, TSS^{+}23, VGG^{+}23, WK21, WCRTT20, WD22, WCP^{+}21b, YLH20, YDG21, Zue22, dBDL^{+}20]. Kelp-associated [KNH24]. Kemp [ARH^{+}20, LJ20]. Kerguelen [BDC^{+}22]. Key [LORB^{+}24, UPK^{+}20, ASPM21, DDG^{+}21, JHDE20, LGGR^{+}22, NSMJ20, PWvdV20, PAT^{+}24, RAF^{+}22]. Keys [VSGD24, GBA^{+}22, RBM22]. keystone [SKHC23, SSM20]. kHz [BMOW21, BMS^{+}21]. kill [FFMR24]. Killer [MBM^{+}20, VBB^{+}21, BMBC21, DFP^{+}21b, DME^{+}23, FPG^{+}21, RSvV^{+}23, WYO^{+}20]. Kitikmeot [HYG^{+}20]. kittiwake [MIL^{+}20, OBW^{+}20]. kittiwakes [BSB^{+}21, EGK^{+}21]. knowledge [Ano23, CGM^{+}24, GAPM24, KCLJ20, RCC^{+}24, SCB^{+}20]. known [HGK^{+}22]. Kodiak [MDK^{+}21]. Kongsfjorden [BLW^{+}24]. Kotzebue [LHM^{+}23]. krempfi [SMA^{+}20]. krill [BBEF^{+}20, FJJ^{+}21, GDAC23, HOK^{+}22, IMM^{+}20, KDB^{+}24, MSR^{+}21b, RBVA23, SBSS21, SBRQ21, WRW20, ZZLS24]. Kuroshio [NTN^{+}24].

laboratory [GON^{+}20]. Labrador [BOY^{+}23, KF20]. Labridae [PSBH21]. Labroides [GBC^{+}20]. Lack [RCB^{+}24a]. Lactation [AAP^{+}23]. Lagenodelphis [CNC^{+}20]. Lagenorhynchus [SBS^{+}22]. Lagodon [LOF^{+}24]. lagoon [APG^{+}22, BPSB^{+}22, GCGCAS^{+}20, TWC22, VGLL^{+}21, WC21]. lagoons [CATF^{+}21]. Lagrangian [QFT^{+}24]. lamellosa [CH20]. Laminaria [KMW^{+}21, LPB^{+}23, WF21]. laminarialean [YDG21]. Lamna [WHB^{+}20]. lance [BJM^{+}22, RBS^{+}24]. lanceolatus [CDC^{+}23]. Land [BAH^{+}22, CSH^{+}21b, GFDN^{+}24, SVB^{+}24, TABM21]. Land-dependent [BAH^{+}22]. land-derived [CSH^{+}21b]. land-sea [GFDN^{+}24]. landing [SSC^{+}23]. landings [Str23]. Landscape [TD21]. Landscape-scale [TD21]. lanternfish [FTC^{+}20]. Laptev [DSB^{+}21]. Large [ILON21, LOS^{+}22, VLG^{+}21, ZCM^{+}21, ACC^{+}20, BVKF23, BBH^{+}20, BNR20, CGL^{+}22, CDS23, FSF^{+}24, FNH^{+}23, GMS^{+}21, HPDG22, JMR21, LFK^{+}21, NGPGO^{+}23, NSPH20, NFDF^{+}20, PS23, PWvdM^{+}23, RBCAV22, SMH^{+}20, SLO^{+}22, SNE^{+}24, THR^{+}23, TdLHL21, XW24]. large- [BVKF23]. large-eddy [BNR20]. Large-scale [ILON21, LOS^{+}22, CGL^{+}22, FNH^{+}23, JMR21, LFK^{+}21, NGPGO^{+}23, NFDF^{+}20, PS23, SLO^{+}22, SNE^{+}24, XW24]. Larger [DL21, SDE^{+}22]. largest [BSRVS22, BDB^{+}20, FJJ^{+}21, HGK^{+}22, MB20, RMR^{+}23, vdVTW^{+}22]. largha [YXB22]. Laridae [GQBB^{+}24]. Larus [EVVMQ^{+}23, JTBS^{+}22]. larvae [BCC20, BAF20, BSF24, CHDG^{+}23, CKCS20, DHJ^{+}22, FLY^{+}24, HKY^{+}20, JPJ^{+}22, JMS^{+}22, KPS^{+}22, LBH^{+}21b, MK20, MAOR^{+}21, NHH^{+}21, NP22, OAM^{+}24, PGM^{+}23, PLH^{+}21, RJE^{+}22, SBMW22, SDM^{+}23, SLJ^{+}20, VMCA^{+}20, WMKV21]. Larval [GAPM24, GSP^{+}20, SJ22, AMUPJR^{+}24, ABH^{+}21, ASH^{+}20, ASBA^{+}20, BLVL24, CAP^{+}21, CSDB22, CCL^{+}22, DBHCO^{+}22, EPB23, FVC^{+}20, FST22, GM20, GSC^{+}22, GSRC20, GWC^{+}21, HIT^{+}22, JDM^{+}20, KCGR20, KPK^{+}24, KWD21, LO20, Lei20, MAPFH22, NDA^{+}22, OAM^{+}24, OAdJA^{+}20, PMH23, QFT^{+}24, QRC^{+}24, RSS20, SRHM21, SCM21, SSBA^{+}20, TKS^{+}20, TCT^{+}23, TTA20, TTJS20, WTK^{+}24, WC21, WRG^{+}24, ZDBS20, vdVWF^{+}24]. Last [KI24]. late [BBGM23, CGW21, LBH^{+}21b]. late-stage [LBH^{+}21b]. laticostata [SB20a]. latissima [BAHG23, DB21, GJK^{+}24, YDG21]. latitude [AC20, MHH^{+}20, MBS^{+}24, ODB^{+}24, RAA^{+}23]. latitudes [BPS^{+}24a, KCRA22]. latitudinal [GLB^{+}20, LSC^{+}21, MDB^{+}21, WBS24]. latro [ACvdC^{+}21]. latus [GFDN^{+}24, MSDY23]. launch [Str23]. Lawrence [BTMS24, FNL^{+}23]. layer [Hew21]. layers [BNR20, COGGH20]. Laysan [GFH^{+}22, MCJ^{+}20]. Leach [HBE22]. lead [Ano24o, HDL^{+}21, SGW^{+}21]. leading [BE21]. leads [DCG^{+}20, HBJ^{+}21, RS21]. Leaf [Rob21, HMKK^{+}20]. Leaf-litter [Rob21]. learning [CCT^{+}23, JDM^{+}20, KKC^{+}24, KTK21, MHH^{+}22, SSF^{+}22]. leatherback [AEHD21, BPS^{+}21]. lecithotrophic [ZDBS20]. LED [Sig23]. legged [BSB^{+}21, EGK^{+}21, MIL^{+}20]. lemuru [BRC22]. length [HMK^{+}22, IMM^{+}20, KWD21, KSW23, MVC^{+}23, SDE^{+}22]. length-dependent [IMM^{+}20]. lenses [CAPS20]. leonina [CGGdD24]. leopardus [CAP^{+}21]. Lepeophtheirus [TBH21]. Lepidodinium [LNZL21]. Lernaeenicus [LLF^{+}20]. lesion [KDLL21]. Lessepsian [ASL^{+}22]. lesser [JTBS^{+}22, LBW21, RWH^{+}24]. Lessons [BLVL24, FFMR24]. Lethal [SRK^{+}23, SW23a]. lethostigma [GVS23]. leucas [EMB^{+}22, LBMF^{+}24, MSH20]. leucorhoa [HBE22]. level [ACG22, CCMRII22, DAT^{+}21, HTGG22, KCB^{+}24, KI24, MTB21, NTN^{+}24, NFDF^{+}20]. levels [JRVE^{+}22, LPJ^{+}20, MTCT^{+}23, RVCT21, ZHG^{+}21]. leveraging [HBR^{+}21]. lewini [EMGMEV^{+}21, GLP^{+}24]. lice [MHF^{+}21, SLBK^{+}20, TBH21]. Life [FS21, AKW^{+}22, AFP^{+}24, Ano23, AAHW20, BRC22, BMK^{+}22, BQM^{+}22, BSH21a, BPS^{+}24a, CRÁ^{+}20, CSK^{+}22, CDS23, DVMH^{+}24, FYM^{+}24, FSDB20, FKS^{+}22, HTT^{+}21, LBH^{+}21a, LSR^{+}24, MHH^{+}20, MLB^{+}21, OÁBP^{+}21, QHT^{+}21, RHSM21, ROMB22, SRS^{+}20, SSK^{+}23, TCP^{+}23, WMKV21, WH24, ZXL^{+}24]. life-history [Ano23, BQM^{+}22, FKS^{+}22, LSR^{+}24, MHH^{+}20]. life-stage [WMKV21]. lifespan [MVSV22]. lifetime [VOP21]. Light [DAT^{+}21, BSF24, BRU^{+}22, CEZH21, CGCC24, DNE^{+}24, GED20, HTGG22, JRVE^{+}22, KSW^{+}21, LMHS21, MSL^{+}21, PLG^{+}23, RHM^{+}22, WCS^{+}21, WS24]. Light-level [DAT^{+}21]. lights [Sig23]. likely [BBGM23]. Limanda [DV20]. limbatus [AKGA^{+}21]. Limfjorden [PLH^{+}21]. limit [BMS^{+}24, BLWJ20]. limitation [BRU^{+}22, DBS^{+}20, KBP22, MRB^{+}24, NHH^{+}21, SSK^{+}23]. limitations [GIDP^{+}21, WCS^{+}21]. Limited [GPMH^{+}24, GS20, HP23, BTDR20, COWM^{+}22a, HHY^{+}23, KKS^{+}20, LHM^{+}23]. limits [DGP22]. limpet [IBC^{+}21, KLA21, MK20, OHBJ20, SB20a]. Limulus [CCW21, TWC20, WLC22]. Lindaspio [YWW^{+}24]. line [BSB^{+}21]. lineage [MWCR20]. lineages [BMM21, WKR21]. linear [RS21]. lines [FPG^{+}21]. lingcod [LBH^{+}21a]. Lingering [SAMP24]. link [KTK21]. linked [DVP^{+}23, SKHC23, SWC^{+}23b]. Linking [BRC22, MDQ^{+}21, SBB^{+}24, BQM^{+}24, LHFD^{+}24]. Links [BBEF^{+}20, Ano23, HW20, MAOR^{+}21, RSvL21]. lion [LNM22]. lionfish [CAPS20, DJR^{+}20, GMS^{+}21, HBJ^{+}21, TGB^{+}20]. lions [BBR24, MSF24]. Lipid [HMT21, LBH^{+}21b, TEK21, ZDBS20, ZPVN23, LH22, SSG^{+}22, TBH21]. Lipid-related [TEK21]. Lipids [MPR^{+}23, OCR^{+}24]. Lipophrys [JMP^{+}24]. lipped [QRC^{+}24, STH21]. Lithothamnion [BG20, BG21]. litter [RDT^{+}23, Rob21]. Little [HPRP21, HTL^{+}20, CKT^{+}24, DMS^{+}21, JCGS22, LKR^{+}21, WSPS^{+}20]. Littoral [Whi20, BS21]. Littoraria [GBA21]. Littorina [RH21, Str23]. live [BAS21]. live-bearing [BAS21]. lived [BBH^{+}20, MSV21]. lively [CRÁ^{+}20]. lividus [CCRP20]. Living [GIDP^{+}21, LHFD^{+}24, MDH^{+}20, SF23]. loading [KBP22]. Lobophora [PGM^{+}23, PCM^{+}20]. lobster [ADBF21, BBRM22, KCB^{+}24, MSDY23, PGW23, SBL^{+}23, VNM^{+}21, WBGC24]. lobsters [CGW21, HO24, JBWS24, LMS^{+}20, SJSR24, TdLHL21, TFH^{+}22]. Local [SCB^{+}20, BMW24, BPS^{+}21, CLD23, CGL^{+}22, MGHH21, SDW20, SRG^{+}23, STY^{+}20, SLJ^{+}20, TMN^{+}24]. localised [SBT24]. localized [APG^{+}22, KDGG24]. locating [KPW^{+}20]. location [Ano23]. locations [COM^{+}24, JW22]. locomotion [WLC22]. loggerhead [CCS^{+}22, CTH^{+}20, MAH^{+}20, MSV21, TTM^{+}22, TTM^{+}24, WHRC22]. loggers [CILGD^{+}22]. logging [CVG^{+}23]. Loligo [CZS^{+}22, GMRMG22]. Long [DV20, HDM^{+}24, HEP^{+}24, JPB21, LAP^{+}24, LBTE^{+}21, MSH20, MBR^{+}21, NHFS22, NTN^{+}24, SHT^{+}22, BBPO21, BSF24, BBH^{+}20, CEZH21, CBB^{+}22a, DKS^{+}24, DWJL^{+}21, DDD^{+}23, GSL^{+}24b, GVAZ^{+}22, GMS^{+}21, HP23, MSV21, SBS^{+}24, SGB^{+}21, SFO^{+}21, TGT^{+}23, dLCN^{+}22]. long-distance [DWJL^{+}21, GMS^{+}21, SGB^{+}21]. long-finned [BBPO21, SBS^{+}24]. long-lived [BBH^{+}20, MSV21]. long-range [GVAZ^{+}22]. long-spined [DKS^{+}24]. Long-term [DV20, HDM^{+}24, HEP^{+}24, JPB21, LAP^{+}24, LBTE^{+}21, MSH20, MBR^{+}21, NHFS22, NTN^{+}24, SHT^{+}22, BSF24, CEZH21, CBB^{+}22a, DDD^{+}23, GSL^{+}24b, HP23, SFO^{+}21, TGT^{+}23, dLCN^{+}22]. longest [HPFB20, TTA20]. longevity [MA20]. longfin [FJB^{+}23, GKK^{+}22]. longimana [vTLG^{+}21]. longirostris [SON^{+}20, SON^{+}20, ZFD^{+}24]. Longshore [DMM^{+}24a]. loop [SOSE22]. Loss [GBE^{+}22, EB21, ENRC^{+}20, FS22, MJ21]. losses [GKK^{+}22]. Low [BG20, BDB^{+}20, HSNT24, KLN^{+}21, PRM^{+}21, RH21, TGB^{+}20, WS24, BG21, CCJ^{+}23, FCV^{+}24, HYG^{+}20, JBG22, KCRA22, LSS^{+}22, NT24, RHN^{+}20, SVT^{+}24, SSFS20, VS23]. low-pH [CCJ^{+}23, JBG22]. low-salinity [HYG^{+}20]. low-seston [SSFS20]. lower [DL21, LGD23, TQG24, RBM22]. lowering [HTL^{+}20]. Luck [WHN^{+}22]. lumpfish [MPF^{+}22]. lunar [BSFM20]. lurida [HGK^{+}24]. Lutjanus [BDS^{+}20, DSK^{+}22, ESG20, HNB^{+}24, MPKS24]. Lysmata [ZLHH20].

Lysmatidae [CCAdA24].

M. [PCLQ^{+}20]. Macaronesian [BSÁG^{+}22]. MacGillivray [DCR^{+}22]. Machine [MHH^{+}22, CCT^{+}23, JDM^{+}20, KTK21, SSF^{+}22]. machine-learning [CCT^{+}23]. mackerel [KCGR20]. Macro [THD^{+}22]. Macro- [THD^{+}22]. macroalga [PCM^{+}20]. macroalgae [DKS^{+}24, QMDGM22]. Macroalgal [IAG^{+}23, BTB^{+}24, BPSB^{+}22, HNKK21, KHOK22, SKI^{+}20, TKP^{+}20]. macrobenthic [BD20, GWA^{+}23, TQG24, TGT^{+}23]. macrocephalus [BBPO21, DFHT23, WMT20]. Macrocystis [BE21, BLWJ20, ENF^{+}22, LDPH24]. macroecological [RPD^{+}23, SWJ^{+}21, SS24]. macrofauna [AWB23, ATR^{+}24, BMO^{+}24, BGGH^{+}22, FP20, MDG^{+}21]. Macrofaunal [CSB24, SNE^{+}24, CIDM23, GD23, dBDL^{+}20]. macrofaunal-sized [CIDM23]. macroinvertebrate [CDC21, COWM^{+}22a, SBB^{+}24, SPFL20]. macroinvertebrates [COWM^{+}22b]. Macromolecular [SNP20]. macroparasites [ATR^{+}24]. Macrophyte [BVD^{+}20]. Macrophyte-derived [BVD^{+}20]. macrorhynchus [BMS^{+}24]. macrotidal [TQG24]. macrozooplankton [BMOW21, CSD^{+}24]. maenas [CRB^{+}21, CGM23]. maerl [QMDGM22]. Magallana [GSP^{+}20, LK21]. Magellan [AU21, HSA^{+}21]. Magellanic [BFQ^{+}23, DRP^{+}20, YSD^{+}21]. magellanicus [DRP^{+}20]. magister [HRH22, LO20, RJE^{+}22]. magnets [CGH^{+}23]. magnitude [DKSS21]. main [VNG^{+}22]. Maine [DV20, DSK^{+}20, JBWS24, OLA^{+}24]. Maintenance [SFO^{+}21]. Major [LCP^{+}20, AR21b, CCS^{+}22, DHF24, EWBH24, HIT^{+}22, MCvdM^{+}24, SMPLF^{+}22, Whi20]. makes [CSK^{+}22]. making [dOFBSdJR^{+}22, vLSB^{+}21]. mako [VDB^{+}24]. maladaptive [ŠGH^{+}21]. male [BBR24, BDB^{+}20, CSN^{+}23, YBSH23, dLCN^{+}22]. Mallotus [MGA^{+}20]. mammal [HLA^{+}21, OJPC23, OCR^{+}24, VH20]. mammals [JSK19, LMRS^{+}24, MdOO^{+}20, RAF^{+}22, SB20b]. management [AFTW24, BBBT^{+}20, CSK^{+}23, DVR^{+}23, DTHC24, FTA^{+}21, GPMH^{+}24, GBA^{+}22, HESF24, HNB^{+}24, HCP^{+}24, HvLM21, HBG^{+}22, JHK^{+}22, LBH^{+}21a, LBMB^{+}20, MSR^{+}21b, MGA^{+}20, MTRM21, MDQ^{+}21, NCZ^{+}21, THR^{+}23, TB21]. manatees [CHD^{+}22]. manatus [CHD^{+}22]. mangle [HBC^{+}20]. Mangrove [FAJAO^{+}21, FP20, KHSS23, SMPLF^{+}22, AZ24, CEAO^{+}22, GBA21, GCCM20, GB21, HBC^{+}20, KVdM^{+}22, LPD^{+}24, MDPL20, Rob21]. Mangroves [AMRL^{+}24, MA20, SBB^{+}24, SBB20, ZBB^{+}23]. manta [ACC^{+}20, FPZJAO^{+}22, HGK^{+}22, MAMF21, VvDRM20]. mapping [MDH^{+}20, MMS^{+}20]. Mar [GCGCAS^{+}20]. margin [GWA^{+}23]. margins [CHD^{+}22]. mariculture [TYL^{+}24]. marina [RDL^{+}22, AKW^{+}22, BP24, CAEG21, CEZH21, DNE^{+}24, MGHH21, ROMB22, SLO^{+}22, SNE^{+}24]. Marine [CKT^{+}24, DVP^{+}23, HLA^{+}21, HvLM21, JPL^{+}24, LMB^{+}20, SAMdL^{+}20, SBJ^{+}24, SPFL20, VLG^{+}21, AHCT22b, AGB24, ACC^{+}20, Ano24o, AC20, AKGA^{+}21, Bai22, BMW24, BH22, BTA^{+}20, BAB^{+}24, BWGH23, BBM^{+}24, BICQG23, BPS^{+}24a, BAH^{+}22, BCP^{+}22, BPA21, BRSD20, CBHM21, CDF^{+}20, CLR^{+}21, CDS20, CBP22, CC22, CP22a, COM^{+}24, CSB24, CBB^{+}22b, CFSH22, DTHC24, DCG^{+}20, DGP22, DSHM20, DBS^{+}20, DHJ^{+}22, DSA^{+}24, DHA^{+}20, EB20, EBL^{+}21, Fai21, FYM^{+}24, FTA^{+}21, FBP^{+}22, FST22, FLBP^{+}21, FLC^{+}24b, FNO^{+}23, GM20, GBF^{+}23, HGR^{+}22, HLCH23, HCP^{+}24, HB24, HHBP23, JMM24, JBG22, JFR^{+}21, JSK19, KSW^{+}21, KNH24, KDS^{+}20, KGM^{+}24, LMRS^{+}24, LPO^{+}21, LR22, Lei20, LMH^{+}24, LCJ22, LORB^{+}24, MdOO^{+}20, MSL^{+}24, MKS^{+}20, MSF24, MGBCGM^{+}22, MBGM21, MMT^{+}24, MBG^{+}23, MCFB20, MSV21, NSMJ20, NJHN24, NBDM^{+}24, NSB^{+}23, OSBA22, OJPC23, OOW^{+}20]. marine [OCR^{+}24, OCGAP^{+}22, PK24, PVL^{+}21, PSP^{+}20, RVSA^{+}21, RRS^{+}23, RMR^{+}23, RB23, RBF^{+}20, RLP^{+}23, RCB^{+}24a, RBS^{+}24, RHN^{+}24, RAF^{+}22, SSFD^{+}23, SW23b, SBB^{+}24, SEJ^{+}24, SMH^{+}20, SAMP24, SBW^{+}22, SLBK^{+}20, SCW^{+}24, SAF^{+}24, STM^{+}24, SGW^{+}21, SFO^{+}21, SDE^{+}21, SRR^{+}20, SB20b, TPZ^{+}21, TABM21, Tha21, TSS^{+}23, VH20, VNM^{+}21, WPG^{+}24, WXZ^{+}24, WS24, WSD^{+}20, WH24, XCT^{+}20, YLH20, YBSH23, ZLHH20, ZPVN23, vLSB^{+}21, BMP^{+}22, OLPGK22, SVT^{+}24]. Marinobacter [WSU21]. marinus [LBW21]. Marion [SSC^{+}23]. maritima [SAH^{+}24]. maritimus [HDL^{+}21, KTR^{+}20, SLB^{+}21]. mark [BLE^{+}22, KSMP20, MJ21, SBL^{+}23]. mark-recapture [BLE^{+}22, KSMP20, MJ21, SBL^{+}23]. marked [WCS^{+}21]. markedly [KHG^{+}21]. markers [COT^{+}21, HMT21, KGG^{+}21, MPR^{+}23, MTK^{+}21, VNG^{+}22]. marking [KSMP20]. marlin [RBF^{+}22]. marsh [CAV^{+}20, DBK^{+}23, GBA21, LPD^{+}24, MGH21, NF23, RLP^{+}23, RB20, SLH^{+}22, TSW^{+}21, XYSF20, ZCM^{+}21]. marshes [RDT^{+}23, RSL^{+}22, SAH^{+}24, WLB23, dlBSL^{+}22]. mask [GLB^{+}20]. mass [BPS^{+}21, BMP^{+}22, Hew21, HHY^{+}23, JPB21, JPL^{+}24, SSM20, SWHE20, WCP^{+}21b]. Massachusetts [CDD^{+}20a, PGW23]. masses [RBM^{+}21]. massive [WIY^{+}23]. Mastigias [DPL^{+}21]. Masturus [CDC^{+}23]. Match [FSN^{+}20]. Material [OJ24, Det24]. Maternal [KWD21, MTB21, CM22, GFDPASR20, LCP^{+}23, SDE^{+}22]. mating [PGW23, WLC22]. matter [BVD^{+}20, HvdSR^{+}21, JCGS22, MSL^{+}21]. matters [TdLHL21]. maturation [ARH^{+}20, DV20]. mature [SBRQ21]. maturity [ADBF21]. Maui [SON^{+}20]. maura [GVOCSGG21]. Mauritanian [DHF24]. Maurolicus [TNMN24]. Maximum [KI24]. may [AVFD^{+}21, BS20b, GFK^{+}23, MVH^{+}21, SWP20, WH24, MMS^{+}22]. Mayen [WIM23]. meadow [BB24, BBM^{+}24, MA20, MBB^{+}23]. meadows [AKW^{+}22, CAEG21, CBJ^{+}23, FHAP20, HO23, RH20, SYMR21, VPH^{+}21]. mean [CCJ^{+}23]. measure [CMvKdR23, GBH^{+}20, LOS^{+}22]. measured [Ano23, BBM^{+}24, HBE22, LNBdB21]. Measurement [Lei20]. measurements [BB24, BSH^{+}21b, BMPD20, EVVMQ^{+}23, MCvdM^{+}24, WD22]. measures [Cha22, KPK^{+}24, RBB^{+}21]. Measuring [ZCF20]. Mechanism [WCS^{+}21, BNR20, JPJ^{+}22, SSM20]. Mechanisms [BE21, DB21, GVS23, GLM^{+}24]. media [NJHN24]. mediate [WSD^{+}20]. mediated [Ano24o, CAV^{+}20, Det24, NCZ^{+}21, RWM^{+}22, SJQ^{+}22, THL^{+}20, WAA^{+}23, WYO^{+}22, WH24]. mediates [KLA21, VAMF22]. Mediterranean [MAOR^{+}21, ASL^{+}22, BML^{+}22, BQM^{+}22, CCR^{+}20, CCS^{+}22, DJR^{+}20, DBHCO^{+}22, DA24, EAS20, FVC^{+}20, HTT^{+}21, LBMSS23, MSDY23, MBG^{+}23, OVÁS^{+}21, OÁBP^{+}21, OCGAP^{+}22, PTLK21, PCLQ^{+}20, SCB^{+}20, TTM^{+}22, TTM^{+}24, VNM^{+}21, VNG^{+}22, VIV^{+}20, vLBGGT22]. Mediterranean-Atlantic [OVÁS^{+}21]. medium [COM^{+}24, SSP^{+}22]. medium-ranging [COM^{+}24]. medium-sized [SSP^{+}22]. medusa [DA24]. Megadyptes [MCFB20, MCC^{+}21]. Megafauna [VST24]. Megaptera [BKR^{+}21, PLLT^{+}21, RFJC^{+}24]. meiobenthic [ACG22]. meiofauna [GB20]. meiofaunal [CIDM23]. meiofaunal- [CIDM23]. melanostictus [NTN^{+}24]. melas [BBPO21, SBS^{+}24]. meleagris [FSS23]. meltwater [MKS^{+}20]. members [SSG^{+}22]. membranacea [AMD22]. Membranipora [AMD22]. Memoriam [LGM21]. Menor [GCGCAS^{+}20]. MEPS [Ano22a, Ano22b, Ano22c, Ano22e, Ano21a]. mercury [ABD^{+}21, BBR24, SLB^{+}21, RJ21]. merguiensis [vdVVC^{+}21]. meridionalization [SCB^{+}20]. Merluccius [GFDPASR20]. Merrill [Ano21g]. mesh [ASS^{+}23, BSP^{+}21]. meso [MPR^{+}23]. meso- [MPR^{+}23]. mesocosm [KBP22]. Mesograzer [FLCQ21]. Mesopelagic [DHF24, KCRA22, ALL^{+}23, BMOW21, DSHF22, DFSH23, MLAS22, TNMN24]. Mesophotic [LS21, EVH^{+}21, HDB21, KETL22, LR22, LFK^{+}21, LMHS21, MPA^{+}22, MMA^{+}22, PAF^{+}22, SLG^{+}24]. Mesoscale [RWM^{+}22, GDG^{+}22, SRR^{+}20]. Mesozooplankton [LHD20, CSH^{+}21a, DFRS^{+}24, GBSS24, KSS^{+}20, SSG^{+}22]. meta [AZ24, CG22, EB20, MDY^{+}21, ZPVN23, dSLD24]. meta-analysis [AZ24, CG22, EB20, ZPVN23, dSLD24]. meta-populations [MDY^{+}21]. metabarcoding [CVG^{+}23, CGS^{+}23, FWA^{+}23, RRS^{+}23, SJS^{+}23]. Metabolic [BQC^{+}22, ASS^{+}21a, ASY^{+}24, LRM^{+}21, TGB^{+}20]. Metabolism [ARB^{+}20, CAEG21, PAT^{+}24]. metacommunities [FBM^{+}24, STM^{+}24]. Metacommunity [WNH^{+}20, STY^{+}20, SFO^{+}21]. metal [WSU21]. metals [BSA^{+}24]. metapopulation [Det24]. metatranscriptomics [SCW^{+}24]. metazooplankton [PK24]. method [DGM^{+}23b]. methods [BPS^{+}24b, CSB24, DPDJ^{+}21, ECC^{+}21, HO23]. methylation [ICIS23]. metric [ACG22]. metrics [BBM^{+}24, CSB24, dACMRZ^{+}20]. metropolis [HGR^{+}22]. Mexican [LCP^{+}20, RFJC^{+}24]. Mexico [BGMAM^{+}21, PTCD22, AHMVM23, ASH^{+}20, BMO^{+}24, ESG20, FHAP20, FPZJAO^{+}22, GSS22, HAB^{+}22, HPH20, LJ20, MDS^{+}23, REGP22, SRS^{+}20, VOP21]. Micro [SWC^{+}23b, BICQG23]. micro-habitat [BICQG23]. Micro-scale [SWC^{+}23b]. microalgae [KSW^{+}21, YHW^{+}23]. microalgal [SNP20]. microbenthic [XYSF20]. Microbial [BAHG23, Hew21, CLV^{+}20, CKM^{+}20, CEAO^{+}22, DPCL^{+}23, HvdSR^{+}21, MMD20, RDT^{+}23, SOSE22]. Microbiome [LRC20, RH23, FWA^{+}23, HMKK^{+}20]. microbiomes [MMLPP^{+}24]. Microbiota [CM22, CLR20]. microchemistry [CWD24, RWR24, TTJS20, XWH^{+}22]. microeukaryotic [GLM^{+}24]. Microhabitat [VK21, RH21, THL^{+}20, THD^{+}22]. microhabitats [BLB23, FSSL^{+}20]. microlepis [GPP^{+}24, LBMB^{+}20]. Micromesistius [MAOR^{+}21, PJA^{+}21]. micropathogens [MTRM21]. microplastic [MAK^{+}22]. Microporella [DL21]. micropredation [ASS^{+}21b]. microsatellites [SHM^{+}20a]. microstructure [JHK^{+}22]. microtidal [COWM^{+}22a, HPH20]. microtopography [SBT24]. microvolume [HKY^{+}20]. mid [LOF^{+}24, LLF^{+}20, NCS^{+}23, PBS22, PDH^{+}24, VGA20, MHÁC^{+}22]. mid-Atlantic [PDH^{+}24, LLF^{+}20, NCS^{+}23, VGA20, MHÁC^{+}22]. mid-oceanic [PBS22]. mid-western [LOF^{+}24]. migrant [GVAZ^{+}22, JSDH20, MMMV22, NSB^{+}23]. migrate [COGGH20]. migrating [ALL^{+}23, ATLT20, BLGM^{+}24, MMT^{+}24, OTD^{+}24]. Migration [GNC^{+}23, JPJH23, ABC^{+}24, Bai22, CDS20, CZS^{+}22, DWJL^{+}21, EBL^{+}21, EGK^{+}21, FJB^{+}23, GBA^{+}22, HNN^{+}24, HPFB20, HMK^{+}22, IMM^{+}20, JW22, LPJ^{+}21, LPO^{+}21, LRM^{+}21, OLA^{+}24, RCC^{+}22, SVB^{+}24, SCM21, SLBK^{+}20, SGB^{+}21, UIS^{+}24]. migrations [CCS^{+}22, FSH^{+}21, MSV21, OvBS^{+}24]. Migratory [CBWH24, LCP^{+}23, AMT^{+}21, BBL^{+}21a, CBTL20, CBB^{+}22b, CGS^{+}23, DWJL^{+}21, HO24, MDY^{+}21, OvBS^{+}24, PHE^{+}20, RB23, WPM22, WLA^{+}21]. mimic [VHG20]. mimicking [BSF24]. mineralization [FCSG21]. miniata [CLL^{+}22]. minimalistic [PHV20]. minimize [WCP^{+}21a]. mining [BSA^{+}24, WIY^{+}23]. Minke [BMS^{+}21]. minnows [EHHF^{+}20]. Mirounga [CGGdD24]. mirror [BQC^{+}22]. mismatch [FSN^{+}20, GDAC23, WPM22]. missing [HW20]. Mississippi [KPK^{+}24]. misuse [AUS^{+}23]. mitchilli [AAHW20]. mitigates [BWC^{+}22]. mitigation [MDW^{+}22]. Mitochondrial [SVT^{+}24, YWW^{+}24, SHM^{+}20a]. mixed [CTH^{+}20, HKK^{+}23]. mixed-effect [CTH^{+}20]. mixing [DBHCO^{+}22, PSÓ^{+}24, WYO^{+}20]. mixotrophic [DPL^{+}21, LNZL21, PLG^{+}23]. mobile [BLB23, EB21, FSSL^{+}20, MW20, SAMdL^{+}20, SRR^{+}20]. Mobula [ACC^{+}20, FPZJAO^{+}22, HGK^{+}22, LGGR^{+}22, VvDRM20]. Model [KPK^{+}24, AU21, BLVL24, DDR^{+}21, GSL20, HSU^{+}21, HTT^{+}21, HHS21, LBW21, LCFJ20, LRD^{+}20, LORB^{+}24, MLP^{+}24, PHV20, PRC^{+}21, RSS20, RL22, RYF^{+}24, SGW^{+}21]. model-data [MLP^{+}24]. modeled [RJE^{+}22]. Modeling [Det24, LTM20, MFC^{+}24, PLLT^{+}21, ZWXC21, BBEF^{+}20, GRB22, HKD^{+}20, OAM^{+}24, PLH^{+}21, RPB21, RHN^{+}24, SWP^{+}24, WMKV21]. Modelled [GWC^{+}21, AFP^{+}24, PSP^{+}20, vLBGGT22]. Modelling [BMPD20, BNR20, Ano23, BMF24, BGMRGM^{+}20, BZT^{+}23, CLV^{+}20, CHDG^{+}23, DIC^{+}22, GFT^{+}21, LBMSS23, NDA^{+}22, RCB24b]. models [BLE^{+}22, BRD20, CAP^{+}21, CTH^{+}20, CMF21, DCR^{+}21, FTC^{+}20, GRTK^{+}21, GKK^{+}22, HvLM21, HBG^{+}22, MTRM21, ORB^{+}21, RCC^{+}24, SJA^{+}21, SAGG^{+}24, WYO^{+}20]. modest [CHJ21]. modification [DNJCH21, MPKS24, RS20a]. modified [MJ23]. modifies [SOSE22]. modify [DFCM21]. Modiolus [SHG^{+}21]. modulate [LT21, SF23, WCRTT20]. modulated [SSBA^{+}20]. modulates [KCLJ20]. mola [CDC^{+}23, NBW^{+}23]. Molecular [VFG24, JPJ^{+}22, VNG^{+}22]. Molidae [ABWJ^{+}20, NBW^{+}23]. molt [DRP^{+}20]. molted [LNBdB21]. molting [JW22]. momus [BSP^{+}21]. Monitoring [MMMV22, BD20, CDD^{+}20a, CCT^{+}23, MSH20, MBR^{+}21, PBAC23, RYF^{+}24, RDL^{+}22, SBFP^{+}24, TGT^{+}23, YOI^{+}23]. monoceros [ZMD^{+}21]. Monodon [ZMD^{+}21]. Monodonta [YMU^{+}21]. monsoon [SST^{+}20a]. Montenegro [CCA^{+}20]. Monterey [SRHM21]. Montipora [HBR^{+}21, RHM^{+}22]. Moorea [BPSB^{+}22, CFE^{+}22, VLPN24]. mordax [DHM21]. morhua [BLW^{+}24, BSM23, CDD^{+}20a, GBG^{+}23, JNJ^{+}20, KSMP20, LCGS^{+}21, MVC^{+}23, SBS20, SJH^{+}22]. morio [MPKS24]. Moroteuthopsis [vTLG^{+}21]. morphological [BQC^{+}22, NP22, YWW^{+}24]. morphology [Ano23, BPS^{+}21, CCC^{+}23, LMHS21]. morphometric [BSH^{+}21b]. morphometrics [CAP^{+}21, LPJ^{+}20]. morphs [CRG^{+}20]. mortality [CRGMA^{+}21, DAB^{+}21, HAB^{+}22, Hew21, JPMG23, JPL^{+}24, KCGR20, LPJ^{+}21, LFA^{+}23, MBW^{+}20, MHAL24, MJ21, PBS^{+}24, SSC^{+}23, SMK23, SLBK^{+}20, SCP^{+}23, SSM20, WCP^{+}21b, ZLL^{+}24, dMGGG21]. Morus [RMASA20, dDW^{+}22]. Mosquitofish [HBC^{+}20]. moss [FLCQ21]. most [ABB^{+}20, BCC20]. mostly [SRG^{+}23]. mother [ASS^{+}21b]. mother-calf [ASS^{+}21b]. motion [KPS^{+}22, ZZYB21]. mouth [WKR21]. Movement [DHM^{+}24, ESG20, APOGG20, Ano23, BBL^{+}21a, BASM^{+}24, BOY^{+}23, BLBBL20, CILGD^{+}22, CDP^{+}21, EMDR22, FPB^{+}21, FLBP^{+}21, FNO^{+}23, FKTK20, GBF^{+}23, HHM^{+}22, HP23, HMW^{+}21, HJG21, KKS^{+}20, KTR^{+}20, KCB^{+}24, LCR^{+}22, LNKJ^{+}22, MGB^{+}22, PSB^{+}23, SON^{+}20, SvBS^{+}20, TCP^{+}23, VOP21, VvDRM20]. Movements [CGW21, AMT^{+}21, ACC^{+}20, ASPM21, BKA^{+}20, CBWH24, CCA^{+}21, CWD^{+}22, CKKL21, DVR^{+}23, DMM^{+}24a, FDL^{+}21, GVAZ^{+}22, GMS^{+}21, HNB^{+}24, HDA^{+}22, LPJ^{+}21, MSC^{+}20, OH22, PBT^{+}21, SDF^{+}21, SDE^{+}21, TCM^{+}23, VDB^{+}24, VBB^{+}21, WdGR^{+}20, ZSWR^{+}22]. moving [Ano24o, HBC^{+}20, LMRS^{+}24]. Mozambican [PAAEV^{+}24]. Mozambique [VvDRM20]. MPA [JS20, OAdJA^{+}20]. mucous [ASS^{+}23, BSP^{+}21]. mucous-mesh [ASS^{+}23, BSP^{+}21]. mud [KLFP23]. mudflats [HMV^{+}22]. muelleri [HMKK^{+}20, TNMN24]. mullet [PCLQ^{+}20]. mulloway [NHFS22]. Mullus [PCLQ^{+}20]. Multi [BLWJ20, CBB^{+}22b, CH20, EMB^{+}22, HHO^{+}23, IMYW24, LMB^{+}20, TWC^{+}23, BBG^{+}22, CCAdA24, CGM^{+}24, CBTL20, FPG^{+}21, GSN^{+}20, HSNT24, RDL^{+}22, TWC22]. Multi-colony [TWC^{+}23, BBG^{+}22]. Multi-cusped [IMYW24]. Multi-decadal [BLWJ20]. multi-hypothesis [CCAdA24]. Multi-proxy [CBB^{+}22b]. Multi-scale [CH20, CCAdA24, FPG^{+}21]. multi-sensor [CGM^{+}24]. multi-site [CBTL20]. multi-species [HSNT24]. multi-spectral [GSN^{+}20]. multi-tiered [TWC22]. Multi-year [EMB^{+}22, HHO^{+}23, LMB^{+}20, RDL^{+}22]. Multidirectional [CCS^{+}22].

Multidisciplinary [CLR^{+}21]. multievent [BLE^{+}22]. Multiple [CDC21, PDFH20, BBH^{+}20, BPS^{+}24b, Cha22, CP22b, EPHS23, FS22, FP20, PDM22, RGK22, SHB^{+}20, SW23a, SRS^{+}20, TKS^{+}20, YSD^{+}21, ZHG^{+}21]. Multiscale [KMW^{+}21]. multispecies [CDF^{+}20, OCFEB21, RCB24b, WSPS^{+}20]. muricid [CH20]. murre [SFMK^{+}23]. murres [FDE^{+}21, RMD24]. mussel [AJH22, ARB^{+}20, BMM21, DBS^{+}20, NT24, PLH^{+}21, QRC^{+}24, RC23, SHG^{+}21, STH21, SGP24]. mussels [BJH^{+}23, BKT^{+}22, HHBP23, MLVJ23, RSL^{+}22, SdFBM24]. Mustelus [PBO^{+}21]. mutualism [dFRvdG^{+}22]. Mycteroperca [GPP^{+}24, LBMB^{+}20]. myctophid [KRG^{+}20]. Myctophidae [ASS^{+}21a]. myctophids [ASS^{+}21a, COGGH20]. mydas [SBC^{+}21]. Myliobatis [DCN^{+}23]. mysticetus [PDF^{+}23]. Mytilus [BVD^{+}20, SSFS20].

N [SST^{+}20a, BCC20, FLC^{+}24b, HRH22, LNBdB21, MBW^{+}20, NYW^{+}22, SCW^{+}24, VS23, ZMD^{+}21, ZBB^{+}23]. N-tracer [SCW^{+}24]. naiveté [DJR^{+}20]. nanoflagellate [CS22]. Nansei [TKS^{+}20]. narwhal [ZMD^{+}21]. Nassau [BBH^{+}20]. nasus [WHB^{+}20]. natal [BLBBL20, CHDG^{+}23, SHM^{+}20a, UIS^{+}24]. natchiq [LHM^{+}23]. native [ASL^{+}22, Ano24o, BBMM20, CCMRII22, DJR^{+}20, HSB20, HWA^{+}22, ILON21, LK21, LCJ22, NBDM^{+}24, SGP24, SE20]. natural [BDS^{+}20, CSK^{+}23, CSWM24, DAB^{+}21, LAD^{+}21, MPKS24, RCES22, RWSR24, SLG^{+}24, VPB^{+}20]. Nature [MSDY23, SLH^{+}22, TWC20]. nature-based [SLH^{+}22]. naucrates [PBKG22]. nauplii [JP20]. Nautilus [CASF21]. near [PBS22, SSBA^{+}20, THK^{+}22, WKR21, WIM23]. nearby [PTCD22]. Nearshore [RJ21, CLR^{+}21, CSK^{+}22, GNC^{+}23, GSS22, OLG^{+}23, SBJ^{+}24]. need [ER20, PK24, TCM^{+}20]. needs [FFMR24]. Negative [SBB20, CAV^{+}20, KLN^{+}21]. negatively [RTT^{+}21, TGB^{+}20]. neglected [PTLK21]. negligible [GVR^{+}20]. Neither [EPB23]. nekton [BMOW21, FWGD^{+}23, ZCM^{+}21]. nematodes [CIDM23]. nemertean [SFL23]. Neocallichirus [SC23]. Neonatal [RHN^{+}20, BSP^{+}22]. Nephrops [VNM^{+}21, VNG^{+}22, ZMC^{+}22]. Neptune [HMT21]. Neptunea [HMT21]. nest [ODB^{+}24, SBC^{+}21, SWHE20]. nesting [CCS^{+}22, RMR^{+}23, SHM^{+}20a, SFMK^{+}23, WCL20, WCP^{+}21a]. nests [YBSH23]. net [ANP^{+}24, GCDA20]. Network [CDF^{+}20, BWC^{+}22, FLBP^{+}21, MSL^{+}24, OAdJA^{+}20, RBFA20]. networks [NSKH23]. neural [NSKH23]. neustonic [COGGH20, GDG^{+}22]. neutral [BVKF23]. newcomer [EAS20]. Newell [MCJ^{+}20]. Newfoundland [KF20, LGD23, LCGS^{+}21, MVH^{+}21]. newly [PSP^{+}20]. next [FOO20]. next-generation [FOO20]. Nggela [THD^{+}22]. Niche [AGB24, GLM^{+}24, BGMAM^{+}21, BBR24, BDC^{+}22, CFFC^{+}23, CS20, DHM^{+}24, DIC^{+}22, DCK^{+}24, DRP^{+}20, DCL21, FTC^{+}20, FNS^{+}20, GGN21, GFT^{+}21, GGCL22, GRB22, HSA^{+}21, HLCH23, ISMM20, LÁGLL20, LBR^{+}24, LNBdB21, MDK^{+}21, MRH20, PDH^{+}24, RMD24, SWC^{+}23a, TPGG^{+}23, YGMNA22]. niches [GQBB^{+}24, HPRP21, LMRS^{+}24, MHÁC^{+}22, MON^{+}24]. nicholsi [KRG^{+}20]. night [BLW^{+}24]. Nine [vdHMA^{+}24, PSP^{+}20]. Ningaloo [LMB^{+}20]. Niño [AMMADDH20, RVCT21]. nitrate [JRVE^{+}22]. nitrification [KBWRV21]. nitrogen [AS20, CEAO^{+}22, DDR^{+}21, ICH^{+}24, JRVE^{+}22, OCR^{+}24, RDT^{+}23, RSL^{+}22, SDW20, SDM^{+}22, SCW^{+}24, SOSE22, WCS^{+}21]. No [DDD^{+}23, ATR^{+}24, DVP^{+}23, LCR^{+}22, MDPL20, PSB^{+}23, VNM^{+}21]. no-take [ATR^{+}24, DVP^{+}23, LCR^{+}22, MDPL20, PSB^{+}23, VNM^{+}21]. nobilis [GCGCAS^{+}20]. Noctiluca [ZSC^{+}22, OÁBP^{+}21]. Nocturnal [AHCT22b]. nodosa [AS20]. noise [YXB22]. noisy [SFMK^{+}23]. noltei [JRVE^{+}22]. noltii [dFRvdG^{+}22]. nomadica [DA24]. nominally [CFFC^{+}23]. Non [CCA^{+}20, CFE^{+}22, GLG21, SST^{+}20a, SGP24, SDF^{+}21, BFQ^{+}23, BBG^{+}22, CWD^{+}22, CGS^{+}23, GBF^{+}23, HSB20, ILON21, RLME^{+}21, RBB^{+}21, RDL^{+}22, RMD24, SAMdL^{+}20, SCP^{+}23, SE20]. non-acute [SCP^{+}23]. Non-breeding [SDF^{+}21, BFQ^{+}23, BBG^{+}22, CGS^{+}23, GBF^{+}23, RLME^{+}21, RMD24]. Non-consumptive [GLG21]. non-focal [CWD^{+}22]. non-food [RBB^{+}21]. non-indigenous [RDL^{+}22]. non-instantaneous [SAMdL^{+}20]. Non-native [SGP24, HSB20, ILON21, SE20]. Non-random [CFE^{+}22]. Non-Redfieldian [SST^{+}20a]. Non-targeted [CCA^{+}20]. nongeniculate [CCC^{+}23]. nor [EPB23]. Nordic [RBM^{+}21]. normalization [LH22]. North [APOGG20, FPB^{+}21, WIM23, FNS^{+}20, GLM^{+}24, OCC23, AKW^{+}22, AHH^{+}20, ARE^{+}24, AMT^{+}21, ABWJ^{+}20, BSM23, BBBT^{+}20, BGDJ20, CDD^{+}20b, CFBPH23, CMvKdR23, DL20, FOO20, FNL^{+}23, GFH^{+}22, GON^{+}20, HEP^{+}24, HBJ^{+}21, HCvdHM23, JFB^{+}23, KHL^{+}22, KTK21, LSKS24, MPR^{+}23, MAPFH22, MAH^{+}20, MDY^{+}21, RSvL21, RWH^{+}24, RBM^{+}21, RYF^{+}24, RRR^{+}23, RSvV^{+}23, STO^{+}23, SSB^{+}23, SGW^{+}21, SDE^{+}22, SDF^{+}21, SDE^{+}21, UIS^{+}24, VDB^{+}24, VIV^{+}20, vESB^{+}23, vdVTW^{+}22, vdVWF^{+}24]. north-east [FNS^{+}20, AHH^{+}20]. northeast [Hew21, JHDE20, KMW^{+}21, SAMP24, SAF^{+}24, TAN^{+}22, ZWXC21, BBL^{+}21b, FTA^{+}21, MBB^{+}23, MDQ^{+}21, WF21]. northeastern [FKTK20, GFDN^{+}24, KNO^{+}21, LNKJ^{+}22, SdGWN21, TQG24, VOP21, WBG^{+}22]. northern [BCC20, BMF24, DdGA^{+}21, DFHT23, DFSH23, DAT^{+}21, ESG20, FBC^{+}24b, GG23, GMB^{+}24, HMT21, HGT^{+}21, HAB^{+}22, HPH20, KCGR20, LJ20, LPJ^{+}21, MGHH21, MTB21, MMA^{+}22, NYW^{+}22, NBL^{+}22, PBS22, SBDM^{+}23, UAL24, WRW20, WIM23, ZZX^{+}20, dDW^{+}22, KDB^{+}24, SSB^{+}20]. Northward [JW22, vdKMC^{+}24, LMRS^{+}24, OLM24]. northward-moving [LMRS^{+}24]. Northwest [PGGE21, PDM22, VPH^{+}21, ABN^{+}21, DSA^{+}24, DHF24, RCB24b, RMD24]. Northwestern [MMB22, FVC^{+}20, FSH^{+}21, FKS^{+}22, MKS^{+}20, SAC24]. norvegicus [VNM^{+}21, VNG^{+}22, ZMC^{+}22]. Norway [BLW^{+}24, MLVJ23, VGLL^{+}21, VNM^{+}21]. Norwegian [FSN^{+}20, FLY^{+}24, HSU^{+}21, HSSN^{+}21, PSÓ^{+}24, VBB^{+}21]. NOTE [WBS24]. novaeangliae [BKR^{+}21, PLLT^{+}21, RFJC^{+}24]. Novaya [EGK^{+}21]. Novel [OMW^{+}20, ANP^{+}24, AFTW24, BBM^{+}24, GB21, HSB20, NSB^{+}23]. NTZ [LCR^{+}22]. Nucella [CH20]. nucleotide [PSÓ^{+}24]. nudibranchs [CCBT^{+}20]. Nui [SON^{+}20]. numbers [MFC^{+}24]. nurseries [BTMS24, CWD24, GPP^{+}24, vdVWF^{+}24]. nursery [AMRL^{+}24, AAFC^{+}21, BSP^{+}22, CLR^{+}21, CMO^{+}20, CLF^{+}20, EMB^{+}22, FHAP20, LBMF^{+}24, OVO^{+}21, RWR24, SOS^{+}21a, TTJS20, VDC^{+}24, Whi20, Whi21, vdVTW^{+}22]. nursing [ZMD^{+}21]. nurture [TWC20]. Nutrient [CBGB24, BDJ^{+}24, DNE^{+}24, HPH20, KBP22, LOS^{+}22, MCJ^{+}20, NYW^{+}22, PBC^{+}22, RHH^{+}20, SWP^{+}24, TYL^{+}24, WSD^{+}20, dFRvdG^{+}22]. nutrients [LAL21, RTT^{+}21]. nutrition [ICH^{+}24]. nutritional [BBRM22, CKCS20, EBL^{+}21, LNZL21, LHHH^{+}20, OYO^{+}20, PLG^{+}23, RHN^{+}20, RHSM21]. NW [LAGGM^{+}21, SMS^{+}24, GGRV23, MAOR^{+}21].

O [LHW^{+}20]. objects [NGPGO^{+}23]. observation [NJHN24]. observational [EB20]. observations [CGM^{+}24, DeC24, HDM^{+}24, LKR^{+}21, SJA^{+}21, SAGG^{+}24]. observed [Sig23]. occasionally [Ano24o]. occupancy [GBA^{+}22]. occupy [PBKG22]. occurrence [BBPO21, CGM^{+}24, PBS22]. occurrences [SMB^{+}20]. occurring [BMKB22, MHN24, MGBCGM^{+}22, RH23, SCM^{+}23, WD22, WF21]. Ocean [ASS^{+}21a, ASBE22, CG22, DHA^{+}20, HHBP23, PDM22, SHH^{+}22, SOSE22, Ano24o, ABWJ^{+}20, ASPM21, BK23, BG21, BBM^{+}24, BMB24, CEZH21, DVMH^{+}24, FLC^{+}24b, GBE^{+}22, HBR^{+}21, HGH^{+}24, HMB^{+}22, HPFB20, JPL^{+}24, KLN^{+}21, KTK21, LK21, LAD^{+}21, LSS^{+}22, MR20, NMB^{+}22, OVÁS^{+}21, PGGE21, PSN^{+}20, PYB^{+}21, RKV^{+}21, SBSS21, SJ22, SJSR24, THK^{+}22, TABM21, TMPPG24, TCY^{+}22, TFH^{+}22, UPK^{+}20, VHG20, VIV^{+}20, WHB^{+}20, WCRTT20, WAA24, YDG21, APOGG20, ABN^{+}21, ASBE22, AAFC^{+}21, BBL^{+}21b, BBBT^{+}20, BPS^{+}24b, CC22, CBB^{+}22b, CBKM20, DMTP22, ECvBL21, FLC^{+}24a, FLC^{+}24b, GFT^{+}21, GSL^{+}24b, GFH^{+}22, GDCP20, GDG^{+}22, Hew21, HEKH24, KRG^{+}20, LHD20, LÁGLL20, LRM^{+}21, MPR^{+}23, MFC^{+}24, NCS^{+}23, OTD^{+}24, PBS22, PS23, PDM22, QHT^{+}21, RBM^{+}21, RBF^{+}22, SKTO23, SAC24, SAF^{+}24, SCM^{+}23, UIS^{+}24, VDB^{+}24, WWH^{+}22, YGMNA22, ZWXC21]. Oceanic [FJB^{+}23, BGMRGM^{+}20, FPZJAO^{+}22, GVAZ^{+}22, HGK^{+}22, MSL^{+}24, OvBS^{+}24, PBS22, RBF^{+}20]. oceanica [AWB23, MDH^{+}20]. Oceanodroma [HBE22]. Oceanographic [ASS^{+}23, LBH^{+}21c, AMUPJR^{+}24, BBEF^{+}20, DPP^{+}21, GRTK^{+}21, GWA^{+}23, LBH^{+}21a, LOF^{+}24, LRD^{+}20, MPB^{+}24, NBL^{+}22, RB22, TMN^{+}24]. oceanography [BBEF^{+}20, GDG^{+}22]. oceans [MMP^{+}22]. ocellatus [BLBBL20]. octocoral [RH23]. Octopus [LÁGLL20, RRS^{+}23]. Oculina [VHG20]. odd [LAD^{+}21]. odd-year [LAD^{+}21]. Odontella [LBB^{+}21]. odontocete [PSP^{+}20]. odontocetes [PBS22]. off [ASS^{+}23, ASS^{+}21b, BBPO21, CDC^{+}23, DHF24, FPG^{+}21, FPB^{+}21, FSK^{+}21, KNO^{+}21, MDK^{+}21, PRM^{+}21, SBDM^{+}23, TKS^{+}20, TSS^{+}23, WSPS^{+}20, WLB^{+}20]. offs [BQM^{+}22, CPH20, HSO^{+}22]. Offshore [AWC21, PTCD22, WBG^{+}22, FDS^{+}21, FHAP20, JTBS^{+}22, LCGS^{+}21, MTRM21, WBGC24]. offspring [DL21, ŠGH^{+}21]. oil [ESG20, LBB^{+}21, MTB^{+}23, MAPFH22, ZHG^{+}21]. old [CEAO^{+}22]. olfactory [KKC^{+}24]. olive [ARE^{+}24]. Olympia [HGK^{+}24]. omega [HBE22]. omega-3 [HBE22]. Ommastrephidae [PS23]. omnivore [PTLK21]. Oncorhynchus [WDJ20]. One [GRB22, BAS21]. ongoing [SEJ^{+}24]. Only [PWvdV20, CHJ21, DVP^{+}23]. Ontogenetic [BKH^{+}24, CDC^{+}23, DFP^{+}21a, EMGMEV^{+}21, HSA^{+}21, HCS^{+}24, ZMD^{+}21, HMW^{+}21, MSV21, NSMJ20, PSN^{+}20, PAT^{+}24, RGK22, SdGWN21, VK21]. ontogenic [ASGI24]. ontogeny [BSF^{+}20, BDS^{+}20, DMM^{+}24b, GGCL22]. open [HBC^{+}20, MKR^{+}20, OBS^{+}22, THK^{+}22, TCY^{+}22, WFS23]. open-coast [OBS^{+}22]. opening [CCL^{+}22]. operated [SAMdL^{+}20]. operation [WBGC24]. Ophiodon [LBH^{+}21a]. ophiuroids [YGMNA22]. Opinion [ER20, Hew21, LS21, OA24, PAF^{+}22, SJA^{+}21, Tha21, TCM^{+}20]. Opportunistic [DFHT23, OTD^{+}24, vdRJL20]. opportunities [DPP^{+}21, TOMDC23]. Opsanus [MPM21]. optima [ZCM^{+}21]. optimisation [CLV^{+}20]. optimize [HGK^{+}24]. Optimizing [RHM^{+}22, WCP^{+}21a]. orbicellid [EVH^{+}21]. orca [WYO^{+}20]. Orcinus [WYO^{+}20]. Order [WKAH22]. organic [ARB^{+}20, BVD^{+}20, HvdSR^{+}21, KBWRV21, LRD^{+}20, MSL^{+}21, SST^{+}20a, SFF^{+}20, SNE^{+}24, UAL24]. organismal [THL^{+}20]. organs [MHF^{+}21]. orientalis [AOO^{+}21, FSK^{+}21, HKY^{+}20]. orientation [CFE^{+}22, MCH^{+}21]. oriented [CHDG^{+}23]. Origin [DHJ^{+}22, CSK^{+}23, EAS20, LCP^{+}23]. origins [UIS^{+}24]. Oscillation [STO^{+}23]. Oslofjorden [MSTH22]. OSMOSE [LORB^{+}24]. Ostrea [HGK^{+}24, LK21]. otariids [CAAG^{+}22]. other [BKR^{+}21, BGDJ20, SMPLF^{+}22]. Otolith [ASS^{+}21a, BTMS24, CWD24, DMTP22, LDW^{+}21, RWR24, XWH^{+}22, AAFC^{+}21, CAP^{+}21, CMdL^{+}20, GPP^{+}24, HKY^{+}20, JHK^{+}22, SRS^{+}20, TTJS20, ZXL^{+}24]. Otolith-based [LDW^{+}21]. Otolith-derived [ASS^{+}21a]. otoliths [KHL^{+}22, KRG^{+}20, KSMP20]. otter [LKR^{+}21, RSEG21]. otters [GTH^{+}20]. out-migrating [ATLT20, BLGM^{+}24]. outbreaks [AMD22, DDB^{+}24]. outer [EVAB^{+}23]. output [HB24, RVCT21, RHN^{+}20]. overcome [MAS^{+}20]. Overcoming [NBW^{+}23]. overfishing [AKU^{+}22]. overgrazing [PTLK21]. overgrowth [OCFEB21]. overlap [BGDJ20, DMBJ21, FLY^{+}24, FDS^{+}21, GQBB^{+}24, LMS^{+}20, LFFW21, MdOO^{+}20, MDW^{+}22, PRCvdK23, PRM^{+}21, RMR^{+}23, RMD24, TPGG^{+}23, WSB^{+}24]. overwinter [RBS^{+}24, WRW20]. overwintering [CLC^{+}23, MAH^{+}20, SRK^{+}23]. oxidative [CCBD^{+}21]. oxidiser [KBWRV21]. Oxygen [KHL^{+}22, BP24, DHF24, GD23, GBW^{+}20, HKY^{+}20, PBC^{+}22, WLB^{+}20]. oxyrinchus [VDB^{+}24]. Oyster [BMO^{+}24, LSSL21, SWP20, THR^{+}23, AR21a, CHDG^{+}23, GSP^{+}20, HO21, HMB^{+}22, HMV^{+}22, KPK^{+}24, LK21, LBR^{+}24, MPM21, MVSV22, OAM^{+}24, PRD^{+}20, PBC^{+}22, RF20, RAHM^{+}21, RS20a, SHZ^{+}20, SHT^{+}22, ZCBC24]. oysters [ATR^{+}24, CLV^{+}20, GCDP21, GLG21, HGK^{+}24].

P [LDM^{+}20, NYW^{+}22, SST^{+}20a, VS23]. Pacific [APOGG20, BBL^{+}21b, DL20, GFH^{+}22, GDG^{+}22, UIS^{+}24, VDB^{+}24, AHCT22a, AHCT22b, AAI^{+}24, ARE^{+}24, ABWJ^{+}20, AOO^{+}21, BPS^{+}21, BLE^{+}22, BGMAM^{+}21, BLA^{+}24, CTTLCG^{+}24, CLD23, CGLC20, CRLC21, CDS20, CIDM23, CHDG^{+}23, CFSH22, DMTP22, EMDR22, EMGMEV^{+}21, FN24, FBM^{+}24, FRO^{+}21, FNH^{+}23, FSH^{+}21, FKS^{+}22, FSK^{+}21, GSP^{+}20, HKY^{+}20, HGH^{+}24, HBE22, Hew21, HIT^{+}22, JHDE20, JETGM24, JFB^{+}23, KNO^{+}21, KHL^{+}22, LH22, LGGR^{+}22, MLE^{+}22, MHH^{+}20, MGA^{+}20, MMD20, dACMRZ^{+}20, MDH^{+}21, MBB^{+}23, MGM^{+}24, OAM^{+}24, PLC^{+}20, RVCT21, RCB^{+}24a, RBS^{+}24, RBC^{+}24, RSvV^{+}23, SKTO23, SCM21, SAMP24, STO^{+}23, SWC^{+}23a, SAF^{+}24, SSB^{+}23, SSG^{+}22, TAN^{+}22, UIS^{+}24, WPG^{+}24, ZWXC21]. pacificus [WGK22]. pack [ORB^{+}21]. pack-ice [ORB^{+}21]. packaging [WYO^{+}22]. pairs [ASS^{+}21b]. Pais [Ano21e]. Palatability [dSLD24, VOM^{+}20]. Palau [CASF21]. pallasii [FN24, HGH^{+}24, MGM^{+}24, RBC^{+}24]. pallida [LPB^{+}23]. palmata [LvAH^{+}24]. Palythoa [WKR21]. Panama [CLR20]. Pandalus [BMF24]. pandemic [LGR^{+}22]. Panulirus [ADBF21, BBRM22]. Papasula [CBWH24]. papua [DPL^{+}21]. Paracentrotus [CCRP20]. paradigm [Whi20]. paradisaea [RB22]. paralarvae [ASIA22, GMRMG22]. paralichthyid [HTGG22]. Paralichthys [GVS23]. parameter [CLV^{+}20]. parameters [Ano23, PHE^{+}20, RPB21, SLO^{+}22]. Paramuricea [SLG^{+}24]. Parapenaeus [ZFD^{+}24]. parasite [BMKB22, DMBB22, OSBA22]. Parasites [JMM24, BBMM20, JSDH20]. parasitic [LCJ22]. parental [dDW^{+}22]. Park [LMB^{+}20, BMP^{+}22]. parma [UBO^{+}24]. parmifera [SOS^{+}21b]. Parrotfish [TB21, RBB^{+}24]. parrotfishes [MdMR24, VK21]. part [RDL^{+}22]. partial [ABC^{+}24, JPMG23]. particle [AM20, GKK^{+}22, KBWRV21, SVS^{+}22]. particle-tracking [GKK^{+}22]. particles [MAK^{+}22, WSU21]. particular [MPF^{+}22]. particulate [BVD^{+}20, SFF^{+}20]. Partitioning [Fai21, AGB24, BLGM^{+}24, BBR24, CRG^{+}20, DCL21, FJJ^{+}21, GGN21, GGCL22, HLCH23, ISMM20, MDK^{+}21, PBRF23, PDH^{+}24, SWC^{+}23a, SHS21, SD22, WSPS^{+}20]. Passive [NJHN24, CDD^{+}20a, CKKL21, DMM^{+}24a, RYF^{+}24, SHS21, SBFP^{+}24, VH20]. past [JBWS24, WFR^{+}23, vdVTW^{+}22]. Patagonia [ARdlB^{+}22, BCC20]. Patagonian [CGM23, LSR^{+}24, VLG^{+}21, dlBSR^{+}20]. Patch [MMP^{+}22, BGGH^{+}22, SF23, STY^{+}20]. pathogen [KCLJ20]. pathways [ARB^{+}20, BCC20, GAPM24, JDDF21, MAK^{+}22, WHRC22]. Patiria [CLL^{+}22]. pattern [DFB^{+}24, FNH^{+}23, HHO^{+}23, UIS^{+}24]. Patterns [HSU^{+}21, KTR^{+}20, TQG24, AHMVM23, AKW^{+}22, AFTW24, AZ24, AFP^{+}24, Ano23, BMF24, BASM^{+}24, CCAdA24, CBWH24, CSB24, COGGH20, DFSH23, ECvBL21, ESG20, FVC^{+}20, Fig21, FWGD^{+}23, FPB^{+}21, GCDP21, GG23, GIDP^{+}21, HGR^{+}22, HHS21, HPFB20, HCS^{+}24, JPB21, KPW^{+}20, LT21, LSR^{+}24, LMB^{+}20, LFFW21, MCF^{+}21, MLB^{+}21, MTK^{+}21, MGB^{+}22, NRE21, PSN^{+}20, RVSA^{+}21, RF20, RBW^{+}20, RBFA20, RWM^{+}22, SMK23, SS24, SCM^{+}23, SON^{+}20, STY^{+}20, THD^{+}22, TABM21, VPB^{+}20, VH20, VvDRM20, VFG24, WXZ^{+}24, WSB^{+}24, ZFD^{+}24, vdWDH^{+}21]. Paul [DCR^{+}22]. pCO [EHHF^{+}20]. pea [SF23]. peak [DDB^{+}24, FLC^{+}24a, LT21, PDF^{+}23]. Pearsonothuria [HP23]. peat [CEAO^{+}22]. peatlands [MSL^{+}21]. Pelagia [OÁBP^{+}21]. Pelagic [Zue22, AMT^{+}21, BSÁG^{+}22, CWD^{+}22, DPP^{+}21, FDF20, FSAG22, FNH^{+}23, FD21, GGCL22, GAPM24, GSRC20, JDDF21, LR22, LBTE^{+}21, MCvdM^{+}24, MAK^{+}22, MTB^{+}23, MIL^{+}20, MLAS22, OCGAP^{+}22, PRCvdK23, RF20, RSS20, RCC^{+}24, SNP20, SB20b, TNS^{+}22, VFG24]. Pelagic-benthic [Zue22]. pelagics [JMM24]. pen [DNJCH21, GVOCSGG21, GVCSO^{+}24]. Penaeus [vdVVC^{+}21]. penetrability [GBH^{+}20]. penguin [CKT^{+}24, HOK^{+}22, MCFB20]. penguins [BFQ^{+}23, CCBD^{+}21, DRP^{+}20, JCGS22, LLJ^{+}20, MAS^{+}20, MTCT^{+}23, MKR^{+}20, MCC^{+}21, RBW^{+}20, YSD^{+}21, YSP^{+}22]. penguins/hoiho [YSP^{+}22]. Peninsula [BGMAM^{+}21, DFP^{+}21b, DME^{+}23, FJJ^{+}21, GSL^{+}24a, GMRMG22, IAG^{+}23, WRW20, ASS^{+}21b, MBW^{+}20]. Pennatulacea [DNJCH21]. Pennellidae [LLF^{+}20]. perceived [HWA^{+}22]. percent [SWHE20]. percoides [KWD21]. Performance [HBG^{+}22, OHBJ20, ASY^{+}24, AAP^{+}23, BSP^{+}22, CBGB24, CLGH24, GM20, GB21, HSBB^{+}21, JCGS22, JH20, KPK^{+}24, NT24, RH21, RS20a, WDP^{+}21, dSGC^{+}23]. perilous [DML^{+}21]. Perils [Str23]. period [DCG^{+}20, DRP^{+}20, PGGE21]. Periodicity [WRS^{+}23]. periods [FYM^{+}24, SCP^{+}23]. periwinkle [GBA21, Str23]. permeability [WSD^{+}20]. Perna [BMM21, DBS^{+}20, STH21]. Persistence [SB20a, AMD22, PCLQ^{+}20]. Persistent [DVR^{+}23, RDL^{+}22, RFJC^{+}24, SWP20]. personatus [BSH21a]. perspective [LOF^{+}24, SWJ^{+}21]. perspectives [MGB^{+}22]. perturbed [SFO^{+}21]. Peru [ASS^{+}23, BASM^{+}24, SBDM^{+}23]. Peruvian [CAAG^{+}22, CILGD^{+}22]. Peterson [LGM21]. Petrarca [KZC23]. petrel [CVG^{+}23, FDS^{+}21, HBE22]. petrels [BGMAM^{+}21, BDC^{+}22, DCK^{+}24]. Petrolo [Ano21f]. pH [CCJ^{+}23, GPCM22, HTL^{+}20, JBG22, YAZA21]. Phaeophyceae [HOK^{+}20, LPB^{+}23]. phalarope [SVB^{+}24]. phase [MG22]. phenolic [MJW24]. Phenological [CFSH22, HOK^{+}20, GBE^{+}22, WPM22]. phenology [DMCB22, GBA^{+}22, ILON21, JCGS22, JFB^{+}23, KKN^{+}20, LPO^{+}21, MLB^{+}21, TTA20, WMKV21]. phenotypes [KDLL21]. Phenotypic [LBR^{+}24, BPA21]. pheromone [ZLHH20]. Philippines [BRC22, CAP^{+}21]. philopatry [GNC^{+}23]. Phoca [YXB22]. phocid [IMYW24, MWCR20]. Phocoena [HJG21, SBS^{+}22]. pholis [JMP^{+}24]. Phoretic [PBKG22]. phosphate [HWT^{+}22, VS23, ZLHH20]. phosphodiesterase [HWT^{+}22]. Phosphomonoesterase [HWT^{+}22]. phosphorus [RWSR24, WCS^{+}21]. photo [MTK^{+}21]. photo-identification [MTK^{+}21]. photobiology [LHHH^{+}20]. photogrammetric [BSH^{+}21b]. photogrammetry [CBB^{+}22a, MDH^{+}20, OCFEB21]. photographic [RDL^{+}22]. photographs [BCP^{+}22]. Photoperiodism [CLC^{+}23]. photosynthesis [FBKG24, LAGGM^{+}21]. Photosynthetic [WD22, WF21, BP24]. photosystem [KSW^{+}21]. Phycis [ASGI24]. phycosphere [YHW^{+}23]. Phylogeography [SAC24]. phylotypes [DPCL^{+}23]. Physeter [BBPO21, DFHT23, WMT20]. Physical [FPZJAO^{+}22, BPSB^{+}22, DFRS^{+}24, ENF^{+}22, LRD^{+}20, MGH21, MTRM21]. Physiological [OVÁS^{+}21, Ano23, BSP^{+}22, CS20, FOM^{+}20, GSN^{+}20, SBSS21, SBRQ21]. physiology [AR21a, BK23, NYW^{+}22, SSFS20]. Phytoplankton [BDJ^{+}24, TYL^{+}24, UAL24, AHH^{+}20, ASBE22, BWRP20, BNR20, CSH^{+}21a, DSB^{+}21, DKSS20, GSN^{+}20, KHG^{+}21, LEG^{+}20, LAGGM^{+}21, PHV20, PRD^{+}20, SDM^{+}22, SCW^{+}24, URSS24, ZWXC21]. phytoplanktonic [CCR^{+}20]. Picochlorum [PLG^{+}23]. picophytoplankton [ZLL^{+}24]. picture [BRD20]. Piece [ER20, Hew21, LS21, PAF^{+}22, SJA^{+}21, Tha21, TCM^{+}20, OA24, PAL^{+}22]. pigeon [JW22]. pigment [LHHH^{+}20]. pigments [WF21]. pileatus [JP20]. pilot [BBPO21, BMS^{+}24, SBS^{+}24, SJQ^{+}22]. pinfish [LOF^{+}24]. pink [LAD^{+}21, RSvV^{+}23]. Pinna [GCGCAS^{+}20, GVCSO^{+}24]. pinnacles [CGH^{+}23]. pinniped [FWWH22]. Pioneer [MMB22]. Pisces [AGIS^{+}22, ASGI24]. piscivore [CBHM21]. piscivorous [RBS^{+}24]. plaice [vdVWF^{+}24]. plain [TNMN24]. planktivorous [ADF20, SW23b]. plankton [BTA^{+}20, DL20, DHA^{+}20, GCDA20, GGRV23, RKV^{+}21]. Planktonic [JKvLP20, EPB23, MAOR^{+}21, MDS^{+}21]. planning [FTA^{+}21]. Plant [MGH21, CAV^{+}20, CCMRII22, HSBB^{+}21, MLE^{+}22, RSL^{+}22, TSW^{+}21]. Plant-fungal [MGH21]. plants [SLH^{+}22]. Plasticity [GQBB^{+}24, BAS21, BBD^{+}22, CCRP20, CEG^{+}21, DPL^{+}21, GFH^{+}22, LBR^{+}24, MCFB20, MCC^{+}21, NP22, SJQ^{+}22, TBS20]. platessa [vdVWF^{+}24]. platforms [ESG20]. platypterus [BGDJ20]. play [SHM^{+}20b]. playback [SLJ^{+}20]. Plectropomus [CAP^{+}21]. Pleistocene [ASL^{+}22]. Pleuronectes [vdVWF^{+}24]. plots [CBJ^{+}23]. plume [ASH^{+}20, SSBA^{+}20]. plunge [Sig23]. plunge- [Sig23]. Pneumatophores [KVdM^{+}22]. Pocillopora [CFE^{+}22, DFB^{+}24, LHW^{+}20]. podocysts [DA24]. Polar [HDL^{+}21, AAP^{+}23, BLW^{+}24, BBL^{+}21a, BAAR20, DDG^{+}21, FWA^{+}23, GBSS24, KTR^{+}20, LAP^{+}24, MSC^{+}22, SLB^{+}21]. poleward [OHBJ20]. policy [DTHC24]. pollock [OLM24]. polyandry [HB24]. polybranchiata [YWW^{+}24]. Polycerella [CCBT^{+}20]. Polychaeta [ARN^{+}22, CGLC20]. polychaete [GDG^{+}22, PGC21]. polymerus [AR21b]. Polymorphic [FNO^{+}23]. polymorphisms [PSÓ^{+}24]. Polynesia [BPSB^{+}22, CFE^{+}22, VLPN24]. polyphemus [CCW21, TWC20, WLC22]. polyps [DA24, ICIS23, LLL20]. pomacentrid [MHH^{+}22]. Pomacentridae [CPH20]. pool [CEAO^{+}22]. pools [LDM^{+}20, SST^{+}20a]. Poor [CRGMA^{+}21, CDD^{+}20b, NGRH20, SMB^{+}20]. Pop [TCM^{+}23]. Pop-up [TCM^{+}23]. Population [ACvdC^{+}21, BKR^{+}21, CNC^{+}20, CDD^{+}20b, DIC^{+}22, MTB^{+}23, MHÁC^{+}22, ZSC^{+}22, AESS24, ABH^{+}21, Ano23, AWC21, BNS^{+}24, CKT^{+}24, CASF21, CLR^{+}21, CPD^{+}20, CHDG^{+}23, CH20, DMCB22, EHLM^{+}23, FTA^{+}21, FNM^{+}21, FDE^{+}21, FCV^{+}24, GCDP21, GWC^{+}21, GB21, HGK^{+}22, HB24, HRC21, JWH^{+}21, JETGM24, KRMR24, KI24, LGGR^{+}22, MPF^{+}22, MTB21, MA20, MDQ^{+}21, NDA^{+}22, PBO^{+}21, PBS^{+}24, PVB^{+}20, PGW23, RLME^{+}21, RCES22, SW23b, SHG^{+}21, SRS^{+}20, SB20a, SVT^{+}24, SBFP^{+}24, SDE^{+}21, WPM22]. population-level [MTB21]. Population-scale [DIC^{+}22]. population-specific [JWH^{+}21]. populations [ABB^{+}20, CGL^{+}22, CLC^{+}23, CKP20, CCP^{+}21, FTA^{+}21, GAFS20, GON^{+}20, HOK^{+}20, ILON21, KDGG24, KTW^{+}22, LAD^{+}21, MDY^{+}21, NSMJ20, OvBS^{+}24, PSÓ^{+}24, PDM22, RPD^{+}23, RBVA23, SWP20, ŠGH^{+}21, SWJ^{+}21, XW24, ZZX^{+}20]. porbeagle [WHB^{+}20]. porbeagles [ABN^{+}21]. porpoise [NRE21, SvBS^{+}20]. porpoises [HJG21, SBS^{+}22, SWC^{+}23b]. Portugal [PRM^{+}21]. portunid [dlBSR^{+}20]. Posidonia [AWB23, MDH^{+}20]. position [BSÁG^{+}22, BPS^{+}24b, CBB^{+}22b, FWWH22, OFSJ20, RDL^{+}21, WFR^{+}23]. Positioning [OH22]. positions [PBKG22]. positive [CAV^{+}20, THL^{+}20, dFRvdG^{+}22]. positively [OFSJ20]. possible [MY21, THK^{+}22]. Post [LPJ^{+}21, ABN^{+}21, DFCM21, DHA^{+}20, FSF^{+}24, FOM^{+}20, FLC^{+}24a, JDM^{+}20, KCGR20, STF^{+}24, YSP^{+}22]. post-bloom [DHA^{+}20]. Post-fledging [LPJ^{+}21, YSP^{+}22]. post-larval [JDM^{+}20, KCGR20]. post-peak [FLC^{+}24a]. post-release [ABN^{+}21]. post-settlement [DFCM21, FSF^{+}24, FOM^{+}20, STF^{+}24]. postcanine [IMYW24]. Potential [BWGH23, SBW^{+}22, WSPS^{+}20, ABH^{+}21, ACA^{+}24, BQM^{+}24, BJH^{+}23, COT^{+}21, FOM^{+}20, GRB22, GED20, GWC^{+}21, JWH^{+}21, LR22, MMLPP^{+}24, NCS^{+}23, OVÁS^{+}21, PFDF^{+}21, RSL^{+}22, SFL23, YRB^{+}20, YSD^{+}21]. potentially [FNH^{+}23, RJ21]. pourtalesii [BKM^{+}21]. poutassou [MAOR^{+}21, PJA^{+}21]. power [MLE^{+}22]. practical [HJG21]. Pre [SSB^{+}23, DRP^{+}20, MTCT^{+}23]. pre-breeding [MTCT^{+}23]. Pre-industrial [SSB^{+}23]. pre-molt [DRP^{+}20]. Predation [BS21, GBA21, Whi21, CBP22, DJR^{+}20, DKAB23, DVP^{+}23, DFCM21, DDG^{+}21, FSF^{+}24, FNM^{+}21, FSAG22, FBC^{+}24c, HHBP23, JJV^{+}21, LK21, MSR^{+}21a, MBGM21, MMT^{+}24, MSC^{+}20, RBB^{+}24, SSBA^{+}20, TFCS22, VFG24, Whi20, dSLD24]. Predator [CKP20, KLFP23, LSKS24, ASBA^{+}20, BSRVS22, BRSD20, COM^{+}24, CSD^{+}24, DHM21, dOFBSdJR^{+}22, FLC^{+}24b, GGCL22, GSL20, GDCP20, GLG21, GMLC22, GBA24, HMB^{+}22, HHY^{+}23, IMM^{+}20, KDS^{+}20, MSH20, MCFB20, NHFS22, PBT^{+}21, RHSM21, RJ21, RB23, RS20a, RCB24b, SHM^{+}20b, SFOPL21, SSM20, TNMN24, WWH^{+}22, YLH20]. predators [APH^{+}23, BBL^{+}21b, CLV^{+}20, GLG21, HGR^{+}22, HHY^{+}23, HO24, LFFW21, MDS^{+}23, MLVJ23, MMA^{+}22, OFSJ20, OCGAP^{+}22, PBRF23, RBS^{+}24, RAF^{+}22, VNG^{+}22]. predatory [CGGdD24, CGH^{+}23, DFP^{+}21a, DSHM20, GAF20, LSS^{+}22, SMH^{+}20, WSB^{+}24]. predict [CMF21, EPB23, GBW^{+}20]. predictability [NT24]. predictable [SRG^{+}23]. predicted [AU21, DdGA^{+}21, NHFS22, PTCD22, SHT^{+}22]. Predicting [CHDG^{+}23, DNJCH21, EVH^{+}21, GED20, JDM^{+}20, LRD^{+}20, NSPH20, RTG^{+}20, dlBSR^{+}20, BSH^{+}21b, BKM^{+}21]. predictions [GSL20]. predictor [ASY^{+}24]. predictors [STH21, SMB^{+}20]. predicts [GWC^{+}21, SBT24, SWP^{+}24]. preference [CBWH24, FSK^{+}21, HYG^{+}20, SWC^{+}23b]. preferences [ASGI24, BGMRGM^{+}20, BLCC^{+}20, LSKS24, VOM^{+}20]. pregnancy [HPRP21]. pregnant [FSH^{+}21]. Preliminary [ABN^{+}21, LBMSS23, MTRM21]. premature [SLBK^{+}20]. presbyter [VMCA^{+}20]. presence [LD24, LRC20]. presents [Ano23]. pressure [FD21, MBGM21, RS21, SSBA^{+}20, TP20, VKdJ24, WLB23]. pressures [CBVA^{+}21, SBB^{+}24]. prevalence [ANP^{+}24, DMBB22, GVCSO^{+}24, KCLJ20, RBCAV22, ZBB^{+}20]. previous [BAF20]. Prey [FSDB20, GG23, ADF20, ASBA^{+}20, BRD20, BBSM^{+}20, CSH^{+}21a, CGGdD24, CGR^{+}20, CS22, DCG^{+}20, FDF20, FOO20, FWA^{+}23, GSL20, GJM^{+}20, GBF^{+}23, HMB^{+}22, HSO^{+}22, HTGG22, KPW^{+}20, LGD23, LFFW21, LSS^{+}22, LSKS24, MSC^{+}22, MCF^{+}21, MDK^{+}21, OBW^{+}20, OJPC23, PTCD22, PWvdV20, RRR^{+}23, SSBA^{+}20, WAA^{+}23, WYO^{+}22, WDJ20, YLH20, dDW^{+}22, vdWDH^{+}21]. preying [DME^{+}23]. primary [BMPD20, ICIS23, JKvLP20, KPW^{+}20, LEG^{+}20, OSJ^{+}22, RSL^{+}22]. priming [GJK^{+}24]. Prince [DMCB22, vTLG^{+}21]. prion [DCR^{+}22]. Prionace [EMDR22, FSH^{+}21, FKS^{+}22]. prions [DCK^{+}24]. Prior [BICQG23]. priorities [PYB^{+}21]. Priority [BVKF23]. probability [GBG^{+}23]. probing [CKM^{+}20]. procedure [LH22]. process [BS21, HvLM21]. process-based [HvLM21]. processes [AS24, Ano23, BQM^{+}24, JNJ^{+}20, KTK21, SBSS21, STY^{+}20, WJ21]. production [Bai22, BSP^{+}21, BB24, BDB^{+}20, BMPD20, CSBM20, CRLC21, GCDA20, GGRV23, GJK^{+}24, GBSS24, JKvLP20, KSS^{+}20, LEG^{+}20, QMDGM22, RNQ^{+}20, RSL^{+}22, RB20, SBS20, SE20, WLB^{+}20, YBSH23, ZCF20]. productivity [AM20, BTDR20, CFSH22, DNE^{+}24, EKRT^{+}20, HPFB20, KPW^{+}20, MIL^{+}20, RVSA^{+}21, SNP20, SSG^{+}22]. productivity-susceptibility [BTDR20]. profiles [ZMD^{+}21]. program [GSL^{+}24b]. Progress [PEBG^{+}20]. progresses [SWHE20]. progression [TBH21]. Projected [FNM^{+}21, DHE^{+}23]. projections [RHN^{+}24]. prokaryote [CS22]. prolifera [AS20, OVÁS^{+}21, WCS^{+}21]. proliferation [OVÁS^{+}21]. prominent [LFA^{+}23]. promising [GJK^{+}24]. promote [JDDF21]. propagation [LT21, RCES22]. Propagule [GB21]. Proper [ZLL^{+}24]. properties [HW20, OAdJA^{+}20]. Proportional [SAH^{+}24]. prospectus [WKAH22]. Protected [SVT^{+}24, ACC^{+}20, AKGA^{+}21, BQM^{+}24, BWC^{+}22, CDF^{+}20, DVP^{+}23, FST22, HCP^{+}24, MSL^{+}24, MBS^{+}24, dACMRZ^{+}20, PSP^{+}20, RMR^{+}23, SW23b]. protection [BWGH23, HESF24, LCR^{+}22, NCS^{+}23, SLH^{+}22]. protein [ZDBS20]. Protist [MMD20, MKS^{+}20]. protists [FMD20]. protogynous [LBMB^{+}20]. prove [ASL^{+}22]. provide [AFTW24, Ano24o, FSF^{+}24, GVAZ^{+}22, KF20, MVH^{+}21, NMB^{+}22]. provides [JPJ^{+}22, MSH20]. province [MFC^{+}24]. provisioning [CM22, SWP^{+}24, ZDBS20]. proximity [LMS^{+}20]. proxy [CBB^{+}22b]. Pseudodiaptomus [NGRH20]. pseudonitzschiae [YHW^{+}23]. Pseudopleuronectes [DFCM21, MLB^{+}21, TFCS22]. Pteroclava [SMA^{+}20]. Pterodroma [CVG^{+}23]. Pterois [HBJ^{+}21, TGB^{+}20]. Pterotracheoidea [AMMADDH20]. PUFA [NGRH20]. PUFA-poor [NGRH20]. puffin [DSK^{+}20]. puffins [RLME^{+}21, RMD24, SWC^{+}23a, SD22]. Puget [ABH^{+}21, EWF^{+}21, WFR^{+}23]. pugettensis [LRC20]. pulses [APG^{+}22, ATLT20]. pump [ALL^{+}23, RL24]. pumping [SVS^{+}22]. purple [BKTN21, BMB24, DMM^{+}24b, KNH24]. purpuratus [DMM^{+}24b, KNH24]. pursuit [GJM^{+}20]. pursuit-diving [GJM^{+}20]. Pusa [BOY^{+}23, LHM^{+}23, VGLL^{+}21]. Putting [BRD20]. puzzle [PAL^{+}22]. pygmy [BBGM23]. pygoscelid [CCBD^{+}21]. pyrifera [BE21, BLWJ20, ENF^{+}22, LDPH24]. Pyrosoma [SSB^{+}20].

qualitative [HO21]. quality [HSO^{+}22, KPK^{+}24, LLJ^{+}20, LHHH^{+}20, RAA^{+}23, SNH^{+}20]. Quantification [PK24]. quantify [BCP^{+}22, CP22b, MDW^{+}22]. Quantifying [CMTP22, DME^{+}23, HvdSR^{+}21, QFT^{+}24, BGG^{+}22, HSBB^{+}21, OA24, SME^{+}23]. quantity [HSO^{+}22, KPK^{+}24]. quarter [JBWS24]. quasi [BDJ^{+}24]. quasi-stationary [BDJ^{+}24]. Quebec [DNE^{+}24]. queen [VSGD24]. Queensland [CMTP22]. quillback [FSDB20]. quinqueradiata [FKTK20].

rabbitfish [VPH^{+}21]. radiation [GCDA20, KCLJ20]. radiatus [LLF^{+}20]. radiocarbon [FLC^{+}24a]. rafting [NF23]. Raja [DHM^{+}24]. Rajidae [SHS21]. Rajiformes [SOS^{+}21b]. random [CFE^{+}22, CKKL21]. Range [AESS24, ASL^{+}22, Ano23, Ano24o, BJH^{+}23, BLWJ20, CKT^{+}24, CHD^{+}22, DFP^{+}21a, DNJCH21, EVAB^{+}23, GVAZ^{+}22, IBC^{+}21, KLA21, LPD^{+}24, MDB^{+}21, MCFB20, MCC^{+}21, OHBJ20, OOW^{+}20, PSBH21, PGGE21, RRS^{+}23, SEJ^{+}24, SBB20, SSG^{+}22, TPGG^{+}23, TWC^{+}23, TFH^{+}22, VDG24, WBS24, WAA24, WRB21, vdKMC^{+}24]. range-edge [CKT^{+}24]. range-expanding [SBB20, WBS24, WRB21]. range-restricted [KLA21]. range-shifting [Ano24o, PSBH21, RRS^{+}23, TFH^{+}22]. ranges [Ano24o, GMS^{+}21, SCM21]. ranging [COM^{+}24, KPW^{+}20]. Rapid [HO21, OCFEB21, VKdJ24]. rapidly [WCP^{+}21b]. rate [DGP22, FBC^{+}24c, HIT^{+}22, SSFD^{+}23, SVS^{+}22, TGB^{+}20]. rates [ASS^{+}21a, AZ24, BVD^{+}20, BAF20, BQC^{+}22, CS22, DPCL^{+}23, DVP^{+}23, DBP^{+}22, FDE^{+}21, GSBO21, IHA^{+}22, JNJ^{+}20, KBWRV21, LJ20, MHN24, MDH^{+}21, RL22, SFL23, STH21, Sig23, TYL^{+}24, VPH^{+}21, ZLL^{+}24]. rather [ZCM^{+}21]. ratio [SST^{+}20a, VS23]. rationale [TOMDC23]. ratios [ARE^{+}24, GVAZ^{+}22, GB20, NYW^{+}22]. ray [DCN^{+}23, FPZJAO^{+}22]. rays [ACC^{+}20, HGK^{+}22, KHSS23, MAMF21, OMW^{+}20, SHM^{+}20b, VvDRM20]. razorbills [LGD23, RMD24, SD22]. re [AZ24, CSB24]. re-established [AZ24]. re-examining [CSB24]. reach [MLVJ23]. reaction [MY21]. reactions [HJG21]. realized [JPMG23, LBR^{+}24]. reared [GMB^{+}24]. rearing [BDC^{+}22, RBW^{+}20]. reassessment [Bai22]. Rebuilding [MBG^{+}23]. recapture [BLE^{+}22, KSMP20, MJ21, SBL^{+}23, WFS23]. recently [OVO^{+}21, VK21]. recipient [Ano24o]. reclassified [OVO^{+}21]. reconstruct [XW24]. Reconstructing [WFR^{+}23]. Reconstruction [HKY^{+}20, EWF^{+}21, KdlVJ^{+}21]. record [CAPS20, RDL^{+}21]. recorded [FKTK20, THK^{+}22]. records [PAL^{+}22]. recovered [KRG^{+}20]. recovering [BKR^{+}21, BFC^{+}20, CPD^{+}20, GTH^{+}20]. Recovery [NF23, AAI^{+}24, BGGH^{+}22, CLGH24, CDS23, GWC^{+}21, KDLL21, KF20, MBR^{+}21, RCES22, SGP24]. recreational [CMTP22, SCB^{+}20]. recruited [HMK^{+}22]. Recruitment [BLBBL20, MRB^{+}24, APG^{+}22, BE21, CMTB^{+}22, FBC^{+}24b, FLY^{+}24, GSS22, HHO^{+}23, KLN^{+}21, KRMR24, KTK21, LT21, LvAH^{+}24, LCT^{+}20, LCGS^{+}21, PLC^{+}20, PRC^{+}21, RCES22, RB20, SBB20, SBRQ21, TCT^{+}23, TR24, VGA20, WRW20, WK21, ZZX^{+}20]. Recruits [VHG20, BRU^{+}22, FOM^{+}20, LO20]. recurrent [BAS23]. recycling [SOSE22]. red [BDS^{+}20, BLBBL20, CAY24, DSK^{+}22, ESG20, HBJ^{+}21, PCLQ^{+}20, SVB^{+}24, EKRT^{+}20, MMA^{+}22, SMA^{+}20]. reddening [HMKK^{+}20]. Redfieldian [SST^{+}20a]. redfish [DSK^{+}20]. redox [YAZA21]. reduce [GGCL22, GFK^{+}23, IHA^{+}22, KRMR24, TdLHL21]. reduced [FD21, LGD23, ODB^{+}24, PDFH20, Whi20, YDG21]. reduces [ASBE22, DBS^{+}20, LSSG24, MSL^{+}21, SCK23, dFRvdG^{+}22]. Reduction [RBS^{+}24, ENF^{+}22]. reductions [WC21]. redundancy [EB21, HSNT24]. redundant [MR20]. Reef [CDS23, HHO^{+}23, dACMRZ^{+}20, MRH20, AMRL^{+}24, ACC^{+}20, ARB^{+}20, ARN^{+}22, BS20a, BGG^{+}22, BQM^{+}24, BSH21a, BFC^{+}20, BBH^{+}20, BQC^{+}22, CTTLCG^{+}24, CSBM20, CBVA^{+}21, CLGH24, CCL^{+}22, CGH^{+}23, DFP^{+}21a, DMBJ21, EVH^{+}21, FBP^{+}22, GAF20, GS20, GRQ21, GSRC20, GSPH20, GMCE^{+}22, GBC^{+}20, GMLC22, HO21, HNN^{+}24, JH20, KKC^{+}24, KKH22, KWD21, LMHS21, LS21, LCR^{+}22, MYSF23, MFC^{+}24, MMA^{+}22, MDH^{+}21, MHK22b, NMAE^{+}22, OFSJ20, PAAEV^{+}24, RNQ^{+}20, REGP22, SDW20, SBMW22, SHM^{+}20b, SME^{+}23, SFLQ23, SDM^{+}23, SdGWN21, SLJ^{+}20, THR^{+}23, TB21, VvDRM20, CSDB22, MHK22b, RCES22, SMH^{+}20, SYMR21]. reef-associated [CBVA^{+}21, DFP^{+}21a, MHK22b]. reef-building [CTTLCG^{+}24, JH20, SDM^{+}23]. Reefs [ATR^{+}24, AKU^{+}22, BMO^{+}24, BS20b, BDS^{+}20, CLD23, CGLC20, CFFC^{+}23, CRLC21, DKAB23, DVP^{+}23, EP23, EKRT^{+}20, Fai21, FBC^{+}24a, FQVMF23, FSSL^{+}20, GMS^{+}21, HMV^{+}22, HBS^{+}21, JHDE20, KETL22, LSSL21, MPA^{+}22, MBS^{+}24, MSPS22, MFCF21, MTSO^{+}24, MdMR24, MBR^{+}21, MHK22b, PTLK21, PAF^{+}22, PBT^{+}21, PCM^{+}20, RS20a, Rus20, RBCwD^{+}20, SBS20, SWP^{+}24, TCY^{+}22, VPB^{+}20, VDC^{+}24, VMCA^{+}20]. reel [ABN^{+}21]. reflect [FAJAO^{+}21, JMM24, MCF^{+}21, RVSA^{+}21]. reflectors [HMT21]. reflects [DSK^{+}20, FNL^{+}23, OBW^{+}20]. refuge [FSF^{+}24, GBA21]. refugia [LR22, ZCBC24]. regime [MCJ^{+}20]. regimes [BBEF^{+}20, EHLM^{+}23, SSG^{+}22]. region [AMUPJR^{+}24, BSÁG^{+}22, DMTP22, JBGA23, KDB^{+}24, NTN^{+}24].

Regional [ARH^{+}20, FLBP^{+}21, HAB^{+}22, ACC^{+}20, DCG^{+}20, STY^{+}20]. Regional-scale [FLBP^{+}21]. regionalization [AC20]. regions [Ano24o, BMF24, HOK^{+}22, SKTO23]. regius [VAMF22]. regression [CMF21]. regulation [CEZH21, EHHF^{+}20, HWT^{+}22]. Reina [NMAE^{+}22]. reinforcement [Whi20]. reintroduction [AJH22]. relate [GJM^{+}20]. related [ASGI24, CCC^{+}23, CMTB^{+}22, DHM21, Fai21, FLCQ21, LRD^{+}20, PSB^{+}23, SOS^{+}21b, TEK21]. relatedness [CMvKdR23, JMP^{+}24]. Relating [BBKW20, OJPC23, RLSO23]. relation [BDJ^{+}24, GWA^{+}23, GBF^{+}23, HGR^{+}22, IMM^{+}20, JFR^{+}21, KHOK22, LAGGM^{+}21, MBW^{+}20, SBS20, SCM^{+}23, SD22, YXB22]. Relationship [LPJ^{+}20, OYO^{+}20, SVS^{+}22, BBSM^{+}20, MHH^{+}22, OLM24, RBFA20, TRK^{+}22, ZZLS24]. Relationships [MHH^{+}20, BBEF^{+}20, BMBC21, HSBB^{+}21, KDB^{+}24, SPFL20, TKP^{+}20, Vir23]. Relative [FTA^{+}22, FBC^{+}24a, HHM^{+}22, PBKG22, QMDGM22, RKV^{+}21, TWC20, ZZYB21]. relaxation [GSN^{+}20]. release [ABN^{+}21, SSM20]. relevant [NP22]. reliable [TCM^{+}20]. remain [CCA^{+}21]. remaining [GAPM24]. remarkable [SLG^{+}24]. remodeling [CBP22]. remote [BWGH23, ECvBL21, FWGD^{+}23, KDS^{+}20, VPH^{+}21, WYZ^{+}20]. remotely [OSJ^{+}22]. removal [AAI^{+}24, FQVMF23, SHM^{+}20b, VGG^{+}23]. renewable [JFR^{+}21, SBW^{+}22]. reorganization [ENRC^{+}20]. reovirus [ZBB^{+}20]. repeatability [GBA^{+}22]. repertoire [HDM^{+}24]. Reply [DFRS^{+}24, Whi21]. reports [FFMR24]. represent [NSKH23]. representing [PHV20]. reproducible [TCM^{+}20]. reproduction [DVMH^{+}24, HTT^{+}21, KDGG24, LSC^{+}21, LLL20, RHM^{+}22, RWM^{+}22, SCK23, WRS^{+}23, ZSC^{+}22]. Reproductive [CPD^{+}20, GVOCSGG21, GCCM20, SFL23, AOO^{+}21, ADBF21, BAS21, FSK^{+}21, HB24, ILON21, JWH^{+}21, KRMR24, LPB^{+}23, MBB^{+}22, MMMV22, MTB21, PDFH20, RVCT21, RHN^{+}20, RHSM21, SHH^{+}22, SF23, WYO^{+}22, ZFD^{+}24, ZZX^{+}20, vdVVC^{+}21]. reptile [FNO^{+}23]. requirements [BAAR20]. requires [ECC^{+}21]. rescue [CS20]. Research [CRÁ^{+}20, ARE^{+}24, Ano24o, DTHC24, EB20, HvLM21, PYB^{+}21, TCM^{+}20, vdHMA^{+}24]. reserve [BWGH23, FBP^{+}22, MDPL20, SAMdL^{+}20]. reserves [ATR^{+}24, DVP^{+}23, MG22, MSDY23, TBH21, VNM^{+}21, YLH20]. residence [AUS^{+}23, KBP22, YAZA21]. Residency [ECvBL21, FPB^{+}21, APOGG20, ACC^{+}20, CWD^{+}22, CKKL21, DCN^{+}23, EMB^{+}22, HGR^{+}22, HCP^{+}24, JW22, LMB^{+}20]. Resident [TFH^{+}22, CMO^{+}20, MMMV22, PBRF23, RDL^{+}21]. residents [LPD^{+}24]. Resilience [DL20, STM^{+}24, WPG^{+}24, AGA^{+}22, BG20, BG21, CDS23, HBR^{+}21, MSH20, RVV^{+}24, VHG20, vdHMA^{+}24]. resiliencies [BGL^{+}21]. Resilient [BWC^{+}22, Mon23]. resistance [AGA^{+}22, WSU21]. resistant [Mon23]. resolution [SMK23, TTJS20]. resolve [WYO^{+}20]. Resolving [CAEG21, MCvdM^{+}24]. Resource [BLGM^{+}24, SD22, BBR24, CCRP20, CRG^{+}20, CAPS20, FJJ^{+}21, MSV21, PWvdM^{+}21]. resources [MGH21, MBG^{+}23, OYO^{+}20, PVL^{+}21, RLP^{+}23, vLSB^{+}21]. Respiration [BAF20, Hew21, LAGGM^{+}21]. respond [MR20]. responded [ATLT20]. responds [MGH21]. Response [MJ23, NBDM^{+}24, VIV^{+}20, AVFD^{+}21, Ano24o, AR21a, BB24, BBM^{+}24, BAS21, BS20b, CFHCS^{+}24, CBP22, CGR^{+}20, CFSH22, CKS^{+}24, DHE^{+}23, FMD20, GD23, HHY^{+}23, ICIS23, JPJ^{+}22, KBP22, LRM^{+}21, LPD^{+}24, LLL20, MDG^{+}21, ODB^{+}24, OBD^{+}21, OOW^{+}20, PGGE21, SSP^{+}22, SLH^{+}22, SLJ^{+}20, WGK22, XMY^{+}22]. Responses [FD21, TSW^{+}21, XWW^{+}21, BMBC21, CCC^{+}23, COM^{+}24, CG22, CKP20, CHJ21, CDP^{+}21, COWM^{+}22b, DBS^{+}20, GCDA20, GON^{+}20, GBA24, HJG21, HWA^{+}22, HPFB20, HRC21, JRVE^{+}22, JTBS^{+}22, KKC^{+}24, KGM^{+}24, LK21, LBB^{+}21, LEG^{+}20, LSS^{+}22, MDM^{+}20, RKV^{+}21, RCB^{+}24a, ŠGH^{+}21, TNS^{+}22, TWC22, UPK^{+}20, VGG^{+}23, VLG^{+}21, WMT20, WYO^{+}22, WBGC24]. resting [CCR^{+}20]. restoration [AJH22, BJH^{+}23, CLR^{+}21, GJK^{+}24, HGK^{+}24, MGHH21, MNG^{+}20, RHM^{+}22, RNQ^{+}20, SNH^{+}20, SLO^{+}22, SNE^{+}24, TWC22, XYSF20, XWW^{+}21, ZCBC24]. restored [STH21, SWP^{+}24]. restores [WCP^{+}21b]. Restricted [AAI^{+}24, KLA21, SVT^{+}24]. restrictions [CCJ^{+}23]. result [CCP^{+}21, PDFH20]. results [BBMM20, KHL^{+}22, RB23]. resuspension [PRD^{+}20, PBC^{+}22]. retention [CHDG^{+}23, QFT^{+}24]. retention-oriented [CHDG^{+}23]. retentive [SRHM21]. return [SLBK^{+}20]. reveal [BTMS24, CVG^{+}23, DFRS^{+}24, DAT^{+}21, FSDB20, GGN21, GWDR20, HDM^{+}24, HLCH23, KPW^{+}20, LKR^{+}21, LPD^{+}24, dACMRZ^{+}20, OLG^{+}23, PAL^{+}22, PLH^{+}21, PSN^{+}20, PBT^{+}21, RHSM21, SVT^{+}24, SHT^{+}22, TCM^{+}23, VH20, WSB^{+}24, ZMD^{+}21]. revealed [CKM^{+}20, CILGD^{+}22, DFSH23, FLBP^{+}21, KCGR20, LMS^{+}20, LORB^{+}24, PBB^{+}20, WD22, ZSWR^{+}22, vTLG^{+}21]. reveals [BGMRGM^{+}20, CDF^{+}20, CDD^{+}20b, GMS^{+}21, HHO^{+}23, HW20, MBR^{+}21, MHH^{+}22, OCFEB21, SCB^{+}20, SJS^{+}23, TWC^{+}23, TMN^{+}24, VFG24, XWH^{+}22]. Review [GAPM24, JMM24, MSR^{+}21b, RCC^{+}24, WH24, FDL^{+}21, WKAH22, ZDBS20, BLVL24, LMH^{+}24]. Rhincodon [ASPM21]. rhinoceros [STO^{+}23, SWHE20]. Rhizophora [HBC^{+}20]. Rhodes [vLBGGT22]. rhodolith [QMDGM22]. rhodoliths [BG20, BG21]. rhomboides [LOF^{+}24]. Rhopilema [DA24]. rhythms [TWC20]. Ría [LAGGM^{+}21]. Riaz [Ano21h]. ribbed [RSL^{+}22]. Rica [BPS^{+}21]. rich [HDM^{+}24]. richness [BGG^{+}22, CCAdA24, FST22, JDM^{+}20, SME^{+}23, TKP^{+}20]. Ridge [MHÁC^{+}22]. ridley [ARE^{+}24, ARH^{+}20, LJ20]. right [ASS^{+}21b, CDD^{+}20b, CBB^{+}22a, FNL^{+}23, HCvdHM23, MBW^{+}20, RYF^{+}24, RRR^{+}23, SDE^{+}22]. Rimicaris [MHÁC^{+}22, ZCB20]. ring [UIS^{+}24]. ringed [BOY^{+}23, KGG^{+}21, LHM^{+}23, VGLL^{+}21]. ringens [CMT^{+}21]. rise [CCMRII22, FBC^{+}24b]. Rising [DA24, MR20]. risk [BLCC^{+}20, CCMRII22, CBP22, CWC^{+}22, GRTK^{+}21, HHBP23, MSC^{+}20, MHDJ21, RCB24b]. risks [HSNT24, RB23]. Rissa [EGK^{+}21, MIL^{+}20, OBW^{+}20]. River [MMS^{+}22, ASBA^{+}20, HPDG22, SSBA^{+}20, WKR21, JPJH23, NMB^{+}22]. river-dominated [ASBA^{+}20, HPDG22]. rivers [TFCS22]. Robotic [MDS^{+}21]. rock [GSP^{+}20, THD^{+}22]. rockfish [ABH^{+}21, BAS21, EPHS23, FSDB20, FST22, QFT^{+}24, WTK^{+}24]. rockpools [MGBCGM^{+}22]. rocks [MCK21]. rocky [ARB^{+}20, BLB23, CFFC^{+}23, CGM23, EB21, GLB^{+}20, HBS^{+}21, JHDE20, OFSJ20, PTLK21, SB20a]. rod [ABN^{+}21]. rod-and-reel [ABN^{+}21]. rodgersii [DKS^{+}24]. Role [MG22, SMS^{+}24, WJ21, ZBB^{+}23, CMdL^{+}20, CGL^{+}22, FHAP20, GED20, HHY^{+}23, JMM24, JFB^{+}23, KPW^{+}20, NT24, PTLK21, RAF^{+}22, RCC^{+}24, SW23b, VDC^{+}24, WCS^{+}21]. roles [MDS^{+}23, WKAH22]. rookery [BDB^{+}20, RMR^{+}23]. roots [IHA^{+}22]. rosaceus [BAS21]. rose [ZFD^{+}24]. rosy [BAS21]. round [BGMAM^{+}21, FTA^{+}21, FDS^{+}21, VH20]. rounding [DME^{+}23]. routes [CDS20, FJB^{+}23, HMK^{+}22, WLA^{+}21]. royi [NGRH20]. rubens [SKHC23]. rugosa [GVCSO^{+}24]. rules [BMK^{+}22]. runoff [CSDB22]. Rushan [WLB^{+}20]. Russia [EGK^{+}21]. Ryukyu [YMU^{+}21].

S [SB20b]. S. [GLP^{+}24]. Saanich [KSS^{+}20]. Sabellaria [ARN^{+}22, VDC^{+}24]. Sabellariidae [ARN^{+}22]. sablefish [CBHM21]. Saccharina [BAHG23, DB21, GJK^{+}24, YDG21]. Saccostrea [GSP^{+}20]. safely [DDB^{+}24]. saida [BLW^{+}24, DDG^{+}21, MSC^{+}22]. sailfin [KI24]. sailfish [BGDJ20]. Saint [DCR^{+}22]. saira [FNH^{+}23]. saithe [MDQ^{+}21]. salar [JWH^{+}21]. salinities [BAF20]. salinity [ATR^{+}24, BKT^{+}22, CCMRII22, CMT^{+}21, CH20, HYG^{+}20, KDGG24, LSS^{+}22, MHN24, MVSV22, RS20a, ŠGH^{+}21, TGB^{+}20, WKR21]. Salmo [JWH^{+}21]. Salmon [ACA^{+}24, AHCT22a, AHCT22b, AGB^{+}23, BAB^{+}24, CSK^{+}23, DCG^{+}20, FYM^{+}24, FSDB20, JWH^{+}21, JPJH23, JNY^{+}20, LH22, LAD^{+}21, MHF^{+}21, NMB^{+}22, RS20b, RSvV^{+}23, SHB^{+}20, SLBK^{+}20, SBJ^{+}24, TBH21, WDJ20, WPM22]. salmonis [TBH21]. Salt [CCMRII22, DBK^{+}23, LPD^{+}24, MGH21, NF23, RLP^{+}23, RDT^{+}23, RSL^{+}22, RB20, SAH^{+}24, TSW^{+}21, XYSF20, dlBSL^{+}22]. Saltmarsh [ZZYB21, WJ21]. Salvador [SC23]. Salvelinus [HHM^{+}22, HYG^{+}20]. Samoa [BGL^{+}21]. sample [HO23]. samplers [CSWM24]. Sampling [WCL20, WCP^{+}21a, EPHS23, ECC^{+}21, HHS21, KCT24]. San [CMO^{+}20, SC23]. sanctuary [RGL24]. sand [BJM^{+}22, BQC^{+}22, CWD24, NHH^{+}21, RBS^{+}24, SF23, SNE^{+}24, UBO^{+}24, VMCA^{+}20]. sand-cap [SNE^{+}24]. sand-dwelling [BQC^{+}22]. sand-smelt [VMCA^{+}20]. sandeel [LBW21, OBW^{+}20, RWH^{+}24]. sandfish [KI24]. Sandwich [FTA^{+}22]. sandy [BNSH21, BSA^{+}24, CGL^{+}22, HW20, IHA^{+}22, LFA^{+}23, RPD^{+}23, TQG24, VGA20]. sandy-beach [BNSH21]. sanguineus [RAA^{+}23]. sapidus [HAB^{+}22, LSSL21, SRK^{+}23, TFCS22, ZBB^{+}20]. sardine [BQM^{+}22, FBC^{+}24c, FYK^{+}20, NTN^{+}24]. Sardinella [BRC22]. Sardinops [NTN^{+}24]. Sargassum [GS20, HOK^{+}20, KKN^{+}20, KHOK22, SE20]. sargus [PSB^{+}23]. Satellite [EMDR22, BZT^{+}23, BASM^{+}24, CBTL20, DGM^{+}23b, HDA^{+}22]. satellite-tagged [HDA^{+}22]. satellites [KBE^{+}22]. Saudi [SMA^{+}20]. saury [FNH^{+}23]. save [SBS20]. Saved [LSSL21]. sayi [KLFP23]. Scalability [HSBB^{+}21]. Scale [CATF^{+}21, ACA^{+}24, ASBA^{+}20, BMW24, BVKF23, BPA21, CCAdA24, CGL^{+}22, CH20, DIC^{+}22, EP23, FPG^{+}21, FLBP^{+}21, FLC^{+}24b, FCV^{+}24, FNH^{+}23, GVR^{+}20, GMS^{+}21, HBS^{+}21, ILON21, JMR21, JMP^{+}24, KCT24, KSW23, LT21, LOS^{+}22, LO20, LFK^{+}21, MDH^{+}20, MTSO^{+}24, MHK22a, MCJ^{+}20, MMS^{+}20, NGPGO^{+}23, NFDF^{+}20, OH22, OWH23, PBAC23, PBS22, PFM^{+}21, PS23, PVB^{+}20, RB22, RBCAV22, SWR23, SHM^{+}20a, SHS21, SDM^{+}23, SWC^{+}23b, SLO^{+}22, SNE^{+}24, SSBA^{+}20, TKS^{+}20, TD21, VDG24, WDJ20, WYZ^{+}20, XW24]. scales [Ano23, BMO^{+}24, BGL^{+}21, CP22b, FP20, KDB^{+}24, MAPFH22, OFSJ20, OÁBP^{+}21, RBB^{+}24, THL^{+}20, THR^{+}23]. scallop [TCT^{+}23]. scalloped [EMGMEV^{+}21, GLP^{+}24]. scallops [GMB^{+}24, KRMR24]. scarring [LMB^{+}20]. scat [KRG^{+}20]. scat-recovered [KRG^{+}20]. scats [TMN^{+}24]. scavengers [APH^{+}23]. scavenging [SSP^{+}22]. scenarios [HBJ^{+}21, NHFS22, YSD^{+}21]. schmardae [OVO^{+}21]. school [DIC^{+}22]. schooling [WAA^{+}23]. Sciaenid [MMS^{+}22, VAMF22]. Sciaenops [BLBBL20]. science [NJHN24, NRE21, PYB^{+}21, vdHMA^{+}24]. scintillans [ZSC^{+}22]. scleractinian [KZC23, NYW^{+}22]. sclerochronology [CTTLCG^{+}24]. scope [ASY^{+}24, CHJ21]. Scotia [ASS^{+}21a, BKR^{+}21]. Scottish [MTRM21]. Scutellastra [SB20a]. scutulata [RH21, Str23]. Scyllarides [MSDY23]. SE [BQM^{+}24, GDG^{+}22]. Sea [ASS^{+}21a, BRC22, BLB23, CSD^{+}24, EGK^{+}21, MVH^{+}21, MKR^{+}20, NMAE^{+}22, ORB^{+}21, RKV^{+}21, RSEG21, VH20, WSD^{+}20, WCP^{+}21b, ZLL^{+}24, AAI^{+}24, ARE^{+}24, APH^{+}23, AGB^{+}23, AEHD21, ARH^{+}20, BKM^{+}21, BLA^{+}24, BBL^{+}21a, BBR24, BNH^{+}21, BSRVS22, BAH^{+}22, BKA^{+}20, BOY^{+}23, BMB24, BPSB^{+}22, CCRP20, CCMRII22, CTH^{+}20, CSWM24, DKAB23, DCR^{+}22, DKSS20, DNJCH21, DMM^{+}24b, DDG^{+}21, EBL^{+}21, ENRC^{+}20, FBM^{+}24, GBW^{+}20, GSL^{+}24a, GRB22, GFDN^{+}24, GB20, GTH^{+}20, GEY^{+}21, GLM^{+}24, HP23, HSSN^{+}21, Hew21, HBG^{+}22, IAG^{+}23, JWH^{+}21, JFB^{+}23, KP21, KTR^{+}20, KRMR24, KI24, KHG^{+}21, LKR^{+}21, LPJ^{+}20, LJ20, LvAH^{+}24, LKB^{+}21, LHM^{+}23, LNM22, LHHH^{+}20, MJ23, MSL^{+}21, MSR^{+}21a, MJJJ20, MVC^{+}23, MSF24, MAH^{+}20, MHF^{+}21, MMB22, MFCF21, MDS^{+}21, OBD^{+}21, PGT^{+}21, PAT^{+}24, RPB21, RBM^{+}21, RWR24, SVB^{+}24, SSP^{+}22, SLBK^{+}20, SRG^{+}23]. sea [SKHC23, SST20b, TBH21, TBS20, TPGG^{+}23, TKM^{+}22, TSS^{+}23, UPK^{+}20, VS23, WRW20, WIY^{+}23, WdGR^{+}20, WK21, WHRC22, WCL20, WCP^{+}21a, WKAH22, WLB^{+}20, ZCB20, ASL^{+}22, AGIS^{+}22, AOO^{+}21, ACG22, BML^{+}22, BD20, BAAR20, BSM23, BMP^{+}22, CCR^{+}20, CAY24, CKM^{+}20, CGR^{+}20, CSK^{+}22, CMvKdR23, DKSS20, DKSS21, EKRT^{+}20, FVC^{+}20, FSN^{+}20, FLY^{+}24, FSK^{+}21, FD21, FKTK20, GG23, GBSS24, GGT^{+}20, GBA24, HOP^{+}20, HSU^{+}21, HYG^{+}20, HGT^{+}21, HLCH23, HEP^{+}24, JKvLP20, KBE^{+}22, KDK^{+}24, KSMP20, KTK21, LEG^{+}20, LCP^{+}23, LSKS24, MAPFH22, MAOR^{+}21, MMA^{+}22, NYW^{+}22, OLM24, OSJ^{+}22, OÁBP^{+}21, OCGAP^{+}22, PSÓ^{+}24, PWvdM^{+}21, PWvdM^{+}23, PJA^{+}21, RSvL21, RWH^{+}24, RLP^{+}23, SFF^{+}20, SMPLF^{+}22, SMA^{+}20, SS24, SGW^{+}21, SOS^{+}21b, SOS^{+}21a, TEK21, TTM^{+}22, TTM^{+}24, URSS24, UAL24, VIV^{+}20, Vir23, WCS^{+}21, YWW^{+}24, ZMC^{+}22, ZFD^{+}24, ZXL^{+}24, vESB^{+}23, vdVTW^{+}22]. Sea [vdVWF^{+}24]. Sea-ice [CSD^{+}24, MKR^{+}20, WRW20]. seabed [RSvL21]. Seabird [DHE^{+}23, KGM^{+}24, PBS^{+}24, WYO^{+}22, ABD^{+}21, AMT^{+}21, Ano23, CC22, CWC^{+}22, CKS^{+}24, DPP^{+}21, DMS^{+}21, DWJL^{+}21, FTA^{+}21, GJM^{+}20, KPW^{+}20, KPTT23, LBTE^{+}21, NCS^{+}23, OvBS^{+}24, ODB^{+}24, PGGE21, PHE^{+}20, RJ21, RDL^{+}21, RWH^{+}24, SAMP24, SBW^{+}22, SHH^{+}22, SGB^{+}21, SFMK^{+}23, TWC^{+}23, WPG^{+}24, WH24]. Seabirds [CBKM20, Ano23, BBG^{+}22, GFK^{+}23, MDY^{+}21, OOW^{+}20, OA24, RVSA^{+}21, RBB^{+}21, Sig23, SDF^{+}21, SDE^{+}21, TCP^{+}23, dSGC^{+}23]. seabream [PSB^{+}23]. seafloor [BSFM20, EPHS23, WIY^{+}23]. Seagrass [HMKK^{+}20, IHA^{+}22, JMR21, AS20, BP24, CFHCS^{+}24, CBJ^{+}23, FHAP20, GSS22, HAB^{+}22, HNN^{+}24, HO23, HDE21, ILON21, JRVE^{+}22, KSW^{+}21, MDH^{+}20, MGHH21, MA20, RSEG21, RH20, RVV^{+}24, RWSR24, SYMR21, STY^{+}20, TD21, TR24, VPH^{+}21, ZZX^{+}20]. seagrasses [LBMSS23, SBB^{+}24]. seal [BHQ^{+}22, FOO20, HPRP21, KdlVJ^{+}21, KGG^{+}21, LHM^{+}23, MTB21, MMT^{+}24, OBD^{+}21, ORB^{+}21, TMN^{+}24, YXB22, dLCN^{+}22]. seals [ATLT20, BKA^{+}20, BOY^{+}23, CGGdD24, CSN^{+}23, IMYW24, LNBdB21, MSC^{+}20, MHDJ21, NBL^{+}22, VGLL^{+}21]. seamount [GAF20, GIDP^{+}21]. Searching [DPP^{+}21]. seas [ARE^{+}24, NRE21, OBD^{+}21, DSB^{+}21, HLA^{+}21, RBM^{+}21]. Seascape [PYB^{+}21, SWP^{+}24, GVR^{+}20, HMW^{+}21, MDPL20, SHB^{+}20, SRR^{+}20, YMU^{+}21]. seascape-scale [GVR^{+}20]. season [DDB^{+}24, FSK^{+}21, HNKK21, JBGA23, MJ23, NM22, PRM^{+}21, PDF^{+}23, RAA^{+}23, RMD24, SBRQ21, WdGR^{+}20, ZBB^{+}20]. Seasonal [ARdlB^{+}22, AKGA^{+}21, BWRP20, BBRM22, CFFC^{+}23, DCN^{+}23, FSAG22, FFT^{+}20, FLC^{+}24b, FSH^{+}21, GFDPASR20, GBA^{+}22, JH20, KHOK22, LFFW21, PCR23, RF20, SGB^{+}21, SAF^{+}24, TBH21, VGLL^{+}21, VLPN24, ZMC^{+}22, ZFD^{+}24, ZSWR^{+}22, vdVVC^{+}21, ACA^{+}24, ADBF21, BSF24, CDC^{+}23, Fig21, GVOCSGG21, GWDR20, HSA^{+}21, HNKK21, HNB^{+}24, IAG^{+}23, KKN^{+}20, LKR^{+}21, LLF^{+}20, MSC^{+}20, PHV20, RVCT21, SJS^{+}23, UBO^{+}24, URSS24, WXZ^{+}24, WRG^{+}24, WFS23, ZWXC21, vdWDH^{+}21]. Seasonality [FBKG24, LAGGM^{+}21, MCJ^{+}20, PRCvdK23]. Seasonally [HPFB20, CHJ21, PBRF23]. seasons [BKTN21, HB24]. seawall [AHCT22a]. seawalls [HTL^{+}20]. Seawater [YBSH23, CKP20, DA24, KBWRV21, KNH24, SBC^{+}21]. seaweed [AS20, DB21, GJK^{+}24, TP20, VDG24]. seaweed-associated [TP20]. seaweeds [BS20b, CBGB24, dOFBSdJR^{+}22, MMP^{+}22, WS24]. Sebastes [BAS21]. Sebastolobus [DMTP22]. Sechura [BASM^{+}24]. Secondary [GBSS24, GDAC23]. Section [Ano23, Ano24o, CRÁ^{+}20, HvLM21, RBCwD^{+}20]. Sediment [GD23, SBT24, WRB21, AZ24, BMK^{+}22, CSBM20, CKM^{+}20, GBH^{+}20, HMV^{+}22, HW20, LRC20, LRD^{+}20, SRK^{+}23, STH21, SOSE22, TD21, TR24, XWW^{+}21]. sedimentary [OJ24]. sedimentation [AKU^{+}22, SKI^{+}20]. sediments [CIDM23, CSB24, IHA^{+}22, KVdM^{+}22, SNE^{+}24]. seed [JMR21, XWW^{+}21]. seedling [GB21, XWW^{+}21]. seeds [BP24, ZZYB21]. seep [ICH^{+}24]. segregation [ABWJ^{+}20, BGMAM^{+}21, CAAG^{+}22, CCA^{+}21, DCK^{+}24, FKS^{+}22, IBC^{+}21, MAS^{+}20, PVB^{+}20, SHS21, TWC^{+}23]. Seines [FWGD^{+}23]. seismic [VKdJ24]. selection [CFHCS^{+}24, CHD^{+}22, GG23, SHS21, SF23, SDM^{+}23, WRB21]. selective [KCGR20, SHZ^{+}20]. selectivity [CSH^{+}21a, YGMNA22]. Senilia [EHLM^{+}23]. senilis [EHLM^{+}23]. sensed [OSJ^{+}22]. Sensitive [SFMK^{+}23, ABB^{+}20, VKdJ24]. Sensitivity [CLF^{+}20, vdWDH^{+}21, ASY^{+}24, BLVL24, BMM21, BJM^{+}22, CTTLCG^{+}24, CATF^{+}21, FCV^{+}24, TEK21, WPM22]. sensor [CGM^{+}24]. sensors [MCL24]. sensory [RS20a]. sentinel [VLG^{+}21]. separated [CLR20, RBVA23]. separation [CFBPH23, HvdSR^{+}21, MRH20]. sequencing [FOO20]. sequestration [LOS^{+}22, PFDF^{+}21]. series [LEG^{+}20, NSM24, TTA20, VIV^{+}20]. Seriola [FKTK20]. serrator [RMASA20]. serve [LR22]. services [SBB^{+}24]. seston [SSFS20]. set [GG23, Sig23]. Setting [MSTH22, ARB^{+}20, Sig23]. settled [VK21]. Settlement [JNJ^{+}20, STF^{+}24, BMB24, DFCM21, DMTP22, FVC^{+}20, FSF^{+}24, FOM^{+}20, GCDP21, JMS^{+}22, KKH22, LT21, PGM^{+}23, SBMW22, SMK23, WRS^{+}23]. seven [PHE^{+}20]. Several [ZCM^{+}21]. severe [WPG^{+}24]. severity [SSFD^{+}23]. sewage [GLB^{+}20]. Sex [ARE^{+}24, CTTLCG^{+}24, PFM^{+}21, RB23, BBR24, CAW^{+}22, COT^{+}21, EBL^{+}21, FKS^{+}22, LLJ^{+}20, LBMB^{+}20, ZLHH20]. Sex-associated [CTTLCG^{+}24]. Sex-biased [PFM^{+}21]. sex-identification [COT^{+}21]. Sex-specific [RB23, BBR24]. Sexual [CCA^{+}21, RNQ^{+}20, ADBF21, PVB^{+}20, RHM^{+}22, SHS21, ZZX^{+}20]. Seychelles [ACvdC^{+}21, EVAB^{+}23]. shade [HBC^{+}20]. shadow [SRHM21]. Shady [MBT^{+}20]. shag [WHN^{+}22]. shags [MDB^{+}21]. shallow [BVD^{+}20, CGCC24, CWD24, DVP^{+}23, EVH^{+}21, ECC^{+}21, GAF20, GSPH20, HESF24, HYG^{+}20, KETL22, Whi21]. shallow-water [CGCC24, GSPH20]. shanny [JMP^{+}24]. shape [BSF^{+}20, LÁGLL20, YGMNA22]. shapes [Ano23, ENF^{+}22, PLG^{+}23]. Shaping [BBD^{+}22, MIL^{+}20, NT24]. share [WLA^{+}21]. shared [GSS22]. Sharing [DMS^{+}21]. Shark [CLM^{+}21, APH^{+}23, ACA^{+}24, ASPM21, BML^{+}22, BLGM^{+}24, CDF^{+}20, CWD^{+}22, CBVA^{+}21, CLF^{+}20, DBP^{+}22, EMB^{+}22, EMGMEV^{+}21, FBP^{+}22, GGN21, HCP^{+}24, KKC^{+}24, KKS^{+}20, LBH^{+}21c, MPB^{+}24, NSPH20, PBO^{+}21, PTCD22, PDH^{+}24, SMB^{+}20, SdGWN21]. sharks [AKGA^{+}21, BGMRGM^{+}20, BSP^{+}22, CBVA^{+}21, CMF21, DVR^{+}23, DIC^{+}22, EMDR22, FBC^{+}24a, FSH^{+}21, FKS^{+}22, HGR^{+}22, JJV^{+}21, KHSS23, LMB^{+}20, LBMF^{+}24, MTK^{+}21, SHM^{+}20b, SAMdL^{+}20, TCM^{+}23, VDB^{+}24, WFS23]. sharksuckers [PBKG22]. sharp [BTB^{+}24]. sharptail [CDC^{+}23]. shearwater [MCJ^{+}20]. shearwaters [BLCC^{+}20, GLW21, PRM^{+}21, WdGR^{+}20]. sheath [BP24]. sheath-covered [BP24]. Sheaves [Whi21]. shedding [DBP^{+}22]. sheepshead [EHHF^{+}20].

Shelf [VLG^{+}21, BGG^{+}22, BMK^{+}22, BMOW21, CIDM23, CKS^{+}24, DHJ^{+}22, DSA^{+}24, DNJCH21, KF20, LRD^{+}20, MSL^{+}21, SIBM20, VBB^{+}21, LSR^{+}24]. shell [BGGH^{+}22, GVOCSGG21, GVCSO^{+}24, LSSL21, RBFA20, SWP20, CBP22]. shield [LSSL21]. shift [BAS23, FNL^{+}23, GSL20, NSMJ20, REGP22]. Shifting [OLM24, Ano24o, GB21, MBGM21, PSBH21, RRS^{+}23, TFH^{+}22]. Shifts [JBWS24, MJW24, AESS24, Ano24o, BS20b, BKH^{+}24, CDC^{+}23, DFP^{+}21a, GGCL22, GBE^{+}22, HOK^{+}20, KLA21, MCJ^{+}20, PGGE21, PAT^{+}24, RH23, SEJ^{+}24, SdGWN21, VK21, WAA^{+}23, WMKV21, WAA24, dDW^{+}22, dLCN^{+}22]. shipboard [BZT^{+}23]. ships [KBE^{+}22]. Shoaling [GMLC22]. shore [CRB^{+}21, CGM23, HW20, RAA^{+}23, TQG24]. shorebird [Bai22, CGS^{+}23]. shorebirds [SNH^{+}20]. shoreline [AHCT22b, SLH^{+}22]. shores [BLB23, EB21, GLB^{+}20, SB20a]. Short [RKV^{+}21, SJQ^{+}22, WMT20, BMS^{+}24, FWA^{+}23, GLW21, JCGS22, LRM^{+}21, SBL^{+}23, TNS^{+}22]. Short-finned [SJQ^{+}22, BMS^{+}24]. short-tailed [GLW21]. Short-term [RKV^{+}21, WMT20, FWA^{+}23, JCGS22, LRM^{+}21, SBL^{+}23, TNS^{+}22]. shorter [PGGE21]. shortfin [VDB^{+}24]. shortspine [DMTP22]. show [HHBP23, ICH^{+}24, MDG^{+}21]. shows [FWA^{+}23]. shrimp [BMF24, SC23, ZFD^{+}24, ZCB20, ZLHH20]. shrimps [MHÁC^{+}22, MON^{+}24]. shuffling [IBRL^{+}22]. Shy [MHAL24]. sight [TNMN24]. sighting [CBB^{+}22a]. sightings [DFHT23, FPZJAO^{+}22]. signals [LCGS^{+}21]. signature [HIT^{+}22]. signatures [CMdL^{+}20, XW24]. significant [CAW^{+}22]. signs [SKHC23]. silicon [RAHM^{+}21]. silver [CDS20]. silvertip [TCM^{+}23]. Similar [PGT^{+}21, dSGC^{+}23, BFC^{+}20, CASF21, CMdL^{+}20, HM23, VPB^{+}20]. similarity [PWvdM^{+}23]. simple [BNR20]. SIMS [HKY^{+}20]. simulated [DGP22, HOK^{+}22, NTN^{+}24, SSFD^{+}23, SRK^{+}23]. Simulating [HRC21, QRC^{+}24]. simulation [BNR20, PRC^{+}21]. simulations [BMB24]. simultaneous [DSK^{+}22]. simultaneously [Ano23]. since [RDL^{+}21]. Singapore [MSL^{+}21]. single [PSÓ^{+}24, STY^{+}20, ZCM^{+}21]. Single-Large-or-Several-Small [ZCM^{+}21]. Singularities [DBHCO^{+}22]. sink [BTMS24, LAL21, QRC^{+}24]. sinking [FCSG21, TYL^{+}24]. siphonophores [MSTH22]. Sipuncula [CGLC20]. Sirenia [WKAH22]. site [ACA^{+}24, BGMRGM^{+}20, CBTL20, CGS^{+}23, DMM^{+}24a, FPZJAO^{+}22, HNB^{+}24, MPB^{+}24, RWR24, WSPS^{+}20]. sites [BBH^{+}20, BHQ^{+}22, CCS^{+}22, CLGH24, SHM^{+}20a, WFS23]. situ [BBEF^{+}20, BB24, BMPD20, BNR20, EKRT^{+}20, GBH^{+}20, GON^{+}20, HDM^{+}24, HLCH23, Lei20, WD22, dBDL^{+}20]. Six [AMT^{+}21]. Size [ADBF21, DSB^{+}21, DFP^{+}21b, SS24, TdLHL21, BJH^{+}23, BAS21, CKT^{+}24, CH20, DFP^{+}21a, EBL^{+}21, GPMH^{+}24, GBSS24, GBG^{+}23, IBC^{+}21, KDLL21, KKH22, KF20, KSS^{+}20, LKB^{+}21, LCFJ20, LAGGM^{+}21, MAK^{+}22, MCK21, MJ21, MPB^{+}24, MDG^{+}21, NSMJ20, OFSJ20, OLG^{+}23, PBO^{+}21, RBS^{+}24, SME^{+}23, TBH21, TYL^{+}24, WTK^{+}24, ZZLS24, ZCM^{+}21]. size-based [MAK^{+}22, MDG^{+}21]. size-dependent [MJ21]. Size-fractionated [DSB^{+}21]. Size-fractioned [SS24]. size-spectra [GPMH^{+}24]. size-structured [NSMJ20]. sized [CIDM23, SSP^{+}22]. Skate [SOS^{+}21a, DHM^{+}24, FNS^{+}20, SOS^{+}21b]. skates [JLT^{+}21, SOS^{+}21a]. Skeletonema [YOI^{+}23]. skewed [NYW^{+}22]. skin [OTD^{+}24]. skuas [GDAC23]. slipper [MK20, MSDY23]. SLOSS [ZCM^{+}21]. Slough [HGK^{+}24]. slow [BNH^{+}21, HP23]. slows [Rob21]. Small [KSW23, MHK22a, SDM^{+}23, BVKF23, BWGH23, BPA21, DPP^{+}21, DGM^{+}23b, DCL21, FDS^{+}21, FNH^{+}23, GAPM24, HLCH23, HCP^{+}24, ISMM20, JMM24, LT21, MCvdM^{+}24, MIL^{+}20, OCGAP^{+}22, PRCvdK23, RSS20, RCC^{+}24, STF^{+}24, TNS^{+}22, TdLHL21, VFG24, Whi21, vLSB^{+}21, ZCM^{+}21]. Small-scale [KSW23, MHK22a, SDM^{+}23, BVKF23, BPA21, LT21]. smaller [FQVMF23]. smelt [GKK^{+}22, VMCA^{+}20]. Smith [Ano22d]. smolts [ATLT20, AGB^{+}23, MMT^{+}24, WPM22]. snail [BPA21, CBP22, HB24]. snails [BH22, JBG22, MJ21, YMU^{+}21]. snapper [BDS^{+}20, CHJ21, DSK^{+}22, ESG20, FHAP20, GBE^{+}22, PTCD22]. snook [AWC21]. snow [LHM^{+}23]. SNPs [PSÓ^{+}24]. soaring [GFK^{+}23, SWR23, vESB^{+}23]. Social [KLA21, NJHN24]. sockeye [JPJH23, WPM22]. soft [BMK^{+}22, DGM^{+}23a, HGT^{+}21, LHFD^{+}24, MHK22b]. soft-bottom [DGM^{+}23a, HGT^{+}21, LHFD^{+}24]. softsediment [GBH^{+}20]. soil [RLP^{+}23, RSL^{+}22]. Solar [KCLJ20]. solaris [DDB^{+}24]. Solomon [THD^{+}22]. soluble [ZLHH20]. solution [OCR^{+}24]. some [MAS^{+}20]. sonar [OLG^{+}23, SME^{+}23]. sooty [BLCC^{+}20, FCV^{+}24]. Sotalia [MCLC^{+}23]. Sound [BRD20, EWF^{+}21, LHM^{+}23, MMS^{+}20, YOI^{+}23, MPM21, MCH^{+}21, ABH^{+}21, DMCB22, KPK^{+}24, WFR^{+}23]. sounder [NGPGO^{+}23]. sounds [BMBC21, MPSH22, MMS^{+}22]. soundscape [MMS^{+}20, SBMW22, SFMK^{+}23]. Soundscapes [GSK^{+}20, VPB^{+}20, SLJ^{+}20, TB21]. source [BTMS24, PDM22, QRC^{+}24]. source-sink [BTMS24, QRC^{+}24]. sources [JHK^{+}22, JRVE^{+}22, SC23]. South [BKR^{+}21, BAS23, HLCH23, MMS^{+}22, PLC^{+}20, DKS^{+}24, GLM^{+}24, WCS^{+}21, AGB24, BDC^{+}22, CDS20, CCA^{+}20, CKCS20, DGM^{+}23a, FSS23, IMM^{+}20, MdOO^{+}20, MFCF21, MLP^{+}24, NYW^{+}22, OCC23, PSBH21, PSP^{+}20, RPD^{+}23, YWW^{+}24]. south-eastern [DKS^{+}24]. Southeast [MSL^{+}21, BGG^{+}22, CKCS20, JPMG23, MBS^{+}24, RSEG21, CMTP22, LKR^{+}21]. southeastern [BMOW21, GLP^{+}24, GRQ21, LTM20, LEG^{+}20, MMS^{+}20]. Southern [ASS^{+}21a, CPD^{+}20, AMUPJR^{+}24, AGB24, ASS^{+}21b, BLWJ20, CGGdD24, CBB^{+}22a, CSN^{+}23, DCN^{+}23, DMM^{+}24a, DFSH23, GAFS20, HO24, LMH^{+}24, LNM22, LSKS24, MBW^{+}20, MVH^{+}21, MPSH22, MLP^{+}24, MRH20, OMW^{+}20, ORB^{+}21, PBAC23, PGW23, RBF^{+}20, SMPLF^{+}22, SYMR21, SAH^{+}24, TABM21, TFCS22, VvDRM20, VDG24, WBGC24, ZPVN23, AMMADDH20, ASBE22, BPS^{+}24b, CC22, FST22, FLC^{+}24b, GWA^{+}23, GNC^{+}23, KRG^{+}20, LRM^{+}21, MJ21, MPB^{+}24, OTD^{+}24, QHT^{+}21, WWH^{+}22]. southwest [EMDR22, PRCvdK23, BLCC^{+}20, DIC^{+}22, GSPH20]. southwestern [APG^{+}22, ASIA22, CBB^{+}22b, CBKM20, DKSS21, GFT^{+}21, GSL^{+}24b, SFLQ23, SCM^{+}23]. sp [LNZL21, PLG^{+}23, STF^{+}24, UPK^{+}20, VKdJ24]. Space [BAAR20, DCN^{+}23, FN24, HNB^{+}24, LBR^{+}24, MJJJ20, TWC^{+}23]. Space-use [BAAR20, MJJJ20]. spaces [EMB^{+}22]. Spain [LAGGM^{+}21]. span [TKP^{+}20]. Sparisoma [VK21]. Spartina [HPH20, NCZ^{+}21, RS21, SAH^{+}24]. spat [STF^{+}24]. Spatial [AFTW24, AAHW20, BQM^{+}24, BMOW21, BBH^{+}20, CFB21, CFBPH23, FRO^{+}21, FDE^{+}21, GDCP20, GSPH20, ISMM20, JNY^{+}20, LMS^{+}20, LCGS^{+}21, MGA^{+}20, MFCF21, OCC23, OBW^{+}20, PGW23, RMR^{+}23, SYMR21, VNM^{+}21, YSD^{+}21, vdVWF^{+}24, BMO^{+}24, BGMRGM^{+}20, CDF^{+}20, Cha22, CP22b, DSB^{+}21, DCR^{+}21, DAT^{+}21, EPB23, FTA^{+}21, FSS23, Fig21, FSAG22, FP20, FKS^{+}22, HMV^{+}22, HBS^{+}21, IBC^{+}21, JMR21, KDB^{+}24, KMW^{+}21, LO20, LBMB^{+}20, MCF^{+}21, MSR^{+}21b, MFC^{+}24, MMB22, MTSO^{+}24, dACMRZ^{+}20, MBR^{+}21, MMS^{+}20, PBS22, PRM^{+}21, PFM^{+}21, PWvdM^{+}23, RB22, RBB^{+}24, RBCAV22, SBMW22, SJS^{+}23, SHS21, SWC^{+}23b, THR^{+}23, TMN^{+}24, WSB^{+}24, WYZ^{+}20, WFS23, vdWDH^{+}21, vdVVC^{+}21]. spatial-seasonal [vdWDH^{+}21]. spatially [KKC^{+}24, SLJ^{+}20]. Spatio [FVC^{+}20, GSBO21, HW20, KPW^{+}20, LNM22, NRE21, TTJS20, WXZ^{+}24, BAB^{+}24, CVP^{+}21, FLY^{+}24, VH20]. Spatio-seasonal [WXZ^{+}24]. Spatio-temporal [FVC^{+}20, GSBO21, HW20, KPW^{+}20, LNM22, NRE21, TTJS20, BAB^{+}24, CVP^{+}21, FLY^{+}24, VH20]. Spatiotemporal [AKW^{+}22, ABWJ^{+}20, BMF24, BRSD20, GRTK^{+}21, MAPFH22, MLB^{+}21, RMD24, TABM21, TRK^{+}22, AFP^{+}24, Ano23, ACG22, FFMR24, FWGD^{+}23, FYK^{+}20, KTK21, LFFW21, PCLQ^{+}20]. spawn [HGH^{+}24, RBC^{+}24]. spawned [BSF24]. spawners [HHO^{+}23]. Spawning [JHK^{+}22, PSB^{+}23, SCM21, CDD^{+}20a, DMCB22, DDB^{+}24, FN24, FSK^{+}21, FYK^{+}20, GBE^{+}22, HNKK21, HNB^{+}24, HIT^{+}22, HMK^{+}22, MRB^{+}24, MY21, SFL23, SBRQ21, THK^{+}22, TTJS20, TNS^{+}22, VAMF22, WSPS^{+}20, vdVWF^{+}24]. Spawning-related [PSB^{+}23]. specialisation [CSN^{+}23, DMBJ21]. specialists [GLM^{+}24]. specialization [LNBdB21, MSV21]. Species [DCR^{+}21, PCM^{+}20, WAA24, YOI^{+}23, AMMADDH20, AM20, AMT^{+}21, ABH^{+}21, Ano24o, APH^{+}23, ABWJ^{+}20, BS20a, BGG^{+}22, BML^{+}22, BRD20, BLGM^{+}24, BTB^{+}24, BBMM20, BAH^{+}22, BWC^{+}22, CCAdA24, CGM^{+}24, CEG^{+}21, CCMRII22, CMdL^{+}20, CTH^{+}20, CFB21, CWC^{+}22, CBTL20, CMTB^{+}22, CKS^{+}24, DJR^{+}20, DPCL^{+}23, DKAB23, DL21, DBP^{+}22, DDG^{+}21, DCL21, EB21, EPB23, FS22, FST22, FNS^{+}20, GGN21, GRB22, GNC^{+}23, GPCM22, HOP^{+}20, HSNT24, HKD^{+}20, IBC^{+}21, JDM^{+}20, JHK^{+}22, JHDE20, JDDF21, JPMG23, JH20, KHL^{+}22, KKC^{+}24, KKH22, KKS^{+}20, KDGG24, LPO^{+}21, LvAH^{+}24, LR22, LORB^{+}24, MDW^{+}22, MBB^{+}22, MJW24, MFC^{+}24, MON^{+}24, MGBCGM^{+}22, MCH^{+}21, MLP^{+}24, MSV21, NCZ^{+}21, OBS^{+}22, OFSJ20, OHBJ20, OH22, OCR^{+}24, PCLQ^{+}20, PWvdV20, PDH^{+}24, PSP^{+}20, RC23, RDL^{+}22, RBM22, SSFD^{+}23, SBT24, SME^{+}23, SKHC23, SOS^{+}21b, SVS^{+}22, SdGWN21]. species [SPFL20, THD^{+}22, TKS^{+}20, TPGG^{+}23, TSW^{+}21, TKP^{+}20, TCM^{+}20, VLPN24, WFR^{+}23, WD22, Whi20, XWH^{+}22, ZHG^{+}21, ZXL^{+}24]. species-area [SPFL20]. Species-specific [PCM^{+}20, YOI^{+}23, SSFD^{+}23]. specific [BBR24, DSA^{+}24, FWWH22, GSBO21, HMT21, HPFB20, JWH^{+}21, KdlVJ^{+}21, MDM^{+}20, PCM^{+}20, RB23, RL22, SSFD^{+}23, SSP^{+}22, SHL^{+}21, XCT^{+}20, YOI^{+}23]. spectra [GPMH^{+}24, KF20, KSS^{+}20, OLG^{+}23]. spectral [GSN^{+}20, VPB^{+}20]. spectrum [MSC^{+}22]. speed [Lei20]. sperm [BBPO21, DFHT23, GWDR20, PAL^{+}22, PVB^{+}20, SBFP^{+}24, WMT20]. Spheniscus [DRP^{+}20]. Sphyrna [EMGMEV^{+}21, GLP^{+}24]. spill [MTB^{+}23]. spills [MLAS22]. spined [DKS^{+}24]. spinner [SON^{+}20]. spiny [ADBF21, BBRM22, TdLHL21]. Spionidae [YWW^{+}24]. Spitsbergen [DB21]. sponge [AGA^{+}22, BKM^{+}21, GMCE^{+}22, HDB21, HDM^{+}24, HvdSR^{+}21, LSC^{+}21, MDH^{+}21, RHH^{+}20, RL24, SRG^{+}23]. sponges [BBKW20, MPA^{+}22, OCFEB21, RBM^{+}21, SMS^{+}24, VOM^{+}20]. sporophyte [MBB^{+}22]. spot [MTK^{+}21]. Spotted [YXB22]. spp [RL22, SCM21, SC23, TGB^{+}20, WDP^{+}21, WF21, YMU^{+}21]. sprat [AU21, AFP^{+}24]. spring [BSF24, BBGM23, DKSS21, GGRV23, MCL24, NMB^{+}22]. spring- [BSF24]. Squalus [ASY^{+}24]. squid [ASIA22, CZS^{+}22, JJV^{+}21, JBGA23, WWH^{+}22, vTLG^{+}21]. squids [PS23, QHT^{+}21]. St [FNL^{+}23]. St. [AEHD21, BTMS24, RH20]. stability [AHH^{+}20, GSL^{+}24a, VGG^{+}23]. Stable [CAPS20, FNL^{+}23, GWDR20, HMW^{+}21, LPD^{+}24, PVB^{+}20, RVSA^{+}21, RDL^{+}21, SDW20, APOGG20, BPS^{+}24b, CVG^{+}23, CKM^{+}20, CCA^{+}21, DCR^{+}22, DSK^{+}22, DSA^{+}24, DFSH23, EVVMQ^{+}23, FS21, FWWH22, GFT^{+}21, GVAZ^{+}22, GSL20, GPP^{+}24, GB20, GDAC23, HRH22, JBGA23, KGG^{+}21, LKR^{+}21, LH22, LWP^{+}24, LNBdB21, MBW^{+}20, OCR^{+}24, OTD^{+}24, PBB^{+}20, PSN^{+}20, PLC^{+}20, PWvdM^{+}21, PWvdM^{+}23, SDM^{+}22, SBS^{+}24, SLB^{+}21, VMDA22, VNG^{+}22, WYO^{+}20, WWH^{+}22, XW24, ZMC^{+}22, dLCN^{+}22, vTLG^{+}21]. stable-isotope [CKM^{+}20]. stage [CGW21, CCR^{+}20, FKS^{+}22, LBH^{+}21b, MHH^{+}20, MLB^{+}21, SSK^{+}23, WMKV21, ZXL^{+}24]. stages [AFP^{+}24, AAHW20, CRÁ^{+}20, CSK^{+}22, CDS23, OÁBP^{+}21, RHSM21, RGK22, SRS^{+}20]. staghorn [KDLL21]. staging [OvBS^{+}24]. stalked [NDA^{+}22]. standardised [OWH23]. stanzas [LEG^{+}20]. star [CLL^{+}22, Hew21, SKHC23, TSS^{+}23]. starfish [CSDB22, DML^{+}21, DDB^{+}24]. stars [BLA^{+}24, LKB^{+}21]. startle [HJG21]. startle-eliciting [HJG21]. starvation [TEK21]. starving [RBB^{+}21]. state [AAP^{+}23, CDD^{+}20b, EBL^{+}21, GSL20, SS24]. states [BKTN21, TTA20]. static [FTA^{+}22]. stationary [BDJ^{+}24]. stations [WYZ^{+}20]. Statistical [BGMRGM^{+}20, Ano23]. status [CCBD^{+}21, MMMV22, SGW^{+}21, XMY^{+}22, ZBB^{+}23]. stealing [HOP^{+}20]. steelhead [MBGM21, MMT^{+}24]. Stegnaster [ZDBS20]. Steller [BBR24, MSF24]. stelleri [MB20]. Stenella [SON^{+}20]. step [GRB22]. steps [SDE^{+}21]. stereo [BWGH23, JS20, SAMdL^{+}20]. stereo-video [JS20]. Stereolepis [BSRVS22]. Sterna [RB22]. stimulated [DL20]. stimulates [EKRT^{+}20]. stimuli [HJG21]. stimulus [KKC^{+}24, MY21]. stingray [OVO^{+}21]. stingrays [ECvBL21, MPSH22, OMW^{+}20]. stipes [KMW^{+}21]. stochastic [AS24]. Stock [PBO^{+}21, AFTW24, CAP^{+}21, DDG^{+}21, HKK^{+}23, HESF24, NTN^{+}24, SFL23]. stocks [AZ24, CMvKdR23, JPJH23, WLB23]. Stomach [PWvdM^{+}23, RRS^{+}23, DSK^{+}22, VMDA22, ZMC^{+}22]. stomachs [WWH^{+}22]. Stomolophus [FSS23]. stone [SBS20]. stopover [DME^{+}23]. storage [CHJ21, LAP^{+}24]. storm [BGMAM^{+}21, GSL20, HBE22, HPH20]. storm-driven [GSL20]. storm-petrel [HBE22]. storm-petrels [BGMAM^{+}21]. storms [NBL^{+}22]. stormy [MY21]. story [GCGCAS^{+}20]. strain [FLCQ21]. Strait [HSA^{+}21, MSL^{+}21, AU21, DFHT23, JNY^{+}20, MCL24, MMMV22, ZZLS24]. stranded [SBS^{+}24]. stranding [GSL^{+}24b, JPB21]. strandings [PAL^{+}22]. strata [KKN^{+}20]. Strategies [CAAG^{+}22, AEHD21, BMK^{+}22, EGK^{+}21, HEKH24, LNZL21, MPA^{+}22, MDW^{+}22, SJQ^{+}22, SON^{+}20, WH24]. strategy [ANP^{+}24, AMT^{+}21, BSH21a, BAAR20, DVMH^{+}24, GFK^{+}23, GJK^{+}24, LPD^{+}24, PLG^{+}23, RHN^{+}20, SGB^{+}21]. stream [JFR^{+}21, WRG^{+}24]. strength [BFC^{+}20, FBC^{+}24b]. strengthening [NSPH20]. Stress [DBS^{+}20, ABB^{+}20, AR21a, BB24, BBM^{+}24, BICQG23, CTTLCG^{+}24, CLD23, CCC^{+}23, HBC^{+}20, ICIS23, MHAL24, MG22, NSM24, ODB^{+}24, OA24, RVV^{+}24, STO^{+}23, TD21, VSG20]. Stress-on-stress [DBS^{+}20]. stressed [COWM^{+}22a]. stresses [MDM^{+}20]. stressful [WGK22]. stressor [MGH21]. stressors [BPA21, CDC21, CSH^{+}21b, PDFH20, RGK22, SW23a]. strike [SHL^{+}21]. striped [RBF^{+}22]. Strong [MDY^{+}21, BPA21, HHBP23, LDW^{+}21, RCB^{+}24a]. strongest [KDB^{+}24]. strongly [BS20a, IHA^{+}22]. Strongylocentrotus [DMM^{+}24b, KNH24]. Structural [RBFA20, CFHCS^{+}24, MSPS22]. Structure [CCR^{+}20, AKW^{+}22, ABH^{+}21, BQM^{+}24, BNSH21, BNS^{+}24, CAP^{+}21, CIDM23, CSK^{+}23, CC22, CDC21, CH20, CSD^{+}24, DL20, DFB^{+}24, DMBB22, EHLM^{+}23, Fig21, FST22, FP20, GBW^{+}20, GRQ21, GSPH20, HSSN^{+}21, JMR21, KMW^{+}21, KLFP23, KI24, LS21, LGGR^{+}22, MGHH21, MBS^{+}24, MHÁC^{+}22, MSV21, MDQ^{+}21, NMAE^{+}22, OJPC23, PBO^{+}21, PWvdM^{+}21, PWvdM^{+}23, PRD^{+}20, RSvL21, RSEG21, RBF^{+}20, RBM22, SRS^{+}20, SPFL20, TMN^{+}24, WWH^{+}22, YGMNA22]. structured [NSMJ20]. structures [MAPFH22, TYL^{+}24, YAZA21]. structuring [ASS^{+}23, CGL^{+}22, WSB^{+}24]. studies [AUS^{+}23, FDL^{+}21, KDS^{+}20, OWH23, PCP^{+}20]. study [BBG^{+}22, CTH^{+}20, CHDG^{+}23, GFDN^{+}24, HHO^{+}23, HRC21, KTK21, LFK^{+}21, LNKJ^{+}22, MJ21, NBL^{+}22, PMH23, PCLQ^{+}20, RJ21, SJ22]. studying [CBB^{+}22b]. stutchburyi [FSF^{+}24]. Styracura [OVO^{+}21]. Sub [SW23a, CAW^{+}22, CLM^{+}21, CLC^{+}23, Fai21, FHAP20, LMRS^{+}24, SBJ^{+}24, SFO^{+}21, UBO^{+}24, URSS24, VGG^{+}23, vTLG^{+}21, MSL^{+}24]. sub-adult [SBJ^{+}24]. sub-Antarctic [VGG^{+}23, vTLG^{+}21, MSL^{+}24]. sub-arctic [SFO^{+}21, CLC^{+}23, LMRS^{+}24, UBO^{+}24, URSS24]. sub-colony [CAW^{+}22]. Sub-lethal [SW23a]. sub-tropical [CLM^{+}21, Fai21, FHAP20]. subantarctic [MCC^{+}21]. subarctic [BK23, BMF24, BG20, BG21, GBG^{+}23, GFH^{+}22, LRD^{+}20, MMD20, NSB^{+}23, WDP^{+}21, ZWXC21]. Subcuticular [MMLPP^{+}24]. sublethal [SRK^{+}23]. submarine [HOK^{+}22, SPFL20]. Submergence [HPH20]. subpopulations [FWA^{+}23]. subregions [SCB^{+}20]. subsidies [GPMH^{+}24]. subsidized [MJ23]. subsidy [MVH^{+}21]. substantial [BHQ^{+}22, MBGM21]. substrate [BS20a, OJ24]. substratum [JMS^{+}22, MJ23]. subsurface [HOK^{+}22]. subtidal [ARB^{+}20, MW20, MMP^{+}22, SKI^{+}20, TKP^{+}20]. Subtropical [DL20, AESS24, CFFC^{+}23, CMF21, FWGD^{+}23, GFH^{+}22, GSRC20, GLM^{+}24, LBH^{+}21b, MdOO^{+}20, MFCF21]. success [ADF20, AJH22, GLW21, GJM^{+}20, LPB^{+}23, MBB^{+}22, MTB21, OBW^{+}20, PGW23, RWH^{+}24, SHH^{+}22, STF^{+}24, TBS20, WHRC22, WHN^{+}22]. Successful [PSBH21, GRB22, PGT^{+}21]. succession [OYO^{+}20, SKI^{+}20, dBDL^{+}20]. suction [WMT20]. Suess [SDW20]. suggest [ASL^{+}22, FCV^{+}24, MDK^{+}21, PVB^{+}20]. suggested [MPR^{+}23]. suggests [GFK^{+}23, GDAC23, RDL^{+}21, SOS^{+}21b]. suitability [DFHT23, HBG^{+}22, LFFW21, RGL24]. suitable [GBE^{+}22, RTG^{+}20, SRG^{+}24, SHT^{+}22]. sulfide [WIY^{+}23, dFRvdG^{+}22]. sulfide-detoxification [dFRvdG^{+}22]. Sulfitobacter [YHW^{+}23]. Sulfur [GVAZ^{+}22, SB20b, TD21]. sulfur-based [TD21]. Sulu [BRC22]. Summer [BLW^{+}24, AMMADDH20, FKTK20, GBSS24, GEY^{+}21, SST^{+}20a, STO^{+}23, XMY^{+}22]. summit [GIDP^{+}21]. sun [BVKF23]. Sundays [BAS23]. sunfish [ABWJ^{+}20, NBW^{+}23, PSN^{+}20]. superba [MSR^{+}21b, SBSS21, WRW20, ZZLS24]. supernumerary [BLA^{+}24]. supply [BWC^{+}22, FDF20, GWC^{+}21, GMB^{+}24]. support [BML^{+}22, CCS^{+}22, SOS^{+}21a]. supported [GSL20]. supports [HTL^{+}20, KGG^{+}21, WSU21]. supratidal [BNSH21, MGBCGM^{+}22]. surf [ECC^{+}21]. surface [DSB^{+}21, DBHCO^{+}22, GEY^{+}21, KPS^{+}22, KDS^{+}20, OÁBP^{+}21, RYF^{+}24, Sig23]. surface-feeding [Sig23]. surfaces [KPS^{+}22]. surmuletus [PCLQ^{+}20]. surprise [ŠGH^{+}21]. surrogates [TGT^{+}23]. survey [BZT^{+}23, BBKW20, PAF^{+}22, PS23, WCL20, WCP^{+}21a]. surveying [DPDJ^{+}21]. surveys [AKGA^{+}21, CASF21, DFHT23, SAMdL^{+}20]. Survival [AGB^{+}23, DVMH^{+}24, ABN^{+}21, BLE^{+}22, DMS^{+}21, DFCM21, FSDB20, GBG^{+}23, HEKH24, KETL22, LSSG24, LMB^{+}20, LBR^{+}24, MBGM21, NMB^{+}22, NBL^{+}22, RHM^{+}22, RSS20, RLME^{+}21, SLH^{+}22, SdFBM24, STF^{+}24, WTK^{+}24]. survive [MDB^{+}21]. survivorship [HBR^{+}21, LDPH24]. susceptibility [BTDR20, MRB^{+}24, MGM^{+}24]. suspended [BVD^{+}20]. suspension [SHZ^{+}20]. suspension-feeding [SHZ^{+}20]. sustainability [PYB^{+}21]. Sustainable [SGW^{+}21]. sustained [TCT^{+}23]. sustains [SMPLF^{+}22]. Svalbard [BLW^{+}24, SJH^{+}22, VGLL^{+}21]. SW [BVKF23, FFMR24, GLB^{+}20]. Swimming [CZS^{+}22, WDP^{+}21, KPS^{+}22, Lei20]. switching [BMKB22, BBMM20]. swordfish [APOGG20, HTT^{+}21, SSB^{+}23]. Sydney [GSP^{+}20]. Symbiodiniaceae [DPL^{+}21, WKR21, WNH^{+}20]. symbiont [DDR^{+}21, HvdSR^{+}21, IBRL^{+}22, NSM24]. symbionts [LCJ22]. symbiosis [MGH21, SMA^{+}20]. symbiotic [DDR^{+}21, ICIS23, LCJ22, XMY^{+}22]. sympagic [SNP20]. sympatric [BGMAM^{+}21, BBG^{+}22, CAAG^{+}22, DFP^{+}21b, ECvBL21, GGN21, HOK^{+}20, MDY^{+}21, PAT^{+}24, PDH^{+}24]. sympatrically [DCK^{+}24]. Sympatry [FJJ^{+}21]. synchronize [DMS^{+}21]. synchronous [GGT^{+}20]. synchrony [ACG22, OBW^{+}20, PSB^{+}23, RLME^{+}21]. syndrome [TSS^{+}23]. synecology [BMK^{+}22]. Synergistic [MVSV22]. syngnathids [PCP^{+}20]. synthesis [RPB21]. System [OH22, BRC22, BMW24, CRB^{+}21, CCJ^{+}23, CKM^{+}20, FSN^{+}20, FLY^{+}24, GSC^{+}22, GSL20, MCvdM^{+}24, PLH^{+}21, SMK23, TYL^{+}24, UBO^{+}24, AMMADDH20, CAAG^{+}22, DSHF22, DHF24, SBDM^{+}23]. Systems [DFSH23].

tītī [FNM^{+}21]. tactics [CGGdD24, FNO^{+}23, WHN^{+}22, dDW^{+}22]. tag [MJ21]. tagged [HDA^{+}22]. tagging [Ano23, GCDP21, MGB^{+}22]. Tags [SWHE20, EMDR22, FKTK20, TCM^{+}23, WMT20]. tailed [EVAB^{+}23, GLW21, KKMW23, WdGR^{+}20]. Taiwan [CDC^{+}23, HMK^{+}22]. take [ATR^{+}24, DVP^{+}23, LCR^{+}22, MDPL20, PSB^{+}23, VNM^{+}21]. Taliepus [BSF^{+}20]. Tampa [HKD^{+}20]. Tanzanian [HMW^{+}21]. target [BBMM20]. targeted [CCA^{+}20]. tarpon [GBA^{+}22]. tau [MPM21]. tautog [TPGG^{+}23]. taxa [MDM^{+}20]. taxa-specific [MDM^{+}20]. Taxonomic [FSSL^{+}20, NMAE^{+}22, TGT^{+}23, AS24, GSN^{+}20, WDJ20]. technique [GSN^{+}20, SCW^{+}24]. techniques [CGS^{+}23, PAF^{+}22]. technologies [BML^{+}22, CGM^{+}24]. technology [Ano23]. Tedania [LSC^{+}21]. teeth [IMYW24, KdlVJ^{+}21, dLCN^{+}22]. telemetry [AUS^{+}23, BBH^{+}20, CBTL20, CKKL21, DMM^{+}24a, ESG20, FDL^{+}21, GMS^{+}21, LMS^{+}20, LCR^{+}22, MDW^{+}22, MJJJ20, NBW^{+}23, OLA^{+}24, SHS21, ZSWR^{+}22]. telemetry-based [FDL^{+}21]. teleost [CWD^{+}22]. Teleostei [DMBJ21]. teleosts [BNH^{+}21, DAB^{+}21]. tell [FSDB20]. temperate [ARB^{+}20, BLB23, BS20b, CRB^{+}21, CRG^{+}20, DCN^{+}23, DKAB23, DDR^{+}21, Fai21, FS22, FSSL^{+}20, GS20, HDB21, HSNT24, HO23, HBS^{+}21, HDE21, HTL^{+}20, KWD21, LBH^{+}21b, MHAL24, MMP^{+}22, PSBH21, PBRF23, PWvdV20, QMDGM22, RH20, Rus20, RBCwD^{+}20, SKTO23, VPB^{+}20, VHG20]. Temperature [DMM^{+}24b, FCSG21, PLG^{+}23, VAMF22, ASPM21, BS20a, BSF^{+}20, BMM21, BAS21, BP24, CGCC24, DVMH^{+}24, DKSS20, DDG^{+}21, Fai21, FBC^{+}24c, FMD20, GBW^{+}20, GM20, GED20, GPCM22, HKY^{+}20, HYG^{+}20, HWA^{+}22, KNO^{+}21, LHD20, LRM^{+}21, LNZL21, LCFJ20, LLL20, LHW^{+}20, MLE^{+}22, MHN24, MHH^{+}20, MVSV22, OLM24, RH21, RC23, RGK22, SBSS21, ŠGH^{+}21, SJ22, SBC^{+}21, SKHC23, THS^{+}24, WKR21, YRB^{+}20, dFRvdG^{+}22, BB24]. Temperature-dependent [DMM^{+}24b, SJ22]. temperature-enhanced [LHD20]. temperature-growth [FMD20]. Temperature-light [PLG^{+}23]. temperatures [DL21, DA24, ENRC^{+}20, GEY^{+}21, IBRL^{+}22, KDGG24, MHF^{+}21]. Temporal [ARN^{+}22, BBPO21, CAV^{+}20, FSS23, HBE22, KDK^{+}24, MPB^{+}24, NHH^{+}21, dBDL^{+}20, AS24, BTA^{+}20, BAB^{+}24, BGMRGM^{+}20, CVP^{+}21, CATF^{+}21, CRB^{+}21, Cha22, CFB21, CFBPH23, FVC^{+}20, FLY^{+}24, GSBO21, HW20, KPW^{+}20, LO20, LNM22, LNBdB21, LCGS^{+}21, MGA^{+}20, NRE21, PBS22, RBW^{+}20, SYMR21, SWC^{+}23b, TTJS20, TMN^{+}24, VPB^{+}20, VH20, WMKV21, WDJ20, WYZ^{+}20, vdVWF^{+}24]. temporally [DIC^{+}22]. tenuicaudatus [DCN^{+}23]. tenuis [STF^{+}24]. term [BSF24, CEZH21, CBB^{+}22a, DV20, DDD^{+}23, FWA^{+}23, GSL^{+}24b, HP23, HDM^{+}24, HEP^{+}24, JCGS22, JPB21, LAP^{+}24, LBTE^{+}21, LRM^{+}21, MSH20, MBR^{+}21, NHFS22, NTN^{+}24, RKV^{+}21, SBL^{+}23, SHT^{+}22, SFO^{+}21, TNS^{+}22, TGT^{+}23, WMT20, dLCN^{+}22]. terminating [MKS^{+}20]. terms [LNKJ^{+}22]. tern [EVVMQ^{+}23, RB22]. terns [FTA^{+}22, WLA^{+}21]. terrain [EPHS23]. terrestrial [CSDB22, CMO^{+}20]. Test [KDS^{+}20, GCDP21, MJK24]. Testing [ABH^{+}21, DDR^{+}21, KKH22, LBMB^{+}20, RPD^{+}23, SSM20]. tests [WIY^{+}23]. testudinatus [CGCC24]. testudinum [PDFH20, TD21]. tethering [DKAB23]. tetracycline [KSMP20]. tetrapods [TABM21]. tetricus [RRS^{+}23]. Texas [LBMF^{+}24]. Thailand [MYSF23]. Thalassarche [MHAL24, SSC^{+}23]. Thalasseus [EVVMQ^{+}23]. Thalassia [PDFH20, TD21]. Thatje [Ano21i]. Thecostraca [KZC23]. their [AMT^{+}21, Ano24o, BBSM^{+}20, CGLC20, CGGdD24, CHDG^{+}23, CHD^{+}22, CKP20, FSDB20, FLY^{+}24, FOM^{+}20, HWT^{+}22, JWH^{+}21, JMM24, KKMW23, KHG^{+}21, LDM^{+}20, MCF^{+}21, MMLPP^{+}24, MDB^{+}21, MCC^{+}21, QHT^{+}21, SCM21, SJQ^{+}22, SPFL20, TYL^{+}24, VHG20, WRB21]. Theme [Ano23, Ano24o, CRÁ^{+}20, HvLM21, RBCwD^{+}20]. theoretical [Ano23]. theory [OAdJA^{+}20]. Thermal [ASY^{+}24, GFK^{+}23, GJK^{+}24, LDPH24, MGBCGM^{+}22, TBS20, vESB^{+}23, AWB23, BMM21, CTTLCG^{+}24, CBGB24, DGP22, FN24, FNS^{+}20, FSK^{+}21, HBC^{+}20, JH20, LPB^{+}23, MDM^{+}20, NT24, RH21, RWM^{+}22, YRB^{+}20]. Thermally [BSP^{+}22, RBVA23]. thermoregulator [MB20]. thick [FDE^{+}21]. thick-billed [FDE^{+}21]. thin [BNR20, DCK^{+}24]. thin-billed [DCK^{+}24]. thorns [CSDB22, DML^{+}21, DDB^{+}24]. thornyhead [DMTP22]. thread [RC23]. threadfin [XWH^{+}22]. threat [PDM22]. threatened [ABH^{+}21, DVR^{+}23, EVH^{+}21, HO24, JLT^{+}21, LCJ22, YRB^{+}20]. Three [OCFEB21, AOO^{+}21, BMKB22, BGMAM^{+}21, Cha22, CWC^{+}22, FMD20, GRB22, HMW^{+}21, ISMM20, JPMG23, dACMRZ^{+}20, PAT^{+}24, SVS^{+}22, SWHE20, VIV^{+}20, VLPN24, YGMNA22]. thresher [KKS^{+}20]. threshold [BBSM^{+}20, BB24, LDPH24]. thresholds [MG22, RRR^{+}23]. throughout [BMF24, GGCL22, MVSV22, NM22, QHT^{+}21]. Thunnus [AAFC^{+}21, AOO^{+}21, BKH^{+}24, FLC^{+}24a, FSK^{+}21, HKY^{+}20, SKTO23, VMDA22]. thynnus [BKH^{+}24, VMDA22]. Tidal [CCJ^{+}23, JFR^{+}21, NCZ^{+}21, dlBSL^{+}22, BD20, GB21, HSBB^{+}21, LBR^{+}24, MCvdM^{+}24, PBC^{+}22, SLH^{+}22, TFCS22, WLB23]. tidally [SSBA^{+}20]. tide [NT24]. tidepool [FS22]. tides [CGCC24, SSBA^{+}20]. tiered [TWC22]. tiger [VOM^{+}20]. Time [LEG^{+}20, NSM24, AAP^{+}23, BSFM20, BGL^{+}21, BMB24, DGM^{+}23a, FN24, KBP22, LÁGLL20, MBB^{+}23, NSB^{+}23, TTA20, VIV^{+}20]. time-of-day [BSFM20]. Time-series [LEG^{+}20, NSM24, VIV^{+}20]. times [RBB^{+}21, SFOPL21, YAZA21]. Timing [AJH22, Bai22, DKSS21, HGH^{+}24, JPJH23, RWH^{+}24, RBC^{+}24, RCC^{+}22, TWC^{+}23]. Timor [BBGM23, RKV^{+}21]. tip [AGB24]. tissue [LNBdB21, PSN^{+}20]. TMEs [HDB21]. toadfish [MPM21]. together [Ano23, Ano24o, DPDJ^{+}21]. Tohoku [AAI^{+}24]. tolerance [CCMRII22, DMM^{+}24b, JBG22, MG22, MGBCGM^{+}22, YRB^{+}20]. tonsa [LSSG24]. tool [NJHN24, RSS20]. tools [KdlVJ^{+}21]. toothed [CFBPH23]. toothfish [FPG^{+}21, LSR^{+}24]. top [CLV^{+}20, CAY24, CSD^{+}24, DFP^{+}21a, GIDP^{+}21, HGR^{+}22, MSH20, MSF24, SHM^{+}20b, SAH^{+}24]. top-down [CAY24, MSF24, SAH^{+}24]. topics [Ano24o, HvLM21]. topographic [ADF20]. topography [LMHS21]. Total [SSG^{+}22, MJW24, SLB^{+}21]. tourism [CWD^{+}22, CCA^{+}20]. toxic [CP22a]. toxin [LSSG24]. Trace [XW24]. tracer [SCW^{+}24]. Trachurus [THS^{+}24]. track [BBH^{+}20, HMW^{+}21]. tracker [DGM^{+}23b]. Tracking [CDS20, FDL^{+}21, MHDJ21, SEJ^{+}24, SDE^{+}21, BZT^{+}23, BASM^{+}24, BBG^{+}22, DCR^{+}22, FLBP^{+}21, GKK^{+}22, LBTE^{+}21, OWH23, SMH^{+}20, TWC^{+}23]. tracks [BSÁG^{+}22]. Tract [HHO^{+}23]. Trade [HSO^{+}22, BQM^{+}22, CPH20, FSK^{+}21]. trade-off [FSK^{+}21]. Trade-offs [HSO^{+}22, BQM^{+}22, CPH20]. Tradeoffs [ROMB22, HSBB^{+}21]. trait [BBS^{+}21, LSKS24, WXZ^{+}24]. trait-based [BBS^{+}21]. Traits [SLH^{+}22, AOO^{+}21, BQM^{+}22, BBD^{+}22, CAV^{+}20, DJR^{+}20, EAS20, FSDB20, FP20, LK21, MDG^{+}21, MA20, SFLQ23]. trajectories [ARH^{+}20, THS^{+}24]. trajectory [TGT^{+}23]. trans [ZBB^{+}20]. trans-hemispheric [ZBB^{+}20]. transboundary [DVR^{+}23, SEJ^{+}24]. Transcriptome [JPJ^{+}22, LBB^{+}21]. Transcriptome-wide [LBB^{+}21]. transect [LHD20]. transforming [COGGH20]. transforms [Rob21]. Transgenerational [LLL20, KNH24, TBS20]. transhemispheric [PHE^{+}20]. transient [CMO^{+}20]. transition [DML^{+}21, Fai21, SFLQ23]. transitional [AGB^{+}23]. transmitters [RS20b]. transparency [KBE^{+}22]. transparent [WSU21]. transplantation [LOS^{+}22]. transplanted [CLGH24]. transplanting [BJH^{+}23]. Transpolar [EGK^{+}21]. transport [BLVL24, HNN^{+}24, OAM^{+}24, SCW^{+}24]. trap [CASF21, DCR^{+}21]. trap-camera [DCR^{+}21]. traps [BS20a, BGG^{+}22, DDD^{+}23]. trawl [BBKW20, YSD^{+}21]. trawling [BBS^{+}21, DNJCH21, MDG^{+}21, vDTS^{+}22]. trawling-induced [BBS^{+}21]. Trebouxiophyceae [PLG^{+}23]. tree [CMF21, GCCM20]. Trematode [SdFBM24, BKT^{+}22]. trend [WCL20]. trends [BTB^{+}24, FBP^{+}22, FLC^{+}24b, GSS22, HBE22, KBE^{+}22, NBDM^{+}24, RSvL21, SW23b, TTA20, TRK^{+}22, TOMDC23]. trevally [DMM^{+}24a, GDCP20]. tri [ZLHH20]. triacantha [LRM^{+}21]. Triakidae [DIC^{+}22]. Trialling [SBC^{+}21]. triaxial [SSF^{+}22]. tributary [HCS^{+}24]. Trichechus [CHD^{+}22]. Tridacna [SCK23]. tridactyla [EGK^{+}21, MIL^{+}20, OBW^{+}20]. trip [MTB21]. trips [BDC^{+}22, KKMW23, RBW^{+}20]. tritonis [SMH^{+}20]. trolling [CLM^{+}21]. Trophic [BSÁG^{+}22, BDS^{+}20, CCRP20, CGS^{+}23, CLL^{+}22, DFSH23, EB20, GLP^{+}24, GSL^{+}24a, GFT^{+}21, KZC23, MdOO^{+}20, MAOR^{+}21, OFSJ20, OCGAP^{+}22, PBRF23, PWvdM^{+}21, PDH^{+}24, RBF^{+}20, SBS^{+}24, WWH^{+}22, XCT^{+}20, BCC20, BBEF^{+}20, BMBC21, BSRVS22, BKH^{+}24, BPS^{+}24b, CCBT^{+}20, CSH^{+}21a, CC22, CBB^{+}22b, CSN^{+}23, CMTB^{+}22, DL20, DSK^{+}22, DPL^{+}21, DRP^{+}20, DSA^{+}24, FS21, FWWH22, GGN21, GQBB^{+}24, GAPM24, GDCP20, HSNT24, HPRP21, HW20, JDDF21, KdlVJ^{+}21, KSW23, LPJ^{+}20, LKB^{+}21, LPD^{+}24, MPR^{+}23, MPA^{+}22, MSL^{+}24, MDS^{+}23, MCJ^{+}20, MRH20, OTD^{+}24, PBKG22, PSN^{+}20, QHT^{+}21, RSEG21, RDL^{+}21, RS20a, RCB^{+}24a, SBS^{+}22, THL^{+}20, TPGG^{+}23, WFR^{+}23, XMY^{+}22, ZHG^{+}21, ZBB^{+}23]. trophic-mediated [THL^{+}20]. Trophoscapes [WSB^{+}24]. Tropical [CGLC20, CRLC21, EMGMEV^{+}21, LGGR^{+}22, CATF^{+}21, CLM^{+}21, Fai21, FHAP20, FSSL^{+}20, GQBB^{+}24, HPH20, HDE21, HTL^{+}20, JDM^{+}20, LSC^{+}21, MSL^{+}21, MA20, MdMR24, NGPGO^{+}23, NGRH20, PAAEV^{+}24, RKV^{+}21, SKTO23, SCM21, SC23, TWC^{+}23, WLB23, WC21, dSGC^{+}23]. Tropicalization [LOF^{+}24]. tropicbirds [EVAB^{+}23, GFK^{+}23]. Trough [BBGM23]. trout [AGB^{+}23, EBL^{+}21, LCP^{+}23, SLBK^{+}20]. truncatus [CCA^{+}20, TPZ^{+}21]. truth [SJA^{+}21]. tshawytscha [WDJ20]. tuaki [FSF^{+}24]. tube [PFM^{+}21]. tube-building [PFM^{+}21]. tuberculosa [WKR21]. tubeworm [KPS^{+}22]. tubulat [AGA^{+}22]. tubulosa [MMLPP^{+}24]. tuna [AAFC^{+}21, AOO^{+}21, BKH^{+}24, DBHCO^{+}22, DDD^{+}23, FLC^{+}24a, FSK^{+}21, HKY^{+}20, HIT^{+}22, MMMV22, SKTO23, TKS^{+}20, UIS^{+}24, VMDA22]. tunas [COT^{+}21, NGPGO^{+}23]. turban [MJ21]. Turbidity [RS20a, BS20b, HTGG22, HCS^{+}24]. turf [FDF20]. turf- [FDF20]. Turkey [CCS^{+}22]. turnover [CSH^{+}21b]. Tursiops [CCA^{+}20, TPZ^{+}21]. turtle [ARH^{+}20, BPS^{+}21, BDB^{+}20, BASM^{+}24, CTH^{+}20, CBJ^{+}23, LWP^{+}24, MSR^{+}21a, RPB21, RH20, SYMR21, SHM^{+}20a, SBC^{+}21, WHRC22, YBSH23]. turtles [ARE^{+}24, APH^{+}23, AEHD21, CCS^{+}22, DGM^{+}23b, GFT^{+}21, GSL^{+}24b, HKK^{+}23, LJ20, MLE^{+}22, MJJJ20, MAH^{+}20, MFCF21, MSV21, PLC^{+}20, RVCT21, RMR^{+}23, TTM^{+}22, TTM^{+}24, VPH^{+}21, WCL20, WCP^{+}21a]. Twenty [HKK^{+}23]. Twenty-year [HKK^{+}23]. Twilight [MDB^{+}21, MMD20]. Two [MdMR24, TCY^{+}22, ASBE22, Ano24o, BMM21, BTMS24, BTB^{+}24, BQC^{+}22, BBG^{+}22, CGM^{+}24, CEG^{+}21, CBJ^{+}23, CCP^{+}21, DBP^{+}22, DWJL^{+}21, DCL21, FWA^{+}23, GAF20, GSC^{+}22, GNC^{+}23, GPCM22, GMLC22, GMB^{+}24, HIT^{+}22, HTGG22, JNJ^{+}20, JH20, KKC^{+}24, KKH22, KVdM^{+}22, LBMF^{+}24, LGGR^{+}22, MDS^{+}23, MPB^{+}24, MKR^{+}20, OBS^{+}22, PCLQ^{+}20, PDH^{+}24, RSS20, RH23, RC23, SCB^{+}20, SWC^{+}23a, SCM^{+}23, SIBM20, TSW^{+}21, UPK^{+}20, VOP21, WPG^{+}24, XWH^{+}22, ZZX^{+}20]. two-fold [DWJL^{+}21]. Two-way [TCY^{+}22]. type [HTGG22, MCK21, OJ24, XWW^{+}21]. types [LFK^{+}21, LRD^{+}20, MFC^{+}24, RS20a]. typus [ASPM21].

U [RBVA23]. ubiquitous [MdMR24]. UK [CWC^{+}22, DNJCH21, NRE21, PRCvdK23]. Ulleung [CKM^{+}20]. Ulva [WCS^{+}21]. unaffected [WF21]. uncalcified [HM23]. Uncertain [HCvdHM23]. uncertainty [BSH^{+}21b, DCR^{+}21, GCGCAS^{+}20, LORB^{+}24]. uncover [SDW20]. undecimalis [AWC21]. underestimated [GFDN^{+}24]. undergoing [NSM24]. underlying [CKM^{+}20]. understand [Ano23, CDP^{+}21]. Understanding [NT24, SAGG^{+}24, SDE^{+}21]. understorey [VGG^{+}23]. understory [BE21, KKN^{+}20, WCRTT20]. underwater [BWGH23, FWGD^{+}23, MDH^{+}20, MTSO^{+}24, VDC^{+}24, WYZ^{+}20, YXB22]. undulata [DHM^{+}24]. undulate [DHM^{+}24]. Unexpected [KP21]. Unified [DAB^{+}21]. unique [FLCQ21, RBFA20]. uniqueness [MTK^{+}21]. United [TTA20]. units [ZZYB21]. unmanned [KDS^{+}20, LS21]. unoccupied [AKGA^{+}21]. Unraveling [CCAdA24, ZXL^{+}24]. Unravelling [CGL^{+}22, HEKH24]. unravels [AFP^{+}24]. unrelated [KTW^{+}22]. unscheduled [Str23]. unusual [CRGMA^{+}21]. upgrades [MCK21]. Upogebia [LRC20]. upper [DSHF22, LFK^{+}21, MMD20]. Uptake [VS23, HvdSR^{+}21, dFRvdG^{+}22]. Upwelling [DSHF22, DFSH23, DHF24, BRC22, CEG^{+}21, CAAG^{+}22, CKM^{+}20, NM22, RWM^{+}22, SRHM21]. upwelling-driven [RWM^{+}22]. Urban [HGR^{+}22, LTM20, PVL^{+}21]. urbanised [DVP^{+}23]. Urbanization [RB20]. urbanized [MCLC^{+}23, TRK^{+}22]. urbanizing [RLSO23]. urchin [AAI^{+}24, BKTN21, BMB24, BPSB^{+}22, CCRP20, DKAB23, DKS^{+}24, DMM^{+}24b, KNH24, LvAH^{+}24, MJ23, MR20, SST20b, UPK^{+}20, WK21, WCP^{+}21b]. Urgent [ER20]. Uria [SFMK^{+}23]. Uridine [ZLHH20]. Uridine-5’-tri-phosphate [ZLHH20]. Ursus [HDL^{+}21, KTR^{+}20, SLB^{+}21]. USA [AAHW20, EWF^{+}21, JBWS24, RH20, SAMP24, AKW^{+}22, APH^{+}23, BGG^{+}22, BBRM22, CCW21, DMBB22, FSS23, FPB^{+}21, GLP^{+}24, HGK^{+}24, HDA^{+}22, HKD^{+}20, HBJ^{+}21, MJ21, MMS^{+}22, MMS^{+}20, PDFH20, PTCD22, SRHM21, SRG^{+}24, SNH^{+}20, TD21, TSS^{+}23, VSGD24]. Use [ANP^{+}24, AUS^{+}23, GCDP21, LCR^{+}22, MJK24, ACA^{+}24, AGIS^{+}22, BSB^{+}21, BAAR20, BZT^{+}23, BKA^{+}20, BSP^{+}22, BOY^{+}23, BGDJ20, CP22b, CAPS20, DCN^{+}23, DIC^{+}22, ECvBL21, FS21, FSK^{+}21, GFK^{+}23, GBF^{+}23, HNB^{+}24, HDL^{+}21, HBG^{+}22, JBWS24, JFR^{+}21, JFB^{+}23, KHSS23, LBH^{+}21c, LTM20, LNBdB21, MCvdM^{+}24, MJJJ20, MDK^{+}21, MCLC^{+}23, NCS^{+}23, NSB^{+}23, OLG^{+}23, PBAC23, PSN^{+}20, PGT^{+}21, PDF^{+}23, RB22, RH21, RLP^{+}23, RBF^{+}22, SHL^{+}21, THD^{+}22, TCY^{+}22, TWC^{+}23, VGLL^{+}21, VK21, VvDRM20, WAA^{+}23, WIM23, YSD^{+}21]. used [DDB^{+}24, MCK21]. Using [BCP^{+}22, CDP^{+}21, GBA24, JBGA23, KTK21, MCF^{+}21, NSKH23, OH22, VNG^{+}22, WYO^{+}20, APOGG20, AAFC^{+}21, BCC20, BBH^{+}20, BBKW20, BNR20, BMB24, CDD^{+}20a, CHDG^{+}23, CG22, CKKL21, DMM^{+}24a, DSA^{+}24, EPHS23, EAS20, FWWH22, FLC^{+}24a, FTC^{+}20, GKK^{+}22, HKY^{+}20, HSU^{+}21, HBG^{+}22, HRC21, JDM^{+}20, KSS^{+}20, LMS^{+}20, MJJJ20, MTSO^{+}24, PSÓ^{+}24, RRS^{+}23, SFL23, SMK23, SHS21, SBFP^{+}24, SSF^{+}22, TTJS20, WYZ^{+}20, ZXL^{+}24]. utilization [LDM^{+}20, PWvdM^{+}21, ZDBS20]. Utilizing [FOO20, CKM^{+}20]. UV [GCDA20, KCLJ20].

Valdés [ASS^{+}21b, MBW^{+}20]. validate [MTK^{+}21]. Validating [LNBdB21]. validation [CBTL20, FLC^{+}24a, KSMP20, LH22]. value [HM23]. values [GBA21, LNBdB21, OCC23, OTD^{+}24, SBS^{+}24]. vanished [MLVJ23]. variabilities [THS^{+}24]. Variability [AGA^{+}22, GGRV23, KRG^{+}20, NM22, BRC22, BWRP20, BMOW21, BDS^{+}20, CRB^{+}21, Cha22, CSH^{+}21b, DSB^{+}21, FSAG22, FLBP^{+}21, GFDPASR20, GSBO21, GDCP20, GVOCSGG21, GBA24, HPRP21, JCGS22, JNY^{+}20, KMW^{+}21, KNH24, LBH^{+}21a, LNM22, MCvdM^{+}24, PHE^{+}20, RB22, RWM^{+}22, SYMR21, SvBS^{+}20, TTA20, TOMDC23, WDJ20, ZWXC21, vdVWF^{+}24, vdVVC^{+}21]. Variable [DCG^{+}20, OBS^{+}22, SLJ^{+}20, CG22, GVS23, HPRP21, MDM^{+}20, ZMD^{+}21]. variables [GVCSO^{+}24, TRK^{+}22]. Variation [BSFM20, CAY24, CGR^{+}20, HB24, JBG22, RBCAV22, ABD^{+}21, AKW^{+}22, AVFD^{+}21, ACC^{+}20, BFQ^{+}23, BTA^{+}20, BSB^{+}21, BKA^{+}20, BBG^{+}22, CAV^{+}20, CFFC^{+}23, CMdL^{+}20, CFB21, CSN^{+}23, CH20, CFSH22, DCN^{+}23, DRP^{+}20, EP23, FWA^{+}23, FDE^{+}21, GBA^{+}22, GWDR20, HAB^{+}22, HW20, JWH^{+}21, JHDE20, LKR^{+}21, LO20, LDW^{+}21, MMB22, MBGM21, MBB^{+}23, NHH^{+}21, OCC23, OLG^{+}23, PMH23, PRC^{+}21, RMASA20, RWSR24, SBMW22, ŠGH^{+}21, SJS^{+}23, SHH^{+}22, SAF^{+}24, TPZ^{+}21, TD21, WPM22, WYZ^{+}20]. Variations [SFF^{+}20, DAT^{+}21, SGB^{+}21, WXZ^{+}24, ZMC^{+}22]. Varied [PGM^{+}23]. variegatus [EHHF^{+}20]. varies [GMLC22, IBC^{+}21]. vary [CCA^{+}21, CHJ21, DBP^{+}22, MHH^{+}20, MSC^{+}20, RBB^{+}24]. Varying [FBP^{+}22, SS24]. Vazella [BKM^{+}21]. vegetation [GVR^{+}20, GBA21, NF23, WAA^{+}23]. vehicle [AKGA^{+}21]. vehicles [KDS^{+}20, LS21]. veined [CZS^{+}22]. Velella [JPB21]. velocity [ENF^{+}22]. VEMCO [OH22]. venerid [RVV^{+}24]. vent [CGCC24, FBM^{+}24]. vents [ZCB20]. verified [LBW21]. versus [CSK^{+}23, JRVE^{+}22, SBS20, THS^{+}24, VSG20]. vertebral [UIS^{+}24]. vertebrate [SJS^{+}23]. Vertical [OAM^{+}24, RBF^{+}22, ASPM21, CZS^{+}22, FKTK20, IMM^{+}20, LRM^{+}21, SFF^{+}20, SBJ^{+}24, TCM^{+}23, VDB^{+}24]. vessel [MPM21, SHL^{+}21]. vessels [LGR^{+}22, MBM^{+}20]. via [HNN^{+}24, SSF^{+}22, XCT^{+}20]. viability [KWD21, PMH23]. viable [ASY^{+}24]. video [BGG^{+}22, BWGH23, FWGD^{+}23, HO21, JS20, SAMdL^{+}20, THK^{+}22, WYZ^{+}20]. video-recorded [THK^{+}22]. view [MTRM21]. viewed [XCT^{+}20]. Vigo [LAGGM^{+}21]. Vinegar [DDB^{+}24]. Virgin [AEHD21]. virginica [AR21a, GCDP21, MVSV22, PRD^{+}20]. Vista [THD^{+}22]. Visual [CFHCS^{+}24, DPDJ^{+}21, MSC^{+}22, PS23, VDC^{+}24]. vital [FDE^{+}21, MHN24]. viviparous [KWD21, RHN^{+}20]. vocalisations [YXB22]. volitans [HBJ^{+}21]. Vortical [JP20]. VPS [OH22]. vs [CCC^{+}23, DVMH^{+}24, NBDM^{+}24, SAMdL^{+}20, SB20b, TWC20]. vulgaris [GMRMG22]. vulnerabilities [ACA^{+}24]. vulnerable [CSK^{+}22, NCS^{+}23]. vulpinus [KKS^{+}20].

Wadden [BD20, JKvLP20, PWvdM^{+}21, PWvdM^{+}23, RLP^{+}23, vdVTW^{+}22]. waist [RBF^{+}20]. walleye [OLM24]. walrus [BLE^{+}22, JETGM24]. wandering [SWR23]. warm [AMUPJR^{+}24, CKS^{+}24, DVMH^{+}24, HDE21, LEG^{+}20, OLM24, OHBJ20, PSBH21, QHT^{+}21, RH20]. warm- [CKS^{+}24]. warm-water [OHBJ20, PSBH21]. warmer [OVÁS^{+}21]. Warming [BPS^{+}24a, GEY^{+}21, MHF^{+}21, SFLQ23, Ano24o, BK23, BG21, BBM^{+}24, BMB24, DGP22, FBKG24, HGH^{+}24, HO24, JPL^{+}24, KNO^{+}21, KGM^{+}24, LK21, LBMSS23, MR20, MW20, MMP^{+}22, NHFS22, NBDM^{+}24, PDM22, SJ22, SOSE22, SJSR24, Tha21, TFH^{+}22, VIV^{+}20, WC21, WAA24]. Warming-induced [SFLQ23]. Warsaw [SRS^{+}20]. Washington [LAD^{+}21, TSS^{+}23]. wasp [RBF^{+}20]. wasp-waist [RBF^{+}20]. wasting [GEY^{+}21, Hew21, SKHC23, TSS^{+}23]. watching [LGR^{+}22]. Water [ASPM21, GDG^{+}22, RBM^{+}21, BS20a, CGCC24, CKS^{+}24, DNE^{+}24, DCR^{+}21, GSPH20, GED20, HBC^{+}20, HWA^{+}22, KBP22, MKR^{+}20, MLAS22, NGRH20, OHBJ20, OJPC23, PSBH21, PLC^{+}20, TP20, VNM^{+}21, VNG^{+}22, XWW^{+}21, ZFD^{+}24, dlBSL^{+}22]. waters [AHMVM23, AGIS^{+}22, AGB^{+}23, BGG^{+}22, BVD^{+}20, CDC^{+}23, CSDB22, Fig21, GCDA20, HO24, JPJH23, LBH^{+}21b, MdOO^{+}20, MAMF21, MMT^{+}24, PRCvdK23, QHT^{+}21, SBJ^{+}24, WXZ^{+}24, YOI^{+}23, dlBSR^{+}20]. Wave [AMD22, LT21, MMP^{+}22, TSW^{+}21]. waves [KCT24, KPS^{+}22]. wavy [MJ21]. way [TCY^{+}22]. weak [OvBS^{+}24]. weaning [NBL^{+}22]. weather [Ano23, MSH20, MY21, RLME^{+}21]. web [CLV^{+}20, CGM23, DSA^{+}24, GSRC20, GB20, IAG^{+}23, LRD^{+}20, MMD20, MLAS22, PWvdV20, PWvdM^{+}23, RCC^{+}24, SDM^{+}22, SMS^{+}24]. webs [Bai22, HBE22, HW20, vdRJL20]. Weddell [ORB^{+}21, VH20]. wedge [CGL^{+}22, RPD^{+}23, WdGR^{+}20]. wedge-tailed [WdGR^{+}20]. weight [KWD21, MVC^{+}23]. Weizhou [NYW^{+}22]. well [Ano23, PLH^{+}21, SRG^{+}23]. well-connected [PLH^{+}21]. well-defined [SRG^{+}23]. were [CASF21]. West [DGM^{+}23a, LBH^{+}21a, CKCS20, PRM^{+}21, CHD^{+}22, FFT^{+}20, GMRMG22, GPP^{+}24, MHN24, MPF^{+}22, OLPGK22]. Western [LMB^{+}20, PSBH21, BBPO21, BGDJ20, CVP^{+}21, CFBPH23, DA24, FYM^{+}24, FOO20, GSL^{+}24a, GDCP20, HDL^{+}21, JKvLP20, KPK^{+}24, KSMP20, LOF^{+}24, MdOO^{+}20, MDS^{+}23, MAH^{+}20, MLVJ23, OÁBP^{+}21, PCLQ^{+}20, PWvdM^{+}21, SKTO23, TTM^{+}22, TTM^{+}24, ANP^{+}24, DBHCO^{+}22, ECvBL21, GCDA20, IAG^{+}23, LBMSS23, MFC^{+}24, RBF^{+}22, SVB^{+}24, SB20a]. wetland [HSBB^{+}21, WAA^{+}23, ZHG^{+}21]. whale [ASPM21, ASS^{+}21b, BBEF^{+}20, BMBC21, BPS^{+}24b, CDD^{+}20b, DFP^{+}21b, GVAZ^{+}22, HSO^{+}22, KTW^{+}22, LGR^{+}22, LMB^{+}20, MBW^{+}20, MTK^{+}21, OYO^{+}20, OTD^{+}24, PBS^{+}24, PVB^{+}20, PDF^{+}23, RFJC^{+}24, RYF^{+}24, RRR^{+}23, SBFP^{+}24, SSF^{+}22, VBB^{+}21, WIM23]. whale-fall [OYO^{+}20]. whale-watching [LGR^{+}22]. whales [AU21, ANP^{+}24, BKR^{+}21, BBPO21, BSFM20, BBSM^{+}20, BMBC21, BMS^{+}24, BMS^{+}21, BZT^{+}23, BBGM23, CPD^{+}20, CGR^{+}20, CRGMA^{+}21, CBB^{+}22a, CFBPH23, DFHT23, DME^{+}23, FPG^{+}21, FPB^{+}21, FNL^{+}23, FFT^{+}20, GWDR20, HSA^{+}21, HDA^{+}22, HCvdHM23, JFB^{+}23, MDK^{+}21, MBM^{+}20, OTD^{+}24, PLLT^{+}21, PAL^{+}22, RBVA23, RSvV^{+}23, RCC^{+}22, SBS^{+}24, SJQ^{+}22, SHL^{+}21, SDE^{+}22, WYO^{+}20, WMT20]. whelk [HMT21]. whelks [BNS^{+}24, WBS24]. where [ER20, MLVJ23]. while [FCV^{+}24, Sig23]. whiptail [OMW^{+}20]. whisker [MWCR20]. White [URSS24, BGMRGM^{+}20, EVAB^{+}23, LBH^{+}21c, MPB^{+}24, PSB^{+}23, SBS^{+}22, SMB^{+}20, WFS23, ACG22, CAY24]. white-beaked [SBS^{+}22]. white-tailed [EVAB^{+}23]. Whitfield [BS21]. whiting [MAOR^{+}21, PJA^{+}21]. wide [BPA21, HGT^{+}21, KPW^{+}20, LBB^{+}21]. wide-dispersing [BPA21]. wide-ranging [KPW^{+}20]. wider [MLAS22]. widespread [CCS^{+}22, DL21]. width [DHM^{+}24]. wild [CS20, GMRMG22, KSMP20, MVH^{+}21, MVC^{+}23]. wildlife [NJHN24]. will [SHM^{+}20b]. William [DMCB22]. wilsoni [ARN^{+}22]. Wind [SWR23, Ano23, HPFB20, JTBS^{+}22, KPTT23, KKMW23, SSC^{+}23, TCP^{+}23, WBG^{+}22, WBGC24, vESB^{+}23]. wing [Ano23]. winner [BKM^{+}21]. winter [ABD^{+}21, BBGM23, CBHM21, CSD^{+}24, DFCM21, FKTK20, GBG^{+}23, HDM^{+}24, HRC21, IMM^{+}20, JW22, MLB^{+}21, MAOR^{+}21, MDB^{+}21, NSM24, NBL^{+}22, RH21, TFCS22]. wintering [DMS^{+}21, LBTE^{+}21, PGGE21, SdFBM24]. Within [BFQ^{+}23, PMH23, AFTW24, BGGH^{+}22, FS22, HMV^{+}22, HB24, HMK^{+}22, JBGA23, LFFW21, MWCR20, MA20, OFSJ20, SW23b, STH21, SFMK^{+}23, STF^{+}24]. Within- [BFQ^{+}23]. within-brood [HB24]. Within-family [PMH23]. within-patch [BGGH^{+}22]. within-stock [AFTW24]. without [CBVA^{+}21]. wood [HSSN^{+}21]. World [MB20, Ano24o, BDB^{+}20, ER20, HGK^{+}22, PGC21, Rus20, RBCwD^{+}20]. worms [CGLC20, PGC21, SFL23]. wrack [GSBO21, WSD^{+}20]. wrasse [HSNT24]. wurdemanni [ZLHH20]. Wysiecki [Ano22e].

Xenograpsus [CGCC24]. Xiphias [APOGG20, HTT^{+}21, SSB^{+}23].

Yaeyama [HKK^{+}23]. Yangtze [XYSF20]. Year [BGMAM^{+}21, FTA^{+}21, FDS^{+}21, VH20, CAW^{+}22, EMB^{+}22, GVS23, HKK^{+}23, HHO^{+}23, KCGR20, LMB^{+}20, LHM^{+}23, LAD^{+}21, MHDJ21, NBL^{+}22, RDL^{+}22, ZPVN23]. Year-round [BGMAM^{+}21, FTA^{+}21, FDS^{+}21, VH20]. years [BBH^{+}20, CCA^{+}21, HB24, MW20]. yellow [MCFB20, MCC^{+}21, YSP^{+}22, WCS^{+}21, ZXL^{+}24]. yellow-eyed [MCFB20, MCC^{+}21, YSP^{+}22]. yellowfin [AAFC^{+}21, DDD^{+}23, FLC^{+}24a, SKTO23]. yellowtail [DV20]. yellowtails [FKTK20]. yezoense [KKN^{+}20]. yields [CASF21, CAPS20]. York [TCT^{+}23]. young [KCGR20, MHDJ21]. young-of-the-year [KCGR20, MHDJ21].

Zalophus [LNM22]. Zamboanga [BRC22]. Zealand [BRD20, CWD24, FSF^{+}24, FNM^{+}21, FJB^{+}23, GAFS20, HDB21, KSW23, MON^{+}24, OCC23, PAL^{+}22, QRC^{+}24, RJ21, RDL^{+}21, SW23a]. Zemlya [EGK^{+}21]. Zero [BAS21]. Ziphius [FPB^{+}21]. zoantharian [KZC23]. zone [ECC^{+}21, Fai21, LCR^{+}22, MMD20, PSB^{+}23, SFLQ23]. zones [COM^{+}24, LGGR^{+}22, TKP^{+}20]. zooplanktivorous [HBS^{+}21, LCFJ20]. zooplankton [BAS23, BBL^{+}21b, DKSS21, Fig21, HHS21, IMYW24, KCT24, LCFJ20, MAK^{+}22, NTN^{+}24, SS24, URSS24]. zooxanthellae [XMY^{+}22]. zooxanthellate [DPL^{+}21]. Zostera [AKW^{+}22, BP24, CAEG21, CEZH21, DNE^{+}24, HMKK^{+}20, ILON21, JRVE^{+}22, MGHH21, ROMB22, SLO^{+}22, SNE^{+}24, ZZX^{+}20, dFRvdG^{+}22, vdHMA^{+}24].

References

Artetxe-Arrate:2021:DYT

[AAFC^{+}21] I. Artetxe-Arrate, I. Fraile, N. Clear, A. M. Darnaude, D. L. Dettman, C. Pécheyran, J. Farley, and H. Murua. Discrimination of yellowfin tuna Thunnus albacares between nursery areas in the Indian Ocean using otolith chemistry. Marine Ecology Progress Series, 673:165–181, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

               URL https://www.int-res.com/articles/meps2021/673/
               m673p165.pdf.

Auth:2020:SEG

[AAHW20] T. D. Auth, T. Arula, E. D. Houde, and R. J. Woodland. Spatial ecology and growth in early life stages of bay anchovy Anchoa mitchilli in Chesapeake Bay (USA). Marine Ecology Progress Series, 651:125–143, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

               int-res.com/articles/meps2020/651/m651p125.pdf.

Agatsuma:2024:RRE

[AAI^{+}24] Y. Agatsuma, T. Abe, E. Inomata, S. Takagi, K. Tanaka, M. Hirotsune, K. Maeda, and M. N. Aoki. Restricted recovery of Eisenia bicyclis kelp forest following sea urchin removal on the Pacific coast of Tohoku, Japan. Marine Ecology Progress Series, 748:17–31, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/748/
m748p017.pdf.

Archer:2023:LPP

[AAP^{+}23] L. C. Archer, S. N. Atkinson, A. M. Pagano, S. R. Penk, and P. K. Molnár. Lactation performance in polar bears is associated with fasting time and energetic state. Marine Ecology Progress Series, 720:175–189, October 5, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m720p175.pdf.

Aalto:2020:APM

[ABB^{+}20] E. A. Aalto, J. P. Barry, C. A. Boch, S. Y. Litvin, F. Micheli, C. B. Woodson, and G. A. De Leo. Abalone populations are most sensitive to environmental stress effects on adult individuals. Marine Ecology Progress Series, 643:75– 85, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/643/m643p075.pdf.

Arai:2024:EGE

[ABC^{+}24] K. Arai, J. E. Best, C. A. Craig, V. Lyubchich, N. R. Miller, and D. H. Secor. Early growth and environmental conditions control partial migration of an estuarine-dependent fish. Marine Ecology Progress Series, 732:149–166, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/732/m732p149.pdf.

Albert:2021:IAV

[ABD^{+}21] C. Albert, V. S. Bråthen, S. Descamps, T. Anker-Nilssen, A. Cherenkov, S. Christensen-Dalsgaard, J. Danielsen, K. E. Erikstad, M. Gavrilo, S. A. Hanssen, H. H. Helgason, J. E. Jónsson, Y. Kolbeinsson, Y. Krasnov, M. Langset, E. Lorentzen, B. Olsen, T. K. Reiertsen, H. Strøm, G. H. Systad, G. Tertitski, P. M. Thompson, T. L. Thórarinsson, P. Bustamante, B. Moe, and J. Fort. Inter-annual variation in winter distribution affects individual seabird contamination with mercury. Marine Ecology Progress Series, 676:243–254, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/676/m676p243.pdf.

Andrews:2021:TPL

[ABH^{+}21] K. Andrews, B. Bartos, C. J. Harvey, D. Tonnes, M. Bhuthimethee, and P. MacCready. Testing the potential for larval dispersal to explain connectivity and population structure of threatened rockfish species in Puget Sound. Marine Ecology Progress Series, 677:95–113, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/677/m677p095.pdf.

Anderson:2021:PEP

[ABN^{+}21] B. N. Anderson, H. D. Bowlby, L. J. Natanson, R. Coelho, E. Cortés, A. Domingo, and J. A. Sulikowski. Preliminary estimate of post-release survival of immature porbeagles caught with rod-and-reel in the Northwest Atlantic Ocean. Marine Ecology Progress Series, 660:153–159, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/660/
m660p153.pdf.

Arostegui:2020:SSO

[ABWJ^{+}20] M. C. Arostegui, C. D. Braun, P. A. Woodworth-Jefcoats, D. R. Kobayashi, and P. Gaube. Spatiotemporal segregation of ocean sunfish species (Molidae) in the eastern North Pacific. Marine Ecology Progress Series, 654:109–125, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m654p109.pdf.

Arfianti:2020:GBM

[AC20] T. Arfianti and M. J. Costello. Global biogeography of marine amphipod crustaceans: latitude, regionalization, and beta diversity. Marine Ecology Progress Series, 638:83–94, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/638/m638p083.pdf.

Arnoldi:2024:SSS

[ACA^{+}24] N. S. Arnoldi, A. B. Carlisle, S. Andrzejaczek, M. R. Castleton, F. Micheli, R. J. Schallert, T. D. White, and B. A. Block. Salmon shark seasonal site fidelity demonstrates the influence of scale on identifying potential high-use areas and vulnerabilities. Marine Ecology Progress Series, 735:125–140, May 2, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/735/m735p125.pdf.

Andrzejaczek:2020:IVR

[ACC^{+}20] S. Andrzejaczek, T. K. Chapple, D. J. Curnick, A. B. Carlisle, M. Castleton, D. M. P. Jacoby, L. R. Peel, R. J. Schallert, D. M. Tickler, and B. A. Block. Individual variation in residency and regional movements of reef manta rays Mobula alfredi in a large marine protected area. Marine Ecology Progress Series, 639:137–153, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/639/m639p137.pdf.

Azovsky:2022:CLS

[ACG22] A. I. Azovsky, E. S. Chertoprud, and L. A. Garlitska. Community-level spatiotemporal synchrony: new metric and application to White Sea meiobenthic harpacticoids. Marine Ecology Progress Series, 698:55–68, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/698/
m698p055.pdf.

Appoo:2021:PDC

[ACvdC^{+}21] J. C. Appoo, E. F. Cagua, J. van de Crommenacker, A. J. Burt, P. Pistorius, F. E. Taylor, H. Richards, P. Haupt, N. Doak, F. Fleischer-Dogley, and N. Bunbury. Population dynamics of the coconut crab Birgus latro on Aldabra Atoll, Seychelles. Marine Ecology Progress Series, 665:103– 113, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/665/m665p103.pdf.

Atherley:2021:SSM

[ADBF21] N. A. M. Atherley, M. M. Dennis, D. C. Behringer, and M. A. Freeman. Size at sexual maturity and seasonal reproductive activity of the Caribbean spiny lobster Panulirus argus. Marine Ecology Progress Series, 671:129–145, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m671p129.pdf.

Aarflot:2020:FSP

[ADF20] J. M. Aarflot, P. Dalpadado, and Ø. Fiksen. Foraging success in planktivorous fish increases with topographic blockage of prey distributions. Marine Ecology Progress Series, 644:129–142, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m644p129.pdf.

Asada:2021:ASL

[AEHD21] A. Asada, S. A. Eckert, W. H. Hagey, and R. W. Davis. Antipredatory strategies of leatherback sea turtles during internesting intervals on St. Croix, US Virgin Islands. Marine Ecology Progress Series, 678:153–170, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/678/
m678p153.pdf.

Adams:2024:REP

[AESS24] D. H. Adams, D. D. Edwards, J. E. Schneider, and A. R. Searles. Range expansion and population shifts of estuarine fishes in a changing subtropical estuary. Marine Ecology Progress Series, 728:221–238, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m728p221.pdf.

Alonso:2024:IEM

[AFP^{+}24] V. A. García Alonso, B. C. Franco, E. D. Palma, M. Pájaro, and F. L. Capitanio. Integration of empirical and modelled data unravels spatiotemporal distribution and connectivity patterns of Fuegian sprat early life stages. Marine Ecology Progress Series, 731:27–49, March 13, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m731p027.pdf.

Allgayer:2024:SPW

[AFTW24] R. L. Allgayer, P. G. Fernandes, J. M. J. Travis, and P. J. Wright. Spatial patterns of within-stock connectivity provide novel insights for fisheries management. Marine Ecology Progress Series, 731:159–178, March 13, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m731p159.pdf.

Abraham:2022:VAC

[AGA^{+}22] A. Clayshulte Abraham, D. J. Gochfeld, B. Avula, K. J. Macartney, M. P. Lesser, and M. Slattery. Variability in antimicrobial chemical defenses in the Caribbean sponge Agelas tubulat: implications for disease resistance and resilience. Marine Ecology Progress Series, 690:51–64, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/690/
m690p051.pdf.

Artero:2023:SAS

[AGB^{+}23] C. Artero, S. D. Gregory, W. A. Beaumont, Q. Josset, N. Jeannot, A. Cole, L. Lamireau, E. Réveillac, and R. B. Lauridsen. Survival of Atlantic salmon and sea trout smolts in transitional waters. Marine Ecology Progress Series, 709: 91–108, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/709/m709p091.pdf.

Alvarez:2024:NPA

[AGB24] C. D. Alvarez, A. R. Giussi, and F. Botto. Niche partitioning among demersal marine fishes at the southern tip of South America. Marine Ecology Progress Series, 739:191– 205, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/739/m739p191.pdf.

Arronte:2022:HUG

[AGIS^{+}22] J. C. Arronte, J. M. González-Irusta, R. Somavilla, J. Fernández-Feijoo, S. Parra, and A. Serrano. Habitat use by Gadiculus argenteus (Pisces, Gadidae) in the Galician and Cantabrian Sea waters (NE Atlantic). Marine Ecology Progress Series, 694:175–191, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/694/m694p175.pdf.

Accola:2022:ACJ

[AHCT22a] K. L. Accola, J. K. Horne, J. R. Cordell, and J. D. Toft. Acoustic characterization of juvenile Pacific salmon distributions along an eco-engineered seawall. Marine Ecology Progress Series, 682:207–220, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/682/m682p207.pdf.

Accola:2022:NDJ

[AHCT22b] K. L. Accola, J. K. Horne, J. R. Cordell, and J. D. Toft. Nocturnal distributions of juvenile Pacific salmon along an eco-engineered marine shoreline. Marine Ecology Progress Series, 687:113–123, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/687/m687p113.pdf.

Allen:2020:ISP

[AHH^{+}20] S. Allen, S. Henson, A. Hickman, C. Beaulieu, P. C. Doncaster, and D. G. Johns. Interannual stability of phytoplankton community composition in the North-East Atlantic. Marine Ecology Progress Series, 655:43–57, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m655p043.pdf.

Ahuatzin-Hernandez:2023:BPH

[AHMVM23] J. M. Ahuatzin-Hernández, J. J. Morrone, and V. M. Vidal-Martínez. Biogeographic patterns of hydrozoans in the Gulf of Mexico and adjacent waters: biological assemblages, beta diversity, and endemism. Marine Ecology Progress Series, 720: 85–94, October 5, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/720/m720p085.pdf.

Alder:2022:TMD

[AJH22] A. Alder, A. Jeffs, and J. Hillman. Timing mussel deployments to improve reintroduction success and restoration efficiency. Marine Ecology Progress Series, 698:69–83, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/698/
m698p069.pdf.

Ayres:2021:SAB

[AKGA^{+}21] K. A. Ayres, J. T. Ketchum, R. González-Armas, F. Galván-Magaña, A. Hearn, F. R. Elorriaga-Verplancken, R. O. Martínez-Rincón, E. M. Hoyos-Padilla, and S. M. Kajiura. Seasonal aggregations of blacktip sharks Carcharhinus limbatus at a marine protected area in the Gulf of California, assessed by unoccupied aerial vehicle surveys. Marine Ecology Progress Series, 678:95–107, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/678/m678p095.pdf.

Akita:2022:IOS

[AKU^{+}22] Y. Akita, T. Kurihara, M. Uehara, T. Shiwa, and K. Iwai. Impacts of overfishing and sedimentation on the feeding behavior and ecological function of herbivorous fishes in coral reefs. Marine Ecology Progress Series, 686:141–157, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/686/
m686p141.pdf.

Allcock:2022:SVP

[AKW^{+}22] K. E. Allcock, S. J. Kamel, P. L. Willeboordse, Z. T. Long, and J. C. Jarvis. Spatiotemporal variation in patterns of genetic diversity, genetic structure, and life history across Zostera marina meadows in North Carolina, USA. Marine Ecology Progress Series, 683:53–66, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/683/m683p053.pdf.

Aksnes:2023:EMM

[ALL^{+}23] D. L. Aksnes, A. S. Løtvedt, C. Lindemann, M. L. Calleja, X. A. G. Morán, S. Kaarvedt, and T. F. Thingstad. Effects of migrating mesopelagic fishes on the biological carbon pump. Marine Ecology Progress Series, 717:107–126, August 24, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m717p107.pdf.

Acquafredda:2020:ESD

[AM20] M. P. Acquafredda and D. Munroe. Effect of species diversity on particle clearance and productivity in farmed bivalves. Marine Ecology Progress Series, 639:107–126, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m639p107.pdf.

Attridge:2022:WEA

[AMD22] C. M. Attridge, A. Metaxas, and D. Denley. Wave exposure affects the persistence of kelp beds amidst outbreaks of the invasive bryozoan Membranipora membranacea. Marine Ecology Progress Series, 702:39–56, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/702/m702p039.pdf.

Aceves-Medina:2020:DAS

[AMMADDH20] G. Aceves-Medina, M. Moreno-Alcántara, R. Durazo, and D. Delgado-Hofmann. Distribution of Atlantidae species (Gastropoda: Pterotracheoidea) during an El Niño event in the Southern California Current System (summer–fall 2015). Marine Ecology Progress Series, 648:153–168, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/648/
m648p153.pdf.

Aguaiza:2024:MNH

[AMRL^{+}24] C. Aguaiza, P. J. Mumby, E. Rastoin-Laplane, O. Aburto-Oropeza, and P. Salinas-De-León. Mangroves as nursery habitats for Galapagos reef fishes. Marine Ecology Progress Series, 747:99–115, October 17, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2024/747/m747p099.pdf.

Amelineau:2021:SPS

[AMT^{+}21] F. Amélineau, B. Merkel, A. Tarroux, S. Descamps, T. Anker-Nilssen, O. Bjørnstad, V. S. Bråthen, O. Chastel, S. Christensen-Dalsgaard, J. Danielsen, F. Daunt, N. Dehnhard, M. Ekker, K. E. Erikstad, A. Ezhov, P. Fauchald, M. Gavrilo, G. T. Hallgrimsson, E. S. Hansen, M. P. Harris, M. Helberg, H. H. Helgason, M. K. Johansen, J. E. Jónsson, Y. Kolbeinsson, Y. Krasnov, M. Langset, S. H. Lorentsen, E. Lorentzen, M. V. Melnikov, B. Moe, M. A. Newell, B. Olsen, T. Reiertsen, G. H. Systad, P. Thompson, T. L. Thórarinsson, E. Tolmacheva, S. Wanless, K. Wojczulanis-Jakubas, J. Åström, and H. Strøm. Six pelagic seabird species of the North Atlantic engage in a fly-and-forage strategy during their migratory movements. Marine Ecology Progress Series, 676:127–144, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m676p127.pdf.

Aceves-Medina:2024:IEC

[AMUPJR^{+}24] G. Aceves-Medina, A. G. Uribe-Prado, S. P. A. Jiménez-Rosenberg, R. Durazo, R. J. Saldierna-Martínez, R. Avendaño-Ibarra, and A. N. Sarmiento-Lezcano. Influence of extreme cold and warm oceanographic events on larval fish assemblages in the southern region of the California Current. Marine Ecology Progress Series, 728:199–219, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/728/
m728p199.pdf.

Anonymous:2020:CCC

[Ano20] Anonymous. Corrigendum: Camps–Castellà et al., 645:67– 82 (2020). Marine Ecology Progress Series, 647:229, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/647/m647p229.pdf. See [CCBT^{+}20].

Anonymous:2021:CBA

[Ano21a] Anonymous. Corrigendum: Baldanzi et al., 646:93–107 (2020) MEPS 662:215. Marine Ecology Progress Series, 662: 215, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/662/m662p215.pdf. See [BSF^{+}20].

Anonymous:2021:CDA

[Ano21b] Anonymous. Corrigendum: Depot et al., 656:75–87 (2020). Marine Ecology Progress Series, 661:243, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m661p243.pdf. See [DSK^{+}20].

Anonymous:2021:CJAb

[Ano21c] Anonymous. Corrigendum: Johnson et al., 641:111–121 (2020). Marine Ecology Progress Series, 664:243, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/664/m664p243.pdf. See [JSDH20].

Anonymous:2021:CJAa

[Ano21d] Anonymous. Corrigendum: Jusufovski et al., 627:207–232 (2019). Marine Ecology Progress Series, 658:253, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/658/m658p253.pdf. See [JSK19].

Anonymous:2021:CPF

[Ano21e] Anonymous. Corrigendum: Pais de Faria et al., 660:189– 201 (2021). Marine Ecology Progress Series, 663:237–238, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/663/m663p237.pdf. See [PVL^{+}21].

Anonymous:2021:CPA

[Ano21f] Anonymous. Corrigendum: Petrolo et al., 678:109–124 (2021). Marine Ecology Progress Series, 679:241, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/679/m679p241.pdf. See [PBO^{+}21].

Anonymous:2021:EMA

[Ano21g] Anonymous. Erratum: Merrill et al., 666:217–229 (2021). Marine Ecology Progress Series, 670:247, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m670p247.pdf. See [MTB21].

Anonymous:2021:ERA

[Ano21h] Anonymous. Erratum: Riaz et al., 654:177–194 (2020). Marine Ecology Progress Series, 672:243–244, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/672/

m672p243.pdf. See [RBW^{+}20].

Anonymous:2021:ET

[Ano21i] Anonymous. Erratum: Thatje, 660:233–240 (2021). Marine Ecology Progress Series, 666:243, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/666/
m666p243.pdf. See [Tha21].

Anonymous:2022:CBAa

[Ano22a] Anonymous. Corrigendum: Barile et al., (2021) MEPS 661:203–227. Marine Ecology Progress Series, 698:199–??, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/698/m698p199.pdf. See [BBPO21].

Anonymous:2022:CBAb

[Ano22b] Anonymous. Corrigendum: Beatty et al., (2022) MEPS 697:167–182. Marine Ecology Progress Series, 699:199–??, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m699p199.pdf.

Anonymous:2022:CHA

[Ano22c] Anonymous. Corrigendum: Howell et al., (2022) MEPS 693:33–51. Marine Ecology Progress Series, 700:197–??, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/700/m700p197.pdf. See [HBG^{+}22].

Anonymous:2022:CSA

[Ano22d] Anonymous. Corrigendum: Smith et al., 683:221–226 (2022). Marine Ecology Progress Series, 688:197, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m688p197.pdf. See [SHT^{+}22].

Anonymous:2022:EWA

[Ano22e] Anonymous. Erratum: De Wysiecki et al., (2022) MEPS 697:81–95. Marine Ecology Progress Series, 699:201–??, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/699/m699p201.pdf. See [DIC^{+}22].

Anonymous:2022:EHA

[Ano22f] Anonymous. Erratum: Hamilton et al., 688:83–98 (2022). Marine Ecology Progress Series, 691:203, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/691/

m691p203.pdf. See [HPDG22].

Anonymous:2023:TSP

[Ano23] Anonymous. This theme section presents current knowledge on how wind shapes the movement, energetics, foraging behaviour and life history of seabirds. global wind and weather patterns are changing, affecting seabirds through direct impacts on flight and indirect effects across a range of spatiotemporal scales from individual wing beats to population processes. advances in tagging technology have allowed animal location, environmental factors and physiological parameters to be measured simultaneously. contributions to this theme section bring together recent developments in biologging, as well as atmospheric modelling and statistical and theoretical ecology to better understand the links between wind and seabird morphology, movement, energetics and life-history characteristics. Marine Ecology Progress Series, ??:??, November 23, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDa

[Ano24a] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDb

[Ano24b] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, March 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDc

[Ano24c] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDd

[Ano24d] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, April 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDe

[Ano24e] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDf

[Ano24f] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDg

[Ano24g] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDh

[Ano24h] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, August 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDi

[Ano24i] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, August 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDj

[Ano24j] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, September 5, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDk

[Ano24k] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, September 19, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDl

[Ano24l] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, October 17, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDm

[Ano24m] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:CDn

[Ano24n] Anonymous. Copyright and disclaimer. Marine Ecology Progress Series, ??:??, November 21, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Anonymous:2024:TSB

[Ano24o] Anonymous. This theme section brings together the current research on two interrelated marine ecological topics:

(1) invasive species in a warming ocean, and (2) climatemediated species range shifts. marine species are moving beyond their native geographical ranges either in response to climate change or by being introduced into new marine regions as a consequence of an increasingly connected world. the arrival of species at new regions can lead to ecological challenges for both the invader and the recipient community. contributions to this theme section provide evidence that global warming, and occasionally environmental degradation, are crucial factors altering the distribution of both marine invasive species and range-shifting species. Marine Ecology Progress Series, ??:??, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Allen:2024:UPN

[ANP^{+}24] J. A. Allen, R. C. Nichols, L. J. Pallin, D. W. Johnston, and A. S. Friedlaender. Use and prevalence of novel bubble-net foraging strategy in Western Antarctic humpback whales. Marine Ecology Progress Series, 743:97–111, August 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/743/m743p097.pdf.

Ashida:2021:DRT

[AOO^{+}21] H. Ashida, Y. Okochi, S. Ohshimo, T. Sato, Y. Ishihara, S. Watanabe, K. Fujioka, S. Furukawa, T. Kuwahara, Y. Hiraoka, and Y. Tanaka. Differences in the reproductive traits of Pacific bluefin tuna Thunnus orientalis among three fishing grounds in the Sea of Japan. Marine Ecology Progress Series, 662:125–138, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/662/m662p125.pdf.

Adjeroud:2022:EHL

[APG^{+}22] M. Adjeroud, C. Peignon, C. Gauliard, L. Penin, and M. Kayal. Extremely high but localized pulses of coral recruitment in the southwestern lagoon of New Caledonia and implications for conservation. Marine Ecology Progress Series, 692:67–79, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m692p067.pdf.

Aoki:2023:FDS

[APH^{+}23] D. M. Aoki, J. R. Perrault, S. L. Hoffmann, J. R. Guertin, A. Page-Karjian, B. A. Stacy, and D. Lowry. Forensic determination of shark species as predators and scavengers of sea turtles in Florida and Alabama, USA. Marine Ecology Progress Series, 703:145–159, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/703/m703p145.pdf.

Acosta-Pachon:2020:ARM

[APOGG20] T. A. Acosta-Pachón, S. Ortega-García, and B. Graham. Assessing residency and movement dynamics of swordfish Xiphias gladius in the Eastern North Pacific Ocean using stable isotope analysis. Marine Ecology Progress Series, 645: 171–185, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/645/m645p171.pdf.

Ashey:2021:EEH

[AR21a] J. Ashey and E. B. Rivest. Effect of environmental history on the physiology and acute stress response of the eastern oyster Crassostrea virginica. Marine Ecology Progress Series, 674:115–130, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m674p115.pdf.

Ayre:2021:GFA

[AR21b] D. J. Ayre and N. Rosser. Gene flow across a major biogeographic barrier is not increasing under climate change for the barnacle Catomerus polymerus. Marine Ecology Progress Series, 669:97–106, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/669/m669p097.pdf.

Attard:2020:MSR

[ARB^{+}20] K. M. Attard, I. F. Rodil, P. Berg, A. O. M. Mogg, M. Westerbom, A. Norkko, and R. N. Glud. Metabolism of a subtidal rocky mussel reef in a high-temperate setting: pathways of organic C flow. Marine Ecology Progress Series, 645: 41–54, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m645p041.pdf.

Arcangel:2022:SCF

[ARdlB^{+}22] A. E. Arcángel, E. A. Rodríguez, P. de la Barra, P. J. Pereyra, and M. Narvarte. Seasonal changes in facilitation between an ascidian and a kelp in Patagonia. Marine Ecology Progress Series, 693:95–106, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/693/m693p095.pdf.

Allen:2024:SRO

[ARE^{+}24] C. D. Allen, L. E. Peavey Reeves, T. Eguchi, S. J. Sawyer, L. T. Ballance, R. L. Pitman, S. L. Martin, T. T. Jones, and J. A. Seminoff. Sex ratios of olive ridley sea turtles in the North Pacific high seas: implications for climate change research. Marine Ecology Progress Series, 748:149–162, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m748p149.pdf.

Avens:2020:RDK

[ARH^{+}20] L. Avens, M. D. Ramirez, A. G. Hall, M. L. Snover, H. L. Haas, M. H. Godfrey, L. R. Goshe, M. Cook, and S. S. Heppell. Regional differences in Kemp’s ridley sea turtle growth trajectories and expected age at maturation. Marine Ecology Progress Series, 654:143–161, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/654/m654p143.pdf.

Aviz:2022:TDS

[ARN^{+}22] D. Aviz, M. Rollnic, I. B. A. Nascimento, L. A. Pinheiro, and J. S. Rosa Filho. Temporal dynamics of a Sabellaria wilsoni (Sabellariidae: Polychaeta) reef on the Brazilian Amazon Coast. Marine Ecology Progress Series, 702:87– 103, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/702/m702p087.pdf.

Alexandre:2020:CNB

[AS20] A. Alexandre and R. Santos. Competition for nitrogen between the seaweed Caulerpa prolifera and the seagrass Cymodocea nodosa. Marine Ecology Progress Series, 648:125– 134, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/648/m648p125.pdf.

Aarnio:2024:ESP

[AS24] S. Aarnio and J. Soininen. Environmental and stochastic processes drive diatom taxonomic and functional temporal beta diversity. Marine Ecology Progress Series, 742:59–71, August 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/742/m742p059.pdf.

Axler:2020:FSL

[ASBA^{+}20] K. E. Axler, S. Sponaugle, C. Briseño-Avena, F. Hernandez, Jr., S. J. Warner, B. Dzwonkowski, S. L. Dykstra, and R. K. Cowen. Fine-scale larval fish distributions and predator– prey dynamics in a coastal river-dominated ecosystem. Marine Ecology Progress Series, 650:37–61, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/650/
m650p037.pdf.

Andrew:2022:OAR

[ASBE22] S. M. Andrew, R. F. Strzepek, O. Branson, and M. J. Ellwood. Ocean acidification reduces the growth of two Southern Ocean phytoplankton. Marine Ecology Progress Series, 682: 51–64, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/682/m682p051.pdf.

Arronte:2024:HPP

[ASGI24] J. C. Arronte, A. Serrano, and J. M. González-Irusta. Habitat preferences of Phycis blennoides (Pisces: Gadiformes): environmental factors and fishing-related ontogenic deepening. Marine Ecology Progress Series, 745:147–163, September 19, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/745/m745p147.pdf.

Axler:2020:CPE

[ASH^{+}20] K. E. Axler, S. Sponaugle, F. Hernandez, Jr., C. Culpepper, and R. K. Cowen. Consequences of plume encounter on larval fish growth and condition in the Gulf of Mexico.

Marine Ecology Progress Series, 650:63–80, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/650/
m650p063.pdf.

Alberto:2022:HGA

[ASIA22] M. L. Torres Alberto, M. Saraceno, M. Ivanovic, and E. M. Acha. Habitat of Argentine squid (Illex argentinus) paralarvae in the southwestern Atlantic. Marine Ecology Progress Series, 688:69–82, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/688/m688p069.pdf.

Albano:2022:ALF

[ASL^{+}22] P. G. Albano, A. Sabbatini, J. Lattanzio, J. F. Päßler, J. Steger, Q. Hua, D. S. Kaufman, S. Szidat, M. Zuschin, and A. Negri. Alleged Lessepsian foraminifera prove native and suggest Pleistocene range expansions into the Mediterranean Sea. Marine Ecology Progress Series, 700:65–78, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m700p065.pdf.

Arrowsmith:2021:WTK

[ASPM21] L. M. Arrowsmith, A. M. M. Sequeira, C. B. Pattiaratchi, and M. G. Meekan. Water temperature is a key driver of horizontal and vertical movements of an ocean giant, the whale shark Rhincodon typus. Marine Ecology Progress Series, 679:101–114, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/679/m679p101.pdf.

Alewijnse:2021:ODF

[ASS^{+}21a] S. R. Alewijnse, G. Stowasser, R. A. Saunders, A. Belcher, O. A. Crimmen, N. Cooper, and C. N. Trueman. Otolithderived field metabolic rates of myctophids (family Myctophidae) from the Scotia Sea (Southern Ocean). Marine Ecology Progress Series, 675:113–131, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/675/m675p113.pdf.

Azizeh:2021:ACB

[ASS^{+}21b] T. R. Azizeh, K. R. Sprogis, R. Soley, M. L. K. Nielsen, M. M. Uhart, M. Sironi, C. F. Marón, L. Bejder, P. T. Madsen, and F. Christiansen. Acute and chronic behavioral effects of kelp gull micropredation on southern right whale mother-calf pairs off Península Valdés, Argentina. Marine Ecology Progress Series, 668:133–148, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/668/m668p133.pdf.

Auch:2023:OSM

[ASS^{+}23] D. Auch, V. Steinen, L. Steckhan, R. Koppelmann, S. Yari, V. Mohrholz, A. Schukat, M. Fernández-Méndez, L. R. Kittu, and M. A. Peck. Oceanographic structuring of the mucousmesh grazer community in the Humboldt Current off Peru. Marine Ecology Progress Series, 725:29–44, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m725p029.pdf.

Andres:2024:TSM

[ASY^{+}24] A. M. Andres, E. Slesinger, R. E. Young, G. K. Saba, V. S. Saba, B. A. Phelan, J. Rosendale, D. Wieczorek, C. F. White, and B. A. Seibel. Thermal sensitivity of metabolic performance in Squalus acanthias: efficacy of aerobic scope as a predictor of viable thermal habitat. Marine Ecology Progress Series, 738:161–185, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/738/m738p161.pdf.

Allegue:2020:HSR

[ATLT20] H. Allegue, A. C. Thomas, Y. Liu, and A. W. Trites. Harbour seals responded differently to pulses of out-migrating Coho and Chinook smolts. Marine Ecology Progress Series, 647:211–227, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/647/m647p211.pdf.

Anchondo:2024:RNT

[ATR^{+}24] Z. B. Anchondo, A. M. Tracy, A. Raza, K. A. Meckler, and M. B. Ogburn. Reefs in no-take reserves host more oysters, macroparasites, and macrofauna than harvested reefs across an estuarine salinity gradient. Marine Ecology Progress Series, 739:65–83, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m739p065.pdf.

Acevedo:2021:EFS

[AU21] J. Acevedo and J. Urbán. Estimates of Fuegian sprat consumption by humpback whales in the Magellan Strait feeding area as predicted by a bioenergetic model. Marine Ecology Progress Series, 657:223–239, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/657/m657p223.pdf.

Appert:2023:UMA

[AUS^{+}23] C. Appert, V. Udyawer, C. A. Simpfendorfer, M. R. Heupel, M. Scott, L. M. Currey-Randall, A. R. Harborne, F. Jaine, and A. Chin. Use, misuse, and ambiguity of indices of residence in acoustic telemetry studies. Marine Ecology Progress Series, 714:27–44, July 13, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/714/m714p027.pdf.

Alsuwaiyan:2021:GVR

[AVFD^{+}21] N. A. Alsuwaiyan, S. Vranken, K. Filbee-Dexter, M. Cambridge, M. A. Coleman, and T. Wernberg. Genotypic variation in response to extreme events may facilitate kelp adaptation under future climates. Marine Ecology Progress Series, 672:111–121, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m672p111.pdf.

Agius:2023:ITE

[AWB23] A. Agius, L. S. Wright, and J. A. Borg. Impacts of thermal effluent on Posidonia oceanica and associated macrofauna. Marine Ecology Progress Series, 707:15–29, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/707/
m707p015.pdf.

Ault:2021:OBC

[AWC21] E. Ault, S. Webb, and D. Cox. Offshore behavioral contingent of an estuarine fish population, common snook Centropomus undecimalis. Marine Ecology Progress Series, 669:

175–189, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/669/m669p175.pdf.

Alongi:2024:BCB

[AZ24] D. M. Alongi and M. Zimmer. Blue carbon biomass stocks but not sediment stocks or burial rates exhibit global patterns in re-established mangrove chronosequences: a metaanalysis. Marine Ecology Progress Series, 733:27–42, April 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2024/733/m733p027.pdf.

Blanchet:2020:SUS

[BAAR20] M. A. Blanchet, J. Aars, M. Andersen, and H. Routti. Spaceuse strategy affects energy requirements in Barents Sea polar bears. Marine Ecology Progress Series, 639:1–19, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bass:2024:IEF

[BAB^{+}24] A. L. Bass, S. C. Anderson, A. W. Bateman, B. M. Connors, M. A. Peña, S. Li, K. H. Kaukinen, D. A. Patterson, S. G. Hinch, and K. M. Miller. Intrinsic and extrinsic factors associated with the spatio-temporal distribution of infectious agents in early marine Chinook and Coho salmon. Marine Ecology Progress Series, 736:107–127, May 23, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m736p107.pdf.

Berg:2020:RRH

[BAF20] F. Berg, L. Andersson, and A. Folkvord. Respiration rates of herring larvae at different salinities, and effects of previous environmental history. Marine Ecology Progress Series, 650:141–152, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/650/m650p141.pdf.

Blondin:2022:LDM

[BAH^{+}22] H. E. REVIEW Blondin, K. C. Armstrong, E. L. Hazen, W. K. Oestreich, B. S. Santos, D. E. Haulsee, C. S. Mikles, C. J. Knight, A. E. Bennett, and L. B. Crowder. Land-dependent marine species face climate-driven impacts on land and at sea. Marine Ecology Progress Series, 699:181–198, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m699p181.pdf.

Boldreel:2023:MDD

[BAHG23] E. H. Boldreel, K. M. Attard, K. Hancke, and R. N. Glud. Microbial degradation dynamics of farmed kelp deposits from Saccharina latissima and Alaria esculenta. Marine Ecology Progress Series, 709:1–15, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Baird:2022:DFA

[Bai22] P. Baird. Diatoms and fatty acid production in Arctic and estuarine ecosystems — a reassessment of marine food webs, with a focus on the timing of shorebird migration. Marine Ecology Progress Series, 688:173–196, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/688/
m688p173.pdf.

Beyer:2021:ZOM

[BAS21] S. G. Beyer, S. H. Alonzo, and S. M. Sogard. Zero, one or more broods: reproductive plasticity in response to temperature, food, and body size in the live-bearing rosy rockfish Sebastes rosaceus. Marine Ecology Progress Series, 669: 151–173, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/669/m669p151.pdf.

Bornman:2023:EZC

[BAS23] E. Bornman, J. B. Adams, and N. A. Strydom. Estuarine zooplankton communities shift but are not depleted during recurrent harmful algal blooms of Heterosigma akashiwo in the Sundays Estuary, South Africa. Marine Ecology Progress Series, 723:19–36, November 23, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/723/m723p019.pdf.

Briggs:2024:JHT

[BASM^{+}24] M. Briggs, J. Alfaro-Shigueto, J. C. Mangel, N. Acuña-Perales, S. Pingo, A. Jimenez, and A. F. Johnson. Juvenile hawksbill turtle Eretmochelys imbricata movement patterns in Sechura Bay, Peru, assessed with satellite tracking. Marine Ecology Progress Series, 738:243–252, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/738/
m738p243.pdf.

Berger:2024:EMRa

[BB24] A. C. Berger and P. Berg. Eelgrass meadow response to heat stress. I. Temperature threshold for ecosystem production derived from in situ aquatic eddy covariance measurements. Marine Ecology Progress Series, 736:35–46, May 23, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m736p035.pdf.

Bryndum-Buchholz:2020:CCI

[BBBT^{+}20] A. Bryndum-Buchholz, D. G. Boyce, D. P. Tittensor, V. Christensen, D. Bianchi, and H. K. Lotze. Climate-change impacts and fisheries management challenges in the North Atlantic Ocean. Marine Ecology Progress Series, 648:1–17, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Brambilla:2022:SCT

[BBD^{+}22] V. Brambilla, M. Barbosa, I. Dehnert, J. Madin, D. Maggioni, C. Peddie, and M. Dornelas. Shaping coral traits: plasticity more than filtering. Marine Ecology Progress Series, 692:53–65, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/692/m692p053.pdf.

Barlow:2020:LTC

[BBEF^{+}20] D. R. Barlow, K. S. Bernard, P. Escobar-Flores, D. M. Palacios, and L. G. Torres. Links in the trophic chain: modeling functional relationships between in situ oceanography, krill, and blue whale distribution under different oceanographic regimes. Marine Ecology Progress Series, 642:207–225, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m642p207.pdf.

Buckingham:2022:IVN

[BBG^{+}22] L. Buckingham, M. I. Bogdanova, J. A. Green, R. E. Dunn, S. Wanless, S. Bennett, R. M. Bevan, A. Call, M. Canham, C. J. Corse, M. P. Harris, C. J. Heward, D. C. Jardine, J. Lennon, D. Parnaby, C. P. F. Redfern, L. Scott, R. L. Swann, R. M. Ward, E. D. Weston, R. W. Furness, and F. Daunt. Interspecific variation in non-breeding aggregation: a multi-colony tracking study of two sympatric seabirds. Marine Ecology Progress Series, 684:181–197, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m684p181.pdf.

Burton:2023:ELF

[BBGM23] C. Burton, P. J. Bouchet, P. Gill, and S. A. Marley. Evidence of likely foraging by pygmy blue whales in the timor trough during the late austral winter and early austral spring. Marine Ecology Progress Series, 718:99–117, September 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2023/718/m718p099.pdf.

Blincow:2020:SEN

[BBH^{+}20] K. M. Blincow, P. G. Bush, S. A. Heppell, C. M. McCoy, B. C. Johnson, C. V. Pattengill-Semmens, S. S. Heppell, S. J. Stevens-McGeever, L. Whaylen, K. Luke, and B. X. Semmens. Spatial ecology of Nassau grouper at home reef sites: using acoustic telemetry to track a large, long-lived epinephelid across multiple years (2005–2008). Marine Ecology Progress Series, 655:199–214, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m655p199.pdf.

Bosley:2020:RGD

[BBKW20] K. L. Bosley, K. M. Bosley, A. A. Keller, and C. E. Whitmire. Relating groundfish diversity and biomass to deepsea corals and sponges using trawl survey catch data. Marine Ecology Progress Series, 646:127–143, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/646/m646p127.pdf.

Biddlecombe:2021:ESI

[BBL^{+}21a] B. A. Biddlecombe, E. M. Bayne, N. J. Lunn, D. McGeachy, and A. E. Derocher. Effects of sea ice fragmentation on polar bear migratory movement in Hudson Bay. Marine Ecology Progress Series, 666:231–241, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/666/m666p231.pdf.

Brodeur:2021:DFP

[BBL^{+}21b] R. D. Brodeur, T. W. Buckley, G. M. Lang, D. L. Draper, J. C. Buchanan, and R. E. Hibpshman. Demersal fish predators of gelatinous zooplankton in the Northeast Pacific Ocean. Marine Ecology Progress Series, 658:89–104, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m658p089.pdf.

Berger:2024:EMRb

[BBM^{+}24] A. C. Berger, P. Berg, K. J. McGlathery, L. R. Aoki, and K. Kerns. Eelgrass meadow response to heat stress. II. Impacts of ocean warming and marine heatwaves measured by novel metrics. Marine Ecology Progress Series, 736:47–62, May 23, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m736p047.pdf.

Blakeslee:2020:HSA

[BBMM20] A. M. H. Blakeslee, R. B. Barnard, K. Matheson, and C. H. McKenzie. Host-switching among crabs: species introduction results in a new target host for native parasites. Marine Ecology Progress Series, 636:91–106, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/636/
m636p091.pdf.

Barile:2021:TAO

[BBPO21] C. Barile, S. Berrow, G. Parry, and J. O’Brien. Temporal acoustic occurrence of sperm whales Physeter macrocephalus and long-finned pilot whales Globicephala melas off western Ireland. Marine Ecology Progress Series, 661:203– 227, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/661/m661p203.pdf. See corrigendum

[Ano22a].

Bishop:2024:FNS

[BBR24] A. M. Bishop, B. D. Barst, and L. D. Rea. Foraging niche, sex-specific resource partitioning, and mercury concentrations of male Steller sea lions in Alaska. Marine Ecology Progress Series, 736:147–165, May 23, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m736p147.pdf.

Butler:2022:SFI

[BBRM22] C. B. Butler, J. Butler, E. P. Ross, and T. R. Matthews. Seasonal and fishery impacts on the nutritional condition of the Caribbean spiny lobster Panulirus argus in Florida, USA. Marine Ecology Progress Series, 696:43–56, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m696p043.pdf.

Beauchard:2021:GAD

[BBS^{+}21] O. Beauchard, A. Brind’Amour, M. Schratzberger, P. Laffargue, N. T. Hintzen, P. J. Somerfield, and G. Piet. A generic approach to develop a trait-based indicator of trawling-induced disturbance. Marine Ecology Progress Series, 675:35–52, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/675/m675p035.pdf.

Benoit-Bird:2020:CTI

[BBSM^{+}20] K. J. Benoit-Bird, B. L. Southall, M. A. Moline, D. E. Claridge, C. A. Dunn, K. A. Dolan, and D. J. Moretti. Critical threshold identified in the functional relationship between beaked whales and their prey. Marine Ecology Progress Series, 654:1–16, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bajo:2020:ETP

[BCC20] A. Bernal Bajo, L. R. Castro, and D. Costalago. Elucidating trophic pathways of the most abundant fish larvae in northern Patagonia using δ ^{13}C and δ ^{15}N isotopes. Marine Ecology Progress Series, 650:253–267, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/650/m650p253.pdf.

Bolt:2022:UBF

[BCP^{+}22] M. H. Bolt, C. T. Callaghan, A. G. B. Poore, A. Vergés, and C. J. Roberts. Using the background of fish photographs to quantify habitat composition in marine ecosystems. Marine Ecology Progress Series, 688:167–172, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m688p167.pdf.

Beukema:2020:HCM

[BD20] J. J. Beukema and R. Dekker. Half a century of monitoring macrobenthic animals on tidal flats in the Dutch Wadden Sea. Marine Ecology Progress Series, 656:1–18, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Booth:2020:LMP

[BDB^{+}20] D. T. Booth, A. Dunstan, I. Bell, R. Reina, and J. Tedeschi. Low male production at the world’s largest green turtle rookery. Marine Ecology Progress Series, 653:181–190, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/653/m653p181.pdf.

Bost:2022:FTI

[BDC^{+}22] C. A. Bost, K. Delord, Y. Cherel, C. M. Miskelly, A. Carravieri, P. Bustamante, J. P. Y. Arnould, and A. Fromant. Foraging trips and isotopic niche of chick-rearing South Georgian diving petrels from the Kerguelen Islands. Marine Ecology Progress Series, 689:169–177, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/689/
m689p169.pdf.

Bendtsen:2024:PCC

[BDJ^{+}24] J. Bendtsen, N. Daugbjerg, R. S. Jensen, M. C. Brady, M. H. Nielsen, J. L. S. Hansen, and K. Richardson. Phytoplankton community changes in relation to nutrient fluxes along a quasi-stationary front. Marine Ecology Progress Series, 727: 67–80, January 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m727p067.pdf.

Brewton:2020:TER

[BDS^{+}20] R. A. Brewton, C. H. Downey, M. K. Streich, J. J. Wetz, M. J. Ajemian, and G. W. Stunz. Trophic ecology of red snapper Lutjanus campechanus on natural and artificial reefs: interactions between annual variability, habitat, and ontogeny. Marine Ecology Progress Series, 635:105–122, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/635/
m635p105.pdf.

Beckley:2021:MLR

[BE21] B. A. Beckley and M. S. Edwards. Mechanisms leading to recruitment inhibition of giant kelp Macrocystis pyrifera by an understory alga. Marine Ecology Progress Series, 657: 59–71, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/657/m657p059.pdf.

Bittick:2020:HSS

[BFC^{+}20] S. J. Bittick, C. R. Fong, R. J. Clausing, J. D. Harvey, T. M. Johnson, T. A. Frymann, and P. Fong. Herbivory strength is similar or even greater in algal- compared to coral-dominated habitats on a recovering coral reef. Marine Ecology Progress Series, 634:225–229, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2019/634/m634p225.pdf.

Barrionuevo:2023:WAC

[BFQ^{+}23] M. Barrionuevo, E. Frere, F. Quintana, J. Ciancio, N. Marchisio, and S. Lisovski. Within- and among-colony variation in non-breeding dispersion of Magellanic penguins breeding along the coast of Argentina. Marine Ecology Progress Series, 721:151–160, October 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/721/m721p151.pdf.

Belanger:2020:LGR

[BG20] D. Bélanger and P. Gagnon. Low growth resilience of subarctic rhodoliths (Lithothamnion glaciale) to coastal eutrophication. Marine Ecology Progress Series, 642:117–132, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/642/
m642p117.pdf.

Belanger:2021:HGR

[BG21] D. Bélanger and P. Gagnon. High growth resilience of subarctic rhodoliths (Lithothamnion glaciale) to ocean warming and chronic low irradiance. Marine Ecology Progress Series, 663:77–97, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/663/m663p077.pdf.

Bubley:2020:IDH

[BGDJ20] W. J. Bubley, B. Galuardi, A. W. Dukes, and W. E. Jenkins. Incorporating depth into habitat use descriptions for sailfish Istiophorus platypterus and habitat overlap with other billfishes in the western North Atlantic. Marine Ecology Progress Series, 638:137–148, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m638p137.pdf.

Bacheler:2022:CVT

[BGG^{+}22] N. M. Bacheler, Z. D. Gillum, K. C. Gregalis, E. P. Pickett, C. M. Schobernd, Z. H. Schobernd, B. Z. Teer, T. I. Smart, and W. J. Bubley. Comparison of video and traps for detecting reef fishes and quantifying species richness in the continental shelf waters of the southeast USA. Marine Ecology Progress Series, 698:111–123, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m698p111.pdf.

Brustolin:2022:ISD

[BGGH^{+}22] M. Colossi Brustolin, R. V. Gladstone-Gallagher, J. Hewitt, A. M. Lohrer, and S. F. Thrush. The importance of shell debris for within-patch heterogeneity and disturbance-recovery dynamics of intertidal macrofauna. Marine Ecology Progress Series, 700:53–64, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/700/m700p053.pdf.

Birkeland:2021:DRC

[BGL^{+}21] C. Birkeland, A. Green, A. Lawrence, G. Coward, M. Vaeoso, and D. Fenner. Different resiliencies in coral communities over ecological and geological time scales in American Samoa. Marine Ecology Progress Series, 673:55–68, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m673p055.pdf.

Bedolla-Guzman:2021:YRN

[BGMAM^{+}21] Y. Bedolla-Guzmán, J. F. Masello, A. Aguirre-Muñoz, B. E. Lavaniegos, C. C. Voigt, J. Gómez-Gutiérrez, L. Sánchez-Velasco, C. J. Robinson, and P. Quillfeldt. Year-round niche segregation of three sympatric Hydrobates storm-petrels from Baja California Peninsula, Mexico, Eastern Pacific. Marine Ecology Progress Series, 664:207–225, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/664/
m664p207.pdf.

Becerril-Garcia:2020:SMR

[BGMRGM^{+}20] E. E. Becerril-García, R. O. Martínez-Rincón, F. Galván-Magaña, O. Santana-Morales, and E. M. Hoyos-Padilla. Statistical modelling reveals spatial, temporal, and environmental preferences for white sharks at an oceanic aggregation site. Marine Ecology Progress Series, 655:171–183, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/655/
m655p171.pdf.

Barnes:2022:CBC

[BH22] A. J. Barnes and J. M. Hill. Copper barriers can cause behavioral artifacts in experiments with marine snails. Marine Ecology Progress Series, 685:127–136, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/685/
m685p127.pdf.

Burns:2022:GSH

[BHQ^{+}22] N. M. Burns, C. R. Hopkins, M. M. Quaggiotto, D. J. Mc-Cafferty, and D. M. Bailey. Grey seal Halichoerus grypus breeding sites contribute substantial carrion biomass to the Firth of Forth. Marine Ecology Progress Series, 691:191– 201, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/691/m691p191.pdf.

Bernal-Ibanez:2023:PSM

[BICQG23] A. Bernal-Ibáñez, E. Cacabelos, E. Quintano, and I. Gestoso. Prior stress by marine heatwaves and micro-habitat fragmentation drive the colonisation of epifaunal assemblages in marine forests. Marine Ecology Progress Series, 722:89–105, November 9, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/722/m722p089.pdf.

Benjamin:2023:DRP

[BJH^{+}23] E. D. Benjamin, A. Jeffs, S. J. Handley, T. A. Toone, and J. R. Hillman. Determining restoration potential by transplanting mussels of different size classes over a range of aerial exposures. Marine Ecology Progress Series, 713:71– 81, June 29, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/713/m713p071.pdf.

Baumann:2022:IHE

[BJM^{+}22] H. Baumann, L. F. Jones, C. S. Murray, S. A. Siedlecki, M. Alexander, and E. L. Cross. Impaired hatching exacerbates the high CO_{2} sensitivity of embryonic sand lance Ammodytes dubius. Marine Ecology Progress Series, 687:147–162, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m687p147.pdf.

Bacus:2023:EOA

[BK23] S. C. Bacus and A. L. Kelley. Effects of ocean acidification and ocean warming on the behavior and physiology of a subarctic, intertidal grazer. Marine Ecology Progress Series, 711: 31–45, May 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/711/m711p031.pdf.

Botha:2020:GVS

[BKA^{+}20] J. A. Botha, S. P. Kirkman, J. P. Y. Arnould, A. T. Lombard, G. J. G. Hofmeyr, M. A. Meÿer, P. G. H. Kotze, and P. A. Pistorius. Geographic variation in at-sea movements, habitat use and diving behaviour of female Cape fur seals. Marine Ecology Progress Series, 649:201–218, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/649/
m649p201.pdf.

Burbank:2024:OSI

[BKH^{+}24] J. Burbank, B. Kelly, A. Hanke, H. P. Benoît, and M. Power. Ontogenetic shifts in isotope-inferred trophic ecology of Atlantic bluefin tuna Thunnus thynnus in Atlantic Canada. Marine Ecology Progress Series, 744:147–160, September 5, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m744p147.pdf.

Beazley:2021:CCW

[BKM^{+}21] L. Beazley, E. Kenchington, F. J. Murillo, D. Brickman, Z. Wang, A. J. Davies, E. M. Roberts, and H. T. Rapp. Climate change winner in the deep sea? Predicting the impacts of climate change on the distribution of the glass sponge Vazella pourtalesii. Marine Ecology Progress Series, 657:1–23, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Baines:2021:PAR

[BKR^{+}21] M. Baines, N. Kelly, M. Reichelt, C. Lacey, S. Pinder, S. Fielding, E. Murphy, P. Trathan, M. Biuw, U. Lindstrøm, B. A. Krafft, and J. A. Jackson. Population abundance of recovering humpback whales Megaptera novaeangliae and other baleen whales in the Scotia Arc, South Atlantic. Marine Ecology Progress Series, 676:77–94, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/676/m676p077.pdf.

Bommarito:2022:CES

[BKT^{+}22] C. Bommarito, M. Khosravi, D. W. Thieltges, C. Pansch, T. Hamm, F. Pranovi, and J. Vajedsamiei. Combined effects of salinity and trematode infections on the filtration capacity, growth and condition of mussels. Marine Ecology Progress Series, 699:33–44, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m699p033.pdf.

Belleza:2021:EDC

[BKTN21] D. F. C. Belleza, Y. Kawabata, T. Toda, and G. N. Nishihara. Effects of dead conspecifics, hunger states, and seasons on the foraging behavior of the purple urchin Heliocidaris crassispina. Marine Ecology Progress Series, 664:133– 148, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/664/m664p133.pdf.

Bickell:2024:ECS

[BLA^{+}24] A. L. Bickell, E. G. Lim, P. Amos, Q. Anderson, D. Currie-Olsen, S. R. Gray, A. Hann, J. Lambe, C. McKinnon, S. E. Sproule, and I. M. Côté. Ecological consequences of supernumerary arms in eastern Pacific sea stars. Marine Ecology Progress Series, 739:147–155, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/739/
m739p147.pdf.

Bedgood:2023:SAM

[BLB23] S. A. Bedgood, S. T. Levell, and M. E. S. Bracken. Sea anemone microhabitats enhance the diversity and biomass of mobile invertebrates on temperate rocky shores. Marine Ecology Progress Series, 715:57–68, July 27, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/715/
m715p057.pdf.

Burnsed:2020:RME

[BLBBL20] S. Walters Burnsed, S. Lowerre-Barbieri, J. Bickford, and E. Hoerl Leone. Recruitment and movement ecology of red drum Sciaenops ocellatus differs by natal estuary. Marine Ecology Progress Series, 633:181–196, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/633/
m633p181.pdf.

Bonnet-Lebrun:2020:HPF

[BLCC^{+}20] A. S. Bonnet-Lebrun, P. Catry, T. J. Clark, L. Campioni, A. Kuepfer, M. Tierny, E. Kilbride, and E. D. Wakefield. Habitat preferences, foraging behaviour and bycatch risk among breeding sooty shearwaters Ardenna grisea in the Southwest Atlantic. Marine Ecology Progress Series, 651:163–181, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m651p163.pdf.

Beatty:2022:EPW

[BLE^{+}22] W. S. Beatty, P. R. Lemons, J. P. Everett, C. J. Lewis, R. L. Taylor, R. J. Lynn, S. A. Sethi, L. Quakenbush, J. J. Citta, M. L. Kissling, N. Kryukova, and J. K. Wenburg. Estimating Pacific walrus abundance and survival with multievent mark-recapture models. Marine Ecology Progress Series, 697: 167–182, September 22, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m697p167.pdf. See corri-

gendum [Ano22b].

Besnard:2024:RPH

[BLGM^{+}24] L. Besnard, G. Le Croizier, F. Galván-Magaña, M. Mathieu-Resuge, E. Kraffe, R. O. Martínez-Rincón, F. Le Grand, A. Bideau, and G. Schaal. Resource partitioning in hammerhead shark species out-migrating from coastal ecosystems in the Gulf of California. Marine Ecology Progress Series, 732: 101–117, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/732/m732p101.pdf.

Barbut:2024:RLC

[BLVL24] L. Barbut, S. Lehuta, F. A. M. Volckaert, and G. Lacroix. REVIEW: Lessons from the calibration and sensitivity analysis of a fish larval transport model. Marine Ecology Progress Series, 731:67–88, March 13, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m731p067.pdf.

Bengtsson:2024:SPN

[BLW^{+}24] O. Bengtsson, C. Lydersen, J. M. WEsławski, J. Berge, and K. M. Kovacs. Summer and polar night diets of polar cod Boreogadus saida and Atlantic cod Gadus morhua in Kongsfjorden, Svalbard, Norway. Marine Ecology Progress Series, 747:117–132, October 17, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m747p117.pdf.

Butler:2020:MDD

[BLWJ20] C. L. Butler, V. L. Lucieer, S. J. Wotherspoon, and C. R. Johnson. Multi-decadal decline in cover of giant kelp Macrocystis pyrifera at the southern limit of its Australian range.

Marine Ecology Progress Series, 653:1–18, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Buchheister:2024:ETD

[BMB24] A. Buchheister, P. McElhany, and E. P. Bjorkstedt. Evaluating the time to detect biological effects of ocean acidification and warming: an example using simulations of purple sea urchin settlement. Marine Ecology Progress Series, 738: 133–149, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/738/m738p133.pdf.

Benti:2021:IBR

[BMBC21] B. Benti, P. J. O. Miller, M. Biuw, and C. Curé. Indication that the behavioural responses of humpback whales to killer whale sounds are influenced by trophic relationships. Marine Ecology Progress Series, 660:217–232, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m660p217.pdf.

Baker:2024:SMN

[BMF24] K. D. Baker, D. R. J. Mullowney, and S. Fulton. Spatiotemporal modelling of northern shrimp Pandalus borealis distribution patterns throughout Canada’s subarctic and Arctic regions. Marine Ecology Progress Series, 740:79–93, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m740p079.pdf.

Beauchard:2022:BSS

[BMK^{+}22] O. Beauchard, S. Mestdagh, L. Koop, T. Ysebaert, and P. M. J. Herman. Benthic synecology in a soft sediment shelf: habitat contrasts and assembly rules of life strategies. Marine Ecology Progress Series, 682:31–50, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/682/
m682p031.pdf.

Barnard:2022:IYE

[BMKB22] R. B. Barnard, C. S. Moore, C. L. Keogh, and A. M. H. Blakeslee. If you encyst: evidence of parasite escape and host switching among three co-occurring crabs. Marine Ecology Progress Series, 697:67–80, September 22, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/697/
m697p067.pdf.

Bargnesi:2022:NTC

[BML^{+}22] F. Bargnesi, S. Moro, A. Leone, I. Giovos, and F. Ferretti. New technologies can support data collection on endangered shark species in the Mediterranean Sea. Marine Ecology Progress Series, 689:57–76, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/689/m689p057.pdf.

Barker:2021:EFT

[BMM21] C. Barker, C. J. Monaco, and C. D. McQuaid. Exposure to fluctuating temperature increases thermal sensitivity in two lineages of the intertidal mussel Perna perna. Marine Ecology Progress Series, 668:85–95, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/668/m668p085.pdf.

Balboa:2024:ORH

[BMO^{+}24] W. Balboa, D. McDonald, Z. Olsen, D. Williford, J. Harper, and J. Anderson. Oyster reefs as habitat for aquatic macrofauna in a Gulf of Mexico estuary: biotic complexity at spatial, ecological, and demographic scales. Marine Ecology Progress Series, 746:49–66, October 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m746p049.pdf.

Blair:2021:SVE

[BMOW21] H. B. Blair, J. L. Miksis-Olds, and J. D. Warren. Spatial variability of epi- and mesopelagic 38 kHz backscatter from nekton and macrozooplankton across the southeastern US shelf break. Marine Ecology Progress Series, 669:33–50, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/669/
m669p033.pdf.

Burn:2022:CIC

[BMP^{+}22] D. Burn, S. Matthews, C. Pisapia, A. S. Hoey, and M. S. Pratchett. Changes in the incidence of coral injuries during mass bleaching across Australia’s Coral Sea Marine Park. Marine Ecology Progress Series, 682:97–109, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/682/
m682p097.pdf.

Bordeyne:2020:MAP

[BMPD20] F. Bordeyne, A. Migné, M. Plus, and D. Davoult. Modelling the annual primary production of an intertidal brown algal community based on in situ measurements. Marine Ecology Progress Series, 656:95–107, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/656/m656p095.pdf.

Boisseau:2021:MWB

[BMS^{+}21] O. Boisseau, T. McGarry, S. Stephenson, R. Compton, A. C. Cucknell, C. Ryan, R. McLanaghan, and A. Moscrop. Minke whales Balaenoptera acutorostrata avoid a 15 kHz acoustic deterrent device (ADD). Marine Ecology Progress Series, 667:191–206, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/667/m667p191.pdf.

Blawas:2024:ADL

[BMS^{+}24] A. M. Blawas, L. E. Miller, J. M. Shearer, W. R. Cioffi, D. L. Webster, Z. T. Swaim, H. J. Foley, D. M. Waples, N. J. Quick, D. P. Nowacek, and A. J. Read. Aerobic dive limit in shortfinned pilot whales Globicephala macrorhynchus: an assessment of behavioral criteria. Marine Ecology Progress Series, 744:161–170, September 5, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/744/m744p161.pdf.

Balbar:2024:CAE

[BMW24] A. C. Balbar, A. Metaxas, and Y. Wu. Comparing approaches for estimating ecological connectivity at a local scale in a marine system. Marine Ecology Progress Series, 731: 51–65, March 13, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m731p051.pdf.

Black:2021:EEE

[BNH^{+}21] J. A. Black, A. B. Neuheimer, P. L. Horn, D. M. Tracey, and J. C. Drazen. Environmental, evolutionary, and ecological drivers of slow growth in deep-sea demersal teleosts. Marine Ecology Progress Series, 658:1–26, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Brereton:2020:MSM

[BNR20] A. Brereton, Y. Noh, and S. Raasch. Modelling a simple mechanism for the formation of phytoplankton thin layers using large-eddy simulation: in situ growth. Marine Ecology Progress Series, 653:77–90, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/653/m653p077.pdf.

Bezuidenhout:2024:HIG

[BNS^{+}24] K. Bezuidenhout, R. Nel, D. Schoeman, W. S. Grant, and L. Hauser. Historical imprints on genetic population structure in direct-developing beach whelks (Bullia digitalis). Marine Ecology Progress Series, 736:63–80, May 23, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/736/
m736p063.pdf.

Bezuidenhout:2021:HDB

[BNSH21] K. Bezuidenhout, R. Nel, D. S. Schoeman, and L. Hauser. Historic dispersal barriers determine genetic structure and connectivity in a supratidal sandy-beach brooder. Marine Ecology Progress Series, 674:1–13, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Brown:2023:DHU

[BOY^{+}23] T. M. Brown, W. R. Ogloff, D. J. Yurkowski, J. Coffey, G. Stenson, and B. Sjare. Divergent habitat use and the influence of sea ice concentration on the movement behaviour of ringed seals Pusa hispida in Labrador, Canada. Marine Ecology Progress Series, 710:137–153, May 4, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m710p137.pdf.

Brodersen:2024:EIT

[BP24] K. E. Brodersen and M. F. Pedersen. Effects of increasing temperature on the photosynthetic activity and oxygen balance of sheath-covered seagrass Zostera marina seeds. Marine Ecology Progress Series, 743:15–24, August 22, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m743p015.pdf.

Bonel:2021:ESI

[BPA21] N. Bonel, J. P. Pointier, and P. Alda. Environmental stressors induced strong small-scale phenotypic differentiation in a wide-dispersing marine snail. Marine Ecology Progress Series, 674:143–162, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/674/m674p143.pdf.

Bandimere:2021:EEM

[BPS^{+}21] A. Bandimere, F. V. Paladino, J. R. Spotila, A. Panagopoulou, J. Diéguez-Uribeondo, and P. Santidrián Tomillo. Effects of egg mass and local climate on morphology of East Pacific leatherback turtle Dermochelys coriacea hatchlings in Costa Rica. Marine Ecology Progress Series, 669:191–200, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/669/m669p191.pdf.

Bernardo:2024:WCL

[BPS^{+}24a] C. P. Bernardo, L. Pecuchet, J. Santos, A. V. Dolgov, M. Fossheim, B. Husson, and R. Primicerio. Warming changes the life history composition of marine fish communities at high latitudes. Marine Ecology Progress Series, 732:119– 133, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m732p119.pdf.

Bury:2024:SOH

[BPS^{+}24b] S. J. Bury, K. J. Peters, A. J. M. Sabadel, K. St John Glew, C. Trueman, M. B. Wunder, M. R. D. Cobain, N. Schmitt, D. Donnelly, S. Magozzi, K. Owen, J. C. S. Brown, P. Escobar-Flores, R. Constantine, R. L. O’Driscoll, M. Double, N. Gales, S. Childerhouse, and M. H. Pinkerton. Southern Ocean humpback whale trophic ecology. I. Combining multiple stable isotope methods elucidates diet, trophic position and foraging areas. Marine Ecology Progress Series, 734:123–155, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m734p123.pdf.

Bulleri:2022:CES

[BPSB^{+}22] F. Bulleri, C. Pozas-Schacre, H. Bischoff, L. Bramanti, S. D’agata, J. Gasc, and M. M. Nugues. Compounded effects of sea urchin grazing and physical disturbance on macroalgal canopies in the lagoon of Moorea, French Polynesia. Marine Ecology Progress Series, 697:45–56, September 22, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/697/m697p045.pdf.

Brandl:2022:MRM

[BQC^{+}22] S. J. Brandl, C. N. Quigley, J. M. Casey, A. Mercière, N. M. D. Schiettekatte, T. Norin, V. Parravicini, and I. M. Côté. Metabolic rates mirror morphological and behavioral differences in two sand-dwelling coral reef gobies. Marine Ecology Progress Series, 684:79–90, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/684/m684p079.pdf.

Beauvieux:2022:EAT

[BQM^{+}22] A. Beauvieux, Q. Queiros, L. Metral, G. Dutto, E. Gasset, F. Criscuolo, J. M. Fromentin, C. Saraux, and Q. Schull. Energy allocation trade-offs between life-history traits in the Mediterranean sardine: an ecophysiological approach. Marine Ecology Progress Series, 701:99–118, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/701/
m701p099.pdf.

Barreto:2024:SSP

[BQM^{+}24] C. R. Barreto, J. P. Quimbayo, T. C. Mendes, C. A. M. M. Cordeiro, and A. A. V. Flores. Spatial structure and potential processes linking fish and benthic communities in a protected reef ecosystem in SE Brazil. Marine Ecology Progress Series, 738:41–55, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/738/m738p041.pdf.

Bagarinao-Regalado:2022:LVE

[BRC22] A. Bagarinao-Regalado and W. L. Campos. Linking variability in early life growth of Sardinella lemuru to changes in habitat conditions in Zamboanga upwelling system, Sulu Sea, the Philippines. Marine Ecology Progress Series, 682: 191–206, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/682/m682p191.pdf.

Bennington:2020:PPP

[BRD20] S. Bennington, W. Rayment, and S. Dawson. Putting prey into the picture: improvements to species distribution models for bottlenose dolphins in Doubtful Sound, New Zealand. Marine Ecology Progress Series, 653:191–204, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/653/
m653p191.pdf.

Brough:2020:SDF

[BRSD20] T. Brough, W. Rayment, E. Slooten, and S. Dawson. Spatiotemporal distribution of foraging in a marine predator: behavioural drivers of hotspot formation. Marine Ecology Progress Series, 635:187–202, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/635/m635p187.pdf.

Brunner:2022:ECC

[BRU^{+}22] C. A. Brunner, G. F. Ricardo, S. Uthicke, A. P. Negri, and M. O. Hoogenboom. Effects of climate change and light limitation on coral recruits. Marine Ecology Progress Series, 690:65–82, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m690p065.pdf.

Bacheler:2020:CRF

[BS20a] N. M. Bacheler and K. W. Shertzer. Catchability of reef fish species in traps is strongly affected by water temperature and substrate. Marine Ecology Progress Series, 642: 179–190, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/642/m642p179.pdf.

Blain:2020:DRF

[BS20b] C. O. Blain and N. T. Shears. Differential response of forestforming seaweeds to elevated turbidity may facilitate ecosystem shifts on temperate reefs. Marine Ecology Progress Series, 641:63–77, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/641/m641p063.pdf.

Baker:2021:PLH

[BS21] R. Baker and M. Sheaves. Predation in littoral habitats is a complex process: Comment on Whitfield (2020). Marine Ecology Progress Series, 662:205–208, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/662/
m662p205.pdf. See [Whi20].

Brahim:2024:BFA

[BSA^{+}24] S. Brahim, T. G. Sobrinho, J. Albino, G. Zanetti, V. S. Amaral, M. T. W. D. Carneiro, J. R. Souza, L. B. Castro, D. Muehe, B. S. Souza, G. P. Brandão, M. Krause, L. H. S. Oliveira, U. P. Coelho, E. E. Silva, I. J. Silva, J. A. Neves, D. Paiva-Tagliari, and K. G. Costa. Benthic fauna along sandy beaches enriched by metals after mining disaster. Marine Ecology Progress Series, 726:181–196, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/726/m726p181.pdf.

Bode:2022:TPD

[BSÁG^{+}22] A. Bode, C. Saavedra, M. Álvarez-González, M. Arregui, M. Arbelo, A. Fernández, L. Freitas, M. A. Silva, R. Prieto, J. M. N. Azevedo, J. Giménez, G. J. Pierce, and M. B. Santos. Trophic position of dolphins tracks recent changes in the pelagic ecosystem of the Macaronesian region (NE Atlantic). Marine Ecology Progress Series, 699:167–180, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m699p167.pdf.

Bertrand:2021:FFL

[BSB^{+}21] P. Bertrand, H. Strøm, J. Bêty, H. Steen, J. Kohler, M. Vihtakari, W. van Pelt, N. G. Yoccoz, H. Hop, S. M. Harris, S. C. Patrick, P. Assmy, A. Wold, P. Duarte, G. Moholdt, and S. Descamps. Feeding at the front line: interannual variation in the use of glacier fronts by foraging black-legged kittiwakes. Marine Ecology Progress Series, 677:197–208, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m677p197.pdf.

Baldanzi:2020:CET

[BSF^{+}20] S. Baldanzi, D. Storch, M. Fusi, N. Weidberg, A. Tissot, S. A. Navarrete, and M. Fernández. Combined effects of temperature and hypoxia shape female brooding behaviors and the early ontogeny of the Chilean kelp crab Taliepus dentatus. Marine Ecology Progress Series, 646:93–107, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/646/

m646p093.pdf. See corrigendum [Ano21a].

Berg:2024:GSA

[BSF24] F. Berg, G. Seljestad, and A. Folkvord. Growth of spring- and autumn-spawned larvae of Atlantic herring Clupea harengus: a long-term experiment mimicking seasonal light conditions. Marine Ecology Progress Series, 741:203–216, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m741p203.pdf.

Barlow:2020:VDB

[BSFM20] J. Barlow, G. S. Schorr, E. A. Falcone, and D. Moretti. Variation in dive behavior of Cuvier’s beaked whales with seafloor depth, time-of-day, and lunar illumination. Marine Ecology Progress Series, 644:199–214, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/644/m644p199.pdf.

Beeken:2021:DLH

[BSH21a] N. S. Beeken, J. D. Selwyn, and J. D. Hogan. Determining the life history strategy of the cryptobenthic reef gobies Coryphopterus hyalinus and C. personatus. Marine Ecology Progress Series, 659:161–173, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/659/m659p161.pdf.

Bierlich:2021:BAP

[BSH^{+}21b] K. C. Bierlich, R. S. Schick, J. Hewitt, J. Dale, J. A. Goldbogen, A. S. Friedlaender, and D. W. Johnston. Bayesian approach for predicting photogrammetric uncertainty in morphometric measurements derived from drones. Marine Ecology Progress Series, 673:193–210, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m673p193.pdf.

Blocker:2023:DDN

[BSM23] A. M. Blöcker, C. Sguotti, and C. Möllmann. Discontinuous dynamics in North Sea cod Gadus morhua caused by ecosystem change. Marine Ecology Progress Series, 713:133– 149, June 29, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m713p133.pdf.

BenTal:2021:HMM

[BSP^{+}21] A. Ben Tal, N. Shenkar, A. Paz, K. Conley, K. Sutherland, and G. Yahel. High mucous-mesh production by the ascidian Herdmania momus. Marine Ecology Progress Series, 663:223–228, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/663/m663p223.pdf.

Bouyoucos:2022:TIP

[BSP^{+}22] I. A. Bouyoucos, C. A. Simpfendorfer, S. Planes, G. D. Schwieterman, O. C. Weideli, and J. L. Rummer. Thermally insensitive physiological performance allows neonatal sharks to use coastal habitats as nursery areas. Marine Ecology Progress Series, 682:137–152, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/682/m682p137.pdf.

Blincow:2022:GAE

[BSRVS22] K. M. Blincow, R. Swalethorp, A. Ramírez-Valdez, and B. X. Semmens. Giant appetites: exploring the trophic ecology of California’ss largest kelp forest predator, the giant sea bass Stereolepis gigas. Marine Ecology Progress Series, 695:157– 171, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m695p157.pdf.

Barton:2020:ITE

[BTA^{+}20] A. D. Barton, F. González Taboada, A. Atkinson, C. E. Widdicombe, and C. A. Stock. Integration of temporal environmental variation by the marine plankton community. Marine Ecology Progress Series, 647:1–16, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Bizien:2024:CTI

[BTB^{+}24] A. Bizien, C. Toumi, A. Boyé, A. Sturbois, M. Le Duff, J. Grall, M. Helias, O. Gauthier, and T. Burel. Contrasted trends of intertidal macroalgal communities and sharp decline of canopy-forming species across two decades. Marine Ecology Progress Series, 742:73–86, August 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/742/
m742p073.pdf.

Baillargeon:2020:IPS

[BTDR20] G. A. Baillargeon, M. F. Tlusty, E. T. Dougherty, and A. L. Rhyne. Improving the productivity-susceptibility analysis to assess data-limited fisheries. Marine Ecology Progress Series, 644:143–156, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/644/m644p143.pdf.

Bassi:2024:OEF

[BTMS24] L. Bassi, R. Tremblay, O. Morissette, and P. Sirois. Otolith elemental fingerprints reveal source-sink dynamics between two Greenland halibut nurseries in the St. Lawrence Estuary and Gulf. Marine Ecology Progress Series, 731:217– 229, March 13, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m731p217.pdf.

Bearham:2020:MDD

[BVD^{+}20] D. Bearham, M. A. Vanderklift, R. A. Downie, D. P. Thomson, and L. A. Clementson. Macrophyte-derived detritus in shallow coastal waters contributes to suspended particulate organic matter and increases growth rates of Mytilus edulis. Marine Ecology Progress Series, 644:91–103, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/644/
m644p091.pdf.

Barbosa:2023:PEE

[BVKF23] A. C. C. Barbosa, C. Vinagre, M. V. Kitahara, and A. A. V. Flores. Priority effects, environmental filtering and neutral coexistence explain large- to small-scale distribution of invasive sun corals in the SW Atlantic. Marine Ecology Progress Series, 707:31–42, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/707/m707p031.pdf.

Bowgen:2022:RPA

[BWC^{+}22] K. M. Bowgen, L. J. Wright, N. A. Calbrade, D. Coker, S. G. Dodd, I. Hainsworth, R. J. Howells, D. S. Hughes, P. Jenks, M. D. Murphy, W. G. Sanderson, R. C. Taylor, and N. H. K. Burton. Resilient protected area network enables species adaptation that mitigates the impact of a crash in food supply. Marine Ecology Progress Series, 681:211–225, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/681/
m681p211.pdf.

Belackova:2023:PPE

[BWGH23] A. Belackova, N. T. Willenbrink, J. M. S. Gonçalves, and B. Horta e Costa. Potential protection effects in a small marine reserve: a demersal community analysis from stereo baited remote underwater video. Marine Ecology Progress Series, 708:79–100, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/708/m708p079.pdf.

Barth:2020:SIV

[BWRP20] A. Barth, R. K. Walter, I. Robbins, and A. Pasulka. Seasonal and interannual variability of phytoplankton abundance and community composition on the Central Coast of California. Marine Ecology Progress Series, 637:29–43, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/637/
m637p029.pdf.

Bortolotto:2023:DHU

[BZT^{+}23] G. A. Bortolotto, A. N. Zerbini, L. Thomas, A. Andriolo, and P. S. Hammond. Distribution and habitat use modelling from satellite tracking data of humpback whales in Brazil agrees with shipboard survey data modelling. Marine Ecology Progress Series, 720:161–174, October 5, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/720/
m720p161.pdf.

Cardenas-Alayza:2022:SSD

[CAAG^{+}22] S. Cárdenas-Alayza, M. J. Adkesson, D. Gutiérrez, H. Demarcq, and Y. Tremblay. Strategies for segregation during foraging in sympatric otariids of the Peruvian upwelling Humboldt Current System. Marine Ecology Progress Series, 702:153–170, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/702/m702p153.pdf.

Camillini:2021:RCM

[CAEG21] N. Camillini, K. M. Attard, B. D. Eyre, and R. N. Glud. Resolving community metabolism of eelgrass Zostera marina meadows by benthic flume-chambers and eddy covariance in dynamic coastal environments. Marine Ecology Progress Series, 661:97–114, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m661p097.pdf.

Cabasan:2021:CIO

[CAP^{+}21] J. P. Cabasan, H. O. Arceo, P. Pata, K. L. Labrador, R. B. Casauay, N. Miller, and M. V. Baria-Rodriguez. Combining information on otolith morphometrics and larval connectivity models to infer stock structure of Plectropomus leopardus in the Philippines. Marine Ecology Progress Series, 679:115–131, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/679/m679p115.pdf.

Curtis:2020:SIA

[CAPS20] J. S. Curtis, M. A. Albins, E. B. Peebles, and C. D. Stallings. Stable isotope analysis of eye lenses from invasive lionfish yields record of resource use. Marine Ecology Progress Series, 637:181–194, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/637/m637p181.pdf.

Carlson:2021:NBP

[CASF21] B. Carlson, M. Awai, W. B. Saunders, and E. C. Franklin. Nautilus belauensis population demographics and trap yields in Palau were similar between surveys in 1982 and 2015. Marine Ecology Progress Series, 670:239–245, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/670/
m670p239.pdf.

Camara:2021:SSE

[CATF^{+}21] E. M. Camara, M. F. Andrade-Tubino, T. P. Franco, L. N. Santos, A. F. G. N. Santos, and F. G. Araújo. Scale sensitivity of environmental effects on the temporal beta diversity of fishes in tropical coastal lagoons. Marine Ecology Progress Series, 658:195–217, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/658/m658p195.pdf.

Canepuccia:2020:TVP

[CAV^{+}20] A. D. Canepuccia, D. Alemany, E. Espinosa Vidal, M. F. Alvarez, and O. O. Iribarne. Temporal variation in positive and negative interactions between marsh herbivores mediated by changes in plant traits. Marine Ecology Progress Series, 634:89–97, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/634/m634p089.pdf.

Cannell:2022:DBB

[CAW^{+}22] B. L. Cannell, P. J. D. Allen, E. M. Wiley, B. Radford, C. A. Surman, and A. R. Ridley. Diet of brown boobies at a globally significant breeding ground is influenced by sex, breeding, sub-colony and year. Marine Ecology Progress Series, 681:227–245, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/681/m681p227.pdf.

Chava:2024:VTC

[CAY24] A. Chava, A. Artemieva, and E. Yakovis. Variation in top-down control of red algae epibiosis in the White Sea. Marine Ecology Progress Series, 744:53–67, September 5, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2024/744/m744p053.pdf.

Christiansen:2022:ECG

[CBB^{+}22a] F. Christiansen, L. Bejder, S. Burnell, R. Ward, and C. Charlton. Estimating the cost of growth in southern right whales from drone photogrammetry data and long-term sighting histories. Marine Ecology Progress Series, 687:173–194, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/687/m687p173.pdf.

Coletto:2022:MPA

[CBB^{+}22b] J. L. Coletto, A. C. Besser, S. Botta, L. A. S. P. Madureira, and S. D. Newsome. Multi-proxy approach for studying the foraging habitat and trophic position of a migratory marine consumer in the southwestern Atlantic Ocean. Marine Ecology Progress Series, 690:147–163, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/690/m690p147.pdf.

Capone:2024:NEC

[CBGB24] H. E. Capone, M. Brandt, P. W. Gabrielson, and J. F. Bruno. Nutrient enrichment can increase the thermal performance of Galápagos seaweeds. Marine Ecology Progress Series, 749:57– 69, November 21, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/749/m749p057.pdf.

Callahan:2021:FWE

[CBHM21] M. W. Callahan, A. H. Beaudreau, R. Heintz, and F. Mueter. First winter energy allocation in juvenile sablefish Anoplopoma fimbria, a fast growing marine piscivore. Marine Ecology Progress Series, 663:145–156, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m663p145.pdf.

Constant:2023:DAG

[CBJ^{+}23] N. Constant, A. B. Bolten, R. A. Johnson, A. M. L. Brooks, and K. A. Bjorndal. Dynamics and aging of green turtle grazing plots at two Caribbean seagrass meadows. Marine Ecology Progress Series, 705:109–125, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/705/m705p109.pdf.

Costa:2020:SEF

[CBKM20] P. L. Costa, L. Bugoni, P. G. Kinas, and L. A. S. P. Madureira. Seabirds, environmental features and the Argentine anchovy Engraulis anchoita in the southwestern Atlantic Ocean. Marine Ecology Progress Series, 651:199–213, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/651/
m651p199.pdf.

Charifson:2022:SRR

[CBP22] D. M. Charifson, P. E. Bourdeau, and D. K. Padilla. Shell remodeling in response to increased risk of predation in a marine snail. Marine Ecology Progress Series, 688:57–67, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/688/m688p057.pdf.

Coffee:2020:WIF

[CBTL20] O. I. Coffee, D. T. Booth, J. A. Thia, and C. J. Limpus. When isotopes fail: importance of satellite telemetry and multi-site validation when estimating the foraging grounds of migratory species. Marine Ecology Progress Series, 633: 197–206, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/633/m633p197.pdf.

Clementi:2021:APR

[CBVA^{+}21] G. M. Clementi, E. A. Babcock, J. Valentin-Albanese, M. E. Bond, K. I. Flowers, M. R. Heithaus, E. R. Whitman, M. P. M. Van Zinnicq Bergmann, T. L. Guttridge, O. R. O’Shea, O. N. Shipley, E. J. Brooks, S. T. Kessel, and D. D. Chapman. Anthropogenic pressures on reef-associated sharks in jurisdictions with and without directed shark fishing. Marine Ecology Progress Series, 661:175–186, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/661/m661p175.pdf.

Chambon:2024:MMD

[CBWH24] J. Chambon, L. Börger, H. Weimerskirch, and J. C. Hennicke. Migratory movements, distribution, habitat preference, and activity patterns of the endangered Abbott’s booby Papasula abbotti. Marine Ecology Progress Series, 743:75– 96, August 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/743/m743p075.pdf.

Cherel:2022:AME

[CC22] Y. Cherel and A. Carrouée. Assessing marine ecosystem complexity: isotopic integration of the trophic structure of seabird communities from the Southern Ocean. Marine Ecology Progress Series, 694:193–208, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/694/m694p193.pdf.

Clarkson:2020:NTT

[CCA^{+}20] J. Clarkson, F. Christiansen, T. Awbery, L. Abbiss, N. Nikpaljevic, and A. Akkaya. Non-targeted tourism affects the behavioural budgets of bottlenose dolphins Tursiops truncatus in the South Adriatic (Montenegro). Marine Ecology Progress Series, 638:165–176, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/638/m638p165.pdf.

Clark:2021:SSG

[CCA^{+}21] B. L. Clark, S. L. Cox, K. M. Atkins, S. Bearhop, A. W. J. Bicknell, T. W. Bodey, I. R. Cleasby, W. J. Grecian, K. C. Hamer, B. R. Loveday, P. I. Miller, G. Morgan, L. Morgan, J. Newton, S. C. Patrick, K. L. Scales, R. B. Sherley, F. Vigfúsdóttir, E. D. Wakefield, and S. C. Votier. Sexual segregation of gannet foraging over 11 years: movements vary but isotopic differences remain stable. Marine Ecology Progress Series, 661:1–16, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Calixto-Cunha:2024:UGS

[CCAdA24] M. Calixto-Cunha, D. F. R. Alves, and A. C. de Almeida. Unraveling global species richness patterns of the Lysmatidae family: a multi-scale and multi-hypothesis ecological approach. Marine Ecology Progress Series, 748:83–98, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/748/m748p083.pdf.

Colominas-Ciuro:2021:DAO

[CCBD^{+}21] R. Colominas-Ciuró, M. Bertellotti, V. L. D’Amico, E. Carabajal, J. Benzal, V. Vidal, M. Motas, M. Santos, N. Coria, and A. Barbosa. Diet, antioxidants and oxidative status in pygoscelid penguins. Marine Ecology Progress Series, 665: 201–216, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/665/m665p201.pdf.

Camps-Castella:2020:ANC

[CCBT^{+}20] J. Camps-Castellà, M. Ballesteros, R. Trobajo, M. Pontes, and P. Prado. Not all nudibranchs are carnivorous: trophic ecology of Polycerella emertoni in the Ebro Delta. Marine Ecology Progress Series, 645:67–82, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/645/m645p067.pdf. See

correction [Ano20].

Caronni:2023:GVN

[CCC^{+}23] S. Caronni, D. G. Calvi, G. Ceccherelli, S. Citterio, M. A. Delaria, R. Gentili, C. Montagnani, A. Navone, P. Panzalis, and D. Basso. Geniculate vs. nongeniculate coralline algae: different responses to stress are also related to morphology. Marine Ecology Progress Series, 705:43–59, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/705/
m705p043.pdf.

Carlton:2023:TRC

[CCJ^{+}23] H. NOTE Carlton, L. Champlin, R. Jeppesen, J. C. Haskins, F. I. Rahman, and E. B. Watson. Tidal restrictions in a central Californian estuarine system are associated with higher mean pH, but increased low-pH exposure. Marine Ecology Progress Series, 703:177–182, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2022/703/m703p177.pdf.

Counsell:2022:IAL

[CCL^{+}22] C. W. W. Counsell, R. R. Coleman, S. S. Lal, B. W. Bowen, E. C. Franklin, A. B. Neuheimer, B. S. Powell, R. J. Toonen, M. J. Donahue, M. A. Hixon, and M. A. McManus. Interdisciplinary analysis of larval dispersal for a coral reef fish: opening the black box. Marine Ecology Progress Series, 684:117–132, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m684p117.pdf.

Castillo:2022:STD

[CCMRII22] J. M. Castillo, G. Curado, A. F. Muñoz-Rodríguez, and M. D. Infante-Izquierdo. Salt tolerance during germination identifies native intertidal plant species at risk under increasing salinity with sea level rise. Marine Ecology Progress Series, 684:57–68, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/684/m684p057.pdf.

Corbeau:2021:DFH

[CCP^{+}21] A. Corbeau, J. Collet, A. Pajot, R. Joo, T. Thellier, and H. Weimerskirch. Differences in foraging habitat result in contrasting fisheries interactions in two albatross populations. Marine Ecology Progress Series, 663:197–208, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/663/
m663p197.pdf.

Casabianca:2020:SED

[CCR^{+}20] S. Casabianca, S. Capellacci, F. Ricci, F. Andreoni, T. Russo, M. Scardi, and A. Penna. Structure and environmental drivers of phytoplanktonic resting stage assemblages in the central Mediterranean Sea. Marine Ecology Progress Series, 639: 73–89, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/639/m639p073.pdf.

Camps-Castella:2020:TPS

[CCRP20] J. Camps-Castellà, J. Romero, and P. Prado. Trophic plasticity in the sea urchin Paracentrotus lividus, as a function of resource availability and habitat features. Marine Ecology Progress Series, 637:71–85, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/637/m637p071.pdf.

Cerritelli:2022:MMM

[CCS^{+}22] G. Cerritelli, P. Casale, D. Sözbilen, S. Hochscheid, P. Luschi, and Y. Kaska. Multidirectional migrations from a major nesting area in Turkey support the widespread distribution of foraging sites for loggerhead turtles in the Mediterranean. Marine Ecology Progress Series, 683:169–177, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/683/
m683p169.pdf.

Cossa:2023:DML

[CCT^{+}23] D. Cossa, M. Cossa, I. Timba, J. Nhaca, A. Macia, and E. Infantes. Drones and machine-learning for monitoring dugong feeding grounds and gillnet fishing. Marine Ecology Progress Series, 716:123–136, August 10, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m716p123.pdf.

Cheng:2021:DJA

[CCW21] H. Cheng, C. C. Chabot, and W. H. Watson III. Distribution of juvenile American horseshoe crabs Limulus polyphemus in the Great Bay Estuary, New Hampshire, USA. Marine Ecology Progress Series, 662:199–203, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/662/
m662p199.pdf.

Cimon:2021:MSD

[CDC21] S. Cimon, A. Deslauriers, and M. Cusson. Multiple stressors and disturbance effects on eelgrass and epifaunal macroinvertebrate assemblage structure. Marine Ecology Progress Series, 657:93–108, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m657p093.pdf.

Chang:2023:OSS

[CDC^{+}23] C. T. Chang, J. C. Drazen, W. C. Chiang, D. J. Madigan, A. B. Carlisle, N. J. Wallsgrove, H. H. Hsu, Y. H. Ho, and B. N. Popp. Ontogenetic and seasonal shifts in diets of sharptail mola Masturus lanceolatus in waters off Taiwan. Marine Ecology Progress Series, 715:113–127, July 27, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/715/
m715p113.pdf.

Caiger:2020:DMA

[CDD^{+}20a] P. E. Caiger, M. J. Dean, A. I. DeAngelis, L. T. Hatch, A. N. Rice, J. A. Stanley, C. Tholke, D. R. Zemeckis, and S. M. Van Parijs. A decade of monitoring Atlantic cod Gadus morhua spawning aggregations in Massachusetts Bay using passive acoustics. Marine Ecology Progress Series, 635:89–103, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m635p089.pdf.

Christiansen:2020:PCR

[CDD^{+}20b] F. Christiansen, S. M. Dawson, J. W. Durban, H. Fearnbach, C. A. Miller, L. Bejder, M. Uhart, M. Sironi, P. Corkeron, W. Rayment, E. Leunissen, E. Haria, R. Ward, H. A. Warick, I. Kerr, M. S. Lynn, H. M. Pettis, and M. J. Moore. Population comparison of right whale body condition reveals poor state of the North Atlantic right whale. Marine Ecology Progress Series, 640:1–16, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Casselberry:2020:NAR

[CDF^{+}20] G. A. Casselberry, A. J. Danylchuk, J. T. Finn, B. M. DeAngelis, A. Jordaan, C. G. Pollock, I. Lundgren, Z. Hillis-Starr, and G. B. Skomal. Network analysis reveals multispecies spatial associations in the shark community of a Caribbean marine protected area. Marine Ecology Progress Series, 633: 105–126, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/633/m633p105.pdf.

Cote:2021:UMD

[CDP^{+}21] D. Cote, J. B. Dempson, M. Piersiak, K. Layton, S. Roul, R. Laing, J. Angnatok, and I. Bradbury. Using movement, diet, and genetic analyses to understand Arctic charr responses to ecosystem change. Marine Ecology Progress Series, 673:135–149, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m673p135.pdf.

Chang:2020:TMM

[CDS20] Y. L. K. Chang, G. Dall’Olmo, and R. Schabetsberger. Tracking the marine migration routes of South Pacific silver eels. Marine Ecology Progress Series, 646:1–12, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Crossett:2023:RRR

[CDS23] D. Crossett, R. Dunmore, and D. Schiel. Reef recovery and resilience: interpreting abiotic effects on critical early life stages of large brown algae. Marine Ecology Progress Series, 710: 27–42, May 4, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/710/m710p027.pdf.

Costa:2022:BCM

[CEAO^{+}22] M. T. Costa, E. Ezcurra, O. Aburto-Oropeza, M. Maltz, K. Arogyaswamy, J. Botthoff, and E. Aronson. Baja California Sur mangrove deep peat microbial communities cycle nitrogen but do not affect old carbon pool. Marine Ecology Progress Series, 695:15–31, August 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Caputo:2021:DPT

[CEG^{+}21] M. Caputo, S. Elwen, T. Gridley, S. A. Kohler, J. P. Roux, P. W. Froneman, and J. J. Kiszka. Dietary plasticity of two coastal dolphin species in the Benguela upwelling ecosystem. Marine Ecology Progress Series, 669:227–240, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/669/
m669p227.pdf.

Celebi-Ergin:2021:IOC

[CEZH21] B. Celebi-Ergin, R. C. Zimmerman, and V. J. Hill. Impact of ocean carbonation on long-term regulation of light harvesting in eelgrass Zostera marina. Marine Ecology Progress Series, 671:111–128, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/671/m671p111.pdf.

Choi:2021:STV

[CFB21] E. S. Choi, L. E. Furtado, and R. S. Burton. Spatial and temporal variation in the species diversity of coastal California fish eggs. Marine Ecology Progress Series, 669:139– 149, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/669/m669p139.pdf.

Cohen:2023:STS

[CFBPH23] R. E. Cohen, K. E. Frasier, S. Baumann-Pickering, and J. A. Hildebrand. Spatial and temporal separation of toothed whales in the western North Atlantic. Marine Ecology Progress Series, 720:1–24, October 5, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Corso:2022:NRO

[CFE^{+}22] J. Corso, B. J. French, C. B. Edwards, N. E. Pedersen, B. J. Zgliczynski, S. Planes, S. Pacala, and S. A. Sandin. Non-random orientation of Pocillopora colonies on forereefs of Moorea, French Polynesia. Marine Ecology Progress Series, 693:177–182, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/693/m693p177.pdf.

Cardozo-Ferreira:2023:SVD

[CFFC^{+}23] G. C. Cardozo-Ferreira, C. E. L. Ferreira, J. H. Choat, T. C. Mendes, R. M. Macieira, C. E. Rezende, J. C. Joyeux, and K. D. Clements. Seasonal variation in diet and isotopic niche of nominally herbivorous fishes in subtropical rocky reefs. Marine Ecology Progress Series, 722:125–143, November 9, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/722/m722p125.pdf.

Castro-Fernandez:2024:VHS

[CFHCS^{+}24] J. Castro-Fernández, H. Hinz, I. Castejon-Silvo, I. Catalán, J. M. Disdier-Gómez, J. Moranta, A. Ruiz-Frau, A. Grau, and J. Terrados. Visual habitat selection by juvenile fish in response to seagrass with different structural complexity. Marine Ecology Progress Series, 738:187–201, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2024/738/m738p187.pdf.

Crossin:2022:PCB

[CFSH22] G. T. Crossin, R. Filgueira, K. R. Studholme, and J. M. Hipfner. Phenological cues to breeding and the differential response of Pacific auks to variation in marine productivity. Marine Ecology Progress Series, 687:163–172, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/687/
m687p163.pdf.

Clements:2022:OAB

[CG22] J. C. Clements and M. N. George. Ocean acidification and bivalve byssus: explaining variable responses using metaanalysis. Marine Ecology Progress Series, 694:89–103, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/694/m694p089.pdf.

Chan:2024:ASW

[CGCC24] B. K. K. Chan, J. C. Guo, C. L. Chen, and T. Y. Chan. Activity of the shallow-water hydrothermal vent crab Xenograpsus testudinatus is affected by tides, water temperature, and light intensity. Marine Ecology Progress Series, 744:69–82, September 5, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/744/m744p069.pdf.

Chevallay:2024:HTS

[CGGdD24] M. Chevallay, C. Guinet, P. Goulet, and T. Jeanniard du Dot. Hunting tactics of southern elephant seals Mirounga leonina and anti-predatory behaviours of their prey. Marine Ecology Progress Series, 736:167–179, May 23, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/736/
m736p167.pdf.

Cresswell:2023:CRP

[CGH^{+}23] B. J. Cresswell, G. F. Galbraith, H. B. Harrison, M. I. Mc-Cormick, and G. P. Jones. Coral reef pinnacles act as ecological magnets for the abundance, diversity and biomass of predatory fishes. Marine Ecology Progress Series, 717:143– 156, August 24, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/717/m717p143.pdf.

Celentano:2022:URL

[CGL^{+}22] E. Celentano, J. Gómez, D. Lercari, A. de Álava, and O. Defeo. Unravelling the role of local and large-scale factors in structuring sandy beach populations: the wedge clam Donax hanleyanus. Marine Ecology Progress Series, 696:29–42, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m696p029.pdf.

Cardona-Gutierrez:2020:BWS

[CGLC20] M. F. Cardona-Gutiérrez and E. Londoño-Cruz. Boring worms (Sipuncula and Annelida: Polychaeta): their early impact on Eastern Tropical Pacific coral reefs. Marine Ecology Progress Series, 641:101–110, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/641/m641p101.pdf.

Cordone:2023:IIG

[CGM23] G. Cordone, D. E. Galván, and F. R. Momo. Impacts of an invasion by green crab Carcinus maenas on the intertidal food web of a Patagonian rocky shore, Argentina. Marine Ecology Progress Series, 713:97–115, June 29, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/713/
m713p097.pdf.

Calonge:2024:IMS

[CGM^{+}24] A. Calonge, J. Goossens, C. Muñiz, J. Reubens, and E. Debusschere. Importance of multi-sensor observations to advance species co-occurrence knowledge: a demonstration of two acoustic technologies. Marine Ecology Progress Series, 727:49–65, January 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m727p049.pdf.

Choy:2020:VDB

[CGR^{+}20] E. S. Choy, C. Giraldo, B. Rosenberg, J. D. Roth, A. D. Ehrman, A. Majewski, H. Swanson, M. Power, J. D. Reist, and L. L. Loseto. Variation in the diet of beluga whales in response to changes in prey availability: insights on changes in the Beaufort Sea ecosystem. Marine Ecology Progress Series, 647:195–210, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/647/m647p195.pdf.

Correia:2023:TEM

[CGS^{+}23] E. Correia, J. P. Granadeiro, B. Santos, A. Regalla, V. A. Mata, and T. Catry. Trophic ecology of a migratory shorebird community at a globally important non-breeding site: combining DNA metabarcoding and conventional techniques. Marine Ecology Progress Series, 705:127–144, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m705p127.pdf.

Carloni:2021:MEB

[CGW21] J. T. Carloni, J. S. Goldstein, and W. H. Watson III. Movements of egg-bearing American lobsters Homarus americanus during late stage brooding and hatching. Marine Ecology Progress Series, 661:163–173, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m661p163.pdf.

Covernton:2020:MSV

[CH20] G. A. Covernton and C. D. G. Harley. Multi-scale variation in salinity: a driver of population size and structure in the muricid gastropod Nucella lamellosa. Marine Ecology Progress Series, 643:1–19, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Chapman:2022:EST

[Cha22] M. G. Chapman. Evaluation of spatial and temporal variability of multiple measures of diversity in three intertidal assemblages. Marine Ecology Progress Series, 693:19–38, July 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cloyed:2022:HSA

[CHD^{+}22] C. S. Cloyed, E. E. Hieb, K. P. DaCosta, M. Ross, and R. H. Carmichael. Habitat selection and abundance of West Indian manatees Trichechus manatus at the margins of their expanding range. Marine Ecology Progress Series, 696:151– 167, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/696/m696p151.pdf.

Clavel-Henry:2023:PER

[CHDG^{+}23] M. Clavel-Henry, T. Dabrowski, R. J. Giesler, T. P. Crowe, and J. M. Yearsley. Predicting the export of retentionoriented larvae from their natal population using coastal features: a modelling study on the Pacific oyster. Marine Ecology Progress Series, 724:81–95, December 7, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m724p081.pdf.

Cook:2021:GES

[CHJ21] D. Cook, N. Herbert, and A. Jerrett. Growth and energy storage responses vary seasonally in the Australasian snapper Chrysophrys auratus with only modest changes in aerobic scope. Marine Ecology Progress Series, 659:199–217, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/659/m659p199.pdf.

Charrier:2023:ICS

[CIDM23] B. R. Charrier, J. Ingels, S. L. Danielson, and S. L. Mincks. Infaunal community structure, diversity, and function in Pacific– Arctic shelf sediments: a comparison of meiofaunal- and macrofaunal-sized nematodes. Marine Ecology Progress Series, 720:95–116, October 5, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m720p095.pdf.

Clark:2022:IBA

[CILGD^{+}22] B. L. Clark, C. Irigoin-Lovera, D. D. Gonzales-DelCarpio, I. Diaz-Santibañez, S. C. Votier, and C. B. Zavalaga. Interactions between anchovy fisheries and Peruvian boobies revealed by bird-borne cameras and movement loggers. Marine Ecology Progress Series, 701:145–157, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/701/m701p145.pdf.

Costalago:2020:GNC

[CKCS20] D. Costalago, Y. Kisten, C. Clemmesen, and N. A. Strydom. Growth and nutritional condition of anchovy larvae on the west and southeast coasts of South Africa. Marine Ecology Progress Series, 644:119–128, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/644/m644p119.pdf.

Cramer:2021:DRB

[CKKL21] A. N. Cramer, S. Katz, C. Kogan, and J. Lindholm. Distinguishing residency behavior from random movements using passive acoustic telemetry. Marine Ecology Progress Series, 672:73–87, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/672/m672p073.pdf.

Cho:2020:AUM

[CKM^{+}20] H. Cho, B. Kim, J. S. Mok, A. Choi, B. Thamdrup, and J. H. Hyun. Acetate-utilizing microbial communities revealed by stable-isotope probing in sediment underlying the upwelling system of the Ulleung Basin, East Sea. Marine Ecology Progress Series, 634:45–61, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/634/m634p045.pdf.

Contolini:2020:PPD

[CKP20] G. M. Contolini, K. J. Kroeker, and E. P. Palkovacs. Predator populations differ in their foraging responses to acute seawater acidification. Marine Ecology Progress Series, 646: 69–78, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/646/m646p069.pdf.

Cushing:2024:DRS

[CKS^{+}24] D. A. Cushing, K. J. Kuletz, L. Sousa, R. H. Day, S. L. Danielson, E. A. Labunski, and R. R. Hopcroft. Differential response of seabird species to warm- and cold-water events in a heterogeneous cross-shelf environment in the Gulf of Alaska. Marine Ecology Progress Series, 737:31–58, June 6, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m737p031.pdf.

Cannell:2024:MHA

[CKT^{+}24] B. L. Cannell, W. L. Kendall, J. A. Tyne, M. Bunce, Y. Hetzel, D. Murray, and B. Radford. Marine heatwaves affect breeding, diet and population size but not body condition of a rangeedge little penguin colony. Marine Ecology Progress Series, 737:193–213, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2024/737/m737p193.pdf.

Coguiec:2023:POB

[CLC^{+}23] E. Coguiec, K. S. Last, J. H. Cohen, L. Hobbs, M. Choquet, E. Ershova, J. Berge, and M. Daase. Photoperiodism and overwintering in boreal and sub-Arctic Calanus finmarchicus populations. Marine Ecology Progress Series, 712:49– 65, June 8, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/712/m712p049.pdf.

Cannon:2023:IBL

[CLD23] S. E. Cannon, A. Liu, and S. Donner. Interactions between local disturbance and climate-driven heat stress on central Pacific coral reefs. Marine Ecology Progress Series, 709: 17–31, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/709/m709p017.pdf.

Crear:2020:SSN

[CLF^{+}20] D. P. Crear, R. J. Latour, M. A. M. Friedrichs, P. St-Laurent, and K. C. Weng. Sensitivity of a shark nursery habitat to a changing climate. Marine Ecology Progress Series, 652: 123–136, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m652p123.pdf.

Clements:2024:CPC

[CLGH24] C. S. Clements, M. C. Ladd, J. P. Gallagher, and M. E. Hay. Coral performance is comparable when transplanted to disparate reef sites despite divergent histories of reef decline and recovery. Marine Ecology Progress Series, 748:53– 67, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/748/m748p053.pdf.

Cryan:2022:TEA

[CLL^{+}22] D. M. Cryan, N. H. N. Low, S. Y. Litvin, C. W. Osenberg, and F. Micheli. Trophic ecology of an abundant kelp forest echinoderm, the bat star Patiria miniata. Marine Ecology Progress Series, 696:57–68, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/696/m696p057.pdf.

Carmody:2021:SDC

[CLM^{+}21] H. Carmody, T. Langlois, J. Mitchell, M. Navarro, N. Bosch, D. McLean, J. Monk, P. Lewis, and G. Jackson. Shark depredation in a commercial trolling fishery in sub-tropical Australia. Marine Ecology Progress Series, 676:19–35, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/676/m676p019.pdf.

Carrier:2020:EES

[CLR20] T. J. Carrier, H. A. Lessios, and A. M. Reitzel. Eggs of echinoids separated by the Isthmus of Panama harbor divergent microbiota. Marine Ecology Progress Series, 648: 169–177, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/648/m648p169.pdf.

Champagnat:2021:MAN

[CLR^{+}21] J. Champagnat, J. B. Lecomte, E. Rivot, L. Douchet, N. Martin, F. Grasso, F. Mounier, P. Labadie, V. Loizeau, N. Bacq, and O. Le Pape. Multidisciplinary assessment of nearshore nursery habitat restoration for an exploited population of marine fish. Marine Ecology Progress Series, 680:97–109, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/680/m680p097.pdf.

Caillibotte:2020:IOT

[CLV^{+}20] R. Caillibotte, Y. Leredde, F. Vidussi, C. Ulses, P. Marsaleix, C. Estournel, and B. Mostajir. Impact of oysters as top predators on microbial food web dynamics: a modelling approach with parameter optimisation. Marine Ecology Progress Series, 641:79–100, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/641/m641p079.pdf.

Carrier:2022:MAE

[CM22] T. J. Carrier and J. S. McAlister. Microbiota associated with echinoid eggs and the implications for maternal provisioning. Marine Ecology Progress Series, 683:67–79, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/683/
m683p067.pdf.

Cavole:2020:REV

[CMdL^{+}20] L. M. Cavole, J. A. Miller, P. Salinas de León, O. Aburto-Oropeza, J. R. Marin Jarrin, and A. F. Johnson. The role of extrinsic variation — cohabiting juvenile fish species exhibit similar otolith elemental signatures. Marine Ecology Progress Series, 646:109–125, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/646/m646p109.pdf.

Cottrant:2021:BRT

[CMF21] E. Cottrant, P. Matich, and M. R. Fisher. Boosted regression tree models predict the diets of juvenile bull sharks in a subtropical estuary. Marine Ecology Progress Series, 659: 127–141, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/659/m659p127.pdf.

Colombano:2020:ETH

[CMO^{+}20] D. D. Colombano, A. D. Manfree, T. A. O’Rear, J. R. Durand, and P. B. Moyle. Estuarine-terrestrial habitat gradients enhance nursery function for resident and transient fishes in the San Francisco Estuary. Marine Ecology Progress Series, 637:141–157, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m637p141.pdf.

Castro:2021:ESC

[CMT^{+}21] L. Castro, V. Morin, O. Tiznado, A. Miranda, S. Soto, and M. Gonzalez. Effects of salinity changes on hatching, hatchinggene expression, and hatching-enzyme expression in anchoveta Engraulis ringens eggs. Marine Ecology Progress Series, 658: 181–194, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/658/m658p181.pdf.

Correia-Martins:2022:FBF

[CMTB^{+}22] A. Correia-Martins, R. Tremblay, B. Bec, C. Roques, A. Atteia, A. Gobet, M. Richard, M. Hamaguchi, T. Miyajima, M. Hori, G. Miron, S. Pouvreau, and F. Lagarde. Failure of bivalve foundation species recruitment related to trophic changes during an extreme heatwave event. Marine Ecology Progress Series, 691:69–82, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/691/m691p069.pdf.

Chong-Montenegro:2022:QHD

[CMTP22] C. Chong-Montenegro, R. H. Thurstan, and J. M. Pandolfi. Quantifying the historical development of recreational fisheries in Southeast Queensland, Australia. Marine Ecology Progress Series, 696:135–149, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/696/m696p135.pdf.

Croll:2023:GME

[CMvKdR23] J. C. Croll, Q. D. Mudde, T. van Kooten, and A. M. de Roos. Growth as a measure for environmental conditions and relatedness of North Sea fish stocks. Marine Ecology Progress Series, 725:57–73, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m725p057.pdf.

Chen:2020:PGD

[CNC^{+}20] I. Chen, S. Nishida, L. S. Chou, T. Isobe, A. A. Mignucci-Giannoni, and A. R. Hoelzel. Population genetic diversity and historical dynamics of Fraser’s dolphins Lagenodelphis hosei. Marine Ecology Progress Series, 643:183–195, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/643/
m643p183.pdf.

Contreras:2020:FPT

[COGGH20] T. Contreras, M. P. Olivar, J. I. González-Gordillo, and P. A. Hulley. Feeding patterns of transforming and juvenile myctophids that migrate into neustonic layers. Marine Ecology Progress Series, 650:239–252, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/650/m650p239.pdf.

Cleasby:2024:FRM

[COM^{+}24] I. R. Cleasby, E. Owen, P. I. Miller, R. J. Jones, L. J. Wilson, and M. Bolton. Functional responses of a medium-ranging marine predator highlight the importance of frontal zones as foraging locations. Marine Ecology Progress Series, 740: 175–191, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/740/m740p175.pdf.

Chiba:2021:EPA

[COT^{+}21] S. N. Chiba, S. Ohashi, F. Tanaka, A. Suda, A. Fujiwara, D. Snodgrass, H. Kiyofuji, K. Satoh, and N. Suzuki. Effectiveness and potential application of sex-identification DNA markers in tunas. Marine Ecology Progress Series, 659:175– 184, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/659/m659p175.pdf.

Cronin-OReilly:2022:CSB

[COWM^{+}22a] S. Cronin-O’Reilly, N. S. Wells, R. McCallum, C. S. Hallett, J. R. Tweedley, F. J. Valesini, and B. D. Eyre. Chronically stressed benthic macroinvertebrate communities exhibit limited effects on ecosystem function in a microtidal estuary. Marine Ecology Progress Series, 701:1–16, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Cronin-OReilly:2022:DDE

[COWM^{+}22b] S. Cronin-O’Reilly, N. S. Wells, R. McCallum, C. S. Hallett, J. R. Tweedley, F. J. Valesini, and B. D. Eyre. Defaunation by deoxygenation: efficacy and divergent responses of estuarine macroinvertebrates. Marine Ecology Progress Series, 701:17–24, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/701/m701p017.pdf.

Clark:2022:ETD

[CP22a] D. L. Clark and J. A. Pechenik. Effects of the toxic dinoflagellate Karenia brevis on the development of the marine gastropod Crepidula fornicata. Marine Ecology Progress Series, 699:45–64, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/699/m699p045.pdf.

Clarkson:2022:CEU

[CP22b] E. Clarkson and J. Beseres Pollack. Cost-effective use of aerial imagery to quantify faunal-habitat associations across multiple spatial scales. Marine Ecology Progress Series, 684:

37–56, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/684/m684p037.pdf.

Chero:2020:RCE

[CPD^{+}20] G. Chero, R. Pradel, S. Derville, C. Bonneville, O. Gimenez, and C. Garrigue. Reproductive capacity of an endangered and recovering population of humpback whales in the southern hemisphere. Marine Ecology Progress Series, 643:219– 227, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m643p219.pdf.

Chase:2020:BTO

[CPH20] T. J. Chase, M. S. Pratchett, and M. O. Hoogenboom. Behavioral trade-offs and habitat associations of coral-dwelling damselfishes (family Pomacentridae). Marine Ecology Progress Series, 633:141–156, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/633/m633p141.pdf.

Catalan:2020:TSR

[CRÁ^{+}20] I. A. Catalán, P. Reglero, I. Álvarez, R. G. Asch, I. A. Catalán, J. Davis, S. Garrido, F. Mueter, P. Olivar, and M. Peck. Theme section: Research on early life stages of fish: a lively field. Marine Ecology Progress Series, 650:1–5, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Campos:2021:BCE

[CRB^{+}21] J. Campos, F. Ribas, A. Bio, V. Freitas, A. T. Souza, and H. W. van der Veer. Body condition and energy content of shore crab Carcinus maenas in a temperate coastal system: temporal variability. Marine Ecology Progress Series, 667: 99–112, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/667/m667p099.pdf.

Cardona:2020:FAR

[CRG^{+}20] L. Cardona, O. Reñones, A. Gouragine, F. Saporiti, A. Aguilar, and J. Moranta. Fishing alters resource partitioning between colour morphs in a temperate coastal fish. Marine Ecology Progress Series, 648:179–190, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/648/
m648p179.pdf.

Christiansen:2021:PBC

[CRGMA^{+}21] F. Christiansen, F. Rodríguez-González, S. Martínez-Aguilar, J. Urbán, S. Swartz, H. Warick, F. Vivier, and L. Bejder. Poor body condition associated with an unusual mortality event in gray whales. Marine Ecology Progress Series, 658: 237–252, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/658/m658p237.pdf.

Cespedes-Rodriguez:2021:GCC

[CRLC21] E. C. Céspedes-Rodríguez and E. Londoño-Cruz. Gross calcium carbonate production in Eastern Tropical Pacific coral reefs (Gorgona Island, Colombia). Marine Ecology Progress Series, 665:37–46, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/665/m665p037.pdf.

Clayton:2020:EPN

[CS20] K. A. Clayton and J. I. Spicer. Evidence for physiological niche expansion of an intertidal flatworm: evolutionary rescue in the wild. Marine Ecology Progress Series, 651: 85–95, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/651/m651p085.pdf.

Corradino:2022:GHN

[CS22] G. L. Corradino and A. Schnetzer. Grazing of a heterotrophic nanoflagellate on prokaryote and eukaryote prey: ingestion rates and gross growth efficiency. Marine Ecology Progress Series, 682:65–77, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/682/m682p065.pdf.

Clinton:2024:MDP

[CSB24] M. E. Clinton, P. V. R. Snelgrove, and A. E. Bates. Macrofaunal diversity patterns in coastal marine sediments: reexamining common metrics and methods. Marine Ecology Progress Series, 735:1–26, May 2, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Castro-Sanguino:2020:DCS

[CSBM20] C. Castro-Sanguino, Y. M. Bozec, and P. J. Mumby. Dynamics of carbonate sediment production by Halimeda: implications for reef carbonate budgets. Marine Ecology Progress Series, 639:91–106, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/639/m639p091.pdf.

Czapanskiy:2024:SIM

[CSD^{+}24] M. F. Czapanskiy, J. A. Santora, K. S. Dietrich, M. A. Cimino, E. L. Hazen, C. S. Reiss, and R. R. Veit. Sea-ice and macrozooplankton distribution as determinants of top predator community structure in Antarctic winter. Marine Ecology Progress Series, 738:57–73, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m738p057.pdf.

Clements:2022:FGB

[CSDB22] M. Clements, P. Selvakumaraswamy, D. Deaker, and M. Byrne. Freshening of Great Barrier Reef waters is deleterious for larval crown-of-thorns starfish, counter to the terrestrial runoff hypothesis. Marine Ecology Progress Series, 696:1–14, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Chen:2021:EPS

[CSH^{+}21a] M. Chen, Y. Si, L. Han, X. Liu, B. Huang, and C. K. Kang. Effect of prey selectivity and trophic cascades induced by mesozooplankton on the dynamics of phytoplankton. Marine Ecology Progress Series, 662:35–51, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/662/
m662p035.pdf.

Clark:2021:ILD

[CSH^{+}21b] D. E. Clark, F. Stephenson, J. E. Hewitt, J. I. Ellis, A. Zaiko, A. Berthelsen, R. H. Bulmer, and C. A. Pilditch. Influence of land-derived stressors and environmental variability on compositional turnover and diversity of estuarine benthic communities. Marine Ecology Progress Series, 666:1–18, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ciannelli:2022:GEN

[CSK^{+}22] L. Ciannelli, E. Smith, K. Kearney, M. Hunsicker, and C. McGilliard. Greater exposure of nearshore habitats in the Bering Sea makes fish early life stages vulnerable to climate change. Marine Ecology Progress Series, 684:91–102, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m684p091.pdf.

Chen:2023:ASN

[CSK^{+}23] E. K. Chen, W. H. Satterthwaite, B. J. Kormos, R. C. Johnson, C. C. Phillis, and S. M. Carlson. Age structure of natural versus hatchery-origin endangered Chinook salmon and implications for fisheries management in California. Marine Ecology Progress Series, 723:37–55, November 23, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/723/
m723p037.pdf.

Cormack:2023:EDS

[CSN^{+}23] A. Cormack, E. Slavich, J. Negrete, H. Bornemann, G. A. Daneri, and T. L. Rogers. Extreme dietary specialisation in adult male southern elephant seals: determining variation between individual trophic diets. Marine Ecology Progress Series, 725:185–199, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m725p185.pdf.

Cunnington:2024:EIS

[CSWM24] A. V. Cunnington, P. Shum, C. S. Wilding, and S. Mariani. Exploring intertidal sea anemones Actinia equina as natural eDNA samplers for coastal biodiversity assessment. Marine Ecology Progress Series, 743:159–165, August 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2024/743/m743p159.pdf.

Chasco:2020:IME

[CTH^{+}20] B. E. Chasco, J. T. Thorson, S. S. Heppell, L. Avens, J. Braun McNeill, A. B. Bolten, K. A. Bjorndal, and E. J. Ward. Integrated mixed-effect growth models for species with incomplete ageing histories: a case study for the loggerhead sea turtle Caretta caretta. Marine Ecology Progress Series, 636:221–234, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/636/m636p221.pdf.

Cabral-Tena:2024:SAD

[CTTLCG^{+}24] R. A. REVIEW Cabral-Tena, J. J. A. Tortolero-Langarica, J. P. Carricart-Ganivet, A. P. Rodríguez-Troncoso, I. Cruz-Ortega, A. L. Cupul-Magaña, E. F. Balart, H. Reyes-Bonilla, and A. López-Pérez. Sex-associated differences in sclerochronology and sensitivity to thermal stress in Caribbean and eastern Pacific reef-building corals. Marine Ecology Progress Series, 743:167–183, August 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/743/
m743p167.pdf.

Campioni:2023:CBL

[CVG^{+}23] L. Campioni, F. Ventura, J. P. Granadeiro, J. Madeiros, C. Gjerdrum, and M. C. Silva. Combining bio-logging, stable isotopes and DNA metabarcoding to reveal the foraging ecology and diet of the endangered Bermuda petrel Pterodroma cahow. Marine Ecology Progress Series, 723:151–170, November 23, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/723/m723p151.pdf.

Calado:2021:IFS

[CVP^{+}21] J. G. Calado, S. N. Veríssimo, V. H. Paiva, R. Ramos, P. T. Vaz, D. Matos, J. Pereira, C. Lopes, N. Oliveira, A. Quaresma, F. R. Ceia, A. Velando, and J. A. Ramos. Influence of fisheries on the spatio-temporal feeding ecology of gulls along the western Iberian coast. Marine Ecology Progress Series, 661:187–201, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/661/m661p187.pdf.

Cleasby:2022:ABR

[CWC^{+}22] I. R. Cleasby, L. J. Wilson, R. Crawford, E. Owen, Y. Rouxel, and M. Bolton. Assessing bycatch risk from gillnet fisheries for three species of diving seabird in the UK. Marine Ecology Progress Series, 684:157–179, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/684/m684p157.pdf.

Clarke:2022:EST

[CWD^{+}22] T. M. Clarke, S. K. Whitmarsh, R. G. Dwyer, V. Udyawer, H. Pederson, and C. Huveneers. Effects of shark tourism on the daily residency and movements of a non-focal pelagic teleost. Marine Ecology Progress Series, 687:133–146, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/687/m687p133.pdf.

Currie:2024:OMI

[CWD24] T. H. Currie, S. R. Wing, and L. M. Durante. Otolith microchemistry identifies shallow, intertidal-dominated estuaries as important nurseries for sand flounder in New Zealand. Marine Ecology Progress Series, 742:113–130, August 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m742p113.pdf.

Cones:2022:SBD

[CZS^{+}22] S. F. Cones, D. Zhang, K. A. Shorter, K. Katija, D. A. Mann, F. H. Jensen, J. Fontes, P. Afonso, and T. A. Mooney. Swimming behaviors during diel vertical migration in veined squid Loligo forbesii. Marine Ecology Progress Series, 691: 83–96, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/691/m691p083.pdf.

Dror:2024:RST

[DA24] H. Dror and D. Angel. Rising seawater temperatures affect the fitness of Rhopilema nomadica polyps and podocysts and the expansion of this medusa into the western Mediterranean. Marine Ecology Progress Series, 728:123–143, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m728p123.pdf.

Dureuil:2021:UNM

[DAB^{+}21] M. Dureuil, W. H. Aeberhard, K. A. Burnett, R. E. Hueter, J. P. Tyminski, and B. Worm. Unified natural mortality estimation for teleosts and elasmobranchs. Marine Ecology Progress Series, 667:113–129, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m667p113.pdf.

Morales-de-Anda:2020:RFF

[dACMRZ^{+}20] D. Morales de Anda, A. L. Cupul-Magaña, F. A. Rodríguez-Zaragoza, C. Aguilar-Betancourt, G. González-Sansón, and A. P. Rodríguez-Troncoso. Reef fish functional composition and metrics reveal spatial differences in three protected islands in the Eastern Pacific. Marine Ecology Progress Series, 635:139–150, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/635/m635p139.pdf.

Dupuis:2021:LLG

[DAT^{+}21] B. Dupuis, F. Amélineau, A. Tarroux, O. Bjørnstad, V. S. Bråthen, J. Danielsen, S. Descamps, P. Fauchald, G. T. Hallgrimsson, E. S. Hansen, M. Helberg, H. H. Helgason, J. E. Jónsson, Y. Kolbeinsson, E. Lorentzen, P. Thompson, T. L. Thórarinsson, and H. Strøm. Light-level geolocators reveal spatial variations in interactions between northern fulmars and fisheries. Marine Ecology Progress Series, 676:159–172, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m676p159.pdf.

Diehl:2021:CCA

[DB21] N. Diehl and K. Bischof. Coping with a changing Arctic: mechanisms of acclimation in the brown seaweed Saccharina latissima from Spitsbergen. Marine Ecology Progress Series, 657:43–57, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/657/m657p043.pdf.

deBettignies:2020:TSM

[dBDL^{+}20] F. de Bettignies, P. Dauby, G. Lepoint, P. Riera, E. Bocher, O. Bohner, C. Broudin, C. Houbin, C. Leroux, S. Loisel, and D. Davoult. Temporal succession of a macrofaunal community associated with kelp fragment accumulations in an in situ experiment. Marine Ecology Progress Series, 656:109– 121, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/656/m656p109.pdf.

Diaz-Barroso:2022:SSM

[DBHCO^{+}22] L. Díaz-Barroso, I. Hernández-Carrasco, A. Orfila, P. Reglero, R. Balbín, M. Hidalgo, J. Tintoré, F. Alemany, and D. Alvarez-Berastegui. Singularities of surface mixing activity in the Western Mediterranean influence bluefin tuna larval habitats. Marine Ecology Progress Series, 685:69–84, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m685p069.pdf.

Dunn:2023:EIS

[DBK^{+}23] R. P. Dunn, T. L. Buck, J. L. Krask, J. B. Stevens, and E. M. Smith. Elevation influences salt marsh crab abundance, diversity, and burrowing. Marine Ecology Progress Series, 704:55–66, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/704/m704p055.pdf.

Dillon:2022:DDS

[DBP^{+}22] E. M. Dillon, A. Bagla, K. D. Plioplys, D. J. McCauley, K. D. Lafferty, and A. O’Dea. Dermal denticle shedding rates vary between two captive shark species. Marine Ecology Progress Series, 682:153–167, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2022/682/m682p153.pdf.

Delorme:2020:SSR

[DBS^{+}20] N. J. Delorme, L. Biessy, P. M. South, L. N. Zamora, N. L. C. Ragg, and D. J. Burritt. Stress-on-stress responses of a marine mussel, Perna canaliculus: food limitation reduces the ability to cope with heat stress in juveniles. Marine Ecology Progress Series, 644:105–117, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/644/m644p105.pdf.

Davis:2020:VPC

[DCG^{+}20] M. J. Davis, J. W. Chamberlin, J. R. Gardner, K. A. Connelly, M. M. Gamble, B. R. Beckman, and D. A. Beauchamp. Variable prey consumption leads to distinct regional differences in Chinook salmon growth during the early marine critical period. Marine Ecology Progress Series, 640:147– 169, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/640/m640p147.pdf.

Delord:2024:ENS

[DCK^{+}24] K. Delord, Y. Cherel, A. Kato, P. Bustamante, and C. Barbraud. Ecological niche segregation between sympatrically breeding blue petrels and thin-billed prions. Marine Ecology Progress Series, 749:159–180, November 21, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/749/
m749p159.pdf.

Durante:2021:INP

[DCL21] C. A. Durante, E. A. Crespo, and R. Loizaga. Isotopic niche partitioning between two small cetacean species. Marine Ecology Progress Series, 659:247–259, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/659/
m659p247.pdf.

Davey:2023:SVS

[DCN^{+}23] J. Davey, T. M. Clarke, Y. Niella, J. D. Dennis, and C. Huveneers. Seasonal variation in space use and residency of the southern eagle ray Myliobatis tenuicaudatus in a temperate ecosystem. Marine Ecology Progress Series, 705:77– 94, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/705/m705p077.pdf.

Doherty:2021:SDM

[DCR^{+}21] B. Doherty, S. P. Cox, C. N. Rooper, S. D. N. Johnson, and A. R. Kronlund. Species distribution models for deep-water coral habitats that account for spatial uncertainty in trapcamera fishery data. Marine Ecology Progress Series, 660: 69–93, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/660/m660p069.pdf.

Delord:2022:SBE

[DCR^{+}22] K. Delord, Y. Cherel, A. Roy, P. Bustamante, K. M. Swadling, H. Weimerskirch, C. A. Bost, and C. Barbraud. At-sea behavioural ecology of the endangered MacGillivray’s prion from Saint Paul Island: combining tracking and stable isotopes. Marine Ecology Progress Series, 697:149–165, September 22, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/697/m697p149.pdf.

Dumas:2024:VIC

[DDB^{+}24] P. Dumas, A. Durbano, B. Bourgeois, H. Gossuin, and C. Peignon. Vinegar injections can be used safely to control outbreaks of crown-of-thorns starfish Acanthaster cf. solaris during the peak spawning season. Marine Ecology Progress Series, 742:87–95, August 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/742/m742p087.pdf.

Dupaix:2023:NEL

[DDD^{+}23] A. Dupaix, L. Dagorn, A. Duparc, A. Guillou, J. L. Deneubourg, and M. Capello. No evidence from long-term analysis of yellowfin tuna condition that drifting fish aggregating devices act as ecological traps. Marine Ecology Progress Series, 711:121–127, May 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/711/m711p121.pdf.

Dupont:2021:ESI

[DDG^{+}21] N. Dupont, J. M. Durant, H. Gjøsæter, Ø. Langangen, and L. C. Stige. Effects of sea ice cover, temperature and predation on the stock dynamics of the key Arctic fish species polar cod Boreogadus saida. Marine Ecology Progress Series, 677:141–159, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/677/m677p141.pdf.

Darby:2021:FDB

[DdGA^{+}21] J. H. Darby, S. de Grissac, G. E. Arneill, E. Pirotta, J. J. Waggitt, L. Börger, E. Shepard, D. Cabot, E. Owen, M. Bolton, E. W. J. Edwards, P. M. Thompson, J. L. Quinn, and M. Jessopp. Foraging distribution of breeding northern fulmars is predicted by commercial fisheries. Marine Ecology Progress Series, 679:181–194, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m679p181.pdf.

DiRoberts:2021:TAN

[DDR^{+}21] L. E. DiRoberts, A. Dudek, N. E. Ray, R. W. Fulweiler, and R. D. Rotjan. Testing assumptions of nitrogen cycling between a temperate, model coral host and its facultative symbiont: symbiotic contributions to dissolved inorganic nitrogen assimilation. Marine Ecology Progress Series, 670:61–74, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m670p061.pdf.

dEntremont:2022:IIA

[dDW^{+}22] K. J. N. d’Entremont, G. K. Davoren, C. J. Walsh, S. I. Wilhelm, and W. A. Montevecchi. Intra- and inter-annual shifts in foraging tactics by parental northern gannets Morus bassanus indicate changing prey fields. Marine Ecology Progress Series, 698:155–170, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/698/m698p155.pdf.

DeCarlo:2024:WFE

[DeC24] T. M. DeCarlo. Are we forgetting early observations of coral bleaching? Marine Ecology Progress Series, 748:191–195, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m748p191.pdf.

Detmer:2024:MGM

[Det24] A. R. Detmer. Modeling grazer-mediated effects of demographic and material connectivity on giant kelp metapopulation dynamics. Marine Ecology Progress Series, 726:49– 69, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/726/m726p049.pdf.

Dennis:2024:IDP

[DFB^{+}24] L. E. Dennis, F. Favoretto, E. F. Balart, A. Munguía-Vega, C. Sánchez-Ortiz, and D. A. Paz-García. Isolation by disturbance: a pattern of genetic structure of the coral Pocillopora grandis in the Gulf of California. Marine Ecology Progress Series, 733:43–57, April 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/733/m733p043.pdf.

Dolan:2021:EFM

[DFCM21] T. E. Dolan, M. G. Frisk, R. M. Cerrato, and A. E. McElroy. Environmental factors modify post-settlement survival and growth of winter flounder Pseudopleuronectes americanus in the absence of predation. Marine Ecology Progress Series, 676:57–75, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/676/m676p057.pdf.

Davidson:2023:OSF

[DFHT23] E. R. Davidson, S. H. Ferguson, J. W. Higdon, and M. A. Treble. Opportunistic sightings from fisheries surveys inform habitat suitability for northern bottlenose whales Hyperoodon ampullatus and sperm whales Physeter macrocephalus in Baffin Bay and Davis Strait, Canadian Arctic. Marine Ecology Progress Series, 723:57–71, November 23, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m723p057.pdf.

Daly:2021:OSH

[DFP^{+}21a] R. Daly, J. D. Filmalter, L. R. Peel, B. Q. Mann, J. S. E. Lea, C. R. Clarke, and P. D. Cowley. Ontogenetic shifts in home range size of a top predatory reef-associated fish (Caranx ignobilis): implications for conservation. Marine Ecology Progress Series, 664:165–182, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m664p165.pdf.

Durban:2021:SBC

[DFP^{+}21b] J. W. Durban, H. Fearnbach, A. Paredes, L. S. Hickmott, and D. J. LeRoi. Size and body condition of sympatric killer whale ecotypes around the Antarctic Peninsula. Marine Ecology Progress Series, 677:209–217, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m677p209.pdf.

Dewar-Fowler:2024:BIR

[DFRS^{+}24] V. Dewar-Fowler, C. Robinson, R. A. Saunders, S. E. Thorpe, E. P. Abrahamsen, and G. A. Tarling. Biological indicators reveal mesozooplankton foray behaviour in dynamic physical environments: Reply to Kaartvedt et al. (2024). Marine Ecology Progress Series, 734:177–180, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/734/
m734p177.pdf. See [KCT24].

deFouw:2022:ITR

[dFRvdG^{+}22] J. de Fouw, K. Rehlmeyer, M. van der Geest, A. J. P. Smolders, and T. van der Heide. Increased temperature reduces the positive effect of sulfide-detoxification mutualism on Zostera noltii nutrient uptake and growth. Marine Ecology Progress Series, 692:43–52, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m692p043.pdf.

Duncan:2023:TEM

[DFSH23] S. E. Duncan, H. O. Fock, A. F. Sell, and W. Hagen. Trophic ecology of mesopelagic fishes in the northern and southern Benguela Upwelling Systems revealed through stable isotope patterns. Marine Ecology Progress Series, 725:75–93, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m725p075.pdf.

DiazdeAstarloa:2023:EII

[DGM^{+}23a] C. Diaz de Astarloa, M. Giorgini, A. Miguez, O. Iribarne, and M. Escapa. Effect of increased inundation time on intertidal ecosystem functions of South West Atlantic soft-bottom environments. Marine Ecology Progress Series, 711:1–16, May 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Diggins:2023:CFS

[DGM^{+}23b] R. L. Diggins, J. Grimm, D. Mendez, K. Jones, M. Hamann, I. Bell, and E. Ariel. Confirmed feasibility of a satellite tracker attachment method on small juvenile hawksbill turtles Eretmochelys imbricata. Marine Ecology Progress Series, 704:119–130, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m704p119.pdf.

DeLeij:2022:FTL

[DGP22] R. De Leij, L. J. Grange, and L. S. Peck. Functional thermal limits are determined by rate of warming during simulated marine heatwaves. Marine Ecology Progress Series, 685:183–196, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/685/m685p183.pdf.

Dorner:2020:OAI

[DHA^{+}20] I. Dörner, H. Hauss, N. Aberle, K. Lohbeck, C. Spisla, U. Riebesell, and S. M. H. Ismar-Rebitz. Ocean acidification impacts on biomass and fatty acid composition of a postbloom marine plankton community. Marine Ecology Progress Series, 647:49–64, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/647/m647p049.pdf.

Davies:2023:SAP

[DHE^{+}23] J. G. Davies, E. M. Humphreys, T. Evans, R. J. Howells, R. O’Hara-Murray, and J. W. Pearce-Higgins. Seabird abundances projected to decline in response to climate change in Britain and Ireland. Marine Ecology Progress Series, 725:121– 140, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/725/m725p121.pdf.

Duncan:2024:MFA

[DHF24] S. E. Duncan, W. Hagen, and H. O. Fock. Mesopelagic fish assemblages in the Mauritanian Upwelling System off Northwest Africa with oxygen as a major driving force. Marine Ecology Progress Series, 733:95–110, April 4, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m733p095.pdf.

Descoteaux:2022:OMI

[DHJ^{+}22] R. Descôteaux, M. Huserbråten, L. L. Jørgensen, P. E. Renaud, R. B. Ingvaldsen, E. A. Ershova, and B. A. Bluhm. Origin of marine invertebrate larvae on an Arctic inflow shelf. Marine Ecology Progress Series, 699:1–17, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Dale:2021:CRH

[DHM21] K. E. Dale, R. Hallisey, and R. S. Mehta. Coloration is related to habitat in Gymnothorax mordax, a kelp forest predator. Marine Ecology Progress Series, 677:67–79, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/677/
m677p067.pdf.

Daban:2024:MED

[DHM^{+}24] P. Daban, A. Hillinger, G. Mucientes, A. Blanco, and A. Alonso-Fernández. Movement ecology determines isotopic niche width in the undulate skate Raja undulata. Marine Ecology Progress Series, 731:147–158, March 13, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m731p147.pdf.

DeWysiecki:2022:PSH

[DIC^{+}22] A. M. De Wysiecki, A. J. Irigoyen, F. Cortés, N. D. Bovcon, A. C. Milessi, N. M. Hozbor, M. Coller, and A. J. Jaureguizar. Population-scale habitat use by school sharks Galeorhinus galeus (Triakidae) in the Southwest Atlantic: insights from temporally explicit niche modelling and habitat associations. Marine Ecology Progress Series, 697:81–95, September 22, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/697/m697p081.pdf. See erratum

[Ano22e].

DAgostino:2020:BTF

[DJR^{+}20] D. D’Agostino, C. Jimenez, T. Reader, L. Hadjioannou, S. Heyworth, M. Aplikioti, M. Argyrou, and D. A. Feary. Behavioural traits and feeding ecology of Mediterranean lionfish and naiveté of native species to lionfish predation. Marine Ecology Progress Series, 638:123–135, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/638/
m638p123.pdf.

Day:2023:EHP

[DKAB23] J. Day, N. A. Knott, D. Ayre, and M. Byrne. Effects of habitat on predation of ecologically important sea urchin species on east coast Australian temperate reefs in tethering experiments. Marine Ecology Progress Series, 714:71– 86, July 13, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/714/m714p071.pdf.

Day:2024:IDC

[DKS^{+}24] J. K. Day, N. A. Knott, D. S. Swadling, D. J. Ayre, M. J. Huggett, and T. F. Gaston. Investigating the diets and condition of Centrostephanus rodgersii (long-spined urchin) in barrens and macroalgae habitats in south-eastern Australia. Marine Ecology Progress Series, 729:167–183, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/729/m729p167.pdf.

Dong:2020:AAT

[DKSS20] K. Dong, Ø. K. Kvile, N. C. Stenseth, and L. C. Stige. Associations among temperature, sea ice and phytoplankton bloom dynamics in the Barents Sea. Marine Ecology Progress Series, 635:25–36, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/635/m635p025.pdf.

Dong:2021:ABT

[DKSS21] K. Dong, K. Ø. Kvile, N. C. Stenseth, and L. C. Stige. Associations between timing and magnitude of spring blooms and zooplankton dynamics in the southwestern Barents Sea. Marine Ecology Progress Series, 668:57–72, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/668/
m668p057.pdf.

Decima:2020:RPT

[DL20] M. Décima and M. R. Landry. Resilience of plankton trophic structure to an eddy-stimulated diatom bloom in the North Pacific Subtropical Gyre. Marine Ecology Progress Series, 643:33–48, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/643/m643p033.pdf.

DiMartino:2021:LOA

[DL21] E. Di Martino and L. H. Liow. Larger offspring associated with lower temperatures across species of Microporella, a widespread colonial invertebrate. Marine Ecology Progress Series, 662:1–13, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

delaBarra:2022:TWE

[dlBSL^{+}22] P. de la Barra, M. W. Skov, P. J. Lawrence, J. I. Schiaffi, and J. G. Hiddink. Tidal water exchange drives fish and crustacean abundances in salt marshes. Marine Ecology Progress Series, 694:61–72, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/694/m694p061.pdf.

delaBarra:2020:PDP

[dlBSR^{+}20] P. de la Barra, G. Svendsen, M. A. Romero, M. S. Avaca, and M. Narvarte. Predicting the distribution of a portunid crab in Patagonian coastal waters. Marine Ecology Progress Series, 638:95–105, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/638/m638p095.pdf.

deLima:2022:ESI

[dLCN^{+}22] R. C. de Lima, J. D. Cebuhar, J. Negrete, A. Ferreira, E. R. Secchi, and S. Botta. Ecosystem shifts inferred from longterm stable isotope analysis of male Antarctic fur seal Arctocephalus gazella teeth. Marine Ecology Progress Series, 695:203–216, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/695/m695p203.pdf.

Duermit-Moreau:2022:CBA

[DMBB22] E. Duermit-Moreau, J. Bojko, and D. C. Behringer. Cyanobacterial blooms alter benthic community structure and parasite prevalence among invertebrates in Florida Bay, USA. Marine Ecology Progress Series, 694:29–44, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/694/
m694p029.pdf.

Doll:2021:HSO

[DMBJ21] P. C. Doll, P. L. Munday, M. C. Bonin, and G. P. Jones. Habitat specialisation and overlap in coral reef gobies of the genus Eviota (Teleostei: Gobiidae). Marine Ecology Progress Series, 677:81–94, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/677/m677p081.pdf.

Dias:2022:IEP

[DMCB22] B. S. Dias, D. W. McGowan, R. Campbell, and T. A. Branch. Influence of environmental and population factors on Prince William Sound herring spawning phenology. Marine Ecology Progress Series, 696:103–117, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m696p103.pdf.

Durban:2023:QSK

[DME^{+}23] J. W. Durban, C. O. Matkin, D. K. Ellifrit, R. D. Andrews, and L. G. Barrett-Lennard. Quantifying a stopover of killer whales preying on gray whales rounding the Alaska Peninsula. Marine Ecology Progress Series, 724:1–15, December 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

deMontaudouin:2021:JGD

[dMGGG21] X. de Montaudouin, S. Grimault, M. Grandpierre, and A. Garenne. Juvenile growth deficit as an early alert of cockle Cerastoderma edule mortality. Marine Ecology Progress Series, 679:85–99, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m679p085.pdf.

Deaker:2021:CDP

[DML^{+}21] D. J. Deaker, B. Mos, C. Lawson, S. A. Dworjanyn, C. Budden, and M. Byrne. Coral defences: the perilous transition of juvenile crown-of-thorns starfish to corallivory. Marine Ecology Progress Series, 665:115–125, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/665/
m665p115.pdf.

Dixon:2024:LMS

[DMM^{+}24a] R. B. Dixon, T. S. Murray, B. Q. Mann, P. D. Cowley, R. Daly, and J. D. Filmalter. Longshore movements and site fidelity of the iconic giant trevally Caranx ignobilis from southern Africa, determined using passive acoustic telemetry. Marine Ecology Progress Series, 729:201–218, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/729/m729p201.pdf.

Duncan:2024:TDH

[DMM^{+}24b] M. I. Duncan, F. Micheli, J. A. Marquez, C. J. Lowe, S. L. Hamilton, and E. A. Sperling. Temperature-dependent hypoxia tolerance of purple sea urchin Strongylocentrotus purpuratus across biogeography and ontogeny. Marine Ecology Progress Series, 739:129–146, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/739/
m739p129.pdf.

Descamps:2021:SWG

[DMS^{+}21] S. Descamps, B. Merkel, H. Strøm, R. Choquet, H. Steen, J. Fort, M. Gavrilo, D. Grémillet, D. Jakubas, K. Jerstad, N. J. Karnovsky, Y. V. Krasnov, B. Moe, J. Welcker, and K. Wojczulanis-Jakubas. Sharing wintering grounds does not synchronize annual survival in a high Arctic seabird, the little auk. Marine Ecology Progress Series, 676:233– 242, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/676/m676p233.pdf.

Dorval:2022:OCI

[DMTP22] E. Dorval, R. D. Methot, I. G. Taylor, and K. R. Piner. Otolith chemistry indicates age and region of settlement of immature shortspine thornyhead Sebastolobus alascanus in the eastern Pacific Ocean. Marine Ecology Progress Series, 693:157–175, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/693/m693p157.pdf.

Davis:2024:ELW

[DNE^{+}24] K. E. Davis, F. Noisette, J. K. Ehn, Z. Z. A. Kuzyk, C. J. Peck, and M. I. O’Connor. Effects of light and water column nutrient availability on eelgrass Zostera marina productivity in Eeyou Istchee, eastern James Bay, Quebec. Marine Ecology Progress Series, 738:103–117, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/738/m738p103.pdf.

Downie:2021:PSP

[DNJCH21] A. L. Downie, T. Noble-James, A. Chaverra, and K. L. Howell. Predicting sea pen (Pennatulacea) distribution on the UK continental shelf: evidence of range modification by benthic trawling. Marine Ecology Progress Series, 670: 75–91, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/670/m670p075.pdf.

Fernandes:2022:ACS

[dOFBSdJR^{+}22] F de O. Fernandes, M. S. Barbosa-Silva, J. F. de J. Resende, M. A. do A. Carneiro, G. O. Longo, and E. Marinho-Soriano. Amphipod choice for seaweeds under predator cues: interactions on decision-making behavior. Marine Ecology Progress Series, 698:85–94, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/698/m698p085.pdf.

Dadon-Pilosof:2023:DCR

[DPCL^{+}23] A. Dadon-Pilosof, K. Conley, F. Lombard, K. R. Sutherland, A. Genin, M. Richter, F. O. Glöckner, and G. Yahel. Differential clearance rates of microbial phylotypes by four appendicularian species. Marine Ecology Progress Series, 706:73–89, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/706/m706p073.pdf.

Dalpaz:2021:BTA

[DPDJ^{+}21] L. Dalpaz, A. D. Paro, F. G. Daura-Jorge, M. Rossi-Santos, T. F. Norris, S. N. Ingram, and L. L. Wedekin. Better together: analysis of integrated acoustic and visual methods when surveying a cetacean community. Marine Ecology Progress Series, 678:197–209, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/678/m678p197.pdf.

Djeghri:2021:HTP

[DPL^{+}21] N. Djeghri, P. Pondaven, F. Le Grand, A. Bideau, N. Duquesne, M. Stockenreiter, S. Behl, J. Y. T. Huang, T. Hansen, S. Patris, G. Ucharm, and H. Stibor. High trophic plasticity in the mixotrophic Mastigias papua — Symbiodiniaceae holobiont: implications for the ecology of zooxanthellate jellyfishes. Marine Ecology Progress Series, 666:73–88, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/666/m666p073.pdf.

DePascalis:2021:SED

[DPP^{+}21] F. De Pascalis, D. Pala, D. Pisu, J. Morinay, A. Benvenuti, C. Spano, A. Ruiu, L. Serra, D. Rubolini, and J. G. Cecere. Searching on the edge: dynamic oceanographic features increase foraging opportunities in a small pelagic seabird. Marine Ecology Progress Series, 668:121–132, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/668/
m668p121.pdf.

Dodino:2020:IAV

[DRP^{+}20] S. Dodino, L. Riccialdelli, M. J. Polito, K. Pütz, and A. Raya Rey. Inter-annual variation in the trophic niche of Magellanic penguins Spheniscus magellanicus during the pre-molt period in the Beagle Channel. Marine Ecology Progress Series, 655:215–225, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/655/m655p215.pdf.

Doherty:2024:CET

[DSA^{+}24] E. A. Doherty, G. Stenson, M. R. Anderson, C. Mahaffey, and R. M. Jeffreys. Critical evaluation of trophic discrimination factors using compound-specific stable isotopes in a Northwest Atlantic shelf marine food web. Marine Ecology Progress Series, 747:19–34, October 17, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/747/m747p019.pdf.

Demidov:2021:SFS

[DSB^{+}21] A. B. Demidov, I. N. Sukhanova, T. A. Belevich, M. V. Flint, V. I. Gagarin, V. M. Sergeeva, E. V. Eremeeva, and A. V. Fedorov. Size-fractionated surface phytoplankton in the Kara and Laptev Seas: environmental control and spatial variability. Marine Ecology Progress Series, 664:59–77, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/664/
m664p059.pdf.

dosSantos:2023:SBP

[dSGC^{+}23] I. dos Santos, A. M. M. Gonçalves, A. R. Carreiro, B. Martins, C. P. Rocha, C. Vieira, D. M. Matos, I. B. Gutiérrez, I. Rodrigues, N. Almeida, J. A. Ramos, V. H. Paiva, and P. M. Araújo. Similar breeding performance despite inter-annual differences in diet composition of seabirds inhabiting a tropical environment. Marine Ecology Progress Series, 725:95– 119, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/725/m725p095.pdf.

Duncan:2022:EDU

[DSHF22] S. E. Duncan, A. F. Sell, W. Hagen, and H. O. Fock. Environmental drivers of upper mesopelagic fish assemblages in the Benguela Upwelling System. Marine Ecology Progress Series, 688:133–152, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m688p133.pdf.

DeRoy:2020:HPE

[DSHM20] E. M. DeRoy, R. Scott, N. E. Hussey, and H. J. MacIsaac. High predatory efficiency and abundance drive expected ecological impacts of a marine invasive fish. Marine Ecology Progress Series, 637:195–208, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/637/m637p195.pdf.

Depot:2020:APD

[DSK^{+}20] K. M. Depot, L. C. Scopel, S. W. Kress, P. Shannon, A. W. Diamond, and K. H. Elliott. Atlantic puffin diet reflects haddock and redfish abundance in the Gulf of Maine. Marine Ecology Progress Series, 656:75–87, 2020. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m656p075.pdf. See

corrigendum [Ano21b].

Dillon:2022:ERS

[DSK^{+}22] K. S. Dillon, C. Slife, B. Kohler, D. Gibson, G. Gray, and J. S. Franks. Evaluation of red snapper Lutjanus campechanus trophic dynamics with simultaneous stomach content and stable isotope analysis. Marine Ecology Progress Series, 699: 117–134, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/699/m699p117.pdf.

daSilva:2024:PAM

[dSLD24] J. H. Garcia da Silva, L. C. Leal, and G. M. Dias. Palatability of ascidians: a meta-analysis of the predation effect on ascidians. Marine Ecology Progress Series, 742:97–112, August 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/742/m742p097.pdf.

Darnaude:2024:IAR

[DTHC24] A. M. Darnaude, S. E. Tanner, E. Hunter, and F. Costantini. INTRODUCTION: Advancing research in marine functional connectivity for improved policy and management. Marine Ecology Progress Series, 731:1–8, March 13, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m731p001.pdf.

DeCelles:2020:LTC

[DV20] G. R. DeCelles and T. Vidal. Long-term changes in the maturation and growth of Cape Cod/Gulf of Maine yellowtail flounder Limanda ferruginea. Marine Ecology Progress Series, 633:169–180, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/633/m633p169.pdf.

DelaCruz-Pino:2024:SVR

[DVMH^{+}24] G. De la Cruz-Pino, E. Velarde, M. Marín-Hernández, E. R. Inzunza, and E. Ezcurra. Survival vs. reproduction: dimorphic life history strategy in Heermann’s gulls during warm ocean temperature anomalies. Marine Ecology Progress Series, 737:183–191, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m737p183.pdf.

DellaMarta:2023:MPA

[DVP^{+}23] I. R. Della Marta, A. Vergés, S. Powell, S. M. Smith, and A. G. B. Poore. Marine protected areas are linked to higher predation rates by fish in shallow urbanised reefs, but only in no-take reserves. Marine Ecology Progress Series, 721:135– 150, October 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m721p135.pdf.

Daly:2023:PTM

[DVR^{+}23] R. Daly, S. K. Venables, T. D. Rogers, J. D. Filmalter, T. N. Hempson, T. S. Murray, N. E. Hussey, I. Silva, M. A. M. Pereira, B. Q. Mann, B. A. S. Nharreluga, and P. D. Cowley. Persistent transboundary movements of threatened sharks highlight the importance of cooperative management for effective conservation. Marine Ecology Progress Series, 720:117– 131, October 5, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/720/m720p117.pdf.

Dufour:2021:TFI

[DWJL^{+}21] P. Dufour, K. Wojczulanis-Jakubas, S. Lavergne, J. Renaud, D. Jakubas, and S. Descamps. A two-fold increase in migration distance does not have breeding consequences in a long-distance migratory seabird with high flight costs. Marine Ecology Progress Series, 676:117–126, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/676/
m676p117.pdf.

Evans:2020:AFM

[EAS20] J. Evans, E. Arndt, and P. J. Schembri. Atlantic fishes in the Mediterranean: using biological traits to assess the origin of newcomer fishes. Marine Ecology Progress Series, 643:133–143, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/643/m643p133.pdf.

Eger:2020:TCC

[EB20] A. M. Eger and J. K. Baum. Trophic cascades and connectivity in coastal benthic marine ecosystems: a meta-analysis of experimental and observational research. Marine Ecology Progress Series, 656:139–152, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/656/m656p139.pdf.

Elsberry:2021:FRB

[EB21] L. A. Elsberry and M. E. S. Bracken. Functional redundancy buffers mobile invertebrates against the loss of foundation species on rocky shores. Marine Ecology Progress Series, 673:43–54, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/673/m673p043.pdf.

Eldoy:2021:ENS

[EBL^{+}21] S. H. Eldøy, X. Bordeleau, M. J. Lawrence, E. B. Thorstad, A. G. Finstad, F. G. Whoriskey, G. T. Crossin, S. J. Cooke, K. Aarestrup, L. Rønning, A. D. Sjursen, and J. G. Davidsen. The effects of nutritional state, sex and body size on the marine migration behaviour of sea trout. Marine Ecology Progress Series, 665:185–200, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/665/m665p185.pdf.

Esmaeili:2021:CAS

[ECC^{+}21] Y. Shah Esmaeili, G. N. Corte, H. H. Checon, T. R. C. Gomes, J. S. Lefcheck, A. C. Z. Amaral, and A. Turra. Comprehensive assessment of shallow surf zone fish biodiversity requires a combination of sampling methods. Marine Ecology Progress Series, 667:131–144, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2021/667/m667p131.pdf.

Elston:2021:RHU

[ECvBL21] C. Elston, P. D. Cowley, R. G. von Brandis, and J. Lea. Residency and habitat use patterns by sympatric stingrays at a remote atoll in the Western Indian Ocean. Marine Ecology Progress Series, 662:97–114, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/662/m662p097.pdf.

Ezhov:2021:TBD

[EGK^{+}21] A. V. Ezhov, M. V. Gavrilo, Y. V. Krasnov, V. S. Bråthen, B. Moe, A. V. Baranskaya, and H. Strøm. Transpolar and bi-directional migration strategies of black-legged kittiwakes Rissa tridactyla from a colony in Novaya Zemlya, Barents Sea, Russia. Marine Ecology Progress Series, 676:189–203, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/676/m676p189.pdf.

Enzor:2020:EPH

[EHHF^{+}20] L. A. Enzor, C. Hankins, M. Hamilton-Frazier, E. Moso, S. Raimondo, and M. G. Barron. Elevated pCO_{2} and hypoxia alter the acid-base regulation of developing sheepshead minnows Cyprinodon variegatus. Marine Ecology Progress Series, 636:157–168, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/636/m636p157.pdf.

El-Hacen:2023:GPS

[EHLM^{+}23] E. M. El-Hacen, S. Y. C. Lemrabott, K. J. Meijer, T. Piersma, L. L. Govers, J. A. van Gils, and H. Olff. Growth and population structure of bloody cockles Senilia senilis at Banc d’Arguin and Bijagós with different environmental conditions and harvesting regimes. Marine Ecology Progress Series, 710: 71–83, May 4, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m710p071.pdf.

El-Khaled:2020:LSE

[EKRT^{+}20] Y. C. El-Khaled, F. Roth, A. Tilstra, N. Rädecker, D. B. Karcher, B. Kürten, B. H. Jones, C. R. Voolstra, and C. Wild. In situ eutrophication stimulates dinitrogen fixation, denitrification, and productivity in Red Sea coral reefs. Marine Ecology Progress Series, 645:55–66, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/645/
m645p055.pdf.

Edwards:2022:MYA

[EMB^{+}22] M. L. Edwards, M. McCallister, L. R. Brewster, C. W. Bangley, T. H. Curtis, M. B. Ogburn, and M. J. Ajemian. Multiyear assessment of immature bull shark Carcharhinus leucas residency and activity spaces in an expansive estuarine nursery. Marine Ecology Progress Series, 695:125–138, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m695p125.pdf.

Elliott:2022:STD

[EMDR22] R. G. Elliott, J. C. Montgomery, A. Della Penna, and C. A. Radford. Satellite tags describe movement and diving behaviour of blue sharks Prionace glauca in the southwest Pacific. Marine Ecology Progress Series, 689:77–94, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m689p077.pdf.

Estupinan-Montano:2021:OFE

[EMGMEV^{+}21] C. Estupiñán-Montaño, F. Galván-Magaña, F. Elorriaga-Verplancken, M. J. Zetina-Rejón, A. Sánchez-González, C. J. Polo-Silva, D. J. Villalobos-Ramírez, J. Rojas-Cundumí, and A. Delgado-Huertas. Ontogenetic feeding ecology of the scalloped hammerhead shark Sphyrna lewini in the Colombian Eastern Tropical Pacific. Marine Ecology Progress Series, 663:127–143, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/663/m663p127.pdf.

Elsmore:2022:MPF

[ENF^{+}22] K. Elsmore, K. J. Nickols, T. Ford, K. C. Cavanaugh, K. C. Cavanaugh, and B. Gaylord. Macrocystis pyrifera forest development shapes the physical environment through current velocity reduction. Marine Ecology Progress Series, 694: 45–59, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/694/m694p045.pdf.

Emblemsvaag:2020:ITD

[ENRC^{+}20] M. Emblemsvåg, I. Núñez-Riboni, H. T. Christensen, A. Nogueira, A. Gundersen, and R. Primicerio. Increasing temperatures, diversity loss and reorganization of deep-sea fish communities east of Greenland. Marine Ecology Progress Series, 654:127–141, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m654p127.pdf.

Edmunds:2023:DSV

[EP23] P. J. Edmunds and C. T. Perry. Decadal-scale variation in coral calcification on coral-depleted Caribbean reefs. Marine Ecology Progress Series, 713:1–19, June 29, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Esser:2023:NLD

[EPB23] E. A. Esser, J. M. Pringle, and J. E. Byers. Neither larval duration nor dispersal distance predict spatial genetic diversity in planktonic dispersing species. Marine Ecology Progress Series, 721:161–167, October 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/721/
m721p161.pdf.

Echave:2023:CCU

[EPHS23] K. B. Echave, J. L. Pirtle, J. Heifetz, and K. Shotwell. Cautious considerations for using multiple covariate distance sampling and seafloor terrain for improved estimates of rockfish density. Marine Ecology Progress Series, 703:125–143, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/703/m703p125.pdf.

Edmunds:2020:OPU

[ER20] P. J. Edmunds and B. Riegl. Opinion piece: Urgent need for coral demography in a world where corals are disappearing. Marine Ecology Progress Series, 635:233–242, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m635p233.pdf.

Everett:2020:MPR

[ESG20] A. G. Everett, S. T. Szedlmayer, and B. J. Gallaway. Movement patterns of red snapper Lutjanus campechanus based on acoustic telemetry around oil and gas platforms in the northern Gulf of Mexico. Marine Ecology Progress Series, 649:155–173, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/649/m649p155.pdf.

Ensanyar-Volle:2023:DFR

[EVAB^{+}23] O. Ensanyar-Volle, J. Appoo, N. Bunbury, G. Clucas, N. Khan, G. Rocamora, C. Sanchez, and A. L. Fayet. Differences in foraging range between white-tailed tropicbirds breeding on inner and outer Seychelles Islands. Marine Ecology Progress Series, 724:141–154, December 7, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m724p141.pdf.

Egan:2021:PDT

[EVH^{+}21] K. E. Egan, T. S. Viehman, D. M. Holstein, M. Poti, S. H. Groves, and T. B. Smith. Predicting the distribution of threatened orbicellid corals in shallow and mesophotic reef ecosystems. Marine Ecology Progress Series, 667:61–81, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m667p061.pdf.

Elias-Valdez:2023:FEC

[EVVMQ^{+}23] A. Elias-Valdez, E. Velarde, A. Medina-Quej, J. M. Castro-Perez, J. Navarro, and R. Rosas-Luis. Feeding ecology of coexisting Heermann’s gull (Larus heermanni) and elegant tern (Thalasseus elegans) chicks, based on stable isotope measurements. Marine Ecology Progress Series, 712:101–111, June 8, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/712/m712p101.pdf.

Ellis:2024:HDD

[EWBH24] G. K. Ellis, J. Weir, S. Bräger, and S. Heinrich. Hector’s dolphin distribution at Kaikōura before and after a major earthquake. Marine Ecology Progress Series, 748:175–190, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/748/m748p175.pdf.

Essington:2021:HRP

[EWF^{+}21] T. Essington, E. J. Ward, T. B. Francis, C. Greene, L. Kuehne, and D. Lowry. Historical reconstruction of the Puget Sound (USA) groundfish community. Marine Ecology Progress Series, 657:173–189, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/657/m657p173.pdf.

Fairclough:2021:PMT

[Fai21] D. V. Fairclough. Partitioning of marine transition zone reefs among temperate, sub-tropical and tropical fishes is related more to depth and habitat than temperature. Marine Ecology Progress Series, 672:175–192, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/672/m672p175.pdf.

Fierro-Arcos:2021:MFA

[FAJAO^{+}21] D. Fierro-Arcos, J. R. Marín Jarrín, O. Aburto-Oropeza, E. S. Harvey, E. Rastoin-Laplane, and P. Salinas de León. Mangrove fish assemblages reflect the environmental diversity of the Galapagos Islands. Marine Ecology Progress Series, 664:183–205, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/664/m664p183.pdf.

Farabaugh:2024:IEF

[FBC^{+}24a] N. F. Farabaugh, M. E. Bond, D. Chapman, E. Clua, A. R. Harborne, M. Heupel, J. J. Kiszka, and M. R. Heithaus. Incorporating environmental factors is critical for determining conservation baselines for relative abundance of sharks on coral reefs. Marine Ecology Progress Series, 736:93–105, May 23, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/736/m736p093.pdf.

Ferreira:2024:ARI

[FBC^{+}24b] A. Ferreira, A. C. Brito, J. L. Costa, V. Brotas, A. Teles-Machado, and S. Garrido. Anchovy on the rise: Investigating environmental drivers of recruitment strength in the northern Canary Current. Marine Ecology Progress Series, 741: 315–330, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m741p315.pdf.

Fonseca:2024:EDC

[FBC^{+}24c] P. Fonseca, M. Barata, S. Castanho, P. Pousão-Ferreira, and S. Garrido. Effect of diet composition and temperature on the gastric evacuation rate of European sardine: implication for egg predation estimates. Marine Ecology Progress Series, 741: 101–112, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m741p101.pdf.

Franke:2024:SIE

[FBKG24] K. Franke, I. Bartsch, U. Karsten, and A. Graiff. Seasonality influences the effect of warming on kelp photosynthesis. Marine Ecology Progress Series, 739:15–29, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/739/
m739p015.pdf.

Fleming:2024:ECP

[FBM^{+}24] B. F. M. Fleming, S. E. Beaulieu, S. W. Mills, O. E. Gaggiotti, and L. S. Mullineaux. Ecological connectivity in Pacific deep-sea hydrothermal vent metacommunities. Marine Ecology Progress Series, 731:267–278, March 13, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m731p267.pdf.

Flowers:2022:VRS

[FBP^{+}22] K. I. Flowers, E. A. Babcock, Y. P. Papastamatiou, M. E. Bond, N. Lamb, A. Miranda, R. Nuñez, J. Valentin-Albanese, G. M. Clementi, M. C. Kelley, and D. D. Chapman. Varying reef shark abundance trends inside a marine reserve: evidence of a Caribbean reef shark decline. Marine Ecology Progress Series, 683:97–107, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/683/m683p097.pdf.

Franco-Cisterna:2021:TEC

[FCSG21] B. Franco-Cisterna, P. Stief, and R. N. Glud. Temperature effects on carbon mineralization of sinking copepod carcasses. Marine Ecology Progress Series, 679:31–45, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m679p031.pdf.

Fromant:2024:FSB

[FCV^{+}24] A. Fromant, J. Collet, C. Vansteenberghe, R. Musseau, D. Filippi, K. Delord, and C. Barbraud. Fine-scale behaviour and population estimates of endangered sooty albatross suggest low exposure (while not excluding high sensitivity) to bycatch. Marine Ecology Progress Series, 749:181–192, November 21, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/749/m749p181.pdf.

Fukuwaka:2021:RPF

[FD21] M. Fukuwaka and N. D. Davis. Responses of a pelagic fish community to reduced fishing pressure in the central Bering Sea. Marine Ecology Progress Series, 673:183–192, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/673/m673p183.pdf.

Frederiksen:2021:SVV

[FDE^{+}21] M. Frederiksen, S. Descamps, K. H. Elliott, A. J. Gaston, N. P. Huffeldt, Y. Kolbeinsson, J. F. Linnebjerg, E. Lorentzen, F. R. Merkel, H. Strøm, and T. L. Thórarinsson. Spatial variation in vital rates and population growth of thick-billed murres in the Atlantic Arctic. Marine Ecology Progress Series, 672:1–13, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Figueiredo:2020:EFC

[FDF20] C. K. Figueiredo, R. C. Duarte, and A. A. V. Flores. Ecosystem functioning of canopy- and turf- forming algae: contrasting supply of invertebrate prey to pelagic consumers. Marine Ecology Progress Series, 647:79–92, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/647/
m647p079.pdf.

Florko:2021:TMD

[FDL^{+}21] K. R. N. Florko, E. R. Davidson, K. J. Lees, L. J. Hammer, M. F. Lavoie, R. J. Lennox, É. Simard, P. Archambault, M. Auger-Méthé, C. W. McKindsey, F. G. Whoriskey, and N. B. Furey. Tracking movements of decapod crustaceans: a review of a half-century of telemetry-based studies. Marine Ecology Progress Series, 679:219–239, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m679p219.pdf.

Fischer:2021:YRO

[FDS^{+}21] J. H. Fischer, I. Debski, D. B. Spitz, G. A. Taylor, and H. U. Wittmer. Year-round offshore distribution, behaviour, and overlap with commercial fisheries of a critically endangered small petrel. Marine Ecology Progress Series, 660: 171–187, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/660/m660p171.pdf.

Fernandez-Figueroa:2024:FKL

[FFMR24] E. G. Fernandez-Figueroa, S. A. Mapes, and S. R. Rogers. Fish kill lessons and data needs: a spatiotemporal analysis of citizen fish kill reports in coastal SW Florida. Marine Ecology Progress Series, 742:21–33, August 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/742/
m742p021.pdf.

Fortune:2020:SDF

[FFT^{+}20] S. M. E. Fortune, S. H. Ferguson, A. W. Trites, B. LeBlanc, V. LeMay, J. M. Hudson, and M. F. Baumgartner. Seasonal diving and foraging behaviour of Eastern Canada–West Greenland bowhead whales. Marine Ecology Progress Series, 643:197–217, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/643/m643p197.pdf.

Fodrie:2020:DNR

[FHAP20] F. J. Fodrie, K. L. Heck, Jr., C. F. T. Andrus, and S. P. Powers. Determinants of the nursery role of seagrass meadows in the sub-tropical Gulf of Mexico: inshore-offshore connectivity for snapper and grouper. Marine Ecology Progress Series, 647:135–147, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/647/m647p135.pdf.

Figueroa:2021:EFZ

[Fig21] D. F. Figueroa. Environmental forcing on zooplankton distribution in the coastal waters of the Galápagos Islands: spatial and seasonal patterns in the copepod community structure. Marine Ecology Progress Series, 661:49–69, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m661p049.pdf.

Franklin:2023:OMR

[FJB^{+}23] P. A. Franklin, D. Jellyman, C. Baker, K. Birnie-Gauvin, P. Jellyman, and K. Aarestrup. Oceanic migration routes and behaviour of the New Zealand longfin eel Anguilla dieffenbachii. Marine Ecology Progress Series, 724:127–139, December 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/724/m724p127.pdf.

Friedlaender:2021:SRP

[FJJ^{+}21] A. S. Friedlaender, T. Joyce, D. W. Johnston, A. J. Read, D. P. Nowacek, J. A. Goldbogen, N. Gales, and J. W. Durban. Sympatry and resource partitioning between the largest krill consumers around the Antarctic Peninsula. Marine Ecology Progress Series, 669:1–16, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Fujinami:2022:NIS

[FKS^{+}22] Y. Fujinami, A. Kurashima, K. Shiozaki, Y. Hiraoka, Y. Semba, S. Ohshimo, H. Nakano, and M. Kai. New insights into spatial segregation by sex and life-history stage in blue sharks Prionace glauca in the northwestern Pacific. Marine Ecology Progress Series, 696:69–84, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/696/
m696p069.pdf.

Furukawa:2020:HVM

[FKTK20] S. Furukawa, A. Kozuka, T. Tsuji, and H. Kubota. Horizontal and vertical movement of yellowtails Seriola quinqueradiata during summer to early winter recorded by archival tags in the northeastern Japan Sea. Marine Ecology Progress Series, 636:139–156, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/636/m636p139.pdf.

Friess:2021:RSV

[FLBP^{+}21] C. Friess, S. K. Lowerre-Barbieri, G. R. Poulakis, N. Hammerschlag, J. M. Gardiner, A. M. Kroetz, K. Bassos-Hull, J. Bickford, E. C. Bohaboy, R. D. Ellis, H. Menendez, W. F.

Patterson III, M. E. Price, J. S. Rehage, C. P. Shea, M. J. Smukall, S. Walters Burnsed, K. A. Wilkinson, J. Young, A. B. Collins, B. C. DeGroot, C. T. Peterson, C. Purtlebaugh, M. Randall, R. M. Scharer, R. W. Schloesser, T. R. Wiley, G. A. Alvarez, A. J. Danylchuk, A. G. Fox, R. D. Grubbs, A. Hill, J. V. Locascio, P. M. O’Donnell, G. B. Skomal, F. G. Whoriskey, and L. P. Griffin. Regional-scale variability in the movement ecology of marine fishes revealed by an integrative acoustic tracking network. Marine Ecology Progress Series, 663:157–177, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/663/m663p157.pdf.

Fraile:2024:AVY

[FLC^{+}24a] I. Fraile, P. L. Luque, S. E. Campana, J. H. Farley, K. Krusic-Golub, N. Clear, J. P. Eveson, I. Artetxe-Arrate, I. Zudaire, H. Murua, and G. Merino. Age validation of yellowfin tuna Thunnus albacares in the Indian Ocean using post-peak bomb radiocarbon chronologies. Marine Ecology Progress Series, 734:91–104, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2024/734/m734p091.pdf.

Friscourt:2024:SOB

[FLC^{+}24b] N. Friscourt, M. A. Lea, Y. Cherel, S. Wotherspoon, E. A. Brewer, W. C. Oosthuizen, P. J. N. de Bruyn, M. Wege, M. E. Goebel, P. N. Trathan, and A. Walters. Seasonal and ocean basin-scale assessment of amino acid δ ^{15}N trends in a Southern Ocean marine predator. Marine Ecology Progress Series, 747:151–169, October 17, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m747p151.pdf.

Flynn:2021:MIU

[FLCQ21] P. Tummon Flynn, K. D. Lynn, D. K. Cairns, and P. A. Quijón. Mesograzer interactions with a unique strain of Irish moss Chondrus crispus: colonization, feeding, and algal condition-related effects. Marine Ecology Progress Series, 669:83–96, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/669/m669p083.pdf.

Ferreira:2024:HST

[FLY^{+}24] A. S. A. Ferreira, Ø. Langangen, N. A. Yaragina, I. P. Prokopchuk, and J. M. Durant. How the spatio-temporal overlap of cod, haddock, and capelin larvae affects their recruitment in the Norwegian–Barents Sea system. Marine Ecology Progress Series, 734:79–90, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/734/
m734p079.pdf.

Franze:2020:CTG

[FMD20] G. Franzè and S. Menden-Deuer. Common temperaturegrowth dependency and acclimation response in three herbivorous protists. Marine Ecology Progress Series, 634:1–13, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ferreira:2024:ETC

[FN24] A. S. A. Ferreira and A. B. Neuheimer. Estimating a thermal constant of spawning to explain spawning time of Pacific herring Clupea pallasii across space and time. Marine Ecology Progress Series, 741:237–249, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/741/
m741p237.pdf.

Fuji:2023:BIP

[FNH^{+}23] T. Fuji, S. I. Nakayama, M. Hashimoto, H. Miyamoto, Y. Kamimura, S. Furuichi, K. Oshima, and S. Suyama. Biological interactions potentially alter the large-scale distribution pattern of the small pelagic fish, Pacific saury Cololabis saira. Marine Ecology Progress Series, 704:99–117, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m704p099.pdf.

Forbes:2023:SIA

[FNL^{+}23] R. A. Forbes, B. J. Nakamoto, N. S. Lysiak, T. Wimmer, and B. Hayden. Stable isotope analysis of baleen from North Atlantic right whales Eubalaena glacialis reflects distribution shift to the Gulf of St. Lawrence. Marine Ecology Progress Series, 722:177–193, November 9, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/722/
m722p177.pdf.

Fletcher:2021:PIC

[FNM^{+}21] D. Fletcher, J. Newman, S. McKechnie, C. Bragg, P. Dillingham, R. Clucas, D. Scott, S. Uhlmann, P. O. Lyver, A. M. Gormley, S. Bull, K. Davis, R. Davis, R. Davis, T. Davis, L. Edwards, J. Kitson, T. Nixon, M. Skerrett, and H. Moller. Projected impacts of climate change, bycatch, harvesting, and predation on the Aotearoa New Zealand tītī Ardenna grisea population. Marine Ecology Progress Series, 670:223– 238, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/670/m670p223.pdf.

Fujita:2023:PFT

[FNO^{+}23] K. Fujita, H. Nishizawa, J. Okuyama, M. Arita, S. Takuma, T. Narazaki, and A. Watabe. Polymorphic foraging tactics in a marine reptile: insight from horizontal movement and dive behavior analyses. Marine Ecology Progress Series, 707:115–129, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/707/m707p115.pdf.

Frost:2020:DTN

[FNS^{+}20] M. Frost, F. C. Neat, D. Stirling, V. Bendall, L. R. Noble, and C. S. Jones. Distribution and thermal niche of the common skate species complex in the north-east Atlantic. Marine Ecology Progress Series, 656:65–74, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/656/
m656p065.pdf.

Foret:2020:IPC

[FOM^{+}20] M. Forêt, F. Olivier, P. Miner, G. Winkler, J. B. Nadalini, and R. Tremblay. Influence of the physiological condition of bivalve recruits on their post-settlement dispersal potential. Marine Ecology Progress Series, 636:77–89, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/636/
m636p077.pdf.

Flanders:2020:UNG

[FOO20] K. R. Flanders, Z. H. Olson, and K. A. Ono. Utilizing nextgeneration sequencing to identify prey DNA in western North Atlantic grey seal Halichoerus grypus diet. Marine Ecology Progress Series, 655:227–240, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/655/m655p227.pdf.

Freitas:2020:MBM

[FP20] R. F. Freitas and P. R. Pagliosa. Mangrove benthic macrofauna: drivers of community structure and functional traits at multiple spatial scales. Marine Ecology Progress Series, 638: 25–38, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/638/m638p025.pdf.

Foley:2021:RMP

[FPB^{+}21] H. J. Foley, K. Pacifici, R. W. Baird, D. L. Webster, Z. T. Swaim, and A. J. Read. Residency and movement patterns of Cuvier’s beaked whales Ziphius cavirostris off Cape Hatteras, North Carolina, USA. Marine Ecology Progress Series, 660:203–216, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m660p203.pdf.

Faure:2021:CTD

[FPG^{+}21] J. Faure, C. Péron, N. Gasco, F. Massiot-Granier, J. Spitz, C. Guinet, and P. Tixier. Contribution of toothfish depredated on fishing lines to the energy intake of killer whales off the Crozet Islands: a multi-scale bioenergetic approach. Marine Ecology Progress Series, 668:149–161, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m668p149.pdf.

Fonseca-Ponce:2022:PED

[FPZJAO^{+}22] I. A. Fonseca-Ponce, A. A. Zavala-Jiménez, O. Aburto-Oropeza, A. Maldonado-Gasca, F. Galván-Magaña, R. González-Armas, and J. D. Stewart. Physical and environmental drivers of oceanic manta ray Mobula birostris sightings at an aggregation site in Bahía de Banderas, Mexico. Marine Ecology Progress Series, 694:133–148, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/694/m694p133.pdf.

Feitosa:2023:GFH

[FQVMF23] J. L. L. Feitosa, L. V. M. V. Queiroz-Véras, M. Maida, and B. P. Ferreira. Going further on herbivore fishing: the removal of smaller fishes from algal-dominated reefs. Marine Ecology Progress Series, 713:117–132, June 29, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/713/
m713p117.pdf.

Fox:2021:SAA

[FRO^{+}21] C. H. Fox, C. Robertson, P. D. O’Hara, R. Tadey, and K. H. Morgan. Spatial assessment of albatrosses, commercial fisheries, and bycatch incidents on Canada’s Pacific coast. Marine Ecology Progress Series, 672:205–222, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m672p205.pdf.

Faletti:2021:LHT

[FS21] M. E. Faletti and C. D. Stallings. Life history through the eyes of a hogfish: trophic growth and differential juvenile habitat use from stable isotope analysis. Marine Ecology Progress Series, 666:183–202, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2021/666/m666p183.pdf.

Fields:2022:FSL

[FS22] J. B. Fields and N. J. Silbiger. Foundation species loss alters multiple ecosystem functions within temperate tidepool communities. Marine Ecology Progress Series, 683:1–19, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Fonseca:2022:SSV

[FSAG22] P. Fonseca, A. D. Silva, M. M. Angélico, and S. Garrido. Seasonal and spatial variability of atlanto-iberian pelagic fish diet with estimates of intraguild predation. Marine Ecology Progress Series, 687:95–111, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/687/m687p095.pdf.

Fennie:2020:PTW

[FSDB20] H. W. Fennie, S. Sponaugle, E. A. Daly, and R. D. Brodeur. Prey tell: what quillback rockfish early life history traits reveal about their survival in encounters with juvenile Coho salmon. Marine Ecology Progress Series, 650:7–18, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/650/
m650p007.pdf.

Fenton:2024:HDL

[FSF^{+}24] K. F. Fenton, N. H. Salmond, S. E. Foreman, J. S. Curtis, T. Jowett, C. Savage, and S. R. Wing. High densities of large tuaki, the New Zealand cockle Austrovenus stutchburyi, provide a post-settlement predation refuge for conspecific juveniles. Marine Ecology Progress Series, 726:85– 98, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m726p085.pdf.

Fujinami:2021:SMP

[FSH^{+}21] Y. Fujinami, K. Shiozaki, Y. Hiraoka, Y. Semba, S. Ohshimo, and M. Kai. Seasonal migrations of pregnant blue sharks Prionace glauca in the northwestern Pacific. Marine Ecology Progress Series, 658:163–179, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/658/m658p163.pdf.

Fujioka:2021:HUA

[FSK^{+}21] K. Fujioka, K. Sasagawa, T. Kuwahara, E. E. Estess, Y. Takahara, K. Komeyama, T. Kitagawa, C. J. Farwell, S. Furukawa, J. Kinoshita, H. Fukuda, M. Kato, A. Aoki, O. Abe, S. Ohshimo, and N. Suzuki. Habitat use of adult Pacific bluefin tuna Thunnus orientalis during the spawning season in the Sea of Japan: evidence for a trade-off between thermal preference and reproductive activity. Marine Ecology Progress Series, 668:1–20, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ferreira:2020:MMD

[FSN^{+}20] A. S. A. Ferreira, L. C. Stige, A. B. Neuheimer, B. Bogstad, N. Yaragina, I. Prokopchuk, and J. M. Durant. Match– mismatch dynamics in the Norwegian–Barents Sea system.

Marine Ecology Progress Series, 650:81–94, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m650p081.pdf.

Faulk:2023:TSD

[FSS23] L. G. Faulk, T. Smart, and J. P. Stone. Temporal and spatial distribution of the cannonball jellyfish Stomolophus meleagris in the South Atlantic Bight, USA. Marine Ecology Progress Series, 717:51–65, August 24, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/717/m717p051.pdf.

Fraser:2020:TCM

[FSSL^{+}20] K. M. Fraser, R. D. Stuart-Smith, S. D. Ling, F. J. Heather, and G. J. Edgar. Taxonomic composition of mobile epifaunal invertebrate assemblages on diverse benthic microhabitats from temperate to tropical reefs. Marine Ecology Progress Series, 640:31–43, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/640/m640p031.pdf.

Freeman:2022:IMP

[FST22] J. B. Freeman, B. X. Semmens, and A. R. Thompson. Impacts of marine protected areas and the environment on larval rockfish species richness and assemblage structure in the Southern California Bight. Marine Ecology Progress Series, 698:125–137, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/698/m698p125.pdf.

Fauchald:2021:YRD

[FTA^{+}21] P. Fauchald, A. Tarroux, F. Amélineau, V. S. Bråthen, S. Descamps, M. Ekker, H. H. Helgason, M. K. Johansen, B. Merkel, B. Moe, J. Åström, T. Anker-Nilssen, O. Bjørnstad, O. Chastel, S. Christensen-Dalsgaard, J. Danielsen, F. Daunt, N. Dehnhard, K. E. Erikstad, A. Ezhov, M. Gavrilo, G. T. Hallgrimsson, E. S. Hansen, M. Harris, M. Helberg, J. E. Jónsson, Y. Kolbeinsson, Y. Krasnov, M. Langset, S. H. Lorentsen, E. Lorentzen, M. Newell, B. Olsen, T. K. Reiertsen, G. H. Systad, P. Thompson, T. L. Thórarinsson, S. Wanless, K. Wojczulanis-Jakubas, and H. Strøm. Year-round distribution of Northeast Atlantic seabird populations: applications for population management and marine spatial planning. Marine Ecology Progress Series, 676:255–276, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m676p255.pdf.

Fijn:2022:RES

[FTA^{+}22] R. C. Fijn, C. B. Thaxter, G. Aarts, J. Adema, R. P. Middelveld, and R. S. A. van Bemmelen. Relative effects of static and dynamic abiotic conditions on foraging behaviour in breeding sandwich terns. Marine Ecology Progress Series, 692:137–150, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m692p137.pdf.

Freer:2020:ECD

[FTC^{+}20] J. J. Freer, G. A. Tarling, M. A. Collins, J. C. Partridge, and M. J. Genner. Estimating circumpolar distributions of lanternfish using 2D and 3D ecological niche models. Marine Ecology Progress Series, 647:179–193, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/647/
m647p179.pdf.

Faillettaz:2020:STP

[FVC^{+}20] R. Faillettaz, R. Voué, R. Crec’hriou, L. H. Garsi, G. Lecaillon, S. Agostini, P. Lenfant, and J. O. Irisson. Spatiotemporal patterns of larval fish settlement in the northwestern Mediterranean Sea. Marine Ecology Progress Series, 650: 153–173, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/650/m650p153.pdf.

Franz:2023:FDM

[FWA^{+}23] M. Franz, L. Whyte, T. C. Atwood, D. Menning, S. A. Sonsthagen, S. L. Talbot, K. L. Laidre, E. Gonzalez, and M. A. McKinney. Fecal DNA metabarcoding shows credible shortterm prey detections and explains variation in the gut microbiome of two polar bear subpopulations. Marine Ecology Progress Series, 704:131–147, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/704/m704p131.pdf.

Flaherty-Walia:2023:SBR

[FWGD^{+}23] K. E. Flaherty-Walia, C. L. Gervasi, J. P. Davis, S. Trabelsi, Z. Fratto, R. O. Santos, P. W. Stevens, R. E. Matheson, Jr., and J. S. Rehage. Seines and baited remote underwater video document consistent spatiotemporal patterns in nekton communities of subtropical coastal habitats. Marine Ecology Progress Series, 722:37–64, November 9, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/722/
m722p037.pdf.

Feddern:2022:RDC

[FWWH22] M. L. Feddern, E. J. Ward, A. J. Warlick, and G. W. Holtgrieve. Recent divergent changes in Alaskan pinniped trophic position detected using compound-specific stable isotope analysis. Marine Ecology Progress Series, 688:153– 166, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/688/m688p153.pdf.

Furuichi:2020:DEC

[FYK^{+}20] S. Furuichi, T. Yasuda, H. Kurota, M. Yoda, K. Suzuki, M. Takahashi, and M. Fukuwaka. Disentangling the effects of climate and density-dependent factors on spatiotemporal dynamics of Japanese sardine spawning. Marine Ecology Progress Series, 633:157–168, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/633/m633p157.pdf.

Farley:2024:CPM

[FYM^{+}24] E. V. Farley, Jr., E. M. Yasumiishi, J. M. Murphy, W. Strasburger, F. Sewall, K. Howard, S. Garcia, and J. H. Moss. Critical periods in the marine life history of juvenile western Alaska chum salmon in a changing climate. Marine Ecology Progress Series, 726:149–160, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m726p149.pdf.

Gandra:2020:IIA

[GAF20] M. Gandra, P. Afonso, and J. Fontes. Intra- and interspecific associations in two predatory reef fishes at a shallow seamount. Marine Ecology Progress Series, 654:93–107, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/654/m654p093.pdf.

Gerrity:2020:EEA

[GAFS20] S. Gerrity, T. Alestra, H. S. Fischman, and D. R. Schiel. Earthquake effects on abalone habitats and populations in southern New Zealand. Marine Ecology Progress Series, 656: 153–161, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/656/m656p153.pdf.

Garrido:2024:RLT

[GAPM24] S. Garrido, M. Albo-Puigserver, and M. Moyano. Review: Larval trophic ecology of small pelagic fishes: a review of recent advances and pathways to fill remaining knowledge gaps. Marine Ecology Progress Series, 741:127–143, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/741/m741p127.pdf.

Gradinger:2020:FAA

[GB20] R. Gradinger and B. A. Bluhm. First analysis of an Arctic sea ice meiofauna food web based on abundance, biomass and stable isotope ratios. Marine Ecology Progress Series, 634: 29–43, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/634/m634p029.pdf.

Grogan:2021:PES

[GB21] S. V. Grogan and S. S. Bell. Propagule establishment and seedling growth in a shifting mangrove population: performance in novel tidal conditions. Marine Ecology Progress Series, 679:59–71, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/679/m679p059.pdf.

Glazner:2021:PRV

[GBA21] R. Glazner, J. Ballard, and A. R. Armitage. Predation refuge values of marsh and mangrove vegetation for the marsh periwinkle Littoraria irrorata. Marine Ecology Progress Series, 672:153–162, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/672/m672p153.pdf.

Griffin:2022:SVP

[GBA^{+}22] L. P. Griffin, J. W. Brownscombe, A. J. Adams, P. E. Holder, A. Filous, G. A. Casselberry, J. K. Wilson, R. E. Boucek, S. K. Lowerre-Barbieri, A. Acosta, D. Morley, S. J. Cooke, and A. J. Danylchuk. Seasonal variation in the phenology of Atlantic tarpon in the Florida Keys: migration, occupancy, repeatability, and management implications. Marine Ecology Progress Series, 684:133–155, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/684/m684p133.pdf.

Gunther:2024:UPD

[GBA24] K. M. Gunther, M. R. Baker, and K. Y. Aydin. Using predator diets to infer forage fish distribution and assess responses to climate variability in the eastern Bering Sea. Marine Ecology Progress Series, 741:71–99, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/741/
m741p071.pdf.

Grutter:2020:ECF

[GBC^{+}20] A. S. Grutter, S. Bejarano, K. L. Cheney, A. W. Goldizen, T. Sinclair-Taylor, and P. A. Waldie. Effects of the cleaner fish Labroides dimidiatus on grazing fishes and coral reef benthos. Marine Ecology Progress Series, 643:99–114, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/643/
m643p099.pdf.

Gokturk:2022:LSO

[GBE^{+}22] E. N. Gokturk, B. S. Bartlett, B. Erisman, W. Heyman, and R. G. Asch. Loss of suitable ocean habitat and phenological shifts among grouper and snapper spawning aggregations in the Greater Caribbean under climate change. Marine Ecology Progress Series, 699:91–115, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m699p091.pdf.

Gulka:2023:AIM

[GBF^{+}23] J. Gulka, A. M. Berlin, K. D. Friedland, A. T. Gilbert, C. Goetsch, W. A. Montevecchi, M. Perry, I. J. Stenhouse, K. A. Williams, and E. M. Adams. Assessing individual movement, habitat use, and behavior of non-breeding marine birds in relation to prey availability in the US Atlantic. Marine Ecology Progress Series, 711:77–99, May 19, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m711p077.pdf.

Geissinger:2023:CSW

[GBG^{+}23] E. A. Geissinger, C. R. Bloom, R. S. Gregory, B. J. Laurel, and P. V. R. Snelgrove. Condition, size, and winter duration affect winter survival probability of juvenile Atlantic cod Gadus morhua in a coastal subarctic ecosystem. Marine Ecology Progress Series, 711:47–60, May 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m711p047.pdf.

Gerwing:2020:ASP

[GBH^{+}20] T. G. Gerwing, M. A. Barbeau, D. J. Hamilton, A. M. Allen Gerwing, J. Sinclair, L. Campbell, M. M. Davies, B. Harvey, F. Juanes, and S. E. Dudas. Assessment of sediment penetrability as an integrated in situ measure of intertidal softsediment conditions. Marine Ecology Progress Series, 648: 67–78, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/648/m648p067.pdf.

Gawinski:2024:SPB

[GBSS24] C. Gawinski, S. L. Basedow, A. Sundfjord, and C. Svensen. Secondary production at the Barents Sea polar front in summer: contribution of different size classes of mesozooplankton. Marine Ecology Progress Series, 735:77–101, May 2, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m735p077.pdf.

Gallo:2020:DOT

[GBW^{+}20] N. D. Gallo, M. Beckwith, C. L. Wei, L. A. Levin, L. Kuhnz, and J. P. Barry. Dissolved oxygen and temperature best predict deep-sea fish community structure in the Gulf of California with climate change implications. Marine Ecology Progress Series, 637:159–180, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/637/m637p159.pdf.

Griffen:2020:REC

[GCCM20] B. D. Griffen, Z. J. Cannizzo, J. Carver, and M. Meidell. Reproductive and energetic costs of injury in the mangrove tree crab. Marine Ecology Progress Series, 640:127–137, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m640p127.pdf.

Garcia-Corral:2020:IUR

[GCDA20] L. S. García-Corral, C. M. Duarte, and S. Agusti. Impact of UV radiation on plankton net community production: responses in Western Australian estuarine and coastal waters. Marine Ecology Progress Series, 651:45–56, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m651p045.pdf.

Gancel:2021:USP

[GCDP21] H. N. Gancel, R. H. Carmichael, J. Du, and K. Park. Use of settlement patterns and geochemical tagging to test population connectivity of eastern oysters Crassostrea virginica. Marine Ecology Progress Series, 673:85–105, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m673p085.pdf.

Gimenez-Casalduero:2020:PNM

[GCGCAS^{+}20] F. Giménez-Casalduero, F. Gomariz-Castillo, F. Alonso-Sarría, E. Cortés, A. Izquierdo-Muñoz, and A. A. Ramos-Esplá. Pinna nobilis in the Mar Menor coastal lagoon: a story of colonization and uncertainty. Marine Ecology Progress Series, 652:77–94, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/652/m652p077.pdf.

Gadeken:2023:SMR

[GD23] K. J. Gadeken and K. M. Dorgan. Sediment macrofaunal response to the diel oxygen cycle. Marine Ecology Progress Series, 703:67–80, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m703p067.pdf.

Grilli:2023:AMB

[GDAC23] M. Graña Grilli, A. Di Virgilio, P. A. E. Alarcón, and Y. Cherel. Apparent mismatch between stable isotopes and foraging habitat suggests high secondary ingestion of Antarctic krill in brown skuas. Marine Ecology Progress Series, 708:163–176, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/708/m708p163.pdf.

Glass:2020:STV

[GDCP20] J. R. Glass, R. Daly, P. D. Cowley, and D. M. Post. Spatial trophic variability of a coastal apex predator, the giant trevally Caranx ignobilis, in the western Indian Ocean. Marine Ecology Progress Series, 641:195–208, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/641/
m641p195.pdf.

Gusmao:2022:WDM

[GDG^{+}22] J. B. Gusmao, O. Diaz, C. Gallardo, V. Hidalgo-Ruz, T. Kiessling, A. Mecho, E. Meerhoff, N. Rozbaczylo, and M. Thiel. Water depth and mesoscale oceanography drive neustonic polychaete assemblages in the SE Pacific Ocean. Marine Ecology Progress Series, 682:123–136, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/682/
m682p123.pdf.

Gonzalez-Espinosa:2020:PCW

[GED20] P. C. González-Espinosa and S. D. Donner. Predicting coldwater bleaching in corals: role of temperature, and potential integration of light exposure. Marine Ecology Progress Series, 642:133–146, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/642/m642p133.pdf.

Groner:2021:WSS

[GEY^{+}21] M. L. Groner, M. E. Eisenlord, R. M. Yoshioka, E. A. Fiorenza, P. D. Dawkins, O. J. Graham, M. Winningham, A. Vompe, N. D. Rivlin, B. Yang, C. A. Burge, B. Rappazzo, C. P. Gomes, and C. D. Harvell. Warming sea surface temperatures fuel summer epidemics of eelgrass wasting disease. Marine Ecology Progress Series, 679:47–58, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/679/
m679p047.pdf.

Gonzalez:2024:IEH

[GFDN^{+}24] J. G. Gonzalez, T. Frédou, P. J. Duarte-Neto, M. Labonne, A. S. Lira, A. P. C. Soares, and A. M. Darnaude. Importance of estuaries for the horse-eye jack Caranx latus in northeastern Brazil: a case study of underestimated connectivity at the land-sea interface. Marine Ecology Progress Series, 731:197– 215, March 13, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m731p197.pdf.

Garcia-Fernandez:2020:SVM

[GFDPASR20] C. García-Fernández, R. Domínguez-Petit, N. Aldanondo, and F. Saborido-Rey. Seasonal variability of maternal effects in European hake Merluccius merluccius. Marine Ecology Progress Series, 650:125–140, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m650p125.pdf.

Gilmour:2022:LAE

[GFH^{+}22] M. E. Gilmour, J. Felis, M. Hester, L. Young, and J. Adams. Laysan albatross exhibit complex behavioral plasticity in the subtropical and subarctic North Pacific Ocean. Marine Ecology Progress Series, 697:125–147, September 22, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m697p125.pdf.

Garde:2023:TST

[GFK^{+}23] B. Garde, A. Fell, K. Krishnan, C. G. Jones, R. Gunner, V. Tatayah, N. C. Cole, E. Lempidakis, and E. L. C. Shepard. Thermal soaring in tropicbirds suggests that diverse seabirds may use this strategy to reduce flight costs. Marine Ecology Progress Series, 723:171–183, November 23, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m723p171.pdf.

Gama:2021:TEJ

[GFT^{+}21] L. R. Gama, M. M. P. B. Fuentes, T. H. Trevizani, F. Pellizzari, G. E. Lemons, J. A. Seminoff, and C. Domit. Trophic ecology of juvenile green turtles in the Southwestern Atlantic Ocean: insights from stable isotope analysis and niche modelling. Marine Ecology Progress Series, 678:139–152, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/678/m678p139.pdf.

Garagouni:2023:PSP

[GG23] M. Garagouni and K. Ganias. Prey selection patterns of bottlenose dolphins depredating bottom-set gillnets in the northern Aegean Sea. Marine Ecology Progress Series, 710:125– 135, May 4, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/710/m710p125.pdf.

Gao:2022:DSN

[GGCL22] X. Gao, Y. Gong, X. Chen, and Y. Li. Dietary shifts and niche partitioning throughout ontogeny reduce intraspecific competition in a pelagic generalist predator. Marine Ecology Progress Series, 692:81–97, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/692/m692p081.pdf.

Galindo:2021:FHT

[GGN21] E. Galindo, A. Giraldo, and A. F. Navia. Feeding habits and trophic interactions of four sympatric hammerhead shark species reveal trophic niche partitioning. Marine Ecology Progress Series, 665:159–175, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/665/m665p159.pdf.

Garcia-Garcia:2023:VPP

[GGRV23] L. M. García-García and M. Ruiz-Villarreal. Variability of plankton production during the spring bloom in NW Iberia. Marine Ecology Progress Series, 708:45–78, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m708p045.pdf.

Gruss:2020:ESC

[GGT^{+}20] A. Grüss, J. Gao, J. T. Thorson, C. N. Rooper, G. Thompson, J. L. Boldt, and R. Lauth. Estimating synchronous changes in condition and density in eastern Bering Sea fishes. Marine Ecology Progress Series, 635:169–185, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/635/
m635p169.pdf.

Gonzalez-Irusta:2021:LTC

[GIDP^{+}21] J. M. González-Irusta, A. De la Torriente, A. Punzón, M. Blanco, J. C. Arronte, R. Bañón, J. E. Cartes, and A. Serrano. Living at the top. connectivity limitations and summit depth drive fish diversity patterns in an isolated seamount. Marine Ecology Progress Series, 670:121–137, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/670/
m670p121.pdf.

Gauci:2024:TPS

[GJK^{+}24] C. Gauci, A. Jueterbock, A. Khatei, G. Hoarau, and I. Bartsch. Thermal priming of Saccharina latissima: a promising strategy to improve seaweed production and restoration in future climates. Marine Ecology Progress Series, 745:59–71, September 19, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m745p059.pdf.

Gulka:2020:ICF

[GJM^{+}20] J. Gulka, E. Jenkins, L. D. Maynard, W. A. Montevecchi, P. M. Regular, and G. K. Davoren. Inter-colony foraging dynamics and breeding success relate to prey availability in a pursuit-diving seabird. Marine Ecology Progress Series, 651:183–198, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/651/m651p183.pdf.

Gross:2022:HDA

[GKK^{+}22] E. Gross, W. Kimmerer, J. Korman, L. Lewis, S. Burdick, and L. Grimaldo. Hatching distribution, abundance, and losses to freshwater diversions of longfin smelt inferred using hydrodynamic and particle-tracking models. Marine Ecology Progress Series, 700:179–196, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2022/700/m700p179.pdf.

Garaffo:2020:FDR

[GLB^{+}20] G. V. Garaffo, E. N. Llanos, M. A. Saracho Bottero, E. Hines, R. Elías, and M. M. L. Jaubet. Functional diversity on rocky shores of the SW Atlantic: sewage effluents influence and mask the effects of the latitudinal gradient. Marine Ecology Progress Series, 648:39–49, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/648/m648p039.pdf.

Gosnell:2021:NCE

[GLG21] J. S. Gosnell, E. A. Levine, and E. M. Goetz. Nonconsumptive effects of predators on oysters differ based on predator identity and duration (but not frequency) of exposure. Marine Ecology Progress Series, 660:95–104, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/660/m660p095.pdf.

Gu:2024:NFA

[GLM^{+}24] Y. Gu, S. Lin, Y. Mo, L. Li, M. Ma, J. Li, S. Lin, H. Yuan, C. Zhu, H. Luo, and W. Zhang. Niche features and assembly mechanisms of microeukaryotic generalists and specialists along a north–south gradient of a subtropical coastal sea. Marine Ecology Progress Series, 742:35–57, August 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2024/742/m742p035.pdf.

Galloway:2024:TEC

[GLP^{+}24] A. S. Galloway, K. Lyons, D. S. Portnoy, A. M. Barker, D. H. Adams, J. Gelsleichter, E. A. Reyier, and B. S. Frazier. Trophic ecology of Carolina Sphyrna gilberti and scalloped S. lewini hammerheads in the southeastern USA. Marine Ecology Progress Series, 743:25–46, August 22, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m743p025.pdf.

Glencross:2021:BSS

[GLW21] J. S. Glencross, J. L. Lavers, and E. J. Woehler. Breeding success of short-tailed shearwaters following extreme environmental conditions. Marine Ecology Progress Series, 672: 193–203, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/672/m672p193.pdf.

Gangur:2020:FFM

[GM20] A. N. Gangur and D. J. Marshall. Facultative feeding in a marine copepod: effects of larval food and temperature on performance. Marine Ecology Progress Series, 652:33– 47, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/652/m652p033.pdf.

Gurr:2024:EFS

[GMB^{+}24] S. J. Gurr, K. McFarland, G. Bernatchez, M. S. Dixon, L. Guy, L. M. Milke, M. E. Poach, D. Hart, L. V. Plough, D. H. Redman, G. Sennefelder, S. Stiles, G. H. Wikfors, D. K. Padilla, and S. L. Meseck. Effects of food supply on northern bay scallops Argopecten irradians reared under two pCO_{2} conditions. Marine Ecology Progress Series, 740:61–78, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m740p061.pdf.

Gonzalez-Murcia:2022:EED

[GMCE^{+}22] S. González-Murcia, A. G. Coppock, M. Ekins, C. N. Battershill, and G. P. Jones. Effects of exposure, depth and aspect on sponge communities on a coral reef. Marine Ecology Progress Series, 685:111–126, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/685/m685p111.pdf.

Guerra:2022:SBC

[GMLC22] A. S. Guerra, D. J. McCauley, D. Lecchini, and J. E. Caselle. Shoaling behavior of coral reef fishes varies between two islands with different predator abundance. Marine Ecology Progress Series, 690:133–145, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/690/m690p133.pdf.

Garcia-Mayoral:2022:DCW

[GMRMG22] E. García-Mayoral, Á. Roura, A. Moreno, and Á. F. González. Diet composition of wild Loligo vulgaris paralarvae along the West Iberian Peninsula coast. Marine Ecology Progress Series, 681:71–85, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/681/m681p071.pdf.

Green:2021:BSA

[GMS^{+}21] S. J. Green, J. K. Matley, D. E. Smith, B. Castillo II, J. L. Akins, R. S. Nemeth, C. Pollock, and K. Reale-Munroe. Broad-scale acoustic telemetry reveals long-distance movements and large home ranges for invasive lionfish on Atlantic coral reefs. Marine Ecology Progress Series, 673:117– 134, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m673p117.pdf.

Gong:2023:MAP

[GNC^{+}23] A. Y. Gong, A. P. Nosal, D. P. Cartamil, J. M. Anderson, L. F. Bellquist, N. J. Ben-Aderet, K. M. Blincow, E. S. Burns, C. Caldow, R. M. Freedman, R. K. Logan, C. G. Lowe, B. X. Semmens, B. S. Stirling, C. F. White, and P. A. Hastings. Migration, aggregation, and philopatry of two nearshore elasmobranch species in the Southern California Bight. Marine Ecology Progress Series, 724:111–125, December 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/724/
m724p111.pdf.

Grear:2020:ECA

[GON^{+}20] J. S. Grear, C. A. O’Leary, J. A. Nye, S. T. Tettelbach, and C. J. Gobler. Effects of coastal acidification on North Atlantic bivalves: interpreting laboratory responses in the context of in situ populations. Marine Ecology Progress Series, 633:89–104, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/633/m633p089.pdf.

Guenther:2022:ETP

[GPCM22] R. Guenther, E. M. A. Porcher, E. Carrington, and P. T. Martone. Effects of temperature and pH on the growth, calcification, and biomechanics of two species of articulated coralline algae. Marine Ecology Progress Series, 700:79–93, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m700p079.pdf.

Gartenstein:2024:LEM

[GPMH^{+}24] S. Gartenstein, A. Pérez-Matus, F. J. Heather, N. Godoy, F. Torres-Cañete, A. M. Catalán, and N. Valdivia. Limited effects of management and ecological subsidies on the size-spectra of kelp forest fish communities. Marine Ecology Progress Series, 732:135–147, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/732/
m732p135.pdf.

Gottesman:2024:DWF

[GPP^{+}24] H. M. Gottesman, W. F. Patterson III, E. B. Peebles, T. S. Switzer, and S. K. Lowerre-Barbieri. Distinguishing West Florida gag (Mycteroperca microlepis) nurseries with otolith stable isotopes. Marine Ecology Progress Series, 749:127–139, November 21, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/749/m749p127.pdf.

Garcia-Quintas:2024:POT

[GQBB^{+}24] A. Garcia-Quintas, P. Bustamante, C. Barbraud, A. Lorrain, D. Denis, and S. Lanco. Plasticity and overlap of trophic niches in tropical breeding laridae. Marine Ecology Progress Series, 742:131–142, August 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m742p131.pdf.

Gimenez:2022:OSA

[GRB22] L. H. Gimenez, R. J. Rivera, and A. Brante. One step ahead of sea anemone invasions with ecological niche modeling: potential distributions and niche dynamics of three successful invasive species. Marine Ecology Progress Series, 690: 83–95, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/690/m690p083.pdf.

Glasgow:2021:EEF

[GRQ21] D. M. Glasgow, M. J. M. Reichert, and J. Quattro. Effects of environmental factors on reef fish assemblage structure in the southeastern US Atlantic. Marine Ecology Progress Series, 671:147–163, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/671/m671p147.pdf.

Gilbert:2021:SMH

[GRTK^{+}21] L. Gilbert, E. Rouby, E. Tew-Kaï, J. Spitz, H. Peltier, V. Quilfen, and M. Authier. Spatiotemporal models highlight influence of oceanographic conditions on common dolphin bycatch risk in the Bay of Biscay. Marine Ecology Progress Series, 679:195–212, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/679/m679p195.pdf.

Ginther:2020:LES

[GS20] S. C. Ginther and M. A. Steele. Limited effects of Sargassum horneri, an invasive alga, on temperate reef fish assemblages. Marine Ecology Progress Series, 643:115–131, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/643/
m643p115.pdf.

Gilson:2021:STV

[GSBO21] A. R. Gilson, D. A. Smale, M. T. Burrows, and N. E. O’Connor. Spatio-temporal variability in the deposition of beach-cast kelp (wrack) and inter-specific differences in degradation rates. Marine Ecology Progress Series, 674:89–102, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m674p089.pdf.

Garcia:2022:EDL

[GSC^{+}22] V. Garcia, H. T. Schilling, D. O. Cruz, S. M. Hawes, J. D. Everett, M. Roughan, A. G. Miskiewicz, E. A. Pakhomov, A. Jeffs, and I. M. Suthers. Entrainment and development of larval fish assemblages in two contrasting cold core eddies of the East Australian current system. Marine Ecology Progress Series, 685:1–18, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Gottesman:2020:SIK

[GSK^{+}20] B. L. Gottesman, J. Sprague, D. J. Kushner, K. Bellisario, D. Savage, M. F. McKenna, D. L. Conlin, E. DiDonato, M. J. Barkaszi, M. B. Halvorsen, and B. C. Pijanowski. Soundscapes indicate kelp forest condition. Marine Ecology Progress Series, 654:35–52, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/654/m654p035.pdf.

Glaspie:2020:PPM

[GSL20] C. N. Glaspie, R. D. Seitz, and R. N. Lipcius. Are predator– prey model predictions supported by empirical data? Evidence for a storm-driven shift to an alternative stable state in a crab-clam system. Marine Ecology Progress Series, 645: 83–90, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m645p083.pdf.

Galloway:2024:TSC

[GSL^{+}24a] A. W. E. Galloway, J. B. Schram, A. T. Lowe, R. Whippo, S. Heiser, K. Iken, J. B. McClintock, A. G. Klein, M. O. Amsler, and C. D. Amsler. Trophic stability and change across a sea ice cover gradient on the western Antarctic Peninsula. Marine Ecology Progress Series, 735:43–61, May 2, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m735p043.pdf.

Gama:2024:DCJ

[GSL^{+}24b] L. R. Gama, J. A. Seminoff, G. E. Lemons, M. M. P. B. Fuentes, F. Pellizzari, M. R. C. Meira-Filho, L. Rosa, G. M. Vélez-Rubio, E. L. da Silveira, and C. Domit. Diet composition of juvenile green turtles in the Southwestern Atlantic Ocean: long-term insights from a beach stranding program. Marine Ecology Progress Series, 727:159–179, January 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/727/m727p159.pdf.

Gorbunov:2020:MSF

[GSN^{+}20] M. Y. Gorbunov, E. Shirsin, E. Nikonova, V. V. Fadeev, and P. G. Falkowski. A multi-spectral fluorescence induction and relaxation (FIRe) technique for physiological and taxonomic analysis of phytoplankton communities. Marine Ecology Progress Series, 644:1–13, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Gibbs:2020:LES

[GSP^{+}20] M. Gibbs, E. Scanes, L. Parker, M. Byrne, W. O’Connor, P. Virtue, and P. Ross. Larval energetics of the Sydney rock oyster Saccostrea glomerata and Pacific oyster Magallana gigas. Marine Ecology Progress Series, 656:51–64, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/656/m656p051.pdf.

Goldsworthy:2020:SAS

[GSPH20] D. S. Goldsworthy, B. J. Saunders, J. R. C. Parker, and E. S. Harvey. Spatial assemblage structure of shallow-water reef fish in Southwest Australia. Marine Ecology Progress Series, 649:125–140, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/649/m649p125.pdf.

Gleiber:2020:FWC

[GSRC20] M. R. Gleiber, S. Sponaugle, K. L. Robinson, and R. K. Cowen. Food web constraints on larval growth in subtropical coral reef and pelagic fishes. Marine Ecology Progress Series, 650:19–36, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/650/m650p019.pdf.

Gorecki:2022:EST

[GSS22] R. Gorecki, M. N. Schrandt, and T. S. Switzer. Evidence of shared trends in juvenile fish recruitment to nearshore seagrass habitats of the eastern Gulf of Mexico. Marine Ecology Progress Series, 694:157–174, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/694/m694p157.pdf.

Grimes:2020:CIR

[GTH^{+}20] T. M. Grimes, M. T. Tinker, B. B. Hughes, K. E. Boyer, L. Needles, K. Beheshti, and R. L. Lewison. Characterizing the impact of recovering sea otters on commercially important crabs in California estuaries. Marine Ecology Progress Series, 655:123–137, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/655/m655p123.pdf.

Garcia-Vernet:2022:SSI

[GVAZ^{+}22] R. García-Vernet, A. Aguilar, J. Zafra, G. Víkingsson, S. D. Halldórsson, and A. Borrell. Sulfur stable isotope ratios provide further insight into movements of the fin whale, an oceanic long-range migrant. Marine Ecology Progress Series, 692:185–194, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/692/m692p185.pdf.

Gomez-Valdez:2024:HIE

[GVCSO^{+}24] M. M. Gómez-Valdez, L. Carvalho-Saucedo, L. Ocampo, J. L. Gutiérrez-González, and D. B. Lluch-Cota. Hermaphroditism and the impact of environmental variables on atresia prevalence in the pen shell Pinna rugosa. Marine Ecology Progress Series, 726:71–83, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/726/
m726p071.pdf.

Gomez-Valdez:2021:RAS

[GVOCSGG21] M. M. Gómez-Valdez, L. Ocampo, L. Carvalho-Saucedo, and J. L. Gutiérrez-González. Reproductive activity and seasonal variability in the biochemical composition of a pen shell, Atrina maura. Marine Ecology Progress Series, 663: 99–113, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/663/m663p099.pdf.

Gagnon:2020:CNS

[GVR^{+}20] K. Gagnon, E. A. Virtanen, P. Rusanen, M. Nurmi, M. Viitasalo, and V. Jormalainen. Cormorants have negligible seascape-scale impacts on benthic vegetation communities. Marine Ecology Progress Series, 654:195–207, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/654/
m654p195.pdf.

Gardner:2023:EMG

[GVS23] S. T. Gardner, A. N. Valenza, and F. S. Scharf. Ecological mechanisms generating variable first-year growth in Paralichthys lethostigma, a coastal flatfish. Marine Ecology Progress Series, 707:77–97, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/707/m707p077.pdf.

Gillett:2023:CMI

[GWA^{+}23] D. J. Gillett, S. B. Weisberg, S. R. Alin, D. Cadien, R. Velarde, K. Barwick, C. Larsen, and A. Latker. Changes in the macrobenthic infaunal community of the Southern California continental margin over five decades in relation to oceanographic factors. Marine Ecology Progress Series, 722:65–88, November 9, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/722/m722p065.pdf.

Gouezo:2021:MLS

[GWC^{+}21] M. Gouezo, E. Wolanski, K. Critchell, K. Fabricius, P. Harrison, Y. Golbuu, and C. Doropoulos. Modelled larval supply predicts coral population recovery potential following disturbance. Marine Ecology Progress Series, 661:127–145, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/661/m661p127.pdf.

Guerra:2020:SIA

[GWDR20] M. Guerra, L. Wing, S. Dawson, and W. Rayment. Stable isotope analyses reveal seasonal and inter-individual variation in the foraging ecology of sperm whales. Marine Ecology Progress Series, 638:207–219, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/638/m638p207.pdf.

Hayes:2022:RVS

[HAB^{+}22] C. T. Hayes, S. B. Alford, B. A. Belgrad, K. M. Correia, M. Z. Darnell, B. T. Furman, M. O. Hall, C. W. Martin, A. M. Mc-Donald, D. L. Smee, and K. M. Darnell. Regional variation in seagrass complexity drives blue crab Callinectes sapidus mortality and growth across the northern Gulf of Mexico. Marine Ecology Progress Series, 693:141–155, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/693/
m693p141.pdf.

Hooks:2024:VPR

[HB24] A. P. Hooks and S. C. Burgess. Variation in polyandry, reproductive output, and within-brood genetic diversity in a marine snail population across seasons and years. Marine Ecology Progress Series, 734:65–77, April 18, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m734p065.pdf.

Hendy:2020:MAT

[HBC^{+}20] I. W. Hendy, O. Burt, S. Clough, L. Young, and S. M. Cragg. Mosquitofish avoid thermal stress by moving from open water to the shade of the mangrove Rhizophora mangle. Marine Ecology Progress Series, 637:103–116, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/637/
m637p103.pdf.

Hebert:2022:TTE

[HBE22] C. E. Hebert, N. M. Burgess, and J. E. Elliott. Temporal trends of essential omega-3 fatty acids in Atlantic and Pacific food webs as measured in eggs of Leach’s stormpetrel Oceanodroma leucorhoa. Marine Ecology Progress Series, 684:199–210, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m684p199.pdf.

Howell:2022:PDS

[HBG^{+}22] K. L. Howell, A. E. Bridges, K. P. Graves, L. Allcock, G. la Bianca, C. Ventura-Costa, S. Donaldson, A. L. Downie, T. Furey, F. McGrath, and R. Ross. Performance of deepsea habitat suitability models assessed using independent data, and implications for use in area-based management. Marine Ecology Progress Series, 695:33–51, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/695/

m695p033.pdf. See corrigendum [Ano22c].

Hunter:2021:GAR

[HBJ^{+}21] M. E. Hunter, C. E. Beaver, N. A. Johnson, E. K. Bors, A. A. Mignucci-Giannoni, B. R. Silliman, D. Buddo, L. Searle, and E. Díaz-Ferguson. Genetic analysis of red lionfish Pterois volitans from Florida, USA, leads to alternative North Atlantic introduction scenarios. Marine Ecology Progress Series, 675:133–151, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m675p133.pdf.

Hancock:2021:CHO

[HBR^{+}21] J. R. Hancock, A. R. Barrows, T. C. Roome, A. S. Huffmyer, S. B. Matsuda, N. J. Munk, S. A. Rahnke, and C. Drury. Coral husbandry for ocean futures: leveraging abiotic factors to increase survivorship, growth, and resilience in juvenile Montipora capitata. Marine Ecology Progress Series, 657:123–133, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/657/m657p123.pdf.

Holland:2021:FSS

[HBS^{+}21] M. M. Holland, A. Becker, J. A. Smith, J. D. Everett, and I. M. Suthers. Fine-scale spatial and diel dynamics of zooplanktivorous fish on temperate rocky and artificial reefs. Marine Ecology Progress Series, 674:221–239, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/674/
m674p221.pdf.

Herrera:2024:HRC

[HCP^{+}24] M. A. Herrera, D. Cardeñosa, Y. P. Papastamatiou, J. Vaudo, C. Bermúdez-Rivas, and M. Shivji. High residency of a critically endangered hammerhead shark to a small area: implications for marine protected area management and design. Marine Ecology Progress Series, 743:47–63, August 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/743/m743p047.pdf.

Hyman:2024:OPJ

[HCS^{+}24] A. C. Hyman, G. S. Chiu, M. S. Seebo, A. Smith, G. G. Saluta, and R. N. Lipcius. Ontogenetic patterns in juvenile blue crab density: effects of habitat and turbidity in a Chesapeake Bay tributary. Marine Ecology Progress Series, 729: 135–150, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m729p135.pdf.

Hutt:2023:UBN

[HCvdHM23] J. C. Hütt, P. Corkeron, J. M. van der Hoop, and M. J. Moore. Uncertain bioenergetics of North Atlantic right whales. Marine Ecology Progress Series, 725:167–184, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m725p167.pdf.

Henderson:2022:BII

[HDA^{+}22] E. E. Henderson, M. Deakos, J. Aschettino, D. Englehaupt, and G. Alongi. Behavior and inter-island movements of satellite-tagged humpback whales in Hawai’i, USA. Marine Ecology Progress Series, 685:197–213, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m685p197.pdf.

Harris:2021:BCC

[HDB21] B. Harris, S. K. Davy, and J. J. Bell. Benthic community composition of temperate mesophotic ecosystems (TMEs) in New Zealand: sponge domination and contribution to habitat complexity. Marine Ecology Progress Series, 671: 21–43, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/671/m671p021.pdf.

Hrebien:2021:IFW

[HDE21] V. Hrebien, E. Deschaseaux, and B. D. Eyre. Isoprene fluxes from warm temperate and tropical seagrass communities. Marine Ecology Progress Series, 676:1–17, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Henderson:2021:PBU

[HDL^{+}21] E. M. Henderson, A. E. Derocher, N. J. Lunn, B. Montpetit, E. H. Merrill, and E. S. Richardson. Polar bear Ursus maritimus use of the western Hudson Bay flaw lead. Marine Ecology Progress Series, 664:227–242, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/664/
m664p227.pdf.

Harrison:2024:LTO

[HDM^{+}24] D. E. Harrison, M. S. Diluvio, E. Matveev, P. V. F. Corrêa, F. C. De Leo, and S. P. Leys. Long-term observations of a sponge in situ reveal a rich repertoire of contractile behaviors, including winter dormancy. Marine Ecology Progress Series, 748:33–52, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m748p033.pdf.

Hirche:2024:FBU

[HEKH24] H. J. Hirche, E. A. Ershova, K. N. Kosobokova, and R. R. Hopcroft. From fringe to basin: unravelling the survival strategies of Calanus hyperboreus and C. glacialis in the Arctic Ocean. Marine Ecology Progress Series, 745:41–57, September 19, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m745p041.pdf.

Hunt:2024:LTC

[HEP^{+}24] G. L. Hunt, G. H. Engelhard, J. K. Pinnegar, J. Barry, B. D. Wigham, and N. V. C. Polunin. Long-term change in a North Sea inshore fish assemblage between 1899–1913 and 2018–2019. Marine Ecology Progress Series, 733:79– 93, April 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m733p079.pdf.

Hansen:2024:ADC

[HESF24] F. T. Hansen, A. C. Erichsen, C. Saurel, and P. S. Freitas. Assessing the demographic connectivity of common cockles in a shallow estuary as a basis for fisheries management and stock protection efforts. Marine Ecology Progress Series, 731: 293–313, March 13, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m731p293.pdf.

Hewson:2021:OPM

[Hew21] I. Hewson. Opinion piece: Microbial respiration in the asteroid diffusive boundary layer influenced sea star wasting disease during the 2013–2014 northeast Pacific Ocean mass mortality event. Marine Ecology Progress Series, 668:231–237, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m668p231.pdf.

Harley:2024:FCT

[HGH^{+}24] J. R. Harley, M. H. Grinnell, K. Hebert, J. Cleary, M. Thompson, and C. N. Rooper. Forecasted changes to the timing of Pacific herring Clupea pallasii spawn in a warming ocean. Marine Ecology Progress Series, 740:95–108, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/740/m740p095.pdf.

Harty:2022:DDW

[HGK^{+}22] K. Harty, M. Guerrero, A. M. Knochel, G. M. W. Stevens, A. Marshall, K. Burgess, and J. D. Stewart. Demographics and dynamics of the world’s largest known population of oceanic manta rays Mobula birostris in coastal Ecuador. Marine Ecology Progress Series, 700:145–159, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m700p145.pdf.

Harris:2024:ECA

[HGK^{+}24] J. Harris, L. Gardner, A. S. Kahn, A. D. Ridlon, and K. Wasson. Experiments in conservation aquaculture to optimize restoration for Olympia oysters Ostrea lurida in Elkhorn Slough, CA, USA. Marine Ecology Progress Series, 734:45– 64, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/734/m734p045.pdf.

Hammerschlag:2022:USR

[HGR^{+}22] N. Hammerschlag, L. F. G. Gutowsky, M. J. Rider, R. Roemer, and A. J. Gallagher. Urban sharks: residency patterns of marine top predators in relation to a coastal metropolis. Marine Ecology Progress Series, 691:1–17, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Haselmair:2021:BWI

[HGT^{+}21] A. Haselmair, I. Gallmetzer, A. Tomašových, A. M. Wieser, A. Übelhör, and M. Zuschin. Basin-wide infaunalisation of benthic soft-bottom communities driven by anthropogenic habitat degradation in the northern Adriatic Sea. Marine Ecology Progress Series, 671:45–65, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/671/
m671p045.pdf.

Hu:2023:OAP

[HHBP23] N. Hu, J. Hollander, C. Brönmark, and A. Persson. Ocean acidification and predation risk, in isolation and in combination, show strong effects on marine mussels. Marine Ecology Progress Series, 721:71–83, October 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/721/
m721p071.pdf.

Hammer:2022:GAC

[HHM^{+}22] L. J. Hammer, N. E. Hussey, M. Marcoux, H. Pettitt-Wade, K. Hedges, R. Tallman, and N. B. Furey. Arctic char Salvelinus alpinus movement dynamics relative to ice breakup in a high Arctic embayment. Marine Ecology Progress Series, 682:221–236, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/682/m682p221.pdf.

Harper:2023:MYC

[HHO^{+}23] L. M. Harper, L. K. Huebner, E. D. O’Cain, R. Ruzicka, D. F. Gleason, and N. D. Fogarty. Multi-year coral recruitment study across the Florida Reef Tract reveals boomor-bust pattern among broadcast spawners and consistency among brooders. Marine Ecology Progress Series, 721:39– 58, October 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m721p039.pdf.

Hjollo:2021:AIZ

[HHS21] S. S. Hjøllo, C. Hansen, and M. D. Skogen. Assessing the importance of zooplankton sampling patterns with an ecosystem model. Marine Ecology Progress Series, 680:163–176, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/680/m680p163.pdf.

Hu:2023:HDL

[HHY^{+}23] N. Hu, Y. Huang, Z. Yu, T. Zhang, D. Liu, and M. Lee. Handling- or digestion-limited predators: the role of body mass and habitat complexity in predator functional response. Marine Ecology Progress Series, 725:1–14, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hiraoka:2022:ABF

[HIT^{+}22] Y. Hiraoka, T. Ishihara, A. Tawa, Y. Tanaka, S. Ohshimo, and Y. Ando. Association between fatty acid signature and growth rate of larval Pacific bluefin tuna in two major spawning grounds. Marine Ecology Progress Series, 689: 127–136, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/689/m689p127.pdf.

Hiley:2021:BRH

[HJG21] H. M. Hiley, V. M. Janik, and T. Götz. Behavioural reactions of harbour porpoises Phocoena phocoena to startle-eliciting stimuli: movement responses and practical applications. Marine Ecology Progress Series, 672:223–241, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/672/
m672p223.pdf.

Hilliard:2020:CSA

[HKD^{+}20] J. Hilliard, D. Karlen, T. Dix, S. Markham, and A. Schulze. Comparative species abundance modeling of Capitellidae (Annelida) in Tampa Bay, Florida, USA. Marine Ecology Progress Series, 653:105–119, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/653/m653p105.pdf.

Hamabata:2023:TYC

[HKK^{+}23] T. Hamabata, M. Kawata, S. Kondo, A. Matsuo, Y. Suyama, K. Suzuki, and K. Kameda. Twenty-year changes in the composition of a mixed stock of foraging green turtles in the Yaeyama Islands of Japan. Marine Ecology Progress Series, 716:93–105, August 10, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2023/716/m716p093.pdf.

Hane:2020:RTE

[HKY^{+}20] Y. Hane, S. Kimura, Y. Yokoyama, Y. Miyairi, T. Ushikubo, T. Ishimura, N. Ogawa, T. Aono, and K. Nishida. Reconstruction of temperature experienced by Pacific bluefin tuna Thunnus orientalis larvae using SIMS and microvolume CF-IRMS otolith oxygen isotope analyses. Marine Ecology Progress Series, 649:175–188, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m649p175.pdf.

Hamilton:2021:MMH

[HLA^{+}21] C. D. Hamilton, C. Lydersen, J. Aars, M. Biuw, A. N. Boltunov, E. W. Born, R. Dietz, L. P. Folkow, D. M. Glazov, T. Haug, M. P. Heide-Jørgensen, L. E. Kettemer, K. L. Laidre, N. Øien, E. S. Nordøy, A. H. Rikardsen, A. Rosing-Asvid, V. Semenova, O. V. Shpak, S. Sveegaard, F. Ugarte, Ø. Wiig, and K. M. Kovacs. Marine mammal hotspots in the Greenland and Barents Seas. Marine Ecology Progress Series, 659:3–28, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

He:2023:SFC

[HLCH23] X. He, M. Lei, F. Cheng, and S. Hu. In situ food compositions reveal niche partitioning in small marine cladocerans and copepods in Daya Bay, South China Sea. Marine Ecology Progress Series, 716:47–61, August 10, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/716/
m716p047.pdf.

Haberman:2023:CCA

[HM23] I. Haberman and P. T. Martone. Calcified coralline algae have similar caloric value to uncalcified algae. Marine Ecology Progress Series, 713:173–179, June 29, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/713/
m713p173.pdf.

Hidalgo:2022:CDE

[HMB^{+}22] F. J. Hidalgo, S. H. Miller, K. Borst, J. Dozier, and D. Breitburg. Context-dependent effects of ocean acidification on the interaction between a crab predator and its oyster prey. Marine Ecology Progress Series, 693:39–54, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/693/
m693p039.pdf.

Hsiung:2022:EEF

[HMK^{+}22] K. M. Hsiung, C. Ma, C. Y. Ko, Y. H. Tseng, Y. C. Kuo, and Y. S. Han. Effects of environmental factors within the spawning area and migration routes on the length of Anguilla japonica glass eels recruited to Taiwan. Marine Ecology Progress Series, 683:109–121, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/683/m683p109.pdf.

Hurtado-McCormick:2020:SLR

[HMKK^{+}20] V. Hurtado-McCormick, T. Kahlke, D. Krix, A. Larkum, P. J. Ralph, and J. R. Seymour. Seagrass leaf reddening alters the microbiome of Zostera muelleri. Marine Ecology Progress Series, 646:29–44, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/646/m646p029.pdf.

Harvey:2021:LMC

[HMT21] H. R. Harvey, R. McMahon, and K. A. Taylor. Lipid markers and compound-specific carbon isotopes as diet and biosynthesis reflectors in the northern Neptune whelk Neptunea heros. Marine Ecology Progress Series, 673:69–84, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m673p069.pdf.

Hogan:2022:IOR

[HMV^{+}22] S. Hogan, E. A. K. Murphy, M. P. Volaric, M. C. N. Castorani, P. Berg, and M. A. Reidenbach. Influence of oyster reefs on infauna and sediment spatial distributions within intertidal mudflats. Marine Ecology Progress Series, 686: 91–106, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/686/m686p091.pdf.

Hernandez:2021:SIT

[HMW^{+}21] M. F. Hernandez, S. R. Midway, L. West, H. Tillya, and M. J. Polito. Stable isotopes track the ontogenetic movement of three commercially important fishes along a coastal Tanzanian seascape. Marine Ecology Progress Series, 670: 139–154, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/670/m670p139.pdf.

Heidmann:2024:DMS

[HNB^{+}24] S. L. Heidmann, R. S. Nemeth, C. R. Biggs, E. Kadison, and B. L. Kojis. Diel movements and space use of Lutjanus analis at a spawning aggregation site, examined to evaluate the efficacy of a seasonal closed area for management. Marine Ecology Progress Series, 739:207–225, July 4, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m739p207.pdf.

Hayakawa:2021:AAA

[HNKK21] J. Hayakawa, K. Nakamoto, M. Kodama, and T. Kawamura. Aggregation of adult abalone Haliotis discus hannai during the spawning season, and its associations with seasonal and interannual changes in the macroalgal community. Marine Ecology Progress Series, 670:105–120, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/670/
m670p105.pdf.

Hendrick:2024:DFM

[HNN^{+}24] G. C. Hendrick, M. D. Nicholson, P. Narvaez, D. Sun, A. Packard, A. S. Grutter, and P. C. Sikkel. Diel fish migration facilitates functional connectivity of coral reef and seagrass habitats via transport of ectoparasites. Marine Ecology Progress Series, 731:249–265, March 13, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m731p249.pdf.

Heggie:2021:RVA

[HO21] K. Heggie and M. B. Ogburn. Rapid video assessment detects qualitative differences in oyster reef habitat. Marine Ecology Progress Series, 667:219–224, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/667/m667p219.pdf.

Henseler:2023:CFM

[HO23] C. Henseler and D. Oesterwind. A comparison of fishing methods to sample coastal fish communities in temperate seagrass meadows. Marine Ecology Progress Series, 715:91– 111, July 27, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/715/m715p091.pdf.

Huizenga:2024:IJL

[HO24] K. Huizenga and C. Oviatt. Inshore juvenile lobsters threatened by warming waters and migratory fish predators in southern New England. Marine Ecology Progress Series, 728: 183–197, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2024/728/m728p183.pdf.

Homma:2020:PSG

[HOK^{+}20] Y. Homma, S. Okuda, M. Kasahara, F. Takahashi, S. Yoshikawa, and S. Uwai. Phenological shifts and genetic differentiation between sympatric populations of Sargassum horneri (Fucales, Phaeophyceae) in Japan. Marine Ecology Progress Series, 642:103–116, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/642/m642p103.pdf.

Hudson:2022:SEA

[HOK^{+}22] K. Hudson, M. J. Oliver, J. Kohut, M. S. Dinniman, J. M. Klinck, M. A. Cimino, K. S. Bernard, H. Statscewich, and W. Fraser. A subsurface eddy associated with a submarine canyon increases availability and delivery of simulated Antarctic krill to penguin foraging regions. Marine Ecology Progress Series, 702:105–122, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m702p105.pdf.

Haase:2020:DDD

[HOP^{+}20] K. Haase, A. Orio, J. Pawlak, M. Pachur, and M. Casini. Diet of dominant demersal fish species in the Baltic Sea: Is flounder stealing benthic food from cod? Marine Ecology Progress Series, 645:159–170, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/645/m645p159.pdf.

Hammond:2023:LLT

[HP23] A. R. Hammond and S. W. Purcell. Limited long-term movement and slow growth of the sea cucumber Pearsonothuria graeffei. Marine Ecology Progress Series, 704:1–14, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hamilton:2022:ECE

[HPDG22] B. R. Hamilton, C. T. Peterson, A. Dawdy, and R. D. Grubbs. Environmental correlates of elasmobranch and large fish distribution in a river-dominated estuary. Marine Ecology Progress Series, 688:83–98, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/688/m688p083.pdf. See

erratum [Ano22f].

Hromadkova:2020:SSR

[HPFB20] T. Hromádková, V. Pavel, J. Flousek, and M. Briedis. Seasonally specific responses to wind patterns and ocean productivity facilitate the longest animal migration on Earth. Marine Ecology Progress Series, 638:1–12, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hill:2020:SNE

[HPH20] J. M. Hill, P. S. Petraitis, and K. L. Heck, Jr. Submergence, nutrient enrichment, and tropical storm impacts on Spartina alterniflora in the microtidal northern Gulf of Mexico. Marine Ecology Progress Series, 644:33–45, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/644/
m644p033.pdf.

Hernandez:2021:LIV

[HPRP21] K. M. Hernandez, W. B. Puryear, J. A. Runstadler, and M. J. Polito. Little interannual variability in gray seal (Halichoerus grypus) trophic niches during pregnancy despite variable environmental conditions. Marine Ecology Progress Series, 667:207–217, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/667/m667p207.pdf.

Huebert:2021:SFP

[HRC21] K. B. Huebert, K. A. Rose, and R. C. Chambers. Simulating fish population responses to elevated CO_{2}: a case study using winter flounder. Marine Ecology Progress Series, 680:137–161, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m680p137.pdf.

Harbison:2022:ACF

[HRH22] T. Harbison, M. Rogers, and S. Henkel. Assessing commercial fishery bait in Dungeness crab Cancer magister feeding ecology: δ ^{13}C and δ ^{15}N stable isotope and gut content analysis. Marine Ecology Progress Series, 700:125–143, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/700/
m700p125.pdf.

Haro:2021:OSA

[HSA^{+}21] D. Haro, P. Sabat, J. Acevedo, J. Capella, B. Cáceres, A. Aguayo-Lobo, and F. Martínez. Ontogenetic and seasonal analysis of the diet and isotopic niche of humpback whales in the Magellan Strait, Chile. Marine Ecology Progress Series, 669:213–226, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/669/m669p213.pdf.

Haram:2020:ENN

[HSB20] L. E. Haram, E. E. Sotka, and J. E. Byers. Effects of novel, non-native detritus on decomposition and invertebrate community assemblage. Marine Ecology Progress Series, 643: 49–61, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/643/m643p049.pdf.

Holmquist:2021:SPT

[HSBB^{+}21] J. R. Holmquist, L. Schile-Beers, K. Buffington, M. Lu, T. J. Mozdzer, J. Riera, D. E. Weller, M. Williams, and J. P. Megonigal. Scalability and performance tradeoffs in quantifying relationships between elevation and tidal wetland plant communities. Marine Ecology Progress Series, 666:57–72, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m666p057.pdf.

Henly:2024:LTR

[HSNT24] L. Henly, J. E. Stewart, J. Newton, and C. Trueman. Low trophic redundancy among temperate wrasse species implies ecosystem risks associated with a multi-species inshore fishery. Marine Ecology Progress Series, 734:105–121, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2024/734/m734p105.pdf.

Hildebrand:2022:TOP

[HSO^{+}22] L. Hildebrand, F. A. Sullivan, R. A. Orben, S. Derville, and L. G. Torres. Trade-offs in prey quantity and quality in gray whale foraging. Marine Ecology Progress Series, 695:189–201, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m695p189.pdf.

Harbour:2021:BCS

[HSSN^{+}21] R. P. Harbour, C. R. Smith, C. Simon-Nutbrown, M. Cecchetto, E. Young, C. Coral, and A. K. Sweetman. Biodiversity, community structure and ecosystem function on kelp and wood falls in the Norwegian deep sea. Marine Ecology Progress Series, 657:73–91, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/657/m657p073.pdf.

Hansen:2021:PEE

[HSU^{+}21] C. Hansen, M. D. Skogen, K. Rong Utne, C. Broms, E. Strand, and S. S. Hjøllo. Patterns, efficiency and ecosystem effects when fishing Calanus finmarchicus in the Norwegian Sea — using an individual-based model. Marine Ecology Progress Series, 680:15–32, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m680p015.pdf.

Howson:2022:FEJ

[HTGG22] U. A. Howson, T. E. Targett, P. A. Grecay, and P. M. Gaffney. Foraging by estuarine juveniles of two paralichthyid flounders: experimental analyses of the effects of light level, turbidity, and prey type. Marine Ecology Progress Series, 695:139–156, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/695/m695p139.pdf.

Hsiung:2020:LEL

[HTL^{+}20] A. R. Hsiung, W. T. Tan, L. H. L. Loke, L. B. Firth, E. C. Heery, J. Ducker, V. Clark, Y. S. Pek, W. R. Birch, A. C. F. Ang, R. S. Hartanto, T. M. F. Chai, and P. A. Todd. Little evidence that lowering the pH of concrete supports greater biodiversity on tropical and temperate seawalls. Marine Ecology Progress Series, 656:193–205, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m656p193.pdf.

Hatzonikolakis:2021:IGR

[HTT^{+}21] Y. Hatzonikolakis, K. Tsiaras, G. Tserpes, S. Somarakis, M. A. St John, P. Peristeraki, D. E. Raitsos, and G. Triantafyllou. Investigating growth and reproduction of the Mediterranean swordfish Xiphias gladius through a full life cycle bioenergetics model. Marine Ecology Progress Series, 680:51–77, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m680p051.pdf.

Hudspith:2021:QSH

[HvdSR^{+}21] M. Hudspith, J. van der Sprong, L. Rix, D. Víg, J. Schoorl, and J. M. de Goeij. Quantifying sponge host and microbial symbiont contribution to dissolved organic matter uptake through cell separation. Marine Ecology Progress Series, 670:1–13, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Hjollo:2021:TSM

[HvLM21] S. S. Hjøllo, S. M. van Leeuwen, and M. Maar. Theme section: Marine research and management topics addressed by process-based ecosystem models. Marine Ecology Progress Series, 680:1–6, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m680p001.pdf.

Hui:2020:STV

[HW20] T. Y. Hui and G. A. Williams. Spatio-temporal variation of sediment properties reveals missing trophic links for depositfeeding crabs in sandy shore food webs. Marine Ecology Progress Series, 654:79–92, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/654/m654p079.pdf.

Howard:2022:EPC

[HWA^{+}22] B. R. Howard, D. T. S. Wong, V. Aguiar, J. Desforges, E. M. Oishi, J. Stewart, and I. M. Côté. Effects of perceived competition and water temperature on the functional responses of invasive and native crabs. Marine Ecology Progress Series, 684:69–78, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/684/m684p069.pdf.

Huang:2022:PPA

[HWT^{+}22] K. Huang, Z. Wang, J. Tan, D. Wang, X. Dai, J. Cen, L. Ou, and S. Lu. Phosphomonoesterase and phosphodiesterase activities and their regulation during dinoflagellate blooms under different external phosphate conditions. Marine Ecology Progress Series, 698:41–54, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/698/m698p041.pdf.

Harris:2020:DTP

[HYG^{+}20] L. N. Harris, D. J. Yurkowski, M. J. H. Gilbert, B. G. T. Else, P. J. Duke, M. M. M. Ahmed, R. F. Tallman, A. T. Fisk, and J. S. Moore. Depth and temperature preference of anadromous Arctic char Salvelinus alpinus in the Kitikmeot Sea, a shallow and low-salinity area of the Canadian Arctic. Marine Ecology Progress Series, 634:175–197, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/634/
m634p175.pdf.

Iken:2023:MIC

[IAG^{+}23] K. Iken, C. D. Amsler, K. B. Gorman, A. G. Klein, A. W. E. Galloway, M. O. Amsler, S. Heiser, R. Whippo, A. T. Lowe, J. B. Schram, Z. X. Schneider, and J. B. McClintock. Macroalgal input into the coastal food web along a gradient of seasonal sea ice cover along the Western Antarctic Peninsula. Marine Ecology Progress Series, 718:1–22, September 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ibanez:2021:GCV

[IBC^{+}21] C. M. Ibáñez, J. Bravo, S. A. Carrasco, M. J. Carter, and M. A. Aguilera. Grazer commensalism varies across the species range edge: host chiton size influences epibiont limpet incidence and spatial segregation. Marine Ecology Progress Series, 674:131–141, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2021/674/m674p131.pdf.

Indergard:2022:AET

[IBRL^{+}22] M. O. Indergard, A. Bellantuono, M. Rodriguez-Lanetty, F. Heng, and M. R. Gilg. Acclimation to elevated temperatures in Acropora cervicornis: effects of host genotype and symbiont shuffling. Marine Ecology Progress Series, 701: 41–65, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/701/m701p041.pdf.

Ishikawa:2024:AAN

[ICH^{+}24] N. F. Ishikawa, C. Chen, R. Hashimoto, N. O. Ogawa, D. Uyeno, and H. Nomaki. Amino acid nitrogen isotopic compositions show seep copepods gain nutrition from host animals. Marine Ecology Progress Series, 727:81–90, January 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/727/m727p081.pdf.

Ishida-Castaneda:2023:CDM

[ICIS23] J. Ishida-Castañeda, A. Iguchi, and K. Sakai. Changes in DNA methylation in response to heat stress and symbiotic conditions in coral primary polyps. Marine Ecology Progress Series, 714:113–117, July 13, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/714/m714p113.pdf.

Infantes:2022:SRS

[IHA^{+}22] E. Infantes, S. Hoeks, M. P. Adams, T. van der Heide, M. M. van Katwijk, and T. J. Bouma. Seagrass roots strongly reduce cliff erosion rates in sandy sediments. Marine Ecology Progress Series, 700:1–12, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ito:2021:LSC

[ILON21] M. A. Ito, H. J. Lin, M. I. O’Connor, and M. Nakaoka. Large-scale comparison of biomass and reproductive phenology among native and non-native populations of the seagrass Zostera japonica. Marine Ecology Progress Series, 675:1–21, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ichii:2020:BLD

[IMM^{+}20] T. Ichii, Y. Mori, K. Mahapatra, P. N. Trathan, M. Okazaki, T. Hayashi, and T. Okuda. Body length-dependent diel vertical migration of Antarctic krill in relation to food availability and predator avoidance in winter at South Georgia. Marine Ecology Progress Series, 654:53–66, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/654/
m654p053.pdf.

Ishihara:2024:MCP

[IMYW24] U. Ishihara, N. Miyazaki, D. J. Yurkowski, and Y. Y. Watanabe. Multi-cusped postcanine teeth are associated with zooplankton feeding in phocid seals. Marine Ecology Progress Series, 729:233–245, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/729/
m729p233.pdf.

Iwahara:2020:SNP

[ISMM20] Y. Iwahara, H. Shirakawa, K. Miyashita, and Y. Mitani. Spatial niche partitioning among three small cetaceans in the eastern coastal area of Hokkaido, Japan. Marine Ecology Progress Series, 637:209–223, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/637/m637p209.pdf.

Jones:2021:ACB

[JBDH21] R. N. Jones, E. G. Brush, E. R. Dilley, and M. A. Hixon. Autumn coral bleaching in Hawai’i. Marine Ecology Progress Series, 675:199–205, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2021/675/m675p199.pdf.

Jellison:2022:VBT

[JBG22] B. M. Jellison, S. C. Bacus, and B. Gaylord. Variation in the behavioral tolerance of congeneric marine snails to lowpH exposure. Marine Ecology Progress Series, 697:57–66, September 22, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/697/m697p057.pdf.

Jones:2023:USI

[JBGA23] J. B. Jones, P. Bustamante, G. Guillou, and A. I. Arkhipkin. Using stable isotope chronologies within squid gladii (Doryteuthis gahi) to evaluate dietary differences by fishing region and season. Marine Ecology Progress Series, 703: 95–108, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/703/m703p095.pdf.

Jarrett:2024:SHU

[JBWS24] R. N. Jarrett II, D. C. Brady, R. R. Wahle, and R. S. Steneck. Shifts in habitat use and demography of American lobsters in coastal Maine (USA) over the past quarter century. Marine Ecology Progress Series, 746:87–98, October 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/746/
m746p087.pdf.

Joly:2022:EEF

[JCGS22] N. B. Joly, A. Chiaradia, J. Y. Georges, and C. Saraux. Environmental effects on foraging performance in little penguins: a matter of phenology and short-term variability. Marine Ecology Progress Series, 692:151–168, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/692/
m692p151.pdf.

Jones:2021:IEE

[JDDF21] A. G. Jones, S. F. Dubois, N. Desroy, and J. Fournier. Intertidal ecosystem engineer species promote benthic-pelagic coupling and diversify trophic pathways. Marine Ecology Progress Series, 660:119–139, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2021/660/m660p119.pdf.

Jaonalison:2020:PSR

[JDM^{+}20] H. Jaonalison, J. D. Durand, J. Mahafina, H. Demarcq, N. Teichert, and D. Ponton. Predicting species richness and abundance of tropical post-larval fish using machine learning. Marine Ecology Progress Series, 645:125–139, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/645/
m645p125.pdf.

Johnson:2024:APC

[JETGM24] D. L. Johnson, J. M. Eisaguirre, R. L. Taylor, and J. L. Garlich-Miller. Assessing the population consequences of disturbance and climate change for the Pacific walrus. Marine Ecology Progress Series, 740:193–211, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/740/
m740p193.pdf.

Joyce:2023:RSI

[JFB^{+}23] T. W. Joyce, M. C. Ferguson, C. L. Berchok, D. L. Wright, J. L. Crance, E. K. Braen, T. Eguchi, W. L. Perryman, and D. W. Weller. The role of sea ice in the distribution, habitat use, and phenology of eastern North Pacific gray whales. Marine Ecology Progress Series, 709:141–158, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/709/
m709p141.pdf.

Johnston:2021:TSU

[JFR^{+}21] D. T. Johnston, R. W. Furness, A. M. C. Robbins, G. A. Tyler, J. McIlvenny, and E. A. Masden. Tidal stream use by black guillemots Cepphus grylle in relation to a marine renewable energy development. Marine Ecology Progress Series, 669:201–212, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/669/m669p201.pdf.

Jurriaans:2020:SAT

[JH20] S. Jurriaans and M. O. Hoogenboom. Seasonal acclimation of thermal performance in two species of reef-building corals. Marine Ecology Progress Series, 635:55–70, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m635p055.pdf.

Jenkinson:2020:BVD

[JHDE20] R. S. Jenkinson, K. A. Hovel, R. P. Dunn, and M. S. Edwards. Biogeographical variation in the distribution, abundance, and interactions among key species on rocky reefs of the northeast Pacific. Marine Ecology Progress Series, 648: 51–65, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/648/m648p051.pdf.

Jenkins:2022:SSC

[JHK^{+}22] G. P. Jenkins, P. A. Hamer, J. A. Kent, J. Kemp, C. D. H. Sherman, and A. J. Fowler. Spawning sources of a coastal fishery species inferred from otolith chemistry and microstructure: implications for management. Marine Ecology Progress Series, 684:103–116, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/684/m684p103.pdf.

Jackson:2021:DFA

[JJV^{+}21] G. D. Jackson, C. H. Jackson, P. Virtue, M. Fluckiger, and P. D. Nichols. Dietary fatty acid analyses of the squid Idioteuthis cordiformis: further evidence for predation on deepwater sharks. Marine Ecology Progress Series, 675:67–79, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/675/m675p067.pdf.

Jacobs:2020:PPP

[JKvLP20] P. Jacobs, J. C. Kromkamp, S. M. van Leeuwen, and C. J. M. Philippart. Planktonic primary production in the western Dutch Wadden Sea. Marine Ecology Progress Series, 639: 53–71, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/639/m639p053.pdf.

Jubinville:2021:DTS

[JLT^{+}21] I. Jubinville, E. Lawler, S. Tattrie, N. L. Shackell, J. Mills Flemming, and B. Worm. Distributions of threatened skates and commercial fisheries inform conservation hotspots. Marine Ecology Progress Series, 679:1–18, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Jacobson:2024:RPS

[JMM24] K. C. Jacobson, D. J. Marcogliese, and K. MacKenzie. Review: Parasites of small pelagics reflect their role in marine ecosystems. Marine Ecology Progress Series, 741:145– 167, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/741/m741p145.pdf.

Jeannot:2024:DCR

[JMP^{+}24] L. L. Jeannot, C. Mouronvalle, C. Peyran, C. Olabarria, M. Beger, A. Blanco, and S. Planes. Deriving connectivity from relatedness: broad-scale isolation-by-distance in the shanny Lipophrys pholis. Marine Ecology Progress Series, 731: 9–26, March 13, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m731p009.pdf.

Jarvis:2021:SSB

[JMR21] J. C. Jarvis, S. A. McKenna, and M. A. Rahseed. Seagrass seed bank spatial structure and function following a large-scale decline. Marine Ecology Progress Series, 665: 75–87, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/665/m665p075.pdf.

Jorissen:2022:HCI

[JMS^{+}22] H. Jorissen, A. Martin, L. Sarribouette, L. Hédouin, and M. M. Nugues. High CO_{2} inhibits substratum exploration and settlement of coral larvae. Marine Ecology Progress Series, 689:47–56, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/689/m689p047.pdf.

Jorgensen:2020:SPI

[JNJ^{+}20] K. E. M. Jørgensen, A. B. Neuheimer, P. E. Jorde, H. Knutsen, and P. Grønkjær. Settlement processes induce differences in daily growth rates between two co-existing ecotypes of juvenile cod Gadus morhua. Marine Ecology Progress Series, 650:175–189, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/650/m650p175.pdf.

Journey:2020:SIV

[JNY^{+}20] M. L. Journey, C. Neville, G. Young, M. Trudel, and B. R. Beckman. Spatial and interannual variability of juvenile Coho salmon growth in the Strait of Georgia (2012–2015). Marine Ecology Progress Series, 646:145–160, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/646/
m646p145.pdf.

Jiang:2020:VFC

[JP20] H. Jiang and G. A. Paffenhöfer. Vortical feeding currents in nauplii of the calanoid copepod Eucalanus pileatus. Marine Ecology Progress Series, 638:51–63, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/638/
m638p051.pdf.

Jones:2021:LTP

[JPB21] T. Jones, J. K. Parrish, and H. K. Burgess. Long-term patterns of mass stranding of the colonial cnidarian Velella velella: influence of environmental forcing. Marine Ecology Progress Series, 662:69–83, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m662p069.pdf.

Jiang:2022:TAP

[JPJ^{+}22] Z. Jiang, S. Ping, C. Jin, C. Tu, and X. Zhou. Transcriptome analysis provides insights into a molecular mechanism of histamine response in the cyprid larvae of Amphibalanus amphitrite. Marine Ecology Progress Series, 681:1–12, 2022.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

James:2023:MTA

[JPJH23] S. E. James, E. A. Pakhomov, B. T. Johnson, and B. P. V. Hunt. Migration timing affects the foraging ecology of Fraser River sockeye salmon stocks in coastal waters of British Columbia, Canada. Marine Ecology Progress Series, 719:93– 108, September 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/719/m719p093.pdf.

Jones:2024:MBM

[JPL^{+}24] T. Jones, J. K. Parrish, J. Lindsey, C. Wright, H. K. Burgess, J. Dolliver, L. Divine, R. Kaler, D. Bradley, G. Sorenson, R. Torrenta, S. Backensto, H. A. Coletti, J. T. Harvey, H. M. Nevins, E. Donnelly-Greenan, D. L. Sherer, J. Roletto, and K. Lindquist. Marine bird mass mortality events as an indicator of the impacts of ocean warming. Marine Ecology Progress Series, 737:161–181, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m737p161.pdf.

Jones:2023:HIP

[JPMG23] N. P. Jones, M. S. Pratchett, R. J. Milligan, and D. S. Gilliam. High incidence of partial colony mortality constrains realized growth for three coral species in southeast Florida. Marine Ecology Progress Series, 721:59–70, October 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/721/m721p059.pdf.

Jimenez-Ramos:2022:DER

[JRVE^{+}22] R. Jiménez-Ramos, B. Villazán, L. G. Egea, R. Cantero, J. L. Pérez-Lloréns, J. J. Vergara, and F. G. Brun. Differential ecophysiological responses to inorganic nitrogen sources (ammonium versus nitrate) and light levels in the seagrass Zostera noltei. Marine Ecology Progress Series, 702:57–70, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m702p057.pdf.

Jaco:2020:EIM

[JS20] E. M. Jaco and M. A. Steele. Early indicators of MPA effects are detected by stereo-video. Marine Ecology Progress Series, 647:161–177, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/647/m647p161.pdf.

Johnson:2020:CME

[JSDH20] D. S. Johnson, J. D. Shields, D. Doucette, and R. Heard. A climate migrant escapes its parasites. Marine Ecology Progress Series, 641:111–121, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/641/m641p111.pdf. See corri-

gendum [Ano21c].

Jusufovski:2019:CBM

[JSK19] D. Jusufovski, C. Saavedra, and A. Kuparinen. Competition between marine mammals and fisheries in contemporary harvested marine ecosystems. Marine Ecology Progress Series, 627:207–232, 2019. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/627/m627p207.pdf. See corrigen-

dum [Ano21d].

Johnston:2022:IAA

[JTBS^{+}22] D. T. Johnston, C. B. Thaxter, P. H. Boersch-Supan, E. M. Humphreys, W. Bouten, G. D. Clewley, E. S. Scragg, E. A. Masden, L. Barber, G. J. Conway, N. A. Clark, N. H. K. Burton, and A. S. C. P. Cook. Investigating avoidance and attraction responses in lesser black-backed gulls Larus fuscus to offshore wind farms. Marine Ecology Progress Series, 686:187–200, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/686/m686p187.pdf.

Johns:2022:NMM

[JW22] M. E. Johns and P. Warzybok. Northward migration, molting locations, and winter residency of California breeding pigeon guillemots Cepphus columba. Marine Ecology Progress Series, 701:133–143, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/701/m701p133.pdf.

Jacobson:2021:GVA

[JWH^{+}21] P. Jacobson, R. Whitlock, M. Huss, K. Leonardsson, J. Östergren, and A. Gårdmark. Growth variation of Atlantic salmon Salmo salar at sea affects their populationspecific reproductive potential. Marine Ecology Progress Series, 671:165–174, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m671p165.pdf.

Kahru:2022:BST

[KBE^{+}22] M. Kahru, H. Bittig, R. Elmgren, V. Fleming, Z. Lee, and G. Rehder. Baltic Sea transparency from ships and satellites: centennial trends. Marine Ecology Progress Series, 697:1–13, September 22, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kaldy:2022:CLR

[KBP22] J. E. Kaldy, C. A. Brown, and S. R. Pacella. Carbon limitation in response to nutrient loading in an eelgrass mesocosm: influence of water residence time. Marine Ecology Progress Series, 689:1–17, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kache:2021:IOP

[KBWRV21] S. Kache, I. Bartl, J. Wäge-Recchioni, and M. Voss. Influence of organic particle addition on nitrification rates and ammonium oxidiser abundances in Baltic seawater. Marine Ecology Progress Series, 674:59–72, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/674/m674p059.pdf.

Konecny:2024:IEI

[KCB^{+}24] C. A. Konecny, D. Cote, J. Broome, J. M. Nicolas, P. M. Regular, A. M. Cook, and F. Hatefi. Influences of environmental and individual-level covariates on movement behaviour in American lobster Homarus americanus. Marine Ecology Progress Series, 729:151–166, February 22, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m729p151.pdf.

Khamassi:2020:EPL

[KCGR20] S. Khamassi, L. Coussau, M. Guillemette, and D. Robert. Evidence of post-larval growth-selective mortality in Atlantic mackerel revealed by the collection of young-of-the-year juveniles ingested by the northern gannet. Marine Ecology Progress Series, 650:95–106, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/650/m650p095.pdf.

Kett:2020:SUR

[KCLJ20] G. F. Kett, S. C. Culloty, S. A. Lynch, and M. A. K. Jansen. Solar UV radiation modulates animal health and pathogen prevalence in coastal habitats — knowledge gaps and implications for bivalve aquaculture. Marine Ecology Progress Series, 653:217–231, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/653/m653p217.pdf.

Kaartvedt:2022:MFH

[KCRA22] S. Kaartvedt, S. Christiansen, A. Røstad, and D. L. Aksnes. Mesopelagic fishes in a hurry at low latitudes. Marine Ecology Progress Series, 694:149–156, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m694p149.pdf.

Kaartvedt:2024:IWF

[KCT24] S. Kaartvedt, S. Christiansen, and J. Titelman. Internal waves and fine-scale zooplankton sampling: Comment on Dewar-Fowler et al. (2023). Marine Ecology Progress Series, 734: 173–175, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2024/734/m734p173.pdf. See reply

[DFRS^{+}24].

Kaplan:2024:HKR

[KDB^{+}24] R. L. Kaplan, S. Derville, K. S. Bernard, E. M. Phillips, and L. G. Torres. Humpback-krill relationships are strongest at fine spatial scales in the Northern California Current region. Marine Ecology Progress Series, 729:219–232, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m729p219.pdf.

Korabik:2024:ITI

[KDGG24] A. R. Korabik, S. M. Dias, G. V. Grisby, and E. D. Grosholz. Increased temperatures impact the reproduction of localized estuarine kelp populations more than salinity or invasive species. Marine Ecology Progress Series, 744:33–51, September 5, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/744/m744p033.pdf.

Keramidas:2024:TEI

[KDK^{+}24] I. Keramidas, D. Dimarchopoulou, N. Kokkos, G. Sylaios, and A. C. Tsikliras. Temporal ecotrophic impacts of fisheries and climate change in the Aegean Sea. Marine Ecology Progress Series, 749:19–45, November 21, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/749/
m749p019.pdf.

Kaufman:2021:IGP

[KDLL21] M. L. Kaufman, M. D’Alessandro, C. Langdon, and D. Lirman. Influences of genotype, phenotypes, and size characteristics on lesion recovery in Caribbean staghorn coral. Marine Ecology Progress Series, 679:213–218, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m679p213.pdf.

Kershaw:2021:CSI

[KdlVJ^{+}21] J. L. Kershaw, C. de la Vega, R. M. Jeffreys, A. K. Frie, T. Haug, C. Mahaffey, C. Mettam, G. Stenson, and S. Smout. Compound-specific isotope analyses of harp seal teeth: tools for trophic ecology reconstruction. Marine Ecology Progress Series, 678:211–225, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2021/678/m678p211.pdf.

Kuhn:2020:TUS

[KDS^{+}20] C. E. Kuhn, A. De Robertis, J. Sterling, C. W. Mordy, C. Meinig, N. Lawrence-Slavas, E. Cokelet, M. Levine, H. Tabisola, R. Jenkins, D. Peacock, and D. Vo. Test of unmanned surface vehicles to conduct remote focal follow studies of a marine predator. Marine Ecology Progress Series, 635:1–7, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Kramer:2022:GSD

[KETL22] N. Kramer, G. Eyal, R. Tamir, and Y. Loya. Growth and survival dynamics of mesophotic coral juveniles in shallow reefs. Marine Ecology Progress Series, 682:237–242, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/682/
m682p237.pdf.

Krumsick:2020:CSS

[KF20] K. J. Krumsick and J. A. D. Fisher. Community size spectra provide indicators of ecosystem recovery on the Newfoundland and Labrador shelf. Marine Ecology Progress Series, 635:123–137, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/635/m635p123.pdf.

Kunisch:2021:IAC

[KGG^{+}21] E. H. Kunisch, M. Graeve, R. Gradinger, T. Haug, K. M. Kovacs, C. Lydersen, Ø. Varpe, and B. A. Bluhm. Ice-algal carbon supports harp and ringed seal diets in the European Arctic: evidence from fatty acid and stable isotope markers. Marine Ecology Progress Series, 675:181–197, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m675p181.pdf.

Kuletz:2024:SRE

[KGM^{+}24] K. J. Kuletz, A. E. Gall, T. C. Morgan, A. K. Prichard, L. B. Eisner, D. G. Kimmel, A. De Robertis, R. M. Levine, T. Jones, and E. A. Labunski. Seabird responses to ecosystem changes driven by marine heatwaves in a warming Arctic. Marine Ecology Progress Series, 737:59–88, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m737p059.pdf.

Kvernvik:2021:ASI

[KHG^{+}21] A. C. Kvernvik, C. J. M. Hoppe, M. Greenacre, S. Verbiest, J. M. Wiktor, T. M. Gabrielsen, M. Reigstad, and E. Leu. Arctic sea ice algae differ markedly from phytoplankton in their ecophysiological characteristics. Marine Ecology Progress Series, 666:31–55, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/666/m666p031.pdf.

Kastelle:2022:OIF

[KHL^{+}22] C. R. Kastelle, T. E. Helser, B. J. Laurel, L. A. Copeman, K. R. Stone, and J. L. McKay. Oxygen isotope fractionation in otoliths: experimental results from four North Pacific and Arctic gadid species. Marine Ecology Progress Series, 686:159–175, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/686/m686p159.pdf.

Kodama:2022:SDG

[KHOK22] M. Kodama, J. Hayakawa, S. Oba, and T. Kawamura. Seasonal dispersal of gammaridean amphipods away from Sargassum beds in relation to macroalgal host defoliation. Marine Ecology Progress Series, 681:117–128, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/681/
m681p117.pdf.

Kanno:2023:MUS

[KHSS23] S. REVIEW Kanno, M. R. Heupel, M. J. Sheaves, and C. A. Simpfendorfer. Mangrove use by sharks and rays: a review. Marine Ecology Progress Series, 724:167–183, December 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/724/m724p167.pdf.

Kurihara:2024:GPS

[KI24] H. Kurihara and M. Ikeda. Genetic population structure and demographic history of the sailfin sandfish Arctoscopus japonicus associated with sea level changes during the Last Glacial Maximum. Marine Ecology Progress Series, 747:133– 149, October 17, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m747p133.pdf.

Kilfoil:2024:DLR

[KKC^{+}24] J. P. Kilfoil, G. Krohn, E. E. G. Clua, S. Planes, K. R. Gastrich, M. R. Heithaus, and A. J. Wirsing. Divergent learning responses to a spatially consistent olfactory stimulus in two reef shark species. Marine Ecology Progress Series, 738: 151–160, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m738p151.pdf.

Kingsford:2022:TCS

[KKH22] M. J. Kingsford, E. A. Krunes, and A. E. Hall. Testing the critical size at settlement hypothesis for two species of coral reef fish. Marine Ecology Progress Series, 681:87– 101, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/681/m681p087.pdf.

Kumagai:2023:BTG

[KKMW23] A. Kumagai, K. Kazama, K. Mikami, and Y. Watanuki. Black-tailed gulls alter their flight height and airspeed according to wind conditions during their coastal commuting trips. Marine Ecology Progress Series, 723:201–212, November 23, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/723/m723p201.pdf.

Kodama:2020:EAP

[KKN^{+}20] M. Kodama, T. Kawamura, K. Nakamoto, N. Ohtsuchi, J. Hayakawa, T. Kanki, T. Kitagawa, and Y. Watanabe. Effect of algal phenology on seasonal dynamics of gammarid assemblages: differences between canopy and understory strata in a Sargassum yezoense bed. Marine Ecology Progress Series, 634:63–76, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/634/m634p063.pdf.

Kinney:2020:CTS

[KKS^{+}20] M. J. Kinney, D. Kacev, T. Sippel, H. Dewar, and T. Eguchi. Common thresher shark Alopias vulpinus movement: Bayesian inference on a data-limited species. Marine Ecology Progress Series, 639:155–167, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/639/
m639p155.pdf.

Kim:2021:SEM

[KLA21] J. J. H. Kim, L. Liggins, and J. D. Aguirre. Social environment mediates habitat shifts in a range-restricted giant limpet. Marine Ecology Progress Series, 666:89–98, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/666/
m666p089.pdf.

Kulp:2023:PDS

[KLFP23] R. E. Kulp, A. V. Lowell, N. J. Floros, and B. J. Peterson. Predator density, not structure, influences intraspecific competition in the mud crab Dyspanopeus sayi. Marine Ecology Progress Series, 709:45–54, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/709/m709p045.pdf.

Kluibenschedl:2021:LIA

[KLN^{+}21] A. Kluibenschedl, M. D. Lamare, W. A. Nelson, N. G. Barr, and C. D. Hepburn. Low irradiance amplifies negative effects of ocean acidification on recruitment of coralline algae communities. Marine Ecology Progress Series, 674:103– 113, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/674/m674p103.pdf.

King:2021:MSV

[KMW^{+}21] N. G. King, P. J. Moore, C. Wilding, H. L. Jenkins, and D. A. Smale. Multiscale spatial variability in epibiont assemblage structure associated with stipes of kelp Laminaria hyperborea in the northeast Atlantic. Marine Ecology Progress Series, 672:33–44, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/672/m672p033.pdf.

Kozal:2024:KAV

[KNH24] L. C. Kozal, J. C. Nelson, and G. E. Hofmann. Kelp-associated variability in seawater chemistry during a marine heatwave event connects to transgenerational effects in the purple urchin Strongylocentrotus purpuratus. Marine Ecology Progress Series, 733:59–77, April 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2024/733/m733p059.pdf.

Kakehi:2021:BTW

[KNO^{+}21] S. Kakehi, Y. Narimatsu, Y. Okamura, A. Yagura, and S. Ito. Bottom temperature warming and its impact on demersal fish off the Pacific coast of northeastern Japan. Marine Ecology Progress Series, 677:177–196, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/677/m677p177.pdf.

Kellogg:2021:UDE

[KP21] C. A. Kellogg and Z. A. Pratte. Unexpected diversity of Endozoicomonas in deep-sea corals. Marine Ecology Progress Series, 673:1–15, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Klein:2024:MEE

[KPK^{+}24] J. C. Klein, E. N. Powell, D. A. Kreeger, X. Zhang, S. M. Pace, K. M. Kuykendall, and R. L. Thomas. Model estimation of eastern oyster larval performance from food quantity and quality measures in western Mississippi Sound. Marine Ecology Progress Series, 745:73–94, September 19, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/745/
m745p073.pdf.

Koehl:2022:ECW

[KPS^{+}22] M. A. R. Koehl, E. Perotti, D. Sischo, T. Hata, and M. G. Hadfield. Effects of currents, waves, and biofilms on motion and surface contacts by tubeworm larvae swimming above or below surfaces. Marine Ecology Progress Series, 686:107– 126, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/686/m686p107.pdf.

Keys:2023:BWF

[KPTT23] D. Z. Keys, P. A. Pistorius, Y. Tremblay, and A. Thiebault. Both wind and flying with conspecifics influence the flight dynamics of a seabird. Marine Ecology Progress Series, 723: 135–149, November 23, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2023/723/m723p135.pdf.

Kane:2020:STP

[KPW^{+}20] A. Kane, E. Pirotta, S. Wischnewski, E. J. Critchley, A. Bennison, M. Jessopp, and J. L. Quinn. Spatio-temporal patterns of foraging behaviour in a wide-ranging seabird reveal the role of primary productivity in locating prey. Marine Ecology Progress Series, 646:175–188, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/646/m646p175.pdf.

Klemmedson:2020:VAS

[KRG^{+}20] A. D. Klemmedson, C. S. Reiss, M. E. Goebel, R. S. Kaufmann, E. Dorval, T. B. Linkowski, and R. Borras-Chavez. Variability in age of a Southern Ocean myctophid (Gymnoscopelus nicholsi) derived from scat-recovered otoliths. Marine Ecology Progress Series, 633:55–69, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/633/
m633p055.pdf.

Kowaleski:2024:EPD

[KRMR24] K. R. Kowaleski, S. A. Roman, R. Mann, and D. B. Rudders. Extreme population densities reduce reproductive effort of Atlantic sea scallops in high-density recruitment events. Marine Ecology Progress Series, 746:67–85, October 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m746p067.pdf.

Krumme:2020:AVA

[KSMP20] U. Krumme, S. Stötera, K. McQueen, and E. Pahlke. Age validation of age 0–3 wild cod Gadus morhua in the western Baltic Sea through mark-recapture and tetracycline marking of otoliths. Marine Ecology Progress Series, 645:141– 158, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m645p141.pdf.

Kwong:2020:CMP

[KSS^{+}20] L. E. Kwong, K. D. Suchy, A. R. Sastri, J. F. Dower, and E. A. Pakhomov. Comparison of mesozooplankton production estimates from Saanich Inlet (British Columbia, Canada) using the chitobiase and biomass size spectra approaches.

Marine Ecology Progress Series, 655:59–75, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/655/
m655p059.pdf.

King:2021:CIP

[KSW^{+}21] O. C. King, R. A. Smith, M. S. J. Warne, J. P. van de Merwe, R. M. Connolly, and C. J. Brown. Combined impacts of photosystem II-inhibiting herbicides and light availability on seagrass and marine microalgae. Marine Ecology Progress Series, 668:215–230, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/668/m668p215.pdf.

Kolodzey:2023:SSD

[KSW23] S. Kolodzey, A. K. Stroh, and S. R. Wing. Small-scale differences in blue cod length distribution, growth, and trophic ecology in New Zealand. Marine Ecology Progress Series, 708:125–142, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m708p125.pdf.

Kuhn:2021:UML

[KTK21] B. Kühn, M. H. Taylor, and A. Kempf. Using machine learning to link spatiotemporal information to biological processes in the ocean: a case study for North Sea cod recruitment. Marine Ecology Progress Series, 664:1–22, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Klappstein:2020:PSI

[KTR^{+}20] N. J. Klappstein, R. R. Togunov, J. R. Reimer, N. J. Lunn, and A. E. Derocher. Patterns of sea ice drift and polar bear (Ursus maritimus) movement in Hudson Bay. Marine Ecology Progress Series, 641:227–240, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/641/m641p227.pdf.

Kucheravy:2022:BCF

[KTW^{+}22] C. E. Kucheravy, M. R. Trana, C. A. Watt, J. D. Roth, G. T. Tomy, W. G. Anderson, and S. H. Ferguson. Blubber cortisol in four Canadian beluga whale populations is unrelated to diet. Marine Ecology Progress Series, 698:171–189, 2022.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m698p171.pdf.

Kristensen:2022:PCB

[KVdM^{+}22] E. Kristensen, T. Valdemarsen, P. C. de Moraes, A. Z. Güth, P. Y. G. Sumida, and C. O. Quintana. Pneumatophores and crab burrows increase CO_{2} and CH_{4} emission from sediments in two Brazilian fringe mangrove forests. Marine Ecology Progress Series, 698:29–39, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m698p029.pdf.

Kolodzey:2021:MAL

[KWD21] S. Kolodzey, S. R. Wing, and L. M. Durante. Maternal age, length, and weight influence larval characteristics and viability in the viviparous temperate reef fish Helicolenus percoides. Marine Ecology Progress Series, 659:185–198, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m659p185.pdf.

Kolbasov:2023:TEC

[KZC23] G. A. Kolbasov, A. K. Zalota, and B. K. K. Chan. Trophic ecology of crustacean endoparasites Petrarca and Baccalaureus (Ascothoracida, Thecostraca) in scleractinian and zoantharian corals. Marine Ecology Progress Series, 715:69–78, July 27, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/715/m715p069.pdf.

Litz:2021:COY

[LAD^{+}21] M. N. C. Litz, M. Agha, A. M. Dufault, A. M. Claiborne, J. P. Losee, and A. J. Anderson. Competition with oddyear pink salmon in the ocean affects natural populations of chum salmon from Washington. Marine Ecology Progress Series, 663:179–195, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/663/m663p179.pdf.

Lozano:2021:SPC

[LAGGM^{+}21] J. Lozano, M. Aranguren-Gassís, E. E. García-Martín, J. González, J. L. Herrera, B. Hidalgo-Robatto, D. Mártinez-Castrillón, M. Pérez-Lorenzo, R. A. Varela, and P. Serret. Seasonality of phytoplankton cell size and the relation between photosynthesis and respiration in the Ría de Vigo (NW Spain). Marine Ecology Progress Series, 664: 43–58, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/664/m664p043.pdf.

Lima:2020:GCC

[LÁGLL20] F. D. Lima, L. E. Ángeles-González, T. S. Leite, and S. M. Q. Lima. Global climate changes over time shape the environmental niche distribution of Octopus insularis in the Atlantic Ocean. Marine Ecology Progress Series, 652:111–121, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/652/m652p111.pdf. See [JSDH20].

Lemley:2021:HAB

[LAL21] D. A. Lemley, J. B. Adams, and J. L. Largier. Harmful algal blooms as a sink for inorganic nutrients in a eutrophic estuary. Marine Ecology Progress Series, 663:63–76, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/663/
m663p063.pdf.

Lacombe:2024:LTS

[LAP^{+}24] R. M. Lacombe, T. C. Atwood, E. Peacock, A. Remili, R. Dietz, C. Sonne, and M. A. McKinney. Long-term storage at −20 ^{o} C compromises fatty acid composition of polar bear adipose biopsies. Marine Ecology Progress Series, 728:75– 80, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/728/m728p075.pdf.

Liang:2021:TWR

[LBB^{+}21] Y. Liang, L. Bretherton, C. M. Brown, U. Passow, A. Quigg, A. J. Irwin, and Z. V. Finkel. Transcriptome-wide responses of aggregates of the diatom Odontella aurita to oil. Marine Ecology Progress Series, 671:67–79, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/671/
m671p067.pdf.

Lam:2021:GVL

[LBH^{+}21a] L. S. Lam, B. L. Basnett, M. A. Haltuch, J. Cope, K. Andrews, K. M. Nichols, G. C. Longo, J. F. Samhouri, and S. L. Hamilton. Geographic variability in lingcod Ophiodon elongatus life history and demography along the US West Coast: oceanographic drivers and management implications. Marine Ecology Progress Series, 670:203–222, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/670/m670p203.pdf.

Leal:2021:LAL

[LBH^{+}21b] I. Leal, K. Bohn, S. J. Hawkins, S. R. Jenkins, A. A. V. Flores, and R. Tremblay. Lipid allocation in late-stage barnacle larvae from subtropical and temperate waters. Marine Ecology Progress Series, 661:147–161, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/661/m661p147.pdf.

Lee:2021:OCA

[LBH^{+}21c] K. A. Lee, P. A. Butcher, R. G. Harcourt, T. A. Patterson, V. M. Peddemors, M. Roughan, D. Harasti, A. F. Smoothey, and R. W. Bradford. Oceanographic conditions associated with white shark (Carcharodon carcharias) habitat use along eastern Australia. Marine Ecology Progress Series, 659:143–159, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m659p143.pdf.

Lowerre-Barbieri:2020:TAA

[LBMB^{+}20] S. Lowerre-Barbieri, H. Menendez, J. Bickford, T. S. Switzer, L. Barbieri, and C. Koenig. Testing assumptions about sex change and spatial management in the protogynous gag grouper, Mycteroperca microlepis. Marine Ecology Progress Series, 639:199–214, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/639/m639p199.pdf.

Lofthus:2024:FIN

[LBMF^{+}24] A. J. Lofthus, M. Beach-Mehrotra, M. R. Fisher, J. R. Wozniak, and P. Matich. Factors influencing the nursery dynamics of bull sharks Carcharhinus leucas in two estuaries along the Texas coast. Marine Ecology Progress Series, 728:43– 58, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/728/m728p043.pdf.

Llabres:2023:EGW

[LBMSS23] E. Llabrés, A. Blanco-Magadán, M. Sales, and T. Sintes. Effect of global warming on Western Mediterranean seagrasses: a preliminary agent-based modelling approach. Marine Ecology Progress Series, 710:43–56, May 4, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/710/
m710p043.pdf.

Lin:2024:PPE

[LBR^{+}24] C. Lin, B. A. Belgrad, C. M. Russell, J. Lunt, and D. L. Smee. Phenotypic plasticity expands oyster survival and realized niche space across tidal elevations. Marine Ecology Progress Series, 734:35–43, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/734/m734p035.pdf.

Leandri-Breton:2021:LTT

[LBTE^{+}21] D. J. Léandri-Breton, A. Tarroux, K. H. Elliott, P. Legagneux, F. Angelier, P. Blévin, V. S. Bråthen, P. Fauchald, A. Goutte, W. Jouanneau, S. Tartu, B. Moe, and O. Chastel. Longterm tracking of an Arctic-breeding seabird indicates high fidelity to pelagic wintering areas. Marine Ecology Progress Series, 676:205–218, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m676p205.pdf.

Langton:2021:VDM

[LBW21] R. Langton, P. Boulcott, and P. J. Wright. A verified distribution model for the lesser sandeel Ammodytes marinus. Marine Ecology Progress Series, 667:145–159, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/667/
m667p145.pdf.

Ljungstrom:2020:BSA

[LCFJ20] G. Ljungström, M. Claireaux, Ø. Fiksen, and C. Jørgensen. Body size adaptions under climate change: zooplankton community more important than temperature or food abundance in model of a zooplanktivorous fish. Marine Ecology Progress Series, 636:1–18, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lunzmann-Cooke:2021:STE

[LCGS^{+}21] E. L. Lunzmann-Cooke, R. S. Gregory, P. V. R. Snelgrove, D. Cote, and C. Fuentes-Yaco. Spatial, temporal, and environmental influences on Atlantic cod Gadus morhua offshore recruitment signals in Newfoundland. Marine Ecology Progress Series, 673:151–164, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m673p151.pdf.

Li:2022:CIS

[LCJ22] J. Li, J. Chapman, and P. Johnson. Co-infesting symbionts on a threatened marine host: evaluating correlations between an introduced parasitic isopod and a native symbiotic clam. Marine Ecology Progress Series, 695:83–93, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/695/
m695p083.pdf.

Labarre:2020:MHA

[LCP^{+}20] D. Labarre, P. Charruau, W. F. J. Parsons, S. Larocque-Desroches, and J. A. Gallardo-Cruz. Major hurricanes affect body condition of American crocodile Crocodylus acutus inhabiting Mexican Caribbean islands. Marine Ecology Progress Series, 651:145–162, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/651/m651p145.pdf.

Losee:2023:MBM

[LCP^{+}23] J. Losee, A. Claiborne, D. Palm, A. Holmgren, A. Anderson, T. Brodin, and G. Hellström. Migratory behavior and maternal origin of anadromous brown trout in the Baltic Sea. Marine Ecology Progress Series, 722:145–155, November 9, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/722/m722p145.pdf.

Lippi:2022:UAT

[LCR^{+}22] D. L. Lippi, M. S. Coxey, J. R. Rooker, S. M. Rezende, M. A. Dance, A. L. B. Gaspar, M. Maida, and B. P. Ferreira. Use of acoustic telemetry to evaluate fish movement, habitat use, and protection effectiveness of a coral reef notake zone (NTZ) in Brazil. Marine Ecology Progress Series, 688:113–131, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/688/m688p113.pdf.

Layton:2020:HFC

[LCT^{+}20] C. Layton, M. J. Cameron, M. Tatsumi, V. Shelamoff, J. T. Wright, and C. R. Johnson. Habitat fragmentation causes collapse of kelp recruitment. Marine Ecology Progress Series, 648:111–123, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/648/m648p111.pdf.

Layko:2024:AED

[LD24] R. B. Layko and M. K. Donovan. Anthropogenic and environmental drivers of Acanthurus achilles presence in Hawai’i. Marine Ecology Progress Series, 740:161–174, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m740p161.pdf.

Labry:2020:EPD

[LDM^{+}20] C. Labry, D. Delmas, B. Moriceau, M. Gallinari, J. Quere, and A. Youenou. Effect of P depletion on the functional pools of diatom carbohydrates, and their utilization by bacterial communities. Marine Ecology Progress Series, 641: 49–62, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/641/m641p049.pdf.

Le:2024:TTF

[LDPH24] D. M. Le, M. J. Desmond, D. W. Pritchard, and C. D. Hepburn. Thermal threshold for fertilisation and gametophyte survivorship of the giant kelp Macrocystis pyrifera. Marine Ecology Progress Series, 734:23–33, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/734/
m734p023.pdf.

Lewis:2021:OBA

[LDW^{+}21] L. S. Lewis, C. Denney, M. Willmes, W. Xieu, R. A. Fichman, F. Zhao, B. G. Hammock, A. Schultz, N. Fangue, and J. A. Hobbs. Otolith-based approaches indicate strong effects of environmental variation on growth of a critically endangered estuarine fish. Marine Ecology Progress Series, 676:37–56, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m676p037.pdf.

Lomas:2020:TSD

[LEG^{+}20] M. W. Lomas, L. B. Eisner, J. Gann, S. E. Baer, C. W. Mordy, and P. J. Stabeno. Time-series of direct primary production and phytoplankton biomass in the southeastern Bering Sea: responses to cold and warm stanzas. Marine Ecology Progress Series, 642:39–54, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/642/m642p039.pdf.

Leis:2020:MSA

[Lei20] J. M. Leis. Measurement of swimming ability in larval marine fishes: comparison of critical speed with in situ speed. Marine Ecology Progress Series, 650:203–215, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/650/
m650p203.pdf.

Lynn:2023:CID

[LFA^{+}23] K. D. Lynn, P. Tummon Flynn, C. Anguita, C. Duarte, and P. A. Quijón. Conspecific interactions as drivers of activity and mortality in a prominent sandy beach amphipod. Marine Ecology Progress Series, 713:83–95, June 29, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/713/
m713p083.pdf.

Livernois:2021:SPH

[LFFW21] M. C. Livernois, M. Fujiwara, M. Fisher, and R. J. D. Wells. Seasonal patterns of habitat suitability and spatiotemporal overlap within an assemblage of estuarine predators and prey. Marine Ecology Progress Series, 668:39–55, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m668p039.pdf.

Leitner:2021:BIL

[LFK^{+}21] A. B. Leitner, T. Friedrich, C. D. Kelley, S. Travis, D. Partridge, B. Powell, and J. C. Drazen. Biogeophysical influence of large-scale bathymetric habitat types on mesophotic and upper bathyal demersal fish assemblages: a Hawaiian case study. Marine Ecology Progress Series, 659:219–236, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m659p219.pdf.

Lescure:2023:IFE

[LGD23] L. Lescure, J. Gulka, and G. K. Davoren. Increased foraging effort and reduced chick condition of razorbills under lower prey biomass in coastal Newfoundland, Canada. Marine Ecology Progress Series, 709:109–123, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/709/
m709p109.pdf.

Lopez:2022:EGD

[LGGR^{+}22] K. E. Rojas López, J. J. Guadalupe, M. Gordillo-Romero, A. Montero-Oleas, D. A. Pazmiño, M. Guerrero, and M. L. Torres. Exploring the genetic diversity and population structure of Mobula birostris in two key aggregation zones in the Eastern Tropical Pacific. Marine Ecology Progress Series, 699:75–89, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m699p075.pdf.

Lenihan:2021:MCH

[LGM21] Hunter S. Lenihan, Jonathan Grabowski, and Fiorenza Micheli. In memoriam: Charles Henry Peterson (1946– 2020). Marine Ecology Progress Series, 659:1, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m659p001.pdf.

Laute:2022:IWW

[LGR^{+}22] A. Laute, T. J. Grove, M. H. Rasmussen, A. Smith, O. Loisa, and M. E. H. Fournet. Impact of whale-watching vessels on humpback whale calling behavior on an Icelandic foraging ground during the Covid-19 pandemic. Marine Ecology Progress Series, 701:159–173, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/701/m701p159.pdf.

Lerner:2022:EVC

[LH22] J. E. NOTE Lerner and B. P. V. Hunt. Experimental validation confirms a carbon stable isotope lipid normalization procedure for Pacific salmon. Marine Ecology Progress Series, 698:191–197, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/698/m698p191.pdf.

Landry:2020:MBT

[LHD20] M. R. Landry, R. R. Hood, and C. H. Davies. Mesozooplankton biomass and temperature-enhanced grazing along a 110^{o}E transect in the eastern Indian Ocean. Marine Ecology Progress Series, 649:1–19, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lavoie:2024:LUA

[LHFD^{+}24] C. Lavoie, K. L. Howland, K. Filbee-Dexter, P. Massicotte, J. Goldsmit, C. W. McKindsey, and P. Archambault. Living under Arctic kelp forests: linking soft-bottom communities to kelp cover in the Canadian Arctic. Marine Ecology Progress Series, 740:1–22, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lund-Hansen:2020:EII

[LHHH^{+}20] L. C. Lund-Hansen, I. Hawes, K. Hancke, N. Salmansen, J. R. Nielsen, L. Balslev, and B. K. Sorrell. Effects of increased irradiance on biomass, photobiology, nutritional quality, and pigment composition of Arctic sea ice algae. Marine Ecology Progress Series, 648:95–110, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/648/m648p095.pdf.

Lindsay:2023:CRS

[LHM^{+}23] J. M. Lindsay, D. D. W. Hauser, A. R. Mahoney, K. L. Laidre, J. Goodwin, C. Harris, R. J. Schaeffer, S. r. Schaeffer R, A. V. Whiting, P. L. Boveng, N. J. M. Laxague, S. Betcher, A. Subramaniam, C. R. Witte, and C. J. Zappa. Characteristics of ringed seal Pusa hispida (‘natchiq’) denning habitat in Kotzebue Sound, Alaska, during a year of limited sea ice and snow. Marine Ecology Progress Series, 705:1–20, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lyndby:2020:ETF

[LHW^{+}20] N. H. Lyndby, J. B. Holm, D. Wangpraseurt, R. Grover, C. Rottier, M. Kühl, and C. Ferrier-Pagès. Effect of temperature and feeding on carbon budgets and O_{2} dynamics in Pocillopora damicornis. Marine Ecology Progress Series, 652:49–62, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m652p049.pdf.

Lamont:2020:GRI

[LJ20] M. M. Lamont and D. Johnson. Growth rates for immature Kemp’s ridley sea turtles from a foraging area in the northern Gulf of Mexico. Marine Ecology Progress Series, 652:145–155, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/652/m652p145.pdf.

Lemasson:2021:DRA

[LK21] A. J. Lemasson and A. M. Knights. Differential responses in anti-predation traits of the native oyster Ostrea edulis and invasive Magallana gigas to ocean acidification and warming. Marine Ecology Progress Series, 665:87–102, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/665/
m665p087.pdf.

LeBourg:2021:IEB

[LKB^{+}21] B. Le Bourg, P. Kuklinski, P. Balazy, G. Lepoint, and L. N. Michel. Interactive effects of body size and environmental gradient on the trophic ecology of sea stars in an Antarctic fjord. Marine Ecology Progress Series, 674:189–202, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/674/
m674p189.pdf.

LaRoche:2021:BOS

[LKR^{+}21] N. L. LaRoche, S. L. King, M. C. Rogers, G. L. Eckert, and H. C. Pearson. Behavioral observations and stable isotopes reveal high individual variation and little seasonal variation in sea otter diets in Southeast Alaska. Marine Ecology Progress Series, 677:219–232, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m677p219.pdf.

Lovy:2020:HSH

[LLF^{+}20] J. Lovy, N. L. Lewis, S. E. Friend, K. W. Able, M. J. Shaw, G. S. Hinks, and P. J. Clarke. Host, seasonal and habitat influences on incidence of Lernaeenicus radiatus (Copepoda: Pennellidae) in the mid-Atlantic Bight. Marine Ecology Progress Series, 642:83–101, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/642/m642p083.pdf.

Lescroel:2020:IID

[LLJ^{+}20] A. Lescroël, P. O.’B. Lyver, D. Jongsomjit, S. Veloz, K. M. Dugger, P. Kappes, B. J. Karl, A. L. Whitehead, R. Pech, T. L. Cole, and G. Ballard. Inter-individual differences in the foraging behavior of breeding Adélie penguins are driven by individual quality and sex. Marine Ecology Progress Series, 636:189–205, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/636/m636p189.pdf.

Lu:2020:TAI

[LLL20] Y. Lu, C. H. Lucas, and A. Loveridge. Transgenerational acclimation influences asexual reproduction in Aurelia aurita jellyfish polyps in response to temperature. Marine Ecology Progress Series, 656:35–50, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/656/m656p035.pdf.

Lester:2020:MYP

[LMB^{+}20] E. Lester, M. G. Meekan, P. Barnes, H. Raudino, D. Rob, K. Waples, and C. W. Speed. Multi-year patterns in scarring, survival and residency of whale sharks in Ningaloo Marine Park, Western Australia. Marine Ecology Progress Series, 634:115–125, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/634/m634p115.pdf.

Lett:2024:RCB

[LMH^{+}24] C. Lett, B. S. Malauene, T. B. Hoareau, D. M. Kaplan, and F. Porri. REVIEW: Corridors and barriers to marine connectivity around southern Africa. Marine Ecology Progress Series, 731:105–127, March 13, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m731p105.pdf.

Lesser:2021:ILM

[LMHS21] M. P. Lesser, C. D. Mobley, J. D. Hedley, and M. Slattery. Incident light on mesophotic corals is constrained by reef topography and colony morphology. Marine Ecology Progress Series, 670:49–60, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/670/m670p049.pdf.

Land-Miller:2024:CFN

[LMRS^{+}24] H. Land-Miller, A. M. Roos, M. Simon, R. Dietz, C. Sonne, S. Pedro, A. Rosing-Asvid, F. F. Rigét, and M. A. McKinney. Comparison of feeding niches between Arctic and northward-moving sub-Arctic marine mammals in Greenland. Marine Ecology Progress Series, 728:163–182, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/728/m728p163.pdf.

Lees:2020:SOP

[LMS^{+}20] K. J. Lees, A. C. Mill, D. J. Skerritt, P. A. Robertson, and C. Fitzsimmons. Spatial overlap, proximity, and interaction between lobsters revealed using acoustic telemetry. Marine Ecology Progress Series, 645:109–124, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/645/
m645p109.pdf.

Lubcker:2021:VUB

[LNBdB21] N. Lübcker, S. D. Newsome, M. N. Bester, and P. J. N. de Bruyn. Validating the use of bulk tissue stable isotope and amino acid δ ^{15}N values measured in molted hair and epidermis of elephant seals to assess temporal foraging niche specialization. Marine Ecology Progress Series, 673: 229–243, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/673/m673p229.pdf.

Lundgreen:2022:EFM

[LNKJ^{+}22] R. B. C. Lundgreen, A. Nielsen, M. Krüger-Johnsen, D. Righton, M. Mion, K. Radtke, M. Plikshs, A. J. Leskelä, J. Raitaniemi, C. A. Griffiths, M. Casini, U. Krumme, and K. Hüssy. Examining fish movement in terms of advection and diffusion: a case study of northeastern Atlantic cod. Marine Ecology Progress Series, 691:115–129, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/691/
m691p115.pdf.

Lowry:2022:STD

[LNM22] M. S. Lowry, S. E. Nehasil, and J. E. Moore. Spatio-temporal diet variability of the California sea lion Zalophus californianus in the southern California Current Ecosystem. Marine Ecology Progress Series, 692:1–21, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Liu:2021:ETM

[LNZL21] K. Liu, H. Y. Ng, S. Zhang, and H. Liu. Effects of temperature on a mixotrophic dinoflagellate (Lepidodinium sp.) under different nutritional strategies. Marine Ecology Progress Series, 678:37–49, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/678/m678p037.pdf.

Lee:2020:FSS

[LO20] E. M. J. Lee and K. G. O’Malley. Fine-scale spatial and temporal genomic variation among Dungeness crab Cancer magister larval recruits in the California Current Ecosystem. Marine Ecology Progress Series, 649:67–81, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m649p067.pdf.

Lefcheck:2024:TMW

[LOF^{+}24] J. S. Lefcheck, R. J. Orth, F. J. Fodrie, D. Gong, M. Faulkner, K. M. Laumann, and C. J. Patrick. Tropicalization of mid-western Atlantic coastal bays by pinfish Lagodon rhomboides: a combined ecological and oceanographic perspective. Marine Ecology Progress Series, 748:137–148, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/748/m748p137.pdf.

Lujan:2024:KSI

[LORB^{+}24] C. Luján, R. Oliveros-Ramos, N. Barrier, P. Leadley, and Y. J. Shin. Key species and indicators revealed by an uncertainty analysis of the marine ecosystem model OSMOSE. Marine Ecology Progress Series, 741:29–46, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/741/
m741p029.pdf.

Lange:2022:LSE

[LOS^{+}22] T. Lange, N. S. Oncken, N. Svane, R. C. Steinfurth, E. Kristensen, and M. R. Flindt. Large-scale eelgrass transplantation: a measure for carbon and nutrient sequestration in estuaries. Marine Ecology Progress Series, 685:97–109, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m685p097.pdf.

Liboureau:2023:ETH

[LPB^{+}23] P. Liboureau, G. A. Pearson, L. Barreto, E. A. Serrao, A. Kreiner, and N. Martins. Effects of thermal history on reproductive success and cross-generational effects in the kelp Laminaria pallida (Phaeophyceae). Marine Ecology Progress Series, 715:41–56, July 27, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/715/m715p041.pdf.

Loesser:2024:SIR

[LPD^{+}24] K. B. Loesser, C. E. Powell, B. Davis, M. M. Baustian, and M. J. Polito. Stable isotopes reveal that foraging strategy dictates trophic response of salt marsh residents to black mangrove Avicennia germinans range expansion. Marine Ecology Progress Series, 729:81–97, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/729/
m729p081.pdf.

Ladds:2020:RBM

[LPJ^{+}20] M. A. Ladds, M. H. Pinkerton, E. Jones, L. M. Durante, and M. R. Dunn. Relationship between morphometrics and trophic levels in deep-sea fishes. Marine Ecology Progress Series, 637:225–235, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/637/m637p225.pdf.

Lane:2021:PFM

[LPJ^{+}21] J. V. Lane, C. J. Pollock, R. Jeavons, M. Sheddan, R. W. Furness, and K. C. Hamer. Post-fledging movements, mortality and migration of juvenile northern gannets. Marine Ecology Progress Series, 671:207–218, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m671p207.pdf.

Langan:2021:CAM

[LPO^{+}21] J. A. Langan, G. Puggioni, C. A. Oviatt, M. E. Henderson, and J. S. Collie. Climate alters the migration phenology of coastal marine species. Marine Ecology Progress Series, 660: 1–18, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Lauer:2022:DDB

[LR22] D. A. Lauer and M. L. Reaka. Depth distributions of benthic and pelagic species highlight the potential of mesophotic and deep habitats to serve as marine refugia. Marine Ecology Progress Series, 700:39–52, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/700/m700p039.pdf.

Li:2020:MAP

[LRC20] C. Li, C. E. Reimers, and J. W. Chapman. Microbiome analyses and presence of cable bacteria in the burrow sediment of Upogebia pugettensis. Marine Ecology Progress Series, 648:79–94, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m648p079.pdf.

Lovvorn:2020:PSO

[LRD^{+}20] J. R. Lovvorn, A. R. Rocha, S. L. Danielson, L. W. Cooper, J. M. Grebmeier, and K. S. Hedstrom. Predicting sediment organic carbon and related food web types from a physical oceanographic model on a subarctic shelf. Marine Ecology Progress Series, 633:37–54, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/633/m633p037.pdf.

Liszka:2021:DVM

[LRM^{+}21] C. M. Liszka, C. Robinson, C. Manno, G. Stowasser, and G. A. Tarling. Diel vertical migration of a Southern Ocean euphausiid, Euphausia triacantha, and its metabolic response to consequent short-term temperature changes. Marine Ecology Progress Series, 660:37–52, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/660/m660p037.pdf.

Lesser:2021:OPM

[LS21] M. P. Lesser and M. Slattery. Opinion piece: Mesophotic coral reef community structure: the constraints of imagery collected by unmanned vehicles. Marine Ecology Progress Series, 663:229–236, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/663/m663p229.pdf.

Lanna:2021:DED

[LSC^{+}21] E. Lanna, D. Santos, V. Carvalho, B. Vilela, G. G. Santos, L. F. Skinner, B. Segal, J. L. Carraro, and U. Pinheiro. Different environmental drivers influence the reproduction of a tropical sponge (Tedania ignis) along a latitudinal gradient. Marine Ecology Progress Series, 667:25–41, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/667/
m667p025.pdf.

Ludwig:2024:PPT

[LSKS24] K. E. Ludwig, A. Singer, I. Kröncke, and A. F. Sell. Predator– prey trait associations and feeding preferences of demersal fishes in the southern North Sea. Marine Ecology Progress Series, 739:173–190, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m739p173.pdf.

Lee:2024:ELH

[LSR^{+}24] B. Lee, F. Skeljo, H. S. Randhawa, W. Brownscombe, and A. Arkhipkin. Early life-history patterns in Patagonian toothfish Dissostichus eleginoides from the Patagonian Shelf. Marine Ecology Progress Series, 726:131–148, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2024/726/m726p131.pdf.

Longmire:2022:BRP

[LSS^{+}22] K. S. Longmire, R. D. Seitz, M. S. Seebo, R. W. Brill, and R. N. Lipcius. Biological responses of the predatory blue crab and its hard clam prey to ocean acidification and low salinity. Marine Ecology Progress Series, 701:67–81, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m701p067.pdf.

Ladds:2024:ATG

[LSSG24] M. Ladds, J. Smith, V. Strohm, and C. J. Gobler. Acartia tonsa grazing on the harmful dinoflagellate Dinophysis acuminata reduces copepod survival and increases extracellular toxin concentrations. Marine Ecology Progress Series, 745:25–40, September 19, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m745p025.pdf.

Longmire:2021:SSO

[LSSL21] K. S. Longmire, R. D. Seitz, A. Smith, and R. N. Lipcius. Saved by the shell: Oyster reefs can shield juvenile blue crabs Callinectes sapidus. Marine Ecology Progress Series, 672:163–173, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m672p163.pdf.

Ladah:2021:IWI

[LT21] L. B. Ladah and F. J. Tapia. Internal wave intensity and angle of propagation modulate small-scale settlement patterns of intertidal barnacles during peak recruitment. Marine Ecology Progress Series, 678:81–93, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/678/m678p081.pdf.

Lodi:2020:MCH

[LTM20] L. Lodi, R. Tardin, and G. Maricato. Modeling cetacean habitat use in an urban coastal area in southeastern Brazil. Marine Ecology Progress Series, 642:227–240, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/642/
m642p227.pdf.

Latijnhouwers:2024:SUD

[LvAH^{+}24] K. R. W. Latijnhouwers, M. van Aalst, J. Huckeba, M. W. Miller, M. J. A. Vermeij, and V. F. Chamberland. The sea urchin Diadema antillarum facilitates recruitment of the Critically Endangered Caribbean coral species Acropora palmata. Marine Ecology Progress Series, 746:35–48, October 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/746/m746p035.pdf.

Long:2024:HAB

[LWP^{+}24] C. A. Long, G. A. J. Worthy, R. Paperno, S. A. Ceriani, and K. L. Mansfield. Harmful algal bloom impacts on juvenile green turtle foraging ecology: insights from stable isotope analysis. Marine Ecology Progress Series, 728:59– 73, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/728/m728p059.pdf.

Mishra:2020:ECM

[MA20] A. K. Mishra and D. Apte. Ecological connectivity with mangroves influences tropical seagrass population longevity and meadow traits within an island ecosystem. Marine Ecology Progress Series, 644:47–63, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/644/m644p047.pdf.

McNeill:2020:FOB

[MAH^{+}20] J. Braun McNeill, L. Avens, A. Goodman Hall, I. Fujisaki, and A. R. Iverson. Foraging and overwintering behavior of loggerhead sea turtles Caretta caretta in the western North Atlantic. Marine Ecology Progress Series, 641:209–225, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/641/
m641p209.pdf.

Mahara:2022:ASB

[MAK^{+}22] N. Mahara, J. J. Alava, M. Kowal, E. Grant, J. L. Boldt, L. E. Kwong, and B. P. V. Hunt. Assessing size-based exposure to microplastic particles and ingestion pathways in zooplankton and herring in a coastal pelagic ecosystem of British Columbia, Canada. Marine Ecology Progress Series, 683:139–155, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/683/m683p139.pdf.

Medeiros:2021:EFI

[MAMF21] A. M. Medeiros, C. Ari, and E. L. A. Monteiro-Filho. Environmental factors involved in breaching behavior of manta rays in estuarine waters. Marine Ecology Progress Series, 674:203–219, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/674/m674p203.pdf.

Mir-Arguimbau:2021:TLB

[MAOR^{+}21] J. Mir-Arguimbau, M. P. Olivar, V. Raya, J. Navarro, and A. Sabatés. Trophic links between blue whiting (Micromesistius poutassou) larvae and the winter planktonic community in the NW Mediterranean Sea. Marine Ecology Progress Series, 670:185–201, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/670/m670p185.pdf.

Mayorga-Adame:2022:SSL

[MAPFH22] C. G. Mayorga-Adame, J. A. Polton, A. D. Fox, and L. A. Henry. Spatiotemporal scales of larval dispersal and connectivity among oil and gas structures in the North Sea. Marine Ecology Progress Series, 685:49–67, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m685p049.pdf.

Massaro:2020:DSA

[MAS^{+}20] M. Massaro, D. G. Ainley, J. A. Santora, P. Quillfeldt, A. Lescroël, A. Whitehead, A. Varsani, G. Ballard, and P. O. B. Lyver. Diet segregation in Adélie penguins: some individuals attempt to overcome colony-induced and annual foraging challenges. Marine Ecology Progress Series, 645:205– 218, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/645/m645p205.pdf.

McIntire:2020:WLC

[MB20] L. C. McIntire and P. E. Bourdeau. World’s largest chiton (Cryptochiton stelleri) is an inefficient thermoregulator. Marine Ecology Progress Series, 652:63–76, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/652/
m652p063.pdf.

Martins:2022:DIR

[MBB^{+}22] N. Martins, L. Barreto, I. Bartsch, J. Bernard, E. A. Serrão, and G. A. Pearson. Daylength influences reproductive success and sporophyte growth in the Arctic kelp species Alaria esculenta. Marine Ecology Progress Series, 683: 37–52, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/683/m683p037.pdf.

Munsch:2023:NPE

[MBB^{+}23] S. H. Munsch, F. L. Beaty, K. M. Beheshti, W. B. Chesney, C. A. Endris, T. G. Gerwing, M. Hessing-Lewis, P. M. Kiffney, J. K. O’Leary, L. Reshitnyk, B. L. Sanderson, and R. K. Walter. Northeast Pacific eelgrass dynamics: interannual expansion distances and meadow area variation over time. Marine Ecology Progress Series, 705:61–75, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/705/
m705p061.pdf.

Moullec:2023:RMM

[MBG^{+}23] F. Moullec, N. Barrier, F. Guilhaumon, M. A. Peck, C. Ulses, and Y. J. Shin. Rebuilding Mediterranean marine resources under climate change. Marine Ecology Progress Series, 708: 1–20, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Moore:2021:EFS

[MBGM21] M. E. Moore, B. A. Berejikian, C. M. Greene, and S. Munsch. Environmental fluctuation and shifting predation pressure contribute to substantial variation in early marine survival of steelhead. Marine Ecology Progress Series, 662:139–156, 2021.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m662p139.pdf.

Mul:2020:KWA

[MBM^{+}20] E. Mul, M. A. Blanchet, B. T. McClintock, W. J. Grecian, M. Biuw, and A. Rikardsen. Killer whales are attracted to herring fishing vessels. Marine Ecology Progress Series, 652: 1–13, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Moritz:2021:LTM

[MBR^{+}21] C. Moritz, S. J. Brandl, H. Rouzé, J. Vii, G. Pérez-Rosales, P. Bosserelle, Y. Chancerelle, R. Galzin, V. Liao, G. Siu, M. Taiarui, M. M. Nugues, and L. Hédouin. Long-term monitoring of benthic communities reveals spatial determinants of disturbance and recovery dynamics on coral reefs. Marine Ecology Progress Series, 672:141–152, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/672/
m672p141.pdf.

Martinez:2024:EAS

[MBS^{+}24] S. Martinez, A. T. F. Bernard, C. W. Speed, B. Q. Mann, J. M. Olbers, J. Q. Maggs, C. Floros, M. G. Meekan, and A. Yon. Elasmobranch assemblage structure on protected high-latitude coral reefs of southeast Africa. Marine Ecology Progress Series, 749:87–107, November 21, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/749/
m749p087.pdf.

Mangan:2020:SBD

[MBT^{+}20] S. Mangan, K. R. Bryan, S. F. Thrush, R. V. Gladstone-Gallagher, A. M. Lohrer, and C. A. Pilditch. Shady business: the darkening of estuaries constrains benthic ecosystem function. Marine Ecology Progress Series, 647:33–48, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/647/m647p033.pdf.

Maron:2020:FAS

[MBW^{+}20] C. F. Marón, S. M. Budge, R. E. Ward, L. O. Valenzuela, M. Di Martino, M. Ricciardi, M. Sironi, M. Uhart, J. Seger, and V. J. Rowntree. Fatty acids and stable isotopes (δ ^{13}C, δ ^{1}5N) in southern right whale Eubalaena australis calves in relation to age and mortality at Península Valdés, Argentina. Marine Ecology Progress Series, 646:189–200, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/646/
m646p189.pdf.

Muller:2021:FAP

[MCC^{+}21] C. G. Muller, B. L. Chilvers, A. Chiaradia, R. K. French, A. Kato, Y. Ropert-Coudert, and P. F. Battley. Foraging areas and plasticity of yellow-eyed penguins Megadyptes antipodes in their subantarctic range. Marine Ecology Progress Series, 679:149–162, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2021/679/m679p149.pdf.

Marchal:2021:UDF

[MCF^{+}21] P. Marchal, P. Cresson, A. Foveau, C. Giraldo, S. Lefebvre, and Y. Vérin. Using the diet of fish to reflect spatial patterns of their benthic prey. Marine Ecology Progress Series, 677:33–49, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/677/m677p033.pdf.

Muller:2020:DPA

[MCFB20] C. G. Muller, B. L. Chilvers, R. K. French, and P. F. Battley. Diving plasticity in the ancestral range of the yellow-eyed penguin Megadyptes antipodes, an endangered marine predator. Marine Ecology Progress Series, 648:191–205, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/648/
m648p191.pdf.

Moyano:2021:EBA

[MCH^{+}21] M. P. Sal Moyano, M. Ceraulo, F. J. Hidalgo, T. Luppi, J. Nuñez, C. A. Radford, S. Mazzola, M. A. Gavio, and G. Buscaino. Effect of biological and anthropogenic sound on the orientation behavior of four species of brachyuran crabs.

Marine Ecology Progress Series, 669:107–120, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/669/
m669p107.pdf.

Morra:2020:SDS

[MCJ^{+}20] K. E. Morra, Y. Chikaraishi, H. F. James, S. Rossman, A. E. Wiley, and P. H. Ostrom. Seasonality of decadal-scale trophic declines and nutrient regime shifts in the Laysan albatross and Newell’s shearwater. Marine Ecology Progress Series, 654:163–175, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m654p163.pdf.

Mamo:2021:EEB

[MCK21] L. T. Mamo, M. A. Coleman, and B. P. Kelaher. Ecological enhancement of breakwater upgrades: size and type of rocks used influence benthic communities. Marine Ecology Progress Series, 661:71–82, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/661/m661p071.pdf.

MacNeil:2024:GSB

[MCL24] L. MacNeil, M. Costa, and J. LaRoche. Glimpsing the 2020 spring bloom in the Strait of Georgia (Canada) with autonomous ferry-based sensors. Marine Ecology Progress Series, 736:181–187, May 23, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m736p181.pdf.

Meirelles:2023:HUG

[MCLC^{+}23] A. C. O. Meirelles, K. F. Choi-Lima, T. M. Campos, E. L. d. A. Monteiro-Filho, and T. M. C. Lotufo. Habitat use of Guiana dolphin Sotalia guianensis in a heavily urbanized embayment. Marine Ecology Progress Series, 722:195–206, November 9, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/722/m722p195.pdf.

Maathuis:2024:RVH

[MCvdM^{+}24] M. A. M. Maathuis, B. Couperus, J. van der Molen, J. J. Poos, I. Tulp, and S. Sakinan. Resolving the variability in habitat use by juvenile small pelagic fish in a major tidal system by continuous echosounder measurements. Marine Ecology Progress Series, 741:169–187, July 30, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m741p169.pdf.

Moe:2021:TFE

[MDB^{+}21] B. Moe, F. Daunt, V. S. Bråthen, R. T. Barrett, M. Ballesteros, O. Bjørnstad, M. I. Bogdanova, N. Dehnhard, K. E. Erikstad, A. Follestad, S. Gíslason, G. T. Hallgrimsson, S. H. Lorentsen, M. Newell, A. Petersen, R. A. Phillips, S. B. Ragnarsdóttir, T. K. Reiertsen, J. Åström, S. Wanless, and T. Anker-Nilssen. Twilight foraging enables European shags to survive the winter across their latitudinal range. Marine Ecology Progress Series, 676:145–157, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m676p145.pdf.

McLaverty:2021:BTB

[MDG^{+}21] C. McLaverty, G. E. Dinesen, H. Gislason, M. E. Brooks, and O. R. Eigaard. Biological traits of benthic macrofauna show size-based differences in response to bottom trawling intensity. Marine Ecology Progress Series, 671:1–19, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Marre:2020:FSA

[MDH^{+}20] G. Marre, J. Deter, F. Holon, P. Boissery, and S. Luque. Finescale automatic mapping of living Posidonia oceanica seagrass beds with underwater photogrammetry. Marine Ecology Progress Series, 643:63–74, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/643/m643p063.pdf.

Mortimer:2021:ESC

[MDH^{+}21] C. Mortimer, M. Dunn, A. Haris, J. Jompa, and J. Bell. Estimates of sponge consumption rates on an Indo–Pacific reef. Marine Ecology Progress Series, 672:123–140, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/672/
m672p123.pdf.

McCarthy:2021:DPA

[MDK^{+}21] A. McCarthy, A. De Robertis, S. Kotwicki, K. Hough, P. Wade, and C. Wilson. Differing prey associations and habitat use suggest niche partitioning by fin and humpback whales off Kodiak Island. Marine Ecology Progress Series, 662:181–197, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/662/m662p181.pdf.

McClanahan:2020:HVT

[MDM^{+}20] T. R. McClanahan, E. S. Darling, J. M. Maina, N. A. Muthiga, S. D’agata, J. Leblond, R. Arthur, S. D. Jupiter, S. K. Wilson, S. Mangubhai, A. M. Ussi, M. M. M. Guillaume, A. T. Humphries, V. Patankar, G. Shedrawi, J. Pagu, and G. Grimsditch. Highly variable taxa-specific coral bleaching responses to thermal stresses. Marine Ecology Progress Series, 648:135–151, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m648p135.pdf.

Moreira:2024:TUP

[MdMR24] A. L. P. Moreira, P. R. de Medeiros, and N. C. Roos. Two ubiquitous parrotfishes exhibit distinct foraging ecologies on tropical Brazilian reefs. Marine Ecology Progress Series, 730: 79–93, March 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/730/m730p079.pdf.

Machado:2020:TOB

[MdOO^{+}20] R. Machado, L. R. de Oliveira, P. H. Ott, M. Haimovici, L. G. Cardoso, L. Milmann, M. A. Romero, R. A. dos Santos, and M. Borges-Martins. Trophic overlap between marine mammals and fisheries in subtropical waters in the western South Atlantic. Marine Ecology Progress Series, 639: 215–232, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/639/m639p215.pdf.

Marley:2020:ENT

[MDPL20] G. S. A. Marley, A. E. Deacon, D. A. T. Phillip, and A. J. Lawrence. Effects of a no-take reserve on mangrove fish assemblages: incorporating seascape connectivity. Marine Ecology Progress Series, 649:141–153, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/649/m649p141.pdf.

Myksvoll:2021:LDC

[MDQ^{+}21] M. S. Myksvoll, J. Devine, M. Quintela, A. J. Geffen, R. D. M. Nash, A. Sandvik, F. Besnier, A. Saha, G. Dahle, E. Jansson, K. Nedreaas, and T. Johansen. Linking dispersal connectivity to population structure and management boundaries for saithe in the Northeast Atlantic. Marine Ecology Progress Series, 680:177–191, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/680/m680p177.pdf.

Morgan:2021:RBD

[MDS^{+}21] S. G. Morgan, C. D. Dibble, M. G. Susner, T. G. Wolcott, D. L. Wolcott, and J. L. Largier. Robotic biomimicry demonstrates behavioral control of planktonic dispersal in the sea. Marine Ecology Progress Series, 663:51–61, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m663p051.pdf.

Marsaly:2023:CER

[MDS^{+}23] B. Marsaly, D. Daugherty, O. N. Shipley, C. Gelpi, N. Boyd, J. Davis, M. Fisher, and P. Matich. Contrasting ecological roles and flexible trophic interactions of two estuarine apex predators in the western Gulf of Mexico. Marine Ecology Progress Series, 709:55–76, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/709/m709p055.pdf.

Madigan:2022:CTF

[MDW^{+}22] D. J. Madigan, B. M. Devine, S. B. Weber, A. L. Young, and N. E. Hussey. Combining telemetry and fisheries data to quantify species overlap and evaluate bycatch mitigation strategies in an emergent Canadian Arctic fishery. Marine Ecology Progress Series, 702:1–17, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Merkel:2021:SMC

[MDY^{+}21] B. Merkel, S. Descamps, N. G. Yoccoz, D. Grémillet, P. Fauchald, J. Danielsen, F. Daunt, K. E. Erikstad, A. V.

Ezhov, M. P. Harris, M. Gavrilo, S. H. Lorentsen, T. K. Reiertsen, G. H. Systad, T. L. Thórarinsson, S. Wanless, and H. Strøm. Strong migratory connectivity across metapopulations of sympatric North Atlantic seabirds. Marine Ecology Progress Series, 676:173–188, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m676p173.pdf.

McClanahan:2024:MSD

[MFC^{+}24] T. R. McClanahan, A. M. Friedlander, P. Chabanet, J. H. Bruggemann, J. Wickel, and M. K. Azali. Modeling the spatial distribution of numbers of coral reef fish species and community types in the Western Indian Ocean faunal province. Marine Ecology Progress Series, 730:59–78, March 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m730p059.pdf.

Mello-Fonseca:2021:SDS

[MFCF21] J. Mello-Fonseca, C. A. M. M. Cordeiro, and C. E. L. Ferreira. Spatial distribution of sea turtles on South Atlantic subtropical reefs. Marine Ecology Progress Series, 678:125– 138, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/678/m678p125.pdf.

Mendt:2022:RIE

[MG22] S. R. Mendt and L. A. Gosselin. Role of initial energy reserves in stress tolerance thresholds during the early benthic phase in intertidal invertebrates. Marine Ecology Progress Series, 686:1–13, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

McGowan:2020:STD

[MGA^{+}20] D. W. McGowan, E. D. Goldstein, M. L. Arimitsu, A. L. Deary, O. Ormseth, A. De Robertis, J. K. Horne, L. A. Rogers, M. T. Wilson, K. O. Coyle, K. Holderied, J. F. Piatt, W. T. Stockhausen, and S. Zador. Spatial and temporal dynamics of Pacific capelin Mallotus catervarius in the Gulf of Alaska: implications for ecosystem-based fisheries management. Marine Ecology Progress Series, 637:117–140, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/637/
m637p117.pdf.

Mion:2022:NPG

[MGB^{+}22] M. Mion, C. A. Griffiths, V. Bartolino, S. Haase, A. Hilvarsson, K. Hüssy, M. Krüger-Johnsen, U. Krumme, R. B. C. Lundgreen, J. Lövgren, K. McQueen, M. Plikshs, K. Radtke, J. Raitaniemi, and M. Casini. New perspectives on Eastern Baltic cod movement patterns from historical and contemporary tagging data. Marine Ecology Progress Series, 689:109–126, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/689/m689p109.pdf.

Miron-Gaton:2022:TTD

[MGBCGM^{+}22] J. M. Mirón-Gatón, M. Botella-Cruz, A. J. García-Meseguer, A. Millán, and J. Velasco. Thermal tolerance differs between co-occurring congeneric beetle species in marine supratidal rockpools. Marine Ecology Progress Series, 681:185–196, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/681/m681p185.pdf.

Moore:2021:PFS

[MGH21] A. F. P. Moore, C. A. Gehring, and A. R. Hughes. Plantfungal symbiosis responds to experimental addition of resources and physical stressor in a salt marsh. Marine Ecology Progress Series, 661:115–125, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/661/m661p115.pdf.

Martinez-Garcia:2021:ALG

[MGHH21] L. Martínez-García, B. Hansson, and J. Hollander. Assessment of local genetic structure and connectivity of the common eelgrass Zostera marina for seagrass restoration in northern Europe. Marine Ecology Progress Series, 664:103– 116, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m664p103.pdf.

Murray:2024:EEP

[MGM^{+}24] C. S. Murray, J. L. Gregg, A. H. Mackenzie, H. T. Jayasekera, S. Hall, T. Klinger, and P. K. Hershberger. Effects of elevated pCO_{2} on bioenergetics and disease susceptibility in Pacific herring Clupea pallasii. Marine Ecology Progress Series, 738:225–242, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m738p225.pdf.

Methou:2022:PSE

[MHÁC^{+}22] P. Methou, I. Hernández-Ávila, C. Cathalot, M. A. Cambon-Bonavita, and F. Pradillon. Population structure and environmental niches of Rimicaris shrimps from the Mid-Atlantic Ridge. Marine Ecology Progress Series, 684:1–20, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Mason:2024:SAT

[MHAL24] C. Mason, A. J. Hobday, R. Alderman, and M. A. Lea. Shy albatross Thalassarche cauta chick mortality and heat stress in a temperate climate. Marine Ecology Progress Series, 737: 137–145, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m737p137.pdf.

Murray:2021:TYY

[MHDJ21] K. T. Murray, J. M. Hatch, R. A. DiGiovanni, Jr., and E. Josephson. Tracking young-of-the-year gray seals Halichoerus grypus to estimate fishery encounter risk. Marine Ecology Progress Series, 671:235–245, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m671p235.pdf.

Medcalf:2021:WTE

[MHF^{+}21] K. E. Medcalf, J. A. Hutchings, M. D. Fast, A. Kuparinen, and S. C. Godwin. Warming temperatures and ectoparasitic sea lice impair internal organs in juvenile Atlantic salmon. Marine Ecology Progress Series, 660:161–169, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m660p161.pdf.

Malick:2020:RBT

[MHH^{+}20] M. J. Malick, M. E. Hunsicker, M. A. Haltuch, S. L. Parker-Stetter, A. M. Berger, and K. N. Marshall. Relationships between temperature and Pacific hake distribution vary across latitude and life-history stage. Marine Ecology Progress Series, 639:185–197, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/639/m639p185.pdf.

Munger:2022:MLA

[MHH^{+}22] J. E. Munger, D. P. Herrera, S. M. Haver, L. Waterhouse, M. F. McKenna, R. P. Dziak, J. Gedamke, S. A. Heppell, and J. H. Haxel. Machine learning analysis reveals relationship between pomacentrid calls and environmental cues. Marine Ecology Progress Series, 681:197–210, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m681p197.pdf.

Momota:2022:SSH

[MHK22a] K. Momota, S. Hosokawa, and T. Komuro. Small-scale heterogeneity of fish diversity evaluated by environmental DNA analysis in eelgrass beds. Marine Ecology Progress Series, 688: 99–112, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m688p099.pdf.

Moynihan:2022:IBS

[MHK22b] J. L. Moynihan, A. E. Hall, and M. J. Kingsford. Interrelationships between soft corals and reef-associated fishes on inshorereefs of the Great Barrier Reef. Marine Ecology Progress Series, 698:15–28, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Makri:2024:IST

[MHN24] M. Makri, P. J. Hansen, and T. G. Nielsen. Impact of salinity and temperature on the vital rates of co-occurring Calanus glacialis and C. finmarchicus from West Greenland. Marine Ecology Progress Series, 729:47–62, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m729p047.pdf.

McKnight:2020:CII

[MIL^{+}20] A. McKnight, D. B. Irons, C. S. Loftin, S. T. McKinney, and B. J. Olsen. Combined influence of intrinsic and environmental factors in shaping productivity in a small pelagic gull, the black-legged kittiwake Rissa tridactyla. Marine Ecology Progress Series, 633:207–223, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/633/m633p207.pdf.

McCann:2021:EGS

[MJ21] B. N. McCann and D. W. Johnson. Estimating growth, size-dependent mortality, and tag loss in a mark-recapture study: demography of wavy turban snails in Southern California, USA. Marine Ecology Progress Series, 659:97–112, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/659/m659p097.pdf.

MacGregor:2023:RDK

[MJ23] K. A. MacGregor and L. E. Johnson. Response to drift kelp is modified by both substratum and season: green sea urchin foraging behaviour in subsidized habitats. Marine Ecology Progress Series, 710:85–105, May 4, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/710/
m710p085.pdf.

Matley:2020:EDD

[MJJJ20] J. K. Matley, J. Jossart, L. Johansen, and P. D. Jobsis. Environmental drivers of diving behavior and space-use of juvenile endangered Caribbean hawksbill sea turtles identified using acoustic telemetry. Marine Ecology Progress Series, 652:157–171, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/652/m652p157.pdf.

Morrissey:2024:UET

[MJK24] S. J. Morrissey, D. R. Jerry, and M. J. Kingsford. Use of eDNA to test hypotheses on the ecology of Chironex fleckeri (Cubozoa). Marine Ecology Progress Series, 728:25–41, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/728/m728p025.pdf.

McCarthy:2024:SSC

[MJW24] G. J. McCarthy, T. W. D. Jowett, and S. R. Wing. Shifts in species composition in kelp forest communities: implications of differences in total phenolic composition among species. Marine Ecology Progress Series, 744:17–31, September 5, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m744p017.pdf.

Meyer-Kaiser:2020:CEB

[MK20] K. Meyer-Kaiser. Carryover effects of brooding conditions on larvae in the slipper limpet Crepidula fornicata. Marine Ecology Progress Series, 643:87–97, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/643/
m643p087.pdf.

Michelot:2020:SIE

[MKR^{+}20] C. Michelot, A. Kato, T. Raclot, K. Shiomi, P. Goulet, P. Bustamante, and Y. Ropert-Coudert. Sea-ice edge is more important than closer open water access for foraging Adélie penguins: evidence from two colonies. Marine Ecology Progress Series, 640:215–230, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/640/m640p215.pdf.

Matsuno:2020:IMD

[MKS^{+}20] K. Matsuno, N. Kanna, S. Sugiyama, A. Yamaguchi, and E. J. Yang. Impacts of meltwater discharge from marine– terminating glaciers on the protist community in Inglefield Bredning, northwestern Greenland. Marine Ecology Progress Series, 642:55–65, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/642/m642p055.pdf.

Morzaria-Luna:2022:IDW

[MLAS22] H. N. Morzaria-Luna, C. H. Ainsworth, and R. L. Scott. Impacts of deep-water spills on mesopelagic communities and implications for the wider pelagic food web. Marine Ecology Progress Series, 681:37–51, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/681/m681p037.pdf.

McManus:2021:SPE

[MLB^{+}21] M. C. McManus, J. A. Langan, R. J. Bell, J. S. Collie, G. Klein-MacPhee, M. D. Scherer, and R. G. Balouskus. Spatiotemporal patterns in early life stage winter flounder Pseudopleuronectes americanus highlight phenology changes and habitat dependencies. Marine Ecology Progress Series, 677:161–175, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/677/m677p161.pdf.

Madrak:2022:EAT

[MLE^{+}22] S. V. Madrak, R. L. Lewison, T. Eguchi, A. P. Klimley, and J. A. Seminoff. Effects of ambient temperature on dive behavior of East Pacific green turtles before and after a power plant closure. Marine Ecology Progress Series, 683: 157–168, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/683/m683p157.pdf.

Muller:2024:CCA

[MLP^{+}24] C. Muller, C. Lett, F. Porri, P. Pattrick, D. Bailey, H. Denis, N. Barrier, W. Potts, and D. M. Kaplan. Coastal connectivity of an abundant inshore fish species: model-data comparison along the southern coast of South Africa. Marine Ecology Progress Series, 731:89–104, March 13, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m731p089.pdf.

Meister:2023:BMW

[MLVJ23] N. Meister, T. J. Langbehn, Ø. Varpe, and C. Jørgensen. Blue mussels in western Norway have vanished where in reach of crawling predators. Marine Ecology Progress Series, 721:85– 101, October 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m721p085.pdf.

Morgulis:2022:BCA

[MMA^{+}22] M. Morgulis, S. Martinez, R. Almuly, S. Einbinder, P. Zaslansky, and T. Mass. Black corals (Antipatharia) of the northern Red Sea: ancient predators of the mesophotic reef. Marine Ecology Progress Series, 688:33–47, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/688/
m688p033.pdf.

Mejia-Mercado:2022:CSV

[MMB22] B. E. Mejía-Mercado and A. R. Baco. Characterization and spatial variation of the deep-sea fish assemblages on Pioneer Bank, Northwestern Hawaiian Islands. Marine Ecology Progress Series, 692:99–118, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/692/m692p099.pdf.

McNair:2020:PGC

[MMD20] H. M. McNair and S. Menden-Deuer. Protist grazing contributes to microbial food web at the upper boundary of the twilight zone in the subarctic Pacific. Marine Ecology Progress Series, 636:235–241, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/636/m636p235.pdf.

Martinez-Moreno:2024:SBM

[MMLPP^{+}24] S. Martínez-Moreno, E. Leon-Palmero, H. J. Pula, A. M. Cabello, I. Ferrera, and I. Reche. Subcuticular and biofilm microbiomes in Holothuria tubulosa and their potential for denitrification. Marine Ecology Progress Series, 736:81–92, May 23, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m736p081.pdf.

Medina:2022:MRS

[MMMV22] A. Medina, A. Magro, D. Macías, and J. L. Varela. Monitoring the reproductive status of resident and migrant Atlantic bluefin tuna in the Strait of Gibraltar. Marine Ecology Progress Series, 697:97–108, September 22, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/697/
m697p097.pdf.

Mulders:2022:PDD

[MMP^{+}22] Y. Mulders, L. Mattio, J. C. Phillips, P. S. Lavery, G. A. Kendrick, and T. Wernberg. Patch dynamics driven by wave exposure in subtidal temperate seaweeds are exacerbated by warming oceans. Marine Ecology Progress Series, 685:

85–95, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m685p085.pdf.

Mueller:2020:SCF

[MMS^{+}20] C. Mueller, A. Monczak, J. Soueidan, B. McKinney, S. Smott, T. Mills, Y. Ji, and E. W. Montie. Sound characterization and fine-scale spatial mapping of an estuarine soundscape in the southeastern USA. Marine Ecology Progress Series, 645:1–23, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Monczak:2022:SCS

[MMS^{+}22] A. Monczak, B. McKinney, J. Souiedan, A. D. Marian, A. Seder, E. May, T. Morgenstern, W. Roumillat, and E. W. Montie. Sciaenid courtship sounds correlate with juvenile appearance and abundance in the May River, South Carolina, USA. Marine Ecology Progress Series, 693:1–17, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Moore:2024:HSP

[MMT^{+}24] M. E. Moore, M. J. Malick, A. C. Thomas, M. M. Klungle, and B. A. Berejikian. Harbor seal predation on migrating steelhead smolts entering marine waters. Marine Ecology Progress Series, 743:139–157, August 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m743p139.pdf.

Meysick:2020:CDE

[MNG^{+}20] L. Meysick, A. Norkko, K. Gagnon, M. Gräfnings, and C. Boström. Context-dependency of eelgrass-clam interactions: implications for coastal restoration. Marine Ecology Progress Series, 647:93–108, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/647/m647p093.pdf.

Montagna:2023:BIR

[Mon23] P. A. Montagna. Benthic infauna are resistant and resilient to hurricane disturbance. Marine Ecology Progress Series, 707: 1–13, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Methou:2024:GCI

[MON^{+}24] P. Methou, N. O. Ogawa, H. Nomaki, N. Ohkouchi, C. Chen, and K. Schnabel. Genetic connectivity and isotopic niches of alvinocaridid shrimps from chemosynthetic habitats in Aotearoa/New Zealand, with a new Alvinocaris species. Marine Ecology Progress Series, 739:85–109, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2024/739/m739p085.pdf.

Macartney:2022:DDT

[MPA^{+}22] K. J. Macartney, M. S. Pankey, A. Clayshulte Abraham, M. Slattery, and M. P. Lesser. Depth-dependent trophic strategies of Caribbean sponges on mesophotic coral reefs. Marine Ecology Progress Series, 693:125–140, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/693/
m693p125.pdf.

McCauley:2024:TOF

[MPB^{+}24] D. J. McCauley, J. K. Parsons, C. A. Braman, J. M. Anderson, J. E. Caselle, E. J. Critchley, A. Glina, F. H. Joyce, C. G. Lowe, S. Mladjov, N. Nathan, P. T. Rex, E. Spurgeon, B. S. Stirling, and H. S. Young. Temporal and oceanographic factors differentially affect two size classes of white shark at a Southern California aggregation site. Marine Ecology Progress Series, 744:101–114, September 5, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/744/
m744p101.pdf.

Merkel:2022:BWG

[MPF^{+}22] F. R. Merkel, S. Post, M. Frederiksen, Z. Bak-Jensen, J. Nielsen, and R. B. Hedeholm. Bycatch in the West Greenland lumpfish fishery, with particular focus on the common eider population. Marine Ecology Progress Series, 702:123– 137, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m702p123.pdf.

Munnelly:2024:BMB

[MPKS24] R. T. Munnelly, B. R. Pittinger, S. F. Keenan, and T. S. Switzer. Benthic modification and biotic associations at natural and artificial habitats excavated by Epinephelus morio and Lutjanus campechanus. Marine Ecology Progress Series, 745:125–145, September 19, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/745/m745p125.pdf.

Mackiewicz:2021:EVS

[MPM21] A. G. Mackiewicz, R. L. Putland, and A. F. Mensinger. Effects of vessel sound on oyster toadfish Opsanus tau calling behavior. Marine Ecology Progress Series, 662:115–124, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/662/m662p115.pdf.

Maar:2023:LMB

[MPR^{+}23] K. Maar, M. Papadimitraki, L. Hernández Ruiz, R. Broughton, and S. H. Jónasdóttir. Lipids in meso- and bathypelagic fishes from the North Atlantic Ocean: dietary inputs suggested from fatty acid trophic markers. Marine Ecology Progress Series, 717:127–141, August 24, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m717p127.pdf.

Mickle:2022:ASI

[MPSH22] M. F. Mickle, R. Pieniazek, J. J. Stasso, and D. M. Higgs. Anthropogenic sounds induce escape behaviour in southern stingrays Hypanus americanus. Marine Ecology Progress Series, 694:125–132, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/694/m694p125.pdf.

Minuti:2020:FRH

[MR20] J. J. Minuti and B. D. Russell. Functionally redundant herbivores: urchin and gastropod grazers respond differently to ocean warming and rising CO_{2}. Marine Ecology Progress Series, 656:239–251, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/656/m656p239.pdf.

Mason:2024:RLI

[MRB^{+}24] E. T. Jarvis Mason, T. V. Riecke, L. F. Bellquist, D. J. Pondella II, and B. X. Semmens. Recruitment limitation increases susceptibility to fishing-induced collapse in a spawning aggregation fishery. Marine Ecology Progress Series, 738:203– 224, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m738p203.pdf.

Munroe:2020:EII

[MRH20] S. E. M. Munroe, C. L. Rigby, and N. E. Hussey. Evidence for inter- and intraspecific trophic niche separation among deepwater elasmobranchs on the southern Great Barrier Reef, Australia. Marine Ecology Progress Series, 636: 107–121, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/636/m636p107.pdf.

Moxley:2020:DSM

[MSC^{+}20] J. H. Moxley, G. Skomal, J. Chisholm, P. Halpin, and D. W. Johnston. Daily and seasonal movements of Cape Cod gray seals vary with predation risk. Marine Ecology Progress Series, 644:215–228, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m644p215.pdf.

Maes:2022:CVD

[MSC^{+}22] S. M. Maes, F. L. Schaafsma, H. Christiansen, B. Hellemans, M. Lucassen, F. C. Mark, H. Flores, and F. A. M. Volckaert. Comparative visual and DNA-based diet assessment extends the prey spectrum of polar cod Boreogadus saida. Marine Ecology Progress Series, 698:139–154, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m698p139.pdf.

Miller:2023:NRF

[MSDY23] E. Miller, E. Spanier, R. Diamant, and R. Yahel. Nature reserves facilitate conservation of the Mediterranean slipper lobster Scyllarides latus. Marine Ecology Progress Series, 714: 57–69, July 13, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/714/m714p057.pdf.

McHuron:2024:EMH

[MSF24] E. A. McHuron, K. L. Sweeney, and B. S. Fadely. Effects of the 2014–2016 marine heatwave on Steller sea lions in the Gulf of Alaska and implications for top-down forcing. Marine Ecology Progress Series, 736:129–145, May 23, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/736/
m736p129.pdf.

Matich:2020:LTM

[MSH20] P. Matich, B. A. Strickland, and M. R. Heithaus. Longterm monitoring provides insight into estuarine top predator (Carcharhinus leucas) resilience following an extreme weather event. Marine Ecology Progress Series, 639:169– 183, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/639/m639p169.pdf.

Martin:2021:DOM

[MSL^{+}21] P. Martin, N. Sanwlani, T. W. Q. Lee, J. M. C. Wong, K. Y. W. Chang, E. W. S. Wong, and S. C. Liew. Dissolved organic matter from tropical peatlands reduces shelf sea light availability in the Singapore Strait, Southeast Asia. Marine Ecology Progress Series, 672:89–109, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m672p089.pdf.

Marina:2024:CNT

[MSL^{+}24] T. I. Marina, I. R. Schloss, G. A. Lovrich, C. C. Boy, D. O. Bruno, F. L. Capitanio, S. M. Delpiani, J. M. Díaz de Astarloa, C. Fraysse, V. A. García Alonso, A. Raya Rey, L. Schejter, M. L. Spinelli, M. Tatián, D. Urteaga, and L. Riccialdelli. Complex network of trophic interactions in Burdwood Bank, a sub-Antarctic oceanic marine protected area. Marine Ecology Progress Series, 736:1–18, May 23, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

McCarthy:2022:IDS

[MSPS22] O. S. McCarthy, J. E. Smith, V. Petrovic, and S. A. Sandin. Identifying the drivers of structural complexity on Hawaiian coral reefs. Marine Ecology Progress Series, 702:71–86, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/702/m702p071.pdf.

Martins:2021:EDH

[MSR^{+}21a] S. Martins, L. Sierra, E. Rodrigues, J. Oñate-Casado, I. Torres Galán, L. J. Clarke, and A. Marco. Ecological drivers of the high predation of sea turtle hatchlings during emergence. Marine Ecology Progress Series, 668:97–106, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/668/
m668p097.pdf.

McBride:2021:RAK

[MSR^{+}21b] M. M. McBride, O. Schram Stokke, A. H. H. Renner, B. A. Krafft, O. A. Bergstad, M. Biuw, A. D. Lowther, and J. E. Stiansen. Review: Antarctic krill Euphausia superba: spatial distribution, abundance, and management of fisheries in a changing climate. Marine Ecology Progress Series, 668: 185–214, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m668p185.pdf.

Martell:2022:SBD

[MSTH22] L. Martell, K. Selsø, J. Titelman, and A. Hosia. Setting the baseline for the dynamics of siphonophores and hydromedusae in Oslofjorden. Marine Ecology Progress Series, 686: 71–89, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m686p071.pdf.

Munoz:2021:CSI

[MSV21] C. C. Muñoz, T. Saito, and P. Vermeiren. Cohort structure and individual resource specialization in loggerhead turtles, long-lived marine species with ontogenetic migrations. Marine Ecology Progress Series, 671:175–190, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/671/
m671p175.pdf.

Merrill:2021:MFT

[MTB21] G. B. Merrill, J. W. Testa, and J. M. Burns. Maternal foraging trip duration as a population-level index of foraging and reproductive success for the northern fur seal. Marine Ecology Progress Series, 666:217–229, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m666p217.pdf. See

erratum [Ano21g].

Marques:2023:PCD

[MTB^{+}23] T. A. Marques, L. Thomas, C. G. Booth, L. P. Garrison, P. E. Rosel, R. Takeshita, K. D. Mullin, and L. Schwacke. Population consequences of the Deepwater Horizon oil spill on pelagic cetaceans. Marine Ecology Progress Series, 714: 1–14, July 13, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

McComb-Turbitt:2023:DED

[MTCT^{+}23] S. P. McComb-Turbitt, G. T. Crossin, M. Tierney, P. Brickle, P. Trathan, T. D. Williams, and M. Auger-Méthé. Diving efficiency at depth and pre-breeding foraging effort increase with haemoglobin levels in gentoo penguins. Marine Ecology Progress Series, 722:1–17, November 9, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Meenakshisundaram:2021:GMV

[MTK^{+}21] A. Meenakshisundaram, L. Thomas, W. J. Kennington, M. Thums, E. Lester, and M. Meekan. Genetic markers validate photo-identification and uniqueness of spot patterns in whale sharks. Marine Ecology Progress Series, 668:177– 183, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/668/m668p177.pdf.

Moriarty:2021:BPM

[MTRM21] M. Moriarty, D. Tulett, B. Rabe, and A. G. Murray. Biophysical models for the management of micropathogens in Scottish aquaculture: a preliminary view to farming further offshore. Marine Ecology Progress Series, 679:133–147, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/679/m679p133.pdf.

Miyajima-Taga:2024:FSS

[MTSO^{+}24] Y. Miyajima-Taga, M. Sato, K. Oi, N. Furuichi, and N. Inoue. Fine-scale spatial distribution of a fish community in artificial reefs investigated using an underwater drone and environmental DNA analysis. Marine Ecology Progress Series, 740:123–144, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2024/740/m740p123.pdf.

McAllister:2023:ESC

[MVC^{+}23] L. T. McAllister, T. E. Van Leeuwen, C. Conway, M. Wälle, and M. Abrahams. Effect of sea cage aquaculture on the length, weight, and condition of wild Atlantic cod Gadus morhua. Marine Ecology Progress Series, 719:109–123, September 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/719/m719p109.pdf.

McAllister:2021:SCA

[MVH^{+}21] L. T. McAllister, T. E. Van Leeuwen, J. M. Hanlon, C. J. Morris, J. Potter, J. Wells, and M. Abrahams. Sea cage aquaculture may provide an energetic subsidy to wild juvenile cod in coastal bays of southern Newfoundland, Canada. Marine Ecology Progress Series, 674:241–255, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/674/
m674p241.pdf.

McFarland:2022:SES

[MVSV22] K. McFarland, J. Vignier, E. Standen, and A. K. Volety. Synergistic effects of salinity and temperature on the eastern oyster Crassostrea virginica throughout the lifespan. Marine Ecology Progress Series, 700:111–124, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/700/
m700p111.pdf.

Mulders:2020:FYG

[MW20] Y. R. Mulders and T. Wernberg. Fifteen years in a global warming hotspot: changes in subtidal mobile invertebrate communities. Marine Ecology Progress Series, 656:227–238, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m656p227.pdf.

McHuron:2020:CWG

[MWCR20] E. A. McHuron, T. Williams, D. P. Costa, and C. Reichmuth. Contrasting whisker growth dynamics within the phocid lineage. Marine Ecology Progress Series, 634:231– 236, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2019/634/m634p231.pdf.

Matsumoto:2021:FRP

[MY21] Y. Matsumoto and K. Yatsuya. Fenton reaction as a possible stimulus to induce spawning in Ezo abalone Haliotis discus hannai during stormy weather. Marine Ecology Progress Series, 669:241–246, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/669/m669p241.pdf.

Mattos:2023:BED

[MYSF23] F. M. G. Mattos, T. Yeemin, M. Sutthacheep, and J. L. L. Feitosa. Biogeographic and environmental drivers of reef fish diversity in coastal islands of the Gulf of Thailand. Marine Ecology Progress Series, 724:33–46, December 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2023/724/m724p033.pdf.

Noe:2024:RNI

[NBDM^{+}24] S. Noè, C. Bommarito, D. M. Díaz-Morales, T. Guy-Haim, K. Ermak, M. Wahl, B. Sures, A. R. Morov, and G. Rilov. Response of native and invasive grazers to global warming: marine heatwaves vs. gradual trends. Marine Ecology Progress Series, 748:69–82, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/748/
m748p069.pdf.

Nur:2022:EWS

[NBL^{+}22] N. Nur, R. W. Berger, D. E. Lee, P. M. Warzybok, and J. Jahncke. Effects of winter storms and oceanographic conditions on survival to weaning: a 37 year study of northern elephant seals on the Farallon Islands. Marine Ecology Progress Series, 691:173–189, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/691/m691p173.pdf.

Nyegaard:2023:OCT

[NBW^{+}23] M. Nyegaard, C. D. Braun, M. Welly, R. Djohani, and M. C. Arostegui. Overcoming challenging telemetry data of giant sunfish Mola alexandrini (Molidae) in Bali, Indonesia. Marine Ecology Progress Series, 722:157–175, November 9, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/722/m722p157.pdf.

Neves:2023:CFH

[NCS^{+}23] V. C. Neves, G. Carroll, W. C. Schäfer, H. F. R. Hereward, and P. Quillfeldt. Consistent foraging habitat use by a vulnerable breeding seabird highlights potential areas for protection in the mid-Atlantic Ocean. Marine Ecology Progress Series, 716:107–121, August 10, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/716/m716p107.pdf.

Ning:2021:TCM

[NCZ^{+}21] Z. Ning, C. Chen, Z. Zhu, T. Xie, Q. Wang, B. Cui, and T. J. Bouma. Tidal channel-mediated gradients facilitate Spartina alterniflora invasion in coastal ecosystems: implications for invasive species management. Marine Ecology Progress Series, 659:59–73, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/659/m659p059.pdf.

Nolasco:2022:BML

[NDA^{+}22] R. Nolasco, J. Dubert, J. L. Acuña, A. Aguión, T. Cruz, J. N. Fernandes, K. J. Geiger, D. Jacinto, G. Macho, D. Mateus, A. Rivera, S. Román, E. Thiébaut, E. Vazquez, and H. Queiroga. Biophysical modelling of larval dispersal and population connectivity of a stalked barnacle: implications for fishery governance. Marine Ecology Progress Series, 694: 105–123, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/694/m694p105.pdf.

Nordio:2023:RSM

[NF23] G. Nordio and S. Fagherazzi. Recovery of salt marsh vegetation after ice-rafting. Marine Ecology Progress Series, 710: 57–70, May 4, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/710/m710p057.pdf.

Norderhaug:2020:ELE

[NFDF^{+}20] K. M. Norderhaug, K. Filbee-Dexter, C. Freitas, S. R. Birkely, L. Christensen, I. Mellerud, J. Thormar, T. van Son, F. Moy, M. Vázquez Alonso, and H. Steen. Ecosystem-level effects of large-scale disturbance in kelp forests. Marine Ecology Progress Series, 656:163–180, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m656p163.pdf.

Navarro-Garcia:2023:ADT

[NGPGO^{+}23] M. Navarro-García, D. Precioso, K. Gavira-O’Neill, A. Torres-Barrán, D. Gordo Gómez, V. Gallego, and D. Gómez-Ullate. Aggregation dynamics of tropical tunas around drifting floating objects based on large-scale echo-sounder data. Marine Ecology Progress Series, 715:129–143, July 27, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/715/
m715p129.pdf.

Nielsen:2020:BAT

[NGRH20] B. L. H. Nielsen, H. V. S. Gréve, T. A. Rayner, and B. W. Hansen. Biochemical adaptation by the tropical copepods Apocyclops royi and Pseudodiaptomus annandalei to a PUFApoor brackish water habitat. Marine Ecology Progress Series, 655:77–89, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/655/m655p077.pdf.

Nicolle:2022:LTI

[NHFS22] P. Nicolle, J. Hughes, A. Fowler, and H. T. Schilling. Longterm increase in growth of an estuarine predator, mulloway Argyrosomus japonicus, predicted to continue under future warming scenarios. Marine Ecology Progress Series, 688:1–17, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Nguyen:2021:TVF

[NHH^{+}21] H. Nguyen, T. Hoang, D. Hawkins, B. J. Allen, and B. Pernet. Temporal variation in food limitation in larvae of the sand dollar Dendraster excentricus. Marine Ecology Progress Series, 665:127–143, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/665/m665p127.pdf.

Nascimento:2024:PCS

[NJHN24] L. S. REVIEW Nascimento, M. Nogueira Júnior, C. S. Hara, and M. Almeida Noernberg. Passive citizen science: social media as a tool for marine wildlife observation. Marine Ecology Progress Series, 740:219–233, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/740/
m740p219.pdf.

Ng:2022:VGJ

[NM22] C. A. Ng and F. Micheli. Variability in grazing on juvenile giant kelp throughout an upwelling season. Marine Ecology Progress Series, 693:83–93, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/693/m693p083.pdf.

Navarro-Martinez:2022:TFA

[NMAE^{+}22] Z. M. Navarro-Martínez, M. Armenteros, L. Espinosa, J. J. Lake, and A. Apprill. Taxonomic and functional assemblage structure of coral reef fishes from Jardines de la Reina (Caribbean Sea, Cuba). Marine Ecology Progress Series, 690:113–132, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/690/m690p113.pdf.

Norrie:2022:FGC

[NMB^{+}22] C. R. Norrie, C. A. Morgan, B. J. Burke, L. A. Weitkamp, and J. A. Miller. Freshwater growth can provide a survival advantage to Interior Columbia River spring Chinook salmon after ocean entry. Marine Ecology Progress Series, 691:131–149, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m691p131.pdf.

Nilsson:2022:ELC

[NP22] P. Nilsson and B. Pernet. Echinoid larvae can express foodconditioned morphological plasticity at ecologically relevant culture densities. Marine Ecology Progress Series, 694:1–12, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Nielsen:2021:STP

[NRE21] K. A. Nielsen, J. R. Robbins, and C. B. Embling. Spatiotemporal patterns in harbour porpoise density: citizen science and conservation in UK seas. Marine Ecology Progress Series, 675:165–180, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/675/m675p165.pdf.

Nordli:2023:BHU

[NSB^{+}23] E. Nordli, J. F. Strøm, T. Bøhn, E. B. Thorstad, R. M. Serra-Llinares, R. Nilsen, and P. A. Bjørn. Behaviour and habitat use of first-time migrant Arctic charr: novel insights from a subarctic marine area. Marine Ecology Progress Series, 709:77–90, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m709p077.pdf.

Ni:2023:UFN

[NSKH23] Y. Ni, L. K. Sandal, S. F. Kvamsdal, and C. Hansen. Using feedforward neural networks to represent ecosystem dynamics for bioeconomic analysis. Marine Ecology Progress Series, 716: 1–15, August 10, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Nakajima:2024:TSD

[NSM24] Y. Nakajima, S. Suzuki, and S. Mitarai. Time-series dynamics in symbiont communities of acroporid corals undergoing cold stress in winter. Marine Ecology Progress Series, 749:71– 85, November 21, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/749/m749p071.pdf.

Narvaez:2020:ICS

[NSMJ20] C. A. Narvaez, B. Sainte-Marie, and L. E. Johnson. Intraspecific competition in size-structured populations: ontogenetic shift in the importance of interference competition in a key marine herbivore. Marine Ecology Progress Series, 649:97–110, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/649/m649p097.pdf.

Niella:2020:PCD

[NSPH20] Y. Niella, A. F. Smoothey, V. Peddemors, and R. Harcourt. Predicting changes in distribution of a large coastal shark in the face of the strengthening East Australian current. Marine Ecology Progress Series, 642:163–177, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/642/
m642p163.pdf.

Nancollas:2024:URL

[NT24] S. J. Nancollas and A. E. Todgham. Understanding the role of low tide habitat, thermal predictability, and food availability in shaping the thermal performance of the California mussel. Marine Ecology Progress Series, 747:61–81, October 17, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/747/m747p061.pdf.

Nishikawa:2024:LTC

[NTN^{+}24] H. Nishikawa, H. Tsujino, S. Nishikawa, S. I. Nakayama, H. Nakano, T. Sugiyama, and Y. Ishikawa. Long-term comparison between the Japanese sardine Sardinops melanostictus stock level and simulated zooplankton density around the Kuroshio region. Marine Ecology Progress Series, 741:271– 287, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m741p271.pdf.

Ning:2022:ENE

[NYW^{+}22] Z. Ning, K. Yu, Y. Wang, F. Wei, Z. Liao, B. Yang, C. Fang, R. Xia, X. Huang, and G. Song. Effects of nutrient enrichment and skewed N:P ratios on physiology of scleractinian corals from Weizhou Island in the northern South China Sea. Marine Ecology Progress Series, 682:111–122, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/682/
m682p111.pdf.

Oswald:2024:OPC

[OA24] S. A. Oswald and J. M. Arnold. Opinion piece: Challenges of quantifying direct heat stress effects of climate change on seabirds. Marine Ecology Progress Series, 737:

25–29, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/737/m737p025.pdf.

Ottmann:2021:APN

[OÁBP^{+}21] D. Ottmann, D. Álvarez-Berastegui, L. Prieto, R. Balbín, F. Alemany, Ø. Fiksen, A. Gordoa, and P. Reglero. Abundance of Pelagia noctiluca early life stages in the western Mediterranean Sea scales with surface chlorophyll. Marine Ecology Progress Series, 658:75–88, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/658/
m658p075.pdf.

Ospina-Alvarez:2020:MND

[OAdJA^{+}20] A. Ospina-Alvarez, S. de Juan, J. Alós, G. Basterretxea, A. Alonso-Fernández, G. Follana-Berná, M. Palmer, and I. A. Catalán. MPA network design based on graph theory and emergent properties of larval dispersal. Marine Ecology Progress Series, 650:309–326, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/650/m650p309.pdf.

Onitsuka:2024:VDP

[OAM^{+}24] G. Onitsuka, K. Abo, T. Matsubara, K. i. Mizuno, S. i. Ikeda, T. Sato, T. Shikata, T. Onduka, and M. Hamaguchi. Vertical distribution of Pacific oyster Crassostrea gigas larvae and modeling larval transport in Hiroshima Bay, Japan. Marine Ecology Progress Series, 740:43–60, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/740/
m740p043.pdf.

Olnes:2021:JBS

[OBD^{+}21] J. Olnes, G. A. Breed, M. L. Druckenmiller, J. J. Citta, J. A. Crawford, A. L. Von Duyke, and L. Quakenbush. Juvenile bearded seal response to a decade of sea ice change in the Bering, Chukchi, and Beaufort seas. Marine Ecology Progress Series, 661:229–242, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2021/661/m661p229.pdf.

Obaza:2022:VFH

[OBS^{+}22] A. K. Obaza, A. Bird, R. Sanders, R. Ware, and D. W. Ginsburg. Variable fish habitat function in two open-coast eelgrass species. Marine Ecology Progress Series, 696:15–27, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Olin:2020:SSB

[OBW^{+}20] A. B. Olin, N. S. Banas, P. J. Wright, M. R. Heath, and R. G. Nager. Spatial synchrony of breeding success in the blacklegged kittiwake Rissa tridactyla reflects the spatial dynamics of its sandeel prey. Marine Ecology Progress Series, 638:177–190, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m638p177.pdf.

Ogilvy:2023:SVI

[OCC23] C. Ogilvy, R. Constantine, and E. L. Carroll. Spatial variation in isotope values of Hector’s dolphins from the north coast of the South Island, New Zealand. Marine Ecology Progress Series, 711:61–75, May 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m711p061.pdf.

Olinger:2021:TCT

[OCFEB21] L. K. Olinger, A. Chaves-Fonnegra, I. C. Enochs, and M. E. Brandt. Three competitors in three dimensions: photogrammetry reveals rapid overgrowth of coral during multispecies competition with sponges and algae. Marine Ecology Progress Series, 657:109–121, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/657/m657p109.pdf.

Ouled-Cheikh:2022:TIS

[OCGAP^{+}22] J. Ouled-Cheikh, J. Giménez, M. Albo-Puigserver, J. Navarro, E. Fernández-Corredor, J. M. Bellido, M. G. Pennino, and M. Coll. Trophic importance of small pelagic fish to marine predators of the Mediterranean Sea. Marine Ecology Progress Series, 696:169–184, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2022/696/m696p169.pdf.

Ouellet:2024:DBI

[OCR^{+}24] J. F. Ouellet, J. Cabrol, È. Rioux, X. Bordeleau, and V. Lesage. Dealing with biases introduced by lipids in stable carbon and nitrogen isotope analyses: a solution based on 28 marine invertebrate, fish, and mammal species. Marine Ecology Progress Series, 738:75–87, June 20, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m738p075.pdf.

Olin:2024:BFR

[ODB^{+}24] A. B. Olin, L. Dück, P. A. Berglund, E. Karlsson, M. Bohm, O. Olsson, and J. Hentati-Sundberg. Breeding failures and reduced nest attendance in response to heat stress in a highlatitude seabird. Marine Ecology Progress Series, 737:147– 160, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m737p147.pdf.

Olson:2020:TPS

[OFSJ20] A. M. Olson, A. Frid, J. B. Q. Santos, and F. Juanes. Trophic position scales positively with body size within but not among four species of rocky reef predators. Marine Ecology Progress Series, 640:189–200, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/640/m640p189.pdf.

Orrell:2022:UVP

[OH22] D. L. Orrell and N. E. Hussey. Using the VEMCO Positioning System (VPS) to explore fine-scale movements of aquatic species: applications, analytical approaches and future directions. Marine Ecology Progress Series, 687:195– 216, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/687/m687p195.pdf.

Orostica:2020:PWW

[OHBJ20] M. H. Oróstica, S. J. Hawkins, B. R. Broitman, and S. R. Jenkins. Performance of a warm-water limpet species towards its poleward range edge compared to a colder-water congener. Marine Ecology Progress Series, 656:207–225, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/656/
m656p207.pdf.

Oxley:2024:MTA

[OJ24] C. C. Oxley and L. J. Jurgens. Material type affects the community composition and abundance of hard-substrate assemblages in a sedimentary Atlantic estuary. Marine Ecology Progress Series, 727:35–47, January 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/727/
m727p035.pdf.

Orphanides:2023:RMM

[OJPC23] C. D. Orphanides, J. M. Jech, D. L. Palka, and J. Collie. Relating marine mammal distribution to water column prey structure derived from echosounding. Marine Ecology Progress Series, 711:101–119, May 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m711p101.pdf.

Ogburn:2024:ATC

[OLA^{+}24] M. B. Ogburn, H. D. Legett, R. Aguilar, N. B. Furey, K. Heggie, M. C. Livernois, and K. D. Richie. Acoustic telemetry captures the full annual migration of alewife between Chesapeake Bay and the Gulf of Maine. Marine Ecology Progress Series, 745:115–124, September 19, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/745/m745p115.pdf.

Olson:2023:FSS

[OLG^{+}23] J. C. Olson, J. S. Lefcheck, M. R. Goodison, A. Lienesch, and M. B. Ogburn. Fish size spectra from imaging sonar reveal variation in habitat use across nearshore coastal ecosystems. Marine Ecology Progress Series, 705:95–108, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/705/
m705p095.pdf.

Oke:2024:STA

[OLM24] K. B. Oke, M. A. Litzow, and F. Mueter. Shifting temperature–abundance relationship for Bering Sea walleye pollock consistent with northward expansion during exceptionally warm conditions. Marine Ecology Progress Series, 749:141–158, November 21, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m749p141.pdf.

Olsen:2022:CFA

[OLPGK22] A. Y. Olsen, S. Larson, J. L. Padilla-Gamiño, and T. Klinger. Changes in fish assemblages after Marine Heatwave Events in West Hawai‘i Island. Marine Ecology Progress Series, 698: 95–109, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m698p095.pdf.

OShea:2020:NID

[OMW^{+}20] O. R. O’Shea, M. H. Meadows, E. E. Wrigglesworth, J. Newton, and L. A. Hawkes. Novel insights into the diet of southern stingrays and Caribbean whiptail rays. Marine Ecology Progress Series, 655:157–170, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/655/m655p157.pdf.

Osborne:2020:BSI

[OOW^{+}20] O. E. Osborne, P. D. O’Hara, S. Whelan, P. Zandbergen, S. A. Hatch, and K. H. Elliott. Breeding seabirds increase foraging range in response to an extreme marine heatwave. Marine Ecology Progress Series, 646:161–173, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m646p161.pdf.

Oosthuizen:2021:HBD

[ORB^{+}21] W. C. Oosthuizen, R. R. Reisinger, M. N. Bester, D. Steinhage, H. Auel, H. Flores, R. Knust, S. Ryan, and H. Bornemann. Habitat-based density models of pack-ice seal distribution in the southern Weddell Sea, Antarctica. Marine Ecology Progress Series, 673:211–227, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/673/m673p211.pdf.

Orenstein:2022:DDC

[OSBA22] E. C. Orenstein, E. Saberski, and C. Briseño-Avena. Discovery and dynamics of a cryptic marine copepod-parasite interaction. Marine Ecology Progress Series, 691:29–40, 2022.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m691p029.pdf.

Orkney:2022:AIP

[OSJ^{+}22] A. Orkney, S. Sathyendranath, T. Jackson, M. Porter, and H. A. Bouman. Atlantic inflow is the primary driver of remotely sensed autumn blooms in the Barents Sea. Marine Ecology Progress Series, 701:25–40, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/701/
m701p025.pdf.

Owen:2024:SOH

[OTD^{+}24] K. Owen, R. M. Thompson, D. Donnelly, M. Noad, S. J. Bury, M. H. Pinkerton, and R. Dunlop. Southern Ocean humpback whale trophic ecology. II. Influence of fasting and opportunistic feeding on skin stable isotope values of migrating whales. Marine Ecology Progress Series, 734:157– 171, April 18, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/734/m734p157.pdf.

Olive:2021:PPC

[OVÁS^{+}21] I. Olivé, E. Varela-Álvarez, J. Silva, E. A. Serrão, and R. Santos. Physiological potential of the chlorophyte Caulerpa prolifera for proliferation across the Mediterranean-Atlantic basins in a warmer ocean. Marine Ecology Progress Series, 668: 73–84, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/668/m668p073.pdf.

OHanlon:2024:APD

[OvBS^{+}24] N. J. O’Hanlon, R. S. A. van Bemmelen, K. R. S. Snell, G. J. Conway, C. B. Thaxter, H. Aiton, D. Aiton, D. E. Balmer, S. A. Hanssen, J. R. Calladine, S. Hammer, S. J. Harris, B. Moe, H. Schekkerman, I. Tulp, and E. M. Humphreys. Atlantic populations of a declining oceanic seabird have complex migrations and weak migratory connectivity to staging areas. Marine Ecology Progress Series, 730:113–129, March 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/730/m730p113.pdf.

OShea:2021:EDN

[OVO^{+}21] O. R. O’Shea, T. E. Van Leeuwen, D. A. O’Brien, L. Arrowsmith, R. McCalman, M. Griffiths, and D. A. Exton. Evidence and description of a nursery habitat for the recently reclassified stingray Styracura schmardae from The Bahamas. Marine Ecology Progress Series, 660:141–151, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/660/
m660p141.pdf.

Orrell:2023:SFD

[OWH23] D. L. REVIEW Orrell, D. Webber, and N. E. Hussey. A standardised framework for the design and application of finescale acoustic tracking studies in aquatic environments. Marine Ecology Progress Series, 706:125–151, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/706/
m706p125.pdf.

Onishi:2020:RBG

[OYO^{+}20] Y. Onishi, T. Yamanaka, K. I. Ozaki, R. Nakayama, S. Shimamura, R. Itami, A. Fukushima, M. Miyamoto, and Y. Fujiwara. Relationship between geochemical environments, nutritional resources, and faunal succession in whale-fall ecosystems. Marine Ecology Progress Series, 636:35–46, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/636/m636p035.pdf.

Puebla-Aparicio:2024:CFA

[PAAEV^{+}24] M. Puebla-Aparicio, C. Ascencio-Elizondo, M. Vieira, M. C. P. Amorim, R. Duarte, and P. J. Fonseca. Characterization of the fish acoustic communities in a Mozambican tropical coral reef. Marine Ecology Progress Series, 727:143– 158, January 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/727/m727p143.pdf.

Pawlik:2022:OPC

[PAF^{+}22] J. R. Pawlik, R. A. Armstrong, S. Farrington, J. Reed, S. Rivero-Calle, H. Singh, B. K. Walker, and J. White. Opinion piece: Comparison of recent survey techniques for estimating benthic cover on Caribbean mesophotic reefs. Marine Ecology Progress Series, 686:201–211, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m686p201.pdf.

Palmer:2022:PPA

[PAL^{+}22] E. Palmer, A. Alexander, L. Liggins, M. Guerra, S. J. Bury, H. Hendriks, K. A. Stockin, and K. J. Peters. A piece of the puzzle: analyses of recent strandings and historical records reveal new genetic and ecological insights on New Zealand sperm whales. Marine Ecology Progress Series, 690:201–217, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m690p201.pdf.

Priester:2024:COS

[PAT^{+}24] C. R. Priester, P. Afonso, C. N. Trueman, G. Menezes, G. Graça, and J. Fontes. Contrasting ontogenetic shifts in habitat and metabolism of three sympatric key deep-sea fishes. Marine Ecology Progress Series, 729:185–199, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m729p185.pdf.

Paitach:2023:CEF

[PBAC23] R. L. Paitach, G. A. Bortolotto, M. Amundin, and M. J. Cremer. Critically endangered franciscana dolphins in an estuarine area: fine-scale habitat use and distribution from acoustic monitoring in Babitonga Bay, southern Brazil. Marine Ecology Progress Series, 707:131–150, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/707/
m707p131.pdf.

Peters:2020:FEC

[PBB^{+}20] K. J. Peters, S. J. Bury, E. L. Betty, G. J. Parra, G. Tezanos-Pinto, and K. A. Stockin. Foraging ecology of the common dolphin Delphinus delphis revealed by stable isotope analysis. Marine Ecology Progress Series, 652:173–186, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/652/
m652p173.pdf.

Porter:2022:ETR

[PBC^{+}22] E. T. Porter, S. Blickenstaff, J. C. Cornwell, M. Jackson, and S. N. Tolbert. Effect of tidal resuspension with oyster biodeposits on nutrient and oxygen dynamics. Marine Ecology Progress Series, 686:37–60, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/686/m686p037.pdf.

Peterson:2022:PSE

[PBKG22] C. T. Peterson, B. A. Bachman, R. T. Kraus, and R. D. Grubbs. Phoretic sharksuckers Echeneis naucrates associated with an elasmobranch host occupy higher relative trophic positions. Marine Ecology Progress Series, 687:125– 132, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/687/m687p125.pdf.

Petrolo:2021:SSE

[PBO^{+}21] E. Petrolo, J. Boomer, J. O’Hare, K. Bilgmann, and A. Stow. Stock structure and effective population size of the commercially exploited gummy shark Mustelus antarcticus. Marine Ecology Progress Series, 678:109–124, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/678/

m678p109.pdf. See corrigendum [Ano21f].

Plumlee:2023:TPA

[PBRF23] J. D. Plumlee, C. Branham, S. J. Ryburn, and F. J. Fodrie. Trophic partitioning among seasonally resident predators in a temperate estuary. Marine Ecology Progress Series, 721:119– 133, October 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/721/m721p119.pdf.

Panicker:2022:FSS

[PBS22] D. Panicker, M. F. Baumgartner, and K. M. Stafford. Finescale spatial and temporal acoustic occurrence of islandassociated odontocetes near a mid-oceanic atoll in the northern Indian Ocean. Marine Ecology Progress Series, 683:195– 208, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m683p195.pdf.

Piotto:2024:SAC

[PBS^{+}24] M. Piotto, I. Barberá, M. Sironi, V. J. Rowntree, M. M. Uhart, M. Agrelo, A. A. Fernández Ajó, J. Seger, and C. F. Marón. Seabird attacks contribute to calf mortality in a whale population. Marine Ecology Progress Series, 746:1–16, October 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Puckeridge:2021:FBM

[PBT^{+}21] A. C. Puckeridge, A. Becker, M. D. Taylor, M. B. Lowry, J. McLeod, H. T. Schilling, and I. M. Suthers. Foraging behaviour and movements of an ambush predator reveal benthopelagic coupling on artificial reefs. Marine Ecology Progress Series, 666:171–182, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/666/m666p171.pdf.

Paradinas:2020:ASP

[PCLQ^{+}20] I. Paradinas, D. Conesa, A. López-Quílez, A. Esteban, L. M. Martín López, J. M. Bellido, and M. G. Pennino. Assessing the spatiotemporal persistence of fish distributions: a case study on two red mullet species (Mullus surmuletus and M. barbatus) in the western Mediterranean. Marine Ecology Progress Series, 644:173–185, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/644/m644p173.pdf.

Puk:2020:SSE

[PCM^{+}20] L. D. Puk, N. Cernohorsky, A. Marshell, J. Dwyer, K. Wolfe, and P. J. Mumby. Species-specific effects of herbivorous fishes on the establishment of the macroalga Lobophora on coral reefs. Marine Ecology Progress Series, 637:1–14, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Planas:2020:EDB

[PCP^{+}20] M. Planas, A. Chamorro, A. Paltrinieri, S. Campos, A. Jiménez, K. Nedelec, and J. Hernández-Urcera. Effect of diet on breeders and inheritance in syngnathids: application of isotopic experimentally derived data to field studies. Marine Ecology Progress Series, 650:107–123, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/650/
m650p107.pdf.

Pierrejean:2023:SEE

[PCR23] M. Pierrejean, M. Cusson, and F. Rossi. Seasonal effects of edge and habitat complexity on eelgrass epifaunal assemblages. Marine Ecology Progress Series, 718:39–52, September 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/718/m718p039.pdf.

Pontbriand:2023:DIB

[PDF^{+}23] T. Pontbriand, G. K. Davoren, S. M. E. Fortune, C. Pomerleau, B. G. Young, and S. H. Ferguson. Differences in individual bowhead whale Balaena mysticetus habitat use, foraging dive depth and diet during the peak feeding season. Marine Ecology Progress Series, 706:109–123, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m706p109.pdf.

Penalver:2020:MSR

[PDFH20] M. M. Peñalver, M. J. Durako, B. T. Furman, and M. O. Hall. Multiple stressors result in reduced reproductive effort by Thalassia testudinum in Florida Bay, USA. Marine Ecology Progress Series, 647:65–78, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/647/m647p065.pdf.

Priester:2024:TEC

[PDH^{+}24] C. R. Priester, J. Dierking, T. Hansen, D. Abecasis, J. M. Fontes, and P. Afonso. Trophic ecology and coastal niche partitioning of two sympatric shark species in the Azores (mid-Atlantic). Marine Ecology Progress Series, 726:113– 130, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m726p113.pdf.

Pratt:2022:OWM

[PDM22] C. Pratt, D. Denley, and A. Metaxas. Ocean warming and multiple source populations increase the threat of an invasive bryozoan to kelp beds in the northwest Atlantic Ocean. Marine Ecology Progress Series, 695:65–81, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/695/
m695p065.pdf.

Pages-Escola:2020:PDE

[PEBG^{+}20] M. Pagès-Escolà, P. E. Bock, D. P. Gordon, S. Wilson, C. Linares, B. Hereu, and M. J. Costello. Progress in the discovery of extant and fossil bryozoans. Marine Ecology Progress Series, 635:71–79, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/635/m635p071.pdf.

Pedersen:2021:CSP

[PFDF^{+}21] M. F. Pedersen, K. Filbee-Dexter, N. L. Frisk, Z. Sárossy, and T. Wernberg. Carbon sequestration potential increased by incomplete anaerobic decomposition of kelp detritus. Marine Ecology Progress Series, 660:53–67, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m660p053.pdf.

Peres:2021:SBD

[PFM^{+}21] P. A. Peres, A. P. Ferreira, G. B. O. Machado, M. Azevedo-Silva, S. G. L. Siqueira, and F. P. P. Leite. Sex-biased dispersal depends on the spatial scale in a tube-building amphipod. Marine Ecology Progress Series, 658:135–148, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/658/
m658p135.pdf.

Pamungkas:2021:BPW

[PGC21] J. Pamungkas, C. J. Glasby, and M. J. Costello. Biogeography of polychaete worms (Annelida) of the world. Marine Ecology Progress Series, 657:147–159, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/657/m657p147.pdf.

Patterson:2021:NRS

[PGGE21] A. Patterson, H. G. Gilchrist, A. Gaston, and K. H. Elliott. Northwest range shifts and shorter wintering period of an Arctic seabird in response to four decades of changing ocean climate. Marine Ecology Progress Series, 679:163–179, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m679p163.pdf.

Page:2023:VEL

[PGM^{+}23] C. A. Page, C. Giuliano, K. Meehan, R. Fisher, C. A. Motti, A. P. Negri, and C. J. Randall. Varied effects of Lobophora chemistry on settlement of larvae from five coral genera. Marine Ecology Progress Series, 717:17–35, August 24, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/717/m717p017.pdf.

Ponchon:2021:SSB

[PGT^{+}21] A. Ponchon, A. Gamble, J. Tornos, K. Delord, C. Barbraud, J. M. J. Travis, H. Weimerskirch, and T. Boulinier. Similar at-sea behaviour but different habitat use between failed and successful breeding albatrosses. Marine Ecology Progress Series, 678:183–196, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m678p183.pdf.

Pugh:2023:SDF

[PGW23] T. L. Pugh, R. P. Glenn, and W. H. Watson III. Spatial differences in female lobster mating success and population characteristics in southern Massachusetts. Marine Ecology Progress Series, 718:69–84, September 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/718/
m718p069.pdf.

Price:2020:CVB

[PHE^{+}20] C. A. Price, K. Hartmann, T. J. Emery, E. J. Woehler, C. R. McMahon, and M. A. Hindell. Climate variability and breeding parameters of a transhemispheric migratory seabird over seven decades. Marine Ecology Progress Series, 642:191–205, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/642/m642p191.pdf.

Piltz:2020:EBM

[PHV20] S. H. Piltz, P. G. Hjorth, and Ø. Varpe. Empirically based minimalistic model for representing seasonal phytoplankton dynamics. Marine Ecology Progress Series, 640:

63–77, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/640/m640p063.pdf.

Post:2021:BWM

[PJA^{+}21] S. Post, S. H. Jónasdóttir, H. Andreasen, A. H. Ólafsdóttir, and T. Jansen. Blue whiting Micromesistius poutassou diel feeding behaviour in the Irminger Sea. Marine Ecology Progress Series, 678:1–16, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Paffenhofer:2024:QMM

[PK24] G. A. Paffenhöfer and M. Köster. Quantification of marine metazooplankton: Do we need to include juveniles? Marine Ecology Progress Series, 739:269–274, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/739/
m739p269.pdf.

Piovano:2020:DRG

[PLC^{+}20] S. Piovano, G. E. Lemons, A. Ciriyawa, A. Batibasaga, and J. A. Seminoff. Diet and recruitment of green turtles in Fiji, South Pacific, inferred from in-water capture and stable isotope analysis. Marine Ecology Progress Series, 640: 201–213, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/640/m640p201.pdf.

Pang:2023:TLS

[PLG^{+}23] M. Pang, K. Liu, Z. Gao, C. K. Kang, and H. Liu. Temperature-light shapes the nutritional strategy of a mixotrophic green alga, Picochlorum sp. GLMF1 (Trebouxiophyceae). Marine Ecology Progress Series, 713:39–53, June 29, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2023/713/m713p039.pdf.

Pastor:2021:ABM

[PLH^{+}21] A. Pastor, J. Larsen, F. T. Hansen, A. Simon, N. Bierne, and M. Maar. Agent-based modeling and genetics reveal the Limfjorden, Denmark, as a well-connected system for mussel larvae. Marine Ecology Progress Series, 680:193–205, 2021.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m680p193.pdf.

Pacheco:2021:MBH

[PLLT^{+}21] A. S. Pacheco, M. A. Llapapasca, N. L. López-Tejada, S. Silva, and B. Alcorta. Modeling breeding habitats of humpback whales Megaptera novaeangliae as a function of group composition. Marine Ecology Progress Series, 666: 203–215, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/666/m666p203.pdf.

Pan:2023:WFV

[PMH23] F. T. C. NOTE Pan, D. T. Manahan, and D. Hedgecock. Within-family variation in larval viability and growth is controlled by different genes: a case study with Crassostrea gigas. Marine Ecology Progress Series, 704:149–153, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/704/
m704p149.pdf.

Politikos:2021:CVA

[PRC^{+}21] D. V. Politikos, K. A. Rose, E. N. Curchitser, D. M. Checkley, Jr., R. R. Rykaczewski, and J. Fiechter. Climate variation and anchovy recruitment in the California Current: a cause-and-effect analysis of an end-to-end model simulation. Marine Ecology Progress Series, 680:111–136, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/680/
m680p111.pdf.

Patel:2023:SDO

[PRCvdK23] R. Patel, S. Roy, E. Capuzzo, and J. van der Kooij. Seasonality of diet overlap among small pelagic fish in the waters southwest of the UK. Marine Ecology Progress Series, 708:101–123, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/708/m708p101.pdf.

Porter:2020:ERE

[PRD^{+}20] E. T. Porter, E. Robins, S. Davis, R. Lacouture, and J. C. Cornwell. Effects of resuspension of eastern oyster Crassostrea virginica biodeposits on phytoplankton community structure. Marine Ecology Progress Series, 640:79–105, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/640/
m640p079.pdf.

Pereira:2021:LSO

[PRM^{+}21] J. M. Pereira, J. A. Ramos, A. M. Marques, F. R. Ceia, L. Krüger, S. C. Votier, and V. H. Paiva. Low spatial overlap between foraging shearwaters during the breeding season and industrial fisheries off the west coast of Portugal. Marine Ecology Progress Series, 657:209–221, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/657/
m657p209.pdf.

Pierucci:2023:ADF

[PS23] A. Pierucci and G. Suaria. Abundance and distribution of flying fishes (Exocoetidae) and flying squids (Ommastrephidae) in the Eastern Atlantic Ocean based on a large-scale visual survey. Marine Ecology Progress Series, 725:45–56, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m725p045.pdf.

Pedaccini:2023:SRF

[PSB^{+}23] M. Pedaccini, I. Sousa, L. Bentes, C. DeBenito-Abelló, P. G. Lino, E. Moland, J. M. S. Gonçalves, and B. Horta e Costa. Spawning-related fish movement synchrony of white seabream Diplodus sargus in a no-take zone. Marine Ecology Progress Series, 707:99–113, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/707/m707p099.pdf.

Parker:2021:SER

[PSBH21] J. R. C. Parker, B. J. Saunders, S. Bennett, and E. S. Harvey. Successful establishment of range-shifting, warm-water Labridae in temperate South Western Australia. Marine Ecology Progress Series, 667:161–175, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/667/m667p161.pdf.

Phillips:2020:BTA

[PSN^{+}20] N. D. Phillips, E. A. Elliott Smith, S. D. Newsome, J. D. R. Houghton, C. D. Carson, J. Alfaro-Shigueto, J. C. Mangel, L. E. Eagling, L. Kubicek, and C. Harrod. Bulk tissue and amino acid stable isotope analyses reveal global ontogenetic patterns in ocean sunfish trophic ecology and habitat use. Marine Ecology Progress Series, 633:127–140, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/633/
m633p127.pdf.

Pampoulie:2024:DPA

[PSÓ^{+}24] C. Pampoulie, A. Slotte, G. J. Óskarsson, G. Ólafsdóttir, J. A. Jacobsen, H. Joensen, S. K. Sigurdsson, S. Sveinsson, L. Andersson, A. K. Daníelsdóttir, and D. Gíslason. Discriminating populations of Atlantic herring mixing in the Norwegian Sea feeding ground using single nucleotide polymorphisms (SNPs). Marine Ecology Progress Series, 739:227–240, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m739p227.pdf.

Purdon:2020:SAN

[PSP^{+}20] J. Purdon, F. Shabangu, M. Pienaar, M. J. Somers, and K. P. Findlay. South Africa’s newly approved marine protected areas have increased the protected modelled habitat of nine odontocete species. Marine Ecology Progress Series, 633:1–21, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Pickens:2022:OSS

[PTCD22] B. A. Pickens, J. C. Taylor, M. D. Campbell, and W. B. Driggers III. Offshore snapper and shark distributions are predicted by prey and area of nearby estuarine environments in the Gulf of Mexico, USA. Marine Ecology Progress Series, 682:169–189, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m682p169.pdf.

Papadakis:2021:NRO

[PTLK21] O. Papadakis, K. Tsirintanis, V. Lioupa, and S. Katsanevakis. The neglected role of omnivore fish in the overgrazing of Mediterranean rocky reefs. Marine Ecology Progress Series, 673:107–116, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/673/m673p107.pdf.

Pirotta:2020:SIS

[PVB^{+}20] E. Pirotta, M. Vighi, J. M. Brotons, E. Dillane, M. Cerdà, and L. Rendell. Stable isotopes suggest fine-scale sexual segregation in an isolated, endangered sperm whale population. Marine Ecology Progress Series, 654:209–218, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/654/
m654p209.pdf.

PaisdeFaria:2021:IMR

[PVL^{+}21] J. Pais de Faria, P. T. Vaz, C. S. Lopes, J. G. Calado, J. M. Pereira, S. N. Veríssimo, V. H. Paiva, A. M. M. Gonçalves, and J. A. Ramos. The importance of marine resources in the diet of urban gulls. Marine Ecology Progress Series, 660:189–201, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/660/m660p189.pdf. See corrigen-

dum [Ano21e].

Poiesz:2021:TSR

[PWvdM^{+}21] S. S. H. Poiesz, J. I. J. Witte, M. T. J. van der Meer, H. W. van der Veer, and K. E. R. Soetaert. Trophic structure and resource utilization of the coastal fish community in the western Wadden Sea: evidence from stable isotope data analysis. Marine Ecology Progress Series, 677:115–128, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/677/
m677p115.pdf.

Poiesz:2023:SCS

[PWvdM^{+}23] S. S. H. Poiesz, J. I. J. Witte, M. T. J. van der Meer, Z. Jager, K. E. R. Soetaert, T. van der Heide, and H. W. van der Veer. Stomach content and stable isotopes illustrate large spatial similarity in the Wadden Sea fish food-web structure. Marine Ecology Progress Series, 707:57–76, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/707/
m707p057.pdf.

Poiesz:2020:OFK

[PWvdV20] S. S. H. Poiesz, J. I. J. Witte, and H. W. van der Veer. Only a few key prey species fuel a temperate coastal fish food web. Marine Ecology Progress Series, 653:153–166, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/653/
m653p153.pdf.

Pittman:2021:SEI

[PYB^{+}21] S. J. Pittman, K. L. Yates, P. J. Bouchet, D. Alvarez-Berastegui, S. Andréfouët, S. S. Bell, C. Berkström, C. Boström, C. J. Brown, R. M. Connolly, R. Devillers, D. Eggleston, B. L. Gilby, M. Gullström, B. S. Halpern, M. Hidalgo, D. Holstein, K. Hovel, F. Huettmann, E. L. Jackson, W. R. James, J. B. Kellner, C. Y. Kot, V. Lecours, C. Lepczyk, I. Nagelkerken, J. Nelson, A. D. Olds, R. O. Santos, K. L. Scales, D. C. Schneider, H. T. Schilling, C. Simenstad, I. M. Suthers, E. A. Treml, L. M. Wedding, P. Yates, and M. Young. Seascape ecology: identifying research priorities for an emerging ocean sustainability science. Marine Ecology Progress Series, 663:1–29, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Quigley:2024:QFR

[QFT^{+}24] L. A. Quigley, P. J. S. Franks, A. R. Thompson, J. C. Field, and J. A. Santora. Quantifying the fates and retention of larval rockfish through Lagrangian analyses. Marine Ecology Progress Series, 749:109–125, November 21, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m749p109.pdf.

Queiros:2021:WCW

[QHT^{+}21] J. P. Queirós, A. Hilário, D. R. Thompson, F. R. Ceia, G. Elliott, K. Walker, Y. Cherel, and J. C. Xavier. From warm to cold waters: new insights into the habitat and trophic ecology of Southern Ocean squids throughout their life cycle. Marine Ecology Progress Series, 659:113–126, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/659/
m659p113.pdf.

Qui-Minet:2022:RCF

[QMDGM22] Z. N. Qui-Minet, D. Davoult, J. Grall, and S. Martin. The relative contribution of fleshy epiphytic macroalgae to the production of temperate maerl (rhodolith) beds. Marine Ecology Progress Series, 693:69–82, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/693/m693p069.pdf.

Quigley:2024:SLD

[QRC^{+}24] C. N. Quigley, M. Roughan, R. Chaput, A. G. Jeffs, and J. P. A. Gardner. Simulating larval dispersal across the distribution of the New Zealand green-lipped mussel: insights into connectivity and source-sink dynamics. Marine Ecology Progress Series, 731:129–145, March 13, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m731p129.pdf.

Reese:2023:ESL

[RAA^{+}23] T. C. Reese, J. Alder, E. G. Asay, A. M. H. Blakeslee, D. Cabrera, L. C. Crane, L. S. Fletcher, E. Pinkston, M. F. Repetto, N. Smith, C. Stancil, C. K. Tepolt, B. J. Toscano, and B. D. Griffen. Effects of season and latitude on the diet quality of the invasive Asian shore crab Hemigrapsus sanguineus. Marine Ecology Progress Series, 704:67–79, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m704p067.pdf.

Rupil:2022:RMK

[RAF^{+}22] G. M. Rupil, R. Angelini, J. L. R. Filho, J. Roman, and F. G. Daura-Jorge. The role of mammals as key predators in marine ecosystems. Marine Ecology Progress Series, 684:211–222, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/684/m684p211.pdf.

Ray:2021:CSC

[RAHM^{+}21] N. E. Ray, A. N. Al-Haj, T. J. Maguire, M. C. Henning, and R. W. Fulweiler. Coastal silicon cycling amplified by oyster aquaculture. Marine Ecology Progress Series, 673:29–41, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m673p029.pdf.

Rudershausen:2020:UIP

[RB20] P. J. Rudershausen and J. A. Buckel. Urbanization impacts on production and recruitment of Fundulus heteroclitus in salt marsh creeks. Marine Ecology Progress Series, 645:187–204, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/645/m645p187.pdf.

Redfern:2022:GAT

[RB22] C. P. F. Redfern and R. M. Bevan. The Arctic tern Sterna paradisaea: consistency and variability in spatial use at a global oceanographic scale. Marine Ecology Progress Series, 691:151–171, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/691/m691p151.pdf.

Rebstock:2023:SSM

[RB23] G. A. Rebstock and P. Dee Boersma. Sex-specific migratory behavior in a marine predator results in higher risks to females. Marine Ecology Progress Series, 725:141–156, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/725/m725p141.pdf.

Roman:2021:DTC

[RBB^{+}21] L. Roman, S. Bryan, N. Bool, L. Gustafson, and K. Townsend. Desperate times call for desperate measures: non-food ingestion by starving seabirds. Marine Ecology Progress Series, 662:157–168, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/662/m662p157.pdf.

Rempel:2024:EDP

[RBB^{+}24] H. S. Rempel, K. N. Bodwin, D. E. Burkepile, T. C. Adam, A. H. Altieri, E. M. Barton, R. L. Francisca, M. C. Goodman, R. J. Lamore, M. Lippert, M. Marroquín, T. C. O’Rourke, P. D. VanderBloomer, and B. I. Ruttenberg. Ecological drivers of parrotfish coral predation vary across spatial scales. Marine Ecology Progress Series, 740:145–160, July 25, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m740p145.pdf.

Rooper:2024:EFA

[RBC^{+}24] C. N. Rooper, J. L. Boldt, J. Cleary, M. A. Peña, M. Thompson, and M. Grinnell. Evaluating factors affecting the distribution and timing of Pacific herring Clupea pallasii spawn in British Columbia. Marine Ecology Progress Series, 741: 251–269, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m741p251.pdf.

Rodriguez:2022:VHI

[RBCAV22] S. M. Rodríguez, J. E. Byers, F. Cerda-Aliaga, and N. Valdivia. Variation in helminth infection prevalence, abundance, and co-infection in an intermediate host across a large spatial scale. Marine Ecology Progress Series, 681:103–116, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/681/m681p103.pdf.

Russell:2020:TSE

[RBCwD^{+}20] Bayden D. Russell, Lissandro Benedetti-Cecchi, Yun wei Dong, Laura J. Falkenberg, and Myron A. Peck. Theme section: The ecology of temperate reefs in a changing world. Marine Ecology Progress Series, 656:89–251, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m656p089.pdf.

Riccialdelli:2020:TSS

[RBF^{+}20] L. Riccialdelli, Y. A. Becker, N. E. Fioramonti, M. Torres, D. O. Bruno, A. Raya Rey, and D. A. Fernández. Trophic structure of southern marine ecosystems: a comparative isotopic analysis from the Beagle Channel to the oceanic Burdwood Bank area under a wasp-waist assumption. Marine Ecology Progress Series, 655:1–27, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Rohner:2022:VHU

[RBF^{+}22] C. A. Rohner, R. Bealey, B. M. Fulanda, C. E. M. Prebble, S. M. Williams, and S. J. Pierce. Vertical habitat use by black and striped marlin in the Western Indian Ocean.

Marine Ecology Progress Series, 690:165–183, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/690/
m690p165.pdf.

Rodrigues:2020:SPC

[RBFA20] G. F. B. Rodrigues, C. S. Ballarin, A. Fransozo, and F. W. Amorim. Structural patterns of a coastal hermit crabgastropod shell interaction network: new insights from a unique relationship. Marine Ecology Progress Series, 640: 117–126, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/640/m640p117.pdf.

Roberts:2021:WMC

[RBM^{+}21] E. M. Roberts, D. G. Bowers, H. K. Meyer, A. Samuelsen, H. T. Rapp, and P. Cárdenas. Water masses constrain the distribution of deep-sea sponges in the North Atlantic Ocean and Nordic Seas. Marine Ecology Progress Series, 659:75–96, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m659p075.pdf.

Ross:2022:EEF

[RBM22] E. P. Ross, J. Butler, and T. R. Matthews. Examining the ecological function of structure: species assemblages at casitas and coral heads in the Lower Florida Keys. Marine Ecology Progress Series, 681:169–183, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m681p169.pdf.

Robinson:2024:ROB

[RBS^{+}24] C. L. K. Robinson, D. F. Bertram, H. Shannon, V. R. von Biela, W. Greentree, W. Duguid, and M. L. Arimitsu. Reduction in overwinter body condition and size of Pacific sand lance has implications for piscivorous predators during marine heatwaves. Marine Ecology Progress Series, 737:89–99, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m737p089.pdf.

Rita:2023:AUI

[RBVA23] D. Rita, A. Borrell, G. Víkingsson, and A. Aguilar. Alkenone U_{37}^{k} index differs between thermally separated populations of fin whales and krill. Marine Ecology Progress Series, 725:157– 165, December 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/725/m725p157.pdf.

Riaz:2020:TBD

[RBW^{+}20] J. Riaz, S. Bestley, S. Wotherspoon, J. Freyer, and L. Emmerson. From trips to bouts to dives: temporal patterns in the diving behaviour of chick-rearing Adélie penguins, East Antarctica. Marine Ecology Progress Series, 654:177– 194, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/654/m654p177.pdf. See [Ano21h].

Roberts:2023:BTA

[RC23] E. A. Roberts and E. Carrington. Byssal thread attachment and growth are not correlated across gradients of temperature and food availability for two congeneric mussel species. Marine Ecology Progress Series, 704:35–54, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/704/
m704p035.pdf.

Robinson:2024:LSR

[RCB^{+}24a] B. Robinson, H. A. Coletti, B. Ballachey, J. L. Bodkin, K. Kloecker, S. B. Traiger, and D. Esler. Lack of strong responses to the Pacific marine heatwave by benthivorous marine birds indicates importance of trophic drivers. Marine Ecology Progress Series, 737:215–226, June 6, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m737p215.pdf.

Rossi:2024:APC

[RCB24b] S. P. Rossi, S. P. Cox, and H. P. Benoît. Absence of predator control increases cod extirpation risk in a Northwest Atlantic ecosystem: inference from multispecies modelling. Marine Ecology Progress Series, 746:99–119, October 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/746/m746p099.pdf.

Russell:2022:BCM

[RCC^{+}22] G. Russell, A. Colefax, F. Christiansen, G. Russell, Z. Fowler, and D. Cagnazzi. Body condition and migration timing of east Australian humpback whales. Marine Ecology Progress Series, 692:169–183, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2022/692/m692p169.pdf.

Ruzicka:2024:RRS

[RCC^{+}24] J. Ruzicka, L. Chiaverano, M. Coll, S. Garrido, J. Tam, H. Murase, K. Robinson, G. Romagnoni, L. Shannon, A. Silva, D. Szalaj, and S. Watari. Review: The role of small pelagic fish in diverse ecosystems: knowledge gleaned from foodweb models. Marine Ecology Progress Series, 741:7–27, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m741p007.pdf.

Roper:2022:CIN

[RCES22] C. D. Roper, E. F. Camp, J. Edmondson, and D. J. Suggett. Combined impacts of natural recruitment and active propagation for coral population recovery on the Great Barrier Reef. Marine Ecology Progress Series, 700:95–109, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/700/
m700p095.pdf.

Rayner:2021:SIR

[RDL^{+}21] M. J. Rayner, B. J. Dunphy, K. Lukies, N. Adams, M. Berg, L. Kozmian-Ledward, M. H. Pinkerton, and S. J. Bury. Stable isotope record from a resident New Zealand seabird community suggests changes in distribution but not trophic position since 1878. Marine Ecology Progress Series, 678:171– 182, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/678/m678p171.pdf.

Rondeau:2022:PDN

[RDL^{+}22] S. Rondeau, D. Davoult, C. Lejeusne, J. M. Kenworthy, O. Bohner, S. Loisel, and R. P. M. Gauff. Persistent dominance of non-indigenous species in the inner part of a marina highlighted by multi-year photographic monitoring. Marine Ecology Progress Series, 690:15–30, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/690/
m690p015.pdf.

Rippel:2023:DEN

[RDT^{+}23] T. M. Rippel, A. L. DeCandia, J. Tomasula, C. L. McIntosh, S. M. Murphy, and G. M. Wimp. Detritivores and exogenous nitrogen influence litter microbial communities in coastal salt marshes. Marine Ecology Progress Series, 716: 17–29, August 10, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/716/m716p017.pdf.

Randazzo-Eisemann:2022:ECA

[REGP22] Á. Randazzo-Eisemann and J. R. Garza-Pérez. The effects of coral assemblage shift on reef functions in Akumal, Mexico. Marine Ecology Progress Series, 695:53–63, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/695/
m695p053.pdf.

Ray:2020:SPB

[RF20] N. E. Ray and R. W. Fulweiler. Seasonal patterns of benthicpelagic coupling in oyster habitats. Marine Ecology Progress Series, 652:95–109, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m652p095.pdf.

Ransome:2024:EDP

[RFJC^{+}24] N. Ransome, A. Frisch-Jordán, T. Cheeseman, J. Calambokidis, A. Kew, O. Titova, O. Filatova, N. R. Loneragan, and J. N. Smith. Environmental drivers of persistent humpback whale Megaptera novaeangliae feeding events in a Mexican breeding area. Marine Ecology Progress Series, 726:161– 179, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/726/m726p161.pdf.

Rogers:2022:DTS

[RGK22] T. L. Rogers, T. C. Gouhier, and D. L. Kimbro. Distinct temperature stressors acting on multiple ontogenetic stages influence the biogeography of Atlantic blue crabs. Marine Ecology Progress Series, 690:97–111, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/690/m690p097.pdf.

Ralph:2024:ECB

[RGL24] G. M. Ralph, J. Gartland, and R. J. Latour. Evaluation of the Chesapeake Bay blue crab sanctuary through habitat suitability. Marine Ecology Progress Series, 739:111– 128, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/739/m739p111.pdf.

Rodriguez:2020:GTH

[RH20] A. R. Rodriguez and K. L. Heck, Jr. Green turtle herbivory and its effects on the warm, temperate seagrass meadows of St. Joseph Bay, Florida (USA). Marine Ecology Progress Series, 639:37–51, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/639/m639p037.pdf.

Reid:2021:LTE

[RH21] H. B. Reid and C. D. G. Harley. Low temperature exposure determines performance and thermal microhabitat use in an intertidal gastropod (Littorina scutulata) during the winter. Marine Ecology Progress Series, 660:105–118, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/660/
m660p105.pdf.

Reigel:2023:MES

[RH23] A. M. Reigel and M. E. Hellberg. Microbiome environmental shifts differ between two co-occurring octocoral hosts. Marine Ecology Progress Series, 720:59–83, October 5, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m720p059.pdf.

Ramsby:2020:DIN

[RHH^{+}20] B. D. Ramsby, J. Heishman, M. O. Hoogenboom, S. Whalan, and N. S. Webster. Dissolved inorganic nutrient enrichment does not affect sponge growth or condition. Marine Ecology Progress Series, 634:77–88, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/634/m634p077.pdf.

Rahnke:2022:OSR

[RHM^{+}22] S. A. Rahnke, J. R. Hancock, N. J. Munk, C. Caruso, and C. Drury. Optimizing sexual reproduction of Montipora capitata for restoration: effects of abiotic conditions and light acclimation on juvenile survival and growth. Marine Ecology Progress Series, 691:41–54, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m691p041.pdf.

Rangel:2020:NNS

[RHN^{+}20] B. S. Rangel, N. E. Hussey, Y. Niella, L. A. Martinelli, A. D. Gomes, and R. G. Moreira. Neonatal nutritional strategy of a viviparous elasmobranch with extremely low reproductive output. Marine Ecology Progress Series, 638:107– 121, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/638/m638p107.pdf.

Rose:2024:ABB

[RHN^{+}24] K. A. REVIEW Rose, K. Holsman, J. A. Nye, E. H. Markowitz, T. N. S. Banha, N. Bednaršek, J. Bueno-Pardo, D. Deslauriers, E. A. Fulton, K. B. Huebert, M. Huret, S. Ito, S. Koenigstein, L. Li, H. Moustahfid, B. A. Muhling, P. Neubauer, J. R. Paula, E. C. Siddon, M. D. Skogen, P. D. Spencer, P. D. van Denderen, G. I. van der Meeren, and M. A. Peck. Advancing bioenergetics-based modeling to improve climate change projections of marine ecosystems. Marine Ecology Progress Series, 732:193–221, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m732p193.pdf.

Rangel:2021:DRB

[RHSM21] B. S. Rangel, N. Hammerschlag, J. A. Sulikowski, and R. G. Moreira. Dietary and reproductive biomarkers in a generalist apex predator reveal differences in nutritional ecology across life stages. Marine Ecology Progress Series, 664:149– 163, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/664/m664p149.pdf.

Rankin:2021:NES

[RJ21] L. L. Rankin and H. P. Jones. Nearshore ecosystems on seabird islands are potentially influenced by invasive predator eradications and environmental conditions: a case study at the Mercury Islands, New Zealand. Marine Ecology Progress Series, 661:83–96, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/661/m661p083.pdf.

Rasmuson:2022:DMD

[RJE^{+}22] L. K. Rasmuson, T. Jackson, C. A. Edwards, K. G. O’Malley, and A. Shanks. A decade of modeled dispersal of Dungeness crab Cancer magister larvae in the California Current. Marine Ecology Progress Series, 686:127–140, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/686/
m686p127.pdf.

Rahlff:2021:STR

[RKV^{+}21] J. Rahlff, S. Khodami, L. Voskuhl, M. P. Humphreys, C. Stolle, P. Martinez Arbizu, O. Wurl, and M. Ribas-Ribas. Short-term responses to ocean acidification: effects on relative abundance of eukaryotic plankton from the tropical timor sea. Marine Ecology Progress Series, 658:59–74, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/658/
m658p059.pdf.

Riisgaard:2022:BMS

[RL22] H. U. Riisgård and P. S. Larsen. Bioenergetic model and specific growth rates of jellyfish Aurelia spp. Marine Ecology Progress Series, 688:49–56, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/688/m688p049.pdf.

Riisgaard:2024:ECS

[RL24] H. U. REVIEW Riisgård and P. S. Larsen. Energy costs of the sponge pump. Marine Ecology Progress Series, 746:141– 151, October 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/746/m746p141.pdf.

Reiertsen:2021:IPS

[RLME^{+}21] T. K. Reiertsen, K. Layton-Matthews, K. E. Erikstad, K. Hodges, M. Ballesteros, T. Anker-Nilssen, R. T. Barrett, S. Benjaminsen, M. Bogdanova, S. Christensen-Dalsgaard, F. Daunt, N. Dehnhard, M. P. Harris, M. Langset, S. H. Lorentsen, M. Newell, V. S. Bråthen, I. Støyle-Bringsvor, G. H. Systad, and S. Wanless. Inter-population synchrony in adult survival and effects of climate and extreme weather in non-breeding areas of Atlantic puffins. Marine Ecology Progress Series, 676:219–231, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m676p219.pdf.

Rinke:2023:CBR

[RLP^{+}23] M. Rinke, K. Lõhmus, D. Pieck, M. Maraun, and S. Scheu. Channelling of basal resources and use of allochthonous marine carbon by soil arthropods of the Wadden Sea salt marsh. Marine Ecology Progress Series, 706:57–71, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/706/
m706p057.pdf.

Rudershausen:2023:RAA

[RLSO23] P. J. Rudershausen, S. M. Lombardo, G. R. Stilson, and M. J. O’Donnell. Relating absolute abundance of an estuarine fish to habitat area in an urbanizing environment. Marine Ecology Progress Series, 719:77–92, September 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/719/
m719p077.pdf.

Rodriguez-Malagon:2020:III

[RMASA20] M. A. Rodríguez-Malagón, L. P. Angel, C. N. Speakman, and J. P. Y. Arnould. Inter- and intra-individual variation in the diet of Australasian gannets Morus serrator. Marine Ecology Progress Series, 636:207–220, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/636/m636p207.pdf.

Runnells:2024:SIN

[RMD24] E. S. Runnells, W. A. Montevecchi, and G. K. Davoren. Spatiotemporal and isotopic niche overlap among Atlantic puffins, razorbills, and common murres during the non-breeding season in the Northwest Atlantic. Marine Ecology Progress Series, 739:241–256, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/739/m739p241.pdf.

Raposo:2023:SDI

[RMR^{+}23] C. Raposo, J. Mestre, R. Rebelo, A. Regalla, A. Davies, C. Barbosa, and A. R. Patrício. Spatial distribution of internesting green turtles from the largest Eastern Atlantic rookery and overlap with a marine protected area. Marine Ecology Progress Series, 703:161–175, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/703/m703p161.pdf.

Randall:2020:SPC

[RNQ^{+}20] C. J. Randall, A. P. Negri, K. M. Quigley, T. Foster, G. F. Ricardo, N. S. Webster, L. K. Bay, P. L. Harrison, R. C. Babcock, and A. J. Heyward. Sexual production of corals for reef restoration in the Anthropocene. Marine Ecology Progress Series, 635:203–232, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m635p203.pdf.

Robertson:2021:LLC

[Rob21] A. I. Robertson. Leaf-litter consumption slows crab growth but transforms mangrove food chains. Marine Ecology Progress Series, 667:83–98, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/667/m667p083.pdf.

Ruesink:2022:TLH

[ROMB22] J. L. Ruesink, B. A. Briones Ortiz, C. H. Mawson, and F. C. Boardman. Tradeoffs in life history investment of eelgrass Zostera marina across estuarine intertidal conditions. Marine Ecology Progress Series, 686:61–70, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/686/
m686p061.pdf.

Ramirez:2021:GSS

[RPB21] M. D. Ramirez, T. Popovska, and E. A. Babcock. Global synthesis of sea turtle von Bertalanffy growth parameters through Bayesian hierarchical modeling. Marine Ecology Progress Series, 657:191–207, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m657p191.pdf.

Risoli:2023:TMH

[RPD^{+}23] M. C. Risoli, A. R. Piola, O. Defeo, D. Luzzatto, E. Celentano, and B. J. Lomovasky. Testing macroecological hypotheses in sandy beach populations: the wedge clam Donax hanleyanus in South America. Marine Ecology Progress Series, 707:43–56, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m707p043.pdf.

Ross:2023:ENA

[RRR^{+}23] C. H. Ross, J. A. Runge, J. J. Roberts, D. C. Brady, B. Tupper, and N. R. Record. Estimating North Atlantic right whale prey based on Calanus finmarchicus thresholds. Marine Ecology Progress Series, 703:1–16, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Ramos:2023:SCC

[RRS^{+}23] J. E. Ramos, Á. Roura, J. M. Strugnell, N. A. Moltschaniwskyj, R. Bargiela, and G. T. Pecl. Stomach content characterisation of the marine range-shifting Octopus tetricus using DNA metabarcoding. Marine Ecology Progress Series, 717: 67–83, August 24, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/717/m717p067.pdf.

Reustle:2020:TSI

[RS20a] J. W. Reustle and D. L. Smee. Turbidity and salinity influence trophic cascades on oyster reefs through modification of sensory performance and facilitation of different predator types. Marine Ecology Progress Series, 639:127–136, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/639/
m639p127.pdf.

Rub:2020:EDB

[RS20b] A. M. Wargo Rub and B. P. Sandford. Evidence of a ‘dinner bell’ effect from acoustic transmitters in adult Chinook salmon. Marine Ecology Progress Series, 641:1–11, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Renzi:2021:IGD

[RS21] J. J. Renzi and B. R. Silliman. Increasing grazer density leads to linear decreases in Spartina alterniflora biomass and exponential increases in grazing pressure across a barrier island. Marine Ecology Progress Series, 659:49–58, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/659/
m659p049.pdf.

Raymond:2021:SOE

[RSEG21] W. W. Raymond, J. B. Schram, G. L. Eckert, and A. W. E. Galloway. Sea otter effects on trophic structure of seagrass communities in southeast Alaska. Marine Ecology Progress Series, 674:37–58, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m674p037.pdf.

Rossi:2022:GGR

[RSL^{+}22] R. E. Rossi, C. A. Schutte, J. Logarbo, C. Bourgeois, and B. J. Roberts. Gulf ribbed mussels increase plant growth, primary production, and soil nitrogen cycling potential in salt marshes. Marine Ecology Progress Series, 689:33–46, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/689/
m689p033.pdf.

Raya:2020:BBM

[RSS20] V. Raya, J. Salat, and A. Sabatés. The box-balance model: a new tool to assess fish larval survival, applied to field data on two small pelagic fish. Marine Ecology Progress Series, 650:289–308, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/650/m650p289.pdf.

Rademaker:2021:CLB

[RSvL21] M. Rademaker, I. M. Smallegange, and A. van Leeuwen. Causal links between North Sea fish biomass trends and seabed structure. Marine Ecology Progress Series, 677:129– 140, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m677p129.pdf.

Ruggerone:2023:DKW

[RSvV^{+}23] G. T. Ruggerone, A. M. Springer, G. B. van Vliet, B. Connors, J. R. Irvine, L. D. Shaul, M. R. Sloat, and W. I. Atlas. From diatoms to killer whales: impacts of pink salmon on North Pacific ecosystems. Marine Ecology Progress Series, 719:1– 40, September 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Runnebaum:2020:PBH

[RTG^{+}20] J. Runnebaum, K. R. Tanaka, L. Guan, J. Cao, L. O’Brien, and Y. Chen. Predicting bycatch hotspots based on suitable habitat derived from fishery-independent data. Marine Ecology Progress Series, 641:159–175, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/641/
m641p159.pdf.

Ramseyer:2021:ENH

[RTT^{+}21] T. N. Ramseyer, A. Tronholm, T. Turner, M. E. Brandt, and T. B. Smith. Elevated nutrients and herbivory negatively affect Dictyota growth dynamics. Marine Ecology Progress Series, 671:81–95, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m671p081.pdf.

Russell:2020:ETR

[Rus20] B. D. Russell. The ecology of temperate reefs in a changing world. Marine Ecology Progress Series, 656:91–94, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m656p091.pdf.

Ramirez:2021:NEC

[RVCT21] H. Ramírez, V. Valverde-Cantillo, and P. Santidrián Tomillo. El Niño events and chlorophyll levels affect the reproductive frequency but not the seasonal reproductive output of East Pacific green turtles. Marine Ecology Progress Series, 659:237–246, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/659/m659p237.pdf.

Ramirez:2021:SIS

[RVSA^{+}21] F. Ramírez, D. Vicente-Sastre, I. Afán, J. M. Igual, D. Oro, and M. G. Forero. Stable isotopes in seabirds reflect changes in marine productivity patterns. Marine Ecology Progress Series, 662:169–180, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/662/m662p169.pdf.

Roman:2024:CDR

[RVV^{+}24] M. Román, E. Vázquez, R. M. Viejo, S. A. Woodin, D. S. Wethey, S. Román, N. Weidberg, J. S. Troncoso, M. M. Mendez, and C. Olabarria. Context-dependent resilience of intertidal seagrass and venerid clams after hyposalinity stress. Marine Ecology Progress Series, 729:117–133, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/729/m729p117.pdf.

Regnier:2024:ETA

[RWH^{+}24] T. Régnier, P. J. Wright, M. P. Harris, F. M. Gibb, M. Newell, D. Eerkes-Medrano, F. Daunt, and S. Wanless. Effect of timing and abundance of lesser sandeel on the breeding success of a North Sea seabird community. Marine Ecology Progress Series, 727:1–17, January 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Roman:2022:MPB

[RWM^{+}22] S. Román, N. Weidberg, C. Muñiz, A. Aguion, E. Vázquez, J. Santiago, P. Seoane, B. Barreiro, R. Outeiral, D. Villegas-Ríos, S. Fandiño, and G. Macho. Mesoscale patterns in barnacle reproduction are mediated by upwelling-driven thermal variability. Marine Ecology Progress Series, 685:153– 170, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/685/m685p153.pdf.

Ryan:2024:OMI

[RWR24] D. Ryan, C. Wogerbauer, and W. K. Roche. Otolith microchemistry to investigate nursery site fidelity and connectivity of juvenile European sea bass in Ireland. Marine Ecology Progress Series, 731:231–247, March 13, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m731p231.pdf.

Roth:2024:IVS

[RWSR24] J. S. Roth, P. C. Wilson, L. Steele, and L. K. Reynolds. Intraspecific variation in seagrass anti-herbivore defenses along a natural gradient in phosphorus availability impacts herbivore feeding decisions. Marine Ecology Progress Series, 727:111– 122, January 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/727/m727p111.pdf.

Roberts:2024:NAR

[RYF^{+}24] J. J. Roberts, T. M. Yack, E. Fujioka, P. N. Halpin, M. F. Baumgartner, O. Boisseau, S. Chavez-Rosales, T. V. N. Cole, M. P. Cotter, G. E. Davis, R. A. DiGiovanni, Jr., L. C. Ganley, L. P. Garrison, C. P. Good, T. A. Gowan, K. A. Jackson, R. D. Kenney, C. B. Khan, A. R. Knowlton, S. D. Kraus, G. G. Lockhart, K. S. Lomac-MacNair, C. A. Mayo, B. E. McKenna, W. A. McLellan, D. P. Nowacek, O. O’Brien, D. A. Pabst, D. L. Palka, E. M. Patterson, D. E. Pendleton, E. Quintana-Rizzo, N. R. Record, J. V. Redfern, M. E. Rickard, M. White, A. D. Whitt, and A. M. Zoidis. North Atlantic right whale density surface model for the US Atlantic evaluated with passive acoustic monitoring. Marine Ecology Progress Series, 732: 167–192, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m732p167.pdf.

Shahdadi:2024:PIB

[SAC24] A. Shahdadi, Z. Azan, and B. K. K. Chan. Phylogeography of the intertidal barnacle Chthamalus barnesi in the northwestern Indian Ocean. Marine Ecology Progress Series, 726:99– 112, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/726/m726p099.pdf.

Simon:2024:SVM

[SAF^{+}24] L. Simon, P. Arcese, C. H. Fox, K. H. Morgan, and S. Wilson. Seasonal variation in marine bird distribution in the northeast Pacific Ocean. Marine Ecology Progress Series, 735: 141–156, May 2, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m735p141.pdf.

Skogen:2024:BGI

[SAGG^{+}24] M. D. OPINION PIECES Skogen, J. M. Aarflot, L. M. García-García, R. Ji, M. Ruiz-Villarreal, E. Almroth-Rosell, A. Belgrano, D. Benkort, U. Daewel, M. Edman, R. Friedland, S. Gao, M. Hill-Cruz, S. S. Hjøllo, M. Huret, J. B. Kellner, S. van Leeuwen, A. Lopez de Gamiz-Zearra, M. Maar, E. A. Mousing, M. A. Peck, A. Pastor Rollan, S. F. Sailley, S. Saraiva, C. Speakman, T. Troost, and V. Ç. Yumruktepe. Bridging the gap: integrating models and observations for better ecosystem understanding. Marine Ecology Progress Series, 739:257–268, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m739p257.pdf.

Smit:2024:PTE

[SAH^{+}24] L. A. Smit, J. B. Adams, S. A. Hawkes, N. Peer, and G. M. Rishworth. Proportional top-down effects of grapsoid crabs on growth of Spartina maritima cordgrass in southern African salt marshes. Marine Ecology Progress Series, 739:49–64, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/739/m739p049.pdf.

Smith:2020:IVN

[SAMdL^{+}20] A. N. H. Smith, D. Acuña-Marrero, P. Salinas de León, E. S. Harvey, M. D. M. Pawley, and M. J. Anderson. Instantaneous vs. non-instantaneous diver-operated stereo–video (DOV) surveys of highly mobile sharks in the Galápagos Marine Reserve. Marine Ecology Progress Series, 649:111–123, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/649/
m649p111.pdf.

Schoen:2024:LIN

[SAMP24] S. K. Schoen, M. L. Arimitsu, C. E. Marsteller, and J. F. Piatt. Lingering impacts of the 2014–2016 northeast Pacific marine heatwave on seabird demography in Cook Inlet, Alaska (USA). Marine Ecology Progress Series, 737:121–136, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m737p121.pdf.

Scheibling:2020:PGP

[SB20a] R. E. Scheibling and R. Black. Persistence of giants: population dynamics of the limpet Scutellastra laticostata on rocky shores in Western Australia. Marine Ecology Progress Series, 646:79–92, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/646/m646p079.pdf.

Szpak:2020:SIA

[SB20b] P. Szpak and M. Buckley. Sulfur isotopes (δ ^{34}S) in Arctic marine mammals: indicators of benthic vs. pelagic foraging. Marine Ecology Progress Series, 653:205–216, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/653/
m653p205.pdf.

Smith:2020:NIE

[SBB20] R. S. Smith, J. A. Blaze, and J. E. Byers. Negative indirect effects of hurricanes on recruitment of range-expanding mangroves. Marine Ecology Progress Series, 644:65–74, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/644/m644p065.pdf.

Santos:2024:LEP

[SBB^{+}24] A. Resende Santos, M. Bento, S. Broszeit, J. Paula, and A. Marçal Correia. Linking ecosystem pressures and marine macroinvertebrate ecosystem services in mangroves and seagrasses. Marine Ecology Progress Series, 732:15–32, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/732/m732p015.pdf.

Smith:2021:TSI

[SBC^{+}21] C. E. Smith, D. T. Booth, A. Crosby, J. D. Miller, M. N. Staines, H. Versace, and C. A. Madden-Hof. Trialling seawater irrigation to combat the high nest temperature feminisation of green turtle Chelonia mydas hatchlings. Marine Ecology Progress Series, 667:177–190, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m667p177.pdf.

Schukat:2023:CBC

[SBDM^{+}23] A. Schukat, M. Bode-Dalby, J. C. Massing, W. Hagen, and H. Auel. Carbon budgets of copepod communities in the northern Humboldt Current System off Peru. Marine Ecology Progress Series, 724:47–65, December 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/724/
m724p047.pdf.

Solsona-Berga:2024:ASW

[SBFP^{+}24] A. Solsona-Berga, K. E. Frasier, N. Posdaljian, S. Baumann-Pickering, S. Wiggins, M. Soldevilla, L. Garrison, and J. A. Hildebrand. Accounting for sperm whale population demographics in density estimation using passive acoustic monitoring. Marine Ecology Progress Series, 746:121–140, October 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m746p121.pdf.

Smith:2024:MBC

[SBJ^{+}24] J. M. Smith, B. J. Burke, D. Jackson, B. Wells, B. Beckman, W. Duguid, T. P. Quinn, and D. D. Huff. Marine biophysical conditions influence the vertical and horizontal distribution of sub-adult Chinook salmon in nearshore marine waters. Marine Ecology Progress Series, 744:133–146, September 5, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/744/m744p133.pdf.

Skerritt:2023:ECD

[SBL^{+}23] D. J. Skerritt, M. C. Bell, K. J. Lees, A. C. Mill, N. V. C. Polunin, and C. Fitzsimmons. Estimating catchability and density of the European lobster Homarus gammarus from continuous, short-term mark-recapture data. Marine Ecology Progress Series, 715:79–89, July 27, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/715/
m715p079.pdf.

Salas:2022:EFD

[SBMW22] A. K. Salas, M. S. Ballard, T. A. Mooney, and P. S. Wilson. Effects of frequency-dependent spatial variation in soundscape settlement cues for reef fish larvae. Marine Ecology Progress Series, 687:1–21, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Steinke:2021:EDP

[SBRQ21] K. B. Steinke, K. S. Bernard, R. M. Ross, and L. B. Quetin. Environmental drivers of the physiological condition of mature female Antarctic krill during the spawning season: implications for krill recruitment. Marine Ecology Progress Series, 669:65–82, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m669p065.pdf.

Schwartzbach:2020:ACG

[SBS20] A. Schwartzbach, J. W. Behrens, and J. C. Svendsen. Atlantic cod Gadus morhua save energy on stone reefs: implications for the attraction versus production debate in relation to reefs. Marine Ecology Progress Series, 635:81–87, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/635/
m635p081.pdf.

Samarra:2022:ITE

[SBS^{+}22] F. I. P. Samarra, A. Borrell, A. Selbmann, S. D. Halldórson, C. Pampoulie, V. Chosson, T. Gunnlaugsson, G. M. Sigurdsson, A. Aguilar, and G. A. Víkingsson. Insights into the trophic ecology of white-beaked dolphins Lagenorhynchus albirostris and harbour porpoises Phocoena phocoena in Iceland. Marine Ecology Progress Series, 702:139–152, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/702/
m702p139.pdf.

Samarra:2024:TEB

[SBS^{+}24] F. I. P. Samarra, A. Borrell, A. Selbmann, V. Chosson, C. Chicco, C. E. Haas, C. J. Basran, A. Smith, B. G. Ovide, G. T. Einarsson, S. D. Halldórson, E. S. O’Brien, A. Eleman, G. M. Sigurdsson, and A. Aguilar. Trophic ecology, based on stable isotope values, of long-finned pilot whales Globicephala melas stranded on the Icelandic coast. Marine Ecology Progress Series, 748:163–174, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/748/
m748p163.pdf.

Saba:2021:CEO

[SBSS21] G. K. Saba, A. B. Bockus, C. T. Shaw, and B. A. Seibel. Combined effects of ocean acidification and elevated temperature on feeding, growth, and physiological processes of Antarctic krill Euphausia superba. Marine Ecology Progress Series, 665: 1–18, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Schenone:2024:SMP

[SBT24] S. Schenone, I. Bartl, and S. F. Thrush. Sediment microtopography predicts localised benthic ecosystem functioning and the effect of species interactions. Marine Ecology Progress Series, 749:47–55, November 21, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m749p047.pdf.

Searle:2022:PCD

[SBW^{+}22] K. R. Searle, A. Butler, J. J. Waggitt, P. G. H. Evans, L. R. Quinn, M. I. Bogdanova, T. J. Evans, J. E. Braithwaite, and F. Daunt. Potential climate-driven changes to seabird demography: implications for assessments of marine renewable energy development. Marine Ecology Progress Series, 690:185–200, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m690p185.pdf.

Seike:2023:IDF

[SC23] K. Seike and H. A. Curran. Interspecies differences in food sources for the tropical callichirid shrimp Neocallichirus spp. on San Salvador Island, Bahamas. Marine Ecology Progress Series, 709:33–44, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m709p033.pdf.

Sbragaglia:2020:LEK

[SCB^{+}20] V. Sbragaglia, J. Cerri, L. Bolognini, B. Dragićević, J. Dulćić, F. Grati, and E. Azzurro. Local ecological knowledge of recreational fishers reveals different meridionalization dynamics of two Mediterranean subregions. Marine Ecology Progress Series, 634:147–157, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2019/634/m634p147.pdf.

Sayco:2023:BRR

[SCK23] S. L. G. Sayco, P. C. Cabaitan, and H. Kurihara. Bleaching reduces reproduction in the giant clam Tridacna gigas. Marine Ecology Progress Series, 706:47–56, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/706/
m706p047.pdf.

Schabetsberger:2021:SML

[SCM21] R. Schabetsberger, Y. L. K. Chang, and M. J. Miller. Spawning migration and larval dispersal of tropical Pacific eels (Anguilla spp.) in the centre of their distribution ranges. Marine Ecology Progress Series, 670:167–184, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/670/
m670p167.pdf.

Sobrero:2023:CDP

[SCM^{+}23] L. Sobrero, L. M. Chiaverano, L. Machinandiarena, P. Martos, C. Derisio, D. Giberto, P. Betti, E. Leonarduzzi, G. Macchi, and A. Schiariti. Contrasting distributional patterns of two co-occurring ctenophores in relation to ichthyoplankton and environmental features in the Southwestern Atlantic Ocean. Marine Ecology Progress Series, 713:55–70, June 29, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/713/m713p055.pdf.

Smith:2023:DCM

[SCP^{+}23] H. A. Smith, C. C. M. Chen, F. J. Pollock, M. Re, B. L. Willis, and D. G. Bourne. Drivers of coral mortality in non-acute disturbance periods. Marine Ecology Progress Series, 717:37– 50, August 24, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/717/m717p037.pdf.

Shih:2024:AIN

[SCW^{+}24] C. Y. Shih, K. Z. Chang, P. L. Wang, J. Chang, and L. K. Kang. Assessing inorganic nitrogen transport in marine phytoplankton assemblages through the ^{15}N-tracer technique and metatranscriptomics. Marine Ecology Progress Series, 726:17– 30, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/726/m726p017.pdf.

Symons:2022:RPA

[SD22] S. C. Symons and A. W. Diamond. Resource partitioning in Atlantic puffins and razorbills facing declining food: an analysis of feeding areas and dive behaviour in relation to diet. Marine Ecology Progress Series, 699:153–165, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/699/
m699p153.pdf.

Strom:2021:TMN

[SDE^{+}21] H. Strøm, S. Descamps, M. Ekker, P. Fauchald, and B. Moe. Tracking the movements of North Atlantic seabirds: steps towards a better understanding of population dynamics and marine ecosystem conservation. Marine Ecology Progress Series, 676:97–116, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m676p097.pdf.

Stewart:2022:LFM

[SDE^{+}22] J. D. Stewart, J. W. Durban, H. Europe, H. Fearnbach, P. K. Hamilton, A. R. Knowlton, M. S. Lynn, C. A. Miller, W. L. Perryman, B. W. H. Tao, and M. J. Moore. Larger females have more calves: influence of maternal body length on fecundity in North Atlantic right whales. Marine Ecology Progress Series, 689:179–189, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m689p179.pdf.

Strom:2021:NBD

[SDF^{+}21] Hallvard Strøm, Sébastien Descamps, Per Fauchald, Børge Moe, Morten Ekker, Kyle H. Elliott, Samantha Patrick, Steve Votier, and Rory P. Wilson. Non-breeding distribution and movements of North Atlantic seabirds. Marine Ecology Progress Series, 676:95–96, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m676p095.pdf.

Selbach:2024:UCI

[SdFBM24] C. Selbach, E. de Framond-Benard, and K. N. Mouritsen. Under the cover of ice: Trematode infections affect survival and growth of wintering mussels. Marine Ecology Progress Series, 740:213–218, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2024/740/m740p213.pdf.

Stoffers:2021:DOH

[SdGWN21] T. Stoffers, M. de Graaf, H. V. Winter, and L. A. J. Nagelkerke. Distribution and ontogenetic habitat shifts of reef– associated shark species in the northeastern Caribbean. Marine Ecology Progress Series, 665:145–158, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/665/
m665p145.pdf.

Sabadel:2022:AAN

[SDM^{+}22] A. J. M. Sabadel, M. Décima, K. McComb, M. Meyers, N. Barr, M. Gall, K. Safi, and C. S. Law. Amino acid nitrogen stable isotopes as biomarkers of coastal phytoplankton assemblages and food web interactions. Marine Ecology Progress Series, 690:1–13, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Speare:2023:SSH

[SDM^{+}23] K. E. Speare, A. Duran, M. W. Miller, H. V. Moeller, and D. E. Burkepile. Small-scale habitat selection by larvae of a reef-building coral. Marine Ecology Progress Series, 724:67– 79, December 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/724/m724p067.pdf.

Sabadel:2020:SIA

[SDW20] A. J. M. Sabadel, L. M. Durante, and S. R. Wing. Stable isotopes of amino acids from reef fishes uncover Suess and nitrogen enrichment effects on local ecosystems. Marine Ecology Progress Series, 647:149–160, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m647p149.pdf.

Sullaway:2020:INN

[SE20] G. H. Sullaway and M. S. Edwards. Impacts of the nonnative alga Sargassum horneri on benthic community production in a California kelp forest. Marine Ecology Progress Series, 637:45–57, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/637/m637p045.pdf.

Sbragaglia:2024:TOT

[SEJ^{+}24] V. Sbragaglia, L. Espasandín, I. Jarić, R. Vardi, F. Ramírez, and M. Coll. Tracking ongoing transboundary marine distributional range shifts in the digital era. Marine Ecology Progress Series, 728:103–114, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/728/
m728p103.pdf.

Souza:2023:AAP

[SF23] J. A. Souza and A. A. V. Flores. Abundance and active patch selection modulate reproductive connectivity and fitness of pea crabs living on sand dollars. Marine Ecology Progress Series, 721:1–15, October 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Sampei:2020:VCD

[SFF^{+}20] M. Sampei, A. Forest, L. Fortier, T. Yamamoto, H. Hattori, and H. Sasaki. Variations in contributions of dead copepods to vertical fluxes of particulate organic carbon in the Beaufort Sea. Marine Ecology Progress Series, 642: 67–81, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/642/m642p067.pdf.

Schneider:2023:RPB

[SFL23] A. K. Schneider, M. C. Fabrizio, and R. N. Lipcius. Reproductive potential of the blue crab spawning stock in Chesapeake Bay across eras and exploitation rates using nemertean worms as biomarkers. Marine Ecology Progress Series, 716: 77–91, August 10, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m716p077.pdf.

Silva:2023:WIC

[SFLQ23] F. C. Silva, S. R. Floeter, M. Lindegren, and J. P. Quimbayo. Warming-induced changes in reef fish community traits in the Southwestern Atlantic transition zone. Marine Ecology Progress Series, 710:107–123, May 4, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/710/
m710p107.pdf.

Smith:2023:SAH

[SFMK^{+}23] A. B. Smith, I. Fischer-McMorrow, Y. Kolbeinsson, M. Rasmussen, M. R. Shero, J. N. McElwaine, O. R. Jones, and T. A. Mooney. Sensitive aerial hearing within a noisy nesting soundscape in a deep-diving seabird, the common murre Uria aalge. Marine Ecology Progress Series, 714:87–104, July 13, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m714p087.pdf.

Stortini:2021:MLT

[SFO^{+}21] C. H. Stortini, K. T. Frank, V. J. Ontiveros, W. C. Leggett, and N. L. Shackell. Maintenance of long-term equilibrium in a perturbed metacommunity of sub-arctic marine fishes. Marine Ecology Progress Series, 675:81–96, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m675p081.pdf.

Stallings:2021:IDE

[SFOPL21] C. D. Stallings, I. M. Freytes-Ortiz, M. M. Plafcan, and C. Langdon. Inducible defenses in an estuarine bivalve do not alter predator handling times and are not affected by climate change. Marine Ecology Progress Series, 679:73–84, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m679p073.pdf.

Seyer:2021:SVM

[SGB^{+}21] Y. Seyer, G. Gauthier, J. Bêty, J. F. Therrien, and N. Lecomte. Seasonal variations in migration strategy of a long-distance Arctic-breeding seabird. Marine Ecology Progress Series, 677:1–16, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Sargac:2021:COR

[ŠGH^{+}21] Z. Šargač, L. Giménez, S. Harzsch, J. Krieger, K. Fjordside, and G. Torres. Contrasting offspring responses to variation in salinity and temperature among populations of a coastal crab: a maladaptive ecological surprise? Marine Ecology Progress Series, 677:51–65, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m677p051.pdf.

Soria:2024:NNB

[SGP24] S. A. Soria, J. L. Gutiérrez, and M. G. Palomo. Non-native barnacle accelerates the recovery from disturbance of intertidal mussel beds. Marine Ecology Progress Series, 732:73– 83, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/732/m732p073.pdf.

Spence:2021:SFC

[SGW^{+}21] M. A. Spence, C. A. Griffiths, J. J. Waggitt, H. J. Bannister, R. B. Thorpe, A. G. Rossberg, and C. P. Lynam. Sustainable fishing can lead to improvements in marine ecosystem status: an ensemble-model forecast of the North Sea ecosystem. Marine Ecology Progress Series, 680:207–221, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/680/
m680p207.pdf.

Sabal:2020:CCS

[SHB^{+}20] M. C. Sabal, E. L. Hazen, S. J. Bograd, R. B. MacFarlane, I. D. Schroeder, S. A. Hayes, J. A. Harding, K. L. Scales, P. I. Miller, A. J. Ammann, and B. K. Wells. California Current seascape influences juvenile salmon foraging ecology at multiple scales. Marine Ecology Progress Series, 634:159–173, 2020.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m634p159.pdf.

Sameoto:2021:CID

[SHG^{+}21] J. A. Sameoto, K. Hall, S. E. Gass, D. Keith, S. Kirchhoff, and C. J. Brown. Conservation implications of demographic changes in the horse mussel Modiolus modiolus population of the inner Bay of Fundy. Marine Ecology Progress Series, 670:93–104, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/670/m670p093.pdf.

Seher:2022:OIE

[SHH^{+}22] V. L. Seher, B. A. Holzman, E. Hines, R. W. Bradley, P. Warzybok, and B. H. Becker. Ocean -influenced estuarine habitat buffers high interannual variation in seabird reproductive success. Marine Ecology Progress Series, 689: 155–167, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/689/m689p155.pdf.

Stepanuk:2021:ASB

[SHL^{+}21] J. E. F. Stepanuk, E. I. Heywood, J. F. Lopez, R. A. Di-Giovanni, Jr., and L. H. Thorne. Age-specific behavior and habitat use in humpback whales: implications for vessel strike. Marine Ecology Progress Series, 663:209–222, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/663/
m663p209.pdf.

Shamblin:2020:GTM

[SHM^{+}20a] B. M. Shamblin, K. M. Hart, K. J. Martin, S. A. Ceriani, D. A. Bagley, K. L. Mansfield, L. M. Ehrhart, and C. J. Nairn. Green turtle mitochondrial microsatellites indicate finer-scale natal homing to isolated islands than to continental nesting sites. Marine Ecology Progress Series, 643: 159–171, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/643/m643p159.pdf.

Sherman:2020:WSR

[SHM^{+}20b] C. S. Sherman, M. R. Heupel, S. K. Moore, A. Chin, and C. A. Simpfendorfer. When sharks are away, rays will play: effects of top predator removal in coral reef ecosystems. Marine Ecology Progress Series, 641:145–157, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/641/m641p145.pdf.

Simpson:2021:HSF

[SHS21] S. J. Simpson, N. E. Humphries, and D. W. Sims. Habitat selection, fine-scale spatial partitioning and sexual segregation in Rajidae, determined using passive acoustic telemetry. Marine Ecology Progress Series, 666:115–134, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/666/
m666p115.pdf.

Smith:2022:LTD

[SHT^{+}22] R. S. Smith, S. Hogan, K. N. Tedford, B. Lusk, M. A. Reidenbach, and M. C. N. Castorani. Long-term data reveal greater intertidal oyster biomass in predicted suitable habitat. Marine Ecology Progress Series, 683:221–226, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_

oa/m683p221.pdf. See corrigendum [Ano22d].

Sakamaki:2020:EOA

[SHZ^{+}20] T. Sakamaki, K. Hayashi, Y. Zheng, M. Fujibayashi, and O. Nishimura. Effects of oyster age on selective suspensionfeeding and the chemical composition of biodeposits: insights from fatty acid analysis. Marine Ecology Progress Series, 644:75–89, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/644/m644p075.pdf.

Sutton:2020:CFD

[SIBM20] L. Sutton, K. Iken, B. A. Bluhm, and F. J. Mueter. Comparison of functional diversity of two Alaskan Arctic shelf epibenthic communities. Marine Ecology Progress Series, 651:1–21, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Sigurdsson:2023:IBR

[Sig23] G. M. Sigurdsson. Increase in bycatch rates of plunge- and surface-feeding seabirds observed while setting and hauling bottom set gillnets equipped with LED lights. Marine Ecology Progress Series, 720:191–194, October 5, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/720/
m720p191.pdf.

Shelley:2022:LFW

[SJ22] C. E. Shelley and D. W. Johnson. Larval fish in a warming ocean: a bioenergetic study of temperature-dependent growth and assimilation efficiency. Marine Ecology Progress Series, 691:97–114, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/691/m691p097.pdf.

Skogen:2021:OPD

[SJA^{+}21] M. D. Skogen, R. Ji, A. Akimova, U. Daewel, C. Hansen, S. S. Hjøllo, S. M. van Leeuwen, M. Maar, D. Macias, E. A. Mousing, E. Almroth-Rosell, S. F. Sailley, M. A. Spence, T. A. Troost, and K. van de Wolfshaar. Opinion piece: Disclosing the truth: Are models better than observations? Marine Ecology Progress Series, 680:7–13, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/680/m680p007.pdf.

Spotowitz:2022:NEE

[SJH^{+}22] L. Spotowitz, T. Johansen, A. Hansen, E. Berg, C. Stransky, and P. Fischer. New evidence for the establishment of coastal cod Gadus morhua in Svalbard fjords. Marine Ecology Progress Series, 696:119–133, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m696p119.pdf.

Shearer:2022:SFP

[SJQ^{+}22] J. M. Shearer, F. H. Jensen, N. J. Quick, A. Friedlaender, B. Southall, D. P. Nowacek, M. Bowers, H. J. Foley, Z. T. Swaim, D. M. Waples, and A. J. Read. Short-finned pilot whales exhibit behavioral plasticity in foraging strategies mediated by their environment. Marine Ecology Progress Series, 695:1–14, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Schiott:2023:EDM

[SJS^{+}23] S. Schiøtt, M. R. Jensen, E. E. Sigsgaard, P. R. Møller, M. d. P. Avila, P. F. Thomsen, and S. Rysgaard. Environmental DNA metabarcoding reveals seasonal and spatial variation in the vertebrate fauna of Ilulissat Icefjord, Greenland. Marine Ecology Progress Series, 706:91–108, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m706p091.pdf.

Sisti:2024:BGB

[SJSR24] A. R. Sisti, B. Jellison, J. D. Shields, and E. B. Rivest. Broodgrooming behavior of American lobsters Homarus americanus in conditions of ocean warming and acidification. Marine Ecology Progress Series, 744:83–99, September 5, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m744p083.pdf.

Smith:2023:ETL

[SKHC23] S. Smith, H. P. Kunc, I. Hewson, and P. C. Collins. Elevated temperature linked to signs associated with sea star wasting disease in a keystone European species, Asterias rubens. Marine Ecology Progress Series, 724:97–109, December 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/724/m724p097.pdf.

Suzuki:2020:EHG

[SKI^{+}20] H. Suzuki, Y. Kubo, E. Inomata, Y. Agatsuma, and M. N. Aoki. Effects of herbivorous gastropod grazing on the sedimentation and succession of subtidal macroalgal assemblages. Marine Ecology Progress Series, 656:123–138, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/656/
m656p123.pdf.

Satoh:2023:CJY

[SKTO23] K. Satoh, S. Katayama, T. Tanabe, and K. Okamoto. Connectivity in juvenile yellowfin tuna Thunnus albacares between temperate and tropical regions of the western Pacific Ocean. Marine Ecology Progress Series, 713:151–172, June 29, 2023.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m713p151.pdf.

Stern:2021:FHB

[SLB^{+}21] J. H. Stern, K. L. Laidre, E. W. Born, Ø. Wiig, C. Sonne, R. Dietz, A. Fisk, and M. A. McKinney. Feeding habits of Baffin Bay polar bears Ursus maritimus: insight from stable isotopes and total mercury in hair. Marine Ecology Progress Series, 677:233–244, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2021/677/m677p233.pdf.

Serra-Llinares:2020:ISL

[SLBK^{+}20] R. M. Serra-Llinares, T. Bøhn, Ø. Karlsen, R. Nilsen, C. Freitas, J. Albretsen, T. Haraldstad, E. B. Thorstad, K. M. S. Elvik, and P. A. Bjørn. Impacts of salmon lice on mortality, marine migration distance and premature return in sea trout. Marine Ecology Progress Series, 635:151–168, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m635p151.pdf.

Sartoretto:2024:EGC

[SLG^{+}24] S. Sartoretto, J. B. Ledoux, E. Gueret, D. Guillemain, C. Ravel, L. Moirand, and D. Aurelle. Ecological and genomic characterization of a remarkable natural heritage: a mesophotic ‘giant’ Paramuricea clavata forest. Marine Ecology Progress Series, 728:85–101, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/728/
m728p085.pdf.

Schoutens:2022:TTM

[SLH^{+}22] K. Schoutens, P. Luys, M. Heuner, E. Fuchs, V. Minden, T. Schulte Ostermann, T. J. Bouma, J. Van Belzen, and S. Temmerman. Traits of tidal marsh plants determine survival and growth response to hydrodynamic forcing: implications for nature-based shoreline protection. Marine Ecology Progress Series, 693:107–124, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/693/m693p107.pdf.

Suca:2020:VSE

[SLJ^{+}20] J. J. Suca, A. Lillis, I. T. Jones, M. B. Kaplan, A. R. Solow, A. D. Earl, S. Habtes, A. Apprill, J. K. Llopiz, and T. A. Mooney. Variable and spatially explicit response of fish larvae to the playback of local, continuous reef soundscapes. Marine Ecology Progress Series, 653:131–151, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/653/
m653p131.pdf.

Steinfurth:2022:IBF

[SLO^{+}22] R. C. Steinfurth, T. Lange, N. S. Oncken, E. Kristensen, C. O. Quintana, and M. R. Flindt. Improved benthic fauna community parameters after large-scale eelgrass (Zostera marina) restoration in Horsens Fjord, Denmark. Marine Ecology Progress Series, 687:65–77, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m687p065.pdf.

Seveso:2020:EGH

[SMA^{+}20] D. Seveso, D. Maggioni, R. Arrigoni, E. Montalbetti, M. L. Berumen, P. Galli, and S. Montano. Environmental gradients and host availability affecting the symbiosis between Pteroclava krempfi and alcyonaceans in the Saudi Arabian central Red Sea. Marine Ecology Progress Series, 653:91– 103, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/653/m653p091.pdf.

Spaet:2020:ECP

[SMB^{+}20] J. L. Y. Spaet, A. Manica, C. P. Brand, C. Gallen, and P. A. Butcher. Environmental conditions are poor predictors of immature white shark Carcharodon carcharias occurrences on coastal beaches of eastern Australia. Marine Ecology Progress Series, 653:167–179, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m653p167.pdf.

Sibley:2023:CIS

[SME^{+}23] E. C. P. REVIEW Sibley, A. S. Madgett, T. S. Elsdon, M. J. Marnane, E. S. Harvey, and P. G. Fernandes. The capacity of imaging sonar for quantifying the abundance, species richness, and size of reef fish assemblages. Marine Ecology Progress Series, 717:157–179, August 24, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m717p157.pdf.

Schlaff:2020:ATL

[SMH^{+}20] A. Schlaff, P. Menéndez, M. Hall, M. Heupel, T. Armstrong, and C. Motti. Acoustic tracking of a large predatory marine gastropod, Charonia tritonis, on the Great Barrier Reef. Marine Ecology Progress Series, 642:147–161, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/642/
m642p147.pdf.

Schrage:2023:FRP

[SMK23] K. R. Schrage and K. S. Meyer-Kaiser. Fine-resolution patterns of fouling community settlement, growth, and mortality using the CATAIN camera system. Marine Ecology Progress Series, 711:17–29, May 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/711/m711p017.pdf.

Sandoval:2022:MCS

[SMPLF^{+}22] L. A. Sandoval, J. E. Mancera-Pineda, J. Leal-Flórez, J. F. Blanco-Libreros, and A. Delgado-Huertas. Mangrove carbon sustains artisanal fish and other estuarine consumers in a major mangrove area of the southern Caribbean Sea. Marine Ecology Progress Series, 681:21–35, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/681/
m681p021.pdf.

Stratmann:2024:RAS

[SMS^{+}24] T. Stratmann, F. J. Murillo, M. Sacau, M. K. Alonso, and E. Kenchington. Role of astrophorina sponges (Demospongiae) in food-web interactions at the Flemish Cap (NW Atlantic). Marine Ecology Progress Series, 729:99–116, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m729p099.pdf.

Steinfurth:2024:MCA

[SNE^{+}24] R. C. Steinfurth, B. Nielsen, L. F. Elsberg, T. Lange, N. S. Oncken, C. O. Quintana, E. Kristensen, and M. R. Flindt. Macrofaunal colonization after large-scale sand-cap and Zostera marina restoration of organic enriched sediments in a Danish fjord. Marine Ecology Progress Series, 729:63–79, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m729p063.pdf.

Smith:2020:BRI

[SNH^{+}20] J. A. M. Smith, L. J. Niles, S. Hafner, A. Modjeski, and T. Dillingham. Beach restoration improves habitat quality for American horseshoe crabs and shorebirds in the Delaware Bay, USA. Marine Ecology Progress Series, 645:91–107, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/645/m645p091.pdf.

Sheehan:2020:MCP

[SNP20] C. E. Sheehan, D. A. Nielsen, and K. Petrou. Macromolecular composition, productivity and dimethylsulfoniopropionate in Antarctic pelagic and sympagic microalgal communities. Marine Ecology Progress Series, 640:45–61, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/640/
m640p045.pdf.

Stack:2020:ISD

[SON^{+}20] S. H. Stack, G. L. Olson, V. Neamtu, A. F. Machernis, R. W. Baird, and J. J. Currie. Identifying spinner dolphin Stenella longirostris longirostris movement and behavioral patterns to inform conservation strategies in Maui Nui, Hawai’i. Marine Ecology Progress Series, 644:187–197, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m644p187.pdf.

Spies:2021:SEN

[SOS^{+}21a] I. Spies, J. W. Orr, D. E. Stevenson, P. Goddard, G. Hoff, J. Guthridge, M. Hollowed, and C. Rooper. Skate egg nursery areas support genetic diversity of Alaska and Aleutian skates in the Bering Sea. Marine Ecology Progress Series, 669:121–138, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m669p121.pdf.

Spies:2021:GEE

[SOS^{+}21b] I. Spies, J. W. Orr, D. E. Stevenson, P. Goddard, G. R. Hoff, J. Guthridge, and C. N. Rooper. Genetic evidence from embryos suggests a new species of skate related to Bathyraja parmifera (Rajiformes: Arhynchobatidae) in the Bering Sea. Marine Ecology Progress Series, 670:155–166, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m670p155.pdf.

Simone:2022:OAM

[SOSE22] M. N. Simone, J. M. Oakes, K. G. Schulz, and B. D. Eyre. Ocean acidification modifies the impact of warming on sediment nitrogen recycling and assimilation by enhancing the benthic microbial loop. Marine Ecology Progress Series, 681: 53–69, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/681/m681p053.pdf.

Stortini:2020:MMS

[SPFL20] C. H. Stortini, B. Petrie, K. T. Frank, and W. C. Leggett. Marine macroinvertebrate species-area relationships, assemblage structure and their environmental drivers on submarine banks. Marine Ecology Progress Series, 641:25–47, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m641p025.pdf.

Sigler:2023:ADS

[SRG^{+}23] M. F. Sigler, C. N. Rooper, P. Goddard, R. Wilborn, and K. Williams. Alaska deep-sea coral and sponge assemblages are well-defined and mostly predictable from local environmental conditions. Marine Ecology Progress Series, 712:67– 85, June 8, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/712/m712p067.pdf.

Schonfeld:2024:HIE

[SRG^{+}24] A. J. Schonfeld, G. M. Ralph, J. Gartland, P. St-Laurent, M. A. M. Friedrichs, and R. J. Latour. Hypoxia influences the extent and dynamics of suitable fish habitat in Chesapeake Bay, USA. Marine Ecology Progress Series, 748:117– 135, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m748p117.pdf.

Satterthwaite:2021:ILD

[SRHM21] E. V. Satterthwaite, J. P. Ryan, J. B. J. Harvey, and S. G. Morgan. Invertebrate larval distributions influenced by adult habitat distribution, larval behavior, and hydrodynamics in the retentive upwelling shadow of Monterey Bay, California, USA. Marine Ecology Progress Series, 661:35–47, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m661p035.pdf.

Saluta:2023:LSE

[SRK^{+}23] G. G. Saluta, G. M. Ralph, K. E. Knick, M. S. Seebo, and R. N. Lipcius. Lethal and sublethal effects of simulated dredged sediment deposition on overwintering blue crabs Callinectes sapidus. Marine Ecology Progress Series, 719:65–75, September 21, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m719p065.pdf.

Svendsen:2020:DDG

[SRR^{+}20] G. M. Svendsen, M. Ocampo Reinaldo, M. A. Romero, G. Williams, A. Magurran, S. Luque, and R. A. González. Drivers of diversity gradients of a highly mobile marine assemblage in a mesoscale seascape. Marine Ecology Progress Series, 638:149–164, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/638/m638p149.pdf.

Sanchez:2020:AOC

[SRS^{+}20] P. J. Sanchez, J. R. Rooker, M. Zapp Sluis, J. Pinsky, M. A. Dance, B. Falterman, and R. J. Allman. Application of otolith chemistry at multiple life history stages to assess population structure of Warsaw grouper in the Gulf of Mexico.

Marine Ecology Progress Series, 651:111–123, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m651p111.pdf.

Skjoldal:2024:SFZ

[SS24] H. R. Skjoldal and E. Sperfeld. Size-fractioned zooplankton biomass in the Barents Sea: macroecological patterns across biogeography, climate, and varying ecosystem state (1989–2020). Marine Ecology Progress Series, 726:31–47, January 11, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m726p031.pdf.

Schram:2020:ADF

[SSB^{+}20] J. B. Schram, H. L. Sorensen, R. D. Brodeur, A. W. E. Galloway, and K. R. Sutherland. Abundance, distribution, and feeding ecology of Pyrosoma atlanticum in the Northern California Current. Marine Ecology Progress Series, 651: 97–110, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/651/m651p097.pdf.

Smith:2023:PIE

[SSB^{+}23] E. A. Elliott Smith, P. Szpak, T. J. Braje, B. Newsom, and T. C. Rick. Pre-industrial ecology and foraging behavior of swordfish Xiphias gladius in the eastern North Pacific. Marine Ecology Progress Series, 711:129–134, May 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/711/
m711p129.pdf.

Swieca:2020:CTF

[SSBA^{+}20] K. Swieca, S. Sponaugle, C. Briseño-Avena, M. S. Schmid, R. D. Brodeur, and R. K. Cowen. Changing with the tides: fine-scale larval fish prey availability and predation pressure near a tidally modulated river plume. Marine Ecology Progress Series, 650:217–238, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/650/m650p217.pdf.

Schoombie:2023:IWC

[SSC^{+}23] J. Schoombie, S. Schoombie, M. Connan, C. W. Jones, M. Risi, K. J. Craig, L. Smith, P. G. Ryan, and E. L. C. Shepard. Impact of wind on crash-landing mortality in grey-headed albatrosses Thalassarche chrysostoma breeding on Marion Island. Marine Ecology Progress Series, 723:213–225, November 23, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/723/m723p213.pdf.

Sweeney:2022:CBW

[SSF^{+}22] D. A. Sweeney, G. S. Schorr, E. A. Falcone, B. K. Rone, R. D. Andrews, S. N. Coates, S. L. Watwood, S. L. DeRuiter, M. P. Johnson, and D. J. Moretti. Cuvier’s beaked whale foraging dives identified via machine learning using depth and triaxial acceleration. Marine Ecology Progress Series, 692:195–208, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m692p195.pdf.

Sahin:2023:HRE

[SSFD^{+}23] D. Sahin, V. Schoepf, K. Filbee-Dexter, D. P. Thomson, B. Radford, and T. Wernberg. Heating rate explains speciesspecific coral bleaching severity during a simulated marine heatwave. Marine Ecology Progress Series, 706:33–46, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m706p033.pdf.

Steeves:2020:EFP

[SSFS20] L. Steeves, T. Strohmeier, R. Filgueira, and Ø. Strand. Exploring feeding physiology of Mytilus edulis across geographic and fjord gradients in low-seston environments. Marine Ecology Progress Series, 651:71–84, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/651/
m651p071.pdf.

Stevens:2022:TLF

[SSG^{+}22] C. J. Stevens, R. Sahota, M. D. Galbraith, T. A. Venello, A. C. Bazinet, M. Hennekes, K. Yongblah, and S. K. Juniper. Total lipid and fatty acid composition of mesozooplankton functional group members in the NE Pacific over a range of productivity regimes. Marine Ecology Progress Series, 687:43–64, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m687p043.pdf.

Sims:2023:CEL

[SSK^{+}23] C. A. Sims, E. M. Sampayo, S. W. Kim, M. M. Mayfield, and J. M. Pandolfi. Coral early life history dynamics: conspecific facilitation or limitation are dependent on distinct life stage interactions. Marine Ecology Progress Series, 720:39– 57, October 5, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/720/m720p039.pdf.

Sullivan-Stack:2020:TER

[SSM20] J. Sullivan-Stack and B. A. Menge. Testing ecological release as a compensating mechanism for mass mortality in a keystone predator. Marine Ecology Progress Series, 637: 59–69, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/637/m637p059.pdf.

Scheer:2022:FFS

[SSP^{+}22] S. L. Scheer, A. K. Sweetman, U. Piatkowski, E. K. Rohlfer, and H. J. T. Hoving. Food fall-specific scavenging response to experimental medium-sized carcasses in the deep sea. Marine Ecology Progress Series, 685:31–48, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/685/
m685p031.pdf.

Sahoo:2020:NRC

[SST^{+}20a] D. Sahoo, H. Saxena, N. Tripathi, M. A. Khan, A. Rahman, S. Kumar, A. K. Sudheer, and A. Singh. Non-Redfieldian C:N:P ratio in the inorganic and organic pools of the Bay of Bengal during the summer monsoon. Marine Ecology Progress Series, 653:41–55, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/653/m653p041.pdf.

Spyksma:2020:ICA

[SST20b] A. J. P. Spyksma, N. T. Shears, and R. B. Taylor. Injured conspecifics as an alarm cue for the sea urchin Evechinus chloroticus. Marine Ecology Progress Series, 641:135– 144, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/641/m641p135.pdf.

Suzuki:2024:SSP

[STF^{+}24] G. Suzuki, S. Tashiro, Y. Fujikura, I. Tanita, Y. Suhara, W. Fujiie, Y. Yonezawa, T. Kanyama, and A. Suto. Settlement success and post-settlement survival of Acropora sp. aff. tenuis spat within a small bay in Japan. Marine Ecology Progress Series, 728:15–24, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/728/
m728p015.pdf.

Sea:2021:EPS

[STH21] M. A. Sea, S. F. Thrush, and J. R. Hillman. Environmental predictors of sediment denitrification rates within restored green-lipped mussel Perna canaliculus beds. Marine Ecology Progress Series, 667:1–13, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Smeti:2024:RCM

[STM^{+}24] E. Smeti, G. Tsirtsis, A. Meziti, K. A. Kormas, D. B. Danielidis, D. L. Roelke, and S. Spatharis. Resilience of coastal marine metacommunities to increasing hydraulic connectivity. Marine Ecology Progress Series, 738:89–101, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/738/m738p089.pdf.

Shimabukuro:2023:ANP

[STO^{+}23] U. Shimabukuro, A. Takahashi, J. Okado, N. Kokubun, J. B. Thiebot, A. Will, Y. Watanuki, B. Addison, S. A. Hatch, J. M. Hipfner, L. Slater, B. A. Drummond, and A. S. Kitaysky. Across the North Pacific, dietary-induced stress of breeding rhinoceros auklets increases with high summer Pacific Decadal Oscillation index. Marine Ecology Progress Series, 708:177–189, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/708/m708p177.pdf.

Strathmann:2023:PDE

[Str23] R. R. Strathmann. Perils of drifting encapsulated embryos of the periwinkle Littorina scutulata from failures at launch and unscheduled landings. Marine Ecology Progress Series, 703:109–124, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/703/m703p109.pdf.

Stark:2020:BSP

[STY^{+}20] K. A. Stark, P. L. Thompson, J. Yakimishyn, L. Lee, E. M. Adamczyk, M. Hessing-Lewis, and M. I. O’Connor. Beyond a single patch: local and regional processes explain diversity patterns in a seagrass epifaunal metacommunity. Marine Ecology Progress Series, 655:91–106, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/655/
m655p091.pdf.

Saalfeld:2024:LSF

[SVB^{+}24] S. T. Saalfeld, M. Valcu, S. Brown, W. English, M. A. Giroux, A. L. Harrison, J. Krietsch, K. Kuletz, J. F. Lamarre, C. Latty, N. Lecomte, R. McGuire, M. Robards, A. Scarpignato, S. Schulte, P. A. Smith, B. Kempenaers, and R. B. Lanctot. From land to sea: the fall migration of the red phalarope through the Western Hemisphere. Marine Ecology Progress Series, 729:1–29, February 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Stalder:2020:IEV

[SvBS^{+}20] D. Stalder, F. M. van Beest, S. Sveegaard, R. Dietz, J. Teilmann, and J. Nabe-Nielsen. Influence of environmental variability on harbour porpoise movement. Marine Ecology Progress Series, 648:207–219, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m648p207.pdf.

Steeves:2022:RBP

[SVS^{+}22] L. Steeves, C. Vimond, T. Strohmeier, S. Casas, Ø. Strand, L. A. Comeau, and R. Filgueira. Relationship between pumping rate and particle capture efficiency in three species of bivalves. Marine Ecology Progress Series, 691:55–68, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/691/m691p055.pdf.

Schwanck:2024:MHR

[SVT^{+}24] T. N. Schwanck, L. F. Vizer, J. Thorburn, J. Dodd, P. J. Wright, D. W. Donnan, L. R. Noble, and C. S. Jones. Mitochondrial haplotypes reveal low diversity and restricted connectivity of the critically endangered batoid population in a Marine Protected Area. Marine Ecology Progress Series, 731: 279–291, March 13, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m731p279.pdf.

Salmond:2023:SLL

[SW23a] N. H. Salmond and S. R. Wing. Sub-lethal and lethal effects of chronic and extreme multiple stressors on a critical New Zealand bivalve under hypoxia. Marine Ecology Progress Series, 703:81–93, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m703p081.pdf.

Saltzman:2023:DRE

[SW23b] J. Saltzman and E. R. White. Determining the role of environmental covariates on planktivorous elasmobranch population trends within an isolated marine protected area. Marine Ecology Progress Series, 722:107–123, November 9, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/722/
m722p107.pdf.

Shoji:2023:ENP

[SWC^{+}23a] A. Shoji, S. Whelan, J. T. Cunningham, S. A. Hatch, Y. Niizuma, C. Nakajima, and K. H. Elliott. Ecological niche partitioning in two Pacific puffins. Marine Ecology Progress Series, 709:125–139, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m709p125.pdf.

Stedt:2023:MSS

[SWC^{+}23b] J. Stedt, M. Wahlberg, J. Carlström, P. A. Nilsson, M. Amundin, N. Oskolkov, and P. Carlsson. Micro-scale spatial preference and temporal cyclicity linked to foraging in harbour porpoises. Marine Ecology Progress Series, 708: 143–161, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m708p143.pdf.

Sun:2020:TBT

[SWHE20] A. Sun, S. Whelan, S. A. Hatch, and K. H. Elliott. Tags below three percent of body mass increase nest abandonment by rhinoceros auklets, but handling impacts decline as breeding progresses. Marine Ecology Progress Series, 643: 173–181, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m643p173.pdf.

Segura:2021:MPF

[SWJ^{+}21] A. M. Segura, R. Wiff, A. J. Jaureguizar, A. C. Milessi, and G. Perera. A macroecological perspective on the fluctuations of exploited fish populations. Marine Ecology Progress Series, 665:177–183, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/665/m665p177.pdf.

Sakowski:2020:OCF

[SWP20] E. G. Sakowski, K. E. Wommack, and S. W. Polson. Oyster calcifying fluid harbors persistent and dynamic autochthonous bacterial populations that may aid in shell formation. Marine Ecology Progress Series, 653:57–75, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/653/
m653p057.pdf.

Stuart:2024:SCM

[SWP^{+}24] C. E. Stuart, L. M. Wedding, S. J. Pittman, J. E. Serafy, A. Moura, A. W. Bruckner, and S. J. Green. Seascape connectivity modeling predicts hotspots of fish-derived nutrient provisioning to restored coral reefs. Marine Ecology Progress Series, 731:179–196, March 13, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m731p179.pdf.

Schoombie:2023:WDE

[SWR23] S. Schoombie, R. P. Wilson, and P. G. Ryan. Wind driven effects on the fine-scale flight behaviour of dynamic soaring wandering albatrosses. Marine Ecology Progress Series, 723: 119–134, November 23, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m723p119.pdf.

Scott:2021:STV

[SYMR21] A. L. Scott, P. H. York, P. I. Macreadie, and M. A. Rasheed. Spatial and temporal variability of green turtle and dugong herbivory in seagrass meadows of the southern Great Barrier Reef (GBR). Marine Ecology Progress Series, 667:225– 231, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/667/m667p225.pdf.

Tavares:2021:SPC

[TABM21] M. Tavares, D. J. Alvares, and M. Borges-Martins. Spatiotemporal patterns of carrion biomass of marine tetrapods at the ocean–land interface on the southern Brazilian coastline. Marine Ecology Progress Series, 672:57–72, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/672/
m672p057.pdf.

Thompson:2022:DIE

[TAN^{+}22] P. L. Thompson, S. C. Anderson, J. Nephin, D. R. Haggarty, M. A. Peña, P. A. English, K. S. P. Gale, and E. Rubidge. Disentangling the impacts of environmental change and commercial fishing on demersal fish biodiversity in a northeast Pacific ecosystem. Marine Ecology Progress Series, 689:137–154, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m689p137.pdf.

Tricas:2021:PSI

[TB21] T. C. Tricas and K. S. Boyle. Parrotfish soundscapes: implications for coral reef management. Marine Ecology Progress Series, 666:149–169, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/666/m666p149.pdf.

Taccardi:2021:SPE

[TBH21] E. Y. Taccardi, I. R. Bricknell, and H. J. Hamlin. Seasonal progression of embryo size and lipid reserves in sea lice Lepeophtheirus salmonis collected from salmon farms. Marine Ecology Progress Series, 664:79–86, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/664/
m664p079.pdf.

Tanner:2020:TET

[TBS20] R. L. Tanner, R. C. K. Bowie, and J. H. Stillman. Thermal exposure and transgenerational plasticity influence embryonic success in a bivoltine estuarine sea hare. Marine Ecology Progress Series, 634:199–211, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/634/m634p199.pdf.

Twist:2020:OPN

[TCM^{+}20] B. A. Twist, C. E. Cornwall, S. J. McCoy, P. W. Gabrielson, P. T. Martone, and W. A. Nelson. Opinion piece: The need to employ reliable and reproducible species identifications in coralline algal research. Marine Ecology Progress Series, 654:225–231, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/654/m654p225.pdf.

Tickler:2023:PAT

[TCM^{+}23] D. M. Tickler, A. B. Carlisle, J. J. Meeuwig, T. K. Chapple, D. J. Curnick, J. J. Dale, M. J. Castleton, R. J. Schallert, and B. A. Block. Pop-up archival tags reveal environmental influences on the vertical movements of silvertip sharks Carcharhinus albimarginatus. Marine Ecology Progress Series, 717:85–105, August 24, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/717/m717p085.pdf.

Thorne:2023:EWM

[TCP^{+}23] L. H. INTRODUCTION: REVIEW Thorne, T. A. Clay, R. A. Phillips, L. G. Silvers, and E. D. Wakefield. Effects of wind on the movement, behavior, energetics, and life history of seabirds. Marine Ecology Progress Series, 723:73–117, November 23, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m723p073.pdf.

Tettelbach:2023:CNY

[TCT^{+}23] S. T. Tettelbach, R. E. Czaja, Jr., H. Tobi, S. W. T. Hughes, B. J. Peterson, S. M. Heck, J. MacGregor, F. DeLany, B. J. Scannell, E. Pales Espinosa, and B. Allam. Collapse of the New York Bay scallop fishery despite sustained larval and juvenile recruitment. Marine Ecology Progress Series, 714: 45–56, July 13, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m714p045.pdf.

Tone:2022:TWH

[TCY^{+}22] K. Tone, W. C. Chiang, H. M. Yeh, S. T. Hsiao, C. H. Li, K. Komeyama, K. Kudo, T. Hasegawa, T. Sakamoto, I. Nakamura, Y. Sakakura, and R. Kawabe. Two-way habitat use between reefs and open ocean in adult greater amberjack: evidence from biologging data. Marine Ecology Progress Series, 699:135–151, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m699p135.pdf.

Timbs:2021:LSV

[TD21] R. Timbs and M. J. Durako. Landscape-scale variation in a sulfur-based sediment stress indicator for the seagrass Thalassia testudinum in Florida Bay, USA. Marine Ecology Progress Series, 670:33–47, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/670/m670p033.pdf.

Tuffley:2021:SML

[TdLHL21] E. J. Tuffley, S. de Lestang, J. How, and T. Langlois. Size matters: large spiny lobsters reduce the catchability of small conspecifics. Marine Ecology Progress Series, 666:99–113, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m666p099.pdf.

Tarrant:2021:LRG

[TEK21] A. M. Tarrant, L. B. Eisner, and D. G. Kimmel. Lipid-related gene expression and sensitivity to starvation in Calanus glacialis in the eastern Bering Sea. Marine Ecology Progress Series, 674:73–88, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/674/m674p073.pdf.

Taylor:2022:BCC

[TFCS22] D. L. Taylor, M. M. Fehon, K. J. Cribari, and A. K. Scro. Blue crab Callinectes sapidus dietary habits and predation on juvenile winter flounder Pseudopleuronectes americanus in southern New England tidal rivers. Marine Ecology Progress Series, 681:145–167, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2022/681/m681p145.pdf.

Twiname:2022:RLD

[TFH^{+}22] S. Twiname, Q. P. Fitzgibbon, A. J. Hobday, C. G. Carter, M. Oellermann, and G. T. Pecl. Resident lobsters dominate food competition with range-shifting lobsters in an ocean warming hotspot. Marine Ecology Progress Series, 685:171– 181, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m685p171.pdf.

Trehern:2020:LSN

[TGB^{+}20] R. H. Trehern, A. Garg, W. B. Bigelow, H. Hauptman, A. Brooks, L. A. Hawkes, and T. E. Van Leeuwen. Low salinity negatively affects metabolic rate, food consumption, digestion and growth in invasive lionfish Pterois spp. Marine Ecology Progress Series, 644:157–171, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/644/m644p157.pdf.

Toumi:2023:TSL

[TGT^{+}23] C. Toumi, O. Gauthier, É. Thiébaut, C. Guedes, and J. Grall. Taxonomic surrogates for long-term macrobenthic community monitoring: an application with community trajectory analysis. Marine Ecology Progress Series, 714:105–111, July 13, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m714p105.pdf.

Thatje:2021:OPC

[Tha21] S. Thatje. Opinion piece: Climate warming affects the depth distribution of marine ectotherms. Marine Ecology Progress Series, 660:233–240, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2021/660/m660p233.pdf. See erra-

tum [Ano21i].

Tanita:2022:MMU

[THD^{+}22] I. Tanita, E. B. Hirohavi, B. A. Diau, K. Masaki, T. Komatsu, and C. Ramofafia. Macro- and microhabitat use patterns of holothurians in Buena Vista, Nggela, Solomon Islands: rock climber and bottom crawler species. Marine Ecology Progress Series, 687:79–93, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/687/m687p079.pdf.

Takeuchi:2022:EDD

[THK^{+}22] A. Takeuchi, T. Higuchi, M. Kuroki, S. Watanabe, M. J. Miller, T. Okino, T. Miwa, and K. Tsukamoto. Environmental DNA detects a possible Japanese eel spawning event near a video-recorded anguillid eel in the open ocean. Marine Ecology Progress Series, 689:95–107, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/689/m689p095.pdf.

Taira:2020:EEA

[THL^{+}20] D. Taira, E. C. Heery, L. H. L. Loke, A. Teo, A. G. Bauman, and P. A. Todd. Ecological engineering across organismal scales: trophic-mediated positive effects of microhabitat enhancement on fishes. Marine Ecology Progress Series, 656:181–192, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m656p181.pdf.

Tracy:2023:ORH

[THR^{+}23] A. M. Tracy, K. Heggie, C. Ritter, D. Norman, R. Aguilar, and M. B. Ogburn. Oyster reef habitat depends on environmental conditions and management across large spatial scales. Marine Ecology Progress Series, 721:103–117, October 19, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m721p103.pdf.

Takahashi:2024:IVG

[THS^{+}24] M. Takahashi, T. Higuchi, K. Shirai, S. i. Ito, and M. Yoda. Interdecadal variabilities in growth and temperature trajectories of Trachurus japonicus juveniles: 1960s–1970s versus 2000s–2010s. Marine Ecology Progress Series, 741:301–313, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/741/m741p301.pdf.

Teramura:2022:AED

[TKM^{+}22] A. Teramura, K. Koeda, A. Matsuo, M. P. Sato, H. Senou, H. C. Ho, Y. Suyama, K. Kikuchi, and S. Hirase. Assessing the effectiveness of DNA barcoding for exploring hidden genetic diversity in deep-sea fishes. Marine Ecology Progress Series, 701:83–98, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/701/m701p083.pdf.

Twist:2020:BSR

[TKP^{+}20] B. A. Twist, A. Kluibenschedl, D. Pritchard, M. J. Desmond, R. D’Archino, W. A. Nelson, and C. D. Hepburn. Biomass and species richness relationships in macroalgal communities that span intertidal and subtidal zones. Marine Ecology Progress Series, 654:67–78, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/654/m654p067.pdf.

Tawa:2020:FSH

[TKS^{+}20] A. Tawa, T. Kodama, K. Sakuma, T. Ishihara, and S. Ohshimo. Fine-scale horizontal distributions of multiple species of larval tuna off the Nansei Islands, Japan. Marine Ecology Progress Series, 636:123–137, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/636/
m636p123.pdf.

Trzcinski:2024:DAS

[TMN^{+}24] M. K. Trzcinski, S. Majewski, C. A. Nordstrom, A. D. Schulze, K. M. Miller, and S. Tucker. DNA analysis of scats reveals spatial and temporal structure in the diversity of harbour seal diet from local haulouts to oceanographic bioregions. Marine Ecology Progress Series, 743:113–138, August 22, 2024.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m743p113.pdf.

Teles-Machado:2024:AOC

[TMPPG24] A. Teles-Machado, S. M. Plecha, A. Peliz, and S. Garrido. Anomalous ocean currents and European anchovy dispersal in the Iberian ecosystem. Marine Ecology Progress Series, 741: 289–300, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/741/m741p289.pdf.

Thorvaldsen:2024:HPS

[TNMN24] K. G. Thorvaldsen, S. Neuenfeldt, P. Mariani, and J. R. Nielsen. Hiding in plain sight: predator avoidance behaviour of mesopelagic Maurolicus muelleri during foraging. Marine Ecology Progress Series, 741:189–202, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/741/
m741p189.pdf.

Tiedemann:2022:ASR

[TNS^{+}22] M. Tiedemann, I. Ndour, F. N. Sow, E. Bagøien, J. O. Krakstad, M. Ostrowski, E. K. Stenevik, T. Ensrud, and S. Isari. Asynchronized spawning responses of small pelagic fishes to a short-term environmental change. Marine Ecology Progress Series, 696:85–102, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m696p085.pdf.

Trenkel:2023:RHI

[TOMDC23] V. M. Trenkel, H. Ojaveer, D. C. M. Miller, and M. Dickey-Collas. The rationale for heterogeneous inclusion of ecosystem trends and variability in ICES fishing opportunities advice. Marine Ecology Progress Series, 704:81–97, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/704/
m704p081.pdf.

Taylor:2020:DWP

[TP20] R. B. Taylor and S. Patke. Decreasing water pressure cues ‘bailout’ in seaweed-associated crustaceans. Marine Ecology Progress Series, 653:121–129, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/653/m653p121.pdf.

Taylor:2023:HTN

[TPGG^{+}23] D. L. Taylor, C. Pearson, L. Green-Gavrielidis, N. V. Hobbs, C. Thornber, G. Cicchetti, A. Gerber-Williams, and M. C. McManus. Habitat and trophic niche overlap among juvenile black sea bass, tautog, and cunner: interspecific interactions amid a species geographic range expansion. Marine Ecology Progress Series, 720:133–159, October 5, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/720/
m720p133.pdf.

Tatsch:2021:BGD

[TPZ^{+}21] A. C. C. Tatsch, M. C. Proietti, R. Zanini, P. F. Fruet, and E. R. Secchi. Beyond genetic differences: epigenetic variation in common bottlenose dolphins Tursiops truncatus from contrasting marine ecosystems. Marine Ecology Progress Series, 671:219–233, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/671/m671p219.pdf.

Tauran:2024:PMD

[TQG24] A. Tauran, N. Quillien, and J. Grall. Patterns in macrobenthic diversity in the lower shore of northeastern Atlantic macrotidal sandy beaches. Marine Ecology Progress Series, 738:21– 40, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/738/m738p021.pdf.

Turrietta:2024:EEF

[TR24] E. Turrietta and M. A. Reidenbach. Edge effects of a fragmented seagrass habitat on flow, bivalve recruitment, and sediment dynamics. Marine Ecology Progress Series, 732:53– 71, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/732/m732p053.pdf.

Trabue:2022:STB

[TRK^{+}22] S. G. Trabue, M. L. Rekdahl, C. D. King, S. Strindberg, S. K. Adamczak, and H. C. Rosenbaum. Spatiotemporal trends in bottlenose dolphin foraging behavior and relationship to environmental variables in a highly urbanized estuary. Marine Ecology Progress Series, 690:219–235, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/690/
m690p219.pdf.

Tolimieri:2023:CKF

[TSS^{+}23] N. Tolimieri, A. O. Shelton, J. F. Samhouri, C. J. Harvey, B. E. Feist, G. D. Williams, K. S. Andrews, K. E. Frick, S. Lonhart, G. Sullaway, O. Liu, H. D. Berry, and J. Waddell. Changes in kelp forest communities off Washington, USA, during and after the 2014–2016 marine heatwave and sea star wasting syndrome. Marine Ecology Progress Series, 703:47–66, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2022/703/m703p047.pdf.

Temple:2021:RTF

[TSW^{+}21] N. A. Temple, E. L. Sparks, B. M. Webb, J. Cebrian, M. F. Virden, A. E. Lucore, and H. B. Moss. Responses of two fringing salt marsh plant species along a wave climate gradient. Marine Ecology Progress Series, 675:53–66, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m675p053.pdf.

Thaxton:2020:CAT

[TTA20] W. C. Thaxton, J. C. Taylor, and R. G. Asch. Climateassociated trends and variability in ichthyoplankton phenology from the longest continuous larval fish time series on the east coast of the United States. Marine Ecology Progress Series, 650:269–287, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/650/m650p269.pdf.

Thomas:2020:STR

[TTJS20] O. R. B. Thomas, K. V. Thomas, G. P. Jenkins, and S. E. Swearer. Spatio-temporal resolution of spawning and larval nursery habitats using otolith microchemistry is element dependent. Marine Ecology Progress Series, 636:169–187, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/636/m636p169.pdf.

Tomillo:2022:ECM

[TTM^{+}22] P. Santidrián Tomillo, J. Tomás, A. Marco, A. Panagopoulou, and G. Tavecchia. Environmental changes in the Mediterranean Sea could facilitate the western expansion of loggerhead turtles. Marine Ecology Progress Series, ??:??, November 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Tomillo:2024:ECM

[TTM^{+}24] P. Santidrian Tomillo, J. Tomás, A. Marco, A. Panagopoulou, and G. Tavecchia. Environmental changes in the Mediterranean Sea could facilitate the western expansion of loggerhead turtles. Marine Ecology Progress Series, 728:145– 161, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/728/m728p145.pdf.

Thomas:2020:RIN

[TWC20] T. N. Thomas, W. H. Watson III, and C. C. Chabot. The relative influence of nature vs. nurture on the expression of circatidal rhythms in the American horseshoe crab Limulus polyphemus. Marine Ecology Progress Series, 649:83– 96, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/649/m649p083.pdf.

Troast:2022:MTA

[TWC22] B. V. Troast, L. J. Walters, and G. S. Cook. A multi-tiered assessment of fish community responses to habitat restoration in a coastal lagoon. Marine Ecology Progress Series, 698:1–14, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Trevail:2023:MCT

[TWC^{+}23] A. M. Trevail, H. Wood, P. Carr, R. E. Dunn, M. A. Nicoll, S. C. Votier, and R. Freeman. Multi-colony tracking reveals segregation in foraging range, space use, and timing in a tropical seabird. Marine Ecology Progress Series, 724:155– 165, December 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/724/m724p155.pdf.

Tian:2024:PSS

[TYL^{+}24] X. Tian, Y. Ye, Y. Li, L. Diao, Y. Wang, X. Yang, and X. Zhang. Phytoplankton size structures and their interactive effects with nutrient conditions on sinking rates in a mariculture system. Marine Ecology Progress Series, 747:35– 48, October 17, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/747/m747p035.pdf.

Uth:2024:PCC

[UAL24] C. Uth, E. Asmala, and A. M. Lewandowska. Phytoplankton community composition as a driver of annual autochthonous organic carbon dynamics in the northern coastal Baltic Sea. Marine Ecology Progress Series, 745:13–24, September 19, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps_oa/m745p013.pdf.

Uboldi:2024:ACE

[UBO^{+}24] T. Uboldi, G. Bridier, F. Olivier, J. Grall, L. Chauvaud, and R. Tremblay. Acclimation capacity of Echinarachnius parma sand dollars enduring a highly seasonal sub-Arctic coastal system. Marine Ecology Progress Series, 747:83–97, October 17, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/747/m747p083.pdf.

Uematsu:2024:DNO

[UIS^{+}24] Y. Uematsu, T. Ishihara, T. Shimose, K. S. Chen, J. A. Mohan, J. R. Rooker, R. J. D. Wells, O. E. Snodgrass, H. Dewar, S. Ohshimo, and Y. Tanaka. Distinct natal origins based on vertebral ring analysis corroborate the migration pattern of Pacific bluefin tuna in the North Pacific Ocean. Marine Ecology Progress Series, 743:65–74, August 22, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/743/
m743p065.pdf.

Uthicke:2020:KBR

[UPK^{+}20] S. Uthicke, F. Patel, S. Karelitz, H. M. Luter, N. S. Webster, and M. Lamare. Key biological responses over two generations of the sea urchin Echinometra sp. a under future ocean conditions. Marine Ecology Progress Series, 637:87–101, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m637p087.pdf.

Usov:2024:JSD

[URSS24] N. Usov, I. Radchenko, V. Smirnov, and A. Sukhotin. Joint seasonal dynamics of phytoplankton and zooplankton in the sub-Arctic White Sea. Marine Ecology Progress Series, 732: 33–51, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/732/m732p033.pdf.

Vieira:2022:TMC

[VAMF22] M. Vieira, M. C. P. Amorim, T. A. Marques, and P. J. Fonseca. Temperature mediates chorusing behaviour associated with spawning in the sciaenid Argyrosomus regius. Marine Ecology Progress Series, 697:109–124, September 22, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/697/m697p109.pdf.

Vogel:2021:KWM

[VBB^{+}21] E. F. Vogel, M. Biuw, M. A. Blanchet, I. D. Jonsen, E. Mul, E. Johnsen, S. S. Hjøllo, M. T. Olsen, R. Dietz, and A. Rikardsen. Killer whale movements on the Norwegian shelf are associated with herring density. Marine Ecology Progress Series, 665:217–231, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/665/m665p217.pdf.

Vaudo:2024:IVH

[VDB^{+}24] J. J. Vaudo, H. Dewar, M. E. Byrne, B. M. Wetherbee, and M. S. Shivji. Integrating vertical and horizontal movements of shortfin mako sharks Isurus oxyrinchus in the eastern North Pacific Ocean. Marine Ecology Progress Series, 732: 85–99, March 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/732/m732p085.pdf.

Ventura:2024:DRS

[VDC^{+}24] D. Ventura, S. F. Dubois, E. Cardella, F. Colloca, F. Tiralongo, A. Rakaj, A. Bonifazi, G. Jona Lasinio, E. Mancini, I. Bertocci, E. Casoli, G. Mancini, T. Valente, A. Belluscio, G. Ardizzone, and M. F. Gravina. Defining the role of Sabellaria alveolata reefs as nursery areas for juvenile fish: first evidence from drone-based imagery and underwater visual census data. Marine Ecology Progress Series, 735:103– 123, May 2, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/735/m735p103.pdf.

Viejo:2024:DFS

[VDG24] R. M. Viejo, M. Des, and D. Gutiérrez. Drivers of the finescale distribution of a canopy-forming seaweed at the southern edge of its range. Marine Ecology Progress Series, 727:91– 109, January 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/727/m727p091.pdf.

vonderHeyden:2024:NDR

[vdHMA^{+}24] S. REVIEW von der Heyden, R. P. Mofokeng, J. B. Adams, G. F. Midgley, D. Pillay, K. M. Watson, B. J. Boshoff, A. Bossert, C. Combrink, T. Engelbrecht, S. Govender, M. F. Mokumo, A. Ndhlovu, and A. Searle. Nine decades of research on Zostera capensis: from foundational science to conservation and resilience. Marine Ecology Progress Series, 748:197– 212, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/748/m748p197.pdf.

vanderKooij:2024:NRE

[vdKMC^{+}24] J. van der Kooij, N. McKeown, F. Campanella, G. Boyra, M. Doray, M. Santos Mocoroa, J. Fernandes da Silva, and M. Huret. Northward range expansion of Bay of Biscay anchovy into the English Channel. Marine Ecology Progress Series, 741:217–236, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m741p217.pdf.

vanderReis:2020:FHF

[vdRJL20] A. L. van der Reis, A. G. Jeffs, and S. D. Lavery. From feeding habits to food webs: exploring the diet of an opportunistic benthic generalist. Marine Ecology Progress Series, 655:107–121, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/655/m655p107.pdf.

vanDenderen:2022:EBT

[vDTS^{+}22] P. D. van Denderen, A. Törnroos, M. Sciberras, H. Hinz, R. Friedland, R. Lasota, M. C. Mangano, C. Robertson, S. Valanko, and J. G. Hiddink. Effects of bottom trawling and hypoxia on benthic invertebrate communities. Marine Ecology Progress Series, 694:13–27, August 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

vanderVeer:2022:CFL

[vdVTW^{+}22] H. W. van der Veer, I. Tulp, J. I. J. Witte, S. S. H. Poiesz, and L. J. Bolle. Changes in functioning of the largest coastal North Sea flatfish nursery, the Wadden Sea, over the past half century. Marine Ecology Progress Series, 693:183–201, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m693p183.pdf.

vanderVelde:2021:SIS

[vdVVC^{+}21] T. D. van der Velde, W. N. Venables, P. J. Crocos, S. Edgar, F. Evans, and P. C. Rothlisberg. Seasonal, interannual and spatial variability in the reproductive dynamics of Penaeus merguiensis. Marine Ecology Progress Series, 658:117–133, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2021/658/m658p117.pdf.

vanderVeer:2024:STV

[vdVWF^{+}24] H. W. van der Veer, J. I. J. Witte, P. Flege, J. van der Molen, and S. S. H. Poiesz. Spatial and temporal variability in larval connectivity of North Sea plaice Pleuronectes platessa between spawning grounds and coastal European nurseries. Marine Ecology Progress Series, 733:111–127, April 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m733p111.pdf.

vandeWolfshaar:2021:SFC

[vdWDH^{+}21] K. E. van de Wolfshaar, U. Daewel, S. S. Hjøllo, T. A. Troost, M. Kreus, J. Pätsch, R. Ji, and M. Maar. Sensitivity of the fish community to different prey fields and importance of spatial-seasonal patterns. Marine Ecology Progress Series, 680:79–95, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/680/m680p079.pdf.

vanErp:2023:TSN

[vESB^{+}23] J. van Erp, E. Sage, W. Bouten, E. van Loon, K. C. J. Camphuysen, and J. Shamoun-Baranes. Thermal soaring over the North Sea and implications for wind farm interactions. Marine Ecology Progress Series, 723:185–200, November 23, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m723p185.pdf.

Verissimo:2024:MBR

[VFG24] A. Veríssimo, P. Fonseca, and S. Garrido. Molecular barcoding reveals patterns of egg predation in small pelagic fish. Marine Ecology Progress Series, 741:113–125, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/741/
m741p113.pdf.

Valenti:2020:JFA

[VGA20] J. L. Valenti, T. M. Grothues, and K. W. Able. Juvenile fish assemblage recruitment dynamics in a mid-Atlantic estuary: before and after Hurricane Sandy. Marine Ecology Progress Series, 641:177–193, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2020/641/m641p177.pdf.

Valdivia:2023:DSR

[VGG^{+}23] N. Valdivia, I. Garrido, I. Gómez, P. Huovinen, and L. M. Pardo. Diverse stability responses of a sub-Antarctic invertebrate understorey community to experimental giant kelp removal. Marine Ecology Progress Series, 716:63–75, August 10, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/716/m716p063.pdf.

Vacquie-Garcia:2021:SHU

[VGLL^{+}21] J. Vacquié-Garcia, C. Lydersen, E. Lydersen, G. N. Christensen, C. Guinet, and K. M. Kovacs. Seasonal habitat use of a lagoon by ringed seals Pusa hispida in Svalbard, Norway. Marine Ecology Progress Series, 675:153–164, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m675p153.pdf.

VanOpzeeland:2020:YRP

[VH20] I. Van Opzeeland and H. Hillebrand. Year-round passive acoustic data reveal spatio-temporal patterns in marine mammal community composition in the Weddell Sea, Antarctica. Marine Ecology Progress Series, 638:191–206, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m638p191.pdf.

Varnerin:2020:RTC

[VHG20] B. V. Varnerin, B. M. Hopkinson, and D. F. Gleason. Recruits of the temperate coral Oculina arbuscula mimic adults in their resilience to ocean acidification. Marine Ecology Progress Series, 636:63–75, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/636/m636p063.pdf.

Virta:2023:DBD

[Vir23] L. Virta. Diversity of benthic diatoms in the Baltic Sea: alpha and beta diversity, environmental drivers, and diversity– biomass relationships. Marine Ecology Progress Series, 712: 35–47, June 8, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/712/m712p035.pdf.

Villarino:2020:RCC

[VIV^{+}20] E. Villarino, X. Irigoien, F. Villate, A. Iriarte, I. Uriarte, S. Zervoudaki, J. Carstensen, T. D. O’Brien, and G. Chust. Response of copepod communities to ocean warming in three time-series across the North Atlantic and Mediterranean Sea. Marine Ecology Progress Series, 636:47–61, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/636/
m636p047.pdf.

Valles:2021:MUR

[VK21] H. Vallès and D. L. Kramer. Microhabitat use by recently settled Sparisoma parrotfishes: ontogenetic shifts and association with algal-gardening damselfishes. Marine Ecology Progress Series, 675:97–112, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/675/m675p097.pdf.

Vereide:2024:CAS

[VKdJ24] E. H. Vereide, B. Khodabandeloo, and K. de Jong. The copepod Acartia sp. is more sensitive to a rapid pressure drop associated with seismic airguns than Calanus sp. Marine Ecology Progress Series, 730:15–30, March 7, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m730p015.pdf.

vanLeeuwen:2022:GMR

[vLBGGT22] S. M. van Leeuwen, J. A. Beecham, L. M. García-García, and R. Thorpe. The Mediterranean Rhodes Gyre: modelled impacts of climate change, acidification and fishing. Marine Ecology Progress Series, 690:31–50, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/690/
m690p031.pdf.

Ventura:2021:DRE

[VLG^{+}21] F. Ventura, P. M. Lukacs, J. P. Granadeiro, R. Matano, and P. Catry. Demographic responses to environmental change of the black-browed albatross, sentinel of the Patagonian shelf large marine ecosystem. Marine Ecology Progress Series, 668:107–120, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/668/m668p107.pdf.

Vizon:2024:SGC

[VLPN24] C. Vizon, L. Lagourgue, C. E. Payri, and M. M. Nugues. Seasonal growth and calcification of three species of crustose coralline algae in Moorea, French Polynesia. Marine Ecology Progress Series, 739:31–48, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m739p031.pdf.

vanLeeuwen:2021:CCM

[vLSB^{+}21] S. M. van Leeuwen, H. Salgado, J. L. Bailey, J. Beecham, J. L. Iriarte, L. García-García, and R. Thorpe. Climate change, marine resources and a small Chilean community: making the connections. Marine Ecology Progress Series, 680:223–246, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m680p223.pdf.

Vicente:2020:CCH

[VMCA^{+}20] P. Vicente, S. Martins-Cardoso, F. Almada, E. J. Gonçalves, and A. M. Faria. Chemical cues from habitats and conspecifics guide sand-smelt Atherina presbyter larvae to reefs. Marine Ecology Progress Series, 650:191–202, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/650/
m650p191.pdf.

Varela:2022:FGA

[VMDA22] J. L. Varela, A. Medina, S. Déniz, and F. J. Abascal. Feeding of Atlantic bluefin tuna Thunnus thynnus around the Canary Islands assessed from stomach content and stable isotope analyses. Marine Ecology Progress Series, 686:177– 186, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/686/m686p177.pdf.

Vigo:2022:UMS

[VNG^{+}22] M. Vigo, J. Navarro, J. Giménez, N. Andón, A. Martínez-Lage, J. B. Company, and G. Rotllant. Using molecular and stable isotope markers to identify the main predators of Nephrops norvegicus in Mediterranean deep-water ecosystems. Marine Ecology Progress Series, 695:95–108, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/695/m695p095.pdf.

Vigo:2021:SEN

[VNM^{+}21] M. Vigo, J. Navarro, I. Masmitja, J. Aguzzi, J. A. García, G. Rotllant, N. Bahamón, and J. B. Company. Spatial ecology of Norway lobster Nephrops norvegicus in Mediterranean deep-water environments: implications for designing no-take marine reserves. Marine Ecology Progress Series, 674:173– 188, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/674/m674p173.pdf.

Vicente:2020:IPF

[VOM^{+}20] J. Vicente, A. Osberg, M. J. Marty, K. Rice, and R. J. Toonen. Influence of palatability on the feeding preferences of the endemic Hawaiian tiger cowrie for indigenous and introduced sponges. Marine Ecology Progress Series, 647:109– 122, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/647/m647p109.pdf.

Vecchio:2021:ICL

[VOP21] J. L. Vecchio, J. L. Ostroff, and E. B. Peebles. Isotopic characterization of lifetime movement by two demersal fishes from the northeastern Gulf of Mexico. Marine Ecology Progress Series, 657:161–172, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/657/m657p161.pdf.

VanHoeck:2020:SNA

[VPB^{+}20] R. V. Van Hoeck, A. B. Paxton, D. R. Bohnenstiehl, J. C. Taylor, F. J. Fodrie, D. P. Nowacek, C. M. Voss, and C. H. Peterson. Soundscapes of natural and artificial temperate reefs: similar temporal patterns but distinct spectral content. Marine Ecology Progress Series, 649:35–51, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/649/
m649p035.pdf.

Vanderklift:2021:HRH

[VPH^{+}21] M. A. Vanderklift, R. D. Pillans, M. Hutton, L. De Wever, G. A. Kendrick, A. Zavala-Perez, A. Vergés, R. Garthwin, D. Oades, P. McCarthy, K. George, T. Sampi, D. George, C. Sampi, Z. Edgar, K. Dougal, and A. Howard. High rates of herbivory in remote northwest Australian seagrass meadows by rabbitfish and green turtles. Marine Ecology Progress Series, 665:63–73, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m665p063.pdf.

Vanharanta:2023:UEP

[VS23] M. Vanharanta and K. Spilling. Uptake of excess phosphate at low inorganic N:P ratio in a coastal sea afflicted with eutrophication. Marine Ecology Progress Series, 718:23–37, September 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m718p023.pdf.

VanAlstyne:2020:IVC

[VSG20] K. L. Van Alstyne, L. Sutton, and S. A. Gifford. Inducible versus constitutive antioxidant defenses in algae along an environmental stress gradient. Marine Ecology Progress Series, 640:107–115, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/640/m640p107.pdf.

Voss:2024:AHI

[VSGD24] J. N. Voss, E. Sandbank, R. A. Glazer, and G. A. Delgado. Aftereffects of Hurricanes Irma and Ian on queen conch Aliger gigas in the Florida Keys, USA. Marine Ecology Progress Series, 733:129–135, April 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m733p129.pdf.

Vallim:2024:MIB

[VST24] A. L. OPINION PIECE Vallim, S. Schenone, and S. F. Thrush. Megafauna: the ignored bioturbators. Marine Ecology Progress Series, 733:137–144, April 4, 2024. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m733p137.pdf.

vanTonder:2021:EML

[vTLG^{+}21] A. van Tonder, N. Lübcker, M. Guerreiro, J. C. Xavier, Y. Cherel, and P. J. N. de Bruyn. Ecology of Moroteuthopsis longimana at the sub-Antarctic Prince Edward Islands, revealed through stable isotope analysis of squid beaks. Marine Ecology Progress Series, 658:105–115, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/658/
m658p105.pdf.

Venables:2020:HUM

[VvDRM20] S. K. Venables, D. I. van Duinkerken, C. A. Rohner, and A. D. Marshall. Habitat use and movement patterns of reef manta rays Mobula alfredi in southern Mozambique. Marine Ecology Progress Series, 634:99–114, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2019/634/m634p099.pdf.

Walker:2023:ICM

[WAA^{+}23] J. E. Walker, C. Angelini, A. H. Altieri, C. W. Martin, and T. Z. Osborne. Influence of climate-mediated shifts in wetland vegetation on prey fish habitat use and schooling dynamics. Marine Ecology Progress Series, 716:31–45, August 10, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2023/716/m716p031.pdf.

Wesselmann:2024:SRS

[WAA24] M. INTRODUCTION Wesselmann, E. T. Apostolaki, and A. Anton. Species range shifts, biological invasions and ocean warming. Marine Ecology Progress Series, 728:81–83, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m728p081.pdf.

Wilber:2022:OWF

[WBG^{+}22] D. H. Wilber, L. Brown, M. Griffin, G. R. DeCelles, and D. A. Carey. Offshore wind farm effects on flounder and gadid dietary habits and condition on the northeastern US coast. Marine Ecology Progress Series, 683:123–138, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m683p123.pdf.

Wilber:2024:ALH

[WBGC24] D. H. Wilber, L. J. Brown, M. Griffin, and D. A. Carey. American lobster Homarus americanus responses to construction and operation of an offshore wind farm in southern New England. Marine Ecology Progress Series, 727:123–142, January 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/727/m727p123.pdf.

Waite:2024:NDA

[WBS24] H. R. Waite, R. A. Beshai, and C. J. B. Sorte. NOTE: Demography across latitudinal and elevational gradients for rangeexpanding whelks. Marine Ecology Progress Series, 728:115– 121, February 8, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/728/m728p115.pdf.

Weinstock:2021:HWA

[WC21] J. B. Weinstock and R. Collin. Hypoxia and warming are associated with reductions in larval bivalve abundance in a tropical lagoon. Marine Ecology Progress Series, 662:85– 95, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/662/m662p085.pdf.

Whiting:2020:SNS

[WCL20] A. U. Whiting, M. Chaloupka, and C. J. Limpus. Sampling nesting sea turtles: impact of survey error on trend detection. Marine Ecology Progress Series, 634:213–223, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2019/634/
m634p213.pdf.

Whiting:2021:SNS

[WCP^{+}21a] A. U. Whiting, M. Chaloupka, N. Pilcher, P. Basintal, and C. J. Limpus. Sampling nesting sea turtles: optimizing survey design to minimize error. Marine Ecology Progress Series, 674:257–270, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/674/m674p257.pdf.

Williams:2021:SUM

[WCP^{+}21b] J. P. Williams, J. T. Claisse, D. J. Pondella II, C. M. Williams, M. J. Robart, Z. Scholz, E. M. Jaco, T. Ford, H. Burdick, and D. Witting. Sea urchin mass mortality rapidly restores kelp forest communities. Marine Ecology Progress Series, 664:117–131, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/664/m664p117.pdf.

Wernberg:2020:DID

[WCRTT20] T. Wernberg, M. Couraudon-Réale, F. Tuya, and M. Thomsen. Disturbance intensity, disturbance extent and ocean climate modulate kelp forest understory communities. Marine Ecology Progress Series, 651:57–69, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m651p057.pdf.

Wang:2021:MMI

[WCS^{+}21] C. Wang, C. Chen, R. Su, Z. Luo, L. Mao, and Y. Zhang. Mechanism for the marked increase of Ulva prolifera in the south Yellow Sea: role of light intensity, nitrogen, phosphorus, and co-limitations. Marine Ecology Progress Series, 671: 97–110, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/671/m671p097.pdf.

White:2022:PCC

[WD22] L. White and D. Davoult. Photosynthetic capacity of cooccurring kelp species revealed by in situ measurements. Marine Ecology Progress Series, 697:31–43, September 22, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/697/m697p031.pdf.

Weimerskirch:2020:SMW

[WdGR^{+}20] H. Weimerskirch, S. de Grissac, A. Ravache, A. Prudor, A. Corbeau, B. C. Congdon, F. McDuie, K. Bourgeois, S. Dromzée, J. Butscher, C. Menkes, V. Allain, E. Vidal, A. Jaeger, and P. Borsa. At-sea movements of wedge-tailed shearwaters during and outside the breeding season from four colonies in New Caledonia. Marine Ecology Progress Series, 633:225–238, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m633p225.pdf.

Weil:2020:FST

[WDJ20] J. Weil, W. D. P. Duguid, and F. Juanes. Fine-scale taxonomic and temporal variability in the energy density of invertebrate prey of juvenile Chinook salmon Oncorhynchus tshawytscha. Marine Ecology Progress Series, 655:185–198, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/655/
m655p185.pdf.

Weidberg:2021:SPS

[WDP^{+}21] N. Weidberg, C. DiBacco, C. Pezzola, E. Rebiffe, and S. L. Basedow. Swimming performance of subarctic Calanus spp. facing downward currents. Marine Ecology Progress Series, 665:47–61, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m665p047.pdf.

Wright:2021:PPC

[WF21] L. S. Wright and A. Foggo. Photosynthetic pigments of cooccurring Northeast Atlantic Laminaria spp. are unaffected by decomposition. Marine Ecology Progress Series, 678:227– 232, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/678/m678p227.pdf.

Welicky:2023:RTP

[WFR^{+}23] R. L. Welicky, M. L. Feddern, T. Rolfe, K. Leazer, A. Moosmiller, E. Fiorenza, K. P. Maslenikov, L. Tornabene, G. W. Holtgrieve, and C. L. Wood. Reconstructing trophic position over the past century for five Puget Sound fish species. Marine Ecology Progress Series, 706:1–15, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Winton:2023:OSC

[WFS23] M. V. Winton, G. Fay, and G. B. Skomal. An open spatial capture-recapture framework for estimating the abundance and seasonal dynamics of white sharks at aggregation sites. Marine Ecology Progress Series, 715:1–25, July 27, 2023. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wyeth:2022:EHA

[WGK22] A. C. Wyeth, D. Grünbaum, and J. E. Keister. Effects of hypoxia and acidification on Calanus pacificus: behavioral changes in response to stressful environments. Marine Ecology Progress Series, 697:15–29, September 22, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m697p015.pdf.

Woehler:2024:RIM

[WH24] E. J. Woehler and A. J. Hobday. Review: Impacts of marine heatwaves may be mediated by seabird life history strategies. Marine Ecology Progress Series, 737:9–23, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m737p009.pdf.

Wang:2020:COC

[WHB^{+}20] Z. Wang, R. Horwitz, H. D. Bowlby, F. Ding, and W. N. Joyce. Changes in ocean conditions and hurricanes affect porbeagle Lamna nasus diving behavior. Marine Ecology Progress Series, 654:219–224, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m654p219.pdf.

Whitfield:2020:LHM

[Whi20] A. K. Whitfield. Littoral habitats as major nursery areas for fish species in estuaries: a reinforcement of the reduced predation paradigm. Marine Ecology Progress Series, 649:219–234, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/649/m649p219.pdf. See comment

[BS21].

Whitfield:2021:PSJ

[Whi21] A. Whitfield. Predation on small juvenile fishes in shallow estuarine nursery areas: Reply to Baker & Sheaves (2021). Marine Ecology Progress Series, 662:209–214, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/662/
m662p209.pdf.

Wilson:2022:LTF

[WHN^{+}22] R. P. Wilson, M. D. Holton, A. Neate, M. Del’Caño, F. Quintana, K. Yoda, and A. Gómez-Laich. Luck and tactics in foraging success: the case of the imperial shag. Marine Ecology Progress Series, 682:1–12, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Whitesell:2022:DIP

[WHRC22] M. J. Whitesell, E. A. Hunter, D. C. Rostal, and J. M. Carroll. Direct and indirect pathways for environmental drivers of hatching success in the loggerhead sea turtle. Marine Ecology Progress Series, 701:119–132, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/701/m701p119.pdf.

Woo:2023:HUN

[WIM23] K. Y. Woo, S. Isojunno, and P. J. O. Miller. Habitat use of the northern bottlenose whale Hyperoodon ampullatus near Jan Mayen, North Atlantic. Marine Ecology Progress Series, 718:119–136, September 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2023/718/m718p119.pdf.

Washburn:2023:IFD

[WIY^{+}23] T. W. Washburn, A. Iguchi, K. Yamaoka, M. Nagao, Y. Onishi, T. Fukuhara, Y. Yamamoto, and A. Suzuki. Impacts of the first deep-sea seafloor massive sulfide mining excavation tests on benthic communities. Marine Ecology Progress Series, 712:1–19, June 8, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Williams:2021:REI

[WJ21] B. L. Williams and D. S. Johnson. Role of ecological interactions in saltmarsh geomorphic processes. Marine Ecology Progress Series, 658:149–161, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/658/m658p149.pdf.

Weitzman:2021:BCS

[WK21] B. Weitzman and B. Konar. Biological correlates of sea urchin recruitment in kelp forest and urchin barren habitats. Marine Ecology Progress Series, 663:115–125, 2021. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m663p115.pdf.

Wirsing:2022:ERI

[WKAH22] A. J. Wirsing, J. J. Kiszka, A. C. Allen, and M. R. Heithaus. Ecological roles and importance of sea cows (order:

Sirenia): a review and prospectus. Marine Ecology Progress Series, 689:191–215, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m689p191.pdf.

Wee:2021:ETS

[WKR21] H. B. Wee, Y. Kobayashi, and J. D. Reimer. Effects of temperature, salinity, and depth on Symbiodiniaceae lineages hosted by Palythoa tuberculosa near a river mouth. Marine Ecology Progress Series, 667:43–60, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2021/667/m667p043.pdf.

Wong:2021:ATC

[WLA^{+}21] J. B. Wong, S. Lisovski, R. T. Alisauskas, W. English, M. A. Giroux, A. L. Harrison, D. Kellett, N. Lecomte, M. Maftei, A. Nagy-MacArthur, R. A. Ronconi, P. A. Smith, M. L. Mallory, and M. Auger-Méthé. Arctic terns from circumpolar breeding colonies share common migratory routes. Marine Ecology Progress Series, 671:191–206, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/671/
m671p191.pdf.

Wu:2020:EOD

[WLB^{+}20] W. Wu, J. Liu, A. F. Bouwman, J. Wang, X. Yin, J. Zang, and X. Ran. Exploring oxygen dynamics and depletion in an intensive bivalve production area in the coastal sea off Rushan Bay, China. Marine Ecology Progress Series, 649: 53–65, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/649/m649p053.pdf.

Waltham:2023:BCS

[WLB23] N. J. Waltham, C. Lovelock, and C. A. Buelow. Blue carbon stocks and cycling in tropical tidal marshes facing grazing pressure. Marine Ecology Progress Series, 717:1–16, August 24, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Watson:2022:IBB

[WLC22] W. H. Watson III, A. G. Lemmon, and C. C. Chabot. The impacts of biomedical bleeding on locomotion and mating behavior in the horseshoe crab, Limulus polyphemus. Marine Ecology Progress Series, 699:65–74, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/699/
m699p065.pdf.

Weigel:2021:ECP

[WMKV21] B. Weigel, J. Mäkinen, M. Kallasvuo, and J. Vanhatalo. Exposing changing phenology of fish larvae by modeling climate effects on temporal early life-stage shifts. Marine Ecology Progress Series, 666:135–148, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m666p135.pdf.

Warren:2020:STR

[WMT20] V. E. Warren, P. J. O. Miller, and P. L. Tyack. Short-term responses of sperm whales Physeter macrocephalus to the attachment of suction cup tags. Marine Ecology Progress Series, 645:219–234, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2020/645/m645p219.pdf.

Wepfer:2020:MES

[WNH^{+}20] P. H. Wepfer, Y. Nakajima, F. K. C. Hui, S. Mitarai, and E. P. Economo. Metacommunity ecology of Symbiodiniaceae hosted by the coral Galaxea fascicularis. Marine Ecology Progress Series, 633:71–87, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m633p071.pdf.

Wagner:2024:RSM

[WPG^{+}24] E. L. Wagner, S. F. Pearson, T. P. Good, P. J. Hodum, E. R. Buhle, and M. B. Schrimpf. Resilience to a severe marine heatwave at two Pacific seabird colonies. Marine Ecology Progress Series, 737:101–120, June 6, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/737/m737p101.pdf.

Wilson:2022:IIP

[WPM22] S. M. Wilson, D. A. Patterson, and J. W. Moore. Intraand inter-population variation in sensitivity of migratory sockeye salmon smolts to phenological mismatch. Marine Ecology Progress Series, 692:119–136, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m692p119.pdf.

Wong:2021:SSR

[WRB21] R. J. Wong, M. S. Roy, and J. E. K. Byrnes. Sediment selection: range-expanding fiddler crabs are better burrowers than their historic-range counterparts. Marine Ecology Progress Series, 674:163–171, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m674p163.pdf.

Weisberg:2024:GSI

[WRG^{+}24] S. J. Weisberg, S. M. Roberts, L. K. Gruenburg, T. G. Schwemmer, T. Menz, I. F. Fenwick, J. A. Nye, and R. G. Asch. Gulf Stream intrusions associated with extreme seasonal fluctuations among larval fishes. Marine Ecology Progress Series, 739:157–172, July 4, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m739p157.pdf.

Weinstock:2023:PCF

[WRS^{+}23] J. B. Weinstock, N. Reyns, M. Swiderski, D. Flerchinger, and J. Pineda. Periodicity in Chthamalus fissus reproduction does not guarantee periodicity in settlement. Marine Ecology Progress Series, 718:53–67, September 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2023/718/
m718p053.pdf.

Walsh:2020:FAK

[WRW20] J. Walsh, C. S. Reiss, and G. M. Watters. Flexibility in Antarctic krill Euphausia superba decouples diet and recruitment from overwinter sea-ice conditions in the northern Antarctic Peninsula. Marine Ecology Progress Series, 642:1– 19, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Wernberg:2024:LLE

[WS24] T. Wernberg and S. C. Straub. Low light exacerbates effects of marine heatwaves on seaweeds. Marine Ecology Progress Series, 747:49–59, October 17, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps_oa/m747p049.pdf.

Wells:2024:TPF

[WSB^{+}24] B. K. Wells, J. A. Santora, J. J. Bizzarro, A. Billings, R. D. Brodeur, E. A. Daly, J. C. Field, K. E. Richerson, and J. T. Thorson. Trophoscapes of predatory fish reveal biogeographic structuring of spatial dietary overlap and inform fisheries bycatch patterns. Marine Ecology Progress Series, 741: 47–70, July 30, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/741/m741p047.pdf.

Wickham:2020:SWD

[WSD^{+}20] S. B. Wickham, N. Shackelford, C. T. Darimont, W. Nijland, L. Y. Reshitnyk, J. D. Reynolds, and B. M. Starzomski. Sea wrack delivery and accumulation on islands: factors that mediate marine nutrient permeability. Marine Ecology Progress Series, 635:37–54, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m635p037.pdf.

Wilson:2020:PGA

[WSPS^{+}20] K. C. Wilson, B. X. Semmens, C. V. Pattengill-Semmens, C. McCoy, and A. Širović. Potential for grouper acoustic competition and partitioning at a multispecies spawning site off Little Cayman, Cayman Islands. Marine Ecology Progress Series, 634:127–146, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m634p127.pdf.

Will:2021:HMR

[WSU21] V. Will, A. Stahl, and M. S. Ullrich. Heavy metal resistance in Marinobacter adhaerens HP15 supports colonization of transparent exopolymer particles during its interaction with diatoms. Marine Ecology Progress Series, 658: 47–57, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/658/m658p047.pdf.

Walsh:2024:DSB

[WTK^{+}24] K. A. Walsh, A. R. Thompson, G. T. Kwan, B. X. Semmens, H. W. Fennie, and R. Swalethorp. Diet and size at birth affect larval rockfish condition and survival. Marine Ecology Progress Series, 745:95–114, September 19, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2024/745/
m745p095.pdf.

Woods:2022:TSG

[WWH^{+}22] B. L. Woods, A. Walters, M. Hindell, A. T. Revill, I. Field, S. A. McCormack, Y. Cherel, and R. Trebilco. Trophic structure of Southern Ocean squid: a cross-basin analysis of stable isotopes in archived beaks from predator stomachs. Marine Ecology Progress Series, 685:137–152, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/685/
m685p137.pdf.

Wang:2024:SSV

[WXZ^{+}24] J. Wang, B. Xu, C. Zhang, Y. Ji, Y. Xue, and Y. Ren. Spatio-seasonal variations in functional trait composition and diversity patterns of marine fish communities in coastal waters. Marine Ecology Progress Series, 744:115–131, September 5, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2024/744/m744p115.pdf.

Warlick:2020:UBS

[WYO^{+}20] A. J. Warlick, G. M. Ylitalo, S. M. O’Neill, M. B. Hanson, C. Emmons, and E. J. Ward. Using Bayesian stable isotope mixing models and generalized additive models to resolve diet changes for fish-eating killer whales Orcinus orca. Marine Ecology Progress Series, 649:189–200, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/649/
m649p189.pdf.

Watanuki:2022:SRR

[WYO^{+}22] Y. Watanuki, M. Yamamoto, J. Okado, M. Ito, and W. Sydeman. Seabird reproductive responses to changing climate and prey communities are mediated by prey packaging. Marine Ecology Progress Series, 683:179–194, 2022. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m683p179.pdf.

Wines:2020:ASS

[WYZ^{+}20] S. L. Wines, M. A. Young, R. Zavalas, J. M. Logan, P. Tinkler, and D. Ierodiaconou. Accounting for spatial scale and temporal variation in fish-habitat analyses using baited remote underwater video stations (BRUVS). Marine Ecology Progress Series, 640:171–187, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/640/m640p171.pdf.

Xing:2020:THC

[XCT^{+}20] D. Xing, B. Choi, Y. Takizawa, R. Fan, S. Sugaya, M. Tsuchiya, N. Ohkouchi, and Y. Chikaraishi. Trophic hierarchy of coastal marine fish communities viewed via compound-specific isotope analysis of amino acids. Marine Ecology Progress Series, 652:137–144, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m652p137.pdf.

Xu:2022:IDR

[XMY^{+}22] S. Xu, Z. Men, K. Yu, H. Chen, H. Mo, B. Chen, and Z. Qin. Interspecific differences in the response of coral trophic status to the decrease in symbiotic zooxanthellae in summer. Marine Ecology Progress Series, 694:73–87, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/694/
m694p073.pdf.

Xuan:2024:TES

[XW24] Z. Xuan and W. X. Wang. Trace elemental and stable isotopic signatures to reconstruct the large-scale environmental connectivity of fish populations. Marine Ecology Progress Series, 730:95–111, March 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/730/m730p095.pdf.

Xuan:2022:OMR

[XWH^{+}22] Z. Xuan, X. Wang, S. Hajisamae, K. W. K. Tsim, J. Yang, and W. X. Wang. Otolith microchemistry reveals different environmental histories for two endangered fourfinger threadfin species. Marine Ecology Progress Series, 700:161– 178, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/700/m700p161.pdf.

Xu:2021:RES

[XWW^{+}21] S. Xu, P. Wang, F. Wang, X. Zhang, X. Song, and Y. Zhou. Responses of eelgrass seed germination and seedling establishment to water depth, sediment type, and burial depth: implications for restoration. Marine Ecology Progress Series, 678:51–61, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/678/m678p051.pdf.

Xu:2020:ESM

[XYSF20] Y. Xu, S. Yao, K. Soetaert, and X. Fan. Effects of salt marsh restoration on eukaryotic microbenthic communities in the Yangtze Estuary. Marine Ecology Progress Series, 638: 39–50, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m638p039.pdf.

Yin:2021:BSC

[YAZA21] H. Yin, R. C. Aller, Q. Zhu, and J. Y. Aller. Biogenic structures and cable bacteria interactions: redox domain residence times and the generation of complex pH distributions. Marine Ecology Progress Series, 669:51–63, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2021/669/
m669p051.pdf.

Young:2023:SIN

[YBSH23] L. R. Young, D. T. Booth, C. E. Smith, and C. A. Madden Hof. Seawater irrigation on nests can increase male marine turtle production. Marine Ecology Progress Series, 718:85– 97, September 7, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m718p085.pdf.

Young:2021:DBO

[YDG21] C. S. Young, M. H. Doall, and C. J. Gobler. Dual benefit of ocean acidification for the laminarialean kelp Saccharina latissima: enhanced growth and reduced herbivory. Marine Ecology Progress Series, 664:87–102, 2021. CODEN MESEDT.

ISSN 0171-8630 (print), 1616-1599 (electronic). URL https:

//www.int-res.com/articles/meps_oa/m664p087.pdf.

Yunda-Guarin:2022:IDF

[YGMNA22] G. Yunda-Guarin, L. N. Michel, C. Nozais, and P. Archambault. Interspecific differences in feeding selectivity shape isotopic niche structure of three ophiuroids in the Arctic Ocean. Marine Ecology Progress Series, 683:81–95, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/683/
m683p081.pdf.

Yang:2023:DID

[YHW^{+}23] L. Yang, T. Hu, S. Wang, Y. Zhang, Q. Li, B. Han, and L. Zheng. Distinct interactions driven by DMSP between different microalgae and the phycosphere bacterium Sulfitobacter pseudonitzschiae H46. Marine Ecology Progress Series, 710: 15–25, May 4, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/710/m710p015.pdf.

Yates:2020:EMR

[YLH20] D. C. Yates, S. I. Lonhart, and S. L. Hamilton. Effects of marine reserves on predator–prey interactions in central California kelp forests. Marine Ecology Progress Series, 655: 139–155, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/655/m655p139.pdf.

Yamazaki:2021:CSG

[YMU^{+}21] D. Yamazaki, O. Miura, S. Uchida, M. Ikeda, and S. Chiba. Comparative seascape genetics of co-distributed intertidal snails Monodonta spp. in the Japanese and Ryukyu archipelagoes. Marine Ecology Progress Series, 657:135–146, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/657/m657p135.pdf.

Yoshida:2023:SSM

[YOI^{+}23] K. Yoshida, H. Ota, T. Iwanaga, A. Yoshitake, T. Mine, M. Omura, and K. Kimura. Species-specific monitoring of Skeletonema blooms in the coastal waters of Ariake Sound, Japan. Marine Ecology Progress Series, 703:31–46, 2023.

CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/703/m703p031.pdf.

Yetsko:2020:GDT

[YRB^{+}20] K. Yetsko, M. Ross, A. Bellantuono, D. Merselis, M. Rodriguez-Lanetty, and M. R. Gilg. Genetic differences in thermal tolerance among colonies of threatened coral Acropora cervicornis: potential for adaptation to increasing temperature. Marine Ecology Progress Series, 646:45–68, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/646/
m646p045.pdf.

Yorio:2021:SUM

[YSD^{+}21] P. Yorio, N. Suárez, P. Dell’Arciprete, C. Marinao, M. E. Góngora, L. Pichegru, L. Prosdocimi, and T. Kasinsky. Spatial use of multiple jurisdictions by Magellanic penguins and assessment of potential conflicts in the face of changing trawl fisheries scenarios. Marine Ecology Progress Series, 658:219– 236, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/658/m658p219.pdf.

Young:2022:CIP

[YSP^{+}22] M. J. Young, P. J. Seddon, K. Pütz, P. Agnew, T. Mattern, R. P. Hickcox, B. C. Robertson, and Y. van Heezik. Conservation implications for post-fledging dispersal of yelloweyed penguins/hoiho. Marine Ecology Progress Series, 695: 173–188, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2022/695/m695p173.pdf.

Yan:2024:MMA

[YWW^{+}24] Y. Yan, M. Wang, X. Wu, H. Wang, Z. Zhong, and C. Li. Mitochondrial and morphological adaptions of Lindaspio polybranchiata (Annelida: Spionidae) in the South China Sea. Marine Ecology Progress Series, 730:43–58, March 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2024/730/m730p043.pdf.

Yang:2022:SSP

[YXB22] L. Yang, X. Xu, and P. Berggren. Spotted seal Phoca largha underwater vocalisations in relation to ambient noise. Marine Ecology Progress Series, 683:209–220, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2022/683/
m683p209.pdf.

Zhao:2020:CSA

[ZBB^{+}20] M. Zhao, D. C. Behringer, J. Bojko, A. S. Kough, L. Plough, C. P. S. Tavares, A. Aguilar-Perera, O. S. Reynoso, G. Seepersad, O. Maharaj, M. B. Sanders, D. Carnales, G. Fabiano, D. Carnevia, M. A. Freeman, N. A. M. Atherley, L. D. Medero-Hernández, and E. J. Schott. Climate and season are associated with prevalence and distribution of transhemispheric blue crab reovirus (Callinectes sapidus reovirus 1). Marine Ecology Progress Series, 647:123–133, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m647p123.pdf.

Zilius:2023:RCH

[ZBB^{+}23] M. Zilius, M. Bartoli, S. Bonaglia, U. Cardini, V. G. Chiozzini, U. Marzocchi, P. C. Moraes, A. Zaiko, and E. S. Braga. Role of crab holobionts in benthic N cycling in mangroves with different trophic status. Marine Ecology Progress Series, 712: 87–99, June 8, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2023/712/m712p087.pdf.

Zbinden:2020:REB

[ZCB20] M. Zbinden and M. A. Cambon-Bonavita. Rimicaris exoculata: biology and ecology of a shrimp from deep-sea hydrothermal vents associated with ectosymbiotic bacteria. Marine Ecology Progress Series, 652:187–222, 2020. CO-DEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (elec-

tronic). URL https://www.int-res.com/articles/meps_
oa/m652p187.pdf.

Zabin:2024:EHI

[ZCBC24] C. J. Zabin, A. L. Chang, J. Blumenthal, and B. S. Cheng. Exploring high intertidal refugia as an approach for the restoration of an intertidal oyster. Marine Ecology Progress Series, 738:119–132, June 20, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps_oa/m738p119.pdf.

Zottoli:2020:MBB

[ZCF20] J. D. Zottoli, J. S. Collie, and M. J. Fogarty. Measuring the balance between fisheries catch and fish production. Marine Ecology Progress Series, 643:145–158, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

URL https://www.int-res.com/articles/meps2020/643/
m643p145.pdf.

Ziegler:2021:IMI

[ZCM^{+}21] S. L. Ziegler, L. R. Clance, A. R. McMains, M. D. Miller, and F. J. Fodrie. Influence of marsh island size on nekton communities: intermediate optima rather than Single-Large-or-Several-Small (SLOSS). Marine Ecology Progress Series, 672:45–56, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/672/m672p045.pdf.

Zamora:2020:LPU

[ZDBS20] L. N. Zamora, N. J. Delorme, M. Byrne, and M. A. Sewell. Lipid and protein utilization during lecithotrophic development in the asteroid Stegnaster inflatus, with a review of larval provisioning in lecithotrophic echinoderms. Marine Ecology Progress Series, 641:123–134, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2020/641/m641p123.pdf.

Zacchetti:2024:SCD

[ZFD^{+}24] L. Zacchetti, E. Fanelli, F. Donato, F. Domenichetti, M. Lauteri, A. Santojanni, and M. Martinelli. Seasonal changes in the dietary patterns and reproductive aspects of deep-water rose shrimp Parapenaeus longirostris in the central Adriatic Sea. Marine Ecology Progress Series, 748:99–116, November 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m748p099.pdf.

Zerebecki:2021:EFS

[ZHG^{+}21] R. A. Zerebecki, A. R. Hughes, J. Goff, T. C. Hanley, W. Scheffel, and K. L. Heck, Jr. Effect of foundation species composition and oil exposure on wetland communities across multiple trophic levels. Marine Ecology Progress Series, 662: 53–68, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2021/662/m662p053.pdf.

Zhang:2020:UTP

[ZLHH20] D. Zhang, X. Liu, M. A. Harley, and J. D. Hardege. Uridine-5’-tri-phosphate is a candidate component of the soluble sex pheromone bouquet in a marine shrimp, Lysmata wurdemanni. Marine Ecology Progress Series, 640:139–146, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2020/640/m640p139.pdf.

Zufia:2024:GMR

[ZLL^{+}24] J. Alegria Zufia, C. P. Laber, C. Legrand, E. Lindehoff, and H. Farnelid. Growth and mortality rates of picophytoplankton in the Baltic Sea Proper. Marine Ecology Progress Series, 735:63–76, May 2, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps_oa/m735p063.pdf.

Zacchetti:2022:SVF

[ZMC^{+}22] L. Zacchetti, M. Martinelli, S. Colella, A. Santojanni, and E. Fanelli. Seasonal variations in the feeding ecology of Nephrops norvegicus in the Adriatic Sea: insights from stomach contents and stable isotope analyses. Marine Ecology Progress Series, 695:109–123, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/695/m695p109.pdf.

Zhao:2021:OPD

[ZMD^{+}21] S. T. Zhao, C. J. D. Matthews, G. K. Davoren, S. H. Ferguson, and C. A. Watt. Ontogenetic profiles of dentine isotopes (δ ^{15}N and δ ^{13}C) reveal variable narwhal Monodon monoceros nursing duration. Marine Ecology Progress Series, 668:163–175, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2021/668/m668p163.pdf.

Zhang:2023:LDS

[ZPVN23] B. Zhang, H. Pethybridge, P. Virtue, and P. D. Nichols. Lipid dynamics in the southern hemisphere: a 30-year meta-analysis of marine consumers. Marine Ecology Progress Series, 710:1– 14, May 4, 2023. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic).

Zhang:2022:PDR

[ZSC^{+}22] Y. Zhang, S. Song, T. Chen, X. Sun, and C. Li. Population dynamics and reproduction of the heterotrophic dinoflagellate Noctiluca scintillans in Jiaozhou Bay, China. Marine Ecology Progress Series, 693:55–68, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2022/693/m693p055.pdf.

Zarrella-Smith:2022:SEM

[ZSWR^{+}22] K. A. Zarrella-Smith, J. N. Woodall, A. Ryan, N. B. Furey, and J. S. Goldstein. Seasonal estuarine movements of green crabs revealed by acoustic telemetry. Marine Ecology Progress Series, 681:129–143, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-

res.com/articles/meps2022/681/m681p129.pdf.

Zuercher:2022:PBC

[Zue22] R. Zuercher. Pelagic-benthic coupling in kelp forests of central California. Marine Ecology Progress Series, 682:79–96, 2022. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599

(electronic). URL https://www.int-res.com/articles/
meps2022/682/m682p079.pdf.

Zhang:2021:MSV

[ZWXC21] H. R. Zhang, Y. Wang, P. Xiu, and F. Chai. Modeling the seasonal variability of phytoplankton in the subarctic northeast Pacific Ocean. Marine Ecology Progress Series, 680: 33–50, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m680p033.pdf.

Zhang:2024:UDC

[ZXL^{+}24] Y. Zhang, M. Xu, Z. Liu, Y. Jin, and S. Li. Unraveling dispersal of a coastal fish species in the juvenile life stage in the Yellow Sea using otolith chemistry. Marine Ecology Progress Series, 740:109–121, July 25, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.

int-res.com/articles/meps2024/740/m740p109.pdf.

Zhang:2024:AFA

[ZZLS24] H. Zhang, G. Zhu, H. Liu, and K. M. Swadling. Autumn food availability in Bransfield Strait for Antarctic krill Euphausia superba and the relationship between body size and fatty acid content. Marine Ecology Progress Series, 730:31– 42, March 7, 2024. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.

com/articles/meps2024/730/m730p031.pdf.

Zhang:2020:DRE

[ZZX^{+}20] X. Zhang, Y. Zhou, S. Xu, P. Wang, P. Zhao, S. Yue, R. Gu, X. Song, S. Xu, J. X. Liu, and X. Wang. Differences in reproductive effort and sexual recruitment of the seagrass Zostera japonica between two geographic populations in northern China. Marine Ecology Progress Series, 638: 65–81, 2020. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps2020/638/m638p065.pdf.

Zhao:2021:SSM

[ZZYB21] Z. Zhao, L. Zhang, L. Yuan, and T. J. Bouma. Saltmarsh seeds in motion: the relative importance of dispersal units and abiotic conditions. Marine Ecology Progress Series, 678: 63–79, 2021. CODEN MESEDT. ISSN 0171-8630 (print), 1616-1599 (electronic). URL https://www.int-res.com/

articles/meps_oa/m678p063.pdf.